A Complete Bibliography of Publications in the Journal of Fish Biology (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

+ [22, 1238, 44, 1711, 801]. ^{−} [693]. ^{13} [1396, 166, 2189, 1668, 1986, 3089]. ^{15} [1396, 166, 2189, 1668, 1095, 1986, 3089]. ^{2+} [584]. _{2} [1618, 2167, 296, 905]. _{3} [1986]. _{4} [1986]. _{C} [667]. _{crit} [2492]. _{e} [667, 2147]. α [214, 2355, 1931]. b [485]. β [910, 214, 9, 1175, 1647, 189]. β_{2} [1956]. β_{A} [3087]. β_{B} [3087]. : [1852, 700, 1597, 426, 865, 619, 1136]. Δ [1986, 1396, 166, 2189, 1668, 1095, 3089]. Q_{10} [2310]. r [1444]. × [1412, 2167, 2505, 2774, 3024, 1006, 2495].

-1 [847]. -adrenoceptor [1956]. -ATPase [584, 1238, 44, 1711, 801]. -dihydroxypregn-4-en-3-one [9, 1175]. -hydroxycorticosterone [2355]. -K [44]. -oestradiol [1647]. -polymorphism [910]. -selection [1444].

/H [801]. /N [2147].

1 [3017]. 10th [2693]. 11-ketotestosterone [241]. 18S [77]. 19-norclerodane [247]. 1930 [1210].

2 [1581]. 2018 [2635]. 20S [115]. 21-trihydroxypregn-4-en-3-one [189]. 21st [790, 442, 464, 488, 507, 527, 551, 570, 861]. 2nd [108, 141].

3 [2878].

4s [1638].

6 [3016, 681]. 61l [1738].

8 [2364]. 8-epidiosbulbin [247].

9 [1529, 909]. 978 [3016, 3017]. 978-1-119-07577-6 [3016]. 978-1-119-15210-1 [3017].

A. [2882]. Aas [671]. abandonment [1418]. Abant [658]. abanticus [658]. abaster [1051]. abdominalis [1171]. abilities [293]. Ability [2067, 2592, 559, 2544, 948]. abiotic [558, 2021]. ablation [619, 1482]. Able [605]. Abnormal [2762]. Abnormalities [2567, 1401, 1742, 706]. abortion [2769]. Abramis [485, 2505, 1295]. Abrolhos [2030]. Absence [1679, 1646, 811]. absorbing [666]. Abu [776]. Abundance [1618, 1052, 779, 2096, 2140, 781, 1591, 1008, 3049, 1183, 174, 370, 50, 2136, 3045, 2718, 1214, 2797, 1658, 2745, 765, 696, 2442, 1018]. abundances [827]. abundant [512]. abuse [2198, 763]. acanthias [792, 760, 1935, 665]. Acanthobrama [1359]. Acanthocybium [1913, 2289]. Acanthopagrus [47, 613, 1161, 2751, 2348, 2679, 1097, 938, 218, 233, 481, 1029, 171, 804]. Acanthuridae [2599, 2748]. Acanthuroidei [1649]. Acanthurus [3063, 2599, 2713]. acarne [427]. accelerating [1674]. acceleration [808]. accelerometer [1808]. accelerometers [2663]. accelerometry [2334]. access [3013]. accessory [2772]. accidents [1229]. acclimated [1262, 1037]. acclimation [2705, 1837, 3075, 1511, 815]. accounting [1821]. accumulation [2117, 409, 97, 2671]. accuracy [2330, 2958, 2816]. accurate [58, 2004]. Acheilognathus [1754]. achieved [2007]. Acid [1328, 1490, 1245, 2398, 2255, 819, 2117, 1420, 2477, 2557, 2112, 1572, 693, 2552, 2113, 230, 2283]. acidic [674, 2867]. acidification [1792, 2707, 2989]. acidified [2440, 558]. acids [2381, 1683, 530]. Acipenser [471, 2786, 2609, 2622, 373, 2119, 610, 687, 2510, 2059, 2061, 835, 1062, 1848, 2889, 2545, 2174, 1853, 279, 1355, 2205, 1037, 2753, 2575, 2888, 2714, 2109, 877, 2715, 2752, 1654, 2890, 42, 1690, 968, 1664]. Acipenseridae [2201, 42]. Acipenseriformes [42]. Acoustic [2935, 961, 2742, 2958, 2340, 2275, 700, 232, 3085, 476, 2080, 1893, 1924, 2814, 2273, 424, 2467]. acoustics [2597]. Acquisition [3079, 971]. Acropomatidae [2138]. Across [2526, 2754, 2043, 1330, 2731, 2421, 2027, 583, 12, 1202, 1691, 1543, 2515, 1806, 2807, 430, 2463, 2879, 2713, 1633, 2678, 2351, 275, 2497, 16, 2406, 1599, 2525, 1938, 834, 3079, 3058, 1439, 1594]. actin [2487]. Actiniaria [2907]. Actinopterygii [3038, 1766, 598, 2128, 229, 300, 2108, 906, 890, 470]. actinospores [299]. action [737, 1824, 286, 408]. actions [1048]. activated [1263]. activation [1488, 11, 3035]. active [992, 1635, 308, 1520]. activin [3087]. activities [1276, 1675, 1960]. activity [2047, 403, 227, 450, 2303, 1944, 2340, 1238, 668, 299, 845, 115, 1711, 3057, 2779, 2595, 1893, 531, 2131, 2497, 1289, 989, 1116, 850, 2973, 1574, 2654]. acuity [653, 2937]. aculeatus [244, 1397, 210, 678, 1163, 583, 2612, 1684, 1493, 246, 1925, 682, 1366, 649, 2384, 947, 1049, 544, 2351, 2850, 892, 1119, 2858, 514, 2995]. Acute [247, 133, 1836, 1963, 837, 1355, 1379, 3075, 2210, 1794]. acutorostrata [382]. acutus [1776]. ad [2157]. Adaptation [2642, 273, 2534, 2149, 647, 330, 360, 2294]. adaptations [13, 2923]. adapted [2880, 574, 2936]. Adapting [1280]. Adaptive [210, 1916, 549, 1281, 303, 2158]. additional [1991, 850]. address [1557, 2538]. adenohypophyseal [2287]. Adhesion [1240]. adipose [2111, 603, 2671]. adjacent [2008, 2679]. adjustment [1844]. administration [1942, 2117]. admixture [1760]. adrenaline [737]. adrenergic [70]. adrenoceptor [1956]. Adriatic [631, 364, 2622, 2597, 678, 2173, 1242]. adscensionis [2458]. adult [1386, 2628, 2742, 282, 2255, 6, 2958, 1631, 1120, 793, 2777, 2816, 2181, 872, 295, 120, 1389, 17, 106, 1430, 2626, 2092, 729, 1915, 2889, 346, 594, 602, 1167, 1578, 2552, 3032, 1333, 84, 877, 1862, 1728]. adulthood [1895]. adults [1858, 2699]. advance [2199]. Advances [498, 2693, 790, 1830, 1816, 1825, 1803, 2196, 1211]. advantage [1041]. advantageous [372]. advantages [1693]. adverse [1900, 379]. advertisement [2997, 2273]. Aegean [2868, 105]. aeglefinus [509, 1752, 534, 2134, 193, 1965, 276]. Aequidens [2175]. aequifasciatus [407]. aequipinnatus [1384]. aeration [433]. Aerial [429, 2937, 1331, 2648, 1332]. aerobic [2487, 2609, 697, 1827, 1149, 1519, 1832, 1358, 1838, 2848, 1829, 2271, 2512]. Aeromonas [668, 2213, 1712]. Aetobatus [2439, 2226, 1861, 1067, 2819]. afer [2032]. affect [2047, 1181, 1245, 659, 1914, 636, 713, 101, 1933, 1248, 393, 2503, 2228, 1655, 1838, 2797, 811, 693, 2699, 2866, 810, 2053, 46, 1102, 173, 2486]. affected [1023, 804, 1712]. affecting [2506, 2308, 531, 2715, 197, 3023]. affects [1776, 1399, 3001, 2524, 1571, 2763, 1535, 597, 2331, 1153, 3032, 514, 708, 1448, 2973, 1657]. afferent [1208]. affine [1659]. affinis [1148, 2565]. affinity [2439, 1956, 1990, 905]. Afghanistan [1726]. AFLP [21]. Africa [1906, 1776, 2121, 1846, 918, 1558, 2799, 1205, 2561, 2387, 2806, 1451, 2855, 1997, 339, 955]. African [624, 2950, 1863, 2382, 1228, 524, 2559, 1205, 1668, 865, 1963, 2374, 2837, 2610, 1651, 1145, 1299, 240, 70, 2098, 629, 173, 2658, 154, 2106]. After [2796, 1629, 2405, 1942, 2664, 2117, 2336, 2258, 3052, 2092, 1837, 2016, 1066, 2814, 517, 1400]. against [1738, 1325, 668, 104, 1574, 1712, 2276, 2961]. agalactiae [2788, 1608]. AGAT [84]. Age [3050, 2624, 807, 1587, 918, 1105, 2853, 2590, 1740, 297, 351, 2747, 2679, 1996, 945, 2032, 3068, 259, 1628, 1678, 421, 1077, 517, 1485, 1604, 1380, 2506, 2528, 1864, 1629, 2301, 566, 760, 1421, 1804, 1353, 404, 899, 3051, 1475, 2579, 56, 761, 1293, 1062, 1305, 196, 120, 1378, 1059, 1260, 1057, 949, 479, 2092, 2176, 1095, 2756, 2641, 423, 1079, 847, 2678, 1087, 2443, 2434, 1009, 1249, 2775, 656, 126, 1709, 2890, 2118]. age-0 [1293, 1059, 1260, 847, 1009, 1249]. age-at-recruitment [2678]. Age-based [3068, 1485, 2176, 1087, 2443]. age-class [949]. Age-related [259]. ageing [759, 297, 2029, 1419, 2259]. Ageneiosus [2231, 2263]. ages [468, 2990, 966]. aggregata [1828]. Aggregating [2409]. Aggregation [1952, 63, 1739, 245, 1809, 2074, 2119, 3082, 2128, 941, 2779, 2361, 2831, 1722, 283, 2886]. Aggregations [1099, 2024, 2581, 449, 2795, 1682, 1576]. aggression [1702, 707, 1963]. aggressive [256, 1484, 2098, 2394, 197, 1282]. aggressiveness [892]. agility [2238]. Aging [2995]. aglomerular [1769]. Agnathan [3079]. Agnathans [183]. agonist [2819]. agonistic [897, 277, 2098]. Agriculture [980]. Água [2422]. Aguirre [466]. Agulhas [2425]. aids [3078]. Air [1337, 1322, 1329, 1335, 1325, 1327, 1837, 1324, 1334, 1323, 1326, 1333]. Air-breathing [1322, 1329, 1335, 1325, 1327, 1837, 1324, 1326]. Airbreathing [794]. akaara [1918, 265]. akallopisos [2535]. Al [2836, 1516, 2711, 1515, 291, 2251, 2250]. Al-Arab [2836]. Al-Zubair [2836]. alalunga [194, 807, 1105]. alanine [3079]. alanyl [3079]. alanyl-alanine [3079]. alarm [2939, 1389, 2258, 2938, 2102]. Alaska [1463, 1757, 1167, 2553, 2225, 2780]. Alaskan [2252]. albacares [3007, 408]. albacore [194, 807, 1105]. albescens [385]. albida [2268]. albinism [1096, 1454]. albino [2548]. albonubes [367]. Albula [253, 2966, 381]. Albulidae [149]. alburnoides [43]. albus [1521, 1837, 650, 2442]. alcian [65]. alcohol [522]. alcohol-fixed [522]. Alcolapia [1337]. Aldrichetta [2574]. Alejandro [2500]. Alepocephalidae [1898]. Alepocephaliformes [2802]. alert [1995]. Alestidae [2950]. aleuticus [1681]. alevin [2931]. alexandrei [2017]. alexandri [2220]. alfonsino [925, 167]. alfonsinos [2913]. alfredi [3071, 198, 973, 2676]. algae [1591, 504]. algal [2929]. alginolyticus [561]. algiricus [1786]. alien [2235, 2569, 1014, 1447]. alimentary [1539]. alizarin [2489]. allele [373, 1888]. alleles [947, 1594]. Allen [1268]. allies [1787]. Allis [1745, 1969, 3076]. allocate [1605]. allocation [1555, 101, 2141, 1637]. allochthonous [2993, 1441, 1865]. Allometric [1358, 401, 2618, 2318]. allometry [2731]. allopaternal [1091]. allopatric [1484]. allopolyploid [151]. allotriploid [3087]. allow [728, 2466]. allows [1899]. allozyme [1569]. allozymes [363]. almighty [2754]. almost [2242]. Aloe [1712]. alone [897]. along [2211, 611, 385, 675, 1666, 701, 940, 1011, 2441, 1784, 1431, 1293, 651, 833, 460, 2018, 1446, 536, 2490, 730, 2912]. Alopias [137, 1587]. Alosa [323, 2615, 1745, 1969, 1136, 3076, 2789, 2584]. alosine [1698]. alpha [1346]. Alphestes [2032]. alpine [936]. alpinus [1506, 1279, 2097, 2264, 1287, 910, 316, 192, 102, 917, 1510, 1431, 603, 177, 880, 1406, 1642, 435, 1715, 1288, 2797, 638, 1865, 1695, 164, 1031]. also [2512]. alter [2931, 2379]. alterations [1431, 2113]. Altered [1436, 2321, 989, 721]. Alternative [2388, 468, 410, 2182, 2720, 102, 759, 1069, 2951, 2155, 2902, 1288, 1626, 1471, 1844]. alters [988, 1799]. Altiplano [1055]. altitude [1719]. Altitudinal [168, 611]. altricial [407]. always [2649]. amago [838]. amalgamation [2844]. amarus [1030, 374, 1382]. Amatitlania [1539]. Amazon [2631, 2514, 1081, 2608, 1989, 2969, 3006, 2640, 2183, 1235, 2997, 205, 2972, 2608, 248, 1987, 2376, 2360, 2723, 1988, 1985, 2867, 2586, 2956, 1984, 1990, 2324, 27]. Amazonia [731]. Amazonian [2650]. amazonicus [2781, 1282]. amberjack [3062]. ambigua [2297, 2168]. amblycephala [2213, 2774]. Amblypharyngodon [1727]. Amblyraja [1589, 1907, 2319]. amblyrhynchos [2891]. ambrosettii [2563]. ambush [1636, 2646]. Ameiurus [1071, 107, 530]. America [2251, 2250, 156, 1973, 2007, 964, 1974, 1998, 940, 1011, 1978, 1975]. American [1550, 150, 2617, 2025, 2615, 1616, 809, 1883, 2187, 90, 3044, 1670, 3088, 2800, 1411, 1049, 1082, 1991, 2263, 944, 815, 863, 2584, 1690]. americanus [59, 2290, 847, 1449, 1575, 944]. Amia [1336]. amino [2255, 530]. amino-acid [2255]. Amitosis [83]. Amitosis-like [83]. ammocoete [2465]. ammocoetes [1842]. Ammodytes [2211, 227, 2362, 1233, 944]. ammonia [1325, 452]. among [657, 2702, 2673, 954, 836, 1899, 1138, 1737, 350, 699, 1008, 102, 77, 2972, 2799, 1408, 1205, 1111, 773, 1420, 828, 894, 560, 2272, 2611, 886, 2032, 2152, 3056, 2374, 2666, 816, 2188, 240, 2978, 824, 2780, 1256, 705, 131, 1537, 1074, 600]. among-population [1256]. amount [3007]. amp [1072]. amp-up [1072]. amphibious [1700, 1324, 303, 1770, 3023]. amphidromous [2261, 619, 3070, 1341]. amphidromy [3070]. amphihaline [2943]. amphipod [2485]. Amphiprion [2075, 1041, 241, 2535, 2907, 1169]. Amphistichinae [552]. amplification [954, 1703, 830, 972]. amplified [613]. amplitude [2096]. ampullae [2513]. anabantoids [2394]. Anabas [2664]. Anablepidae [2577, 1894]. anadromous [2326, 268, 2308, 2662, 1163, 583, 1427, 1236, 1723, 1131, 3010, 2756, 2616, 3076, 2507, 1471, 2858, 2525, 2584, 1513]. Anadromy [3084, 2988, 1433]. Anaecypris [99]. anaemia [1772, 856]. anaerobic [46]. anaesthesia [2773, 932]. anaesthetics [3058]. Anal [2171, 1840, 1579]. Anal-fin [2171]. analis [1164, 163]. analyses [1112, 613, 817, 3066, 2071, 2018, 2411, 2434, 16, 2589, 1277, 578, 2145]. Analysing [1045, 873]. Analysis [2128, 1111, 299, 3067, 1350, 2156, 157, 2121, 1412, 210, 2508, 2643, 2955, 1528, 769, 398, 1947, 2808, 2910, 718, 2664, 457, 2770, 662, 561, 2319, 687, 1147, 1781, 2630, 1606, 395, 717, 124, 841, 341, 1585, 2058, 174, 196, 1114, 2309, 148, 939, 592, 360, 581, 1596, 1608, 1641, 843, 2539, 2964, 2374, 2232, 2202, 1104, 799, 585, 1651, 1288, 1677, 1292, 2417, 1079, 1913, 333, 1628, 213, 2678, 20, 1890, 2648, 365, 82, 1729, 1750, 1198, 1375, 884, 1307, 1055, 523, 2715, 2721, 3009, 369, 146, 2548, 3004, 480, 2961, 1425]. analytical [762]. Anarhichas [1306, 115]. Anatolia [2542]. anatomical [1697]. Anatomy [2772, 2959, 2075]. Ancestry.fish [2987]. anchoita [1620]. anchoveta [1311, 423, 2678]. Anchoviella [2608]. anchovies [1791]. anchovy [3057, 469, 661, 2568, 214, 1578, 136]. Ancient [624, 522, 1055]. ancistrin [2586]. Ancistrini [2509]. ancistroides [1703]. ancylorhynchus [2002]. Andaman [2588, 2380]. andamanica [2380]. Andean [1980, 1979, 349]. Andes [1308, 168]. androdioecious [670]. androgenesis [2167]. andruzzii [2287]. anemonefish [1041, 241, 2535, 2907, 1169]. anemones [2907]. aneuploid [336]. angel [123, 1454, 1908, 2418]. angelfish [2088]. angelshark [1360]. angiotensin [286]. angled [253]. angler [2042]. angler-led [2042]. anglerfish [1630, 1214]. Anglerfishes [113, 644]. anglers [1556]. angling [715, 1254, 2297, 1735, 1413, 1532, 599, 159]. Angolan [2559]. Anguilla [866, 719, 2882, 1705, 1855, 1351, 1629, 580, 361, 1616, 2451, 2759, 448, 1883, 473, 2496, 2579, 2187, 402, 864, 2590, 2727, 2639, 90, 853, 1962, 2353, 1085, 2267, 2489, 1711, 426, 865, 1411, 3026, 2115, 1495, 1674, 1038, 862, 515, 2872, 863, 199, 1197]. anguillicaudatus [1340, 2477, 2411, 2363]. Anguillicola [1351]. Anguillicoloides [580]. Anguillidae [2686]. Anguillids [1215]. anguilliform [2127, 2267, 2870]. Anguilliformes [385, 1755, 2686]. Angular [1785, 2629, 2331]. Animal [1933, 1928]. animals [1820, 1874]. Anisakis [1186]. anisotropic [1939]. annectens [70]. annotata [351]. Annotated [895, 1879, 2249, 2248]. annotation [2829]. annotations [1042]. Announcement [184, 740, 860, 2572, 2603, 357, 390]. Annual [283, 2119, 2403, 3088, 2136, 3070, 2678, 240, 421, 2580, 2259, 2981, 339, 2486]. Annually [626]. ano [287]. ano-genital [287]. anomalies [325]. anomaly [1868]. Anostomidae [1235, 714, 1890, 1650]. antaly [247]. Antarctic [1704, 2356, 2766, 828, 2066, 1003, 2478]. Antarctica [1003, 297]. antarcticum [1003]. antenna [1458]. antennae [2696]. Anthozoa [2907]. Anthropocene [1976, 2899, 632, 2538]. anthropogenic [631, 2262, 2026, 2877, 2479]. anthropogenically [1953]. antibacterial [1729]. antibodies [2276]. antibody [1738]. anticoagulants [3089]. Antidiuretic [286]. antinutritional [692]. antioxidant [1068]. antioxidation [1244]. antipredator [1361, 661, 2286, 1945, 1243]. antitropical [849]. antiviral [1729, 437]. anxiety [2657]. apart [282]. apela [2714]. apex [1923]. Aphanius [1125, 1627, 1714]. Aphanopus [2758, 1123]. apical [2086]. aplacental [1396]. Aplocheilidae [2839]. aplochelioid [2349]. Aplochiton [1022]. apoda [1779]. Apogonidae [2754, 2054]. apparatus [679, 1660, 1846]. apparent [2814]. appearance [1953]. appetite [2722]. appetite-regulation [2722]. appetitive [2827, 2992]. applicability [2575]. Application [335, 2312, 1123, 2203, 2420, 338, 783, 1811, 1147, 780, 2933, 2295, 467, 497, 2014, 1968, 2276]. Applications [2661, 694, 263, 329, 768, 2888, 76, 2196, 2765]. applied [65, 1557, 2557]. apply [2669]. apportioning [1832]. appraisal [2472, 2029, 2539]. approach [410, 2782, 1199, 271, 1759, 1930, 2952, 518, 841, 445, 761, 589, 376, 135, 1079, 762, 2009, 935, 2133, 1547, 2691]. approaches [2156, 75, 2789]. aprion [2754]. apron [2424]. Apteronotidae [1202]. Apteronotus [1378]. Aquaculture [983, 2708, 1231, 2630, 1280, 1329, 958, 1651, 2616, 2396, 716, 4, 3024, 2898, 2369, 2431, 2765, 182, 108]. aquaculture-produced [1280]. aquarium [2004, 2905]. Aquatic [2507, 156, 1973, 2569, 161, 1974, 158, 2906, 1915, 1983, 1982, 1980, 1821, 1829, 810, 1819, 1820, 1332, 2937]. Arab [2836]. Arabia [1359]. Arabian [2836, 1234, 2817, 340, 575, 1971, 3055, 2730]. arachidonic [1245]. aragonite [2889]. Araguaia [2509]. Arancibia [2500]. Arapaima [1988]. arch [836]. archaeological [2918, 2109]. Archimedean [2598]. Archipelago [245, 1391]. architectural [2114]. architecture [2111, 404, 1957]. Archival [880, 282, 234, 2181, 1000]. archived [522, 1410]. Archoplites [1546]. Arctic [2893, 1506, 1279, 2097, 2264, 1287, 1589, 910, 1414, 1858, 316, 192, 102, 917, 1510, 1431, 851, 603, 177, 880, 1642, 435, 1895, 1715, 1288, 2797, 638, 1865, 1695, 164, 2793, 1031, 909, 362]. arcticus [19]. arcuata [520]. arcuatus [2956]. area [963, 598, 2303, 2790, 1611, 2464, 2019, 2186, 2712, 1406, 347, 207, 1483, 1458, 2283]. Areas [3016, 2435, 2941, 1802, 1214, 1891, 2015, 193, 801, 578, 2690]. areolatus [1087]. argentatus [105]. argenteus [2011, 3013]. argentimaculatus [1678]. Argentina [375, 2391, 320, 1999]. Argentine [1766, 648, 2273, 1249]. Argentinean [675, 470]. argentinensis [1024]. argentiventris [343]. arginase [531]. arginine [1017, 1313]. Argus [1887, 1581, 720, 2049, 71]. Argyrosomus [1846, 1280, 1651]. Arhynchobatidae [2371]. arid [176, 2013, 2012]. ariid [2191]. Ariidae [176, 1142]. arising [1558]. Ariza [461]. Arkansas [1883]. arm [2419]. armatus [2745, 2331]. armed [2745]. armorhead [2056]. armoured [2587, 1788, 2563, 2358]. aromatic [44, 1838]. arowana [2650]. Arratia [112]. array [90]. arrival [835]. arrow [541]. ART [2155]. artedi [2090, 1033]. Artediellus [1010]. Artemia [1768]. arterial [435]. Artificial [2881, 2625, 1679, 713, 2016, 1284, 2965, 369]. artificially [1310, 3026]. artisanal [1993, 1617, 1219]. Ascension [2458]. asexual [646]. Asi [426]. Asia [1920]. Asian [2882, 632, 1837, 1547]. Asparagopsis [3063]. aspect [587]. Aspects [907, 1992, 901, 94, 2588, 123, 63, 1600, 1672, 2592, 2174, 1274, 1932, 2347, 2418, 593]. asperrimus [2370, 1640]. asphyxia [2726]. assay [1801, 1915, 1442]. Assemblage [2870, 779, 1479, 1759, 769, 1632, 168, 2010, 1977, 2020, 2018, 1424, 1985, 1007, 1912, 434, 1814]. assemblages [1661, 364, 1976, 2013, 2026, 1113, 1933, 2186, 347, 1021, 1950, 433, 2805, 2022, 1297, 1075, 3034, 2798, 383, 1979, 1050, 2407, 305]. Assembling [2161]. assembly [1056, 2664, 2829]. assess [2597, 170, 2350, 1897, 1375]. assessed [2439, 2275, 1532, 1136]. Assessing [2140, 1345, 1218, 2061, 751, 474, 2934, 2578, 2040, 2565, 1027, 802, 1916, 1777, 2614, 688, 2028, 1050, 356]. Assessment [2591, 59, 2700, 2753, 1066, 2157, 582, 2140, 1146, 1710, 1863, 1592, 2044, 780, 2379, 467, 2561, 1615, 2016, 2015, 2404, 2352, 2858, 1485, 298, 2793]. assessments [1845, 2148]. assign [996]. assignment [571]. associated [1741, 2391, 2943, 1254, 879, 2260, 347, 1532, 1053, 1641, 571, 2216, 1297, 2648, 997, 1604, 2493, 677, 2752]. Association [1398, 2682, 1839, 979, 707, 2788, 2076, 1337, 2852, 2377, 2419, 2589, 416, 641]. associations [1199, 130, 26, 172, 2134]. associative [1639]. assuming [2176]. assumptions [768, 2202]. Astatotilapia [486]. asterias [2042, 294, 1812]. Asteriscus [948, 339]. astounding [2128]. astra [459]. astroblepid [349]. Astyanax [2617, 1315, 2060, 2566, 2822, 1763, 2422]. Asymbolus [1164]. asymmetric [1221, 1250, 1383]. asymmetrical [2387]. asymmetry [1756]. asynchronously [727]. at-sea [1892]. At-vessel [2607]. Atherina [1491, 939]. Atherinella [2047]. atherinid [1942]. Atherinidae [98]. atherinoides [2929]. atherinopsid [1461, 1960]. Atherinopsidae [2665, 304, 317]. Atlantic [977, 3050, 290, 2581, 2758, 1452, 1475, 395, 543, 1851, 328, 1012, 189, 906, 311, 672, 1741, 3038, 2459, 712, 2869, 400, 2523, 2811, 643, 2156, 2532, 889, 444, 2326, 2708, 484, 2122, 2728, 123, 2724, 1713, 1882, 1271, 2057, 2742, 926, 2717, 2517, 2427, 2790, 2598, 2705, 282, 364, 2024, 1590, 1481, 1857, 1056, 2567, 482, 2421, 2526, 2773, 446, 378, 814, 1935, 2839, 1660, 1047, 1445, 221, 2119, 2518, 1592, 516, 2117, 567, 1345, 1892, 1649, 687, 1811, 301, 974, 2137, 1514, 316, 2044, 748, 1833, 1147, 1519, 511, 2061, 2072, 875]. Atlantic [899, 988, 2952, 2389, 1577, 2126, 2825, 2524, 355, 1306, 943, 1508, 2320, 1238, 1507, 2816, 2087, 1735, 2181, 2304, 174, 2656, 2128, 2941, 1667, 120, 3003, 1571, 2527, 393, 376, 1966, 1420, 871, 2064, 2763, 1792, 2800, 130, 949, 1192, 2081, 2651, 2626, 3052, 187, 1065, 1053, 2159, 929, 1662, 562, 2545, 945, 556, 1419, 2971, 617, 2408, 852, 927, 1107, 2371, 1498, 2707, 2600, 3081, 300, 2996, 2779, 1242, 895, 1530, 2550, 2112, 2698, 417, 1815, 1045, 2080, 2554, 279, 1509, 1501, 1262, 318, 2949, 1786, 2265, 1472, 2547, 1932, 660, 1184]. Atlantic [856, 2016, 1010, 554, 2497, 95, 1535, 1500, 2699, 1064, 1562, 1529, 2099, 2507, 2033, 555, 2030, 2757, 2967, 477, 1154, 904, 1502, 1850, 2466, 2694, 1560, 1511, 1692, 803, 2390, 2418, 1965, 878, 128, 2274, 1153, 2903, 869, 2283, 1123, 1006, 1290, 1505, 37, 2706, 159, 2157, 1487, 1967, 2110, 959, 223, 1622, 251, 1349, 1654, 1856, 1868, 2582, 1499, 1243, 605, 671]. atlanticus [1970, 64, 1010, 2180, 1277]. Atlantoraja [814]. Atlas [141, 1372, 2472, 2070, 2984, 859]. atoll [2966]. atp1a1a.2 [2813]. ATPase [584, 1238, 44, 1711, 801]. atpase6 [2364]. atpase6/8 [2364]. Atractoscion [961]. atresia [1901]. atricauda [2031]. atromaculatus [1073, 2739]. Atrophacanthus [2865]. attached [2067]. attachment [1808, 2680]. attack [2334]. attacking [1859]. attenuated [997]. attract [1090]. attraction [2459, 2187, 513]. attracts [1395]. attributes [160, 2170]. Auchenipteridae [1773, 2282, 2236, 2263]. audax [2483, 960]. auditing [2161]. augmentation [1102]. Aulopareia [912]. aurata [797, 521, 608, 2868, 1313, 452, 44, 1843, 2445]. auratus [825, 715, 1457, 2167, 1871, 201, 2774, 653, 3087, 283]. aureus [1412, 2302, 997]. auriga [1661, 2404]. aurita [2121, 2845, 1899]. aurolineatum [1489]. Austin [3017]. austral [3025]. Australasian [155]. australasica [2716]. Australia [1210, 2045, 576, 1809, 1164, 396, 2844, 377, 1916, 1183, 519, 915, 269, 1630, 1740, 351, 2168, 1646, 129, 430, 2467, 1297, 3074, 798, 1077, 363, 633, 298, 758, 283]. Australian [1210, 1267, 601, 2165, 2191, 1416, 614, 243, 379, 333, 239, 1887, 609, 1301, 1289, 365, 615, 454]. australis [2655, 849, 853, 1089, 1962, 1886]. Australoheros [2306]. Austrolebias [2486]. autecology [255]. author [1360]. authorship [528]. autochthonous [1865]. Automated [2052, 1366]. autonomous [1805]. autonomously [23]. autotetraploids [2774]. autumn [346]. autumnalis [362]. availability [976, 655, 2087, 2120, 1658, 2667, 193, 1025, 1844]. Aversive [2077, 2827, 2992]. avoid [504]. avoidance [1746, 1631, 469, 1535]. Avoiding [2493]. away [2493]. awoara [1918, 736]. axelrodi [1987]. axial [598]. axis [2399, 2444, 2774]. Ayala [109]. Azores [1177, 2378, 1345, 895, 2913]. Azov [2789].

B [1210, 1126, 462, 140, 110, 261, 1211, 568, 588, 3017, 109, 183, 439, 672, 1098, 717, 47]. Ba [619, 1136]. Bacajá [1989]. bach [1917]. Back [311, 2816, 1806]. back-calculated [2816]. backcalculating [479]. backcalculation [1162]. background [293, 1821]. bacteria [2354]. bacterially [923]. Bactericidal [668]. bad [897, 2660]. badge [723]. baerii [2575, 2714]. baffled [1316]. Bagenal [508]. Bahamas [1810, 1807, 1961, 2480]. Bahia [1315]. Baird [1315]. bairdii [2130]. bait [1933]. baited [2459, 1933]. bajad [3045]. Balaenoptera [382]. balance [1328]. Balancing [924]. Balearic [2859, 1480]. Balistes [2506, 2808]. Balitora [100]. Balitoridae [100]. Balkan [85]. Balkans [657, 16]. ballan [2116, 902, 1417, 3037, 224, 1135]. Ballistic [1859]. Baltic [342, 2070, 2914, 2301, 1344, 1916, 1588, 2125, 3081, 2550, 1901, 536, 1038, 1605, 738, 1948, 136, 517, 230, 37, 959]. bamboo [1645, 2833, 1156]. bambooshark [1854]. ban [786]. Banc [1905]. band [3029, 2990, 252]. banded [1011, 2833]. banding [1755]. bandings [736]. Banggai [812]. Bangladeshi [2820]. Bank [1449, 2425, 1487]. bar [1125, 1913]. bar-coding [1913]. barb [836, 1951, 813, 950]. barbata [564, 2724]. Barbatula [188]. barbel [1016, 1063]. barbels [1880, 2293]. barberinus [2243]. Barbus [1880, 1048, 1016, 1063]. barcode [3027, 2673, 2161]. barcode-based [2673]. barcoding [1766, 2950, 1720, 535, 2856, 1442, 3055, 2230, 384, 2838]. Bariche [1209]. Barkin [1173]. barn [1299]. barotrauma [844, 609]. barracudas [1422]. barramundi [802, 837, 1676]. barrier [2336, 3052, 2703]. barriers [2935, 328]. basal [1668]. base [693, 1328]. based [657, 2221, 2121, 2673, 62, 1056, 1558, 1933, 2566, 1410, 1523, 329, 613, 386, 1969, 2176, 2232, 2535, 3068, 86, 716, 135, 1628, 1988, 1087, 2443, 621, 2829, 1154, 2504, 1485, 338, 2589, 548]. baseline [3012]. baselines [3064, 2479]. basic [1129]. Basics [788]. Basin [2647, 1627, 1081, 1906, 331, 375, 3006, 2619, 631, 1773, 2282, 1315, 2558, 2395, 2509, 678, 99, 2843, 222, 1992, 2621, 601, 2549, 1981, 1982, 886, 1459, 248, 2376, 1393, 2563, 2400, 2437, 2606, 2956, 839, 1604, 1541, 2324, 2743, 2495, 2422, 3080, 2953, 1983, 2723, 1480, 1198, 2073, 2969, 2514, 438]. basins [2364, 1205, 1978, 935, 1841, 2586]. basis [700, 1675]. basking [2581, 3088]. bass [797, 2551, 841, 1555, 934, 12, 2200, 452, 455, 1655, 2951, 1259, 346, 1253, 599, 1748, 2583, 1872, 3061, 609, 1647, 615, 2305, 1375, 2691, 1418]. Bassanago [385]. bassi [2425]. Bassogigas [420]. batch [1461]. batfish [1189]. bath [2257]. bathyal [420]. Bathygobius [1025]. bathymetry [1923]. Bathyraja [1166, 2063, 2408, 1757]. batoid [2856, 758]. Batoidea [1207, 757, 2033]. batoids [2248, 2045, 2036, 1879, 2249, 1380]. Batrachoididae [1868]. Batrachoidiformes [1868]. Bay [1632, 396, 1196, 2011, 449, 114, 1021, 2021, 1007, 2119, 1352, 1154]. Bayesian [2952]. be [675, 2242, 2994, 922, 2126, 1640, 2387, 2357, 1255, 2369]. beach [2717, 2019, 1408, 3069, 1381, 2670]. beach-spawning [1408, 3069, 2670]. beaches [2026]. bearded [2534, 1370, 1043]. beardless [2649]. bearer [1252]. bearing [967, 2585]. Beaufort [1131]. Becoming [2238]. Becva [1618]. bed [721]. beds [1402, 1294, 347, 1940, 578]. been [2007]. Before [2016, 1666, 3052, 1497]. Behavior [2985]. Behaviour [1279, 2264, 550, 2275, 1915, 2050, 596, 821, 1952, 2668, 256, 582, 2742, 2043, 2854, 282, 1808, 498, 2031, 63, 1718, 1237, 1011, 988, 134, 3063, 991, 524, 943, 1248, 1187, 1735, 2181, 1408, 1469, 1603, 2942, 984, 1413, 941, 51, 828, 1835, 1532, 1192, 2081, 2184, 2629, 292, 1969, 2804, 2545, 1136, 1939, 3047, 808, 1453, 469, 2174, 1831, 2574, 2698, 1361, 3002, 1573, 1073, 2409, 2293, 1674, 811, 1501, 2583, 846, 1673, 2192, 1078, 1658, 1484, 2652, 639, 1676, 2648, 1500, 2787, 2831, 515, 2957, 2098, 2881, 2695, 2363, 1805, 1370, 987, 885, 2676, 1375]. behaviour [1368, 869, 1945, 2210, 1247, 618, 1730, 2650, 1967, 2110, 197, 2973, 1000, 3028, 1282, 956, 438]. Behavioural [2596, 392, 2892, 565, 2994, 903, 1634, 579, 1715, 1190, 476, 1037, 162, 1171, 2043, 2264, 2255, 572, 3051, 2190, 1389, 1561, 1915, 1250, 300, 279, 688, 2325, 1635, 1379, 942, 1850, 1169]. behaviours [277, 3026, 2286, 2394, 2334]. beholder [2883]. being [2493]. belizanus [1252]. belly [1612]. Belonesox [1252]. belonging [1443, 2740]. Beloniformes [1859, 2153]. below [1480]. beluga [968, 2764]. benefit [2529]. benefits [1222, 2254]. Benguela [2534, 1709]. Benin [1205]. Benthic [2759, 769, 2189, 3013, 2513, 2735, 1548, 2974, 1744, 2748, 193, 2755, 2493]. benthic-fish [769]. benthivorous [1159]. benthophagous [1166]. Benthophilinae [1151]. Benthophiloides [2649]. benthos [953]. bentincki [1186]. bergi [1640]. bergii [2471]. bergylta [2116, 902, 1417, 3037, 224, 1135]. Bering [1934]. Beringian [1319]. Beryx [925, 395, 167, 2913]. best [2928, 2764]. bester [968]. Betta [1334, 217]. Better [897, 2238, 2994, 2524, 1251]. between [1170, 866, 1222, 2708, 1765, 2203, 2182, 2638, 2123, 354, 85, 1203, 2526, 2451, 1398, 2625, 268, 2060, 2308, 2874, 1613, 655, 2382, 2117, 1141, 2429, 1679, 992, 1228, 1795, 3022, 876, 2126, 503, 2719, 474, 2727, 849, 2190, 2379, 1221, 1069, 2200, 1179, 1634, 541, 2165, 3088, 891, 1777, 1476, 1295, 118, 3059, 147, 958, 2120, 1151, 1358, 1352, 1963, 2423, 1538, 965, 947, 1748, 1674, 846, 2028, 318, 847, 163, 2962, 2885, 806, 2497, 946, 95, 218, 2947, 2653, 650, 2377, 950, 2954, 615, 696, 904, 2390, 1945, 37, 2565, 2110, 1885, 1874]. between [1862, 480, 2118, 1383, 1243]. Beyond [2198]. Biannual [1709]. bias [565, 3049]. biased [3071, 2641]. biases [975]. bibarbatus [2534, 1370, 1043]. bibliographic [157]. bicephalic [1124]. bichir [2937]. bicirrhosum [2650]. bicolor [719, 2882, 846, 2882, 1762]. bicoloratus [931, 1390]. bicornis [1318]. bidirectional [615]. Biennial [758]. biflagellate [2837]. big [939]. big-scale [939]. bigeye [290, 1176]. bigger [2524]. bighead [2936]. Bight [136]. bigscales [1530]. bilineata [967]. bilineatus [1199]. Bill [606]. billfishes [169, 2685]. Billion [1463]. Billion-dollar [1463]. bimaculatus [2887]. Bimini [1810, 1961]. bimodal [1826, 1981, 1328, 654]. binding [1514, 1956]. bio [2015]. bio-indicator [2015]. bioaccumulation [1989]. bioacoustics [2877]. Biochemical [1680, 885]. Biodiversity [1976, 549, 1533, 161, 158, 2901, 635, 2341, 1975, 2251, 1139, 2250]. bioelectrical [1585, 806]. bioenergetic [394, 275, 2085]. bioenergetics [917, 1835]. Biogeographic [614, 1140]. biogeographical [2248, 1879, 2249]. biogeography [567, 1633]. bioindication [180]. Biological [1108, 2468, 2056, 754, 1718, 2428, 753, 229, 1024, 641, 628, 1971, 1839, 2845, 2844, 1345, 1120, 651, 2740, 2100, 1655, 35, 938, 1186, 1671]. biologically [1930, 1520]. biologists [2689]. Biology [463, 487, 506, 526, 1373, 749, 1286, 140, 761, 983, 1791, 2894, 1099, 746, 183, 1872, 1174, 2104, 441, 2658, 1875, 489, 907, 1234, 2034, 2588, 792, 745, 1271, 755, 2483, 194, 3025, 378, 814, 231, 1046, 1907, 1600, 1652, 2068, 748, 91, 294, 2135, 2299, 925, 1743, 2389, 340, 575, 2304, 2531, 351, 2747, 1553, 960, 691, 901, 2348, 1474, 1909, 2449, 401, 919, 2371, 1757, 2723, 352, 2821, 121, 3046, 1087, 1012, 1614, 2037, 2948, 747, 2033, 1803, 1176, 2139, 418, 1043, 2855, 3021, 741, 2765, 921, 1106, 142, 2457, 2658]. bioluminescence [3090]. biomagnification [1987, 1989]. biomarker [850]. biomarkers [1490, 204, 2112, 844, 2885]. Biomass [3014, 2106, 2463]. Biomass-dependent [3014]. Biomechanics [647, 679]. biopsy [1253]. biosphere [1402, 635]. Biosynthesis [2355, 2381]. biotic [558]. birostris [2429, 1699, 973, 2676]. births [2229, 1861]. Biscay [1352, 1154]. bisexual [540]. Bite [3002, 2496, 2866]. bitterling [1030, 374, 1754]. bivalve [2061]. bivius [2493]. Biwa [1595, 2594]. black [701, 2551, 47, 3011, 613, 2751, 232, 2348, 2679, 911, 1097, 1071, 107, 844, 2667, 972, 2867, 122, 2318, 730, 530, 171, 804, 2276]. black-bass [2551]. black-chinned [730]. black-faced [2318]. black-spot [232]. blackfish [1319]. blackmouth [1278]. blackspot [204, 1274]. blackspotted [459]. blackstripe [1540, 1579]. Blacktip [2852, 1076, 1810, 915, 2165, 2680, 2436, 241, 1058]. Bleeker [375]. bleekeri [2730]. blennies [287, 801, 2838, 140]. Blenniidae [2327, 801, 2838]. Blennioidei [105]. blenny [1889, 2684, 2327, 1281, 1922, 2932, 2318]. blind [2050, 950]. Bloch [2724]. Blood [1254, 620, 559, 2323, 927, 1615, 2703, 1990, 905, 3089]. blood-O [905]. bloom [97]. blotched [2322]. Blue [786, 2122, 65, 1701, 2910, 1398, 1460, 849, 2302, 1797, 830, 1473, 430, 2879, 53, 2954, 997, 1782]. blueberry [2004]. bluefin [238, 364, 1590, 482, 2681, 2340, 134, 553, 436, 1009, 618]. bluegill [249, 21, 1586, 445, 451, 1375, 2860]. bluegills [31]. bluemouth [352, 2307]. bluenose [297]. bluespine [1485]. bluespotted [2189]. blunt [2774]. blunthead [2808]. boardmani [1164]. boat [1413]. Bodianus [231]. bodied [2303, 98, 1677, 2721]. bodies [624]. Body [1503, 2786, 2677, 2131, 410, 777, 2296, 448, 31, 1845, 2137, 1914, 1030, 934, 503, 589, 1681, 2977, 376, 1111, 281, 590, 2046, 1771, 2136, 512, 2544, 2744, 1358, 734, 2718, 1495, 1615, 2546, 2065, 847, 1082, 806, 2005, 95, 1799, 2377, 2983, 1119, 1025, 2438, 2981, 3021, 3058, 1036, 425, 1002]. body-element [2137]. bogaraveo [204, 1274]. bogue [1953]. Bolbometopon [2361]. Boldness [892, 1702, 583, 1925, 1532]. Boleophthalmus [1333]. bomb [297]. bonapartii [2048]. bonariensis [1942, 2399, 1194, 304, 320, 428]. bonasus [2885]. bone [2508, 1856]. bonefish [253, 2966]. bonefishes [149]. bonga [2828]. bongo [1588]. bonnethead [2783, 2129, 1475]. Bony [183, 1315, 2815]. bonytongue [1205, 2837]. Book [208, 289, 2631, 1550, 1373, 112, 1210, 1267, 113, 1426, 955, 908, 605, 1018, 108, 1126, 142, 550, 462, 140, 110, 1265, 1463, 261, 141, 549, 812, 690, 983, 1172, 1209, 1269, 1268, 1270, 1371, 1372, 3016, 525, 1070, 1173, 604, 831, 1099, 1211, 568, 182, 956, 588, 3017, 2070, 859, 1127, 109, 2314, 183, 1174, 739, 2822, 2893, 439, 9, 671, 139, 138, 189, 356, 111, 311, 438, 548, 672, 389, 980, 981, 982, 1264, 1266, 1098, 1404, 1374, 1874, 1873, 2453, 461, 569]. Boon [356]. Boops [1953]. boosts [903]. boreal [558, 89]. borealis [450]. Boreogadus [1414, 1858, 1895]. borne [922]. Borsa [2249]. bosc [1445]. Bosnia [1675, 1201]. Bostrichthys [296]. both [403, 560, 1281, 1714, 2504]. Boulengerella [2360]. bound [2716]. boundaries [724]. boundary [268]. bovinus [1753]. bowfin [1336]. boxes [180]. boyeri [939]. Bracegirdle [508]. Brachyplatystoma [1666, 2972]. brachypomus [2953, 2871]. Brachyrhaphis [1644, 403]. brackish [574, 1510, 1085, 2390, 399]. Brain [989, 8, 2509, 988, 1567, 894, 2501, 18, 1885, 789, 1717]. brain-cell [2501]. Brain-encysting [989]. Braithwaite [525]. Brama [1089, 1886, 485, 2505, 1295]. bramble [1234]. branchial [2813, 1156]. branchialis [1399]. brasiliensis [2047, 2576, 1613, 1258, 341, 2885, 2037, 625]. Brauner [1265, 1269]. brauneri [2649]. Brazil [2282, 1666, 2378, 2619, 2034, 3027, 2673, 1773, 1438, 2024, 1315, 2031, 2395, 2339, 2185, 2839, 2019, 567, 2843, 222, 2853, 28, 250, 1081, 1474, 2029, 3047, 248, 1987, 2371, 2003, 1530, 1393, 2563, 1617, 2887, 2038, 2400, 2437, 1217, 1012, 2404, 2606, 2867, 2033, 1110, 670, 1841, 839, 2223, 2073, 1997, 2495, 2358, 2422, 3080]. Brazilian [1989, 2047, 2002, 979, 2017, 157, 1235, 385, 2576, 2027, 2770, 2429, 341, 1348, 2463, 3061, 2014, 2037, 2009, 2969]. breadth [2183, 1408]. Breaking [1323]. breaks [724]. Bream [983, 797, 1199, 608, 47, 2868, 334, 485, 1089, 2751, 1313, 452, 2827, 2505, 1295, 2774, 1447, 2445, 171, 804, 2691]. breath [2960, 2946]. breathe [1323]. breathers [1829]. breathing [1335, 1325, 1337, 1322, 1329, 1327, 1837, 1324, 1334, 1326]. breeders [2532]. breeding [1778, 2819, 366, 118, 1104, 1243]. Brevimyrus [2171]. brevirostris [1954, 1808, 3050, 378, 253, 767, 1961, 2814]. brevirostrum [610, 2174, 279, 1355]. Brevoortia [945]. bridled [1199]. brief [1452, 1818]. brier [1761]. brill [652, 1293, 627]. brink [1052]. bristly [1108]. Britain [1874, 14, 2314, 1264, 1404]. British [290, 291, 1839, 2874, 1593, 1233, 1687, 1812, 778, 2912]. britskii [2834]. broad [449, 851, 495, 505, 504]. broad-nosed [495, 505, 504]. broad-scale [449]. Broader [2496]. broadfin [2468]. broadnose [306, 2518, 764]. broken [1403]. bronze [734, 2562]. brood [260, 2076, 2623, 191, 1418]. brood-pouch [191]. brooding [1142]. broodstock [1642]. brook [2760, 1488, 662, 2308, 1746, 1631, 2719, 1585, 1842, 2046, 579, 1104, 213, 1549, 806, 1492, 1471, 46, 3028, 165]. Brotula [2724]. Brown [348, 612, 672, 611, 1569, 2405, 85, 667, 2597, 1563, 1468, 316, 940, 2677, 2988, 1746, 1631, 2142, 962, 2833, 1120, 917, 2186, 2336, 2440, 700, 51, 880, 817, 2930, 1723, 2740, 1131, 845, 1836, 2159, 460, 14, 2761, 2470, 924, 1924, 1185, 1484, 2989, 1284, 406, 2084, 2653, 2978, 2881, 935, 331, 1198, 839, 2525, 2672, 1006, 1967, 857, 2973, 79, 1779, 416, 25]. brown-banded [2833]. brown-marbled [79]. brownbanded [1645]. brucus [1234]. bruniensis [2135]. BRUV [1933]. Brycon [256, 1672, 2592, 2781, 2005, 1282]. Bryconalestes [2950]. Bryconidae [2781, 2005]. bubble [2394]. bucchichi [1884]. bucculentus [758]. Buchmann [1127]. budgets [318]. building [1227]. bull [2818, 282, 234, 1593, 1806, 1891, 3046, 796, 1094]. bullhead [479, 1071, 107, 530]. Bullockia [3005]. bumper [1064]. bumphead [2361]. burbot [2591, 3035, 1959, 893, 845, 2132]. bureschi [16]. Burrowing [2746]. burrows [1333]. burst [1508, 1436, 1256]. burst-swim [1508]. burtoni [486]. Bush [130, 949]. butcheri [2751, 233, 481, 171, 804]. Butis [852]. butleri [600]. butter [2491, 573]. butterfly [2959, 1303, 921]. butterflyfish [1075]. butterflyfishes [285]. Büyük [2542]. by-catch [1054, 2045, 821, 2404, 1654]. Bycatch [2911]. Bythaelurus [1108, 254].

C [142, 1265, 549, 983, 1269, 1372, 831, 1211, 568, 439, 356, 1098, 2508, 1396, 166, 2189, 1852, 1597, 1312, 1867, 2083, 1668, 1986, 3089, 1244]. C-terminal [1244]. C-type [1312, 1867, 2083]. C. [2165]. C.-H [1372]. C.-W [1372]. C9 [668]. C9-deficient [668]. Ca [584, 700, 426, 865, 619, 1136]. Caatinga [2673]. caballus [820]. cachara [2001]. caeca [621]. caerulea [3025]. caerulescens [1778, 1595, 2594, 2980]. caeruleum [2078, 1398, 2516]. caeruleus [2618]. cage [803]. cage-cultured [803]. cages [2708]. Cairns [289]. calcareous [201]. calcarifer [802, 837, 1676]. calcea [1761]. calcein [1284, 71]. calcein-marked [1284]. calcite [2201]. Calcium [1488, 1263, 1379, 3035, 833, 2889, 3053, 2265, 2201]. Calcium-activated [1263]. Calcium-dependent [1379]. calculated [2816]. Caledonia [1630, 3068]. calibrated [3049]. California [3015, 2812, 2959, 1643, 1732, 1870, 2269, 2744, 1725, 3033, 211, 2218, 2670, 1885]. californica [1454, 1908]. Callichthyidae [928, 2587, 2956]. calling [1936]. Callionymidae [1517]. Callionymus [1517, 896]. Callorhinchus [2990]. calls [2997, 1752, 2273]. calretinin [1710]. calva [1336]. camera [2854, 2958]. camouflage [1534]. campechanus [2423]. Campogramma [2488]. camtschaticum [909, 664]. Can [68, 1093, 2529, 2387, 2357, 2148, 1255, 2913, 2967, 1162, 882, 324, 922, 3049, 2379, 2298, 559, 880, 724, 1329, 1562, 2369]. Canada [3083, 1233, 2055, 382, 638, 863, 2893]. canadensis [2999]. Canadian [398, 516, 851]. canadum [478]. canal [1377]. canaliculatus [623, 2381]. canals [2982]. Canaria [2902]. Canary [1786]. candidate [2396]. canescens [254]. canicula [3078, 941, 2601, 1000, 2721]. Cannibalism [2470, 2717, 1663, 2076, 1928, 648, 436]. cannibalistic [1658, 1676]. cannonball [2682]. canopy [1566]. canosai [2273]. cantharus [701]. Canyon [1868]. capability [345]. capacities [2947]. capacity [1508, 1680, 2827, 2578, 317, 1505]. capanema [731]. Cape [1685, 586, 2806, 721]. capelan [616]. capelanus [616]. capelin [2717, 3069]. capensis [2625, 1685, 586, 1451, 1709, 3077]. capriscus [2506, 2808]. capsule [1023]. capsules [254]. captive [2346, 2238, 2097, 2908, 2729, 2229, 3086, 1067, 2819, 1561, 2557, 1936, 2907, 2077, 1243]. captive-reared [2238, 3086, 1561]. captivity [2264, 1785, 2557]. Capture [956, 1339, 2208, 1532, 2629, 57, 927, 1190, 1262, 2699, 181, 1341, 1860, 857]. captured [2825, 1725, 2769]. captured-induced [2769]. captures [2900]. capurrii [431]. carangid [820]. Carangidae [2488, 1460, 1064]. Carangoides [3045]. Caranx [979, 1460, 820]. Carassius [1457, 2167, 337, 1040, 229, 2774, 637, 3087, 825]. Carbohydrates [666]. carbon [770, 518, 2048, 592, 2652, 665, 2867, 2210, 1451]. carbonate [2889, 2201]. Carcass [1724, 786]. carcasses [251]. carcharhinid [2165]. Carcharhinidae [2468, 124, 761, 1961, 1094]. Carcharhiniformes [124, 2983, 907, 2468, 2790, 1164, 254, 519, 1971, 1007]. Carcharhinus [2372, 271, 282, 234, 1076, 1734, 2260, 1810, 761, 915, 2165, 729, 1806, 2680, 2436, 2852, 241, 1058, 2891, 1124, 1932, 3046, 2648, 2040, 2077, 796, 1094]. Carcharias [3002, 1809, 2416, 1693, 2744, 1805]. Carcharodon [1809, 2416, 1693, 2744, 1805]. Cardiac [856, 1386, 2487, 737, 584, 1414, 1149, 3053]. cardinal [1987]. Cardinalfish [812, 923, 2054, 2124]. cardinalis [2348]. cardio [1336]. cardio-respiratory [1336]. cardiomyogenic [23]. Cardiorespiratory [1772, 1399, 898, 1837]. care [12, 828, 1554]. Careproctus [308]. Caribbean [1140, 1611, 2023, 2262, 2129, 2179, 1978, 2108, 2418, 1534]. Carmine [2807]. Carnegiella [914]. carnivorous [2220]. Carolina [2385, 2852]. carophila [2299]. carotenoid [1279, 101]. carotenoids [1683]. carp [1570, 1947, 1944, 718, 2809, 1998, 266, 1237, 717, 97, 913, 269, 2058, 2778, 2266, 116, 1261, 1365, 668, 1867, 2302, 2671, 1137, 948, 1999, 808, 716, 688, 332, 1638, 531, 2774, 1881, 637, 1729, 1241, 437, 2589, 3087, 2936, 2344, 1547, 677, 1875, 339, 825, 146, 826, 1106, 1002, 1964, 2961]. carpio [2221, 1570, 1947, 1944, 216, 718, 1998, 2086, 2167, 1237, 97, 913, 269, 116, 1261, 1365, 2302, 1137, 948, 1999, 808, 531, 1881, 1729, 2344, 677, 339, 369, 146, 1002, 1392, 1964, 2961]. carps [216]. Carrier [549]. carrying [292]. cartilaginous [1533, 789]. case [1766, 782, 210, 1705, 1479, 1056, 2576, 1767, 2530, 1833, 1454, 1880, 1280, 472, 2819, 1642, 949, 1693, 2657, 419, 1893, 2648, 2230, 2954, 1277, 2106]. cases [1096, 1164, 617, 2408, 2493]. Caspian [438, 1086, 840, 935]. caspicus [963, 840]. Casselman [289]. castelnaui [814]. Castle [2127]. Castro [466]. catadromous [2882, 837, 615]. Catalina [2269]. cataloguing [465]. cataractae [2224]. cataracts [104]. Catch [1234, 2588, 2777, 785, 781, 1863, 2518, 1054, 1345, 1735, 1961, 2045, 775, 1473, 821, 2404, 159, 1654]. catch-and-release [1735]. Catch-related [2777]. catches [565, 2472, 1617, 1198, 2654]. catchment [2741, 2081, 251, 1103]. catchment-scale [1103]. catchments [1883, 2525]. categorize [1348]. catfish [2647, 2321, 1203, 2091, 2220, 1130, 2587, 1606, 932, 2312, 2083, 1797, 1788, 1963, 2481, 2563, 2050, 2231, 3072, 2402, 1100, 2263, 990, 173, 2397, 2495, 1607, 2358, 2548]. catfishes [176, 823, 2191, 1142, 2479, 591, 349]. Catla [266]. Catostomus [2055]. catshark [907, 1108, 1278, 254, 2601, 2983, 1000]. catsharks [3078, 941]. caudal [137, 1137, 2595, 2088, 2605]. caught [2588, 1491, 2980, 2906, 1535, 810]. Caulerpa [2638]. Caulolatilus [2738]. causal [889]. cause [2566, 314]. cause-and-effect [314]. caused [114]. Causes [1216, 2526, 867, 2952, 531, 637]. cautus [1734]. cave [2287, 2414, 2215, 950, 634]. cave-dwelling [2215]. cavefish [2104]. Cavefishes [2822]. Caves [2822]. cavity [1574]. cd [909]. Cd8 [1005]. cDNA [1340]. Cech [739]. celebesensis [2590, 2727, 864]. Celebration [1404]. celibela [248]. Cell [792, 1520, 694, 1470, 1710, 2687, 2227, 717, 1567, 2415, 443, 2295, 1137, 2501, 1195, 1262, 265, 2575, 1529, 1568, 2888, 2088, 1394, 79, 2605, 1968, 1717]. Cell-penetrating [1520]. cells [792, 1665, 266, 2086, 574, 2287, 1313, 2884, 1682, 705, 2476, 2175, 2584, 425, 1520]. Cellular [452]. Celtic [2070, 2917, 2447, 1182]. cemali [2851]. Census [667, 1933]. Centracanthidae [386]. Central [940, 1011, 2594, 2591, 2435, 168, 937, 1048, 1987, 103, 895, 270, 1754, 1913, 1633, 995, 1110, 1906, 434, 1868, 2617, 3027, 1842, 2857]. centrarchids [2683, 30]. centrina [1785, 2629, 2100]. Centromochlinae [1773]. Centromochlus [1773]. Centrophoridae [753]. Centrophorus [2449, 1622]. Centropomidae [2389]. Centropomus [1140, 2019, 1937, 2389, 2745, 642]. Centropyge [1762]. Centroscyllium [3011]. Centroscymnus [541]. centuries [2441]. century [790, 442, 464, 488, 507, 527, 551, 570, 861]. cepedianum [1670]. cepedianus [306, 2518, 764]. Cephalic [1194]. Cephalopholis [91, 2049]. Cephaloscyllium [766, 2229]. cephalus [521, 2955, 510, 1784, 1731]. cerebral [1794]. cerebralis [299]. cernuus [1117]. Cerrado [3027]. cestode [2405, 1681]. Cetichthys [906]. Cetomimidae [906]. Cetopsidae [2647]. Cetopsis [2647]. Cetorhinus [2581, 3088]. cettii [2494]. cf [2242, 371, 2968, 818]. CFT [1093]. cgnrhII [1448]. Chaetodontidae [285]. chain [1239, 2381, 1987]. chains [457]. chalcogrammus [2963, 2177]. chalk [2527, 1500]. challenge [561, 994, 1966, 1711, 865, 993, 1073, 1355, 1881, 689]. Challenges [442, 464, 488, 507, 527, 551, 570, 2817, 1386, 1271, 1231, 1974, 1046, 632, 2531, 1978, 861, 571, 1825, 638, 956]. chalumnae [39]. chambers [2960]. chameleon [2306]. chance [2974]. Chang [1372]. Change [1053, 2649, 2805, 323, 2528, 2326, 2149, 1294, 322, 545, 1845, 3007, 632, 316, 2630, 324, 636, 1187, 1365, 1851, 447, 130, 320, 455, 2032, 2810, 1135, 315, 328, 314, 1219, 723, 2084, 317, 2209]. Changes [1732, 301, 2778, 1866, 435, 2010, 225, 2090, 2845, 631, 2450, 450, 2296, 2398, 2880, 2031, 2543, 1998, 2994, 574, 134, 341, 864, 3011, 2448, 559, 1477, 342, 2529, 1512, 1424, 1336, 1888, 191, 2748, 2114, 1522, 345, 1025, 2714, 164, 2284, 2438, 1885, 1513]. Changing [2842, 2536, 2766, 2537, 343, 130, 327, 326, 995]. Channa [2643, 2322, 1946]. channel [1606, 2016, 2917, 1352, 1473]. channels [1263]. Chapada [1315]. Chapalichthys [711]. char [2797]. characid [2003]. Characidae [2000, 2621, 2003, 2324, 2743, 3006, 2617, 2782, 1315, 2558, 1672, 2592, 222, 248, 2004, 2401, 1987, 1393, 2666, 2400, 2437, 1110, 1841, 2422]. Characidium [371, 3080]. characids [2183, 2360]. characiform [132, 625]. Characiformes [1393, 2437, 2005, 3006, 2002, 2619, 2617, 2950, 2000, 1235, 2997, 1315, 2558, 2395, 1088, 2514, 222, 2621, 2953, 1081, 2484, 248, 2004, 2401, 2376, 2003, 2233, 2887, 2400, 2437, 1890, 1110, 1841, 839, 1537, 2324, 2743, 1650, 458, 2422, 3080]. characin [1181, 1453, 2471]. characins [2170]. Characodon [1302]. character [2587]. characterisation [2116, 2175]. characteristic [2324]. Characteristics [727, 1518, 3079, 2832, 1283, 2959, 1602, 2056, 2844, 2119, 2966, 1286, 1303, 295, 3030, 1052, 1747, 374, 2505, 2360, 1897, 1812, 1182, 2525, 2716, 752, 1103, 698]. Characterization [457, 2581, 1450, 2482, 1005, 2058, 2213, 1137, 1213, 2714, 736, 1448, 1440, 1964, 694, 954, 1942, 1697, 1611, 1340, 1457, 1738, 266, 1718, 1540, 1606, 1258, 717, 1567, 1703, 1114, 2083, 833, 443, 1848, 581, 1494, 2295, 2501, 1195, 830, 1707, 265, 2006, 625, 1529, 642, 1492, 380, 1568, 2088, 1650, 1394, 2444, 2878, 1349, 1648, 1717, 2961]. Characterizations [1638]. characterize [1940, 883]. characterized [681]. Characterizing [195, 938]. characters [676, 2484]. charr [1506, 1279, 2097, 2264, 1287, 910, 662, 2308, 258, 316, 192, 2383, 102, 2719, 917, 1510, 1431, 2336, 603, 177, 880, 1642, 435, 1715, 1288, 638, 1865, 1695, 46, 164, 2793, 1031, 1648]. Chasmistes [226]. Chassot [291]. Chaunacidae [1630, 2960]. Chaunacops [1630]. check [1721]. Checklist [2657, 2673, 2248, 1879, 2249]. Cheilinus [63, 237]. Cheilodactylidae [2844]. Cheilodactylus [2844]. Chelyabinsk [598]. Chemical [253, 2489, 2939, 1528, 2187, 732, 927, 904]. Chemicals [2469]. Chemistry [2199, 2017, 1524, 701, 1023, 1691, 2198, 2436, 562, 2197, 948, 1107, 1615, 882, 883, 2196]. Chemosensory [1090, 1395]. Chesapeake [1632]. Chiasmodontidae [3090, 2212]. Chile [1533, 3025, 64, 1186, 1886]. Chilean [651, 1207, 952, 1055]. chilensis [1882, 755, 2729]. Chiloscyllium [1645, 2833, 1854]. Chimaera [2421, 2299, 556]. chimaeras [2248, 1879, 2249]. chimaerid [1487]. Chimaeridae [782, 2565]. chimaerids [2299]. chimaeroid [556]. chimaeroids [3036]. China [2348, 2679, 775, 2414, 2215, 2490, 877, 2645, 733]. chinensis [2490]. Chinese [836, 81, 1567, 1062, 1848, 2490, 877, 2444, 634, 1664]. chinned [730]. Chinook [2760, 1503, 1528, 2296, 3001, 2931, 101, 2254, 1034, 78, 325, 1102]. chionaraia [3061]. Chionodraco [2829]. Chirostoma [1461, 1960]. chitinolytic [1960]. chloride [1665, 574, 1262, 1263]. chloroculus [1614]. Chlorophthalmidae [1318]. Chlorophthalmus [1318]. Chloroscombrus [1064]. Choerodon [2544, 232, 2979]. Choice [1032, 367, 2492, 1034, 288, 2907, 415, 596, 2674, 2274, 1657]. choices [2681, 1388]. choked [2016]. Cholinergic [70]. Chondrichthyan [617, 782, 614, 155, 742]. chondrichthyans [2991, 2926]. Chondrichthyes [2248, 2249, 2034, 782, 755, 1907, 124, 2100, 556, 2371, 1012, 2983, 2493, 2425, 2380, 1094, 1879]. chondrocranium [2826]. Chondrostoma [61]. chooses [2965]. choosiness [2516]. chorion [1706]. choruses [2997]. chromis [2378, 2378, 2970, 2540, 827]. Chromosomal [371, 1755, 77, 458]. chromosome [858, 714, 736]. chromosomes [795, 76]. Chronic [2866, 1719, 842]. chronological [2204]. chronometer [297]. Chrysobrycon [2324]. Chrysophrys [1871, 2827]. chrysopterus [241]. chrysorhynchus [1060]. chrysurus [1064, 593]. chuatsi [457, 1567, 1005, 2878, 3004]. chub [2346, 2955, 510, 1073, 2739]. chubsucker [3083]. Chum [874, 2255, 335, 2704, 1789, 1594, 2118]. Chuzenji [1400]. Cichla [2339, 2152]. Cichlasoma [1470]. cichlid [728, 1470, 24, 524, 624, 486, 1539, 2306, 147, 666, 2152, 160, 415, 2618, 2098, 2073, 2175]. Cichlidae [2339, 1258, 1343, 2073]. Cichlids [1905, 1141]. cinereus [2267]. circadian [1049]. circuli [354]. circumstances [2699]. circumtropical [2051, 1760]. circumvents [2674]. cirratus [2293, 3060]. cirrhatus [545]. Cirrhinus [2266, 716]. cirrosa [2396, 1924]. cisco [2090, 1033, 2225, 362]. Cl [693]. clade [1145]. Clair [2752]. Clarias [2150, 1963, 173, 1997]. ‘Clarification [290]. Clariidae [2397]. Clark [2368, 2367]. clarkii [1236, 2747, 2756, 815, 1504, 1236, 2756, 1504]. class [2788, 949, 467, 150, 924, 910]. classes [1030, 1348]. classic [865]. classical [1272]. classification [377, 2198, 2052]. clavata [900]. clcn2c [2813]. Cleaner [2765, 846]. Cleaning [2556, 2436]. clear [2930]. client [846]. clients [2190]. Climate [323, 2630, 328, 314, 2528, 1294, 322, 316, 324, 636, 1187, 2941, 130, 320, 327, 326, 321, 315, 1219, 2806, 317, 2209]. climbing [2664, 888]. clines [1599, 2344]. clingfish [2450, 2998]. clinid [2585]. Clinidae [105]. Clinitrachus [105]. clip [2434]. clock [3053, 1049]. clonal [2575]. clonal-cell [2575]. clones [292]. Cloning [909, 1340, 561, 717, 1312, 2083, 581, 18, 214, 1241, 1931, 2548]. Close [1222, 722]. closed [2708, 582, 1214, 1828]. closely [1540, 1476, 1347, 2565]. closure [2205]. clown [1041, 1169]. clownfish [241]. Clupea [1580, 58, 1795, 2963, 394, 1363, 929, 1662, 2949, 1605, 1535, 230]. clupeaformis [2641, 2553, 1910, 1750]. Clupeidae [520, 2133, 669]. Clupeiformes [2762, 1620, 2608, 2222, 2789, 2855, 669]. clusters [1896]. Clypeobarbus [1906]. Co [145, 1736, 2183, 1580, 2056, 1846, 1795, 512, 2272, 2797, 88, 2454, 905]. co-culture [2454]. co-existence [2272, 2797]. Co-expression [1736]. co-occurrence [2183]. co-occurring [1580, 2056, 1846, 512, 88]. Co-ordination [145]. co-reared [1795]. Coad [261]. coarse [1102]. Coast [1012, 907, 1108, 2211, 385, 1894, 701, 2101, 940, 1011, 2441, 2320, 519, 652, 1293, 651, 1233, 1740, 2045, 2857, 2649, 2387, 799, 536, 2490, 2855, 283, 1654]. Coastal [1551, 3047, 1741, 1177, 2017, 306, 1705, 1855, 1479, 1438, 781, 1076, 2759, 1468, 567, 1892, 2359, 1149, 2607, 1734, 2256, 934, 1236, 2173, 2825, 474, 449, 269, 2799, 17, 3088, 1731, 1192, 1021, 1476, 2629, 2071, 2018, 1145, 160, 2022, 1991, 273, 2947, 365, 1775, 1458, 2139, 2223, 305, 1504, 1172]. coasting [2860]. coasts [675, 696]. cobia [478]. Cobitidae [336, 353, 2411]. Cobitis [81]. cocoa [2491, 573]. Cocos [274]. Cod [322, 511, 1741, 686, 2459, 712, 1239, 1713, 1310, 1285, 2742, 926, 1056, 1047, 1414, 1858, 974, 2044, 1795, 2952, 651, 120, 376, 1420, 1895, 187, 2066, 2125, 264, 927, 3081, 476, 2550, 1045, 1167, 379, 1184, 2497, 538, 2166, 878, 1153, 189, 1349]. coding [1913]. Codium [3077]. codling [2576]. coefficient [2860, 1594]. coelacanth [39]. coelenterate [1239]. coelestis [2300]. coeliacomesenteric [2703]. coelolepis [541]. Coexistence [2761, 2666]. coexisting [427, 205, 2170]. coffinfishes [2960]. cognatus [1681]. cognition [986]. cohesion [878]. Coho [1764, 1429, 2701, 2994, 2275, 894, 38, 1536, 2578, 2352]. cohort [899]. Coilia [3057]. coincide [2269]. coioides [2847, 561, 1312, 1596, 843, 1200, 212]. cold [580, 59, 2257, 1207]. cold-monomictic [580]. coldwater [2954]. Colin [1099]. collected [2464, 1732, 1588]. collection [2434]. collective [2795, 1535]. Collichthys [2884]. colliei [782]. Colombia [2647, 3006, 2621, 168, 2324]. colonies [827]. Colonization [1273, 2378, 1084, 1152, 14, 2352]. colonizations [601]. colonorum [615]. Colorado [815]. colorimetric [1120]. Colossoma [2867]. colour [897, 2681, 2285, 2220, 202, 623, 2921, 1361, 1362, 1624, 1119, 2858, 3008, 3058]. colouration [1398, 2612]. Coloured [3051, 2725]. Colours [2921]. Columbia [1429, 1593, 1233]. column [2396, 706, 968]. combination [2489]. Combined [592]. Combining [1880, 1095]. combtooth [2838]. come [2530, 2974]. coming [2160]. Comment [2248, 290, 2737, 1516, 2736, 2711, 993, 1515, 2368, 2251, 2250, 2249, 2367]. comments [385, 1315, 1304, 567, 1308, 818, 2608, 2549, 2004, 2401, 1757, 2003, 731, 2400, 1247, 2380]. commercial [2588, 781, 2199, 2046, 2447, 778, 2911, 803]. commercially [1301, 2040, 1309]. commissioning [1226]. Common [1598, 137, 1140, 1864, 598, 829, 1570, 1947, 1353, 2162, 1944, 216, 718, 1998, 1054, 1237, 97, 913, 896, 269, 543, 116, 1261, 1365, 2165, 2302, 2505, 1909, 1206, 948, 2561, 1999, 808, 2293, 3060, 1881, 3075, 1729, 2344, 677, 339, 146, 1002, 1964, 2961, 2807]. commonly [3058]. commons [2452]. communal [2885]. Communication [635, 1098]. communities [2638, 2330, 1345, 324, 474, 2448, 651, 250, 1612, 2712, 628, 2350, 1133, 131, 2253]. Community [396, 631, 675, 1531, 560, 2136, 342, 724, 2071, 2374, 1538, 3000, 1988, 1980, 1213, 255, 1116, 2284, 1180, 708, 2715]. community-based [1988]. Compagno [1270]. company [897]. Comparative [1086, 1164, 2808, 1716, 1755, 1046, 1443, 2013, 2833, 12, 543, 1904, 2268, 2401, 1377, 2480, 1714, 1633, 1165, 2555, 1309, 1160, 2771, 2305, 499]. compare [3009]. compared [1858, 1408, 2159, 41, 2654]. Comparing [333, 2525, 948]. Comparison [2708, 1864, 2491, 2598, 2854, 1660, 951, 127, 915, 953, 2350, 1853, 2363, 2177, 203, 1862, 178, 2625, 2844, 2513, 1612, 2505, 2115, 1818, 318, 1924, 36, 2881, 2995]. Comparisons [1845, 1008, 836, 1438, 1916, 2646, 1691, 773, 1777, 1518]. compatibility [560, 218]. Compensatory [963, 1016, 2578]. competition [655, 636, 50, 406, 2787]. competitive [1667]. complete [2566, 1960, 2694, 19]. Complex [2130, 2577, 2817, 1437, 1470, 43, 1937, 330, 2051, 1916, 1357, 1880, 246, 613, 1161, 939, 843, 2357, 2553, 1150, 2332, 1162, 1502, 1537, 530, 1002, 540]. complexes [151, 938]. complexity [686, 2238, 2330, 735, 565, 1433, 847, 1153, 2973]. complicated [2576]. component [709]. components [301, 933, 3089]. composed [2201]. Composition [250, 1217, 1234, 2588, 631, 2615, 2398, 1623, 31, 1445, 1600, 2137, 2370, 1011, 1914, 962, 663, 1183, 151, 17, 1680, 2302, 2505, 1950, 2477, 2010, 619, 2546, 776, 1362, 846, 1913, 1991, 765, 806, 842, 653, 2667, 136, 193, 1898, 778, 1775, 255, 230, 2283]. compositions [56, 512, 88, 2870, 1139]. Comprehensive [461, 2161]. compressa [363]. compressed [343]. comprised [1690]. comprising [1131]. compromise [1093, 402]. compromised [2981]. computational [2270, 1718]. computed [759, 2599]. concentration [819, 1800]. concentrations [1524, 1795, 3012, 2566, 1400]. concentricus [2428]. concept [1827, 640, 875]. conceptionis [1263]. concern [1546]. Concerns [289]. concolor [1643]. concurrent [2831]. Condition [2770, 1465, 2523, 2786, 31, 1030, 1681, 1486, 1691, 1771, 467, 2744, 1536, 2550, 1016, 2131, 95, 2775, 708, 171]. conditioning [2992, 424]. conditions [2622, 69, 2593, 2737, 1555, 853, 2258, 1337, 2736, 546, 1900, 2907, 811, 2430, 307, 1605, 683, 1793, 1346, 1197]. cone [666]. confined [271]. confirm [1737, 3061]. confirmed [2241, 1096, 2820, 2099]. confirms [828]. conflicts [1558]. confluence [1985]. confluentus [2818, 1593, 1891]. Confocal [1547]. conformer [2815]. conforming [966]. Confronting [548]. congeneric [2239, 659]. congeners [1476, 2608, 2549]. Congo [1541]. Congress [1212, 139]. Congridae [385]. Congriscus [2127]. congruent [2305]. Connecticut [1498]. Connecting [726]. connection [400]. connective [2784]. Connectivity [449, 1737, 1627, 3088, 2738, 848, 1019, 496]. Consecutive [1861].

Consequences [2613, 323, 867, 330, 1561, 1610, 187, 1900, 3073]. Conservation [152, 463, 487, 506, 526, 289, 1995, 812, 1896, 956, 2160, 441, 139, 494, 356, 1404, 2143, 3077, 2817, 2156, 2532, 156, 1973, 745, 361, 2007, 2199, 1817, 153, 161, 1227, 1602, 1974, 749, 158, 640, 2035, 2494, 632, 783, 2991, 1992, 780, 876, 1762, 774, 1978, 1983, 1048, 1551, 1496, 1982, 1675, 2003, 332, 938, 1816, 2158, 213, 2205, 635, 2036, 2479, 639, 2014, 1975, 2251, 747, 637, 1803, 1198, 159, 2250, 784, 1027, 155, 741, 3028, 154, 2690, 549]. conservatism [2731]. conserved [2884]. conserving [1215]. consider [2913]. consideration [2348]. considerations [2149, 571, 2864, 850]. considered [2839]. Consistency [2593, 2737, 2736, 2406]. Consistent [2466, 1169, 2192]. conspecific [2187, 2306, 2258, 51, 2670, 513]. conspecifics [1966, 367, 811, 1153, 41]. Constance [270, 1679, 1410, 1624]. constant [1506, 1820]. constituents [2204]. constraint [1432]. constraints [394, 2062, 2085]. construction [1666, 3052, 416]. consume [309]. consumption [1696, 2880, 2214, 1364, 2917, 1480, 1818, 1355, 2289, 1184, 2675, 990, 1820, 2582]. contain [2889]. containment [2708]. contaminant [2720]. contamination [598, 2975, 1838]. Contemporary [2553, 466, 2602, 863, 496]. content [2615, 926, 453, 1508, 1670, 817, 3010, 1615, 125, 230, 1375]. Contents [185, 358, 3067, 854, 3066, 1028, 225]. context [1788, 1422, 540]. contexts [2043, 583]. continental [156, 2335, 1297, 930, 2912]. continuous [2665, 2258, 443]. continuum [1784]. Contraception [2819]. contracting [23]. contrast [1955]. Contrasting [1112, 1637, 795, 684, 2525]. contrasts [2901]. contribute [3053, 1263, 2148]. Contribution [1691, 2528, 1670, 1430, 2989, 458]. contributions [2918, 2146, 55, 1957]. Control [3017, 2211, 2391, 286, 7, 951, 2061, 395, 2353, 264, 2232, 2691, 2153, 1499]. control-region [264]. controlled [2622, 2847, 2112, 1244]. controls [2813, 2]. controversy [1802]. conventional [766]. conversions [3033]. convict [1539, 2713]. Cookie [2554]. Cookie-cutter [2554]. cool [2994, 1678]. cool-water-range [1678]. Cooperation [1228, 1222]. copepod [1399]. copepods [2503]. Coping [116, 173, 2846, 1881]. copper [51, 1379]. Coptodon [1905]. copulatory [751]. coral [1350, 587, 2067, 976, 179, 1353, 2655, 1356, 2330, 2093, 409, 2599, 2735, 1187, 1762, 1305, 2942, 2075, 2069, 1152, 1189, 953, 2801, 1132, 1673, 1078, 971, 827, 1075, 2556, 2798, 2652, 1609, 1938, 1534, 285, 1938]. coral-dominated [1075]. coral-dwelling [179]. coral-fish [1673]. coral-reef [587, 976, 2655, 2330, 2093, 2599, 2942, 1152, 1189, 1078, 971, 2652, 1609]. coralgrouper [1087]. corallivores [1075]. Corallivory [52, 179]. coregonid [1679, 1610]. Coregonus [1849, 2090, 29, 350, 699, 1357, 1410, 851, 1117, 381, 2641, 1033, 1240, 2553, 1910, 2225, 936, 307, 1750, 399, 362]. Coreius [734, 2562]. Cormack [2465]. corniger [1318]. cornutus [2807]. Correct [465, 2247]. correction [1585, 665]. correlate [1591, 1936, 199]. correlated [922, 2846, 115]. Correlates [2546, 2718, 344, 2131]. Correlation [590, 2510]. correlations [2043, 608, 735]. correspond [2990]. correspondence [899]. Corrigendum [143, 209, 312, 440, 673, 704, 832, 957, 1128, 1155, 1320, 1972, 2105, 2519, 2632, 2895, 3041]. corroborating [339]. corruscans [1203]. Corsica [1569]. Corsican [1281]. corticoid [1513]. cortisol [68, 2939, 2491, 922, 1365, 1313, 1963, 2862, 1847, 241, 1512, 2311, 1073, 2848, 2863, 164, 1885]. coruh [2851]. corvina [2375]. Corydoras [2587, 2956]. Corynopoma [1181]. Coryphaena [2029, 1617]. Coryphaenoides [2409]. cosmopolitan [665]. cost [59, 2680]. Costa [2617, 940, 1011, 724]. costs [2516]. Cottidae [277, 2835]. Cottiformes [2082]. Cottunculus [1481]. Cottus [1618, 1681, 2130, 479, 277, 2835]. could [2306, 2332]. Coulter [955]. count [2211]. Counterintuitive [2656]. counts [2990, 915]. course [1047, 1449, 1282]. court [834]. courting [2957]. Courtship [2676, 277]. covariation [404]. cover [2330, 1591, 913, 1566, 721]. cownose [2885]. CpG [1291]. Cranial [999]. crassus [1351, 580]. crayfish [1610]. creatine [84]. Creatures [111]. creek [1073, 2739]. creeks [2013, 2010, 2120, 2745, 383]. Cren [641]. Crenicichla [1138]. Crenuchidae [371, 3080]. crescent [2082]. cresthead [1598]. crisis [775]. criteria [1120]. Critical [1491, 67, 3067, 2472, 1062, 3003, 3066, 1416, 2310, 2694, 869]. critically [1807, 1686, 2820, 2602]. croaker [2884, 996, 35, 2112, 2273, 1394, 1736]. Croatia [1927]. crocea [996, 1394, 1736]. crocodile [754, 1996, 121]. crocodilus [1407]. cromis [911, 2667]. cross [954, 1703, 2975, 830, 972]. cross-amplification [1703, 830, 972]. cross-contamination [2975]. cross-species [954]. crosses [1679]. crowding [1023]. Crozier [606]. crucian [2774, 637, 3087, 825]. cruentatus [1760]. Cruriraja [2380]. crusma [2970]. crustacean [821]. crustaceans [1943]. Cryopreservation [2072]. Cryptic [2917, 2108, 914, 2950, 1601, 2129, 1216, 2165, 1467, 2801, 2411, 1724, 2294, 384]. Cryptobenthic [735]. Cryptocaryon [1736]. crysos [979, 1460]. CT [3005]. CT-scan [3005]. CT1 [84]. CTAB [1651]. ctenii [388]. Ctenogobius [809]. Ctenoluciidae [2360]. Ctenopharyngodon [2809, 717, 2058, 2778, 668, 1867, 2671, 1638, 1241, 437, 2589, 826, 1106]. cuckoo [919]. cue [2102, 1657]. cues [2939, 564, 253, 302, 835, 2187, 732, 1389, 2258, 888, 904, 1090, 1395, 1171]. cultivated [2110]. Culture [812, 981, 908, 716, 2575, 2454]. culture-based [716]. cultured [2832, 792, 2059, 104, 2402, 280, 803, 1793, 850, 1520, 1690]. cumulative [873]. cuphead [2408]. curema [388]. curemai [1092]. Curimatidae [2514, 2376]. Curimatopsis [2514, 2376]. curimbas [1092]. Current [463, 487, 506, 526, 49, 3, 441, 2941, 1263, 3033, 741, 634]. currents [1705]. curvatures [1697]. cusk [1623, 420]. cusk-eels [420]. cutaneous [1053, 1332]. cutter [2554]. cutthroat [1236, 2747, 2756, 815, 1504]. cuvier [977, 2372, 1892, 1740]. cuvieri [2360]. cyanellus [1586]. cyanobacterial [97]. cycle [976, 1407, 2141, 200, 732, 757, 224, 1980, 1979, 2218, 758, 547]. Cycleptus [53]. cycles [521, 343, 233]. cycling [195, 2552]. cyclone [232]. Cyclopteropsis [1640]. Cyclopterus [2832, 2154, 1423]. cyclostomes [286]. cylinder [684]. Cymatogaster [1828]. Cymodocea [1781]. Cynoglossidae [1467, 2826]. Cynoglossus [2826, 82, 380, 1568, 2476, 1931]. Cynoscion [2335, 2375, 2219, 1061, 467, 1446]. cyp19a [1457, 1291]. cyp19a1a [2237]. cyp1A [850]. cyprinid [2725, 43, 918, 1778, 2594, 2272, 557, 36, 2504]. Cyprinidae [881, 1437, 598, 61, 1880, 2287, 1359, 1951, 1477, 229, 1754, 1906, 2542, 1541]. cyprinids [65, 1421, 368, 683, 1906]. Cypriniformes [267, 836, 336, 2287, 2411, 2739, 2504, 1541, 733]. Cyprininae [1906]. Cyprinodon [1753]. cyprinodont [1627]. Cyprinodontidae [1055]. Cyprinodontiformes [711, 2577, 1894, 2839, 1896, 937, 103, 1004, 1770]. Cyprinus [2221, 1570, 1947, 1944, 216, 718, 1998, 2086, 2167, 1237, 97, 913, 269, 116, 1261, 1365, 2302, 1137, 948, 1999, 808, 531, 1881, 1729, 2344, 677, 339, 369, 146, 1002, 1392, 1964, 2961]. cysteine [1244]. Cytochemical [2620]. cytochrome [485]. Cytogenetic [714, 2108, 523, 890]. cytogenetics [1755, 458]. cytoglobin [1606]. cytokine [566]. cytometric [802]. cytometry [2614]. cytotoxic [1574]. Czech [1618].

D [289, 1018, 1126, 110, 1269, 1270, 3016, 859, 739, 138, 438, 1266, 799, 839]. D-loop [799, 839]. dace [2224]. dactyliophorus [500]. dactylopterus [352, 352, 2307]. daddies [2585]. daemelii [1183]. Dahl [2816]. Dahl-Lea [2816]. Daily [2980, 2678, 40, 2955, 1367, 1721, 2266, 1731, 592, 1480, 1512, 720, 2289, 1578]. Dallia [1319]. Daly [633]. dam [800]. damage [2598, 1068, 1389, 2258, 1066]. damage-released [1389, 2258]. damaged [2726]. dams [1666, 260, 1527]. damsel [2300]. damselfish [2970, 2540, 827, 2787, 1639, 2992]. damselfishes [300]. danger [1583]. Daniels [908]. Danio [1702, 2487, 68, 2939, 2813, 1798, 2785, 1443, 1469, 1114, 2415, 2529, 2804, 1251, 2292, 1358, 1838, 1917, 3053, 2131, 1066, 2552, 1347, 228, 84, 1958]. Danube [557, 2419]. Danubian [2740]. d’Arguin [1905]. dariensis [2006]. dark [2725, 1928, 1790, 2473, 2341, 2936]. dark-adapted [2936]. darter [1091, 1250, 2516, 1383]. darters [2078, 1787]. Darwin [2796, 2473]. dash [2243]. dash-and-dot [2243]. Dasyatidae [2853, 773, 1132]. Dasyatis [2123, 1602, 772]. data [643, 970, 271, 2042, 533, 781, 879, 1345, 1892, 1811, 3049, 1314, 1516, 2559, 1630, 1040, 1969, 1565, 2176, 1095, 1515, 808, 1971, 2108, 397, 1714, 630, 2462, 1492, 1201, 885, 255, 2504, 1938, 656, 1958, 890, 1594]. data-deficient [1971]. data-storage [879]. date [1360]. dates [945]. David [641, 894]. Davis [200]. Day [998, 1702, 343, 1510, 204, 1963, 2694]. day-length [343]. day-long [204]. days [2664]. Dead [872]. Deania [753, 1761]. Debelius [141]. debris [2262, 2010, 1204]. decadactylus [395, 2913]. Decadal [2748, 1518]. Decadal-scale [2748]. decadence [11]. decades [1961, 2805]. decay [131]. decision [2459, 1148]. decisions [1418]. decline [361, 1117, 1242, 637]. declines [327]. Deconstructing [2956]. decreased [274, 1990]. decreases [2516, 2065, 2485]. Deep [113, 2809, 1530, 2912, 1108, 1271, 580, 2717, 1396, 769, 1623, 1276, 2299, 395, 1760, 2063, 1612, 3036, 2257, 2960, 2738, 1214, 1480, 2797, 1275, 1815, 2409, 1217, 1790, 1273, 1682, 1887, 308, 906]. Deep-sea [1530, 1396, 769, 1623, 1276, 2299, 395, 1612, 3036, 2257, 2960, 1480, 308]. Deep-water [2912, 1271, 2717, 2063, 2738, 1214, 1815, 2409, 1217, 1682, 1887, 906]. deepwater [2101, 1970]. defence [1325]. Defences [109, 1068]. Defenses [568]. deficiency [1412]. deficient [1971, 1856, 668]. define [1182]. defined [761]. definitions [1817, 868]. deformities [1292, 280, 2445]. deformity [2339, 2599]. deglaciated [2352]. degradation [2735, 2552]. degree [2700]. degrees [2047]. DEHP [2481]. Dekeyseria [2586]. del [1894]. delaisi [2318]. delay [2881]. Delayed [2751, 1900, 3073]. deleterious [997]. Delimitation [2789, 1685]. delimitations [2673]. delineate [948]. delineations [1796]. delipidation [2508]. delivery [1520]. delta [2023, 2262, 3086]. demersal [2025, 1271, 1054, 1276, 1189, 927, 895, 2022, 95, 2021]. demineralization [2508]. Demographic [2811, 1199, 2446, 1485, 298, 2223]. Demographics [2817, 1883]. demography [1810, 273, 1309]. demonstrate [1143]. demonstrated [1820]. densities [2735, 3024, 1885]. Density [2031, 2186, 686, 444, 1618, 2625, 1956, 1508, 50, 1682, 1535, 1154, 128, 1782, 2981, 126, 1036]. Density-dependent [2031, 2186, 2981]. dental [2006]. dentary [2203]. dentatus [195, 898, 705]. dentex [1864, 1864, 1286, 2288]. Dentition [676, 2424, 2319]. deoxynucleotidyl [2309]. Dependence [2012, 584, 1782]. dependency [737, 3035, 1734]. dependent [2459, 40, 2031, 1178, 2510, 2403, 2186, 3014, 2751, 1017, 2800, 649, 415, 1379, 654, 228, 2981, 857, 165, 2654]. depends [546, 618]. depicting [2328]. depleted [1432]. Deposition [251, 1629, 2266, 2477, 1936, 1578, 252, 339]. depositional [2084]. depots [467]. deprivation [402, 688, 307]. Depth [2134, 735, 2435, 713, 2018, 1893, 2354, 436, 2377, 2928, 1898, 1604]. depths [106]. Derivation [266, 2575]. derived [781, 1567, 3059, 1195, 265, 2774, 1529, 2507, 2088, 1290]. Dermal [1800]. describe [1547]. described [2002, 1569, 61, 1391]. Description [2319, 1467, 2414, 2749, 2645, 2577, 2811, 1590, 2282, 2509, 1088, 2587, 2051, 2075, 2212, 243, 586, 1004, 420, 2263, 591, 1906, 2445, 2138]. descriptions [246, 644, 1161, 2267, 2004]. desert [1302]. desiccation [2403]. Design [1819, 2442]. Deslorelin [2819]. DeSombre [1173]. despite [2165]. destinations [346]. destructive [1300, 1949]. details [2111]. detectable [992]. detected [21, 1255, 1243]. detecting [961, 2089, 1594]. Detection [808, 944, 69, 854, 2933, 1442, 1700, 1893, 1106]. determinant [1847]. determinants [3056, 349, 144]. Determination [1378, 1822, 2665, 2785, 196, 1253]. determine [2783, 2615, 453, 697, 2952, 366]. determined [766, 282, 234, 226, 232, 2071, 3026]. determines [2909, 976, 1357]. Determining [69, 1229, 1930, 2394, 2525]. deterrent [2935]. deterrents [2934]. detritivore [66]. detritivorous [2748, 1986, 885, 2113]. Detroit [26].

Devario [1384]. Developing [1546]. Development [694, 954, 836, 338, 1846, 1258, 1585, 1762, 443, 2484, 1500, 380, 2898, 1511, 1290, 2276, 2834, 710, 2722, 1285, 2764, 2622, 2399, 29, 1188, 902, 1023, 2510, 2592, 511, 2677, 2931, 1567, 1994, 1995, 127, 1521, 2778, 486, 1564, 2700, 1561, 2267, 187, 1797, 893, 292, 1581, 1142, 1918, 1442, 2826, 2401, 2557, 2781, 1509, 2623, 2644, 529, 1843, 3000, 2236, 2580, 1449, 191, 2347, 280, 478, 2114, 2669, 2169, 1370, 2318, 1750, 2088, 2273, 1840, 1333, 706, 618, 1793, 1156, 2001, 1390, 2890, 3004, 2486, 237, 439]. development-related [1521]. Developmental [2880, 968, 2832, 1412, 1335, 404, 2977, 2678, 2618, 2237]. developments [876, 984]. devices [1821]. devidei [2606]. devil [2242, 2779, 1207]. devilray [2968]. Dhabi [776]. Di-2-ethylhexyl [2481]. diadromous [876, 848, 1342, 2107, 883]. diagnosis [1906, 3017]. diagnostic [2324]. Diagramma [575]. Diamantina [1315]. diameter [62]. diamond [1117]. Diaphus [1060]. Diapoma [2743]. Dicentrarchus [797, 452, 1655, 1259, 419, 1292, 2583, 1647]. dicephalia [1094]. dicephalic [432]. Dicephalous [2180]. dicer [1241]. dichromatic [2437]. dichromatism [2437]. dictionary [2144]. DIDSON [582, 2958]. dieffenbachii [1629, 2353]. Diel [879, 767, 2353, 1595, 1167, 990, 1912, 1116, 2253, 2090, 195, 1011, 1959, 394, 2779, 811, 2745, 283, 916]. diel-cycling [195]. diel-migrating [1959]. Diet [1278, 520, 1663, 1445, 534, 1011, 202, 771, 1573, 163, 1003, 240, 1986, 900, 136, 824, 1159, 2438, 627, 1439, 1234, 2047, 1350, 1166, 1344, 1138, 1590, 453, 119, 52, 1602, 1660, 770, 1613, 55, 2998, 1592, 192, 1311, 427, 2142, 2288, 663, 1015, 2312, 474, 1681, 1233, 1089, 1612, 2503, 483, 187, 560, 2120, 1028, 172, 2071, 1016, 456, 846, 1913, 105, 1656, 239, 537, 1948, 1566, 1009, 454, 931, 959, 2715, 2721]. Diet-tissue [1986]. Dietary [1636, 2370, 101, 896, 817, 1132, 2667, 646, 926, 1245, 1112, 377, 2117, 940, 1409, 1431, 1508, 1408, 773, 512, 2477, 2704, 693, 88, 478, 2653, 1282, 1712]. Diets [179, 1632, 2359, 1970, 2544, 2447, 1054, 2646, 2800, 2917, 2477, 2951, 1698, 1986, 501]. differ [966, 1624]. difference [2727]. Differences [166, 2117, 1795, 850, 1385, 2702, 2041, 1752, 1580, 583, 1588, 2646, 2639, 1523, 1603, 1317, 1085, 1143, 1963, 2849, 2192, 36, 477, 2166, 705, 1197]. Different [1549, 2754, 2047, 1506, 1503, 1470, 468, 1751, 2214, 1636, 2056, 1168, 1746, 2026, 932, 1507, 2799, 2827, 2626, 512, 1608, 2477, 1655, 2807, 1107, 602, 2350, 2595, 1016, 2554, 1255, 2005, 280, 2132, 193, 3024, 1820, 499, 612, 905, 369]. Differential [1524, 1867, 2801, 2774, 1881, 2354, 3087, 212, 2237, 2664, 819, 619, 927, 199]. differentially [1713]. differentiate [354]. differentiated [3083]. differentiating [699, 1111]. Differentiation [1521, 303, 2150, 80, 795, 3048, 2059, 1035, 711, 849, 1205, 2402, 328, 621, 950, 144, 437, 2273, 1537, 2001, 890]. differing [2818]. differs [1503, 2060, 118, 2497]. digestion [66, 67]. Digestive [2340, 3057, 963, 2722, 1112, 2620, 402, 1388, 1522, 2871]. digestive-trait [1112]. digests [2670]. dihydroxypregn [9, 1175]. dimensional [2197, 1955]. dimensions [1220, 1780]. dimerus [1470]. dimidiatus [2190]. diminutive [2397]. dimorphic [2558, 2285, 1798, 147, 2983, 1291]. dimorphism [244, 1125, 1640, 1723, 1582, 682, 1417, 695, 2641, 1247, 2380]. Dining [887]. Diodon [647]. Diogenichthys [3054]. Diogo [461]. Dios [2549]. Dioscorea [247]. dioxide [2652, 2210]. dipeptide [3079]. Diphyllobothrium [2405]. Diplectrum [235]. Diplodus [2559, 2712, 419]. diploid [127, 1484, 3087, 2053]. diploids [1032]. diprosopus [2180]. Dipturus [2217, 2729, 771, 1573, 1887]. Direct [1419, 201, 2854, 953]. directed [203]. directing [888]. direction [2137]. directional [1482]. directions [2921]. disalvoi [2438]. discarding [778, 2911]. discards [1217]. discharge [1571, 299, 128]. discharges [2113]. disconnected [2419]. Discopyge [2424]. discovered [429, 1602]. Discovery [1686, 1056, 2809, 335, 2829, 1210]. discrete [2913]. Discriminating [797, 523, 1123]. Discrimination [386, 1403, 1899, 2048, 35, 1352, 1886, 2430, 1986, 530]. discus [407]. discussion [2531, 2749, 912]. disease [1916, 2691]. Diseases [3017]. dispersal [1949, 2200, 2055, 2974, 397, 307, 1341, 215, 41, 2886]. Dispersion [1998, 2517, 1059]. displaced [2124]. Displacement [2336]. display [897]. dissection [1262]. dissolved [1068, 274, 106, 2867]. Dissostichus [2356, 537, 2478]. distance [2296, 448, 2701, 992, 287, 1806, 3057, 78, 131]. distances [2886]. Distichodontidae [2233]. distinct [3048, 2137, 1017, 1549, 1624, 1537]. Distinction [1420]. distinguish [2165, 1107]. distinguished [2357]. Distinguishing [373]. distortions [1185]. distressed [2550]. distributed [1737, 2107]. Distribution [1351, 1611, 1018, 1183, 2741, 17, 2563, 1174, 1447, 2252, 3021, 438, 425, 2211, 582, 2865, 1271, 1344, 1666, 868, 290, 1710, 1580, 2027, 2262, 2011, 735, 1892, 2044, 899, 341, 1784, 2639, 2799, 175, 1923, 1040, 232, 1052, 791, 1417, 1788, 321, 1097, 2879, 973, 1058, 1274, 1815, 18, 660, 2036, 2062, 1979, 233, 349, 1201, 2755, 2021, 654, 1050, 2332, 2714, 1541, 1487, 2476, 1440, 804, 2793]. Distributions [1199, 2681, 451, 1152, 2870, 537]. Disturbance [2997, 1362, 2015, 2273]. disturbances [1984]. disturbed [1027]. disturbs [2481]. diterpene [247]. ditremum [2405]. Diurnal [2655, 227, 2179, 393]. diver [2854, 2044, 953]. Divergence [2308, 657, 85, 1141, 1163, 2559, 624, 1917, 1281, 2152, 1796, 947, 816, 2351, 612, 1055, 2565]. Divergent [37, 1013, 2060, 577, 552]. Diverse [866, 1632]. diversification [371, 2349, 458]. Diversity [113, 153, 1018, 1236, 2825, 347, 1950, 2164, 886, 346, 779, 2950, 2150, 964, 2248, 453, 1429, 1450, 2256, 800, 1410, 1306, 168, 2008, 1462, 1905, 1053, 1983, 2159, 264, 1788, 2824, 628, 1433, 2879, 1815, 2821, 2417, 2553, 967, 3072, 2022, 1217, 2989, 914, 2305, 670, 2407, 1879, 2249, 1876, 1031, 622, 2153]. divine [2810]. diving [2181, 2746]. division [83]. dmrt1 [2059, 1848, 1291]. DNA [2987, 1766, 657, 2950, 3027, 2150, 1720, 2673, 1437, 1300, 535, 1947, 453, 216, 81, 522, 2856, 2868, 2312, 1410, 849, 334, 3014, 1597, 386, 1442, 714, 2982, 1891, 2232, 799, 1651, 2161, 1913, 2225, 20, 3055, 1066, 2230, 1051, 363, 384, 1375, 225, 2838, 76, 952, 2752, 212, 622, 2835, 19]. Do [2047, 2228, 966, 1655, 1376, 2467, 2944, 2990, 732, 366, 676, 1668, 505, 1263, 2098, 2209, 525]. Doce [3005]. Dodge [994]. Dodson [139]. doellojuradoi [1907, 2319]. Does [1181, 2097, 659, 819, 1563, 1914, 1591, 962, 1248, 2087, 393, 684, 1831, 2080, 693, 827, 2867, 810, 1256, 405, 1616, 2931, 1933, 402, 133, 1771, 1838, 417, 273, 2652, 2699, 2669, 2053, 1102, 713, 483, 173]. dog [1338, 163, 45]. dogfish [792, 760, 1935, 2135, 3011, 541, 665, 1761, 2911, 2721]. dojo [2411]. doliatus [1144]. Dolina [2422]. dollar [1463]. Dolly [1524]. dolomieu [841, 12, 2200, 455, 1615, 2305]. dolphinfish [2029, 1617]. Domenici [588]. domestic [2719, 2053]. domesticated [1764, 2087]. Domestication [1020]. dominance [603, 2306, 2535, 417, 2805, 284]. dominant [998, 2610, 885]. dominate [1075]. dominated [2142, 1075]. Donor [2187]. dopamine [7]. Dorosoma [1670, 92]. dorsal [760, 1305, 2302, 1419, 2891, 1579]. dorsal-fin [760]. dorsalis [2585]. dorsomaculatus [222]. doses [1866]. dot [2243]. dotted [818]. double [2599, 826]. double-stranded [826]. double-tailed [2599]. down [1898]. Downstairs [1763]. downstream [400, 864, 2727, 2933, 433, 1500, 1290, 1505, 2329, 1527, 2102, 1499]. Drag [2860, 1545]. dragged [2493]. dragonet [1517, 896]. drain [1883]. drainage [601, 557, 2887, 3005, 30, 2780, 2851]. drainages [269]. draughtboard [766]. dredge [201]. dredging [2025]. drift [226, 1595, 2102]. drifting [1873]. driven [2020, 762]. driver [1444]. Drivers [2952, 2140, 2179, 327, 2852, 542]. driving [2735]. drought [804]. drought-affected [804]. drum [1557, 911, 1936, 2396, 2065, 1656, 1924, 2667]. dry [55, 1016]. DST [879]. dual [582]. dualism [990]. duck [2231]. duct [118]. due [598, 612]. Dules [1661, 2404]. dumeril [123, 2418]. dumerili [3062]. dummy [498]. Dunckerocampus [500]. duplication [571]. duration [2701, 696]. Durbin [2004]. during [2832, 2722, 1386, 727, 2742, 2264, 2955, 354, 2399, 1710, 534, 1732, 1514, 2994, 2510, 2059, 2592, 992, 2677, 2931, 574, 134, 1593, 991, 2141, 874, 1735, 2778, 3078, 175, 559, 880, 1680, 756, 277, 1836, 1084, 958, 1911, 3057, 1104, 1789, 2644, 1355, 1484, 842, 1500, 1980, 191, 2552, 2114, 144, 1511, 869, 67, 1793, 1332, 2329, 959, 1349, 2584, 1482, 2237, 539, 1931, 480]. Dusky [1639, 1739, 254, 1188, 1651, 1025]. duskytail [2730]. dwarf [2869, 1260, 2948]. dwarfgoby [2810, 2094]. dwelling [179, 645, 966, 2544, 2215, 2470, 965, 1865, 2653, 552]. dyes [2624, 1768]. Dynamic [2086, 1534, 1824, 2874]. dynamics [1952, 2270, 1353, 1718, 330, 2416, 920, 2795, 3066, 3030, 1091, 2300, 1366, 374, 2136, 467, 1977, 313, 497, 2610, 3076, 1369, 2289, 842, 2730, 1985, 2507, 1342, 125, 1428, 2715, 2478]. Dysalotus [2212].

E.U [2162]. E2 [1647]. eagle [2439, 2226, 1861, 1067, 2819]. Early [710, 446, 473, 278, 2320, 1434, 1142, 2623, 407, 738, 2618, 1903, 1390, 2346, 2722, 2450, 1245, 1367, 855, 819, 2482, 2059, 1672, 1795, 574, 134, 1994, 1293, 872, 1957, 2266, 1598, 1583, 2267, 118, 3010, 1034, 1292, 1509, 1501, 1572, 2753, 554, 892, 481, 2114, 1370, 987, 1256, 1692, 203, 1102, 2329, 1197, 3054, 480, 405, 462]. early-life [1957]. early-stage [1102]. ears [1415, 285]. earth [2341]. easier [151]. East [1268, 389, 2882, 643, 484, 2122, 306, 2450, 1479, 755, 1481, 2526, 2019, 1345, 1811, 748, 2853, 28, 519, 1740, 351, 2045, 937, 895, 1815, 1045, 1786, 2806, 1154, 159, 1487, 2855, 2799, 624, 2374]. east-central [937]. Easter [2438]. easterly [973]. Eastern [1209, 2360, 1042, 1234, 887, 779, 2733, 2728, 2640, 2673, 364, 2339, 2910, 254, 2839, 770, 2129, 2126, 761, 915, 269, 114, 1630, 1934, 2130, 250, 347, 2168, 2029, 1909, 556, 1419, 1352, 757, 731, 270, 1617, 379, 1446, 1991, 1932, 2745, 2556, 3074, 1887, 1722, 2033, 2914, 2178, 1447, 670, 298, 758, 1307, 1123, 1180, 1860, 1622, 2358, 2729]. eat [398]. eating [147]. eats [2670]. Ebert [1270]. Echeneis [2733]. Echinorhinidae [1234]. Echinorhinus [1234]. Eco [588, 2170, 1428, 2559]. Eco-Ethological [588]. Eco-evolutionary [1428, 2559]. Eco-morphological [2170]. ecocline [1613, 2014]. Ecological [780, 3056, 1660, 1194, 350, 2689, 791, 691, 2373, 2718, 1816, 1220, 2565, 1664, 1126]. ecologically [329, 1376]. Ecologist [508]. ecologists [685]. Ecology [2069, 2315, 2922, 138, 2658, 137, 1741, 1696, 2640, 745, 1929, 1882, 1438, 1590, 2146, 482, 1478, 392, 809, 770, 749, 2461, 1445, 1613, 377, 1834, 940, 1407, 2260, 1743, 518, 2920, 2173, 1187, 915, 652, 1559, 773, 2298, 1791, 2894, 483, 822, 342, 966, 772, 326, 1725, 198, 2938, 300, 2985, 2923, 1926, 599, 1754, 2480, 2050, 750, 554, 2259, 1803, 1272, 2252, 1060, 1912, 1770, 1139, 1180, 1875, 671, 1550, 1267, 605, 690, 672]. Ecomomy [981]. ecomorphology [1759]. economic [2991, 327]. economical [1056]. ecoregion [2673, 2806]. ecoregions [1991]. ecosystem [1438, 640, 2766, 2173, 2472, 2902, 3033, 2534, 2030, 399, 2912, 548]. Ecosystem-based [548]. ecosystems [745, 1612, 1982, 2022, 2798, 2690]. ecotypes [1288]. ectoparasitic [2704]. Ectopic [2943]. Ecuadorian [1813]. Ed [3016]. eddsi [100]. Eddy [1126]. Edge [289, 401, 2139]. Edited [289, 112, 1267, 955, 908, 605, 1018, 1126, 550, 462, 140, 1265, 549, 812, 690, 983, 1269, 1270, 525, 1070, 604, 831, 1099, 1211, 568, 182, 956, 588, 859, 1127, 109, 183, 739, 439, 671, 139, 138, 356, 311, 438, 548, 672, 389, 980, 981, 982, 1266, 1098, 461, 569]. Edition [1172, 2193, 2068, 108, 141, 2570]. Editorial [1129, 2457, 2635]. eDNA [2089, 1801, 2434]. Eds [3017]. eel [866, 719, 1705, 1855, 563, 580, 361, 2759, 448, 473, 2496, 2579, 2187, 1521, 541, 2267, 2066, 1837, 1711, 426, 865, 2387, 3026, 2115, 1495, 1674, 515, 2815, 578, 863]. eelgrass [504]. eelpouts [1589]. eels [2882, 1629, 1616, 1623, 1755, 1883, 402, 864, 2590, 2727, 2639, 90, 3044, 853, 2741, 2353, 1085, 2489, 1411, 420, 1038, 2872, 1197, 289]. Effect [2528, 1702, 1316, 444, 2640, 2362, 2303, 1580, 855, 1364, 913, 1413, 658, 44, 57, 1034, 2901, 2550, 3062, 1619, 1828, 125, 2775, 815, 963, 737, 68, 293, 1470, 926, 411, 2609, 829, 2220, 2662, 59, 1735, 941, 2064, 1313, 1545, 50, 949, 204, 893, 845, 2803, 2174, 433, 2902, 1362, 2028, 408, 1472, 2547, 1336, 314, 653, 538, 3075, 252, 1229, 2695, 997, 2485, 2936, 319]. effective [2532, 667, 851, 2982, 379, 2148, 3072]. Effectiveness [854, 1305, 475]. Effects [2321, 400, 2523, 1166, 2025, 2183, 2508, 1146, 2189, 2773, 1478, 3035, 3031, 1054, 2785, 692, 1833, 1409, 1224, 3063, 3012, 1555, 97, 793, 934, 355, 343, 2219, 1510, 1508, 589, 1062, 1493, 853, 51, 1365, 232, 838, 1747, 553, 840, 2176, 2804, 1581, 1977, 38, 734, 1963, 2707, 3081, 2574, 2704, 92, 1073, 1509, 1501, 2265, 1082, 1078, 936, 307, 2848, 609, 2787, 1566, 2957, 721, 2703, 128, 1782, 164, 1080, 2706, 1967, 1794, 3089, 126, 305, 1607, 686, 1415, 2528, 40, 2845, 80, 1421, 498, 1353, 1384, 1068, 31, 2892, 6, 69, 986]. effects [316, 572, 1519, 324, 2072, 170, 1113, 2375, 917, 2297, 334, 1238, 1681, 2186, 2181, 2440, 1441, 177, 1680, 873, 1561, 3020, 2927, 1792, 2046, 894, 2489, 455, 1565, 1134, 2436, 562, 772, 321, 1433, 2698, 456, 379, 89, 318, 201, 1381, 314, 2479, 3034, 1683, 247, 1763, 1980, 2410, 2354, 2850, 436, 2352, 3024, 325, 2814, 1119, 1750, 1025, 2903, 2210, 1159, 298, 1990, 2912, 2776, 171, 916, 3008, 3058, 165, 1621, 2118, 2871]. efficacy [1316, 620, 206, 2614, 2029]. Efficiency [1594, 2854, 69, 2415, 135, 1362, 857]. efficient [1409, 2489]. effluent [544]. effort [781, 1345, 1961, 1473]. Egg [2717, 254, 2763, 2408, 2516, 2697, 256, 1490, 926, 2398, 1164, 819, 1023, 1914, 341, 503, 2063, 486, 573, 828, 2064, 1895, 911, 617, 1936, 2949, 180, 2850, 436, 502, 2283, 2493, 1090, 1395, 1000, 2835, 513, 42]. egg-guarding [828]. egg-laying [2063, 1000]. egg-spots [486]. eggs [710, 1310, 1285, 1871, 1620, 47, 962, 1785, 343, 854, 1610, 1651, 1240, 1707, 201, 2670, 804, 1400]. EIA [241]. eight [1203, 676, 625, 1492]. Einum [671]. ejaculate [1683]. ejaculates [2849]. Elacatinus [1609]. elasmobranch [2189, 1804, 372, 780, 1785, 1303, 2977, 2819, 756, 743, 776, 762, 744]. Elasmobranchii [1533, 432, 2859, 1733, 772, 456, 2038, 211, 2037, 2234, 2424, 1207]. Elasmobranchs [463, 487, 506, 526, 441, 410, 2856, 2208, 751, 768, 472, 791, 790, 2202, 788, 1742, 1815, 750, 1887, 1803, 784, 741, 2769]. elastomer [2465, 1684]. electric [2925, 991, 2772, 2353, 731, 2037, 2618]. electro [1253]. electro-immobilization [1253]. Electrocommunication [991]. electrofishing [3049, 1072]. electrogenesis [2925]. electrolocation [992]. electromagnetic [876]. Electroreception [2925, 2926]. electroreceptors [791]. electrosensory [2833]. Electroshock [62, 2613]. Electroshock-induced [62, 2613]. eleginoides [2356, 537, 2478]. Eleginops [651, 2932]. element [2508, 2137, 962]. Elemental [2430, 2203, 2202, 2546, 2406, 882]. elemental-isotopic [2546]. Elements [2202]. eleotrid [1921]. Eleotridae [852]. elephantfish [2990]. Eleutheronema [430, 2328]. elevata [1599]. Elevated [29, 2210, 204, 2929, 2652, 2706, 319]. elevation [1308]. elicit [1684]. elongata [2291]. elongate [1275]. elongation [500]. elongatus [1134, 53, 2230]. Elopomorpha [2686]. elucidation [1425]. elusive [2461, 532]. Elwin [2956]. embayment [474, 449, 283, 1601]. embayments [764]. Embiotoca [2957, 122]. Embiotocidae [552]. embryo [62, 1396, 2968, 1847, 1124, 2580, 2180, 1074, 1728, 2486]. embryogenesis [958, 2584, 480]. embryonal [472]. Embryonic [902, 2557, 2970, 1750, 1330, 29, 266, 404, 2931, 177, 893, 2884, 191, 2218, 1333, 1102]. embryos [62, 1330, 1023, 2415, 1900, 3073, 495, 2038, 247, 2062, 997, 1063]. emerald [2929]. emergence [1170, 2818, 1700, 1063]. Emerging [1221, 1708, 1231, 1743, 775, 1651]. emigrating [2353]. emigration [1572]. Emirate [776]. emission [3014]. emperor [1254, 340, 96, 1157]. emphasis [1556, 156, 1286, 1348, 969]. Empirical [978, 806, 685, 1594]. enable [1651]. encaustus [711]. enclosed [2011]. encoded [1738]. encounter [397, 1535]. encountering [1674]. encounters [68, 2098]. encourage [904]. encrasicolus [136]. Encyclopedia [739]. encysting [989]. end [1488, 1344, 1976]. endangered [2221, 2024, 2625, 99, 2551, 1807, 519, 3083, 1909, 2327, 2718, 1686, 1754, 379, 1897, 2471, 1763, 2114, 2965, 2716, 1243, 3077, 2820, 2602]. Endemic [1927, 2388, 881, 2235, 3050, 1319, 2624, 2093, 1601, 2428, 1057, 557, 607, 1796, 2666, 1887, 538, 1722]. endemism [964, 99]. Enderby [1371]. endocrine [286, 51, 2535, 87, 683]. endocrinology [493]. endoreic [601]. endothermy [3007]. endurance [2492]. enemy [2683]. energetic [2680, 2547]. energetics [820, 2828, 12, 1836]. Energy [2141, 323, 2296, 309, 58, 876, 1902, 1615, 1830, 1240, 1818, 1637, 990, 1154]. enforcement [786]. engage [1227]. England [2527, 1449]. English [2917, 1352, 1473, 637, 2911]. Engraulidae [1620, 2120, 2608]. Engraulis [1311, 1620, 469, 423, 661, 2568, 214, 1578, 2678, 136]. enhance [161]. enhanced [894, 2934, 1383]. enhancement [647, 1557, 260, 1461, 908]. enhancements [1558, 1554]. enhances [2238, 1429]. enigmatic [2387]. Enobarbus [353]. Enormous [2960]. enriched [769]. enrichment [3039, 2804, 1562, 2967]. ENSO [1980]. entails [11]. enter [2459]. Entosphenus [602]. entry [2696, 835, 1721]. Envelope [1310]. envelopes [42]. environment [631, 564, 3001, 31, 2988, 59, 2883, 1902, 3086, 1434, 1034, 1411, 163, 87, 2390, 2525, 612, 2973]. Environmental [986, 1126, 835, 2179, 1561, 1851, 2982, 594, 1334, 89, 1984, 686, 2211, 2845, 2918, 80, 805, 2149, 2203, 1827, 2027, 2220, 69, 632, 1833, 1407, 2142, 2375, 295, 329, 3014, 1603, 2008, 2927, 2933, 894, 1900, 2804, 460, 2852, 1891, 1473, 2205, 2188, 2016, 3034, 236, 1881, 1562, 1829, 1229, 1050, 2085, 27, 959, 2715, 1392, 362]. environmentally [1866]. environments [2754, 2017, 2569, 1076, 1974, 1636, 1691, 1021, 1073, 1284, 2534, 2776, 834, 3008]. enzymatic [845]. enzyme [1276, 3057, 241, 1522]. enzymes [2476]. EPA [2671]. epaulette [2833]. eperlanus [2914]. ephemeral [3000]. ephippid [897]. Ephippion [2728]. epidemics [1916]. epidermal [974, 3090, 337, 299, 2975]. epidermis [2086, 2347]. epidiosbulbin [247]. Epinephelidae [91, 1115, 1918, 2831]. epinephelids [2032]. Epinephelus [2722, 1739, 2024, 2847, 561, 1188, 1183, 1312, 1545, 1596, 1115, 1918, 2501, 843, 1200, 2557, 2458, 265, 1941, 2730, 3024, 1526, 736, 1534, 79, 212]. epipelagic [1180]. epiphanes [2810]. epiphyte [2103]. episcopi [403]. episodes [2269]. episodic [1792]. eponym [2144]. Eptatretus [679, 545, 2830, 257, 243, 3079]. equal [1533]. equation [978]. equations [1129, 510, 1585]. equatorial [2057, 2779]. eques [839]. equipment [2492]. equipped [1458]. Equulites [2230]. era [632]. eradication [2235, 1093]. Erdmann [1268]. erectus [1769]. eregoodootenkee [2968]. Eric [641].

Eridacnis [907]. Erie [1604]. Erimyzon [3083]. erinacea [792, 252]. EROD [850]. Erratum [72, 220, 262, 725, 1466, 1688, 1877, 2840, 3018, 73, 2633, 1689, 74]. error [2259]. Erythrinidae [1537]. erythrinus [427]. erythrocytes [83]. escape [195, 2331]. escapement [361]. Esociformes [1319]. Esox [1743, 2777, 2930, 546, 1915, 2761, 811, 89, 1774, 738, 181, 2334, 3009]. Essential [1265, 2821, 1314, 1516, 1515, 2704]. EST [380]. EST-SSR [380]. established [3061, 1077, 3015]. Establishment [2415, 2295, 2381, 2501, 1195, 265, 1529, 1568, 2888, 1394, 2605, 1968, 1717, 2405, 2233]. Este [1894]. estimate [1109, 1216, 120, 1495, 325]. Estimated [996, 2990, 2283, 1307, 2937]. estimates [2506, 3049, 1293, 2492, 2148, 2259, 2860, 2147, 2995, 2890]. Estimating [2532, 2663, 1811, 2048, 1028, 761, 497, 1214, 1830]. estimation [1864, 728, 1105, 2853, 1305, 2176, 1628]. estimators [1594]. estor [1461, 1960]. ESTs [82]. estuaries [1445, 969, 2315, 3070, 577, 1695, 154, 1653, 2106]. Estuarine [605, 943, 2017, 2007, 1524, 2339, 1613, 2019, 1937, 2389, 1101, 969, 2751, 880, 756, 1950, 2020, 1699, 848, 2934, 163, 383, 2014, 2015, 218, 1301, 233, 481, 721, 2021, 2223, 2110, 1072, 764, 1439, 2253, 1621]. Estuarization [2022]. Estuary [2836, 382, 863, 1140, 2025, 582, 2096, 809, 176, 2429, 2013, 1348, 90, 1430, 1531, 2610, 2417, 2188, 2036, 2667, 2012, 615, 2113, 2284, 1518, 27, 804, 1429]. estuary-resident [1430]. Etelinae [1860]. ethanol [2189, 1477]. Etheostoma [2078, 1398, 1787, 1091, 1250, 1361, 2516]. Etheostomatinae [2285]. Ethical [391, 2864]. ethinyl [1866]. Ethiopia [881, 1437, 1796]. Ethmalosa [2828]. Ethological [588]. ethylhexyl [2481]. Etmopteridae [1612]. Etmopterus [1612]. Etroplus [694, 160]. Etrumeus [2855]. Euchiloglanis [1494]. Eucinostomus [2011]. Eudontomyzon [1842]. Eumicrotremus [1640]. Eupallasella [1450]. Eurasian [2301]. Europe [2061, 2941, 1842, 1040, 460, 270, 930, 319]. European [1849, 797, 2918, 1705, 580, 361, 2591, 1456, 2405, 2955, 1899, 1125, 373, 2759, 448, 1660, 2461, 473, 2496, 879, 2061, 1357, 1224, 1030, 510, 402, 2312, 3044, 196, 1603, 3088, 374, 2489, 1655, 426, 1259, 2857, 1459, 1675, 1411, 2115, 1495, 1240, 1674, 1292, 2642, 1071, 2583, 2161, 3072, 1628, 1872, 1658, 1038, 995, 1624, 1799, 515, 136, 1647, 2419, 1201, 2914, 1122, 990, 656, 41, 2329, 434, 1027, 496]. Euryhaline [1269, 22, 2062, 1945, 730]. euthanasia [2323]. Euthanization [566]. euthanized [2598]. Eutrophication [1444]. evacuation [712, 1943, 1047, 684, 2046]. evaluate [1415, 1954, 2862, 1671, 298]. evaluated [2364, 885]. Evaluating [1685, 1305]. Evaluation [2783, 2543, 2224, 2964, 1512, 3024, 803, 354, 2599, 299, 2489, 2744, 607, 1083, 1220, 2888, 1346, 2052]. Evans [690]. event [114]. events [1177, 145, 1293, 1598, 544, 1888, 1980, 1428, 1000]. Everard [1404, 1264]. Everyone [1070, 2124]. Evidence [2791, 2326, 782, 2024, 2874, 713, 1542, 601, 1117, 1792, 130, 495, 2020, 2561, 1987, 2107, 2390, 1840, 796, 3084, 2150, 267, 468, 769, 179, 286, 2665, 2770, 2494, 951, 1601, 2607, 395, 3090, 1314, 1516, 2190, 1745, 370, 2740, 1041, 2055, 1151, 15, 1515, 2172, 1382, 3060, 279, 315, 36, 2479, 122, 1110, 1989, 217, 1938, 1844, 2355, 470, 2815]. evidences [1639]. Eviota [2810, 2094, 1938]. Evolution [1787, 690, 150, 2473, 2858, 1770, 1749, 210, 2146, 2285, 2925, 1443, 743, 1928, 2104, 862, 552, 272]. Evolutionarily [2884]. Evolutionary [1432, 8, 1665, 1743, 1760, 2559, 1559, 839, 1428, 490]. evolving [552]. Evynnis [2348]. Examination [1476, 2460, 1965]. examine [1852, 2259]. Examining [1179, 2653, 2806]. example [639, 2915]. Excess [1184, 2582]. exchange [2200, 1332]. exchanger [22, 801]. exclusive [2991]. exclusively [1357, 1785, 2594]. excretions [253]. Excretory [1325]. exercise [161, 1254, 46, 2582]. exhaustive [1254]. exhibiting [1069]. existence [2272, 2797]. exit [1583]. Exogenous [1647, 1095]. exoskeletons [1047]. exotic [700, 129]. expanding [2423, 1892]. expansion [2728, 267, 1643, 23, 2446, 852, 1459]. expectations [1998]. expenditure [2296]. experience [68, 1228]. experiences [1556, 1226]. experiment [879, 204, 1535, 834]. Experimental [2665, 1382, 279, 1821, 515, 2903, 2224, 917, 3003, 1655, 1073, 2430, 307, 856, 1197]. experiments [594, 878]. explain [1237, 827]. explained [1837]. Explaining [2551, 1103]. explains [2459]. exploitation [1222, 2068, 2897, 294, 2977, 1671]. exploited [631, 687, 2966, 2173, 753, 577, 2040, 752, 1465]. exploration [403]. Exploratory [2043, 1469]. explore [2497]. explored [2587]. Exploring [411, 451, 2975, 2384, 970]. Explosive [1797]. exposed [1746, 452, 435, 661, 1436]. Exposure [47, 133, 2652, 2102, 403, 2720, 429, 2664, 2931, 3012, 2258, 1866, 44, 1331, 2481, 1381, 544, 1379, 2354, 125, 2053, 2726]. express [1503, 612]. expresses [1529]. Expression [1178, 2059, 1711, 1793, 84, 539, 1728, 2321, 1121, 2487, 694, 2643, 2628, 1942, 566, 2764, 1570, 1947, 2943, 718, 1340, 1457, 457, 561, 2510, 1409, 1443, 1606, 717, 1005, 1521, 2058, 329, 2309, 1312, 447, 1867, 1017, 2373, 2213, 360, 1848, 1581, 1596, 2477, 843, 1917, 1097, 2710, 2595, 837, 1262, 1960, 2774, 2575, 1729, 1241, 144, 2589, 3087, 730, 1026, 1448, 2878, 1291, 212, 1958, 1513, 146, 826, 2237, 1736, 1931, 2548, 680, 1964]. expression-based [329]. expressions [1129]. exquisitus [659]. Extended [597]. extension [2327]. extensions [2386, 1530, 16]. Extensive [1593, 2950, 2304, 758]. extensively [2284]. extent [512, 637]. External [1284, 2, 2303, 1808, 2680, 803]. externally [71]. extinct [2839]. extinction [153]. extirpation [1052, 1057]. extraction [2189, 522, 1651]. Extraordinary [977, 113, 1860]. Extreme [670, 1640, 2960, 544, 1888, 1552]. extremely [99, 1900]. extrinsic [1166]. extrusion [1160]. eye [2832, 1245, 522, 2612, 2883, 2795, 104]. eyed [2977, 2574]. eyes [2930].

F [1126, 550, 461, 1594, 2167]. fabp2 [718]. fabricii [3011]. face [499]. faced [2318]. facetus [2306]. facial [1130]. facilitate [1331, 2797]. facilitates [1354]. Facing [1386, 2531, 1057]. Factor [1374, 590, 87, 437, 2961]. factor-I [87]. Factors [2506, 560, 197, 3023, 1166, 785, 81, 2308, 692, 2142, 2048, 10, 791, 558, 594, 89, 2553, 967, 88, 3034, 1986, 2021, 1050, 2792, 2715]. Facultative [3070, 2229, 1854]. faecal [1375]. faeces [2312, 2306]. Fahay [605]. Failure [1227, 110, 831, 1777]. falciformis [2260, 761]. Falkland [1222]. fallax [323, 1136]. fallout [1400]. false [1041, 1169]. familiar [2962]. familiarity [707, 941, 34]. families [2844, 1409, 2274]. Family [1318, 2420, 1636, 1443, 822, 2257, 1115, 2970, 36, 2748, 317, 2082, 149, 890]. Fancisco [2003]. fanskate [2048]. FAO [980]. far [2649, 298, 1874]. far-eastern [298]. farm [2597, 2320, 2157]. Farmed [1692, 2869, 1170, 797, 926, 1857, 2340, 572, 1510, 173]. farmer [2787]. farms [364, 1893]. Farrell [1265, 1269, 739]. farsicus [1627]. fasciatum [626, 2001]. fasciatus [1125, 1714]. fast [457, 1628, 1471, 1607]. fast-start [1607]. fasting [1942, 2714, 1448, 1607]. fat [2117, 1508, 642, 230]. fathead [533, 1481, 1068, 447]. fatigue [2562]. Fatty [2112, 1572, 2113, 1490, 2398, 819, 1420, 2477, 2381, 2557, 1683, 230, 2283]. Fatty-acid [2112, 1572, 2398, 819, 2477, 2557, 230, 2283]. fauna [2025, 2427, 1781, 28, 1995, 2313, 614, 2284, 27]. faunal [1531]. faunas [1978]. Favonigobius [659]. features [2221, 2620, 175, 2357, 1483]. Fecundity [1901, 2783, 889, 93, 2290, 2550, 352, 95, 2177]. fed [2046, 2477, 2951, 1282]. feed [2869, 1409, 402, 2087, 1668, 1509, 1472, 1471]. Feeding [3067, 2301, 1438, 809, 1613, 2260, 205, 2173, 167, 1061, 773, 483, 822, 342, 1733, 772, 382, 135, 1362, 2289, 648, 554, 1522, 1060, 1249, 137, 963, 686, 1181, 1278, 712, 1943, 887, 564, 1590, 1611, 2031, 2910, 679, 52, 31, 1660, 1445, 2119, 1563, 377, 1846, 534, 372, 940, 1011, 2592, 2135, 1593, 3063, 1015, 1388, 2963, 2646, 896, 1785, 2087, 652, 1408, 187, 500, 1095, 953, 808, 1480, 2112, 417, 3002, 1573, 1016, 1727, 2781, 1501, 2480, 2050, 2400, 1724, 2170, 1484, 2745, 1605, 900, 136, 1122, 2113, 1375, 203, 2903, 1159]. feeding [475, 2283, 1180, 499, 578, 2650, 2855, 959, 1547, 1607]. Feel [525]. Female [412, 2641, 352, 1605, 2321, 1042, 1181, 2264, 1701, 2483, 2847, 2117, 2873, 101, 503, 343, 864, 127, 1303, 2076, 756, 1806, 1753, 2095, 1250, 1847, 224, 3062, 3069, 2774, 3024, 683, 2674, 1176, 3087, 1840, 916, 698, 1298, 1862, 480]. Female-biased [2641]. females [51, 207, 1251, 505, 2674]. Femtosecond [619]. feral [1179, 2726]. feral-pig-damaged [2726]. ferrugineus [1385, 2388, 2908]. fertilization [29, 1680, 2674, 705, 223]. fertilized [1310, 1651]. fertilizing [1749, 1896]. fever [903]. few [1653, 1874]. fgeneoxl [1581]. fibres [1797]. fidelity [249, 1523, 2184, 2756, 2172, 3037]. field [858, 338, 620, 1949, 2089, 1801, 2689, 394, 2165, 1134, 279, 2462, 117, 2755, 1719, 2631, 1209]. fields [876, 1900, 2077]. fifth [2193]. Figaro [1164]. fight [834]. fighter [217]. filament [299]. filamentosus [1860]. filamentous [504]. filiformis [2638]. filter [1547]. filter-feeding [1547]. fimbriata [2828]. Fin [2750, 137, 2950, 777, 587, 647, 760, 2111, 2686, 2990, 603, 2379, 1305, 2302, 1137, 1419, 2891, 2595, 1853, 2205, 2434, 786, 2171, 2088, 1840, 2204, 2605, 3009]. fin-clip [2434]. fin-ray [2205]. fin-to-carcass-mass [786]. final [559]. find [2124]. Finding [888, 2373, 2493]. findings [2241]. Fine [1101, 501, 1476, 1143, 3037]. Fine-scale [1101, 501, 1476, 1143, 3037]. Finland [2914, 331, 1198]. Finley [831]. finned [3090]. finning [786]. fins [1315, 1308, 2396, 706, 1579, 272, 968]. fire [2551]. firemouth [666]. First [2241, 484, 2564, 2790, 2123, 1894, 2339, 2908, 2729, 432, 1460, 2758, 1454, 2335, 2820, 2688, 1745, 388, 2075, 1151, 2857, 1459, 586, 2974, 1124, 2460, 107, 1299, 1786, 1064, 930, 1722, 1370, 2757, 1039, 1854, 431, 45, 906, 1094, 669, 2495, 890, 912, 2002, 1096, 2622, 2391, 2991, 1257, 2079, 3090, 2320, 3078, 1544, 187, 417, 973, 2784, 1391, 705, 203, 1207, 664, 1243]. first-feeding [187, 417]. Fish [442, 464, 488, 507, 527, 551, 570, 156, 1973, 2146, 2023, 508, 2457, 324, 1357, 969, 1232, 2799, 2008, 525, 831, 1099, 1211, 3017, 861, 1991, 739, 1297, 1975, 2251, 633, 1527, 27, 2658, 1849, 8, 1121, 928, 3067, 712, 2987, 998, 1556, 1316, 2025, 2508, 1765, 582, 2096, 2238, 1271, 728, 1231, 1479, 2007, 62, 1402, 245, 1294, 976, 2427, 769, 2638, 2140, 2854, 1300, 2944, 354, 2663, 364, 235, 1632, 535, 398, 1421, 1533, 2239, 1976, 2464, 2655, 2943, 457, 98, 2330, 2502, 11, 2060, 640, 2662, 2958, 2089, 522, 1054, 2093]. fish [820, 1345, 2224, 66, 1414, 2386, 1014, 986, 1109, 918, 2551, 2916, 1168, 302, 1833, 2426, 1781, 2904, 1258, 2013, 1992, 2924, 713, 2026, 1113, 1286, 1567, 2593, 2737, 409, 1008, 24, 1933, 2126, 3049, 800, 28, 260, 1388, 2566, 2537, 49, 2735, 2287, 56, 711, 1306, 1005, 2297, 334, 3064, 2186, 1902, 114, 1762, 732, 486, 872, 274, 1305, 250, 329, 1463, 175, 2942, 726, 601, 854, 116, 25, 984, 1564, 1691, 2075, 2531, 1413, 2503, 168, 1462, 1543, 2751, 2935, 1978, 2415, 873, 1925, 2736, 3039, 1835, 3043, 347, 822, 133, 1610, 684, 838]. fish [2492, 2933, 1896, 2906, 2446, 3085, 1904, 1416, 1646, 2975, 206, 560, 1771, 2917, 1687, 342, 2257, 724, 1950, 1152, 147, 1189, 1581, 1444, 1148, 327, 1977, 1442, 2020, 2824, 1700, 628, 2071, 2004, 2172, 2155, 2901, 2610, 1548, 2982, 2738, 1831, 3, 433, 731, 2849, 241, 2550, 2923, 1538, 2018, 2463, 2801, 2805, 2350, 3070, 2934, 1727, 2409, 2666, 315, 2821, 846, 397, 2538, 1079, 75, 985, 2028, 1424, 1893, 265, 3033, 2533, 1673, 2022, 1483, 1825, 3000, 1133, 660, 635, 2580, 1157, 2877, 577, 630, 273, 2922, 821, 3034, 2798, 995, 2652, 639, 2784, 1890, 383, 2447]. fish [87, 1609, 1979, 387, 1985, 2947, 625, 1986, 233, 481, 1213, 914, 2585, 2406, 2419, 597, 2954, 1230, 2755, 2522, 987, 2864, 1341, 2661, 810, 1050, 885, 1162, 2107, 255, 1692, 2088, 670, 203, 2613, 1159, 1116, 2407, 1346, 1080, 1537, 1990, 2284, 434, 1650, 2981, 708, 2726, 2792, 162, 1072, 2605, 305, 154, 1653, 3008, 2765, 3058, 2584, 1103, 1439, 2253, 912, 1717, 424, 2486, 237, 2250, 1269, 568, 588, 2070, 1127, 109]. fish-cell [1717]. fish-oil [1230]. fish-shoal [2593, 2737, 2736]. fish-stock [2140]. Fish. [1173]. FishBase [970]. Fisheries [463, 487, 506, 526, 1226, 1212, 774, 831, 441, 139, 2209, 548, 569, 1222, 2588, 1708, 2918, 1929, 2057, 1557, 2199, 2162, 2536, 1863, 392, 161, 1227, 2899, 158, 2035, 1345, 783, 1558, 2607, 2630, 1228, 1410, 2472, 3064, 2304, 1221, 2689, 1225, 1554, 326, 2803, 2901, 2148, 2158, 1220, 314, 1219, 1223, 2443, 638, 747, 2009, 778, 2911, 159, 2912, 741, 2685, 2145, 2452, 3040, 1654, 604, 982, 110, 1266]. fishers [2900]. Fishery [982, 1226, 785, 1804, 1993, 1054, 1147, 1228, 449, 3030, 2045, 2348, 1565, 716, 1635, 1217, 2205, 1215, 2913, 2914, 2178, 1198].

Fishes [152, 1, 794, 112, 2193, 1210, 1426, 605, 2537, 261, 1172, 1268, 1726, 1983, 2946, 956, 859, 183, 2658, 389, 981, 1404, 1876, 1177, 1696, 1415, 2631, 1556, 3027, 2640, 2673, 2149, 587, 858, 178, 2067, 976, 1930, 179, 1533, 453, 1751, 1822, 1817, 1824, 867, 868, 396, 659, 2536, 1325, 33, 153, 1478, 2285, 795, 1827, 620, 1356, 1716, 392, 1227, 2027, 975, 1949, 268, 1974, 2035, 735, 567, 372, 1276, 632, 1652, 565, 2359, 330, 1558, 2897, 1364, 145, 1916, 2256, 170, 47, 991, 3090, 2103, 876, 2920, 13, 1994, 2825, 77, 589, 1187, 2179]. fishes [2379, 1584, 543, 1317, 1559, 1059, 1927, 2069, 1561, 1434, 2927, 2228, 1435, 1777, 2136, 26, 360, 34, 1322, 1329, 1327, 1826, 2272, 865, 1442, 953, 129, 321, 313, 886, 2423, 2921, 1324, 2374, 1214, 2862, 895, 391, 1830, 1756, 1480, 2546, 1926, 2934, 1275, 1323, 1818, 2926, 1823, 2368, 2161, 2028, 2314, 1326, 333, 720, 1381, 2216, 1078, 1633, 971, 2188, 88, 421, 2877, 314, 1273, 1255, 239, 3055, 2462, 1182, 1165, 2806, 942, 946, 95, 2893, 1301, 2863, 2230, 2012, 2507, 646, 2555, 2928, 9, 363, 1328, 1150, 2030, 71, 1342, 696, 1989, 1154, 2369]. fishes [125, 882, 2815, 2, 2746, 1912, 131, 1302, 884, 1277, 2223, 1465, 542, 1984, 1730, 76, 1547, 2147, 742, 2106, 2143, 2453, 1621, 789, 2367, 634, 1267, 1126, 462, 141, 690, 1269, 1371, 1372, 439, 1264, 1550]. Fishing [157, 708, 1556, 2918, 1218, 1216, 2852, 2902, 1512, 1635, 2404, 1229, 298, 2654]. fishmeal [1230]. fishy [2842]. fit [2533]. fitness [608, 1561, 2907]. fitting [2259]. Five [1158, 1649, 2997, 1751, 2731, 2364, 2627, 2531, 1477, 1480, 2666, 1132, 3033, 935, 1522]. fixation [1750]. fixed [522]. fjord [516, 2435]. fjords [2963]. flagella [11]. Flank [2494]. flash [1248]. flatfish [263, 908, 1224, 2387, 1682, 824, 1458]. Flatfishes [2068, 2567, 1598]. flats [1187]. flavescens [2469, 93, 2596, 1914, 451, 1143, 1800, 834, 126]. flavipinnis [2762]. flesus [1344, 2642]. flexibility [1354, 1774]. flexuosa [386]. flight [1299]. floating [2696, 2260, 1204, 180]. flock [1796]. flood [2640, 2350, 1888]. flood-plain [2350]. floodplain [2239, 1205, 1985, 1912]. floodplains [172, 1988]. Florida [2385, 1557, 1555, 578, 2453]. floridanus [1555]. flounder [1193, 195, 898, 692, 59, 1101, 1598, 148, 1032, 1641, 2642, 847, 1449, 1575, 705, 475, 931, 1247, 1291, 425, 1390, 1717]. Flow [2614, 2754, 80, 1384, 802, 2677, 851, 579, 2055, 2610, 1674, 1763, 1829, 1828, 633, 2703, 1819, 165]. flow-cytometric [802]. flows [804]. fluctuations [40, 1758, 1833]. fluid [410, 1713, 2270, 1944, 1718, 2306, 1680, 2850, 1574]. flukes [104]. Flume [46, 610, 2224]. fluorescence [1755, 1284, 736]. fluorescent [2487, 2624, 367, 1768]. fluvial [664]. fluviatilis [2235, 80, 2301, 1456, 2214, 286, 202, 1916, 636, 1117, 1634, 2930, 50, 1532, 1889, 1151, 57, 558, 1281, 135, 89, 536, 1549, 1624, 1799, 515, 67, 656]. Flyingfish [1180]. focal [2410]. focus [2248, 1804, 2781, 1879, 2249]. foetus [1094]. Fold [2806]. foliaceus [2127]. Follicle [2227, 727, 2643, 2687]. follicle-stimulating [2643]. follicles [238]. follower [1287]. following [2235, 2543, 561, 301, 1113, 1745, 1866, 1095, 927, 2507]. Fonchiiloricaria [591]. fontinalis [2760, 1488, 662, 2308, 1746, 1631, 2719, 1585, 2046, 579, 1104, 806, 1471, 46, 3028, 165]. Food [2742, 427, 1925, 1480, 399, 980, 1696, 1181, 2528, 1239, 1618, 1231, 1456, 769, 398, 520, 655, 1414, 3013, 202, 1469, 2373, 1053, 1694, 2272, 1987, 417, 1016, 688, 318, 936, 163, 88, 307, 639, 2675, 284, 2507, 1865, 169, 368, 1965, 171]. food-limited [417]. food-mimicking [1181]. food-web [284, 2507]. foodweb [3066]. foothills [645]. for-hire [1229]. forage [1221, 2917, 3033, 1154]. forage-fish [3033]. Foraging [394, 546, 1952, 2655, 2288, 2544, 1939, 599, 1078, 1724, 2653, 2957, 1912, 857, 126]. force [2372]. forces [889, 2496]. forecast [1223]. foreign [2575]. forest [2183, 2993, 2313, 1977, 89, 2839]. form [1120]. formalin [1712]. formalin-killed [1712]. Formation [2784, 664, 42, 1397, 1144, 3029, 2059, 34, 720, 1709, 1856]. formed [722]. Formiga [2437]. Formosa [1758, 205, 292]. formosum [235]. formosus [595]. forms [85, 1035, 682]. forsteri [2574, 1289]. fossilis [2321, 336, 2271]. found [1542]. foundation [670]. four [1350, 2640, 1445, 2089, 2298, 1904, 150, 818, 1071, 1638, 2263, 95, 1026, 2476]. Fourteen [929]. Fourth [1172, 139]. Fowler [1270]. foxfish [231]. foxl2 [2237]. fractionation [1396, 2968, 310]. fractionize [376]. fraenum [1787]. fragment [613, 150]. fragment-length [613]. fragmentation [2336]. fragments [216]. framework [2442]. France [630]. Francis [3044, 1172]. Francisco [2606, 1720]. Fraser [872]. free [2795, 579, 1693, 1865, 1063]. free-ranging [2795, 1693]. freezing [3081]. French [2326, 973]. frenchii [231]. Frenetic [11]. frequencies [2778, 1888]. frequency [582, 2663, 829, 260, 1900, 2290, 408, 944, 2886, 1862]. frequently [2133]. Fresh [356, 2702, 574, 1037, 995, 2390, 2706]. freshets [1911]. Freshwater [1855, 1929, 1267, 1426, 2035, 1978, 859, 2062, 1839, 1497, 2658, 981, 1264, 2690, 569, 1849, 866, 998, 1556, 157, 3027, 1705, 2673, 62, 1550, 1524, 1319, 1993, 2664, 33, 809, 2220, 1194, 2089, 678, 532, 1109, 2916, 1168, 2904, 2256, 1784, 49, 1995, 601, 1317, 1564, 1378, 2313, 1543, 1085, 2191, 833, 1416, 1889, 426, 1444, 327, 1442, 619, 129, 1281, 433, 1727, 716, 379, 333, 1633, 693, 1991, 635, 2351, 3034, 2462, 1890, 2806, 233, 481, 363, 1122, 629, 1522, 1162, 2088, 203, 2223, 2872, 1650, 2726, 162, 1653, 804, 2143, 2969, 641, 261, 2314]. Friedland [139]. frigidus [1589]. frillgoby [1025]. fringed [163]. fringing [1385]. frontal [2303]. Frontiers [2196, 1803]. frozen [1477]. Fruit [2993]. fry [949, 1789, 2967, 1256, 2390, 1827]. FSBI [2539]. Fsh [3062]. Fsrp [2878]. Fsrp-3 [2878]. fuelled [2993]. Fukushima [1400]. full [1096, 1037]. full-strength [1037]. Fully [1270, 882]. fulvescens [2786, 835, 2889, 2205, 2715, 2752]. fulvidraco [2083, 2481, 2548]. function [471, 2487, 1399, 286, 545, 1414, 402, 447, 347, 5, 3053, 788, 2460, 2821, 985, 787, 2703, 2878, 1885, 1779, 272, 2253]. Functional [8, 2060, 137, 400, 210, 1533, 581, 1596, 1917, 2996, 2022, 3034, 1182, 2347, 1154, 2407, 514, 789, 2961]. functions [709, 1244]. Fundulus [1330, 22, 1540, 2323, 18, 2062, 2848, 842, 3075, 989, 3032, 1885, 1579, 1621]. Fundy [2119]. furcatus [1797]. furcifer [1014]. furnieri [1112, 35]. further [586]. furzeri [1121]. fuscescens [623]. fuscoguttatus [79, 3024]. fuscus [1919, 1639, 2992, 1025]. Future [1708, 2918, 1271, 2007, 2536, 324, 3064, 1675, 2921, 1045, 2912]. Fyzul [3017].

G [112, 955, 142, 140, 110, 1268, 568, 588, 109, 183, 739, 439, 438, 2943]. G-protein [2943]. G. [2282]. Gadidae [616, 1452]. Gadiformes [470, 643, 2066]. gadoid [450]. gadoids [1943, 1582]. Gadus [189, 1741, 2459, 712, 1399, 1713, 1310, 1285, 2742, 926, 322, 1056, 1047, 534, 974, 2044, 511, 1795, 2952, 2963, 120, 376, 1420, 1476, 187, 2125, 264, 927, 3081, 476, 2550, 585, 1045, 1167, 1744, 1184, 2497, 2166, 878, 2177, 1153, 1349]. Gaffney [2736]. Gaidropsarus [1452]. Galapagos [1722]. Galaxias [1246, 901, 597, 2107]. galaxiid [2235, 429, 1331]. galaxiids [1332]. galectin [2872]. Galeocerdo [977, 2372, 1892, 1740]. Galeorhinus [2790, 1737, 432, 1811, 1542]. Galeus [1278, 2180, 2790, 1737, 432, 1811, 1542]. Galician [484]. galilaeus [424]. Gambia [172]. Gambusia [2391, 937, 1148]. Game [1404]. Gamete [218]. gametogenesis [2584]. GAMT [84]. Ganges [2602]. gangeticus [2602]. ganglion [1130]. gap [992]. gaps [157, 2531, 635]. gar [695, 1775]. García [109]. García-Ayala [109]. garfish [370]. gariepinus [2150, 1963, 173, 1997]. garmani [1060]. gas [106, 1332]. Gasterosteus [244, 1397, 210, 678, 1163, 583, 2612, 1684, 1493, 1925, 682, 1366, 649, 2384, 947, 1049, 544, 2351, 2850, 892, 1119, 2858, 514, 2995]. gastric [712, 1943, 1047, 684, 2046]. gastrically [2626]. gastrointestinal [833, 2476]. gattorugine [1922]. gayi [952]. GCRV [1106]. Gdańsk [517]. gear [2654]. gears [1635]. Gelanoglanis [2282]. gelatinous [1239]. Gender [1417]. Gene [2764, 851, 360, 2878, 1415, 694, 80, 2628, 608, 1570, 1947, 718, 1340, 1457, 1738, 2488, 2364, 2510, 1409, 1443, 717, 2778, 329, 1896, 2055, 2477, 150, 843, 2710, 837, 1262, 1049, 1960, 1763, 2575, 1729, 144, 437, 730, 1448, 681, 2584, 212, 1958, 1513, 539]. gene-rich [608]. genera [2056]. general [2773, 63]. Generalist [454, 1921, 285, 1439]. Generation [82, 2167, 2320, 1854]. generational [2064]. generations [705, 1243]. generic [712, 43, 2991, 912]. genes [2146, 2059, 1443, 1606, 2988, 1521, 2058, 871, 447, 2964, 2951, 2322, 333, 2774, 3087, 2872, 850, 2157, 1026, 1958, 2548]. Genetic [657, 964, 271, 576, 85, 2364, 1163, 687, 395, 2154, 1627, 674, 1810, 1314, 1516, 1035, 1410, 334, 3083, 1957, 1114, 1205, 2740, 1143, 2505, 147, 1641, 1048, 1675, 2152, 2232, 1288, 2879, 270, 3060, 1004, 1446, 2417, 536, 967, 2713, 1083, 816, 2479, 365, 2352, 950, 363, 2780, 2166, 1198, 1537, 199, 622, 2835, 2153, 685, 2811, 2150, 2149, 1954, 2421, 2526, 1767, 662, 2151, 951, 404, 1147, 1916, 2873, 2868, 2777, 2559, 465, 1762, 2130, 1131, 2515, 996, 2738, 1433, 799, 585, 1796, 716, 2553, 2148, 3072, 1714, 328, 211, 2989, 236, 406]. genetic [538, 2555, 2178, 215, 2305, 331, 1946, 670, 1027, 496, 1425, 1515]. Genetically [1624, 667, 2494, 2527, 1392]. Genetics [2972, 1749, 1570, 2162, 2382, 783, 1880, 269, 491, 332]. genital [287]. Genome [2159, 1958, 2156, 1056, 571, 1002]. Genome-wide [2159, 1958, 2156]. genomes [2146, 1441, 2305]. Genomic [263, 75, 2987, 608, 151, 2158]. genomics [2156, 2149, 329, 491, 2160]. genotype [2778, 1626]. genotypes [1608]. Genotypic [607, 1849]. Genotyping [1946, 335, 2585]. Genus [1174, 2617, 866, 964, 179, 2880, 2282, 1894, 1319, 2185, 2509, 1304, 2382, 616, 1308, 124, 472, 753, 1905, 2066, 2684, 1788, 714, 2212, 2233, 420, 1078, 1165, 2263, 591, 2983, 1302, 2872, 2324, 2743, 1055, 2380, 2397, 2138, 890]. geochemical [2084]. geochemistry [2200]. geoffroy [1195]. Geographic [2211, 2728, 852, 294, 341, 2515, 1891, 2555]. Geographical [1761, 1636, 1147, 370]. geographically [920]. Geography [1049]. geomagnetic [3073]. Geometric [1157, 1759, 1880]. geometries [684]. geometry [2005]. Geophagus [1258]. George [483, 2515, 2310]. Georges [1449]. Georgia [2356]. georgii [954, 2268]. Georgiy [2281]. Germ [705, 2884, 2584]. germ-line [2584]. German [1556, 2009]. Gerreidae [2011, 1525, 2014]. Géry [1393]. gestation [756]. GH [1237]. GH-transgenic [1237]. Ghats [1951]. ghrelin [1097]. giant [2722, 2127, 54, 2710, 1699, 2361, 237]. gibberfishes [1530]. gibbosus [636, 3012, 1028, 965, 319]. gibbus [3068]. gibel [825]. gibelio [337, 229, 825]. Gila [2357]. Gilbert [2368]. gill [836, 2773, 2664, 1238, 1065, 2323, 1711, 2960, 693, 1547]. gillnets [1050]. gills [1665, 1335, 304, 22, 1262, 801]. Gilthead [983, 797, 608, 2868, 1313, 452, 44, 1843, 2445]. ginbuna [1040]. Girard [1315]. Girardinus [3058]. girdle [2686, 2749]. Girella [56, 1619, 1599]. girellids [1591]. gizzard [1670]. glacial [267, 1589, 1643, 2055, 14, 16]. glaciations [2383]. glands [118]. Glandulocaudini [2000]. glanis [2312, 3072, 990]. glass [402, 2489]. glasseye [1760]. glauca [2122, 1701, 830, 1473, 786]. Glaucosomatidae [15]. glaucostigma [1733]. glaycos [2488]. glesne [1158]. Glimpses [111]. Global [910, 2472, 1230, 2344, 777, 161, 2630, 3064, 1215, 1019]. globally [1737]. Glossamia [2754]. glossodonta [2966]. glucocorticoid [3012]. glucocorticoids [5]. glue [1403]. glycoconjugates [2808]. glycoprotein [214]. Glycoproteins [304]. Glyphis [2602]. glyphosate [2566]. glyphosate-based [2566]. Gnathopogon [1778, 1595, 2594, 2980]. GnRH [6, 2819]. GnRHa [3062]. Go [221, 2944, 2467, 1497]. Goals [2898]. goatfish [2243]. goatfishes [3030]. gobies [1086, 659, 676, 2611, 303, 2556, 2534]. Gobiesocidae [2450]. Gobiesox [2998, 32]. gobiid [972, 912]. Gobiidae [2420, 98, 809, 1445, 2426, 1348, 1884, 676, 1151, 2417, 303, 2294, 972, 1938, 2438, 1657, 912]. Gobiiformes [2420]. gobio [479]. Gobiodon [179]. gobioid [2291, 888]. Gobioidei [619, 2294]. Gobiosoma [1445]. Gobius [1884, 972]. Gobiusculus [834]. goby [293, 2720, 1146, 809, 1445, 2261, 2076, 2298, 691, 1151, 2095, 1001, 2649, 1724, 972, 1370, 517, 215, 1604, 1938, 1043, 434, 834, 1657, 513, 2654]. goldband [2688]. golden [2297, 2168, 2295, 2114, 2605, 1968]. goldfish [2167]. Goldspotted [2443]. goliath [2024, 894]. gomesi [371]. Gompertz [656]. Gonad [2094, 2059, 1777, 2552, 2237]. Gonadal [511, 711, 2402, 237, 1521, 2884, 1253, 2132, 144]. gonadotrophin [2321, 2819]. gonadotrophin-releasing [2819]. gonadotropic [2643]. gonadotropin [18]. gonadotropin-releasing [18]. gonads [2810, 1682, 2328]. Goncalves [140]. Gonorynchiformes [461]. González [1516, 1515]. good [2660]. goodei [2034, 2681, 436]. Goodeidae [711, 1302]. Goosefish [2290]. gorbuscha [737, 6, 278, 2976, 2118]. Gordon [1232]. Goriup [3016]. Gorringe [1487]. gouldingi [1672, 2592]. gradient [1330, 12, 2806, 275, 2497, 131, 730, 1518, 1439]. gradients [611, 2027, 1431, 1543, 460, 2018]. gradual [840]. grahami [1337]. Gran [2902]. Grande [461]. grandis [1330, 2323, 2062, 2848]. Grants [2539]. graphic [2979]. graphicus [2979]. grass [2809, 717, 2058, 2778, 668, 1867, 2671, 1638, 1241, 437, 2589, 826, 1106]. Graus [1619]. gravel [1170, 645, 2901, 1102]. gravel-dwelling [645]. gravity [1310]. Grayling [19, 196, 1603, 845, 1185, 365, 41, 2329, 1027]. graysby [91]. graytail [2408]. Grazing [1453]. Great [326, 1910, 2385, 1897, 3074, 2237]. greater [3062]. Greece [229, 939]. green [323, 2487, 820, 1591, 1586, 2820, 1686, 1037, 2753, 2175]. greeneye [1614]. Greenland [2811, 2154, 200, 746, 1682]. Greenwoodochromini [1343]. grenadier [3057, 384]. grey [2506, 576, 2808, 1784, 2891, 117, 1439, 2937]. griseocauda [2408]. griseus [117, 1439]. gross [2509, 2871, 1769, 1539]. Grosso [2437, 2422, 3080]. ground [1249, 2775]. groundfish [2963]. grounds [1590, 652, 2063, 1240]. groundwater [2142]. groundwater-dominated [2142]. group [2285, 655, 610, 652, 1293, 1148, 57, 2574, 2535, 878]. group-living [1148, 2535]. grouper [2722, 1739, 2024, 1738, 2847, 561, 1188, 1312, 1596, 2501, 843, 1200, 1939, 2710, 2458, 265, 1941, 2730, 2831, 1722, 2049, 3024, 1526, 736, 79, 212]. groupers [1918, 621, 1576, 1534]. grouping [1154]. groups [2383, 1933, 1727, 1633, 1182, 122]. Growth [2733, 1629, 1294, 2720, 1942, 231, 2990, 2029, 2477, 1536, 417, 1659, 1941, 2675, 2259, 2333, 2085, 963, 1385, 977, 1013, 2506, 2869, 1170, 40, 1776, 444, 256, 2469, 1864, 509, 1283, 1616, 1524, 468, 2609, 533, 760, 3050, 2576, 2624, 807, 1764, 338, 1580, 98, 3029, 31, 855, 2220, 1429, 3031, 802, 1732, 1587, 1811, 1514, 918, 2044, 1588, 1795, 59, 2952, 1475, 97, 2853, 2345, 2375, 56, 2219, 1306, 1510, 2590, 1238, 2087, 1354, 1293, 2181, 853, 174, 2261, 2275, 1740, 297, 2266, 1378, 393, 351, 177, 2747, 1962, 1085, 2302, 590, 1423, 2980, 658, 187]. growth [894, 592, 115, 1915, 2679, 2176, 2804, 649, 966, 1996, 996, 1024, 460, 38, 2756, 945, 1001, 401, 1034, 1955, 2032, 1461, 2707, 1495, 344, 3076, 1369, 1016, 1853, 1509, 315, 318, 1940, 1628, 936, 1678, 2678, 2753, 1605, 236, 2848, 842, 87, 2578, 1077, 481, 2978, 2271, 1647, 2697, 3024, 517, 2318, 1750, 1162, 125, 128, 2903, 1604, 3032, 475, 225, 1782, 656, 41, 164, 2344, 2981, 2886, 1380, 339, 698, 1298, 1862, 1390, 1197, 1244, 1392, 2118, 362]. Growth-band [2990, 3029]. growth-enhanced [894]. growth-hormone [1238, 2578]. growth-promoting [1244]. growth-related [996]. grunion [2670]. grunt [1349]. grunter [3013]. grunters [55]. Guadalquivir [99]. Guadalupe [2564]. Guadiana [30]. Guam [420]. Guangxi [2414, 733]. guarding [828, 2155]. gudgeon [734, 688, 2504, 2562]. gueldenstaedtii [2059, 42, 1690]. Guerrero [2822]. guggenheim [1577]. Guianas [2631]. guianensis [2429]. guichenoti [734, 2562]. guidance [3051]. Guide [1426, 138, 2631, 1209, 2313, 2689, 1687, 1876, 1270]. guideline [685]. guild [1632]. guilds [2359, 1348, 2963, 1527]. Guinea [257, 2387, 431]. Guinean [1088, 172]. guineensis [172]. guitarfish [378, 1011, 1733, 1909, 1012, 1813, 2218]. guitarfishes [3050]. Guizhou [2645]. Gulf [2506, 2733, 1158, 1590, 1883, 1643, 340, 575, 388, 2836, 1052, 1057, 342, 2323, 3045, 2423, 1698, 2232, 1932, 211, 1449, 2870, 3055, 2062, 2848, 2218, 2831, 1229, 2914, 517, 2407, 431, 593, 1576, 1094, 3015, 2812, 1868, 912]. gulper [2449, 1622]. gummy [1887]. gunnel [2082]. Günther [2066]. guppies [207, 1780]. guppy [1179, 1635, 1683, 190, 2767]. guppyi [385]. gut [535, 2482, 1670, 1323, 125, 428]. guts [3067]. guttatus [2853, 1793]. guttifer [2728]. guttulatus [498, 1339, 2644]. Gymnocephalus [1117]. Gymnocharacinus [2471]. Gymnocypris [332]. Gymnothorax [563, 532, 2269]. Gymnotidae [731]. Gymnotiformes [2772, 1202, 731]. Gymnotus [731]. Gymnura [2959, 1303, 921]. Gymnuridae [2959]. gynogenesis [2167, 2614, 369]. gynogenetic [205, 1032]. gyre [221]. Gyrodactylus [1093].

H [208, 112, 908, 22, 141, 1372, 801, 2258, 1355, 1711, 801]. H. [1459, 2936, 2565]. H.-P [112]. Habitat [234, 2019, 1937, 1842, 729, 172, 300, 2294, 2419, 169, 1153, 1921, 672, 569, 1920, 156, 1973, 1199, 1616, 1528, 1611, 1623, 2759, 2597, 1602, 735, 687, 565, 350, 2079, 1101, 341, 1784, 355, 1807, 1523, 1959, 2336, 3083, 1584, 2261, 295, 483, 756, 1052, 366, 512, 2544, 2020, 1699, 259, 279, 2028, 847, 2188, 1619, 2062, 1525, 2014, 2015, 2748, 2928, 1604, 60, 1579, 305, 3077]. habitat-driven [2020]. habitat-generalist [1921]. Habitat-related [1921]. habitat-specialist [1921]. habitats [866, 306, 1524, 2717, 2625, 2060, 1636, 2359, 1218, 1354, 2799, 1934, 2200, 483, 1978, 1085, 1983, 1142, 619, 2801, 2350, 1133, 303, 1075, 3034, 2965, 1730, 1072]. habits [398, 2998, 2135, 771, 167, 1733, 772, 2050, 2400, 2170, 1522, 1965]. haddock [509, 1752, 2134, 193, 1965, 276]. hadiyahensis [1359]. Haematological [1092]. haematology [1355]. Haematophagous [1122]. haemoglobins [13]. haemolytic [1772]. Haemulidae [3038]. Haemulon [1489, 2304, 1078]. hagfish [679, 545, 2830, 257, 586, 2377, 3079]. hagfishes [243]. hake [879, 1228, 2857, 1628, 1658, 648, 1249, 952, 1709, 470]. hakes [1766, 1685]. half [82, 380, 1568, 2476, 1931]. half-smooth [82, 380, 1568, 2476, 1931]. halibut [446, 2151, 200, 856, 1682, 803]. Halieutichthys [246]. hall [887]. halleri [756, 211]. hamatus [2829]. hamlet [2448, 2032]. hammerhead [2385, 2057, 3029, 1587, 2873, 519, 1897, 3074]. Hamon [139]. handling [301, 92, 190, 2334, 1794, 1885]. Handy [1126]. haplochromine [624, 486]. hapuku [710]. harbour [1471]. Hardness [2053]. harengus [230, 1580, 58, 1795, 394, 1363, 929, 1662, 2949, 1605, 1535]. harlequin [1286]. Harmonizing [158]. harsher [2776]. harvest [159]. harvested [56]. harveyi [539]. Hastings [140]. hatch [2832, 945]. hatched [892]. hatcheri [1194]. Hatcheria [2091]. hatchery [2696, 2786, 354, 1563, 1468, 802, 1555, 260, 278, 2219, 334, 1603, 1061, 1966, 232, 838, 2159, 1498, 476, 1789, 1376, 1509, 1501, 661, 2568, 1562, 2286, 2967, 1560, 475, 706, 41, 37, 2973]. hatchery-reared [354, 802, 278, 2219, 1603, 1061, 1966, 232, 2159, 476, 1789, 1376, 1509, 1501, 661, 2568, 2967, 1560, 475, 706, 41]. Hatchie [1601]. hatching [1285, 2729, 47, 2977, 455, 421, 2062, 597]. hatchling [1490, 573]. Hatoma [308]. haunts [2341]. Hawai’i [3030, 1845, 2049]. Hawaiian [888, 1391]. hawaiiensis [1391]. Head [1085, 244, 1245, 2943, 446, 2496, 2295, 2575, 642, 1394, 1574]. head-kidney [2575]. Head-shape [1085]. headed [2180]. heads [894]. headwater [3027, 2551, 2555]. headwaters [1841]. healing [2750, 602]. health [2773, 2297, 553, 688, 2866, 1465, 2691]. hearing [293, 2922, 285]. heart [2487, 697, 3075, 70, 79]. heart-cell [79]. hearts [1858]. heat [2880, 561, 1606, 2964, 837, 539]. heat-shock [561]. Heated [636]. heavily [687, 2173]. Heincke [2377]. Heithaus [549]. held [602]. Helicolenus [352, 2307]. Hellenic [2100]. help [2867, 2332]. Hemibrycon [2621]. hemiclonal [1435]. Hemigrammus [3006]. Hemiodontidae [1081, 2887]. Hemiodus [1081, 2887]. Hemiramphidae [2153]. Hemiscyllium [2833]. Hemisorubim [2620]. henlei [940]. henni [2005]. hepatic [1443, 2477]. hepatic-lipid [2477]. hepatocyte [10, 2381]. Hepsetidae [1088]. Hepsetus [1088, 2382]. Heptapteridae [375, 2843, 2749, 890]. Heptapterus [375, 2843, 890]. Heptranchias [2859]. herbicide [2566]. herbivore [2996, 1599]. herbivorous [1144, 535, 2027, 1388, 56, 1453]. Heritable [3048]. Heritage [2188, 2429]. hermaphroditic [2791, 1749, 235, 231, 1521, 1777, 1918]. Hermaphroditism [2243, 2762, 2328]. heroes [313]. Heroic [311]. herring [1580, 58, 1795, 2963, 394, 929, 1662, 2949, 1605, 1535, 230, 2855]. Herzegovina [1675, 1201]. heterenchelyids [2387]. heterochromatin [795]. heterochrony [2401]. heteroclitus [22, 18, 842, 3075, 3032, 1621]. heterodon [1083]. Heterodontus [1490, 2041, 2962, 298]. heterogeneity [40, 350]. heterogeneous [3084, 1661]. heterolepis [1083]. Heteromycteris [1156]. Heteropneustes [2321]. Heteropriacanthus [1760]. heterospecific [513]. Heterotis [1205, 2837]. Heterozygosity [608, 2776]. Hexagrammos [1435, 366]. Hexanchidae [2859]. hexanchiform [94]. Hexanchiformes [745]. hexapterus [1233]. Hg [1396]. hidden [1360]. hiding [1093]. hierarchical [325]. hierarchy [2535, 1029, 633]. High [2150, 1245, 2306, 2738, 1682, 331, 1237, 1409, 1735, 872, 2415, 452, 133, 1131, 1837, 426, 224, 1240, 32, 1425, 1719]. High-density [1682]. high-energy [1240]. high-feed [1409]. high-intensity [133]. high-resolution [1425]. high-temperature [1837]. high-throughput [1131]. Higher [2977, 2641, 2512]. Highest [2518]. highfin [1941]. highland [1478]. highlands [103]. highlight [1881]. highlighting [271]. highlights [384]. Highly [2239, 3071, 631, 268, 66, 1008, 2446, 886, 714]. Hilborn [1070, 1070]. hildebrandi [1601]. hillstream [100]. hilsa [2943]. Himalayan [645]. Himantura [459]. Himi [1754]. hind [2458]. hindgut [3079]. Hippel [550]. Hippocampinae [1758]. hippocampus [2461, 496, 1920, 2564, 498, 2625, 1339, 2461, 1769, 1314, 1516, 920, 1523, 500, 1204, 1515, 2718, 2623, 2644, 1786, 2948, 2965, 1171, 496, 3077]. hippoglossoides [200, 1682]. Hippoglossus [446, 2151, 856, 803]. hippurus [2029, 1617]. hire [1229]. Hirundo [1299]. Hisonotus [2606]. hispanica [99]. hispidus [1108]. Histochemical [2476, 2808]. histochemistry [304, 2871]. histocompatibility [910, 330, 1916, 1546, 843]. Histological [2644, 1932, 2328, 1769, 529, 2114, 2445]. histology [1539, 2871]. histology/histochemistry [2871]. historic [841]. Historical [2900, 496, 601, 3060, 967, 2240, 1633, 1541]. histories [468, 1716, 879, 519, 1430, 865, 1165, 1887, 1162, 1277, 672]. History [1210, 462, 812, 313, 1839, 2346, 1140, 2882, 631, 2918, 2217, 2865, 2203, 235, 1804, 31, 2844, 1429, 2457, 2461, 1852, 2966, 170, 1432, 2054, 1431, 920, 851, 1069, 459, 1598, 1851, 1835, 891, 1565, 426, 3045, 1725, 1411, 3056, 1242, 3068, 132, 1058, 1274, 1818, 160, 1754, 2458, 1897, 938, 2205, 1812, 2753, 275, 2443, 30, 1609, 1761, 2404, 2049, 2954, 1471, 1903, 2525, 752, 2223, 2204, 37, 1422, 188, 2511]. HNF [10]. hoc [2157]. Hogfish [2099]. Hokkaido [700, 2177]. Holbox [765]. holbrooki [2391]. hold [2946]. holding [2786, 1746, 2960, 92]. holistic [2718, 2009]. Holliday [110]. holocanthus [647]. Holocene [2109]. holocephalan [2687, 3059]. Holocephali [2560, 1533, 2421, 556]. holubi [721]. Home [563, 2172, 2067]. homeostasis [2867, 1265]. homeward [874]. Homing [2124, 878, 2255, 2763, 2184, 870, 2940]. Homogeneity [1661, 928, 670]. homologs [1917]. homologue [1597]. Homology [112, 2772, 2749]. homonycteris [2299]. homoplasy [411]. homozygosity [2785, 1032]. Honduran [103]. Honduras [1196]. Hong [2679]. honmoroko [1595, 2980, 2594]. hook [927]. hookear [1010]. hooks [1315]. Hoplias [54, 1537, 458]. Hoplosternum [928]. Hoplostethus [1970, 64, 1277]. horizontal [1152, 2363]. horkelii [3050]. hormone [2643, 1942, 1764, 1514, 3012, 1238, 2275, 2819, 590, 38, 417, 18, 214, 87, 2578, 2863, 2307, 698, 1298, 1862, 1931]. hormone-binding [1514]. hormone-transgenic [1764]. horse [663]. hospes [388]. Host [2907, 2140, 732, 104, 560, 904, 1080]. hotspot [1139]. hotspots [3088]. hottentotus [2559]. hound [2042, 294, 1812]. houting [307]. Howes [1232]. Hox [1896]. HPG [2774, 2444]. HPG-axis [2774]. hualiensis [669]. hubbsi [891, 648, 1249, 2775]. Hucho [295, 1459]. Human [2517, 2047, 2487, 1893, 1220]. Humane [178]. humic [202, 50]. humpback [3068, 2553]. humphead [63, 2309, 539, 237]. Humphries [1267]. hunger [1702]. Huso [2237, 968]. hybrid [2764, 1203, 43, 1880, 2087, 938, 615, 3024, 1006, 2495, 968, 1425]. Hybridization [650, 216, 1755, 2683, 2719, 1035, 1295, 1151, 1786, 1537, 736, 212]. Hybrids [2118, 662, 2167, 2320, 2165, 2505, 218, 825, 1390, 480]. hydraulic [1930]. hydro [875, 2353]. hydro-electric [2353]. hydrocarbons [44, 1838]. hydrochloride [419]. Hydrocynus [1299]. hydrodynamic [1718, 2026, 1545, 2923]. hydroelectric [1666, 2662, 3052, 1424]. hydrographic [175]. hydrographically [1661]. Hydrolagus [1042, 782, 2299, 2565]. Hydrological [1980, 1979]. hydrologically [239]. hydrology [78]. hydropeaking [1967]. hydrophila [668, 2213]. hydrophone [90]. hydropower [2698, 1502]. hydrostatic [855, 204]. hydrothermal [308]. hydroxycorticosterone [2355]. Hynes [2669]. hyoid [836]. Hypanus [2853]. hyperborea [1589]. hypereutrophic [1634]. hyperiid [946]. Hyperoglyphe [297]. hyperoxia [435, 842]. hyperoxic [1337]. hypersaline [1696, 2667, 2012]. hyperutilization [1402]. Hyphessobrycon [2558, 2395, 2484, 2004, 2400, 2437, 1841]. hypo [1505]. hypo-osmoregulatory [1505]. Hypomesus [3086]. Hypophthalmichthys [2936]. Hypoplectrus [2448]. Hyporhamphus [370, 2153]. hypostoma [2242]. Hypostominae [2509]. Hypostomus [1703, 1788, 2358]. hypotheses [2420, 1432, 2202]. hypothesis [1766, 221, 732, 762]. hypothesis-driven [762]. Hypoxia [1335, 524, 1748, 2065, 227, 195, 2880, 829, 898, 1827, 545, 435, 2325, 1336, 596, 842, 1829, 1828, 2512]. hypoxia-adapted [2880]. Hypoxia-induced [1335]. Hypoxic [1584, 114, 2323, 1184, 2534]. Hypseleotris [688, 363]. Hysteronotus [2003]. HZ08 [1106].

Ia [150]. Ibanaiko [1595]. Iberian [43, 99, 965, 30]. iberus [1714]. ice [579, 313, 1865]. ice-free [579, 1865]. Iceland [947]. Icelandic [2952, 3036, 2240, 1160]. Ichthyobodo [2704]. ichthyofauna [2021, 1518]. ichthyofaunas [1139]. ichthyological [2427]. Ichthyology [1210, 157, 466, 629]. ichthyoplankton [1720, 2262, 1021, 2010]. ICP [619]. ICP-MS [619]. Ictaluridae [1601, 107]. Ictalurus [1606, 1797, 1615]. idella [2809, 717, 2058, 2778, 668, 1867, 2671, 1241, 437, 2589, 826, 1106]. idellus [1638]. Identification [1415, 2385, 785, 535, 1570, 1947, 718, 1738, 974, 1934, 2309, 1209, 871, 1435, 1687, 813, 843, 2322, 933, 830, 1145, 844, 621, 1729, 149, 1026, 826, 3027, 582, 1720, 271, 1203, 1869, 216, 2116, 22, 1147, 2439, 2226, 2287, 915, 1685, 3078, 1317, 388, 360, 1596, 1097, 763, 497, 2232, 86, 3037, 2601, 1186, 1910, 82, 2743, 825, 1958, 146, 1425]. identified [2156, 1871, 1598]. identifies [2884]. identify [1620, 1314, 1516, 2312, 151, 1515, 808, 803]. Identifying [1261, 2611, 884, 1000, 2584, 1032, 1157]. identity [1766, 2508, 728, 702, 2546]. idus [1768]. Ifn [2961, 1529]. IGF1 [1524]. IgM [2276]. igneus [2484]. II [549, 286, 910]. IIA [2788]. IIB [843]. III [1241]. Ikawai [981]. ikei [379]. ilisha [3084, 2943]. illuminate [2735]. illuminating [2341]. illumination [2665]. Illustrated [1270]. image [1114, 2648, 2052]. imaging [1969]. immature [1701, 2117, 1848]. immediately [1152]. Immersion [932]. immobilization [1253]. immune [1493, 553, 958, 1626, 1575, 1241, 909, 1712, 1964]. immune-relevant [1626]. immunity [1729, 437]. immunoassay [241]. Immunocytochemical [2287, 2175]. immunohistochemical [642, 428]. immunohistochemistry [1262]. immunolocalization [801]. Immunological [447]. Immunology [109]. immunoreactivities [296]. immunostimulants [553]. immunosuppression [531]. Impact [1374, 743, 1473, 2016, 540]. Impacts [2989, 2767, 2630, 1187, 2472, 2877, 2404]. impair [2652]. impaired [2929]. impairs [29]. impedance [1585, 806]. imperialis [1649]. impingement [46]. implant [2465, 1684, 1346]. Implantation [573, 3031]. implanted [853, 3085, 2814]. implementation [1226]. implication [1246]. Implications [509, 33, 577, 317, 1665, 805, 1294, 785, 2526, 2759, 392, 1602, 1660, 1345, 294, 988, 340, 575, 2297, 2304, 174, 2977, 2300, 691, 500, 958, 2290, 2516, 2311, 2293, 716, 688, 938, 4, 2205, 635, 88, 2753, 1336, 786, 2259, 2178, 2107, 1198, 159, 496, 789, 1392, 448]. Importance [639, 60, 271, 52, 876, 207, 479, 342, 1095, 2546, 1881, 1821, 1332, 578, 2721]. important [366, 1847, 2970, 1301, 1309]. impose [1221]. impounded [2272]. impoundment [1113, 433, 1289, 339]. Imprint [1479]. imprinting [870, 2940]. improve [1563, 2036, 1562, 2890]. improved [2524, 1837]. Improvements [1220]. improving [875, 2933]. inaccessible [1277]. inactive [1925]. inanga [2107]. Inari [331, 1198]. inbred [997]. inbreeding [1594]. incidence [1936]. incidences [698]. including [290, 2488, 761, 1518, 1425]. inclusion [1712]. incompatibilities [1679]. Incomplete [1250, 891]. incorporating [479]. incorporation [2725, 2507]. increase [2469, 2405, 2977, 840, 2967]. Increased [403, 274, 1837, 1990]. increases [403, 62, 1488, 1469, 2311, 2022, 67, 1482, 1392, 362]. Increasing [2485]. increment [2955, 2266, 1079, 720]. increments [1421, 1367, 1721, 2261, 1378, 2980]. Incubation [2486, 1330, 702, 2729, 177, 2763, 597]. indagator [906]. independently [2990]. indeterminate [93]. Index [186, 359, 338]. India [907, 1108, 1479, 1951]. Indian [2874, 274, 266, 2260, 1517, 1760, 761, 799, 586, 1860]. indicate [2972, 2900, 2769]. indicated [1193, 1524, 430, 2678]. indicates [687, 2171]. indication [1059]. indications [1729]. indicative [1966, 844]. indicator [1065, 2015, 27]. indicators [58, 1615, 848, 1465]. Indices [544, 58, 1615]. Indies [1268, 389]. indigenous [2811, 347, 822, 838, 716]. indirect [2854, 201]. indirectly [2866, 708]. Individual [685, 256, 2041, 3028, 2523, 3071, 728, 1752, 1236, 355, 1785, 3078, 1261, 151, 1028, 38, 497, 1757, 1831, 86, 947, 135, 2192, 2406, 2466, 2284, 514]. individual-based [135]. Individually [1532, 803]. individuals [2719, 2272, 2978]. Indo [2517, 1755, 1161, 2289, 2230, 1447, 2133, 155]. Indo-Australasian [155]. Indo-West [1161, 2133]. Indonesian [1600]. induce [565, 732, 133]. induced [1335, 62, 2264, 3012, 104, 1382, 693, 1185, 856, 723, 2613, 1428, 2769]. induces [1053]. Inducible [1121]. inducing [9]. Induction [2132, 2346, 2614, 71, 850]. industry [1222, 1228, 2906, 1297, 1229]. inermis [2817]. inexpensive [585]. infected [1092, 2345, 989, 212]. infecting [1080]. Infection [1681, 1351, 2405, 1030, 560, 813, 2616, 1674, 407, 1638, 826, 1736]. inference [1954]. Inferred [1019, 81, 701, 770, 395, 334, 915, 2130, 148, 15, 2374, 1714, 2885, 363, 2835, 2153]. infiltration [2543]. inflammatory [903]. inflation [2960]. inflow [233]. inflows [481]. Influence [2869, 227, 2330, 175, 3003, 2168, 2323, 2539, 1626, 1893, 88, 383, 2283, 1579, 804, 1705, 249, 1616, 566, 533, 2663, 1751, 1047, 610, 707, 2142, 2103, 1410, 2087, 1902, 1406, 1387, 187, 1900, 3073, 2544, 3010, 2120, 326, 3057, 2616, 417, 599, 78, 1673, 1656, 1658, 273, 233, 481, 478, 2021, 542, 1418, 405]. influenced [998, 1618, 2137, 2223]. Influences [712, 87, 2487, 2264, 702, 2998, 2026, 2187, 28, 1089, 2008, 1434, 1851, 2064, 2492, 1363, 1148, 1789, 2188, 2062, 236, 70, 422, 362]. influencing [785, 791, 560, 558, 594, 2553]. inform [1816]. Information [2163, 2194, 2195, 2206, 2207, 2245, 2246, 2278, 2279, 2316, 2317, 2342, 2343, 2365, 2366, 2392, 2393, 2412, 2413, 2432, 2433, 2455, 2456, 2474, 2475, 2498, 2499, 2521, 2541, 2573, 2604, 2634, 2659, 2692, 2709, 2734, 2768, 2794, 2823, 2841, 2861, 2875, 2896, 2919, 2945, 2986, 3019, 3042, 3065, 2520, 728, 2569, 1384, 170, 2247, 1148, 1010]. informative [1131]. infrastructure [1930, 2933]. ingens [2564]. ingestion [1053, 2857]. ingressing [945]. inhabiting [1781, 2747, 801, 1730]. iniae [1026]. initial [2770, 1593, 2404, 2230, 869]. initiation [923]. Inland [982, 1227, 158, 1593, 1999, 2649, 935]. inlet [1595]. innate [958]. inner [2119]. innervation [1208]. innesi [2401]. innominatus [771]. innovation [2898]. inodorus [1846]. input [1845]. inputs [1230]. inseminating [2000]. inshore [2122, 450, 715, 1345, 1687]. insight [1905, 1565, 2784, 1043]. Insights [2121, 2421, 2197, 1725, 8, 792, 2569, 1112, 879, 2828, 2942, 1559, 1969, 2460, 630, 2196]. inspection [535]. inspirational [25]. instability [1482]. Institutional [1564]. insular [1150]. insulin [590, 87]. insulin-like [590, 87]. intake [1925, 1694]. intakes [875]. Integrated [2559, 2696, 1557, 2848, 159]. integrating [985, 1220, 2789]. Integration [2230, 1827, 2661]. Integrative [2782, 2133, 398, 1995, 529, 1881]. integument [2943]. intensity [1351, 133, 207, 406]. intensive [2454]. Intentional [2674]. Inter [1849, 1713, 3088, 36, 1518]. inter-annual [3088]. inter-decadal [1518]. inter-family [36]. Inter-population [1713]. Interaction [2999, 2110, 1238, 958]. Interactions [2429, 2190, 2423, 318, 636, 841, 991, 876, 1431, 1584, 1667, 1221, 541, 558, 1538, 2409, 2556, 284, 514]. Interactive [171, 1683]. Interannual [2356, 3037, 1814]. interbreeding [2157]. Interdecadal [663, 2284]. interference [1022, 1241]. interferon [2961]. Intergeneric [591]. interior [2252]. interjurisdictional [2205]. interleukin [1200]. interleukin-1 [1200]. interlinked [1730]. intermedia [2236]. intermediate [2758, 3059, 1459]. intermedius [1437, 2758]. intermittent [1829, 1828, 1819]. intermittent-flow [1829, 1828, 1819]. internal [2]. International [2292, 2693, 356, 237]. Interplay [514]. interpopulation [1450]. interpreting [1835]. interrenal [1715]. interruptus [1546]. intersex [2688]. intersexual [1317, 1757]. interspecies [1425]. Interspecific [2727, 2639, 1652, 1431, 2614, 1151, 2666, 1451]. Interstitial [1063]. intertidal [2502, 1591, 2184, 1531, 1324, 2172, 1255, 810, 2485, 801, 2438]. interval [1779]. intestinal [2808, 2306, 1511, 2703]. Intra [1599, 1849, 2666]. intra-interspecific [2666]. Intra-population [1599, 1849]. intracoelomic [3031]. intramuscular [2543, 2784]. Intraspecific [2754]. intrinsic [1556, 1166, 1679, 3090, 1392]. Introduced [1411, 1916, 3010, 1538, 2049, 319]. introduction [2517, 2391, 1745, 160, 1541]. introductions [49, 1559]. Introductory [108, 2145]. introgression [1194, 1441, 615, 2157]. Introgressive [2719, 2683]. invading [381]. invalidates [1010]. Invasion [1999, 2637, 580, 2530, 2551, 1354, 2711, 691, 2636, 517]. invasions [2061, 995, 1997]. Invasive [2569, 2471, 2654, 1146, 2339, 1869, 1339, 1998, 2683, 2856, 66, 269, 1378, 2298, 1067, 1151, 852, 2152, 585, 1897, 816, 1066, 798, 1077, 2936, 199, 825]. invertebrates [2493]. investigate [1528, 1880, 2583]. Investigating [942]. investigation [2722, 2771, 2718, 1829, 252, 2903]. Investment [118, 2264, 2770, 164, 1844]. inviable [480]. involved [2764, 2665, 2213, 1241, 437]. involves [1455]. Involving [1226, 3070]. ion [1488, 1328, 1511]. ion-transporting [1511]. Ionian [2868]. ionic [2867]. ipophrys [2172]. irak1 [1729]. Iran [1627, 950, 935]. Iranocypris [950]. Iraq [2836, 261]. iratapuru [1081]. Ireland [1351, 1147, 1293, 130, 949, 14, 2976]. irf9 [2961]. iridovirus [1738, 79, 2605, 212]. Irish [1801, 652, 2761]. iron [1412]. irritans [1736]. Isacia [1263]. ISBN [3016, 3017]. ishikawae [838]. Isistius [2554]. Island [2564, 2386, 2269, 1950, 765, 912, 2356, 2900, 2996, 660, 2458, 2438]. Islands [1809, 1196, 274, 2712, 349, 1139, 1177, 1222, 785, 2093, 2051, 2771, 973, 1786]. Isle [1705, 2616]. Isles [1687, 778, 2912]. islets [1730]. isoforms [22]. isogenic [951]. isolated [678, 624, 2996, 1288, 1031]. Isolation [2116, 1540, 1703, 1494, 625, 972, 1492, 1650, 1648, 256, 1300, 2451, 1478, 24, 366, 1250, 2555, 1383]. Isosmotic [1664]. isotocin [1313]. Isotope [2200, 2508, 1396, 769, 2968, 2189, 1623, 2910, 770, 1112, 518, 841, 2379, 2741, 817, 2071, 2374, 2018, 1677, 2870, 2406, 884, 1277, 1307, 1451, 578, 310, 3089, 2721, 3009]. isotope-ratio [2406]. isotopes [166, 1138, 975, 1852, 2048, 994, 768, 1028, 993, 665, 689, 2283]. Isotopic [1751, 2203, 592, 1668, 2546, 1898]. Issue [2163, 2194, 2195, 2206, 2207, 2245, 2246, 2278, 2279, 2316, 2317, 2342, 2343, 2365, 2366, 2392, 2393, 2412, 2413, 2432, 2433, 2455, 2456, 2474, 2475, 2498, 2499, 2521, 2541, 2573, 2604, 2634, 2659, 2692, 2709, 2734, 2768, 2794, 2823, 2841, 2861, 2875, 2896, 2919, 2945, 2986, 3019, 3042, 3065, 2520, 2658]. issues [1227]. istiophorid [954, 2232, 2685]. Istiophorus [1307]. Isurus [1767, 1718, 2092, 1725]. itajara [2024]. Italy [510, 1048, 2332]. Itasenpara [1754]. Iteroparity [2732, 2107]. Ituglanis [1304].

J [289, 142, 140, 1265, 549, 690, 1269, 1371, 1173, 1211, 109, 739, 671, 139, 356, 548, 461]. jack [820, 3014, 281, 2879, 345]. jacks [2488]. Jackson [1490, 2041, 2962]. jacksoni [2957, 122]. jamaicensis [2480]. Jamieson [183]. Janeiro [567]. Japan [2211, 700, 838, 1595, 2594, 1754, 2411, 696, 2166, 308, 2177, 1400]. Japanese [2221, 1920, 1178, 692, 2579, 2741, 17, 1085, 1040, 2267, 581, 1032, 1641, 3057, 469, 3026, 2222, 1512, 661, 2568, 214, 705, 475, 1291, 425, 622, 2835, 1197, 1717]. japonica [2451, 2579, 1085, 2267, 3026, 1197]. japonicus [2865, 1846, 1130, 518, 3014, 281, 469, 1651, 661, 2568, 214, 345, 1156]. Jari [1081]. Javaés [2509]. jaw [2612, 603, 1408]. jellyfish [994, 2917, 993, 2682, 469, 946, 689]. Jennings [1098]. Jenyns [2796, 597]. Jenynsia [2577, 1894]. jig [927]. jocu [1338, 163, 45]. John [1232, 3016, 3017, 1405]. Jolles [2737]. Jolly [2465]. Jones [2669]. Jonsson [672, 672]. Jordan [606, 2384]. jordani [2831]. José [466]. Journal [2457, 2658]. Journals [1374]. Journey [955]. July [2635]. June [226]. junior [2127, 1343]. jurisdictional [1027]. Juruena [2437]. justification [391]. Juvenile [3013, 2594, 963, 1042, 2832, 686, 2668, 444, 1283, 1503, 2786, 1616, 1294, 2955, 1528, 195, 715, 2705, 1808, 2399, 2596, 2759, 338, 3001, 3029, 1188, 253, 534, 767, 3007, 583, 2044, 2370, 2013, 899, 1030, 518, 134, 1101, 2187, 2345, 278, 1508, 1062, 393, 17, 1061, 941, 546, 3030, 2827, 204, 44, 1797, 2680, 34, 1134, 2477, 467, 2744, 1536, 734, 1204, 1259, 2693, 2481, 1382, 2704, 2112, 423, 279, 2430, 1628, 1355, 2236, 1472, 1578, 1037, 2131, 2962, 307, 236, 2497, 1676, 1284, 406, 2012, 2695, 3024, 1497, 122, 128, 1249, 1912, 2085, 931]. juvenile [618, 1793, 1730, 428, 171, 677, 1794, 1856, 1282, 1862, 2562, 1664, 1243]. Juveniles [859, 897, 1245, 1367, 855, 3031, 1937, 59, 1120, 3014, 1598, 281, 64, 1142, 945, 300, 1016, 1376, 1744, 2547, 163, 1157, 2286, 225, 1346, 578].

K-dependent [2510]. Kajikia [2483, 960, 2268]. Kamchatka [2383]. kamoharai [754, 1996, 121]. Kapoor [140, 568, 588, 109, 439]. Kareius [1390]. karst [2414]. karstic [1035]. karyomorphs [1537]. karyotype [731, 736, 890]. Karyotypic [1788, 928]. Kasai [1906]. Katsuwonus [2791, 2141, 799]. kauderni [812]. kaupii [541]. Kawakami [142]. Keeling [274]. keeping [2533]. kelberi [2339, 2152]. Kelloggella [2438]. kelp [2501]. kelt [1572]. Kemp [569]. Kendall [462]. Kerguelen [2478]. keta [2255, 335, 874, 2704, 1789, 1594, 2118]. ketotestosterone [241]. Key [2890, 2617, 2845, 1414, 2531, 1598, 2549, 2666, 2538, 2540, 2743, 1244]. keystone [2766]. Khodorevskaya [438]. Khor [2836]. kidney [471, 286, 1769, 39, 1711, 2295, 2575, 642, 1394, 1574]. Kiel [136]. kill [274]. killed [1712]. killifish [2681, 2839, 707, 2403, 2323, 240, 2062, 2848, 3075, 436, 989, 1885]. killifishes [2164, 1714, 2349, 1770]. killing [178]. kimberleyensis [918]. Kin [707, 722]. Kinematics [1836, 2746, 2303, 820, 499]. king [848, 483, 2515, 2310]. Kingdom [2905]. kingfish [2178, 1036, 1482]. kingsleyae [1088]. Kinship [1104]. kishinouyei [1494]. kisspeptin [8]. kisutch [1764, 1429, 2701, 2994, 2275, 1430, 894, 38, 1536, 2578, 2352]. kit [241]. Klemetsen [671]. Klimley [1373]. Kner [1235]. knife [731]. knifefishes [2772]. Knight [1404]. Knodus [222]. Knoll [308]. know [2209, 1070, 1371]. knowledge [1271, 49, 2531, 3, 635, 634]. known [2003, 2801, 2730, 2586]. Knysna [3077]. kob [1846, 1651]. koi [2086, 2167, 913, 1137]. koilomatodon [852]. Kokanee [1840, 916, 622]. Kong [2679]. Kopf [1211]. Korea [148]. krameri [1675, 2419, 1201]. kranponhah [2002]. Kryptolebias [1749, 2944, 707, 1866, 1896, 670, 3023]. kuhlii [2968, 2189, 351]. Kuiter [141, 208]. Kunz [439]. Kuwait [776]. KwaZulu [2855]. KwaZulu-Natal [2855]. Kwilu [1906]. kype [603]. Kyphosidae [1619]. kyphosids [1591].

L [1099, 1266, 2172, 1282]. L-tryptophan [1282]. L. [2591, 85, 2952, 269, 2941, 1270, 452, 435, 1729, 2967, 1692, 1505]. labelling [419]. Labeo [1541]. Labeobarbus [881, 1437, 918, 1796]. Labidochromis [2618]. Laboratory [2089, 1895, 1570, 2593, 2737, 853, 2258, 2736, 1134, 2529, 2804, 594, 1369, 1829, 117, 1346, 228, 2102]. laboratory-reared [1369]. Labrador [2055]. labrax [797, 452, 1655, 1259, 419, 1292, 2583, 1647]. labrid [231, 512, 966, 418]. Labridae [119, 63, 2099, 2746]. labrids [2544]. Labroides [2190, 846]. Labrus [2116, 902, 1417, 3037, 224, 1135]. labyrinth [2234, 2394]. lacerdae [54]. Lachnolaimus [2099]. Lack [2962, 1010, 3008, 1737, 2377]. lacking [1308, 2004]. lacrymatus [2787]. lactate [1254]. lacustrine [1923, 817, 2081, 3000]. lacustris [2060, 2566]. laeta [2082]. laevifrons [1619]. laevigatus [484]. laevis [1257, 2128, 2337]. Lagocephalus [2817, 484]. Lagodon [2398]. Lagoon [1993, 1758, 1595, 1219, 1696, 998, 2256, 1731, 2018, 2016, 1992]. lagoonal [1950]. lagoons [98, 1991]. Lake [1923, 1595, 2205, 236, 284, 580, 2786, 667, 819, 1450, 1163, 192, 350, 835, 1035, 3083, 2656, 1957, 1413, 1634, 1962, 2353, 2930, 2651, 313, 344, 2641, 2797, 965, 1910, 3000, 936, 1985, 2881, 935, 1050, 1750, 331, 1198, 2716, 2715, 1090, 2752, 881, 819, 1679, 1410, 2130, 2594, 147, 229, 2889, 270, 1796, 332, 415, 1624, 1343, 1604, 2397, 1400]. lake- [965]. lake-bound [2716]. lake-resident [667, 1163]. lake-run [331, 1198]. Lake-specific [236]. lakes [2239, 975, 1022, 2383, 202, 102, 50, 320, 558, 2901, 2761, 89, 2881, 125, 1031, 326, 607, 1910]. lalandi [2178, 1036, 1482]. lalandii [2770]. Lamiopsis [2468]. Lamna [3082]. lamnid [2234]. Lamniformes [2234]. Lampanyctus [1407]. Lampetra [964, 286, 1117, 1401, 348, 213, 1549, 515, 1492]. lamprey [286, 2993, 1117, 2258, 1842, 709, 602, 1549, 2347, 515, 1492, 2259, 2102, 909]. lampreys [1665, 964, 1852, 1069, 382, 213, 1122, 2252, 60]. Lampridiformes [1174]. lance [1233, 944]. lanceolatus [2722, 3024]. Land [517, 2944, 1751, 2546]. land-use [2546]. landings [776]. landlocked [2869, 1509, 1501]. Landmark [550]. landscapes [2525]. lane [163]. langsdorfii [1040]. lantern [1612]. lanternfishes [1060]. laparoscopy [1253]. lapillus [948]. Large [800, 834, 1140, 1629, 2705, 1589, 2581, 350, 1224, 2777, 449, 2999, 2656, 1486, 1634, 2081, 2651, 1904, 1797, 996, 1725, 1891, 78, 2648, 387, 1050, 2329, 2981, 1394, 2835, 1736]. large-egg [2835]. large-scale [1140]. largemouth [918, 1555, 934, 455, 734, 2951, 1253, 599, 1375, 2562, 1418]. larger [1744, 3069]. largest [918, 387, 2030, 339]. Larimichthys [996, 1736]. larva [1300, 423]. Larvae [859, 471, 710, 2450, 2391, 482, 1846, 802, 1858, 1491, 97, 1348, 2799, 2942, 854, 2503, 1598, 1545, 1416, 206, 592, 1152, 1655, 1442, 945, 1939, 3026, 2232, 2710, 1651, 1369, 661, 407, 1673, 201, 1825, 936, 1157, 280, 2947, 181, 2271, 1768, 203, 225, 2329, 1160, 1063, 804, 1036, 2120]. Larval [2834, 2464, 1188, 54, 2781, 1483, 2236, 2318, 2832, 2722, 2609, 2955, 2399, 2998, 226, 2510, 1852, 2592, 1588, 134, 3063, 2993, 2261, 175, 2075, 451, 590, 1595, 2066, 893, 455, 2824, 2557, 2970, 1659, 1292, 2417, 529, 1381, 1656, 971, 2753, 2062, 478, 365, 2259, 597, 2780, 696, 215, 1162, 1912, 1793, 1156, 2102, 2792, 2715]. Laser [1262, 1693, 619, 387]. Laser-capture [1262]. last [2383, 3066]. late [2655, 1738, 1260, 118, 1459, 892]. late-summer [2655]. latency [2497]. Lateolabrax [518, 2444]. later [1434]. lateral [210, 1489, 1377, 2923, 2018, 2932, 1921]. lateral-line [2923, 1921]. Lateralisation [2098, 2883]. laterality [1179]. Lates [802, 837, 1676]. laticeps [766]. Latimeria [39]. latipes [1178, 1017, 247, 680, 1728]. latipinna [205, 922, 2595, 1945]. latitude [1356, 512, 224]. latitudes [966, 2641]. latitudinal [611, 675, 12, 1599, 2139]. Latridae [2844]. latus [1161]. Lauer [1098]. Laurentian [326, 1910]. lavaretus [1849, 29, 350, 1357, 1117, 1240, 936, 399]. law [2377]. Lawrence [382, 863, 1052, 1057]. layer [2086, 3004, 2871]. layers [2464]. laying [2063, 1000]. Lea [2816]. lead [1920, 819]. leader [1287]. leadership [2742]. leading [284]. leads [1402, 47]. leaping [2668]. learn [68, 1329]. learned [1225, 2369]. learning [2078, 2742, 1287, 3001, 2537, 2827, 1639, 345, 1657]. least [1601, 585, 2225, 839]. least-invasive [585]. leatherjacket [2511]. Lebanese [1909, 2449]. Lebiasinidae [839]. Lebranche [2430]. lectin [1312, 2083]. lectin-like [1312]. lectins [1867]. led [2042, 3051]. left [147]. leg [2380]. leg-skate [2380]. Legacy [550, 2383]. legislation [777]. Legumin [1093]. Leiarius [2495]. lemon [1954, 1808, 253, 767, 1961, 2814]. Length [3033, 2523, 1316, 2326, 1765, 1421, 2958, 610, 1357, 56, 343, 613, 479, 2176, 1095, 1641, 3081, 2550, 2131, 2694, 656, 46, 423]. length-at-age [656]. lengths [2816, 499]. leniusculus [1610]. lentic [1586, 2723, 2350]. lentiginosus [659]. Leon [1753]. leopard [2185, 1195]. leopardus [1353, 1305]. leopoldi [2624]. Lepeophtheirus [2616, 904, 164]. Lepidion [643]. Lepidocephalichthys [353]. Lepidorhombus [1206]. Lepisosteus [695, 1775]. Lepomis [249, 31, 636, 3012, 21, 1586, 445, 1015, 451, 1028, 965, 1375, 2860, 2210, 319]. Leporinus [1235, 714, 1890, 1650]. leptocephali [2870]. leptocephalus [2127]. Leptocottus [2331]. Leptonotus [2971]. Leptopanchax [2839]. leptorhynchus [1378]. Lernaeocera [1399]. less [309]. Lessepsian [2230]. Lesser [2242, 227, 2362, 378, 2721]. lessons [1225, 161]. lethal [2615, 587, 1305, 2204, 3009]. Lethenteron [2780, 2252, 909, 664]. lethostigma [1101]. Lethrinidae [1046, 1157]. Lethrinus [1254, 1046, 340, 96]. leucas [282, 234, 1806, 3046, 796, 1094]. Leuciscidae [2739]. Leuciscus [510, 1768]. leucocytes [531, 1574]. leucomaenis [2336, 1376, 1648]. Leucoraja [792, 1052, 1057, 919, 252, 2355]. leucorhynchus [1813]. Leuresthes [1408, 2670]. levators [1156]. level [403, 1702, 1592, 2738, 273, 331, 2223]. levels [2321, 1245, 2303, 566, 692, 1514, 1409, 101, 343, 1238, 590, 2515, 2477, 1512, 2311, 3062, 2265, 1255, 2284]. lewini [3029, 1587, 2040]. lewisi [2747]. Lh [3062]. liberty [2092, 2814]. library [2161]. lice [2668, 2345, 904]. Lichtfouse [1374]. lidocaine [2543]. Lieschke [142]. Life [1140, 2217, 2865, 235, 2461, 2966, 462, 2054, 2403, 459, 3045, 1242, 1754, 2458, 2341, 30, 1609, 1887, 2404, 2394, 752, 2204, 1422, 188, 2511, 311, 672, 1385, 2346, 631, 468, 2622, 1804, 1716, 2770, 2844, 1429, 2482, 2137, 1168, 1852, 1672, 170, 1432, 574, 2187, 278, 1431, 2320, 920, 851, 519, 1293, 732, 3078, 1957, 1069, 1598, 1430, 1851, 1835, 1565, 426, 1725, 1411, 3056, 3068, 132, 1058, 1274, 160, 1897, 938, 1940, 1078, 2205, 1812, 2753, 1165, 275, 2443, 1761, 738, 2049, 2954, 987, 1256, 1471, 1903, 203, 2525, 869, 1277, 37, 2329, 1552, 1197]. Life [1018]. Life-history [1140, 2217, 2966, 1242, 2458, 30, 1609, 2204, 1422, 188, 2511, 631, 1804, 2844, 1429, 1852, 170, 1432, 851, 1069, 1598, 1835, 1565, 1411, 3056, 1058, 1274, 160, 1897, 1812, 2753, 275, 1761, 2954, 2525, 37]. life-phase [1385]. life-stage [203]. lifestyle [2228]. ligand [1956]. Light [2931, 564, 1976, 457, 3051, 2825, 3043, 666, 2824, 827, 436, 2936, 3008]. light-trap [2825]. lightly [752]. lights [2934]. like [266, 2058, 1597, 1312, 590, 2213, 1638, 87, 83, 1380, 146, 1964]. Limanda [655]. limbata [2378]. limbatus [1810, 915, 2165, 2680, 2436, 2852]. Limia [1716]. liming [2440]. limitation [639]. limitations [2202]. limited [2055, 417, 2085, 2210]. limits [1414, 1858, 1784, 1678, 2471, 815]. Limnochromini [1343]. Lin [1372]. lindbergi [1640]. line [2880, 1567, 1489, 2415, 443, 2295, 1137, 2381, 2501, 1377, 2923, 265, 2932, 1529, 997, 1568, 2888, 2088, 1921, 1394, 79, 2605, 2584, 1968, 1717]. lineage [1601, 624, 2740, 2672]. lineages [2987, 1569, 1760, 891, 2611, 552]. linear [913, 1763]. lineata [2577]. lineatum [2427]. lineatus [1092, 885, 2113]. linebelly [3090]. lined [1769]. lines [694, 951, 1195, 2575]. lingcod [1134]. link [337, 1179, 946, 2653, 502, 548]. linkage [335]. linked [1925, 315]. Linking [193]. Liobagrus [523]. lionfish [1869, 2530]. lionfishes [2338]. liorus [226]. Liparidae [1391]. lipase [1443]. lipc [1443]. Lipid [467, 665, 2615, 611, 926, 2189, 1409, 118, 2671, 2477, 3010, 2481, 1615]. lipofuscin [409]. Lipophrys [1367, 2485]. lipopolysaccharide [1638]. lipoprotein [692, 1443]. list [2991, 895]. listed [2429]. literature [2880, 2180]. litter [2561]. little [792, 2449, 252, 1622]. littoral [1953, 1538, 2540, 125]. littorale [928]. litulon [1707]. lituratus [1485]. live [1158, 2598, 1252, 967, 2124, 2585]. live-bearer [1252]. live-bearing [967, 2585]. livebearer [724]. livebearing [646]. lived [1121, 2140, 231, 2300, 2511]. livelihoods [2906]. liver [1572]. Living [1209, 2139, 2248, 1148, 2535, 1480, 2907, 1879, 2249]. Liza [521, 1474, 3047, 2430]. lizardfishes [2051]. loach [81, 100, 645, 2477, 2215, 2411, 16, 2363, 2851, 188]. load [1741, 1398, 2103]. loading [693]. loads [1254]. Local [1705, 28, 2580, 1899, 330, 449, 651, 1363, 360, 273, 2166, 128, 542]. localities [1308]. locality [2004]. localization [2058, 2222, 214]. locate [505]. Location [2579, 166, 574, 2649, 973]. locations [608]. lochs [2793]. loci [954, 608, 1737, 1540, 1258, 1703, 1762, 1131, 1494, 830, 972, 1492, 1650, 1648]. Locomotion [588, 1326]. locomotor [1080, 2654]. locomotory [2781]. locus [77, 1938, 1594]. Logan [994]. logger [808]. Long [1506, 631, 1645, 2662, 2735, 1238, 1441, 3085, 1936, 1424, 1505, 362, 961, 2950, 2326, 2217, 2296, 2140, 231, 301, 800, 1510, 1303, 1961, 1366, 187, 204, 1806, 2381, 1331, 313, 2644, 2466, 499, 2425, 2511]. long-chain [2381]. long-distance [1806]. long-fin [2950]. long-lived [2140, 231, 2511]. long-nosed [2217]. long-period [800]. long-snouted [2644, 2425]. long-tailed [1303]. Long-term [1506, 631, 1645, 2662, 2735, 1238, 1441, 3085, 1936, 1424, 1505, 362, 961, 2326, 301, 1510, 1961, 1366, 187, 1331, 313, 2466]. longa [772]. longer [2607, 2051]. longer-term [2607]. longevity [1954]. longfin [1629]. longfinned [865]. Longhurst [604]. longipinnis [2950, 1754]. longispinis [891]. longline [2057, 785, 781, 927]. longlining [1961]. longnose [2224, 2853, 2055, 695, 1775]. longnosed [1573]. longsnout [2623]. looking [3049]. loop [799, 803, 839]. Lophiidae [644, 1707]. Lophiiformes [246, 644, 1630, 2960, 1707]. Lophiosilurus [2220]. Lophius [1706, 2290, 1214, 1707]. Lorenzini [2513]. lori [1609]. loricariid [3056]. Loricariidae [2185, 2509, 1703, 2549, 1788, 2723, 591, 2606, 2586, 2358]. Loricariinae [2549, 591]. lose [409]. loss [1402, 2566, 2473, 1418]. Lota [2591, 3035, 1959, 893, 845, 2132]. lotic [1586]. loud [1415]. louse [2668, 164]. Low [1149, 1168, 2566, 485, 2515, 2490, 564, 1356, 1409, 1543, 1900, 115, 264, 1919, 3072, 239, 2354, 1566, 944, 670, 32, 2209, 3008]. low-frequency [1900, 944]. low-latitude [1356]. Lowe [1839]. Lowe-McConnell [1839]. lower [2617, 226, 603, 2741, 30, 1903, 2442, 283, 1088]. lowland [1842, 1904, 374, 2272, 348]. lpl [1443]. Ltd [3016, 3017]. Lucania [2681, 436]. lucidus [2884]. lucioperca [2482, 1916, 2930, 1260, 401, 654]. lucius [1743, 2777, 2930, 546, 1915, 2761, 811, 89, 1774, 738, 181, 2334, 3009]. Luer [439]. Luis [2500, 466, 2822]. luminous [923, 2054]. Lumpfish [2832, 2154, 1423]. lumpsuckers [1640]. lumpus [2832, 2154, 1423]. Lunar [93, 2998, 2120, 976, 835, 2, 283]. lunar-related [2]. lungfish [1289, 70]. luochengensis [2414]. Lupinoblennius [2327]. lupus [1306]. luteinizing [2643, 214]. lutjanid [2688]. Lutjanidae [163]. Lutjanus [2017, 1338, 343, 2309, 2423, 3068, 1678, 163, 117, 1793, 45, 1439, 539]. Luvaridae [1649]. Luvarus [1649]. Luxilus [2807]. Luzzatto [1516]. Lycodes [1589]. lyra [896].

M [289, 112, 605, 110, 1265, 549, 983, 1172, 1209, 1268, 1211, 183, 138, 356, 980, 981, 1264, 1266, 1404, 1480]. M. [2676]. maccoyii [553]. Maccullochella [379, 686, 538]. macdonaldi [2375]. Macedonia [85]. mackerel [576, 1643, 663, 3014, 281, 1242, 2879, 345]. macloviana [1166]. maclovinus [651, 2932]. Macquaria [2297, 2168, 609, 615, 2716]. Macquarie [2716]. macraei [2091]. Macrhybopsis [2346, 1903]. Macrobenthic [655]. macrocephalus [1310, 1285, 264, 1167, 2166]. macrochir [848]. macrochirus [249, 31, 21, 1586, 445, 1015, 451, 1375, 2860, 2210]. Macrodon [2031]. Macroecological [1781]. Macrohabitat [26]. macrolepis [3038]. macroparasites [1624]. Macropharyngodon [1195]. macrophyte [721]. macropomum [2867]. macroscopic [1777]. macrostigma [2672]. macrotidal [2745]. Macrourus [384]. maculata [2322]. maculatus [901, 1941, 597, 2107, 2444]. maculosus [2757]. Madagascar [1806]. MADAM [2009]. made [1912]. Madeira [2400, 1666]. Madre [2549]. madtom [1601]. maena [386]. Magadi [1337]. Magdalena [2621, 1981]. maggot [1065]. Magnetic [1489, 1900, 2077]. Magnetoreception [2924]. Mahachai [217]. Maharlu [1627]. mahseer [2114]. main [2200, 1641, 2447]. main-river [2200]. Maine [1698, 1449]. mainstay [466]. maintain [2867]. maintained [2804]. maintenance [2149, 122]. majalis [1621]. Major [330, 910, 266, 1916, 114, 1546, 843, 2863, 1695, 47, 334, 2827, 2348]. Makaira [2910]. Makeyeva [859]. making [1148]. mako [1767, 1718, 2092, 1725]. Malaŵi [1004]. malabaricus [1537, 458]. Maladaptation [1560]. Malaysia [719, 719, 3084, 1946]. Malaysian [153]. maldivensis [2127]. Maldives [2676]. Maldonado [1894]. Male [505, 683, 2388, 1181, 249, 2264, 2182, 1752, 216, 2847, 1398, 2285, 2117, 524, 355, 603, 1303, 1202, 1069, 51, 559, 1017, 207, 118, 1918, 288, 2849, 776, 3037, 2623, 3069, 1919, 2774, 9, 3024, 1119, 904, 2858, 197, 223, 480, 540]. male-specific [216]. male-to-female [2847]. males [793, 2076, 292, 1251, 229, 2095, 412, 415, 2674, 1025, 834, 197]. malformations [598]. Mallotus [2717, 3069]. malma [1524]. Malta [2629]. man [1912, 1705]. man-made [1912]. Management [110, 548, 1140, 1222, 2156, 2569, 785, 1804, 2199, 2146, 2162, 2526, 2759, 392, 161, 1046, 1845, 783, 1147, 780, 340, 575, 1995, 2304, 1221, 774, 1143, 2348, 1496, 326, 2901, 716, 938, 2158, 1988, 635, 3034, 1223, 638, 2913, 2178, 1198, 159, 494, 199, 2690, 2691, 496, 831, 1099, 569, 3016]. Managing [1558, 361, 1554, 2533]. mandarin [457, 1005]. mandarinfish [3004]. mandi [2834]. mandi-pintado [2834]. mandibular [836, 808]. Mangrove [1402, 1749, 2944, 707, 1734, 1866, 2010, 2120, 1678, 163, 2745, 383, 2009, 1730, 3023]. mangrove-fringed [163]. mangrove-related [2009]. manipulating [970]. Manipulation [2547, 2491]. manner [966]. Manta [973, 3071, 198, 1699, 2676, 2429, 198, 1699, 973]. Manual [2958]. manufactured [589]. many [2585]. Maori [981]. Mapping [2036, 2197, 714, 1526]. maps [1218]. maraena [365]. marble [1888, 1039]. marbled [1711, 2095, 2325, 1247, 79]. march5b [1736]. Marches [1048]. Margaret [25]. Marginal [1079, 3070]. marginatoides [523]. marginatus [1739, 1188, 2557]. mariae [1842, 798, 1077]. marianae [1602]. Marine [2435, 462, 141, 2689, 3016, 604, 1496, 956, 2467, 1133, 1722, 2431, 3071, 745, 535, 1894, 221, 2119, 473, 1563, 687, 2359, 2426, 2079, 1721, 876, 260, 1784, 2735, 2524, 2472, 2275, 2846, 2712, 880, 1406, 1583, 1435, 2446, 1693, 1142, 2381, 1802, 1909, 1698, 1756, 2902, 1926, 2805, 2926, 2161, 1263, 265, 1825, 2905, 577, 2351, 3055, 554, 1682, 2648, 2410, 218, 2748, 365, 1039, 1599, 882, 131, 2283, 884, 1465, 283, 2147, 405, 2893, 2453]. marine-entry [1721]. marine-fish [2805]. marinus [2362, 1801, 2993, 1389, 2258, 709, 382, 2347, 1122, 60, 2102]. mark [2226, 1423, 1779]. mark-recapture [1423]. marked [1284]. markedly [2056]. marker [148, 1662, 1186, 2829]. markers [2221, 2156, 1203, 1147, 21, 335, 2868, 1546, 929, 1032, 933, 750, 625, 82, 380, 1946, 959, 622, 2153]. markets [976, 1230]. Marking [1211, 2489, 206, 71, 1768]. marks [2125]. marlin [2483, 2910, 960, 2268]. Marmara [105]. marmorata [2959, 2451, 793, 864, 2590, 2727, 1711, 865]. marmoratus [1749, 2944, 2998, 707, 1866, 1896, 2095, 978, 1888, 1039, 2815, 2495, 3023]. Marquesas [2093, 2051, 973]. marquesi [2969]. marsh [1084]. Marshall [785]. Martin [1150]. masculinization [2665]. Masking [2933]. maskray [2189]. masou [838, 480, 2700, 2732, 1512, 197, 698, 1298, 1862, 480]. Mass [2361, 1768, 2326, 777, 510, 376, 2489, 206, 1148, 978, 1358, 734, 1548, 3081, 2949, 3033, 1082, 786, 125, 1123]. mass-marking [2489]. massive [1428]. masu [2700, 2732, 1512, 197, 698, 1298, 1862, 480]. match [2382]. matchings [1484]. mate [1701, 367, 1034, 2274, 1844, 540]. material [1489, 465, 201]. materials [2725]. Maternal [3010, 1656, 422, 2770, 572, 1914, 1957, 472, 2064, 958, 95, 2218, 1844]. maternity [756, 1919]. mathematical [1129]. Mathis [568]. Mating [1701, 1753, 2388, 249, 2264, 1069, 709, 207, 147, 2155, 1288, 1626, 492, 502, 2676, 2674, 2394, 540]. Mato [2422, 3080, 2437]. matrinxã [1282]. matrix [255]. matrotrophic [758]. matter [1785, 1771, 2080]. matters [1074]. Maturation [96, 448, 1732, 404, 2059, 1357, 1306, 1510, 559, 1363, 2849, 1509, 1501, 2132, 9, 1599, 2333, 276, 916, 94, 2478]. maturation-inducing [9]. mature [355, 51, 189, 197]. maturing [1572]. Maturity [1057, 468, 1845, 1587, 918, 1475, 200, 3011, 1777, 776, 2641, 2171]. Mauritania [1905]. mawsoni [537]. Maximum [820, 831, 697, 2242, 2228, 966, 1823, 2368, 1829, 3075, 2367]. maximus [1710, 2581, 652, 1597, 1680, 3088, 1901, 2354, 2099, 225, 706, 627, 1244]. may [675, 1237, 2242, 1640, 2866, 1692]. Mayak [598]. Maylandia [412].

McConnell [1839]. McCormick [1269]. McDowall [981]. McIntyre [1018]. McKenzie [1211]. McQueen [1426]. mda5 [1964]. meadows [1781]. meagre [2597, 1924]. meagres [1280]. meal [2046]. meals [684]. meaningful [1227]. measured [2340]. Measurement [1823, 410, 587, 1824, 2958, 619, 1259, 2368, 2367]. measurements [3003, 3081, 2311, 1818, 3058]. Measures [1575, 1745, 3043, 2228, 2310]. Measuring [922, 2492, 1771, 1826, 2862, 1825, 1829, 2863, 2854, 1486, 387]. mechanics [1408, 2302]. Mechanism [1295, 2830, 1240]. mechanisms [889, 3053, 1240, 985, 1511, 870, 2940]. mechano [476]. mechano-acoustic [476]. Mechanosensory [2932]. medaka [1178, 247, 680, 1728]. Medes [2712]. media [2486]. median [647]. mediated [2939, 1285, 330, 302, 51]. medirostris [1037, 2753]. Mediterranean [2712, 1209, 2857, 2241, 1952, 2421, 1660, 2494, 1014, 2426, 1407, 2126, 2859, 347, 1884, 676, 2159, 2611, 1909, 1419, 1480, 1714, 328, 2540, 1223, 1064, 2947, 2954, 1447, 45, 2838, 1422]. medium [1223]. medium-term [1223]. meeki [2015, 2346, 666]. meet [1834]. Meeting [1231]. meets [1802]. Megachasma [3021]. megafauna [1217]. Megalobrama [2213, 2774]. megalostomus [2003]. megamouth [3021]. megrim [1206]. Meiobenthos [368]. melanin [1800]. melanochir [370]. Melanogrammus [509, 1752, 534, 2134, 193, 1965, 276]. melanophasma [1042]. melanophore [2466]. melanops [844]. melanopterus [1076, 241, 1058, 2648]. melanostictus [2222]. melanostomus [293, 2720, 1146, 1630, 691, 1151, 1001, 1724, 517, 1604, 513, 2654]. melanotheron [730]. melas [1071, 107, 530]. melastomus [1278]. melding [1027]. meleagris [2682]. Melichthys [2996]. melt [1425]. member [149]. members [2357]. Membrane [3053]. membras [230]. Memory [2827, 2992]. Menderes [2542]. Mendosoma [2427]. menhaden [945]. mentella [1160]. menu [2650]. Mercury [1989, 1396, 1987]. meridionalis [1607]. meristic [470]. Merlangius [534, 1694, 1948, 2675]. merlangus [534, 1694, 1948, 2675]. Merlucciidae [1685]. Merluccius [1766, 879, 1228, 1685, 2857, 1628, 1658, 648, 1637, 1249, 2775, 1451, 952, 1709, 470]. Merry [221]. Merry-Go-Round [221]. Mesa [1211]. Meseguer [109]. mesocosm [1073]. mesocosms [1188, 2275]. mesohabitat [255]. mesopelagic [1765, 1790]. Mesophotic [2030, 2798]. mesopredators [1350]. mesopredatory [2173]. Mesopristes [3013]. mesotrophic [1634]. Mesozoic [112]. messenger [2964]. Meta [2647]. metabarcoding [2312, 2982]. Metabolic [1330, 898, 1817, 1141, 2680, 2663, 2880, 1822, 1827, 3031, 543, 3003, 1041, 2228, 1915, 966, 1838, 734, 2807, 2310, 1831, 2311, 1823, 2368, 2131, 1829, 1277, 2973, 1712, 2367]. metabolism [1325, 2220, 1833, 934, 2477, 1358, 2481, 1830, 1369, 1082, 2848, 46, 84]. metabolism-related [2477, 84]. metabolites [3063]. metaense [1100]. metallicus [3058]. Metals [1265]. metamorphosed [1122]. metamorphosis [59, 2579, 1640, 1931]. Metazoan [651, 1612, 35, 1886]. meter [1615]. methadone [2892]. methanesulphonate [932]. method [2615, 1300, 69, 2089, 2856, 759, 2226, 297, 1378, 151, 2165, 1532, 2228, 1366, 2323, 1495, 585, 1651, 1512, 2192, 2669, 71, 1106, 1425]. methodologies [868, 1008]. methodology [3003, 2769]. Methods [142, 3067, 238, 178, 566, 1817, 1216, 768, 2689, 873, 2827, 1777, 2570, 927, 2350, 1853, 609, 1821, 298, 76]. methyl [1989]. methylation [1291, 2752]. Metriaclima [728]. metric [1486]. metrics [2663]. metropolitan [630]. Meuschenia [2511]. Mexican [3029, 2164, 635, 2104, 1302]. Mexico [2822, 1094, 2506, 2733, 2564, 1158, 1590, 1883, 2101, 466, 2428, 388, 1564, 2423, 2232, 1633, 1932, 2870, 765, 1229, 593, 1576, 3015, 2812, 1868, 937]. MH [2274, 910]. mh-II [910]. MHC [1916, 2788, 150, 947, 924]. MI [2752]. Mice [2650]. Michigan [2130, 819]. Micro [1697, 759, 2023, 2599]. Micro-anatomical [1697]. Micro-computed [759, 2599]. micro-tidal [2023]. Microanatomical [2349]. Microanatomy [1303, 2293]. microarray [329, 212]. microbiota [2482, 1053]. microcephalus [746]. microchemical [2204]. microchemistry [3084, 2828, 2261, 2205, 1342]. microchromosome [292]. microchromosome-carrying [292]. microcosms [417]. microcystin [97]. microenvironment [1023]. Microgobius [2015]. Microhabitat [1657, 28]. microhabitats [2012]. microincrement [1578]. microlepis [2000, 147]. Micromesistius [849, 2954, 1782]. Micronesia [1309]. microocellata [2977]. microphthalmum [2033]. microplastics [2857]. MicroPlex [1131]. Micropogonias [1112, 35, 2112]. microps [2738]. Micropterus [2551, 841, 1555, 934, 12, 2200, 455, 2951, 1253, 1615, 599, 2305, 1375, 1418]. microRNAs [2809, 2964]. Microsatellite [1737, 1301, 369, 657, 954, 608, 411, 1203, 1056, 373, 1147, 1540, 1258, 1703, 2868, 1762, 1131, 148, 1494, 929, 1032, 830, 625, 972, 1492, 2585, 82, 1946, 1650, 622, 1648]. microsatellites [1685, 996]. microscopic [1777]. microscopy [1547]. Microsporidia [813]. microstructure [1623, 226, 120, 1598, 1628, 2678]. microtidal [1518]. microwear [1397, 2460]. micrura [921]. Mid [2779, 2673, 1308]. Mid-Atlantic [2779]. mid-elevation [1308]. mid-north-eastern [2673]. midbrain [409]. middle [1773, 2360]. midshipman [1663, 2849, 1868]. midslope [1761]. midwater [1996]. midwinter [579]. migrate [1497]. migrating [2615, 2628, 1120, 402, 2590, 1959, 1792, 1505]. Migration [442, 464, 488, 507, 527, 551, 570, 518, 106, 2651, 1474, 1136, 861, 1926, 916, 400, 1386, 866, 2326, 1855, 2702, 1245, 2296, 2042, 867, 868, 448, 221, 473, 1563, 1468, 879, 516, 2137, 1149, 2701, 875, 2079, 864, 2336, 1735, 872, 2656, 394, 873, 1387, 2763, 1731, 2933, 2594, 1806, 888, 3057, 346, 2600, 1104, 1789, 1167, 1572, 2547, 236, 862, 1500, 738, 2881, 869, 2716, 870, 1006, 1290, 863, 2110]. Migrations [877, 438, 2090, 1193, 881, 887, 675, 874, 1981, 1045, 1256, 882, 884, 959, 883]. Migratory [700, 2258, 426, 1498, 885, 2882, 1503, 450, 2997, 2943, 2727, 943, 2935, 130, 1192, 2081, 865, 3047, 886, 1288, 2698, 1501, 2583, 1038, 625, 2881, 1342, 1650, 60, 197, 2752, 405]. Migratory-stage [2258]. MIK [1588]. Milagrosa [2422]. mild [435]. miles [2338]. milii [2990]. milk [1776]. millennia [862]. Miller [462]. million [862]. Mimagoniates [2000]. mimetic [897]. mimicking [1181]. mimicry [2093]. Mind [992, 634]. mineralization [1856]. miniature [2282, 2558, 2437, 3080]. miniaturization [2401]. miniatus [1254]. minibarcode [3027]. minimal [992]. minimally [1378, 1897]. minimized [922]. minimum [1954]. minke [382]. minnow [267, 722, 2405, 1068, 1450, 565, 447, 1763]. minnows [533, 1903]. minor [115]. minutus [118]. mirror [2098]. misconception [2956]. Misgurnus [1340, 336, 2477, 2411, 2271, 2363]. Misidentification [2685]. misidentified [2133]. Mismanagement [604]. mismatch [3022, 1560]. missing [732]. Mississippi [1601, 1547, 1868]. Missouri [1903, 2442]. Mitcheson [1099]. Mitigation [2440, 3043, 609]. mitochondria [2613]. Mitochondrial [1437, 1319, 849, 2055, 2225, 1051, 657, 2150, 81, 2364, 1409, 334, 613, 386, 891, 939, 2232, 799, 333, 1714, 20, 363, 952, 622, 2835, 19]. mitotic [858, 1032]. mixed [687, 1147, 684, 1223, 1198]. mixed-stock [687, 1147, 1198]. mixing [1821]. mixtures [530]. mmp2 [2595]. mmp9 [2595]. Mo [2711]. moara [1115, 2501]. mobile [3071]. Mobula [3071, 2968, 2242, 2779, 2676, 1207]. Mobulidae [749]. mode [1724, 502]. model [712, 62, 1557, 1056, 1178, 2665, 1743, 2853, 2941, 3039, 684, 649, 397, 1223, 665, 1341]. Modelling [1943, 1835, 394, 639]. models [2093, 2465, 2524, 1585, 2807, 2259, 325, 2755, 2661, 1162, 2177, 656]. moderate [227]. modern [745, 742, 955]. modes [2653, 857]. modification [1956, 305]. modified [1953, 1651, 2284, 1392]. modifies [1647]. modulates [1713]. modulation [1334]. Moenkhausia [248, 2401, 999]. mohnikei [1920]. mojarra [2011]. mokarran [2385, 1897, 3074]. Mola [2241, 2441, 994, 993, 2240, 689, 1727]. Molecular [1481, 1203, 1457, 561, 22, 1606, 142, 717, 13, 1312, 823, 2083, 1848, 581, 818, 18, 214, 2411, 750, 2006, 16, 1241, 217, 839, 458, 2715, 552, 825, 146, 1931, 2548, 2961, 2241, 3084, 792, 2865, 2420, 783, 1601, 3022, 1880, 2740, 1596, 15, 1651, 2108, 1707, 1890, 2434, 1110, 2504, 2789, 2565, 1860, 1106, 182]. molecule [2961]. molecules [1005, 84]. Molidae [2128, 2803, 2337, 1790]. molitorella [2266, 716]. molitrix [2936]. molluscs [975]. molly [205, 1945, 600]. momentum [647]. monacanthid [2511]. monandric [231]. Monitoring [2916, 138, 2532, 2696, 805, 687, 1366, 2982, 1936, 2467, 1924, 630, 2273]. monoamine [989]. monoaminergic [1794]. monoamines [2661]. monocle [1199]. monoclonal [2276]. Monogenea [407]. monomictic [580]. monophasic [408]. monophyly [1890]. Monopterus [1521, 1837]. Monostichodus [2233]. monotypic [2488]. monsoon [175]. monsoonal [1479]. monstrosa [2421]. Montana [1891]. Monterey [1643]. Monthly [2625]. months [2079]. Moon [1089, 2010, 383]. moonlight [874]. moorei [256]. moray [563, 1755, 532, 2269]. mordax [2269, 284]. Moreton [396]. morhua [1741, 2459, 712, 1399, 1713, 2742, 926, 322, 1056, 1047, 534, 974, 2044, 511, 1795, 2952, 120, 376, 1420, 1476, 187, 2125, 927, 3081, 476, 2550, 585, 1045, 1744, 1184, 2497, 878, 1153, 1349, 189]. Moridae [643]. morio [1526]. Mormyridae [991, 2171]. Morone [150, 346, 1748]. Moronidae [681]. Morpho [2114, 1727]. morpho-groups [1727]. Morpho-histological [2114]. Morphocytochemical [642]. Morphogenesis [1710, 711, 1521, 1756, 2402]. morphogenetic [935]. Morphological [2221, 2622, 1601, 1672, 2513, 864, 1401, 50, 1281, 1352, 2923, 1742, 965, 688, 1071, 1707, 529, 191, 2347, 478, 1110, 477, 1302, 600, 470, 2754, 2241, 1246, 1720, 85, 1481, 1194, 1112, 3048, 3022, 1120, 445, 2559, 2771, 2505, 1641, 1675, 933, 2357, 1004, 2411, 1932, 2170, 1182, 2947, 217, 2504, 2789, 698, 2890, 3004]. morphologically [818]. Morphology [292, 244, 1086, 1503, 1470, 2182, 2720, 2296, 2420, 1164, 2509, 33, 679, 1798, 2060, 2686, 2382, 1130, 1450, 287, 1769, 988, 574, 2833, 1015, 54, 1539, 2942, 2298, 894, 3059, 2268, 1033, 693, 1020, 2094, 597, 2171, 2746, 1921, 1160, 2118, 2871]. Morphometric [1586, 1759, 58, 1672, 1314, 1516, 1143, 1417, 2744, 1515, 1615, 2006, 470]. morphometrics [1880, 281, 1157, 612]. Morphometry [2293, 1899]. morphospace [2503, 3054]. morphospaces [2450]. morphs [2681, 917, 623, 118, 147, 947, 2006, 1624, 950]. mortality [62, 785, 819, 1237, 2208, 2607, 1030, 1216, 56, 872, 559, 3043, 2626, 1382, 1659, 2080, 2265, 2967, 1428, 2110]. mosis [1971]. mosquito [2391]. mosquitofish [937]. mossambica [865]. mossambicus [66, 798]. most [2304, 973, 2669, 1653]. Mother [1396, 2968, 310]. mothers [572]. motility [1713, 11, 3035, 1450, 2566, 845, 2265]. motivation [1181]. mottled [2130, 623, 865]. mound [2725]. mound-nesting [2725]. mountain [2747, 579, 806]. mounting [1712]. Mouth [2865, 2754, 2096, 1531, 147, 1142, 821]. move [904]. Movement [766, 176, 3071, 563, 234, 2119, 1937, 1248, 1134, 2176, 808, 848, 1677, 762, 233, 2695, 1497, 1695, 2390, 1368, 2329, 1063, 764, 3023]. movements [282, 1734, 1101, 1236, 2972, 90, 1962, 2353, 1904, 2168, 2014, 877, 1499]. moving [640]. Mozambique [198, 586, 240, 421]. mRNA [1386, 566, 590, 1581, 1097, 2589, 2714, 2237]. mRNAs [2321]. MS [619]. msDNA [2153]. MSY [2915]. mt [681]. mt-nd [681]. mtDNA [395, 1040, 264, 1459, 914, 2490, 2153]. much [2209]. mucosa [2620]. Mucosal [2872, 2871]. Mucous [1670, 2191, 1262, 2476, 425]. mucous-cell [1262]. mucus [2856, 301, 974, 299, 2975, 1053, 3009]. mud [2266, 852, 716]. mudfish [1779]. mudminnow [1675, 2419, 1201]. mudskipper [1333]. Mugil [521, 1784, 388, 1731, 1474, 3047, 2430, 431]. Mugilidae [521, 1084, 1474, 431]. mullet [1952, 1784, 1731, 1668, 1352, 2574, 2430]. Mulloidichthys [3030]. Mullus [1952, 1352]. Multi [2488, 1938, 1046, 2853, 709, 2678, 2886]. multi-annual [2678]. multi-component [709]. Multi-gene [2488]. Multi-locus [1938]. multi-model [2853]. multi-species [1046, 2886]. multicolor [524]. multicuspid [2400]. multidens [2688]. multifaceted [271]. Multifractal [1939]. multilineatus [468]. Multiple [81, 2229, 709, 223, 93, 1557, 917, 1542, 1934, 601, 2055, 2671, 1919, 333, 213, 3046, 2753, 2040, 2555, 2406, 2674, 1026, 2052]. multiple-spawning [93]. multiplex [1131]. multisensory [564]. Multispecies [1356, 1667, 2915]. multivariate [36, 885]. mummichog [18, 842]. mummichogs [3032]. Muraenesox [2267]. Muraenidae [1755]. Muraenoclinus [2585]. Muraenolepididae [2066]. Muraenolepis [2066]. muricatum [2361]. Murray [686, 538]. Muscle [787, 1156, 2869, 1239, 457, 545, 3007, 1276, 2379, 1797, 115, 1572, 842, 2613, 2283, 1080, 3009, 1400]. muscle-infecting [1080]. muscles [400]. Musculoskeletal [500]. Museum [1839, 2332]. Musick [549, 139]. mussel [2597]. Mustelus [2042, 940, 294, 2561, 1971, 1812, 1887, 2040]. mutant [813]. mutants [1114]. mutton [2032, 163]. Mycteroperca [2831, 1722]. Myctophidae [1765, 1659, 1783, 3054]. Myctophum [1659]. Myeloid [437]. mykiss [1170, 727, 1283, 566, 2543, 309, 584, 1023, 951, 404, 1409, 1772, 903, 23, 127, 106, 299, 104, 833, 1095, 3010, 1911, 2964, 1433, 1847, 2311, 2470, 1436, 1572, 693, 596, 1379, 1020, 2471, 2784, 2695, 2697, 2703, 2085, 2053, 850, 905, 83, 1513, 1712, 2512, 480, 1719]. Myleus [2619]. Myliobatidae [2034]. Myliobatiformes [2959, 772, 456, 1207]. Myliobatis [2034]. Myliobatoidei [2853, 773]. Mylonas [983]. myoglobin [1606]. myops [2051]. myosin [457]. myostatin [1793]. myotomal [3007]. Myrophis [578]. mystery [2905]. Myxine [586]. Myxinidae [257, 243, 586, 2377]. Myxobolus [299].

N [667, 672, 556, 2508, 1396, 166, 2189, 667, 1852, 1668, 1095, 1986, 2147, 3089]. naccarii [2622]. naevus [919]. naked [1445, 2471]. namaycush [819, 1957, 344, 1090, 1395]. name [2004]. names [2247]. Nannostomus [839]. nanodon [591]. nanomaterials [589]. nanonocticolus [2282]. Narcine [2037]. Narcinidae [2037, 2424]. narcosis [2613]. narinari [2439, 2226, 1861, 1067]. Naso [1485]. Nassau [1534]. nasus [851, 3082, 3057]. Natal [2855, 2770, 1107, 2940]. National [245, 1905, 630]. nationally [1117]. Nations [2898]. Native [2760, 657, 2391, 1976, 2089, 1558, 1916, 636, 49, 2186, 2336, 1059, 2935, 1646, 1295, 129, 1411, 2152, 1538, 2797, 2470, 965, 1674, 160, 2563, 1071, 816, 2806, 30, 2867, 637, 434, 1547, 2495]. nattereri [1202]. Natural [1431, 812, 2436, 2535, 1135, 1839, 1098, 210, 2517, 2625, 1429, 1555, 114, 1745, 3086, 1192, 2651, 3052, 2505, 2272, 1034, 75, 1436, 1284, 2965, 1119, 683, 633, 2274, 1537, 2715, 1027]. naturally [1092, 1035, 1966, 989, 223]. nature [2452]. naucrates [2733]. navigation [726]. NCAM [2373]. nd [681]. Near [2038, 3071, 309, 1468, 2494, 2971, 2160, 765]. near-shore [3071]. Near-term [2038]. near-threatened [2494]. Nearctic [1633]. nearshore [1734, 2900, 729, 2374, 1483, 2253]. Nebrius [2908]. nebulosus [340, 2219, 1061, 96, 1446]. née [25, 2313]. need [161, 687, 1995, 1143, 2009]. needed [1845]. Needs [1070, 1231, 2348]. Negaprion [1954, 1808, 253, 767, 1961, 2814]. negative [1200, 3026, 2814]. negative-phototactic [3026]. negatively [1399, 115, 2797]. Negro [1987, 1985, 839]. neighbouring [2525]. nektonic [887]. Nemacheilidae [645, 2414, 2215, 2851, 2645, 733]. Nematocharax [2782]. Nematoda [1351]. nematode [580, 1080]. Nemipteridae [1199]. Nemipterus [1447]. Neoceratodus [1289]. Neochanna [1779]. Neoditrema [1574]. Neoechinorhynchus [1092, 2871]. Neogobius [293, 2720, 1146, 691, 1151, 1001, 1724, 517, 1604, 513, 2654]. neoliberalism [2452]. neon [2300]. neonatal [1544]. neophobia [2078]. Neosebastes [2646]. Neotoca [967]. neotropical [156, 1720, 1981, 1424, 1986, 2969, 1138, 2239, 795, 2364, 1258, 1994, 2136, 1977, 886, 816, 1100, 2479, 1980, 76, 2486]. Neotrygon [2189, 351, 773]. Nepal [100]. Nepalese [645]. Neptune [1809]. nerka [1386, 2668, 2702, 2628, 702, 258, 1149, 1721, 1593, 2345, 1956, 872, 559, 1583, 236, 1256, 1840, 916, 1074, 622]. Nerophis [504]. nervous [1734]. Nest [249, 1091, 2172, 2155, 415, 1025, 416, 1418]. nest-fidelity [2172]. nest-guarding [2155]. Nest-site [249]. nesting [2725, 2076, 288]. nests [2725, 1690]. net [1715]. Netherlands [1225]. nets [1588]. Network [3016, 2593, 2737, 2736, 2784]. neural [2111, 986, 2175]. neuro [2075]. neuro-anatomy [2075]. Neurobehavioural [1412]. Neuroecology [789]. Neuroendocrine [7, 428]. neuroendocrinology [2661]. neuromasts [1482]. neurons [1130]. Neuropeptide [1097, 1440]. neurosciences [2661]. neurotransmitters [1794]. neutral [1916]. neutrophil [447]. Newaj [3017]. Newaj-Fyzul [3017]. newborn [2122]. Newfoundland [1460, 2320, 1476, 2134]. newly [429, 61, 1700, 1077, 1122]. NGOs [1228]. Nhamundá [2514]. NHE2 [22]. Niche [3054, 2183, 2239, 1785, 2190, 817, 2373, 2272, 2120, 313, 1744, 1003, 1525, 646, 454, 627]. niches [534, 2447]. Nielsen [139]. niger [2996, 972, 2171]. night [998, 1963, 3008]. nigra [720, 1619, 71, 501]. nigricans [2910]. nigrofasciata [1539]. nigrofasciatus [2486]. nigrofuscus [2599]. Nile [40, 2880, 855, 2892, 574, 2788, 288]. niloticus [40, 1412, 2880, 855, 2892, 574, 2788, 1205, 1608, 2837, 288]. nine [2350]. Ningaloo [2467]. Niño [2269]. niphonius [2446]. nitric [531]. nitrogen [1325, 770, 2048, 592, 1451]. No [2051, 21, 1041, 2377, 2814, 1844, 1224, 2748]. no-take [1224, 2748]. nobilis [961, 2936]. nocturnal [1953]. nodosa [1781]. noise [293, 876, 1413]. Noisy [2876, 285]. nomenclature [465]. nominally [2027]. Non [587, 1265, 2797, 2351, 439, 3009, 2760, 2615, 1456, 1300, 2391, 1056, 1949, 1339, 2089, 2856, 1558, 636, 49, 2186, 2336, 1305, 1059, 1067, 347, 822, 1646, 1411, 1453, 848, 2470, 965, 1674, 160, 2563, 1071, 1066, 30, 2410, 2204, 434, 2495, 1513, 664]. non-anadromous [1513]. non-destructive [1300, 1949]. non-diadromous [848]. Non-Essential [1265]. non-focal [2410]. non-herbivorous [1453]. non-indigenous [347, 822]. non-invasive [1339, 2856, 1067, 1066]. Non-lethal [587, 3009, 2615, 1305, 2204]. non-model [1056]. Non-native [2797, 2760, 2391, 2089, 1558, 636, 49, 2186, 2336, 1059, 1646, 1411, 2470, 965, 1674, 160, 2563, 1071, 30, 434, 2495]. Non-parallel [2351]. non-parasitic [664]. non-planktonic [1456]. Non-Teleost [439]. nonylphenol [793]. norclerodane [247]. Norin [2367]. Normal [598, 2548]. normalisation [2951]. normalization [2322, 1958]. Norman [2127]. normoxia [596, 2512]. normoxic [1337, 2323]. norms [1074]. North [1550, 605, 2070, 375, 643, 2121, 484, 2122, 2728, 2673, 2717, 1481, 1611, 2526, 2019, 1345, 1811, 748, 1407, 1517, 24, 2853, 2126, 28, 915, 250, 2712, 351, 1978, 2029, 229, 556, 895, 1815, 1045, 1617, 1786, 1297, 95, 2033, 1213, 325, 215, 1154, 1123, 159, 1487, 1622, 1868, 2358, 509, 1271, 964, 807, 770, 2581, 221, 534, 1452, 1475, 395, 3082, 543, 175, 1420, 1851, 2800, 1582, 1467, 150, 1206, 1049, 1913, 328, 307, 193, 824, 2109, 2915, 815, 906, 595, 1690]. north-central [1868]. north-east [643, 484, 2122, 1481, 2526, 2019, 1345, 1811, 748, 2853, 28, 351, 895, 1815, 1045, 1786, 1154, 159, 1487]. north-eastern [2728, 2673, 2126, 915, 250, 2029, 556, 1617, 2033, 1123, 1622, 2358]. north-west [2121, 484, 2717, 1407, 24, 2712, 229, 95, 325, 215]. north-western [375, 1517, 1978, 1297, 1213]. Northern [2941, 130, 949, 2506, 2211, 631, 2385, 1158, 450, 1344, 576, 1590, 2597, 473, 377, 534, 567, 1224, 2777, 2999, 1951, 2348, 2679, 2971, 1206, 2423, 430, 105, 1919, 3072, 239, 798, 1077, 1213, 193, 633, 331, 1198, 1709]. Northernmost [1296, 1622, 484, 385]. norvegicus [1160]. Norway [473, 2435, 1571, 477]. Norwegian [2711, 2441, 2636]. nose [2729]. nosed [2217, 495, 505, 504]. notatus [2182, 1663, 1540, 2849, 1579]. Notes [719, 2037, 2428, 351, 132, 973, 1622]. Nothing [2672, 2416]. Nothobranchius [1121, 2403, 1004, 240, 421]. Notorynchus [306, 2518, 764]. notothenioid [13]. notothenioids [828]. Notropis [267, 2807, 2929, 1083]. Noturus [1601]. nov [1088, 2414, 2215, 2212, 1004, 2504, 2425]. novacula [166, 119]. Novel [2832, 1565, 1338, 694, 954, 2043, 1567, 1956, 1354, 1114, 2165, 1137, 2583, 972, 380, 2888, 149, 79, 2355]. novemaculeata [609, 615]. novo [1056, 2664, 2829]. nubilus [267]. Nuclear [939, 598, 10, 1714, 83]. nucleotide [1056, 335, 2159]. Nui [2974]. null [1594]. number [2532, 1246, 1421, 2494, 791, 947]. Nunavut [638]. nuptial [1398, 2285, 1119]. nurse [2908]. nurseries [2012]. nursery [1704, 767, 2963, 652, 1934, 1142, 1686, 1483, 2014, 2885, 1249, 2775]. nutrient [1670, 495, 1566]. nutrients [2507]. nutrition [1231]. nutritional [1852, 472, 2550, 2218]. NW [2836]. nyererei [1141].

O [908, 671, 1618, 905]. O. [818]. Oarfish [1174, 1158]. Oates [142]. Ob [1459]. obesus [2517, 290, 1455, 1176]. Obituary [1405]. objectives [1557, 158]. objects [3013, 992, 2260]. obligate [994, 993, 689]. obscurus [2040]. Observation [3082, 1630, 813, 1299, 1786, 1487]. Observations [2122, 2687, 3036, 2836, 1423, 2100, 2410, 2855, 1108, 2468, 1704, 1158, 2854, 2729, 754, 2299, 394, 2629, 953, 2802, 117, 1805, 2329, 2387]. observed [1202, 684, 2917, 827]. observer [1892]. observers [2685]. obstacle [1192]. obstructions [309]. obtaining [858, 620]. obtusidens [1650]. occasional [1177]. occulta [1360]. occupancy [1091, 1891]. occupying [1636]. Occurrence [2383, 1348, 1807, 1260, 1699, 280, 1207, 719, 484, 306, 2183, 2517, 1809, 385, 99, 336, 1810, 1923, 1788, 1206, 2976, 2216, 2377, 2109, 434, 1103]. occurring [1580, 2056, 1846, 1035, 512, 895, 88]. Ocean [290, 2874, 1783, 2441, 994, 1806, 993, 2707, 2240, 1790, 325, 1497, 1272, 689, 2984, 3038, 977, 1042, 643, 484, 2122, 2724, 1882, 2450, 2057, 2717, 2517, 2427, 755, 1481, 3050, 2567, 2483, 807, 3029, 378, 814, 1935, 254, 2729, 770, 2581, 567, 1345, 1892, 1649, 1414, 2758, 748, 2260, 1452, 1475, 1517, 395, 1760, 1577, 2825, 761, 114, 644, 274, 2304, 543, 175, 1161, 1851, 1467, 556, 1419, 2971, 617, 2408, 852, 300, 757, 895, 586, 1530, 973, 1815, 1045, 2554, 420, 1786, 1913, 328, 1932, 660, 2745, 1012, 1813, 2099, 2033]. Ocean [2178, 384, 2757, 1154, 2133, 1176, 1123, 159, 906, 1487, 1860, 595, 1622]. Oceanic [1174, 1158, 2900, 2996, 328, 1298, 113]. oceanica [347]. oceanographic [1483]. Oceans [282, 1018, 2536, 1669, 282]. ocellar [1489]. ocellaris [2075, 1041, 2907, 1169]. ocellata [1052, 1057, 2355, 912]. ocellatum [2833]. ocellatus [1557, 2819, 1936, 2065, 1656, 670]. oct4 [1520]. octogenarian [2956]. Ocular [1868]. Ocyurus [593]. odacine [418]. Odax [418]. Odontesthes [1942, 675, 2399, 1194, 304, 320, 1024, 428]. odorant [1263]. odours [513]. oedema [819]. oestradiol [1866, 1647]. oestrogen [1178]. oestrogen-related [1178]. oestrogens [1436]. off [1108, 1704, 2121, 897, 1438, 282, 2910, 3025, 2101, 1732, 1311, 915, 1233, 1740, 2779, 586, 1530, 420, 1617, 2583, 1217, 2462, 3074, 2134, 2177, 298, 1451, 2855]. offs [1221, 1835, 2954, 203, 1945, 1439, 548]. offshore [2464, 1730]. offspring [819, 572, 2064, 310, 369]. ogac [1476]. Ogcocephalidae [246]. Oh [1497]. Ohrid [607]. oil [3082, 1581, 2477, 2704, 1297, 1230, 2407]. oke3 [1594]. Okhotsk [1207, 1296]. Okinawa [308]. old [1766, 1118]. older [1153]. Olfactorily [2939]. Olfactorily-mediated [2939]. Olfactory [2628, 530, 1191, 2943, 2306, 51, 871, 296, 60]. olfactory-mediated [51]. olfax [1722]. oligohaline [98]. oligotrophic [936]. olisiponensis [61]. olivaceus [692, 148, 1032, 1641, 705, 475, 1291, 425, 1717]. olive [148]. olmstedi [1091]. Olson [110]. Oman [1517]. omanensis [1517]. ommaturus [215]. omnivorous [2103, 2996, 721]. omnivory [239]. Onchorhynchus [702, 1504]. Oncorhynchus [1170, 1386, 2668, 2760, 737, 727, 2702, 1283, 2628, 1503, 1528, 566, 2296, 697, 2543, 1764, 2255, 3001, 309, 1429, 6, 584, 258, 1023, 951, 404, 1149, 1834, 2701, 2994, 1409, 2931, 1721, 1593, 101, 335, 2345, 1236, 1772, 1956, 903, 278, 23, 127, 874, 872, 2275, 2700, 2747, 559, 106, 1430, 299, 104, 1583, 838, 833, 894, 2254, 1095, 3010, 1911, 38, 1536, 2756, 2964, 1034, 1433, 1847, 2976, 2704, 259, 1789, 2732, 1512, 2311, 2470, 78, 1436, 1572, 693, 596, 1379, 1020, 2471, 2784, 236, 2578, 2352, 2695, 2697, 325, 1497, 1256, 2703, 2085, 2053, 1840]. Oncorhynchus [870, 2940, 815, 1102, 850, 905, 83, 916, 197, 1074, 1513, 622, 698, 1298, 1862, 1712, 2512, 480, 1594, 2118, 1719]. one [449, 9, 1175, 189, 1106]. one-step [1106]. ongoing [2055]. ongus [1545]. only [2811, 1749, 1763]. onset [2579, 1095, 1334]. Ontogenesis [2445]. Ontogenetic [2731, 679, 3007, 940, 1311, 574, 134, 1523, 1640, 1033, 2028, 1009, 345, 2450, 1138, 2910, 457, 377, 192, 2416, 486, 772, 456, 2802, 2014, 1775, 931, 1451]. ontogenic [1522]. Ontogeny [1416, 469, 661, 2568, 1036, 2955, 1710, 446, 54, 2287, 2200, 817, 958, 2484, 96, 2396, 2644]. oocyte [3076, 1449, 2169]. oocytes [1690]. oogenesis [2115, 2552]. opalescens [556]. open [2708, 98, 1021, 1806, 1152, 2898]. opening [2016]. operated [1158, 2044]. operations [2025, 2767]. opercular [2445]. operculum [803]. Ophichthidae [578]. Ophidiidae [2724, 1747, 420]. Ophidiiformes [2724]. Ophidion [1747, 504]. Ophioblennius [2684]. Ophiodon [1134]. opportunism [2544, 1774]. opportunistic [2113, 2650]. opportunists [1395]. Opportunities [442, 464, 488, 507, 527, 551, 570, 1231, 2531, 861]. Opposite [1635]. opsin [2222]. optic [2373]. Optimal [1013, 1225, 1939]. Optimization [2288, 190]. optimum [697]. options [638, 784]. optokinetic [2937]. Opulent [1669]. Orange [1596, 2847, 1398, 561, 1970, 1312, 207, 64, 3045, 843, 1200, 241, 1277]. Orange-spotted [1596, 2847, 561, 1312, 3045, 843, 1200]. orangefin [241]. orangespine [1485]. orbicular [1189]. orbicularis [1189]. order [745, 1532]. orders [2731]. ordination [145]. Orectolobiformes [1600]. Orectolobus [472]. oregano [2704]. Oregon [1503]. Oreochromis [40, 1412, 2880, 855, 2892, 66, 574, 2788, 2302, 1608, 288, 798, 997, 1026]. oreos [1970]. Oreosomatidae [1970]. Orestias [1055]. organ [1335, 1178, 2560, 589, 2394]. organic [2867]. organism [1056, 1743]. organisms [1553, 2906, 1693, 1819]. organization [1130, 2013]. organs [2772, 1259]. oriental [1340]. orientalis [134, 1009, 618]. orientation [2944, 2942]. origin [43, 624, 160, 1549, 555, 37, 1654, 405]. original [2332]. origins [2055, 14, 490]. Orinoco [2631, 2953, 1983, 1982, 2586]. Orkla [1571]. ornament [1181]. ornamental [2904, 913, 2906, 914, 2369, 2431]. ornatus [363]. orthologue [1241]. orthology [571]. Oryzias [1178, 1017, 247, 680, 1728]. oscillators [800]. Osmerus [284, 2914]. osmolality [2214]. Osmoregulation [1507, 39, 840, 2706]. osmoregulatory [471, 1519, 559, 3059, 1505]. osmosensing [145]. osmotic [145, 71]. osseus [695, 1775]. ossification [836, 1846]. Ostariophysan [461, 2938]. Ostariophysi [2395, 100, 645, 2484, 2401, 1393, 2887, 2050]. Osteichthyes [119, 520, 353, 1987, 1274]. Osteoglossidae [2837]. Osteoglossum [2650]. osteological [2282, 280, 470]. Osteology [2082, 999]. Ostorhinchus [818]. Other [983, 954, 1286, 830, 2970, 972, 2419, 494, 1594]. Othos [1286, 2288]. Otolith [2182, 2955, 1662, 2669, 1342, 3084, 2017, 2121, 1086, 1765, 1864, 1629, 1524, 2720, 1528, 1367, 1623, 855, 701, 2828, 226, 2126, 2261, 700, 120, 2741, 1691, 1598, 2198, 281, 2197, 426, 865, 619, 948, 1136, 1955, 1107, 1756, 720, 1628, 1100, 2678, 1898, 882, 1123, 883, 2196, 1709, 2052, 1372]. otolith-zone [1709]. Otoliths [2889, 797, 1403, 2199, 2576, 1721, 1462, 2771, 2980, 2125, 206, 2029, 2611, 1578, 2201, 1709, 339]. otter [2045]. otter-trawl [2045]. our [2693]. Oust [2668]. out-migration [1789]. Out-of-season [1175]. outcome [2777]. outcrop [2779]. outmigration [1593]. output [2107]. outputs [1230]. Ova [889, 1706, 422]. Ovarian [2850, 2001, 1574, 1690, 238, 727, 1713, 1680, 5, 1581, 502, 1568]. ovaries [189]. ovary [1521]. ovatus [2295, 2605, 1968]. over-winter [1382]. overdominance [2274]. overexploitation [2900]. overexploited [1776]. Overfishing [1070]. overgrown [504]. overlap [1613, 474, 1476, 2120, 240, 646, 824, 627]. overlapping [537]. overlooked [1785, 1260, 2649]. overripening [2850]. Overturf [182]. overview [156, 745, 795, 748, 2008, 2877]. Overwinter [1030, 1914]. overwintering [1471]. oviparous [1623]. oviposition [436]. ovulating [727]. ovulation [2516]. Oxford [3016, 3017]. oxidative [1068, 2665, 3007]. oxide [410, 531]. Oxudercinae [303]. Oxygen [1834, 1818, 1355, 2880, 2214, 1068, 1168, 1364, 1508, 274, 1337, 435, 1826, 1184, 1336, 1898, 1820, 2815, 67, 32, 2209, 1990, 2582]. oxygeneios [710]. Oxynoemacheilus [16, 2851]. Oxynotidae [1785, 2100]. Oxynotus [2135, 1785, 2629, 2100]. oxyrinchus [687, 279, 1654, 2217, 1767, 2119, 1718, 687, 2092, 2545, 1725, 1573, 279, 307, 1654]. Oxytetracycline [2092, 419]. Ozark [267]. Ozarkian [267].

P [1373, 112, 1267, 605, 140, 1265, 690, 1269, 3016, 1099, 956, 588, 859, 1127, 739, 356, 438, 1404, 569]. P. [3050, 1916]. pāroko [2438]. pacemaker [3053]. pachygaster [2808]. Pacifastacus [1610]. Pacific [807, 2729, 770, 114, 644, 175, 1161, 1467, 1913, 1813, 2133, 595, 1042, 2211, 2564, 2450, 1310, 1285, 2517, 282, 755, 697, 2483, 2910, 679, 3029, 1755, 254, 770, 258, 1732, 2151, 2386, 1834, 2830, 1454, 940, 1011, 1105, 2966, 134, 2963, 1233, 264, 1433, 757, 602, 2985, 259, 2384, 973, 1167, 420, 2289, 2745, 2556, 1908, 2230, 1722, 2178, 325, 1009, 1497, 215, 1447, 2166, 1176, 2177, 618, 870, 1180, 3079, 600]. pacifica [2882]. pacificus [770, 1029]. packages [2303]. paddlefish [2203, 1082, 2201, 1690]. paedomorphic [2291]. Pagellus [427, 204, 1274]. Pagrus [715, 47, 334, 2348, 201, 653, 280, 283]. Pain [525]. painted [575]. Pair [1144, 1163, 1484]. pair-wise [1484]. paired [647]. paired-fin [647]. pairs [1169]. Paiva [2327]. paivai [2327]. palaeontology [743]. palatable [1591]. palatal [2560]. Paleozoic [267]. pallasii [2963]. pallid [650, 2442]. pallidus [2565]. pampean [320]. Pampus [2490]. Pamvotis [229]. Panama [2617, 91]. panamensis [91]. Panaqolus [2723]. Panbiogeographical [124]. Pandalus [450]. Pandemic [110]. pandus [2646]. pangasiid [823]. Panjet [65]. Panjet-applied [65]. Panmixia [805, 2446]. Pantanal [2060]. Papers [1212, 550]. papilla [287]. Papua [257]. Pará [1841]. Parablennius [1922]. Paracheirodon [2401, 1987]. paracrine [87]. Paradigm [640]. paradise [888]. paradox [323]. paradoxus [1685, 1451]. Paraguacu [1315]. Paraguai [2422]. Paraguay [3080]. Paraíba [250]. Paralepididae [1765]. Paralichthys [195, 898, 692, 1101, 148, 1032, 1641, 705, 475, 1291, 425, 1717]. Paraliparis [1391]. parallel [947, 2351, 552]. parallelus [1937, 642]. Parameterization [656, 712]. parameters [2845, 1935, 1732, 2045, 996, 1495, 1908, 1380]. Paraná [2495, 1720, 2563, 2223, 2743]. Paranthias [1014]. Parapercis [2655]. paraphyly [2838]. Parasite [2712, 971, 1741, 1661, 675, 1398, 330, 205, 1030, 1681, 651, 370, 1612, 104, 560, 1382, 131, 1885, 199]. parasite-induced [104, 1382]. parasite-mediated [330]. Parasites [129, 568, 848, 1127, 947, 1193, 2140, 560, 1771, 1028, 1536, 35, 430, 1674, 1886, 946, 2354, 1119]. Parasitic [2503, 1399, 1080, 664]. parasitism [2719, 1117, 415, 2871]. parasitological [2241]. parent [2505]. parentage [996, 1651]. Parental [702, 177, 249, 12, 828, 1034, 2674]. pargo [2348]. parietal [1010]. Park [245, 1905, 2467]. parmifera [1757]. Parodontidae [2006]. parr [2532, 2994, 2677, 355, 2087, 51, 1107, 2693, 477, 2694, 1692, 197]. Parrish [1266]. parrotfish [2388, 2771, 2361, 2866, 1309, 2333]. parrotfishes [1845]. Parsons [1211]. part [270, 637]. Parthenogenesis [1455, 2229, 626, 1854]. parthenogenetic [1861]. Partial [867, 868, 2090, 334, 1493]. partially [3010]. particular [1286, 1348, 969]. partitioning [2874, 427, 2190, 2272, 2666, 1132, 163, 88, 2885, 169, 1180, 3054]. Partnership [1226]. partnerships [1652]. parturition [761]. Parupeneus [2243]. parvalbumin [1710]. parvipinnis [2375, 989, 1885]. pass [1316, 2530, 3049]. Passage [3044, 2598, 2224, 1059, 2600]. Passing [1966, 3052]. Passive [2273, 766, 2696, 2597, 1635, 1924, 2848, 2654]. passive-integrated [2696]. Past [2921, 2918, 875, 2537, 984, 1462]. Patagonia [1194, 2035, 1995, 2971]. Patagonian [1516, 1515, 887, 779, 1239, 2427, 2356, 2091, 975, 1022, 1314, 2932, 1249, 1814, 2478]. patagonicus [1314, 1516, 1204, 1515]. Patagonotothen [1239]. patch [2179]. patches [501]. paternal [59, 495, 191]. Paternity [1067, 1542, 2561, 3046, 2040, 223, 1425]. Pathobiology [1127]. pathogenic [539]. Pathogens [568, 2961]. pathology [133]. pathway [1059, 2486]. pathways [1519, 1191]. Patos [1993, 1992]. pattern [400, 227, 1443, 1286, 913, 2058, 1261, 1848, 628, 2616, 2595, 1714, 1628, 502, 764, 539]. patterning [1125]. Patterns [2810, 2710, 765, 1208, 1140, 766, 3071, 1855, 249, 563, 1629, 1479, 354, 1809, 93, 234, 1752, 1076, 2625, 1356, 2011, 1112, 176, 1937, 2093, 767, 226, 2873, 700, 2656, 281, 130, 1021, 1595, 960, 1531, 560, 614, 1753, 966, 1024, 1981, 1001, 886, 3056, 1498, 2463, 344, 86, 1677, 688, 1078, 1217, 2188, 2678, 2745, 995, 1979, 738, 492, 2552, 2618, 2555, 2406, 2178, 1497, 778, 2021, 2911, 1522, 1695, 1050, 2305, 2466, 2858, 1272, 1007, 1116, 2344, 84, 916]. Patzner [140]. Paulo [1012]. paulus [277]. Pavlidis [983]. Pavlov [859, 859, 438]. PCR [1300, 1871, 2964, 2951, 2322, 1910, 20]. PCR-ready [1300]. peacock [2152, 2049]. peaking [1832]. pearly [166, 119]. pectinata [1807]. pectinatus [2389]. pectinirostris [1333]. pectoral [587, 2686, 1797, 2205, 2749]. peelii [686, 538]. pejerrey [2399, 320, 1024, 428]. pejus [1832]. Pelagic [1189, 364, 1587, 2594, 342, 1707, 1213, 696, 1903, 1272, 434]. pelagic-spawning [1903]. pelagicus [1587]. pelagios [3021]. pelamis [2791, 2141, 799]. pellets [2046]. Pellona [2762]. pellucida [2349]. Pelvic [272, 1308]. pencilfish [839]. penetrating [1520]. penetration [2372]. Peninsula [1704, 2066, 965, 30]. Peninsular [719]. Pentacerotidae [2056]. peptide [1520]. Perca [2469, 2235, 80, 2301, 1456, 2214, 2596, 202, 1916, 1914, 636, 1634, 2930, 451, 50, 1532, 1143, 57, 558, 135, 89, 536, 1800, 1624, 1799, 67, 656, 126]. perceive [1153]. percellens [3050, 1012]. perception [994, 993, 2923, 689]. perch [2469, 80, 2301, 1456, 2664, 2214, 2596, 202, 1916, 1914, 636, 1567, 2297, 1634, 2930, 451, 50, 1143, 2168, 1546, 57, 558, 135, 89, 536, 1800, 1624, 1799, 615, 122, 1828, 67, 2716, 656, 2444, 126]. percichthyid [454]. percid [2272]. Percidae [1787, 2285, 1250, 1361]. perciform [837]. Perciformes [2396, 809, 2011, 1460, 2364, 1014, 1258, 2389, 2054, 334, 1161, 483, 1884, 1115, 1918, 2327, 852, 2907, 816, 163, 3061, 2787, 972, 2757, 1343, 431, 2838, 2138]. percnurus [1450]. percobromus [2807]. Percomorphacea [2212]. percula [241]. perfect [2382]. performance [2487, 2303, 1798, 2308, 3031, 1468, 610, 2224, 1555, 343, 1510, 1508, 2087, 2627, 3086, 1061, 1434, 2492, 553, 1416, 1836, 1837, 2477, 1655, 1034, 2174, 2310, 1831, 3002, 1853, 2929, 661, 333, 515, 2271, 1256, 1850, 2613, 1080, 1607, 1621, 405]. performing [1966]. peri [936]. peri-alpine [936]. period [166, 2729, 2892, 1732, 800, 2484, 423, 811]. periodic [1326]. periodicities [1016]. Periodicity [3029, 1079, 1029, 283]. periods [337, 863]. perioperative [2892]. Periophthalmus [2294]. peripheral [3072]. periprandial [2714]. Perissodus [147]. perlo [2859]. permanently [402]. permeable [268]. Perriere [568]. perryi [295]. Persian [471, 2510]. persicus [471, 2510]. persistence [1248]. personality [2041, 1179]. perspective [1832, 2711, 2906, 327, 2636, 1924, 2540, 742, 3016, 588, 356]. Perspectives [1804, 1708, 1272, 76, 2912, 1027]. Peru [1311, 2608, 2549, 591]. Peruvian [2608, 2803, 1578]. petenense [92]. Peter [2280, 424]. Pethia [1951]. Petroleuciscus [2542]. Petromyzon [1801, 2993, 1389, 2258, 709, 382, 2347, 1122, 60, 2102]. petromyzontid [1401, 664]. Petromyzontidae [964]. Petromyzontiformes [1665, 964, 1069, 1492]. PGE [296]. pH [1746, 1631, 1543, 2053, 1990]. Phallichthys [1165]. phase [1385, 2622, 335, 1089, 1955, 1078, 383, 554, 2582]. phases [2010]. Phelps [1211]. phenology [2818, 2326]. phenomenon [1260]. phenotype [2931, 838, 187, 1831, 1856]. phenotypes [1503, 102, 1202, 2752]. Phenotypic [677, 1849, 951, 350, 699, 1627, 1281, 607, 1560]. phenylhydrazine [856]. pheromone [709]. Philippines [2735, 2798, 669]. philopatry [2873, 796]. Pholidae [2082]. pholis [2172, 2082, 1367, 2485]. Phosichthyidae [1765]. phosphorus [1856]. phosphorus-deficient [1856]. Photo [1869, 497, 2601, 2226, 3037]. Photo-identification [1869, 497, 2601, 2226, 3037]. photogenic [2802]. photogrammetry [1693, 387]. Photographic [1426, 2439, 763, 1922]. photographs [2900]. photography [1248]. photoneuroendocrine [3]. Photoperiod [2694, 1499, 2321, 1238, 252]. Photoperiodic [1461, 1049]. photoperiods [1470]. photoreceptor [3004]. Photoreceptors [2930, 2931]. phototactic [3026]. Phoxinus [722, 2405]. Phreatichthys [2287]. Phreatobius [2050]. phthalate [2481]. Phyllopteryx [497]. Phylogenetic [61, 791, 935, 1055, 411, 1304, 1606, 1452, 823, 939, 676, 581, 2611, 3056, 591, 1860]. phylogenetics [2865]. Phylogeny [15, 183, 964, 2488, 2684, 818, 743, 2082, 2746, 552, 681, 19, 112]. Phylogeographic [14, 724, 2713, 16, 1055, 1027]. Phylogeographical [213]. Phylogeography [678, 332, 2739, 1437, 1319, 81, 2225, 363]. Physical [1477, 111, 1433, 2028]. Physiological [344, 1850, 870, 2562, 662, 2308, 2892, 59, 1684, 864, 435, 579, 2923, 688, 2053]. physiologist [25]. Physiology [1, 1126, 549, 1269, 1279, 2609, 2705, 1570, 1817, 2220, 6, 1834, 2072, 2375, 12, 589, 1187, 984, 1866, 1329, 562, 1881, 987, 1770, 739]. phytophagous [2002]. phytoplankton [1727]. piabanha [1672]. Piaractus [2953, 2871]. picarel [386]. pickers [1408]. picote [411]. picta [351, 2586]. pictum [575]. picturatus [2879]. piecewise [1930]. Piedmont [1980]. piers [2852]. Pietsch [113, 389]. pig [2726]. pigment [2116, 1114]. pigmentation [1279, 2220, 546, 552]. Pike [1743, 2382, 2777, 2930, 546, 2267, 1461, 2761, 2360, 811, 89, 1774, 1960, 738, 181, 2334, 3009]. pike-characids [2360]. pikeperch [2482, 2930, 1260, 401, 654]. Pilastro [690]. pilchardus [1899, 1660, 592, 1190, 1369]. pilot [879, 3086]. Pimelodella [2749]. Pimelodidae [2834, 1666, 2620]. Pimelodus [2834]. Pimephales [533, 1068, 447]. pineal [1470]. pinfish [2398]. Pink [2976, 2637, 737, 6, 278, 2711, 2636, 201, 2118]. pintado [2834]. Pioneering [25]. pipefish [495, 505, 2971, 1919, 1940, 1051, 501, 504]. pipefishes [489, 3022, 503, 491, 1145, 720, 191, 502, 499, 490]. Pires [222, 1393, 2395]. Pisces [1516, 1515, 2450, 1437, 1716, 63, 55, 2335, 1314, 386, 1115, 328, 1683, 1165, 1160]. piscivore [2235]. piscivores [1698]. piscivorous [2640, 453, 2288, 684, 2446, 654, 2334]. piscivory [817, 381]. pit [2901, 69, 853, 1317, 1458]. PIT-tagged [1458]. PIT-tagging [1317]. Pitcairn [2386]. pittieri [2401]. Pituitary [2399, 1313]. pivotal [2961]. place [1287, 2124]. placement [2488, 912]. placentatrophic [310]. places [1549, 1790, 1497]. Plagioscion [2364, 816]. plagiosum [1854]. Plagiotremus [2190]. plaice [338]. plain [2350, 2223]. plainfin [1663, 2849]. planeri [1401, 213, 1492]. planet [2341]. planktivore [3071]. planktivores [1362]. planktivorous [381, 135]. plankton [1442, 138, 1873]. Planktonic [1456]. planning [1027]. plant [598, 1933]. plant-based [1933]. Plasma [2214, 1514, 241, 1524, 1944, 2116, 692, 343, 1238, 1512, 3062]. Plastic [2984, 2010]. Plasticity [2087, 2497, 1335, 951, 986, 1627, 1959, 1889, 1281, 1453, 1548, 985]. Plata [2008, 2113]. Platax [1189]. Plateau [267, 2478]. platei [1246]. plates [210]. platessa [338, 655]. platform [3082, 1131]. platforms [2407]. Platichthys [1344, 2642, 931, 1390]. platorynchus [2205, 650]. Platycephalidae [1636, 2646]. platynemum [2972]. platyodus [2400]. platypterus [1307]. platyrhynchos [2620]. Platytroctidae [2802]. plays [2671, 1244]. pleco [2185, 2723]. Plectorhinchus [3038]. plectrodon [1868].

Plectroglyphidodon [2787]. Plectropomus [1353, 1305, 1087]. Pleistocene [1459]. Pleistophora [813]. Plerocercoids [2405]. Pleuragramma [1003]. pleuriticus [815]. Pleuronectes [338, 655]. Pleuronectidae [1247]. Pleuronectiformes [1467, 2826, 1206, 1247]. Pliocene [24, 1938]. ploidy [913, 337, 1455, 369]. plotosid [2191]. Plotosus [1130]. plumbeus [2372, 1932, 2077]. plumierii [2304]. plummeting [1832]. plzfa [1728]. Poecilia [205, 922, 1179, 207, 292, 594, 2595, 1635, 1683, 646, 190, 1945, 1780, 2767, 600]. Poeciliid [690, 3058]. Poeciliidae [1716, 937, 1252, 1683, 1165]. poeciliids [2391, 2883]. Poeciliopsis [724]. poecilopus [1618]. poecilura [1303]. Pogonias [911, 2667]. Pogonophryne [828]. points [1664]. poisonous [1067]. Poland [107]. polar [299, 2257]. policies [2902]. policy [2162, 161, 1496, 1802, 2898]. political [1227]. Poll [1343]. Pollachius [2169, 1349]. pollack [1349]. pollock [2963, 2177, 1463]. pollutant [885]. pollution [810, 2984]. pollux [2835]. poly [1597]. polyandry [1767, 2254, 3046, 2585]. polycyclic [44, 1838]. Polydactylus [848]. polygamous [502]. polymorphic [1703, 1762, 1202, 1494, 830, 625, 1650]. polymorphism [1056, 910, 330, 335, 613, 1085, 1031]. polymorphisms [2788]. polymorphs [2889, 2201]. Polynesia [973]. Polyodon [2203, 1082, 1690]. polyommata [1887]. polyphenic [1015]. polyploidy [2167]. Polyprion [710]. Polypterus [2937]. polyunsaturated [2381, 1683]. polyuranodon [532]. Pomacanthidae [2757]. Pomacanthus [2757]. Pomacentridae [300, 2970, 2907, 2787]. Pomacentrus [2300]. Pomatomidae [328]. Pomatomus [328]. Pomatoschistus [2426, 118, 2095]. pomfret [2490]. pompano [2295, 2605, 1968]. pompanos [2488]. pond [1118, 2477, 41]. ponds [1456]. Pontecorvo [110]. ponyfish [2230]. pool [396, 1478]. pools [28, 2438]. poor [1231, 1605]. poorly [2587, 3049, 2003, 2801, 2730, 2586]. poorly-explored [2587]. poorly-known [2801]. Pop [2181, 282, 234, 2583, 2462]. pop-off [282, 2583, 2462]. Pop-up [2181, 234]. popular [2004]. Population [1749, 80, 1076, 2074, 701, 1602, 1883, 2129, 1196, 1577, 920, 269, 1962, 148, 2446, 1909, 2449, 799, 1058, 2642, 379, 211, 2730, 215, 2305, 952, 1654, 1849, 3084, 2811, 2156, 2150, 1403, 306, 782, 1713, 2149, 2591, 2296, 1146, 576, 1737, 667, 3025, 2625, 1935, 1339, 1758, 1429, 2119, 1643, 2151, 2494, 583, 1149, 350, 1916, 2966, 2868, 1810, 1306, 851, 3083, 2130, 1401, 370, 1179, 1052, 2300, 1143, 374, 2505, 1363, 2055, 229, 3010, 1662, 264, 3047, 497, 2290, 1242, 848, 2879, 716, 1446, 78, 536, 2713, 2148, 1100, 3061, 2989, 1888, 1763, 538, 365]. population [1201, 2780, 1256, 1119, 2442, 1447, 1599, 2490, 331, 670, 2716, 2284, 223, 1027, 496, 2835, 664, 2153]. populations [1386, 657, 2532, 1705, 2702, 582, 2182, 1899, 1737, 1869, 2308, 2662, 678, 258, 2137, 2465, 1427, 1224, 713, 1432, 2154, 2868, 2537, 56, 2559, 2972, 334, 1667, 2700, 1644, 838, 2800, 320, 2159, 558, 1048, 948, 1459, 2152, 14, 2616, 2470, 1071, 1049, 2538, 75, 89, 3072, 816, 371, 2533, 1940, 1988, 630, 2479, 2989, 2351, 273, 1077, 935, 2178, 1019, 2522, 2669, 1775, 730, 627, 2981, 1779, 1074, 1594, 600]. porbeagle [3082]. porcupine [2370]. pores [1377]. porgy [2348, 280]. Porichthys [2182, 1663, 2849, 1868]. Poromitra [1530]. porosus [1124]. Port [1490, 2041, 2962]. portable [2783, 69, 1253]. portend [3028]. portending [11]. Portugal [1758, 159]. Portuguese [701, 2335, 541, 213]. portusjacksoni [1490, 2041, 2962, 298]. Posidonia [347]. position [3007, 2331, 959]. positive [1298]. positively [1925, 1536]. Possible [2873, 1704, 245, 988, 130, 2409]. Post [267, 1643, 475, 41, 2832, 2708, 2770, 372, 1679, 2208, 751, 2079, 1721, 3063, 2219, 2629, 2055, 1184, 554, 2084, 16, 2903, 869, 46, 1006, 2973, 2582]. post-capture [2629]. post-copulatory [751]. post-depositional [2084]. post-exercise [46, 2582]. post-feeding [372]. Post-glacial [267, 1643, 2055, 16]. post-hatch [2832]. post-hypoxic [1184]. post-larval [3063]. post-natal [2770]. Post-release [475, 41, 2208, 2219, 2973]. post-smolt [869, 1006]. post-smolts [2708, 2079, 1721, 554, 2903]. post-zygotic [1679]. postglacial [2130, 14]. postlarvae [585]. postspawning [337, 1423]. postures [1545]. pot [2551]. Potamodromous [1981, 2935]. potamodromy [2988]. Potamotrygon [2624, 2969]. Potamotrygonidae [2969]. Potential [876, 2942, 712, 1415, 2786, 2203, 2526, 253, 655, 1345, 2061, 751, 3051, 170, 1187, 1200, 2290, 408, 2062, 2015, 95, 2178, 215, 2107, 2177, 784, 3028]. potentially [1057]. potentials [737]. Potosí [2822]. pots [2459]. pouch [2623, 191]. pouliulii [2212]. poutassou [849, 2954, 1782]. power [875, 1685]. Poyato [461]. Poyato-Ariza [461]. pp [3016, 3017]. practical [2569, 2691, 908]. practice [958]. Pragmatic [1832]. Prakash [980]. prasinus [563]. prawn [1054]. Pre [611, 2918, 6, 988, 2081, 1160]. pre-adult [6]. pre-extrusion [1160]. pre-smolt [988]. pre-spawning [2081]. Pre-winter [611]. precarious [3083]. Precision [1421, 2816]. precocious [11]. predating [1299]. Predation [1610, 1013, 68, 2469, 2123, 253, 1237, 565, 302, 1684, 394, 1583, 3085, 2554, 1362, 2337, 275, 2967, 1153, 2715, 677, 1392]. Predator [2078, 68, 1882, 2043, 3012, 2126, 2646, 1584, 1221, 1923, 854, 1389, 50, 1057, 1915, 476, 2801, 665, 1535, 944, 2715, 1090, 1709, 1439]. predators [403, 887, 1636, 732, 3085, 661, 689, 708, 1395, 994, 993, 568]. predatory [450, 2179, 1486, 2652, 1182, 1805, 2485]. predict [2528, 1923, 559, 684, 1715, 1831, 2442]. Predictability [1469, 1667]. Predicted [321, 2928]. Predicting [2755, 316]. Prediction [245, 1120, 322]. predictions [889, 2902]. predictive [835]. predictor [712]. predicts [2786, 1532, 1915, 2600]. prefer [2760, 412]. Preference [1746, 2264, 2596, 1631, 896, 2095, 1250, 847, 2962, 117, 122, 3032, 1945, 540]. preferences [1556, 1112, 707, 1523, 1842, 3039, 1037, 1850]. Preferential [2725, 1490]. preferred [207]. pregnancy [2783]. pregnant [1806, 1919]. Preliminary [533, 308, 1590, 1718, 775]. Preline [2708]. premating [366]. premature [872]. premaxillary [1787]. prenanti [1448, 1440]. preoptic [1017]. preovulatory [683]. preparedness [1850]. presbyter [1491]. Presence [2427, 1646, 1665, 367, 2165, 1204, 2852, 693]. present [1271, 1462, 2921]. presented [596]. preservation [975]. preserved [1300, 1477]. preserving [465]. presettlement [2216]. Prespa [607]. pressure [1013, 855, 204, 2085]. pressures [631]. preta [2073]. prevalence [1351]. prevent [433]. prevention [2704]. previous [68, 2056]. previously [2242, 776]. previtellogenic [1690]. Prey [1047, 275, 1676, 181, 1775, 686, 712, 564, 655, 66, 565, 2370, 1795, 2126, 474, 1584, 546, 451, 684, 50, 1057, 3085, 2120, 57, 1698, 811, 1913, 1658, 2667, 1566, 193, 2485, 654, 1154, 2334, 1180, 2715, 857, 126, 2721]. preying [450, 2471]. preys [1589]. Priacanthidae [1760]. Price [3016, 3017]. pricklefishes [1530]. prickly [2135]. Primary [1918, 3076, 1002, 2219, 1856]. primer [683]. primers [1301]. Priming [104]. primitive [745, 1205, 3059]. principles [1129]. Pringle [356]. Prionace [2122, 1701, 830, 1473, 786]. prior [249, 1563, 1152]. priori [699]. Pristidae [2038]. Pristigasteridae [2762]. pristine [863]. pristiophorid [2293]. Pristiophoridae [2293]. Pristiophorus [2293, 3060]. Pristipomoides [2688]. Pristis [1807, 2820, 1686, 2038]. proactive [2846]. probability [2699]. probably [800]. probes [854]. problem [862]. problems [639]. procedure [2247]. procedures [2219, 2436]. Proceedings [139]. process [1177, 1245, 922, 1225, 628]. Processes [956, 2722, 322, 2687, 2227, 1055]. Prochilodus [1092, 885, 2113]. produce [2097]. produced [728, 1280, 1966, 1349]. production [1739, 647, 1738, 713, 343, 1571, 2191, 277, 1668, 911, 2642, 1924, 531, 2831, 1175, 502, 1526, 2792, 2118]. productivity [322, 1671, 2106]. productus [2218]. professionals [1708, 1098]. profile [1942, 1178, 97, 2557]. profiles [2255, 1420, 1136, 2535, 1572, 2307, 1026, 212, 826]. profiling [1606, 2328]. Profundulidae [103]. Profundulus [103]. progenitor [2175]. progeny [702]. progestogens [531]. programme [2042, 2819, 1243]. programmes [1067]. progress [329, 750]. progressive [2325, 596]. project [2009]. projects [1226]. Prolactin [2813]. Proliferation [1275, 2235, 1710]. Prolongation [2267]. Prolonged [402, 1416]. promelas [533, 1068, 447]. Promise [329]. Promised [517]. promising [2769]. promote [1227, 2537]. promoted [350, 2113]. promoters [1291]. promoting [1244]. propeller [2662, 3044]. Proper [2434]. properuptus [818]. propofol [932]. proportions [1276, 2983]. Proposed [510, 2796]. propulsive [647]. Proscylliidae [907]. prospects [1317, 750]. protandrous [2328]. proteasome [115]. protect [104]. Protected [3016, 1611, 2712, 1802, 2690]. protection [1930, 2748, 1127]. protein [2487, 2943, 1738, 561, 1514, 1312, 225, 1520, 539, 1244]. proteins [2510]. proteolytic [1944]. Proteomic [1608, 480]. Proterorhinus [434]. Protocol [1292, 1949, 522, 802]. Protocols [142, 2570]. protogynous [2847, 231, 1521, 1135, 1576]. Protomyctophum [1783]. Protopterus [70]. Prototroctes [365]. protozoans [2704]. Protracted [2398]. Provenance [1074]. provide [728, 2828, 1565]. provides [951, 368]. providing [1041]. Province [598, 2645, 3061]. provisioning [495]. Provo [226]. Proximate [2615]. proximity [2616, 810]. przewalskii [332]. PSA [2373]. PSA-NCAM [2373]. Psammobatis [2371]. PSATs [2583]. PSCF [443]. Pseudacanthicus [2185]. pseudoalbino [1245]. Pseudobatos [3050, 2218]. Pseudocarcharias [754, 1996, 121]. Pseudocrenilabrus [524]. pseudogene [698, 1298, 1862]. pseudogene-positive [1298]. Pseudophycas [564]. Pseudoplatystoma [1203, 1100, 2001, 2495]. Pseudopleuronectes [59, 1598, 2325, 847, 1449, 1575, 1247]. Pseudosciaena [1394]. Pseudoscopelus [3090]. Psychrolutidae [1481, 2257]. Pterapogon [812]. Pterois [2338, 1869, 2530]. Pterophyllum [2088]. Pterygoplichthys [2563]. public [1227, 302]. Publication [1360]. Published [3016, 3017]. pufferfish [647, 2808, 2894, 581]. pughead [2339]. pulcher [1870]. pullus [418]. pulse [991]. pulse-type [991]. pulses [2640]. pum1 [680]. pum2 [680]. pumpkinseed [636, 3012, 1028, 965, 319]. punctata [2643, 3025]. punctatum [1645, 2833]. punctatus [1606, 483, 2515, 2310, 1615, 2443, 2504, 578]. Pundamilia [1141]. Punta [1894]. puntazzo [419]. Puntius [836, 443, 813]. pupfish [1753]. pupping [2790, 1686]. pure [2494]. purpose [2533]. purse [1190]. purse-seine [1190]. Purus [1988]. putative [1880, 2815, 2102]. putitora [2114]. pygmy [907, 2968, 277]. pyloric [621]. pyrolytic [1838]. pyrosequencing [1650].

Q [1211]. Qatar [776]. Qeshm [912]. Qinghai [332]. qPCR [1442, 1958, 1425]. quadruncus [937]. quality [926, 2398, 1509, 180, 2015, 2667, 1154, 27, 171, 1439, 138]. Quantification [238, 1956, 472, 1442, 515]. quantify [953, 2334]. Quantifying [2103, 1698, 1473, 159, 361, 445, 2192]. Quantitative [1350, 662, 873, 911, 2319, 1442, 2964, 2951, 2322, 1292, 2540, 191]. quantity [1509, 1439]. Queensland [1164, 2045, 3060, 363]. quelen [932]. questions [161, 2538]. quick [585]. quitzeoensis [411].

R [208, 108, 1126, 140, 141, 549, 1268, 1070, 1173, 1211, 1174, 139, 311, 438, 981, 982, 461, 2145, 970]. R. [57, 2885]. rabbitfish [1144, 2421]. races [523]. raceway [2708]. Rachycentron [478]. radcliffei [907]. radiation [598, 303, 1619]. Radio [1603, 602]. radio-tagged [602]. Radio-telemetry [1603]. Radiocaesium [1400]. radiocarbon [297]. Rafinesque [447]. Rafting [1204]. Rainbow [309, 2078, 1170, 727, 2543, 584, 1023, 951, 404, 1409, 1772, 903, 23, 127, 299, 104, 833, 1095, 1911, 2964, 1433, 1847, 2516, 2311, 2470, 1436, 693, 596, 1020, 2471, 2784, 284, 2703, 2053, 850, 905, 83, 1513, 2512, 480, 1719]. raised [256, 1284]. raises [2685]. Raja [748, 940, 2977, 900, 1671, 547]. rajid [1096]. Rajidae [755, 2729, 1907, 2607, 2048, 778]. rajids [1111]. Rajiformes [2729, 211, 2380]. Raju [2127]. rakers [1547]. ramada [521]. Ramnogaster [520]. ramp [1316]. ramsayi [1239]. ranched [1192, 553]. Randall [1447, 1860]. randalli [1447]. Randallichthys [1860]. random [147, 2176, 1050]. random-at-age [2176]. random-at-length [2176]. range [1920, 2728, 563, 2301, 1892, 2386, 1295, 401, 2327, 2423, 1530, 2970, 3070, 2713, 1678, 637, 16, 2755, 834]. range-expanding [2423]. ranging [2795, 1693]. rank [1017]. rank-dependent [1017]. ransonnetii [1574]. Ranzania [1257, 2128, 2337]. Rapa [2974]. rapacious [381]. Rapfish [1220]. Rapid [2235, 1056, 66, 1940, 723, 2981, 1392, 697, 632, 1354, 1053, 2982, 1744, 1220, 298, 1106, 1425]. rapidly [2516, 552]. rapids [1666]. rare [1849, 1177, 1117, 2602, 2100, 229, 1206, 1891, 291]. rarely [2299]. rarest [1874]. rasile [1193]. rasptail [940]. rate [1397, 2469, 697, 1822, 1817, 3031, 802, 404, 1588, 1293, 2977, 2046, 2228, 115, 1915, 57, 734, 2807, 3053, 2311, 1818, 1823, 2368, 1893, 1919, 2580, 2131, 1986, 3075, 2973, 1712, 2367]. rates [2523, 1283, 2663, 205, 47, 922, 2048, 543, 3003, 1925, 2680, 2430, 1825, 2289, 1829, 2084, 2866, 2886, 1885, 1197]. ratfish [782]. rather [2786, 2526, 2131]. ratio [587, 1778, 2136, 2804, 1641, 594, 1185, 2406, 2147]. ration [2869, 1469, 2289]. ratios [2508, 777, 2189, 770, 1852, 2379, 700, 2741, 426, 865, 786, 1451]. ray [3071, 2959, 748, 3090, 793, 2226, 1861, 1303, 2819, 919, 198, 1699, 2779, 1853, 2480, 2205, 2037, 900, 1671, 2424, 2171, 1207, 921, 1089]. ray-finned [3090]. Rays [2244, 783, 2242, 2370, 2439, 2513, 1067, 763, 2657, 973, 2885, 2676, 2204, 1373]. razorfish [166, 119]. rDNA [77]. Re [2233, 963, 1954, 1745, 3088]. Re-establishment [2233]. re-evaluate [1954]. re-feeding [963]. re-introduction [1745]. re-sighting [3088]. Reach [1103]. reaches [2741]. reaction [1074]. reactive [2846, 1337]. reactivity [1715]. readings [1864]. ready [1300]. Real [1871, 1910, 2964, 2951, 2322, 2147]. Real-time [1871, 1910, 2964, 2951, 2322]. really [532, 2524, 2209]. reared [400, 2523, 2708, 2238, 2786, 354, 1188, 802, 1795, 278, 2219, 1510, 1603, 3086, 1061, 1966, 1561, 232, 2159, 3026, 1498, 476, 1789, 1369, 1016, 1376, 1509, 1501, 661, 2568, 1472, 2967, 1560, 475, 706, 41, 2476]. Rearing [3001, 1429, 1506, 2609, 533, 1555, 1508, 1895, 1034, 2557, 280, 2695, 612, 171, 2973, 1243]. rearrangements [77]. Reassembling [1403]. Reassessment [760, 2713]. recapture [2042, 1339, 2226, 196, 1423, 1341, 1779]. recaptured [1629]. recaptures [2967]. recently [1602, 2839, 2352]. receptor [2058, 1597, 871, 590, 296, 2213, 1200, 1638, 146, 1931, 1964]. receptors [2643, 1942, 2943, 1178, 1513]. reciprocal [1390]. recirculating [3024]. Recognition [3078, 2078, 2043, 616, 2466]. recognized [1700]. recolonization [2130]. recombinant [2878]. recommendations [3067, 465]. Reconciling [139]. Reconfirming [43]. Reconsidering [2686]. reconstruct [1283, 2332]. Reconstructing [2261, 255, 2203, 865]. reconstruction [2221, 411, 2234]. record [2241, 979, 484, 2564, 580, 2123, 2391, 385, 1894, 2339, 2908, 1304, 432, 2758, 1257, 2335, 2859, 2820, 1544, 388, 2602, 2857, 2649, 586, 1124, 107, 1064, 930, 1722, 2180, 2757, 1039, 1854, 2109, 1207, 1296, 431, 45, 906, 1094, 1622, 2495, 912]. recorders [2204]. recording [961]. recordings [2230]. Records [2441, 644, 1042, 1920, 2724, 2101, 1460, 567, 1649, 2991, 2629, 2066, 64, 1530, 973, 1922, 1786, 2240, 3061, 1370, 1487]. recovering [1429]. Recovery [1988, 784, 1808, 1732, 844, 2989]. recreational [1556, 1557, 2199, 2899, 158, 1345, 2900, 2901, 1229, 159]. recreational-fisheries [2901]. recreationally [1301]. recruit [1658]. Recruitment [98, 3030, 1705, 2502, 1610, 315, 2188, 2678, 2753, 365, 325]. recruits [483]. Recurrent [1130]. recurring [2908, 626]. recycles [1239]. Red [1276, 1952, 1557, 2681, 47, 334, 367, 2827, 2489, 2348, 1352, 2423, 1936, 3068, 265, 2065, 1656, 1678, 2774, 280, 3087, 1526, 552, 1385, 2388, 2074, 1257, 2291, 744]. red-spotted [265]. redd [348]. redds [962, 223]. redefined [1143]. Redescription [1318, 2327, 2003, 2586, 2380, 1235, 1884, 2953, 132]. redeye [999]. Rediscovery [2839, 1359, 132]. redlip [2684, 2333]. redthroat [1254]. reduce [2977, 3043, 2967, 1692]. Reduced [2345, 3072, 1298, 1056, 1563, 47, 559, 2641]. reduces [573, 2102]. Reduction [682, 1243, 579, 1455, 2703]. reductions [2866]. redundancy [2374]. Reef [1816, 660, 1385, 1350, 3071, 587, 245, 2067, 2517, 976, 2854, 2655, 1076, 1356, 2330, 2874, 735, 2093, 2428, 713, 409, 2599, 2735, 2825, 1187, 1762, 1305, 250, 2942, 2075, 2503, 1553, 1435, 1152, 1189, 966, 2544, 953, 198, 2921, 2738, 241, 2974, 2801, 1058, 2891, 1673, 1078, 971, 88, 1455, 1075, 2556, 2798, 2652, 1609, 2648, 2012, 2030, 2866, 1938, 542, 1730, 708, 1534, 237, 1268, 1099]. reef-dwelling [966, 2544]. reef-fish [245, 2854, 2798, 708]. reefs [2638, 2027, 324, 1591, 713, 2179, 2304, 1553, 2069, 953, 2824, 2463, 1132, 2216, 285]. refeeding [1942, 2714]. Referees [48, 413, 703, 1044, 1625, 2277, 2571, 1321, 1878]. reference [39, 2964, 2951, 2322, 2161, 1156, 1958]. refined [2267]. reflect [483, 676, 2098, 1256]. reflects [2667]. refrigerated [190]. refuge [1911, 307, 2331]. refuges [1584]. refugia [2941, 14, 2797, 3028]. refugial [2055]. refugium [267]. Regalecidae [1174, 1158]. Regalecus [1158, 1174]. regalis [2335, 467]. regarding [290]. regenerating [2595]. regeneration [2750]. regime [2087, 1583, 633, 2903]. regimen [1472]. regimes [1506, 2818, 2807, 1501, 1844]. Region [1048, 2096, 85, 1636, 395, 1646, 264, 2232, 799, 1754, 27, 155, 2153]. Regional [2646, 865, 1677, 3007, 542]. regionalization [1647]. regionally [2755]. regions [1017, 2515, 426, 816, 2555, 477, 2858, 3058]. regius [1280]. regression [1930]. regulate [89]. regulated [1597, 106, 886, 30]. regulating [2671]. Regulation [2612, 2722, 2643, 1450, 10, 3, 1901, 693, 2589, 2815, 2981, 1440, 1931]. regulations [1845]. regulator [1200]. regulators [2815]. regulatory [2961]. reidi [500, 2623]. Reinhardtius [200, 1682]. Reis [2250]. relate [1536, 2400]. related [1385, 166, 2809, 1178, 1227, 2701, 2592, 1540, 1703, 2777, 603, 1521, 1598, 1476, 2168, 2373, 512, 2477, 996, 423, 259, 2009, 1347, 2, 850, 1921, 2565, 84, 1002]. relation [2655, 1450, 1407, 864, 106, 451, 1085, 1352, 2535, 661, 544, 1800, 1850, 1291]. relational [1639]. Relationship [1674, 846, 847, 2947, 696, 1170, 256, 3007, 503, 2719, 2186, 472, 2484, 1358, 2218, 1860, 1885, 237]. Relationships [1765, 2203, 2638, 2126, 95, 461, 1294, 411, 1287, 1798, 1194, 1304, 2364, 61, 3022, 823, 1057, 676, 911, 2744, 1107, 3033, 1812, 806, 591]. Relative [781, 207, 2264, 2735, 1183, 2087, 1378, 2010, 1695, 1050, 2442, 542]. relatively [1103]. Relatives [549, 1408, 790, 208]. release [715, 1563, 2208, 922, 2219, 2297, 1735, 1544, 1365, 2629, 1142, 2848, 475, 41, 2973, 1885]. released [2469, 1280, 1389, 2258, 609, 2967, 1560]. releasing [2819, 18]. relevance [1804, 1048, 1823, 2368, 475, 1247, 1055, 2367]. relevant [1140, 777, 1930, 329, 1866, 1626]. reliability [1148, 1154]. reliable [728, 585]. relict [1319, 1541]. remain [962]. Remarkable [928, 2558, 3005]. Remarkably [264, 1589]. remarks [1530]. remnant [1083]. remodelling [856]. remote [1109, 2966, 1933, 953]. remotely [1158]. removal [1977, 927]. removes [2668]. Renal [2191]. rendering [1955]. renewable [876]. reovirus [1867, 1638, 2589, 1106]. repeat [1065, 1648]. Repeatability [1368]. repeated [2773, 1066]. repetitive [714]. replenishment [449]. Reply [2737, 1516, 994, 2368, 291, 2251, 2249]. report [979, 2790, 2688, 329, 2819, 1537, 669]. reported [2242]. Reporting [414]. representation [1056]. represented [2654]. representing [1304]. Reproduction [2502, 2231, 1813, 2733, 2908, 7, 2297, 10, 1745, 1435, 2257, 172, 3, 4, 1932, 1605, 1941, 1599, 308, 708, 94]. Reproductive [2346, 1, 2034, 2000, 123, 521, 2296, 2959, 2420, 755, 2483, 194, 2091, 3025, 2451, 378, 814, 1935, 1907, 91, 294, 2135, 925, 2389, 340, 575, 2304, 2846, 960, 2045, 2348, 919, 198, 2371, 1757, 757, 2360, 2723, 121, 1812, 3046, 1087, 1012, 1614, 1908, 2218, 2831, 798, 2033, 2049, 493, 1029, 1119, 1176, 2418, 593, 418, 921, 1383, 2754, 1234, 2182, 2720, 533, 2356, 1353, 231, 6, 1407, 1286, 2141, 355, 127, 915, 1303, 351, 51, 2747, 1866, 828, 366, 296, 1474, 2484, 1981, 401, 1259, 2155, 2290, 224, 352, 1754, 1626, 415, 750]. reproductive [2443, 2037, 2948, 1671, 1370, 810, 2107, 317, 2, 1485, 2139, 796, 758, 164, 1043, 228, 1527, 283, 1000, 1171, 547, 600, 183]. reproductively [1288, 189]. Republic [1618, 1351]. repulsion [2459]. requirements [316, 2271]. rerio [1702, 2487, 68, 2939, 2813, 1798, 2785, 1443, 1469, 1114, 2415, 2529, 2804, 1251, 2292, 1358, 1838, 1917, 3053, 2131, 1066, 2552, 228, 84, 1958]. Research [1099, 2539, 182, 2007, 178, 263, 161, 872, 2292, 2921, 391, 2009, 2864, 744, 803]. reserve [1402, 2748, 1722]. reserves [2735, 118, 635]. Reservoir [1059, 2235, 1528, 1113, 1994, 2600, 1424, 434]. reservoirs [2591, 800, 1260]. Residence [226, 1898]. Residency [516, 1695, 1076, 1937, 2988, 1196]. resident [667, 98, 2308, 1163, 1427, 1430, 1131, 682, 2732, 2188, 233, 481, 2858]. Resilience [715, 1427, 1432, 1431, 2629]. resistance [2788, 104, 597, 2589]. Resolution [613, 335, 2336, 1685, 1425]. Resource [2885, 3074, 1098, 1031, 2786, 659, 2874, 1555, 2379, 1033, 53, 2666, 368, 3028]. Resources [1209, 982, 1239, 263, 2916, 2993, 2900, 1605, 1865]. respiration [1826, 1825, 1821, 1328, 1333, 1332]. respiratory [2781, 1336]. respirometers [1820]. Respirometry [2311, 1821, 1829, 1828, 1819]. respond [874]. responding [2258]. Response [1778, 1181, 68, 2939, 564, 1092, 195, 1808, 2255, 1384, 2214, 545, 662, 302, 1606, 2677, 2883, 903, 2219, 2087, 1493, 1603, 1389, 1867, 553, 2213, 1889, 1711, 1963, 837, 688, 1049, 1483, 1638, 1336, 36, 406, 2354, 2748, 633, 1780, 514, 1026, 909, 1712, 1736, 2582, 1964, 2961, 2937, 1243]. Responses [1827, 2026, 32, 1386, 468, 1330, 702, 898, 565, 145, 1631, 1684, 1772, 2403, 51, 452, 579, 1097, 1715, 1190, 476, 2325, 1379, 942, 275, 1881, 1535, 1229, 3024, 2331, 2077, 987, 683, 2053, 1159, 2706, 826, 2562]. responsible [2899, 1185]. responsive [1241]. resting [435, 1748, 2973]. restoration [2759]. restored [756]. restraint [1715]. Restricted [1734, 920]. restriction [1472]. results [62, 2056, 308, 2157, 414]. Retention [65, 424, 1490, 853, 2827, 2626, 3085, 1284, 2992]. reticulata [1179, 207, 594, 1635, 1683, 190, 1780, 2767]. reticulatum [1203, 2495]. reticulum [584]. retina [666, 2222]. Retinal [1843, 3004]. retinoic [2552]. retinoic-acid [2552]. retinoid [101]. Retraction [219, 242]. Retrospective [174, 376]. retrospectively [882]. return [2523, 1430, 1806, 2699]. reveal [2673, 1701, 880, 1040, 1668, 619, 2970, 16, 1938, 578]. revealed [3084, 2270, 2042, 1623, 2873, 2741, 817, 3085, 2018, 3037, 1677, 2411, 2205, 1898, 878, 1368, 2407, 1277, 952, 212, 622, 1594]. reveals [2782, 2910, 3083, 2200, 2795, 3088, 1961, 1936, 1262, 2417, 213, 2585, 217, 2956, 839, 2715, 552, 2721]. reversal [2847, 1206, 1185]. reverse [2090, 394, 1518]. Review [208, 289, 2631, 1550, 1373, 112, 1210, 1267, 113, 1426, 955, 908, 605, 1018, 108, 1126, 142, 550, 462, 140, 110, 1265, 246, 1463, 1161, 261, 141, 549, 812, 690, 983, 1172, 1209, 1269, 1268, 1270, 1371, 1372, 3016, 525, 1070, 1173, 604, 831, 1099, 1211, 568, 182, 956, 588, 3017, 2070, 859, 243, 1127, 109, 2314, 183, 1174, 739, 2822, 2893, 439, 9, 671, 1906, 139, 138, 189, 356, 111, 311, 438, 548, 672, 389, 980, 981, 982, 1264, 1266, 1098, 1404, 1374, 1874, 1873, 2453, 461, 569, 3067, 777, 2007, 392, 616, 783, 2208, 1364, 1743]. review [1452, 872, 1401, 984, 873, 691, 2032, 2212, 788, 2437, 747, 2009, 2180, 744, 1997, 742, 890, 461, 982, 9]. revised [1906]. revises [2133]. Revision [420, 2138, 2577, 1088, 2051, 607, 2263]. revisited [2184, 419, 2694]. revisiting [1453]. RFID [1458]. rfishbase [970]. RFLP [20]. Rhabdosargus [721]. Rhadinocentrus [363]. Rhamdia [932]. Rhamphichthyoidea [2772]. rheocharis [2000]. Rhincodon [979, 1611, 2074, 1196, 774, 1544, 2836, 775, 2682, 765, 930, 387, 747, 1019, 1296, 3015, 2812]. Rhine [1745]. Rhinichthys [2224]. rhinobatid [2045]. Rhinobatidae [1733, 1012]. Rhinobatos [1733, 1909, 1012, 1813]. Rhinogobius [2076]. Rhinopetitia [1393]. Rhinoptera [2885]. Rhizoprionodon [1776, 2770, 1592, 124]. Rhodeus [1030, 374, 1382]. rhodomelas [308]. rhomboides [2398]. rhombus [652, 1293, 627]. Rhynchoconger [385]. rhythms [1480, 2]. ria [2231, 1758]. ribbontail [907]. ribonuclease [1241]. Rica [940, 1011, 2617, 724]. ricefield [1521]. rich [608]. Richard [2280]. Richards [739]. richardsoni [2063]. richness [2382, 1109, 1536, 1991, 3034]. ridens [2034]. Ridge [2779]. Rift [624]. Rig [1964]. Rig-I-like [1964]. Riga [342]. right [147]. right-mouth [147]. riisei [1181]. Rim [1433]. ring [1629]. ringens [1311, 423, 1578, 2678]. rings [2086]. Rio [1393, 2422, 2619, 1773, 1315, 2558, 2395, 2185, 2509, 567, 2514, 1081, 1987, 2003, 1393, 2887, 2400, 2437, 3005, 839, 2743, 2073, 3080, 375, 2008, 2113]. Riparian [55, 1977, 318, 1566]. rise [1231, 273]. Risk [1237, 68, 2786, 153, 565, 302, 1916, 1684, 780, 2403, 2131, 2726, 677, 1392]. Risk-taking [1237, 2786, 2131]. Risks [2391, 1693]. Rissik [138]. River [1618, 1503, 2282, 1666, 2843, 222, 90, 1983, 2549, 2610, 1498, 2360, 2563, 1988, 2780, 1989, 633, 1906, 877, 2752, 2495, 1930, 2526, 286, 2262, 2958, 2364, 632, 835, 1593, 2777, 2999, 2379, 1117, 2200, 1205, 1059, 1978, 106, 1192, 2933, 1904, 374, 2272, 886, 2231, 78, 239, 30, 515, 2978, 2835, 405, 3006, 323, 2624, 99, 1601, 226, 2621, 872, 1745, 1571, 2741, 130, 949, 26, 426, 2608, 1981, 172, 1982, 30, 1903, 2442, 2504, 1841, 2542, 2851, 815, 2324, 1547]. Riverine [1468, 295, 2639, 948, 1433, 180, 255, 1139, 60]. Rivers [349, 55, 2440, 2627, 1842, 1981, 1281, 2666, 2005, 365, 935, 2252, 2329, 1985]. rivulatus [2175]. Rivulidae [1896, 2486]. rivulus [1749, 2944, 1866, 3023]. RNA [2664, 338, 1241, 225, 826]. RNA-based [338]. RNAs [2809, 2964]. roach [963, 598, 202, 1354, 2930, 840, 2505, 1900, 3073, 1295, 57, 558, 135, 89, 1175]. roar [2338]. Robert [1405]. Roberts [1174]. robust [802, 2442]. robusta [2357]. robustness [2677]. rochei [1747]. rock [1239, 651, 2458]. Rockall [1815]. rockfish [844, 2276]. rockfishes [891]. rocky [2638, 2502, 2027, 2019, 1591, 953, 2172, 2540, 801, 806]. rod [666, 2928]. rodlet [1682]. rogersi [757, 456]. role [137, 631, 745, 1231, 2140, 2162, 795, 2988, 1994, 2735, 1835, 347, 2666, 946, 1979, 9, 1729, 2394, 2525, 1244]. Roles [10, 1333, 2942, 2671, 1050, 909]. Roques [245]. rosae [1541]. rose [1793]. Rosemary [1839]. Ross [1003]. rostral [2460]. rostrata [1616, 1883, 2187, 90, 1411, 863]. rostratum [2549]. rostrum [2460]. rotenone [1093]. rough [2629, 1562]. roughshark [1785]. roughy [1970, 64, 1277]. Round [513, 2121, 2845, 293, 2720, 1899, 1146, 756, 691, 1001, 211, 1724, 517, 1604, 2855, 2654, 221]. roundel [547]. roundscale [954, 2268]. route [516, 2355]. routes [346]. routine [687, 934, 734, 1082, 2131]. ROV [2407]. roving [846]. Roznovská [1618]. RT [1958]. RT-qPCR [1958]. Ruban [438]. rubiginosus [1164]. rubripes [590, 581, 1520]. rubrolabiatus [2844]. rubroviolaceus [2333]. ruffe [1117]. rule [2384]. run [2523, 515, 331, 1198]. run-time [2523]. runner [1460]. running [1801]. rupestris [2409]. Russia [598, 859]. Russian [2059, 42, 1690]. rustica [1299]. ruthenus [2888, 968]. Rutilus [963, 598, 354, 202, 1354, 2930, 840, 2505, 1900, 3073, 1295, 57, 558, 607, 135, 89, 1175]. rutrum [2371].

S [1426, 462, 1269, 1270, 1173, 859, 671, 111, 438, 548, 2489]. S. [2672, 1006]. s.l [1186]. sac [2931, 893]. Saccodon [2006]. saccular [1208]. saccus [1710].

Sacramento [1546]. saddled [1481]. sagamianus [3090]. Sagamichthys [2802]. sagax [1732]. Sagittal [1756, 1086, 2611, 2201]. Saharan [24]. saida [1414, 1858, 1895]. sailfin [205, 1945]. sailfish [1307]. saithe [2169]. sajori [2153]. Sakhalin [295]. salar [2869, 1415, 400, 2523, 2811, 2156, 2532, 889, 444, 2326, 2708, 2696, 2598, 2705, 1857, 2526, 2773, 221, 516, 2117, 301, 2137, 1514, 316, 1833, 1147, 572, 1519, 2072, 875, 899, 988, 1432, 2524, 355, 943, 1508, 2320, 1238, 1507, 2816, 2087, 1735, 2181, 174, 2656, 2941, 1667, 3003, 1571, 2527, 393, 1966, 871, 1387, 1851, 2064, 2763, 1792, 2800, 130, 949, 1192, 2081, 2651, 2626, 3052, 1065, 1053, 562, 1107, 1498, 2707, 2600, 2698, 417, 2080, 1509, 1501, 1262, 318, 2265, 1472, 2547, 554, 1500, 2699, 1562, 1529, 2286, 555, 2967, 477, 904, 1502]. salar [1850, 2466, 2694, 1560, 1511, 1692, 2390, 128, 2274, 2903, 869, 2283, 1006, 1290, 1505, 37, 2706, 2157, 1967, 2110, 959, 223, 251, 1856, 2582, 1499, 1425, 1243, 405]. Salaria [1889, 1281]. salaris [1093]. sale [2905]. Salí [375]. salina [1768]. salinities [2214]. Salinity [117, 131, 3032, 1506, 1616, 2596, 1364, 934, 2375, 278, 1691, 658, 840, 2489, 1711, 2707, 2642, 1037, 1850, 1945, 730, 1518, 1621]. Salish [2756]. Salmincola [1065]. Salminus [625]. Salmo [2869, 1415, 400, 657, 2523, 2811, 2156, 2532, 889, 444, 2326, 2708, 1618, 2696, 611, 1569, 2491, 2405, 2598, 2705, 85, 1857, 2526, 2773, 667, 221, 1093, 1563, 1468, 516, 2117, 2494, 301, 2137, 1514, 2435, 1022, 316, 1833, 1147, 572, 1519, 2677, 2988, 2072, 875, 2079, 899, 1746, 1631, 988, 1432, 2142, 962, 1120, 1035, 2524, 355, 917, 1801, 943, 1508, 2320, 1238, 1507, 2816, 2087, 2186, 2336, 1735, 2181, 2440, 174, 700, 2656, 573, 2941, 1667, 3003, 1571, 2527, 393, 51, 880, 817, 1406, 1966, 871, 1387, 1723, 1851, 2064, 2763, 2740, 1131]. Salmo [1792, 2800, 130, 949, 1192, 2081, 2651, 2626, 3052, 658, 1118, 1065, 845, 1053, 1836, 2159, 562, 460, 978, 1107, 1498, 2707, 14, 2600, 2761, 2698, 2616, 417, 2470, 2080, 348, 1509, 1501, 1262, 924, 318, 2265, 1472, 2547, 1185, 1484, 2989, 1888, 554, 1284, 1500, 2699, 1562, 406, 2084, 1529, 2286, 2653, 2978, 2881, 935, 555, 2967, 1039, 477, 2332, 904, 1502, 1850, 2466, 2694, 1560, 1511, 1692, 331, 1198, 2390, 128, 2274, 2903, 2525, 869, 2283, 2672, 1006, 1290, 1505, 37, 2706, 612, 2157, 1967, 2110, 959, 857, 2973, 223, 251, 1856, 416]. Salmo [2582, 1499, 1425, 1243, 405]. salmoides [934, 455, 2951, 1253, 599, 1375, 1418]. Salmon [2976, 671, 672, 2869, 400, 2523, 2637, 2811, 2668, 2156, 2532, 889, 2760, 444, 2326, 737, 2708, 2696, 2628, 1503, 1528, 2296, 2598, 2705, 702, 1857, 697, 2526, 2773, 1764, 2255, 3001, 1429, 6, 221, 516, 258, 2117, 301, 2137, 1514, 1149, 1834, 316, 2701, 2994, 1833, 1147, 1519, 2072, 875, 2931, 899, 988, 1721, 1593, 101, 335, 2345, 1956, 278, 2524, 355, 943, 1508, 2320, 1238, 1507, 2816, 2087, 874, 1735, 2181, 2440, 872, 174, 2275, 2656, 2941, 1667, 3003, 1571, 2527, 393, 2700, 559, 1966, 871, 1851, 2064, 2711, 2763, 1583, 838, 1792, 2800]. salmon [130, 949, 1192, 2081, 2651, 2626, 3052, 894, 1065, 1053, 2254, 562, 38, 1536, 1034, 1107, 1498, 2693, 2707, 2600, 2985, 2704, 2636, 2698, 417, 259, 1789, 2732, 1512, 2080, 1509, 1501, 1262, 78, 318, 2265, 1472, 2547, 236, 554, 1500, 2699, 1562, 2578, 1529, 2352, 325, 555, 1497, 2967, 477, 1256, 904, 1502, 1850, 2466, 2694, 1560, 1511, 1692, 2390, 128, 2274, 2903, 1840, 869, 2283, 870, 1006, 1290, 1505, 1102, 37, 2706, 2157, 1967, 2110, 959, 916, 197, 223, 1074, 251, 1856, 698, 1298, 1862, 2582, 480, 1594, 2118, 1499, 1243, 311]. salmoneus [1065]. salmonicida [1712]. Salmonid [180, 2111, 1427, 1626, 4, 2160, 36, 1566, 1368, 815, 480, 569]. Salmonid-egg [180]. Salmonidae [2383, 2672, 1648]. salmonids [2528, 444, 3031, 3048, 674, 1441, 2315, 1400, 19, 1404]. Salmoninae [571]. salmonis [2704, 2616, 904, 164]. salpa [1388]. salt [1084, 693, 1037]. salt-marsh [1084]. saltatrix [328]. saltwater [2705, 1355, 71]. Salvelinus [1506, 2818, 1279, 2760, 2097, 2264, 1524, 1488, 1287, 910, 662, 2308, 819, 258, 316, 192, 2383, 1746, 1631, 1593, 102, 2719, 917, 1510, 1585, 1431, 2336, 603, 1957, 177, 880, 1406, 1642, 435, 2046, 579, 1715, 1891, 1104, 1288, 344, 2797, 1376, 806, 638, 1865, 1471, 1695, 46, 164, 1090, 1395, 3028, 1031, 165, 1648]. same [1636, 1668]. sample [2975, 2434]. sampled [1512, 1709]. sampler [397]. samples [410, 2532, 620, 1410, 1442, 2332, 2157]. Sampling [1891, 1146, 2663, 1421, 1949, 2856, 1493, 1225, 585, 1853, 1066, 2204, 1072, 3008, 3009]. San [2822]. sanctaefilomenae [999]. sand [1491, 1233, 939, 3002, 944, 2746, 552]. sand-diving [2746]. sand-dwelling [552]. sand-smelt [1491]. sandbar [2372, 1932, 2077]. sandeel [227, 2362]. Sander [2482, 1916, 3051, 841, 2999, 2930, 1260, 401, 1615, 2929, 654, 422]. sandy [2026, 1381]. sandyback [758]. sanguijuela [2050]. sanguineus [2309, 539]. sanjaymoluri [1951]. Santa [2269]. São [2003, 1720, 1012, 2606]. sapidissima [2615, 2584]. sarcoplasmic [584]. Sardina [1899, 1660, 592, 1190, 1369]. sardine [1899, 1732, 341, 2222]. sardinella [2121, 2845, 1899, 669, 2121, 2845, 1899, 341, 2133, 669, 2225]. sardines [1660, 1791, 1190]. Sardinops [1732, 2222]. sargus [2712, 2712]. Sarotherodon [730, 424]. Sarpa [1388]. Satellite [76, 282, 234, 2181, 2583]. Satiation [2046]. saturation [1618]. Saudi [1359, 3045, 3055]. sauger [2999]. Saunders [1210]. Saurogobio [2504]. Saving [2158]. Savody [1099]. sawfish [1807, 2820, 1686]. Sawfishes [2270]. sawshark [2460, 2293]. sawsharks [3060]. saxatilis [346, 1748]. scabbardfish [2758]. scabbardfishes [1123]. scaber [2511]. scalare [2088]. Scale [1899, 354, 1252, 764, 1140, 1705, 1864, 1283, 2086, 1147, 1101, 102, 913, 1410, 449, 651, 196, 1261, 1476, 1143, 939, 147, 1655, 2803, 86, 3037, 811, 638, 2748, 501, 2903, 2283, 3040, 3009, 1103]. scale-age [196]. scale-cover [913]. scale-eating [147]. Scale-stacking [1252]. scaleless [332]. Scaleprinting [86]. scales [797, 2816, 1378, 2029, 2268, 2595, 2084, 2406, 1029, 1247, 968]. scaling [679, 789]. scalloped [3029, 1587]. scan [3005, 1002]. Scanning [1458]. scaphiops [2408]. Scaphirhynchus [2205, 650, 2442]. Scaridae [2771]. scarring [765]. scars [1701]. Scarus [1385, 2388, 2333]. scat [1581]. Scatophagus [1581]. scattering [2464]. Scavenging [541, 771]. scenario [2530]. scheme [2539, 2698]. schilbid [823]. Schindleria [2291]. Schindleriidae [2291]. Schistocephalus [1681]. Schizothoracinae [2627].

Schizothorax [1448, 1440]. schlegeli [47, 2276]. schlegelii [2348, 2679, 1097]. Schlupp [690]. schmidti [643]. schnakenbecki [2802]. Schneider [2724]. schoenleinii [232]. School [1535, 2742, 1384, 1542, 2574]. Schooling [618, 1384, 134, 116, 592, 469, 2192, 1078, 1500]. Schrank [110]. schrenki [1598]. Schroederichthys [2493]. Schultze [112]. Sciadicleithrum [407]. Sciaena [2597, 1924]. sciaenid [543]. Sciaenidae [1846, 2364, 2335, 2396, 816, 2273, 2052]. sciaenids [1208]. Sciaenops [1557, 1936, 2065, 1656]. Science [289, 1222, 1226, 157, 1496, 1802]. scientific [178, 1098, 1374]. scientists [1228, 2689, 1554]. Scolopsis [1199]. Scomber [1242]. Scomberomorus [576, 1643, 2446]. Scombriformes [2758]. Scombropidae [17]. scombropids [17]. scombrus [1242]. scope [2609, 697, 1827, 1149, 1832, 1838, 1829, 2512]. Scophthalmidae [1206]. Scophthalmus [1710, 652, 1293, 1597, 1680, 1901, 2354, 225, 706, 627, 1244]. Scorpaena [2108]. Scorpaenidae [2646, 891]. Scorpaeniformes [891, 2108, 1391]. Scotia [1783]. Scotland [1849, 2526, 102, 2616, 2793]. Scottburgh [2855]. scotti [828]. Scottish [2637, 889, 350, 2157, 251]. Screening [216, 697]. screens [2934]. screw [2598]. scuba [2854]. sculpin [2130, 277, 1010, 2331, 2835]. sculpins [1481, 1681, 2184]. scutes [968]. Scyliorhinidae [766, 1164, 254, 2983, 2493]. Scyliorhinus [3078, 941, 2601, 2983, 1000, 2721]. Sea [983, 342, 831, 2070, 1003, 1783, 2223, 311, 2691, 797, 887, 2668, 2708, 2696, 608, 1396, 769, 1623, 735, 1093, 1130, 1892, 1276, 923, 2435, 2299, 2079, 395, 47, 2868, 2345, 334, 2993, 2181, 2258, 1612, 3036, 1966, 709, 1387, 1313, 452, 2827, 1131, 2257, 1655, 1464, 2960, 1259, 1530, 1480, 1275, 382, 2583, 1872, 1273, 273, 2347, 1647, 1122, 904, 308, 2445, 164, 1422, 60, 2102, 2444, 1552, 1385, 2388, 2241, 1952, 1234, 2817, 2588, 631, 509, 2301, 1344, 364, 1611, 2074, 2597, 2023, 2262, 678, 534, 1014, 1257, 1407, 1916, 2291]. Sea [2173, 2859, 3082, 1934, 2712, 347, 1582, 1884, 676, 2917, 2348, 2679, 2611, 2756, 1909, 1419, 1206, 2857, 3081, 1242, 2550, 1480, 1901, 536, 328, 307, 1038, 1223, 2447, 1182, 2730, 1064, 738, 1948, 136, 193, 824, 935, 2914, 2954, 517, 744, 1447, 2418, 2109, 2915, 1207, 1296, 45, 37, 2789, 1422, 959, 113, 1209, 2070]. Sea-level [2223]. seabass [961, 518, 1292]. seabasses [681]. seabed [1218]. seabream [701, 613, 2712, 204, 44, 2348, 2679, 1097, 1274, 1843]. seadragons [489, 497, 491]. Seagrass [1388, 1402, 1294, 1781, 2103, 483, 2374, 1940, 501, 578, 1730]. seagrasses [535]. Seahorse [1758, 1920, 2564, 2625, 1339, 2461, 1769, 1314, 1516, 920, 1523, 1515, 2718, 2623, 2644, 2948, 2965, 496, 3077]. seahorses [489, 1863, 1248, 1565, 1204, 491, 494, 1171, 490, 208]. Seamount [2063, 895]. search [144]. Searching [2899, 2742]. seas [2977, 2868, 2100, 105]. seascape [1101]. season [2264, 1488, 2398, 2727, 1680, 118, 512, 2544, 1865, 1175, 1349]. Seasonal [1385, 2090, 306, 450, 2042, 1752, 192, 374, 2136, 1548, 2779, 1361, 53, 2188, 2014, 1985, 2948, 931, 3015, 1513, 1193, 1809, 93, 2839, 1613, 1810, 1062, 295, 2353, 2484, 344, 1512, 318, 1865, 2021, 1007, 2085, 2855]. seasonality [2262, 1789, 1449]. seasonally [1580, 98, 343]. seatrout [2219, 1061, 1446]. seaward [1387, 1731]. Seawater [1006, 866, 2702, 2664, 1514, 1238, 1507, 274, 1966, 1747, 1034, 2707, 1262, 693, 1511, 1290, 2706]. Sebastes [891, 844, 1160, 2276]. Sebastidae [1160]. Seber [2465]. second [2167, 2068, 2570, 1854, 705, 1243]. second-generation [1854]. secondary [2484, 3076, 1840]. secretion [1313]. secretory [3059]. sediment [396]. sedimentary [2929]. Sedlo [895]. seed [1949]. Seeking [1790, 2652, 60]. seen [2299]. seeps [998]. segments [666]. segregated [2239]. Segregation [1809, 2173, 1476, 729, 2020, 2018, 2294]. seine [1190, 2654]. seismic [1415]. Seitz [1211]. Selachii [1908]. selected [204, 326, 3069, 88, 1212]. Selection [2951, 1616, 520, 2759, 2597, 202, 751, 1959, 1584, 1559, 366, 360, 1444, 599, 53, 279, 924, 2028, 1635, 2473, 1619, 1676, 117, 654, 399]. selective [889, 984, 1727, 2085]. selectivity [2998, 2370, 896, 1566, 1159]. self [1749, 1896, 1225]. self-fertilizing [1749, 1896]. self-sampling [1225]. semelparous [3076]. semen [1450]. semi [2708, 2011, 176, 2013, 1555, 3086, 1034, 303, 2012]. semi-arid [176, 2013, 2012]. semi-closed [2708]. semi-enclosed [2011]. semi-natural [1555, 3086, 1034]. semi-terrestrial [303]. Semicossyphus [1870]. semifasciatus [576]. semifasciolatus [836]. semilaevis [2826, 82, 380, 1568, 2476, 1931]. semilunaris [434]. seminal [1944]. semiochemical [299]. Semotilus [1073, 2739]. Senegal [529]. senegalensis [1245, 1697, 529, 1191, 1794]. Senegalese [1245, 1697, 1191, 1794]. senegalus [2937]. senescence [559]. senior [1318]. sense [60]. sensitivity [1218, 2306, 2806, 2928, 2936]. sensor [1947]. Sensory [2938, 2923, 2940, 2920]. sensu [964]. separated [1707]. separating [236]. separation [992]. Sequence [2222, 334, 714, 799, 571, 1763, 1492, 1938]. sequences [2150, 836, 373, 2364, 465, 386, 2232, 952, 2835, 19]. Sequencing [2829, 2664, 2809, 1459]. sequentially [1777, 1918]. Serbia [1675, 1201]. Serendipitous [3088]. series [2549, 1750, 1031]. Seriola [3062, 2178, 1036, 1482]. Seriolella [3025]. serotypes [1608]. serranid [1183, 2730]. Serranidae [1661, 1014, 1286, 2288, 2448, 3061, 2404, 736]. Serranus [3061]. Serrasalmidae [2002, 2619, 2953]. serrasalmids [2002]. sertanejo [2358]. serum [1340, 668]. serve [2093]. services [640]. set [1415, 1203, 3003]. set-up [3003]. setting [2097, 3053]. settlement [534, 2825, 2269, 2216, 2947]. settlement-stage [2825, 2216]. setup [1819]. seven [2221]. sevengill [306, 2518, 2859, 764]. several [607, 1897, 2947, 517]. severe [227]. Severn [323]. sewage [2113]. Sex [3071, 1344, 1317, 1731, 1694, 1197, 166, 2809, 2847, 1178, 795, 2665, 1845, 2785, 1778, 922, 21, 711, 1523, 2842, 479, 2136, 2804, 2032, 1253, 594, 776, 1135, 2402, 1185, 2850, 2957, 2307, 144, 1007, 164, 1579, 2654, 1425, 1809]. sex-based [1523]. sex-dependent [2654]. sex-ratio [1185]. sex-related [2809]. Sex-specific [1344, 1731, 1694, 21, 1809]. sexes [2182]. Sexing [287, 1417]. Sexual [244, 1125, 200, 1723, 1582, 695, 456, 492, 1247, 468, 2764, 1845, 1357, 205, 751, 1286, 2639, 1640, 682, 1417, 2484, 772, 96, 2437, 646, 2171, 1840, 2380, 1291].

Sexually [1798, 3069, 2983, 2558, 2285, 2810, 2437]. shad [323, 3084, 2615, 2828, 1670, 1745, 1969, 1136, 92, 3076, 2584]. Shade [3039]. shading [318]. shadowy [2341]. Shake [1300]. Shallow [1796, 2122, 767, 2999, 114, 26, 1467, 264, 1953, 2216, 1133, 2798, 2015, 1050, 2490, 1709]. shallow-water [767, 1467, 1709]. shanny [1367, 2172, 2485]. Shape [1889, 2121, 2450, 1245, 2955, 2677, 2416, 102, 1111, 1085, 1662, 1352, 1292, 1100, 1799, 1123, 2438, 1579, 1002]. shape-related [1002]. shaped [81, 967, 654]. shapes [1020, 131]. Shaping [1759]. Sharing [2450]. shark [137, 1234, 2468, 2783, 979, 766, 1776, 123, 2385, 306, 2372, 1645, 2057, 777, 1438, 2517, 2790, 2042, 785, 1954, 1701, 781, 1737, 1076, 2074, 3029, 2770, 770, 2581, 2874, 754, 2518, 1592, 1718, 1892, 2227, 1811, 1734, 1454, 940, 294, 2873, 2416, 124, 2833, 759, 1810, 2859, 474, 761, 3082, 1542, 1740, 774, 1544, 1612, 2602, 2795, 2165, 3088, 1961, 2092, 2100, 3059, 1806, 1996, 2436, 775, 2682, 2449, 1725, 746, 2561, 2852, 830, 241, 1473, 1971, 2974, 3002, 2891, 1124, 1677, 2554, 1897, 121, 1932, 3046, 1455, 1887, 626, 1761, 930, 1908, 2040, 747, 786]. shark [1019, 1805, 2814, 2418, 796, 1296, 1094, 3021, 3089, 595, 1380, 3015, 2812, 1622, 764]. shark-fin-to-body-mass [777]. shark-like [1380]. Sharks [1373, 2144, 977, 2588, 2122, 745, 1490, 2248, 1396, 2303, 1809, 282, 1808, 2041, 1611, 234, 781, 2908, 253, 767, 783, 2129, 2260, 1196, 2173, 519, 915, 472, 1486, 753, 2836, 2165, 729, 2629, 3059, 2680, 1693, 2744, 763, 1377, 2657, 1058, 2467, 1816, 2962, 765, 3074, 2648, 387, 2180, 2234, 787, 2077, 2911, 2139, 310, 2886, 1879, 2249, 94, 549, 1270]. sharpnose [2770, 1592, 2859]. sharptooth [1963]. Shatt [2836]. sheephead [1870]. Shelf [887, 779, 1239, 2030, 1297, 1213, 1814]. Shellfish [3017]. shelter [3039, 1041, 1472, 2652]. shelter-seeking [2652]. Sheltering [3020]. Shelton [311]. shi [2396, 1924]. Shield [2002]. shift [1623, 3013, 2416, 817, 1084, 315, 1744, 1898, 1009, 1775, 931]. shifting [3064]. Shifts [1095, 1138, 2910, 640, 377, 192, 940, 202, 1053, 2120, 1033, 2028, 995, 1525, 1865, 2253]. shiner [2929, 1828, 2807]. shiners [1083]. Shining [3043, 2824]. shoal [2593, 2737, 2736, 1366, 34, 1362]. Shoaling [367, 122, 2648]. shoals [722, 1532, 1953]. shock [561, 539]. shopper [2905]. shore [3071, 2502, 1150, 159]. Short [1091, 1331, 1346, 1121, 448, 11, 2461, 301, 1149, 1510, 1401, 2300, 2489, 1073, 1940, 2652, 2053, 496, 282]. short- [301]. short-day [1510]. short-lived [1121, 2300]. short-snouted [2461, 496]. Short-term [1091, 1346, 11, 2489, 1073, 2652, 2053, 282]. shortage [2528]. shortening [2491]. shortfin [1767, 1718, 2375, 853, 2092, 1725]. shortnose [610, 2174, 279, 1355]. shotgun [1650]. should [1554]. shovelnose [2205, 2218, 650]. show [271]. shows [1603]. shrimp [450, 170, 1217]. shufeldti [809]. Shulman [2281]. Shyness [583, 1779]. Siberia [1459]. Siberian [2575, 2714]. sibling [256, 1141, 122]. Sicilian [2857]. sicklefin [2346, 2779]. Sicydiinae [619]. Sicyopterus [888, 619]. side [1928, 2419]. Siganidae [2443]. Siganus [1144, 623, 2381, 2443]. sight [634]. sighting [3088, 2974]. sightings [1790, 3015]. sign [2974, 2774, 3024]. signal [2925, 145, 2306, 1610]. signals [476, 1255, 492]. signature [592, 1095, 839]. signatures [2203, 1668, 2084, 199]. significance [1817, 2494, 1664]. Significant [2386]. silico [2643]. silky [2260, 761]. Sillaginodes [483, 2515, 2310]. Siluriformes [1304, 2587, 2647, 375, 928, 2834, 1773, 2282, 1666, 2185, 1601, 2620, 2843, 1308, 1788, 2050, 2236, 3005, 2263, 591, 2606, 2586, 2749, 2956, 2397, 890, 2358]. Silurus [2312, 3072, 990, 1607]. Silver [1562, 1855, 3025, 2011, 473, 3013, 932, 2353, 2650, 2936]. silver-stage [1855]. silverfish [2766, 1003]. silverside [2047, 1461, 1960]. silversides [1194, 317]. Silvinichthys [1308]. similar [2760, 1684, 818, 2286]. similarities [2970, 3005]. Similarity [675, 131]. Simon [2039]. simple [1869, 2089, 2529, 1106]. simplex [1186]. simplified [1495]. simulated [1684, 3073, 1911, 1190, 201, 1635, 1535, 1820]. simulation [284]. Simultaneous [2076, 1425]. simultaneously [235, 2354]. since [256, 2516]. sinensis [81, 1062, 296, 1848, 877, 1664]. Singapore [160, 212]. single [2532, 1300, 1056, 395, 335, 3049, 2740, 2561]. single-nucleotide [1056]. single-pass [3049]. Siniperca [457, 1567, 1005, 2878, 3004]. Siphamia [923, 2054]. sires [260]. sister [1644, 1250]. Site [2439, 2184, 249, 2067, 2681, 835, 1248, 1523, 2756, 1686, 3037, 2831, 3015, 2188]. site-attached [2067]. sites [1739, 245, 2997, 1661, 1356]. situ [1704, 1158, 1755, 2489, 1487, 736]. six [1438, 1755, 2771, 1416]. Sixth [1212]. Size [2459, 40, 302, 1588, 2756, 1727, 654, 1007, 228, 2886, 165, 2506, 2869, 1765, 582, 728, 2786, 2264, 411, 2362, 1701, 781, 1353, 667, 448, 610, 2242, 1030, 2579, 1008, 2126, 934, 1348, 503, 2524, 1183, 851, 1062, 3011, 573, 2977, 1962, 1387, 1723, 2064, 590, 1057, 1582, 2136, 1417, 512, 2544, 57, 2718, 2032, 2574, 2535, 2546, 776, 2641, 2080, 379, 2148, 3072, 2065, 847, 1010, 1676, 95, 653, 2957, 2377, 2697, 1119, 2442, 1025, 2085, 1307, 2981, 3021, 1579]. size-at-maturity [3011]. size-based [2535]. Size-dependent [2459, 40, 654, 228, 165]. Size-mediated [302]. Size-selective [1727, 2085]. size-specific [2080]. size-structure [1307]. sized [2626, 2272]. sizes [1845, 932, 281, 2046, 966, 1820]. Skadar [607]. Skate [779, 1704, 1166, 792, 1882, 2217, 1589, 814, 2729, 1907, 2319, 940, 771, 2063, 2977, 1052, 1057, 2408, 1757, 1573, 1887, 252, 2380, 2355, 547, 1814]. skates [2856, 2607, 1934, 2657, 778]. Skeletal [2826, 2396, 706, 2508, 842, 280, 2234, 1856]. Skeletogenesis [2510, 999]. skeleton [836, 598]. skiffia [967]. skilletfish [32]. skills [1656]. skin [2220, 1718, 452, 788]. skipjack [2791, 2141, 799]. skull [1100]. skunk [2535]. Skurdal [671]. sky [2795]. Skye [2616]. sledge [1458]. sleeper [770, 852]. Sleeping [2683]. slender [1257, 2337, 2230]. slickheads [1898]. slipping [1190]. slope [1316, 1814, 779]. slopes [1407, 2912]. Slow [1776, 3007, 1628]. slow-oxidative [3007]. Small [2539, 638, 2329, 2096, 2067, 2303, 2809, 448, 98, 2359, 2551, 1147, 170, 713, 1778, 1593, 2103, 1035, 3078, 2977, 941, 2747, 1962, 374, 579, 2272, 558, 2803, 346, 2779, 2698, 1677, 1727, 2601, 2666, 811, 89, 1888, 2525, 41, 2157, 1000, 3040]. small-bodied [2303, 98, 1677]. small-eyed [2977]. Small-scale [638, 2803, 811, 3040]. small-sized [2272]. small-spotted [3078, 2601]. smaller [2095]. smallest [2426]. smallmouth [841, 12, 2200, 455, 2305]. smallnose [2048]. smalltail [1124]. smalltooth [1807]. smell [2942, 505]. smelt [1491, 3086, 939, 284, 2914]. smiliogastrin [1906]. Smiliogastrini [1906]. smitti [675]. smolt [2696, 2701, 2994, 875, 988, 2524, 2816, 1571, 2700, 2693, 2600, 1509, 2694, 869, 1006]. smoltification [1514, 1511, 2693]. Smolting [1504, 2702]. smolts [400, 2523, 2708, 2598, 1563, 516, 1519, 2079, 1721, 1593, 1956, 943, 1507, 2656, 1966, 1792, 130, 2651, 562, 1498, 2707, 2698, 2080, 1501, 554, 1500, 2699, 2881, 1502, 1850, 2903, 1006, 1290, 1505, 2706, 2110, 1499]. smooth [2042, 294, 2873, 771, 1812, 82, 380, 1568, 2476, 921, 1931]. smooth-hound [2042, 294, 1812]. smoothhound [940, 2561, 1971]. snailfish [1391, 308]. snakehead [2643, 2322, 1946]. snapper [2017, 1338, 715, 1871, 343, 2688, 2309, 2423, 3068, 201, 1678, 653, 117, 1793, 593, 45, 283, 1860, 1439, 539]. snappers [163]. Sneak [1069, 2674]. sneaker [1025]. Snell [1859]. snook [1140, 2389, 2745, 642]. snout [500, 2774, 499, 1968]. snouted [2461, 2644, 2425, 496]. snow [3061]. SNP [2221, 82, 2589]. SNP-based [2589]. snubnose [1787]. Social [1017, 34, 985, 415, 3040, 2459, 2742, 2883, 2593, 2737, 991, 603, 1902, 941, 2736, 1148, 2409, 2962, 834, 540]. sociality [2661]. Socially [2847, 723]. Socio [327]. Socio-economic [327]. sockeye [2668, 2628, 702, 1149, 1721, 1593, 2345, 1956, 872, 559, 1583, 236, 1256, 1074]. Socotra [1950]. soft [396, 2721]. soft-bodied [2721]. software [3078]. solandri [1913, 2289]. solar [47]. sole [1245, 829, 1697, 2826, 2325, 529, 82, 380, 1568, 1156, 1191, 2476, 1794, 1931]. Solea [1245, 829, 1697, 529, 1191, 1794]. Solomon [2771]. Soluble [1200]. Solutions [110]. somatic [2869, 1085, 1986]. Some [2299, 1129, 1765, 2838]. somitogenesis [2267]. Somniosus [770, 746]. sonar [582, 133, 1969]. Sons [3016, 3017]. Soon [2313]. sophore [443]. sorrah [729]. Sotalia [2429]. souffia [270]. Sound [1739, 277, 1924, 2464, 1747, 911, 2831, 1526]. sound-scattering [2464]. Sounds [728, 961, 1415, 2097, 1673, 2877, 944, 1349]. soup [2984]. source [449, 365]. Sources [1820, 2587, 1852, 2904, 1216, 1668]. South [1809, 2386, 1183, 2251, 2250, 2855, 907, 1108, 1234, 2378, 2882, 1086, 306, 1882, 2450, 2057, 1479, 2427, 2790, 755, 2567, 2339, 378, 814, 1935, 254, 701, 1758, 2839, 2844, 2518, 567, 1649, 2129, 2389, 1760, 1577, 1627, 2825, 2320, 2304, 700, 2128, 175, 2168, 1474, 2971, 129, 617, 2408, 2371, 300, 1530, 2357, 2554, 1446, 1932, 660, 2016, 30, 2099, 2178, 2757, 670, 758, 1860, 977, 2385, 156, 2025, 1973, 2007, 2356, 3050, 1125, 2729, 1974, 1998, 1846, 2758, 1105, 1228, 2966, 2559, 114, 644, 1978, 2348, 2679, 2561, 2852, 2610, 1651, 1991, 1975]. South [2806, 2263, 1012, 2030, 1451, 1997, 154, 2106, 339, 1814]. south-east [2882, 306, 2450, 1479, 755]. south-eastern [1234, 2339, 254, 2839, 2129, 2168, 1446, 1932, 2178, 670, 758, 1860, 2729]. south-west [907, 1108, 2427, 2790, 2567, 814, 701, 2518, 1649, 2825, 2971, 617, 2408, 2371, 30]. south-westerly [175]. south-western [1882, 2057, 378, 1935, 2844, 567, 2389, 1760, 1577, 1627, 2304, 700, 2128, 129, 300, 1530, 2357, 2554, 660, 2016, 2099, 2757]. Southeast [1920]. Southern [1209, 1783, 1385, 657, 2034, 1438, 2024, 1164, 2031, 1998, 1907, 2319, 1022, 918, 1558, 1308, 1101, 1784, 340, 575, 651, 1089, 370, 2700, 1052, 1057, 553, 426, 2744, 1725, 401, 3047, 198, 1145, 3060, 1886, 211, 1449, 2134, 2404, 824, 935, 363, 2418, 2109, 593, 1576, 159, 1055, 627, 1607, 733, 384]. southernmost [2724, 1304]. southward [1998, 346]. southwestern [2542]. soybean [2477]. sp [2530, 1088, 1681, 2076, 813, 2414, 2215, 556, 2212, 1004, 1988, 2504, 2425, 2871, 2241, 217]. space [318]. Spain [484, 1549, 786]. spandrels [2158]. spangled [340, 96]. Spanish [1643, 1213]. spanning [2441]. spans [1940]. sparid [47, 1029]. Sparidae [1871, 334, 1161, 2679, 1274, 938, 218, 283, 983]. sparids [427]. Sparkling [284]. Sparus [797, 608, 2868, 1313, 452, 44, 1843, 2445]. spate [1192]. spathula [2203, 1082, 1690]. Spatial [1741, 2017, 1353, 2027, 2262, 2011, 899, 341, 1233, 370, 1107, 2463, 1274, 1617, 1100, 2798, 2021, 1451, 416, 680, 249, 2702, 364, 1666, 1580, 1356, 3001, 819, 2142, 663, 1431, 651, 2008, 451, 1021, 1476, 729, 1917, 2902, 1058, 456, 1913, 2216, 239, 2014, 2948, 1639, 501, 1342, 1050, 345, 1526, 1657]. Spatial-temporal [680, 1476, 1917]. Spatio [1668, 1658, 2178, 2478, 811, 2036]. Spatio-temporal [1668, 1658, 2178, 2478, 811, 2036]. spatiotemporal [1311, 2952, 960]. spawn [448]. spawned [2557]. spawners [2326, 2283]. Spawning [881, 2818, 2828, 260, 2800, 366, 1969, 2257, 1099, 1289, 2716, 1576, 863, 961, 2346, 1170, 2791, 1739, 1488, 2717, 245, 1344, 2997, 2362, 166, 2398, 2024, 93, 1590, 2681, 1356, 63, 835, 962, 341, 343, 200, 337, 449, 1735, 1408, 2076, 2751, 1680, 1731, 2081, 960, 2168, 1065, 1363, 3057, 1461, 2290, 352, 1240, 2409, 348, 2642, 3069, 1449, 2361, 2831, 515, 481, 1722, 2670, 597, 1903, 1247, 1967, 60, 283, 1090, 1395, 197, 223, 1349, 2478, 319]. spawning-related [2168]. spawning-site [835]. Spawning-stock [2716]. Spawning-strategy-dependent [2800]. spearfish [954, 2268]. spearfishing [2866]. Special [1974, 2658, 1156]. specialisations [2960]. Specialist [1075, 1959, 381, 454, 1921]. specialization [2544, 2653, 1921]. speciation [2129, 1055, 664]. specie [2806]. Species [1556, 1138, 2382, 1600, 776, 1174, 3000, 3034, 778, 1139, 2647, 375, 3006, 928, 2577, 2002, 2619, 2617, 881, 998, 2000, 1096, 1086, 2782, 897, 2183, 2673, 954, 777, 1235, 1773, 450, 62, 271, 1294, 2997, 1438, 2427, 2123, 2569, 1481, 2282, 1315, 2558, 1203, 2239, 1751, 2395, 2339, 2185, 2509, 373, 98, 2502, 100, 645, 2023, 268, 2056, 69, 2874, 1046, 2662, 1304, 2089, 1846, 1088, 616, 2093, 1141, 1892, 1163, 1414, 2386, 2514, 1109, 1679, 2428, 2242, 1168, 2426, 2587, 2766, 2051, 2843, 1540, 1308, 257, 222, 1357, 1517, 2291, 1703, 61]. species [2621, 1008, 991, 2126, 1348, 849, 915, 2186, 246, 644, 1685, 329, 2941, 2627, 1923, 386, 1644, 17, 1161, 1111, 773, 2298, 1059, 2771, 1543, 2935, 1337, 3039, 2740, 347, 822, 1951, 2267, 2228, 1435, 366, 1091, 1904, 2505, 1416, 1477, 1884, 676, 1646, 2953, 2136, 1081, 937, 342, 1546, 1668, 512, 966, 2544, 1151, 150, 1115, 2608, 1442, 2549, 619, 1536, 1909, 2796, 2971, 557, 248, 2004, 1250, 3056, 103, 1891, 757, 2376, 830, 2003, 3, 731, 241, 243, 1651, 1538, 2546, 2974, 2357, 2797, 1796, 2396, 1393, 2666, 716, 2554, 420, 348, 2108, 2887, 1132]. species [2400, 2411, 3033, 1991, 2540, 2905, 88, 240, 2294, 577, 36, 291, 2447, 3005, 1637, 2263, 2648, 2410, 537, 591, 2947, 2040, 2404, 2606, 972, 798, 2867, 481, 824, 2377, 2419, 615, 2755, 1347, 1122, 1110, 2749, 384, 217, 1391, 2133, 1906, 2674, 1854, 1309, 2139, 1841, 203, 2956, 1938, 2542, 2851, 1346, 1537, 1990, 2324, 2743, 2073, 310, 2789, 1487, 2565, 1547, 162, 2886, 595, 1653, 2431, 2397, 2645, 1103, 2138, 3054, 2358, 733, 1383, 2422, 3080, 2969, 470, 2057, 1046, 983]. Species-specific [1556, 1138, 777, 1442]. specific [2078, 1556, 1310, 777, 1344, 1809, 1138, 1824, 216, 1178, 2991, 2048, 21, 1731, 1694, 2254, 818, 1442, 3026, 2849, 2112, 2080, 2642, 1262, 236, 2009, 959]. specificity [1944]. specimen [432]. specimens [522, 1477]. speciose [1180]. speckled [1733, 578]. Spectracanthicus [2509]. spectral [2931, 653, 2928]. speculation [862]. speed [820, 1062, 3003, 57, 469, 2310, 2547]. speeds [1491]. spelling [1360]. Sperm [1857, 2849, 1713, 1645, 3035, 1450, 2072, 2566, 118, 845, 2561, 2642, 1066, 190]. sperm-duct [118]. spermary [1520]. spermatozoa [1488, 11, 2265]. spermatozoon [2837]. Sphoeroides [2808]. Sphyraenidae [822, 1422]. Sphyrna [2783, 2385, 2057, 3029, 1587, 2873, 2129, 1475, 1897, 3074, 2040]. Sphyrnidae [519]. Spicara [386]. Spikefish [2865]. spinax [1612]. spine [1545, 1010]. spined [244, 210, 678, 1163, 583, 2612, 1684, 1493, 1925, 682, 1366, 649, 2384, 947, 1049, 544, 2351, 2850, 892, 1119, 514, 2995]. spinefoot [623, 2381, 2443]. spines [760, 2990, 1305, 2302, 1419]. spiny [792, 760, 81, 1935, 665]. spinyhead [2884]. splendens [2839, 925, 167, 1334, 2913]. splitfin [411]. Spondyliosoma [701]. Spontaneous [2167, 3035, 1574]. spoons [1562]. spot [2494, 1489, 232, 2466]. spotback [814]. spots [486, 207]. spottail [729]. spotted [782, 2643, 2847, 561, 793, 2226, 2219, 1861, 2336, 3078, 623, 1061, 941, 1067, 1312, 115, 1581, 1596, 3045, 2381, 843, 1200, 2601, 1446, 265, 1887, 1793, 834, 1000, 2721, 1648]. spottedtail [215]. spp [2211, 697, 1666, 373, 1706, 258, 1834, 2551, 1228, 3022, 1880, 1035, 1410, 2639, 2190, 2448, 1408, 3030, 1048, 1214, 607, 259, 1145, 2554, 688, 240, 862, 218, 646, 2780, 1497, 1903, 2252, 2109, 870, 2940, 1026]. spp. [1591, 1123]. sprat [1588]. Sprattus [1588]. spread [2638, 691]. spreads [858]. spring [203]. Springs [1753]. spurdog [1614, 2425, 595]. Squalidae [1614, 2425, 595]. squaliform [94]. Squaliformes [2425, 1234, 1785, 1612, 1614, 595]. Squalius [2955, 43]. Squalus [792, 760, 1935, 665, 1614, 752, 2425, 595]. squamation [1377]. squamosissimus [2364, 816]. squaretail [1087]. Squatina [123, 1454, 1360, 1577, 1908, 2418]. Squatinidae [1577, 1908]. Squatiniformes [1577]. Squid [2337, 1222]. Sr [700, 426, 865, 619, 1136]. SSP [1529]. SSP-9 [1529]. SSR [380]. St [1052, 1057, 382, 863, 424]. St. [2752]. stability [782, 2868, 2084]. Stable [769, 2910, 768, 689, 578, 2721, 2508, 2203, 166, 1138, 1623, 975, 770, 1112, 1852, 518, 841, 2048, 2379, 2741, 817, 1028, 2071, 2374, 2018, 1677, 2870, 665, 1898, 2283, 884, 1277, 1307, 1451, 3089, 3009, 994, 993]. Stable-isotope [769, 578, 2721, 1623, 770, 1112, 518, 841, 2379, 2741, 2374, 1677, 2870, 884, 1277, 3089]. stable-isotopic [2203]. stacking [1252]. stage [256, 1855, 2137, 2187, 1120, 2825, 278, 2403, 90, 2258, 2216, 2947, 203, 869, 1102, 1197]. stages [2832, 2482, 1168, 1672, 1795, 574, 134, 2066, 2678, 2753, 987, 2329, 2237, 3054]. staghorn [2331]. staging [1750]. stakeholders [1226]. Standard [2807, 1369, 1822, 510, 978, 1829]. standardized [1915, 2192]. standardizing [3066]. standing [2463]. starry [2042, 294, 1812, 1390]. start [1607]. startle [1384, 1780]. startled [829]. starvation [963, 597]. State [567, 250, 982, 580, 3059, 2547, 916, 3080]. station [3052]. Statistical [980, 1079, 414]. statistics [2198]. Status [463, 487, 506, 526, 289, 441, 2783, 881, 745, 533, 43, 2991, 748, 337, 1902, 329, 435, 296, 2348, 2164, 1999, 927, 607, 1815, 415, 723, 1575, 935, 2914, 217, 2912, 155, 741, 154, 1171, 422, 2685, 2793]. status-dependent [415]. stay [2467]. steady [580, 1482]. stealth [2270]. Steelhead [1712, 1283, 106, 1911, 1433, 1572, 2695, 2697, 2085]. Stegastes [1639, 2992]. Stegostoma [626]. steindachner [2514]. stellatus [1390]. Stellifer [1613]. stem [266, 1459, 2175]. stem-like [266]. stem-progenitor [2175]. stenolepis [2151]. step [664, 1106]. Stephanoberycoidei [1530]. stepwise [596]. stereo [2044]. stereo-video [2044]. stereological [238]. sterlet [2888, 968]. Sternarchogiton [1202]. steroid [922, 2850, 9, 2307]. steroidogenesis [727]. steroidogenic [2355]. steroids [164]. Stevardiinae [2000, 2621, 2003, 2324, 2743]. Stevens [739]. stew [1300]. Stickleback [550, 244, 210, 678, 1163, 583, 2612, 1684, 1493, 682, 1366, 649, 2384, 947, 1049, 2351, 2850, 2669, 1119, 2858]. sticklebacks [1925, 544, 892, 514, 2995]. Stickney [108]. Stigler [980]. Stigmatopora [720, 71, 501]. stimpsoni [888]. stimulating [2643]. stimulation [408, 1638]. stimuli [447, 2933, 2923, 2936]. stimulus [293]. sting [1947, 1067]. stingaree [758]. stinging [2321]. stingray [2624, 1602, 2428, 2853, 1132, 211, 2969]. stingrays [2034, 2856, 756]. stochastic [1223, 656]. Stock [1557, 430, 2914, 959, 2121, 727, 2140, 1845, 908, 2828, 1732, 687, 1147, 395, 260, 1685, 1441, 35, 1352, 933, 1186, 1886, 2148, 1198, 2716, 752, 2157, 199, 2685]. stock-enhancement [260]. Stock-specific [959]. stocked [2741, 1376, 1038, 1290]. stocking [1556, 354, 2777, 1559, 1564, 838, 949, 379, 180, 2989, 538, 3024, 475]. Stocks [1193, 438, 797, 1580, 1420, 2949, 1549, 1449, 2913, 622]. stoichiometry [1138]. stomach [3067, 453, 2340, 854, 817, 3066, 1028, 1375]. Stomiidae [1765]. Stomolophus [2682]. stone [931, 2851, 188, 1390]. stop [2258]. storage [1645, 2189, 879, 2436, 467, 2561, 2265, 2462, 190]. stored [58, 2125, 1333]. stores [611, 1902]. storing [620]. storms [3073]. story [1463]. stoutii [679, 2830, 3079]. Straight [504]. Straight-nosed [504]. strain [2221, 1106, 2512]. strains [662, 932, 2159, 1608, 2053, 1026, 1513]. Strait [200]. stranded [2944, 2836, 826]. stranding [114]. strandings [2812]. Strangomera [1186]. strategies [1013, 2000, 1227, 1852, 1388, 1069, 116, 1331, 1288, 2458, 1637, 798, 203, 173]. strategy [2346, 1278, 866, 1421, 2091, 2141, 1634, 2800, 2649, 1626, 648, 900, 1671, 454, 1249]. stratification [1363, 433]. Stream [806, 2725, 1751, 1478, 3048, 2551, 302, 2186, 2076, 682, 579, 1977, 321, 1498, 2546, 417, 2732, 2470, 1073, 965, 348, 1572, 1500, 1980, 1979, 1865, 2653, 2881, 2858, 1116, 41, 2940, 1103, 165]. stream-dwelling [2470, 965, 1865, 2653]. stream-maturing [1572]. stream-reared [1498]. stream-resident [682, 2858]. streambed [1911]. streamflow [2168]. streams [3027, 2183, 2142, 2527, 2747, 168, 1107, 213, 2005, 1566, 2352, 1159]. streamscape [2352]. strength [1037, 1535]. strengthen [2898]. Streptococcus [2788, 1608, 1026]. Stress [2369, 926, 702, 2665, 662, 1254, 301, 145, 1606, 1684, 922, 3012, 2219, 2846, 1642, 3020, 958, 2964, 1963, 2862, 92, 837, 36, 544, 2354, 2863, 3024, 2661, 32, 2706, 959, 1794, 539, 1719]. stress-coping [2846]. stress-induced [3012]. Stressed [572, 2354, 3028]. stressful [922]. Stressor [1847]. Stressors [5, 2753]. stretch [1666]. stretches [2272]. striata [1946]. striatus [1235, 1534]. strictly [1122]. stricto [964]. strigata [914]. striped [1952, 2483, 1731, 960, 1352, 346, 1748, 1946]. strobe [2934]. Strong [716, 2077]. Stronger [2776, 2496]. strongly [116]. strontium [2489]. Structural [2973, 2238, 1528, 2330, 2810, 2967]. Structure [545, 2256, 2963, 39, 2302, 779, 2121, 306, 582, 598, 576, 1632, 675, 2421, 2526, 396, 1384, 1076, 3025, 2625, 2074, 701, 2828, 2151, 2873, 2129, 1196, 2593, 2737, 2154, 1577, 2853, 2868, 1810, 337, 969, 3083, 2130, 250, 2008, 2736, 347, 148, 2136, 2055, 500, 264, 1048, 2071, 14, 799, 788, 2574, 1936, 430, 2879, 1058, 2805, 270, 716, 379, 2821, 1446, 2028, 78, 536, 2713, 1714, 1100, 660, 211, 421, 2479, 2870, 3034, 538, 1201, 363, 2780, 502, 215, 2305, 1447, 2490, 1946, 1912, 1307, 952, 2878, 1027, 622, 2835, 272, 2253, 2153, 42, 968]. structured [2972]. Structures [801, 713, 23, 2781, 1297, 2695, 2995]. structuring [1849, 350, 370, 2515, 331]. strumosus [32]. student [1829]. studies [8, 3067, 805, 2140, 1804, 1949, 1339, 1443, 1743, 2126, 2735, 1762, 329, 726, 1423, 1131, 1004, 762, 2204, 890]. study [1415, 782, 210, 1705, 1479, 1056, 1718, 1833, 572, 2465, 2013, 1880, 917, 1410, 3086, 472, 1571, 1642, 949, 592, 1134, 1693, 1596, 1891, 2112, 2601, 1893, 529, 2905, 2648, 2947, 2230, 2098, 1946, 1277, 499, 428, 2106, 1719]. studying [1869, 1317, 2381, 2115]. stumpnose [721]. stunting [33]. Sturgeon [2201, 471, 2750, 2786, 2609, 2622, 2119, 610, 687, 2510, 2059, 2061, 835, 1062, 2614, 1848, 2545, 2174, 1853, 1355, 2205, 1037, 2753, 2575, 650, 2442, 2714, 877, 2715, 2752, 1654, 2237, 2890, 42, 1690, 968, 1664, 2889]. Sturgeons [438, 2764, 373, 279]. sturio [2061]. Sturisoma [2549]. Stygichthys [132]. style [1881]. styles [2846]. sub [2356, 1421, 809, 1345, 2242, 350, 474, 2889, 346, 1007, 2478]. sub-adult [2889, 346]. sub-Antarctic [2356, 2478]. sub-sampling [1421]. sub-species [2242]. sub-structuring [350]. sub-tidal [1345]. sub-tropical [809, 474, 1007]. subantarctic [2932]. subarctic [192]. subfamily [552]. subgenus [1716]. subject [2353, 716]. Sublethal [2297]. submerged [1994, 721]. submersible [1630]. Submissions [528]. Suboptimal [2978]. subpolar [221]. subsea [876]. subsequent [68, 1290]. subsidence [1779]. subsidy [258]. subspecies [1010]. substances [2938]. substantial [2405]. substrata [366, 602]. Substrate [1944, 2609, 1911]. Substratum [288, 1787, 2174, 1090, 1395]. substructure [1143]. Subsurface [1805]. subterranean [132]. Subtle [2151]. subtractive [216, 212]. subtropical [2025, 1937, 2256, 2018, 3070, 2666, 2188, 2036]. subunit [2321, 214]. subunits [2943]. Success [1777, 2792, 249, 355, 1354, 2846, 873, 546, 455, 2155, 315, 1626, 180, 1038, 481, 2957, 597, 1119, 228, 223, 126]. Successful [522, 448, 1093, 2751, 2600]. successfully [3078]. succession [3000, 2352]. sucetta [3083]. such [2960]. sucker [226, 2055]. suckermouth [2563]. suckers [53]. suckleyi [752]. Suction [1408]. Sufflogobius [2534, 1370, 1043]. suggest [670, 2355]. suggested [2108]. suggests [365, 2838]. suitable [2322, 1958]. suite [577]. Sulawesian [1139]. summary [2693]. Summer [2272, 195, 898, 2655, 1260, 705, 203, 165]. sun [1790]. sunfish [1257, 2441, 21, 1586, 994, 1028, 993, 2337, 2240, 689, 2860]. sunfishes [2803, 1790]. sunlight [1337]. superficial [2086, 1482]. superimposition [348]. superior [2512]. supersaturation [106]. supper [3066]. supplementation [1488, 1581, 2704, 478, 1282]. supply [1469, 318, 936, 1230]. support [361, 2844, 1228, 2559, 147, 2960, 1991, 1223, 2377]. supports [607, 217]. suppress [1376]. suppresses [1574]. suppression [216, 212]. Suprelorin [2819]. suratensis [694, 160]. surf [1381]. Surface [1484, 555, 963, 1310, 2086, 1693, 2648, 425]. surface-associated [2648]. surfacing [2545]. surfperch [552, 1574]. surgeonfish [93, 2599, 2713]. surgeonfishes [2748]. Surgical [602]. surgically [853]. surmuletus [1952, 1352]. surrounding [474]. surveillance [942]. survey [2880, 1008, 775, 2292, 1214, 3005, 1229, 555]. surveys [1415, 2330, 2044, 2648, 2755]. Survival [2891, 1502, 1285, 3031, 516, 2607, 1795, 2524, 2219, 943, 1681, 1735, 2181, 3044, 853, 1957, 1543, 1387, 1835, 1423, 2651, 187, 2489, 455, 562, 1001, 1334, 417, 1789, 1240, 1656, 2062, 236, 1562, 2271, 2259, 317, 1692, 869, 2776, 1074, 165, 2486]. survive [1323]. surviving [945]. survivorship [2804, 1102]. Susan [2094]. susanae [2094]. susceptibility [2788, 2415, 2295, 1195, 1624, 1671, 2589, 79, 2605, 1968]. susceptible [1916]. sustainability [2909, 1776, 2007, 2591, 2904, 2537, 1067, 2906]. Sustainable [831, 2898, 1222, 2918, 2899, 820, 2916, 2897, 294, 1228, 1995, 1223]. Sustained [1326, 1519, 713, 1336]. Suthers [138]. Sutton [1266]. swallower [3090]. swallows [1299]. swamp [2313, 1837, 2815, 2150]. Sweden [683]. Swedish [536, 825]. sweet [2172]. sweetlips [575]. swellshark [2229]. swim [1508, 2613]. swimbladder [580, 1036]. swimmer [2238, 647]. swimmers [1251]. Swimming [2627, 1327, 1386, 647, 2303, 2663, 448, 1798, 309, 2308, 3031, 610, 820, 2224, 1519, 1491, 2592, 1062, 872, 3003, 2836, 2492, 1416, 1836, 57, 1939, 469, 2174, 2310, 1748, 1674, 1853, 1436, 2652, 2947, 515, 787, 2363, 1256, 1607, 1482, 1621, 2562]. swordtail [1181]. sws1 [2222]. Symbiosis [923, 1652]. symbiotically [2054]. Sympatric [336, 1644, 2451, 2502, 2874, 1163, 1679, 2383, 3022, 102, 2513, 2173, 1035, 915, 773, 366, 1668, 14, 300, 1288, 947, 1132, 240, 1525, 1637, 646, 950, 3054]. sympatry [659, 1383]. Symphurus [1467]. Symphysodon [77, 407]. Sympterygia [2048]. synagris [163]. Synaphobranchus [541]. Synbranchus [2815]. sync [2696]. synchrony [421]. synchrotron [1955]. syndrome [819]. syndromes [392, 1169]. Synechogobius [215]. synergistic [6, 1441]. syngnathid [720]. Syngnathidae [1516, 1515, 489, 1314, 2971, 491, 2623, 492, 493, 494, 490]. syngnathids [498]. Syngnathinae [2971]. Syngnathus [3022, 495, 505, 1145, 1919, 191, 1051, 504]. synonym [1318, 2127, 353, 1343]. synonymous [1640]. synonyms [623]. synonymy [1481]. synophthalmia [1454]. synthesis [793, 2552, 189]. synthetic [1436]. system [961, 2708, 2339, 2833, 2825, 2777, 624, 1205, 106, 2927, 296, 500, 958, 1200, 1253, 2723, 433, 2923, 2231, 1960, 1336, 2016, 87, 1637, 1575, 2932, 1819, 1921, 428, 1547, 2715]. Systematic [3022, 2420, 629]. systematics [2494, 1174]. Systematist [1232]. systems [8, 156, 1586, 1933, 589, 30, 3024, 2898].

T [113, 1127, 1174, 139, 389, 1266, 1098, 461]. T.Enderby [1371]. tabasco [2880]. Tachysurus [2083, 2481, 2402, 2548]. Tactic [2254, 1025, 2849]. Tactic-specific [2254, 2849]. tactics [2388, 2182, 2720, 1069, 2155, 1288]. tactile [661]. tadpole [2649]. taeniatus [1022]. taeniolatus [497]. Tag [2626, 2042, 1808, 2340, 2181, 853, 2275, 196, 2680, 1134]. tag-recapture [2042, 196]. tagged [2626, 602, 2080, 1458]. Tagging [1211, 685, 766, 2042, 282, 498, 879, 1811, 1684, 1493, 1544, 3037, 2848, 2699, 2814, 878, 1317]. tags [498, 234, 651, 853, 880, 3085, 35, 2583, 1893, 2462, 803, 1346, 1000]. tailed [2599, 1303]. taimen [295, 1459, 1459]. Taiwan [1372, 2910, 81, 2359, 175, 774, 1544, 1307, 669]. Tak1 [1596]. Takahashi [1343]. take [3067, 1224, 2748]. taken [2607, 2045, 2911]. Takifugu [590, 581, 1520]. taking [2786, 1237, 2131]. tale [659]. Tamaulipas [2822]. tambaqui [2867]. Tana [881, 1796]. tang [93]. Tanganyika [147, 415, 1343, 2397]. Tanichthys [367]. tank [2529]. tankei [2723]. tannery [2767]. Tapajós [2619, 2395, 222, 1393, 2887, 2558, 1841]. TaqMan [854]. tarapacana [2779, 1207]. target [589]. targeted [2156, 1008, 88]. targets [1574]. tarpon [2389]. task [1657]. tasks [1639]. Tasmania [306]. taste [1130, 1785, 2942, 2927]. Tatia [2236]. taurine [478]. taurus [3002]. tawny [2908]. taxa [699, 818, 607, 2838]. taxiformis [3063]. Taxonomic [616, 2263, 1766, 2673, 868, 2991, 2247, 607, 1010]. taxonomical [1622]. taxonomy [2494, 1530, 742]. teacher [25]. technique [858, 1956, 1214, 1897, 1220]. techniques [1146, 535, 620, 2115, 1913, 1087, 280, 1375, 1266]. technology [1072, 1211]. tectum [2373]. Teeth [2372, 1305, 2400]. Telemetry [1134, 2303, 69, 2972, 2275, 700, 1603, 232, 2626, 3085, 1830, 1368]. telencephalon [1548]. Teleocichla [2073]. Teleost [439, 2731, 2665, 7, 22, 395, 854, 1205, 2846, 2373, 958, 2381, 848, 1263]. teleosteans [2101]. Teleostei [1174, 881, 2782, 521, 2865, 267, 2420, 385, 2808, 2339, 1304, 616, 304, 2426, 91, 2587, 2592, 1452, 1517, 2291, 1760, 1703, 61, 1880, 711, 1685, 17, 1951, 1081, 1474, 2414, 1151, 2215, 619, 1675, 714, 2837, 1530, 2222, 1659, 2396, 2781, 107, 2400, 371, 2437, 2005, 1064, 2404, 2171, 2504, 1938, 2672, 2542, 2851, 801, 2073, 578, 2789, 1055, 1422, 552, 2645, 912]. teleosts [685, 1765, 10, 5, 1522, 272]. Teles [2395, 1393, 222]. telescope [2427]. Telestes [270]. tell [324, 2298, 724]. Telmatochromis [415]. temminckii [2468]. Temperate [605, 2638, 1870, 1168, 1591, 518, 2941, 2503, 3085, 2610, 3, 224, 1135, 2970, 2417, 720, 88, 2978, 696, 1050, 1599, 203, 1207, 1116, 1518, 1422, 1860, 681]. Temperature [1285, 584, 316, 1789, 4, 1185, 857, 2321, 2528, 1506, 444, 737, 1283, 1294, 533, 1330, 697, 29, 1068, 2665, 545, 3035, 879, 767, 2340, 1914, 934, 1585, 1238, 274, 393, 1365, 2064, 2763, 1747, 893, 845, 115, 455, 649, 1837, 1581, 57, 38, 734, 3057, 2807, 3053, 92, 1082, 936, 2497, 2134, 252, 2485, 2903, 32, 1782, 815, 1990, 2706, 171, 165, 319]. Temperature-dependent [857, 649]. Temperature-induced [1185]. Temperature-mediated [1285]. temperatures [2346, 1386, 2760, 898, 2655, 2214, 1507, 1735, 1062, 2977, 177, 1406, 2132, 905]. Template [672]. Temporal [2702, 364, 119, 1666, 1959, 3011, 2448, 559, 1021, 1531, 1001, 2902, 1913, 1888, 842, 2084, 1526, 1349, 2017, 2643, 1356, 2011, 427, 2142, 2868, 1233, 2008, 1476, 1668, 3045, 911, 772, 1917, 1107, 456, 811, 1617, 1079, 2216, 2036, 1658, 239, 995, 2178, 1029, 501, 1050, 633, 164, 2478, 680]. temporarily [582]. ten [3056]. tench [20]. tension [435, 67]. tensions [905]. Tenualosa [3084, 2943]. tenue [3077]. tenuis [2670]. terapontid [55]. Terapontidae [55]. terapontids [377]. teratologies [1401]. term [961, 1506, 2326, 631, 1645, 282, 11, 2662, 301, 2607, 2735, 1510, 1238, 1441, 1961, 1091, 1366, 3085, 187, 2489, 1331, 313, 1936, 1073, 2038, 1424, 2652, 1223, 2466, 2053, 1346, 1505, 362]. terminal [2309, 1244]. terraenovae [1592]. terrestrial [1330, 1326, 303, 3023, 2944]. terrestrially [597]. territorial [417, 415, 2787]. Territoriality [1922]. territory [2394, 2874]. terror [2175]. tessellated [1091]. test [1591, 1008, 1966, 1715, 2080, 2462]. tested [2796]. testes [1944, 118, 292]. testicular [2888]. Testing [6, 1651, 1683, 1890, 2274, 2420, 2524, 1067, 3039, 1079]. testis [2643]. testosterone [6]. tests [2098]. testudineus [2664]. tetra [2950, 2060, 2566, 2004, 1987, 999, 1648]. tetra-repeat [1648]. Tetracycline [2125]. tetradactylum [430, 2328]. tetradecylthioacetic [2117]. Tetragonopterus [1110]. Tetraodontidae [484, 2728]. Tetraodontiformes [2865, 484, 2808]. tetraploid [336]. Tetrapturus [954, 2268]. tetrazona [813]. texana [547]. Text [108]. Thai [2588]. Thailand [823, 217].

Thalassenchelys [2127]. Their [208, 549, 138, 111, 694, 256, 1665, 954, 1899, 263, 11, 662, 253, 2242, 1540, 2788, 2141, 1238, 2190, 651, 732, 1408, 3078, 351, 1966, 1867, 451, 2267, 2505, 1295, 426, 790, 467, 2484, 2946, 948, 346, 1104, 2535, 973, 1323, 1638, 1549, 218, 778, 705, 1247, 618, 41, 1080, 173, 1026, 1291, 1876, 1664]. Theme [1324]. theory [966]. there [1976, 161, 1117, 1480, 315, 421, 406, 2694, 1653]. Thermal [1642, 1583, 1363, 2271, 1277, 2726, 2818, 40, 702, 1764, 1414, 1858, 1959, 3020, 313, 433, 92, 259, 2797, 2970, 2471, 2806, 3075, 1347, 3032, 815, 3028, 1074]. thermal-habitat [1959]. thermally [3028]. thermophilic [1422]. thermoregulation [162]. thermotaxis [372]. These [3049]. thiamine [819, 422]. thing [2227, 2842]. things [2974]. thinking [2898]. Thiony [2039]. Third [2649]. Thorichthys [666]. thornback [900]. thorny [1907, 2319]. threadfin [2348, 92, 430, 848, 1447]. threat [1362, 1153]. threatened [3050, 2494, 1183, 2747, 1565, 2036, 1888, 365, 2755, 237]. threats [881, 1920, 632]. Three [1569, 1961, 1955, 2376, 244, 1086, 1618, 210, 2057, 1421, 1481, 3050, 2056, 662, 2662, 678, 258, 1163, 583, 2137, 2013, 2612, 1684, 2513, 1493, 1667, 1161, 773, 1925, 1582, 1435, 366, 682, 1366, 649, 512, 966, 619, 241, 2384, 2805, 947, 1049, 18, 1633, 88, 240, 36, 544, 2351, 3034, 2040, 2850, 892, 2555, 193, 477, 1119, 1903, 2858, 131, 1346, 1576, 1780, 514, 2995, 2582]. Three-dimensional [1955]. three-phase [2582]. three-spined [244, 210, 678, 1163, 583, 2612, 1684, 1493, 1925, 682, 1366, 649, 2384, 947, 1049, 544, 2351, 2850, 892, 1119, 514, 2995]. threespine [2858]. thresher [137, 1587]. threshold [1779]. thresholds [1930, 3026, 1850]. thrombocytes [1338]. throughout [2702, 2398, 2970, 2347]. throughput [1131]. thrust [647]. Thunnus [238, 364, 1590, 482, 290, 194, 807, 3007, 2340, 1105, 134, 553, 1419, 408, 1009, 1176, 618]. Thymallinae [19]. Thymallus [196, 1603, 845, 1185, 41, 2329, 1027, 19]. thymus [1647]. thynnus [238, 364, 1590, 482, 2340, 1419]. thyroid [2399, 1931]. thyroid-hormone [1931]. thyroxine [1512]. tiburo [2783, 2129, 1475]. Tidal [2096, 2745, 2023, 1429, 2662, 1345, 2013, 28, 383, 233]. Tidal-diel [2745]. tidepool [659]. tidepools [2028]. tiger [977, 2372, 1892, 1740, 813, 2723, 3002, 1100]. tigerfish [1299]. tightrope [629]. tilapia [40, 2880, 855, 2892, 574, 2788, 1337, 2302, 1608, 172, 288, 798, 997, 730, 172, 798, 1077, 2454]. tilstoni [271, 915, 2165]. Time [1287, 1387, 630, 1449, 1702, 2523, 2696, 2264, 2362, 516, 1871, 226, 2964, 2951, 2322, 2516, 2595, 318, 1893, 1910, 2265, 171, 1885, 1282, 2486]. Timecourse [2169]. times [1549, 995]. Timing [2600, 276, 835, 1778, 761, 1293, 2700, 2751, 2763, 1363, 346, 1847, 236, 916]. Timothy [508]. Tinbergen [550]. Tinca [20]. Tiny [2147]. Tissue [975, 2829, 1138, 2543, 2048, 1493, 768, 1420, 2671, 2112, 2784, 806, 1986, 2714, 1440, 3009, 1968]. Tissue-based [2829]. tissue-specific [2048, 2112]. tissues [2809, 2379, 2802, 1986]. tlr [826]. tlr1 [2058]. tlr2 [2058]. tlr7 [2589, 1736]. toadfish [2182]. tobianus [227]. Tocantins [1773, 2282]. tolerance [2880, 429, 1764, 1168, 674, 278, 1238, 1837, 1034, 2707, 92, 1748, 2065, 1347, 1828, 815, 1006, 1290, 2512]. tolerances [2806]. Toll [146, 2058, 1597, 2213, 1638]. Toll-like [146, 2058, 1597, 2213, 1638]. Tometes [2002]. tomography [759, 2599, 1955]. tompot [1922]. tomtates [1489]. Tone [2741]. tones [944]. tongue [2826, 82, 380, 1568, 2476, 1931]. tonguefish [1467]. Too [448]. Tool [1870, 2979, 697, 1869, 1339, 3051, 1067, 865, 1107, 2982, 2601, 1886, 1157, 2015]. toolbox [2533]. tools [1766, 1620, 1314, 1516, 1317, 1515, 2557]. Tooth [1397, 2416, 541, 2460]. toothcarp [1125]. toothfish [2356, 537, 2478]. toothless [2772]. top [1882, 1915, 708, 1709]. tope [2790, 1737, 432, 1811]. topminnow [1540, 1579]. Tor [2114]. torafugu [590]. Torpedo [793]. totoaba [2375, 2375]. tourism [2410, 3015]. toxic [2894, 247]. toxicants [987]. toxicity [1325]. Toxicology [1265, 694, 2295]. toxin [97]. trace [675, 962]. Trachinocephalus [2051]. Trachinotus [2295, 2605, 1968]. Trachurus [663, 3014, 281, 2879, 345]. Tracing [2987]. track [882]. Tracking [1038, 1780, 766, 69]. tract [2959, 2620, 1539, 833, 2476]. tracts [1303]. Trade [2954, 1945, 1863, 2904, 1221, 1835, 775, 2004, 2905, 203, 1439, 237, 548]. Trade-offs [2954, 1945, 1221, 1835, 203, 1439, 548]. Traditional [955]. trahira [54]. trail [1874, 581]. training [400, 2742, 2827, 2582]. trait [2754, 1112, 2560]. traits [1140, 631, 2217, 2140, 2356, 1857, 699, 1914, 2320, 1835, 207, 2100, 229, 1444, 3045, 996, 1641, 1024, 401, 1411, 1242, 1058, 2970, 938, 3069, 1683, 1761, 2947, 2954, 810, 37, 1422, 188, 2511, 1169, 1002]. trajectories [256, 2576, 38]. Trans [2064, 1965]. trans-Atlantic [1965]. Trans-generational [2064]. Transcriptome [2444, 2664]. transcriptomic [2722]. transcriptomics [2829]. transduction [145, 1263, 1191]. transfection [2415]. transfer [1396, 1514]. transferase [2309]. transferred [1507]. transferrin [1340, 974]. transfers [1441]. transformation [2994, 2693, 2802, 2694]. Transforming [2452]. transgenerational [206]. transgenesis [417]. transgenic [1121, 1764, 2275, 367, 38, 2578, 1237]. transient [2179, 2188]. Transition [955, 1202, 423, 2025, 2421, 2399, 534, 1189, 2810, 276]. transitions [1609, 2890]. translationally [1244]. translocated [1083]. translocation [717, 681]. transmission [1384]. transmitted [560, 1536]. transmitter [3031, 2814]. transmitters [2626, 3085, 2080]. transmontanus [2609, 1853, 2890]. transmural [408]. transpacificus [3086]. transparency [1800]. transparent [813]. transponder [2696, 2848]. transport [1508, 833, 1336, 3079]. transporting [1511]. Trap [1779, 565, 2825]. Trap-shyness [1779]. trapped [942]. trawl [1054, 1228, 2045, 1217, 821, 555]. trawled [2101]. trawling [1054, 170]. treatment [1510, 391, 531, 2850]. treatments [1093]. trees [1994]. trematodes [732, 989]. trends [3049, 174, 1961, 3045, 1913, 630, 629, 1599, 37, 3015]. trevally [3045]. tri [1648]. tri- [1648]. Triacanthodidae [2865]. Triaenodon [2517, 1455]. Triakidae [2790, 1737, 432, 1971]. trials [803]. Triangle [1938]. triannual [757]. tributaries [1618, 2130, 2003]. tributary [1429, 2200]. tricaine [932]. Trichiuridae [2758]. Trichodina [2704]. Trichomycteridae [1304, 1308, 3005]. Trichomycterus [3005]. Trichopodus [1335, 1334]. trichopterus [1335, 1334]. tricuspidata [56]. tridentatus [602]. triggerfish [2506, 2808, 2996]. triggering [2694]. triggers [274]. Triglidae [2646]. trihydroxypregn [189]. trimethylamine [410]. Trindade [1150]. triostegus [3063, 2713]. triploid [336, 511, 127, 292, 1484, 2053, 905, 83]. Triploidy [988, 1537]. Triplophysa [2414, 2215, 2645, 733]. Tripterygion [2318]. Trisopterus [616]. tritiated [1956]. Trophic [1696, 1882, 482, 377, 1022, 1407, 1354, 1084, 1838, 2071, 2374, 1538, 2018, 1774, 1525, 1307, 1730, 998, 2640, 2183, 1438, 1632, 2239, 659, 1478, 52, 770, 1592, 2359, 2013, 1933, 2173, 969, 1977, 772, 2430, 1003, 1255, 2885, 2447, 946, 1277, 959]. trophic-guild [1632]. trophic-niche [2183]. trophically [560, 1536]. Trophodynamics [1783]. tropical [1696, 928, 719, 3084, 2882, 2025, 2464, 809, 2011, 1046, 176, 1613, 55, 1054, 820, 2916, 302, 2013, 1348, 2825, 474, 864, 2590, 2727, 2639, 2379, 1305, 2128, 1378, 232, 729, 1531, 2120, 2423, 1461, 757, 848, 2974, 2022, 660, 2745, 2556, 2014, 1722, 2033, 2021, 1342, 696, 1162, 1007, 2726, 389]. tropicalization [1177]. Tropics [1646, 271]. troubled [572]. Trough [1815]. Trout [672, 1170, 657, 2818, 2760, 727, 611, 1569, 1488, 2405, 85, 1353, 667, 2543, 309, 819, 584, 1093, 1563, 1468, 1023, 2494, 951, 404, 2435, 316, 1409, 2677, 2988, 2079, 1746, 1631, 2142, 1593, 962, 1120, 1236, 1772, 903, 1035, 23, 917, 1585, 127, 2186, 2336, 2440, 1305, 700, 1957, 51, 2747, 880, 817, 1723, 2740, 299, 104, 1131, 2046, 833, 658, 1118, 579, 845, 1836, 2159, 1095, 1911, 460, 2756, 2964, 1891, 14, 2761, 1433, 1847, 1104, 2985, 344, 2311, 2470, 348, 924, 1436, 1572, 693, 1185, 596, 1484, 2989, 1888, 1020, 2471, 2784, 806, 1284, 406]. trout [2084, 2653, 2978, 2881, 935, 1039, 1471, 2703, 331, 1198, 2053, 2525, 2672, 815, 1006, 850, 612, 905, 1967, 83, 1090, 857, 2973, 3028, 165, 1513, 416, 1712, 2512, 480, 1504, 1719]. true [2956]. truncation [732]. trunk [1482]. trutta [658, 657, 1618, 611, 1569, 2491, 2405, 85, 667, 1093, 1563, 1468, 2435, 1022, 316, 2677, 2988, 2079, 1746, 1631, 2142, 962, 1120, 917, 1801, 2186, 2336, 2440, 700, 573, 51, 880, 817, 1406, 1723, 2740, 1131, 1118, 845, 1836, 2159, 460, 14, 2761, 2616, 2470, 348, 924, 1185, 1484, 2989, 1284, 406, 2084, 2653, 2978, 2881, 935, 2332, 331, 1198, 2525, 2672, 1006, 612, 1967, 857, 2973, 416, 1425]. truttae [2704]. tryptophan [1282]. tschudii [2424]. tshawytscha [2760, 1503, 1528, 2296, 3001, 2931, 101, 2254, 1034, 78, 325, 1102]. tube [2871]. tubelip [52]. tubenose [434]. tubeshoulder [2802]. tubifer [2054]. tumour [1244]. tuna [238, 2791, 785, 364, 1590, 482, 290, 3007, 2340, 134, 2141, 553, 799, 2554, 408, 291, 1009, 1176, 618]. tunicate [1239]. tunnel [1976]. turbid [1076, 181]. turbiditaxis [2568]. turbidity [686, 2929, 135, 661, 514, 126]. turbine [2598, 2662]. turbines [3044]. turbot [1710, 652, 1597, 1680, 1901, 2354, 225, 706, 627, 1244]. Turkey [426, 2542, 2851]. turnover [2048, 2112, 2430, 1986]. turns [381]. turrubarensis [724]. tuskfish [232, 2979]. twaite [323, 1136]. Twenty [691]. twice [2242]. twin [997]. Two [1235, 2843, 1760, 623, 1841, 2002, 1386, 238, 2000, 897, 2450, 450, 178, 898, 2239, 1666, 659, 98, 268, 1883, 2874, 3031, 1846, 3048, 2683, 2386, 427, 1916, 992, 1540, 2299, 2441, 1030, 47, 102, 991, 2173, 1035, 917, 2727, 851, 915, 246, 644, 386, 753, 1612, 3036, 177, 3039, 1867, 891, 2827, 2267, 2800, 2515, 1476, 2136, 1668, 966, 2197, 1918, 1195, 948, 300, 243, 2805, 2934, 1727, 1818, 2949, 1893, 1714, 720, 1549, 577, 2479, 2885, 2006, 1637, 1887, 862, 191, 798, 1077, 2180, 2234, 950, 935, 615, 363, 1110, 502, 2305, 1060, 2139]. two [839, 1139, 499, 310, 834, 523, 162, 552, 339, 1621, 3054, 1383, 470]. two-dimensional [2197]. two-headed [2180]. twoline [967]. type [293, 1764, 1798, 991, 1933, 1312, 367, 1867, 2083, 818, 38, 2549, 2004, 2174, 1529, 801, 2835]. types [3048, 851, 768, 948, 18, 1635, 2476]. typhle [495, 505, 504]. typhlops [132, 950]. typology [969]. typus [979, 1611, 2074, 1196, 774, 1544, 2836, 775, 2682, 765, 930, 387, 747, 1019, 1296, 3015, 2812]. tyr [2548]. tyrannus [945]. tyrp1 [2548]. Tyrrhenian [2857].

U [1070, 2492]. U.K [3016, 3017]. U.K. [323, 1226, 636, 876, 896]. U.S. [1932, 555, 1654]. U.S.A. [1503, 1632, 398, 1883, 2129, 2130, 321, 1498, 1446, 1094]. UAV [2648]. ucayalensis [2231]. ugly [2660]. UK [2526, 2906]. Ukrainian [1842]. ultra [2114]. Ultrasonography [2769, 2783]. Ultrasound [1259]. ultrastructural [1769, 642]. Ultrastructure [471, 1338, 2592, 1310, 1706]. ultraviolet [2222]. Ulva [1591]. Umbra [1675, 2419, 1201, 2597, 1924]. Umbridae [1675]. Umbrina [2396, 1924, 2273]. unanticipated [1559]. uncertainties [2673]. uncertainty [2685]. unchanged [962]. uncoupling [1856]. uncovers [2950]. undecimalis [1140, 2019]. under-represented [2654]. understand [987]. Understanding [1554, 2533, 2905, 2522, 3066, 743, 2693, 2538, 1816, 314, 750, 637, 458]. underwater [1109, 1933, 1248, 942, 1805]. undescribed [1057]. undesirable [3043]. undulata [748, 1671]. undulate [748, 1671]. undulatory [647]. undulatus [63, 2112, 237]. UNESCO [1402]. Unexpected [2095, 2972]. unfamiliar [2962]. UNH159 [997]. unicornfishes [1485]. unicornis [1485]. Unilateral [1482]. uninfluential [1103]. Unintentional [1559]. Unique [3059, 2527, 2960]. uniquely [1145]. unisexual [540]. unit [1961, 1473, 2616, 2913, 255]. United [2905, 2898]. units [1143, 1655, 213, 839]. unknown [608]. unmanaged [2901]. unmanned [2648]. Unprecedented [2976]. unravel [1162]. Unravelling [2576, 1953]. unregulated [1593]. unreported [776]. unrestrained [435]. unstructured [365]. untold [1463]. Unusual [2101, 1303, 292, 731, 132, 1537]. unvegetated [483, 2028]. unveils [2817]. Up-and-down [1898]. up-regulated [1597]. update [391, 1150]. Updated [2673, 2991, 2684]. updates [1920]. upland [251]. uploading [2247]. upon [204, 1473, 2554]. Upper [1764, 1414, 1858, 1347, 2558, 1679, 1992, 2621, 1410, 92, 2563, 2437, 2798, 2749, 815, 1547, 3080, 1720, 2495]. upriver [1149]. Upstream [1113, 1192]. uptake [1834, 1826, 693]. Upward [2865]. Upward-Mouth [2865]. upwelling [1255, 1637, 2534]. urbanised [27]. urbanization [1159]. urchin [923]. urchins [735]. urea [2189]. urgent [1995]. Urobatis [2428, 756, 2480, 211]. Urogymnus [2370]. urolophid [2045]. Urolophus [758]. Urophycis [2576]. urotensin [1313]. Urotrygon [757, 456, 2033]. Urotrygonidae [757, 2480, 2033]. Uruguay [1894, 2843]. USA [2752]. usage [2188]. Use [2856, 409, 1911, 1182, 2783, 563, 1528, 234, 659, 1339, 253, 2019, 1870, 2435, 2916, 2897, 2079, 1101, 1348, 1784, 1807, 2379, 651, 2261, 295, 120, 880, 1406, 756, 2198, 1131, 729, 2373, 813, 2292, 1148, 619, 3047, 763, 300, 1699, 391, 433, 2546, 259, 1033, 2934, 2583, 318, 577, 307, 2979, 2462, 3074, 1865, 2419, 850, 883, 1657, 3028]. used [675, 2126, 3058]. useful [699, 1107, 2015, 1547]. uses [1751, 2801]. Using [1283, 1528, 1954, 697, 2597, 1109, 1852, 2465, 1620, 170, 397, 2648, 1671, 325, 987, 2898, 2442, 298, 2334, 1072, 2156, 2696, 582, 608, 2270, 760, 535, 1138, 1203, 234, 2199, 2624, 373, 216, 2958, 2089, 522, 2364, 1345, 2224, 1892, 2044, 1147, 287, 2439, 1120, 2868, 2599, 1956, 90, 2261, 196, 297, 1261, 388, 1378, 1067, 2819, 2827, 1693, 996, 1028, 1419, 808, 1955, 714, 933, 1891, 799, 1913, 2205, 942, 1087, 1829, 1492, 2755, 1805, 1458, 1050, 1768, 255, 656, 1277, 2204, 1650, 1000, 1709]. Ussuri [2402]. ussuriensis [2402]. Utila [1196]. utilities [1556]. utility [961]. Utilization [453, 2615, 258, 483, 1016, 1818]. Utilizing [1375]. UV [47, 1619]. UV-B [47]. uyato [2449, 1622].

V [112, 908, 1268, 525, 1266, 2598, 1933, 28, 948, 1818, 2180, 1828, 1711]. V-H [1711]. Vacchi [2560]. vacuolar [801]. vacuolar-type [801]. vadigo [2488]. Vagelli [812]. Vaillant [2233]. vairone [270]. Valenciennes [3005, 428]. valid [616, 1115]. Validated [1475, 2624]. Validating [339, 1651]. Validation [1367, 1721, 2226, 196, 2266, 206, 720, 1578, 685, 2615, 1629, 2955, 1804, 338, 2089, 1871, 1105, 1801, 297, 2092, 241, 2112]. Value [356, 3077]. values [3089, 3040]. Valuing [2538]. Vanishing [3064, 1302]. var [2774, 3087]. Varden [1524]. Vari [2280]. Variability [2845, 77, 2216, 1154, 2017, 2356, 364, 1311, 1540, 393, 592, 2436, 2611, 1352, 1675, 1071, 1483, 1658, 2798, 30, 1761, 1051, 2775, 1451, 1814]. variabilum [407]. Variable [1543, 538, 698, 1795, 3069, 239, 2886]. variables [712, 400, 1407, 1603, 559, 2927, 2323, 757, 1615, 1473, 1671]. variably [3053]. variance [1748, 223]. Variation [1246, 102, 1015, 1784, 56, 2379, 296, 1148, 460, 2595, 2949, 2806, 2928, 230, 730, 2752, 2754, 657, 2523, 3071, 210, 1713, 2149, 1569, 1285, 1481, 1353, 1751, 81, 1194, 819, 1112, 1601, 427, 294, 1916, 2952, 2142, 3022, 1586, 445, 674, 485, 1233, 370, 376, 168, 50, 1915, 865, 1281, 1831, 965, 2666, 1617, 2158, 1100, 2580, 239, 236, 406, 2948, 2780, 1256, 2166, 2490, 1820, 1302, 542, 2855, 959, 677, 1349, 1439, 600]. Variations [2882, 1665, 281, 2005, 998, 119, 1636, 663, 1107, 1324, 2178, 501, 2139, 1912, 2223]. varies [2182, 1634]. varii [2647]. various [1673]. Varley [25]. vary [2720, 2056, 2544]. varying [202, 135]. vasa [2884, 2584]. vasculature [191]. vasculosus [1710]. vasotocin [1017, 1313]. vaterite [2889, 2201]. Vaz [1150]. vector [1768]. vegetable [2491]. vegetated [2028]. vehicle [2648, 1805]. vehicles [1158, 2491]. velar [545]. velezi [940]. velocity [1889]. velvet [1612]. vendace [381]. Venezuela [245, 2242, 2991]. Venice [1219]. vent [308]. ventilation [829, 2830]. Ventilatory [2325, 2960]. ventricular [1956, 408]. ventriosum [2229]. vera [1712]. verification [245, 1079, 1709]. Verilus [2138]. versicolor [923]. Verspoor [139]. versus [2735]. vertebrae [760, 3029]. vertebral [2211, 1246, 1697, 915, 2436, 2396, 252, 706, 968]. vertebrate [8, 1749, 1854]. Vertical [1623, 1152, 1272, 2407, 2090, 1125, 2681, 879, 726, 394, 1167, 2870, 2363]. very [231]. vessel [2607]. via [2854, 2229]. viability [177, 1847, 2516]. viable [343]. vibrations [2934]. Vibrio [561, 539]. vicariance [1938]. Victoria [298]. victoriae [524]. victorious [412]. video [2854, 1109, 2044, 1933, 953, 942]. videos [2407]. VIE [1684, 1493]. views [2052]. villosus [2717, 3069]. vincentae [2971]. viral [1947, 1195]. virens [2169]. viridis [827]. virology [694]. Virtual [2234]. virus [1597, 2295, 2575, 826, 1968]. viruses [2415]. Visibility [811, 3008]. Visible [1684, 2465, 2125, 1346]. Vision [790, 2373, 2921, 618]. visits [2179]. Visual [2929, 653, 535, 2330, 2883, 546, 1214, 661, 1913, 2936, 2937]. visualizing [970]. visually [1871]. vital [419]. vitamin [2510]. vitellogenesis [727]. vitellogenin [793, 343]. vitreus [3051, 841, 2999, 1615, 2929, 422]. vitro [792, 23, 1313, 833, 2671, 1990, 905]. vittatus [1299, 415]. Vitule [2251]. viviparous [1623, 711, 206, 758, 2769, 1574]. viviparus [805]. vivo [2496, 1114, 2671, 813, 531]. vocal [2182]. Vogel [111]. vol [2569]. volcanic [1133]. Volga [438, 1459, 654]. volgensis [654]. volitans [2338, 1869]. volitional [2398]. Volume [185, 358, 568, 109, 186, 359]. Voluntary [1700]. Vondracek [1098]. voucher [465]. vtg [2059]. Vulnerability [2057, 320, 1219, 2551, 753, 599, 2443]. vulnerable [2518, 1940]. vulpes [253]. vulpinus [137].

W [113, 908, 605, 462, 110, 1372, 439, 389, 714]. wahoo [1913, 2289]. Wales [1183]. Walker [1267]. walkeri [1887]. Walking [629]. Wallacea [1139]. Walleye [841, 3051, 2963, 2999, 2929, 2177, 422]. Walter [606]. ward [756]. warehou [3025]. warm [1386, 756, 2610]. warm-temperate [2610]. warmest [324]. Warming [1799, 1386, 509, 2941, 2977, 1836, 321, 2797, 2806, 3075, 1422]. was [2839, 2846]. wastewater [544]. Watanabe [908]. Water [1023, 1571, 1406, 562, 138, 2760, 444, 1283, 1271, 2203, 2717, 1894, 29, 767, 202, 2677, 574, 922, 674, 1510, 1801, 1735, 2063, 624, 1691, 1889, 1152, 1467, 3057, 1700, 1107, 2738, 1214, 1815, 2409, 1362, 180, 1082, 1217, 1678, 1037, 1800, 1682, 1887, 2354, 2667, 2867, 181, 436, 2390, 2706, 399, 906, 2912, 1709, 165]. water-borne [922]. waterbody [1912]. waterfall [888]. waterfalls [1763]. Waters [605, 356, 1108, 979, 2882, 484, 2122, 1158, 1479, 2427, 398, 1660, 158, 1460, 1600, 1734, 2952, 2389, 2335, 2820, 1542, 1233, 175, 17, 3036, 2008, 880, 729, 2629, 26, 2679, 3045, 1909, 2449, 1999, 1411, 2876, 224, 3060, 160, 2240, 2289, 1812, 3055, 995, 1213, 2177, 1207, 1160, 825, 470, 290, 291]. waterway [78]. wattersi [2403, 1004]. wave [125]. Wavelength [3026]. Wavelength-specific [3026]. weakfish [467]. weakly [991]. weather [544, 2363]. weatherfish [1340, 336, 2271]. web [769, 1414, 284, 2507]. Webster [2736, 2737]. weedy [497]. weight [2303]. Weigmann [2248]. Welcomme [982]. welfare [2529, 2292, 1472, 2369, 173]. well [761, 1966, 607]. wellbeing [3040]. Welsh [2911]. West [1205, 907, 1108, 2121, 484, 2717, 2427, 2790, 2567, 814, 701, 2518, 1649, 1407, 24, 2825, 652, 1293, 2712, 229, 2971, 617, 2408, 2371, 420, 2806, 30, 95, 2033, 325, 215, 283, 1776, 1863, 2154, 1161, 1668, 2387, 2098, 2133]. westerly [175]. Western [605, 2467, 3038, 375, 2241, 1952, 3006, 1704, 2724, 1882, 2057, 964, 1590, 3050, 2483, 378, 1935, 2844, 2581, 1445, 2101, 567, 2758, 2359, 2260, 1475, 2389, 1517, 1760, 1577, 1627, 2859, 644, 2304, 700, 2128, 543, 175, 1978, 2066, 2917, 724, 1467, 2608, 1419, 129, 852, 1459, 1891, 300, 586, 1530, 1473, 1480, 2357, 2879, 2554, 1049, 660, 1297, 2016, 1012, 1948, 2099, 1213, 136, 2757, 1176, 1604, 595, 1951]. westslope [2747]. westward [1998]. wet [55, 1646]. wetlands [2060, 1429, 2726]. Whale [2812, 2647, 979, 1611, 2074, 1196, 774, 1544, 2836, 775, 2682, 2467, 765, 930, 387, 747, 1019, 944, 1296, 3015]. whalefish [906]. whalefishes [1530]. whales [382]. whenever [398]. Where [2944, 2904, 2674]. wherever [398]. whiffiagonis [1206]. whipray [459]. White [115, 961, 2609, 1809, 3025, 457, 522, 2416, 2336, 623, 2712, 1067, 1693, 2381, 2268, 1853, 1572, 842, 1805, 1648, 2890, 1523, 2718, 2965]. white-eye [522]. white-spotted [2336, 623, 1067, 2381, 1648]. whitefish [1849, 29, 1357, 1410, 851, 1117, 2641, 1240, 2553, 1910, 936, 1750]. whitei [920, 1523, 2718, 2965]. whitemouth [35]. whitesnout [1813]. whitespotted [2439, 1854]. whitetip [2517, 1455]. whiting [849, 483, 2515, 1694, 2310, 1948, 2675, 2954, 1782]. Who [2909]. Whole [2137, 571, 2416, 1615]. wide [2156, 1040, 2159, 1958]. widely [2107]. Widespread [239, 928, 2560, 828, 2976, 2739]. Wild [2906, 797, 1556, 2523, 2668, 2532, 2696, 1503, 468, 354, 364, 1857, 2655, 1764, 1798, 1491, 1280, 2345, 2719, 1956, 278, 334, 2320, 2087, 1603, 2741, 367, 1561, 130, 1192, 1895, 2980, 3085, 894, 1915, 2159, 38, 476, 1830, 1512, 1073, 1376, 924, 2479, 1535, 2699, 1562, 2286, 2669, 2967, 810, 1560, 1692, 1946, 1368, 2053, 730, 41, 1006, 1290, 228, 37]. wild-caught [1491, 2980, 1535, 810]. wild-reared [2523]. wild-type [1764, 1798, 367, 38]. wildlife [2410]. Wiley [3016, 3017]. will [1227, 636]. Willamette [1503]. Wilson [112]. wind [451, 1323]. window [2701, 1859]. Winter [2142, 611, 59, 2963, 1959, 880, 1052, 1057, 1382, 847, 1449, 1575, 2355]. winter-specialist [1959]. Wisconsin [321, 284]. wise [1484]. Within [1050, 582, 2742, 964, 1636, 2874, 3090, 732, 1205, 2819, 822, 2744, 2623, 2949, 2885, 637, 2780, 501, 1471, 1180, 1860, 223, 804, 19]. Within-lake [1050]. without [1771, 2891, 2473]. wobbegong [472]. wolf [1306, 1537]. wolf-fish [1306]. wolffish [115]. wonders [1873]. wongratanai [2855]. Woo [1127]. Wood [1265]. Wootton [1405]. Workable [110]. worked [2915]. workflow [2156]. Workshop [2693]. World [1018, 1212, 1270, 139, 980, 982, 1749, 2248, 2193, 324, 2537, 2244, 2315, 1991, 387, 1879, 2249, 2429, 2188, 317, 141, 1873]. Worlds [111]. worm [578]. worst [2530]. worst-case [2530]. wound [602]. wrasse [2116, 63, 902, 1870, 1417, 1195, 3037, 224, 1135, 237]. wrasses [52, 2746]. Writing [1374, 1129].

xavante [1763]. xenogeneic [1574]. xichouensis [2215]. Xingu [2185, 2073, 2624, 2360, 1841]. Xiphioidei [169]. Xiphophorus [468]. Xyrichtys [166, 119]. xyster [1011]. Xystreurys [1193].

Yala [2150]. yam [247]. Yáñez [2500]. Yangtze [2504, 877]. year [444, 2385, 1490, 2301, 1456, 566, 655, 3051, 1030, 652, 1293, 3078, 1571, 1059, 394, 1642, 1260, 949, 1084, 460, 847, 2885, 1575, 1009, 125, 1249, 2775, 1912, 1153, 1290, 434, 126, 362]. year-group [652]. Yearbook [980]. years [1629, 899, 1113, 550, 2448, 984, 2092, 691, 2125, 2796, 862, 2814, 517]. yellow [2469, 563, 93, 2681, 2596, 2729, 1914, 90, 853, 451, 2083, 1143, 996, 2152, 2481, 2574, 2480, 1800, 2618, 736, 1394, 126, 1736, 2548]. yellow-eyed [2574]. yellow-nose [2729]. yellow-stage [90]. yellowfin [3007, 408]. yellowfish [918]. yellownose [1882]. yellowtail [2566, 343, 2178, 593, 1036, 1482]. Yemen [1950]. Yield [831]. yokohamae [2325, 1247]. yolk [819, 2931, 893]. yolk-sac [2931, 893]. you’ll [1497]. young [444, 2385, 1490, 1456, 1666, 59, 2440, 394, 1084, 460, 2885, 1575, 368, 125, 1153, 1290, 2886, 1709, 362]. young-of-the-year [444, 2385, 1490, 1456, 394, 1084, 460, 1575, 125, 1290, 362]. young-of-year [2885]. YOY [1865]. Yukon [2780]. Yunnan [2215].

Zaccone [568, 109]. Zale [1266]. zander [1916]. Zapteryx [3050, 378, 1011]. Zar [2198]. Zealand [1371, 710, 1629, 545, 1542, 327, 1921, 1426, 1172]. Zearaja [1882, 755]. zebra [626, 728, 1022]. Zebrafish [2570, 1702, 2487, 68, 2939, 2813, 1798, 2785, 1443, 1469, 1114, 2415, 2529, 2804, 1251, 2292, 1358, 1838, 1917, 3053, 2131, 1066, 2552, 228, 84, 1958, 142]. zebrao [412]. Zebrasoma [93]. zeehaani [1622]. Zhang [2368]. zijsron [2820, 1686]. Zoarces [805]. Zoarcoidei [2082]. zona [2349]. zonation [2407]. zone [2421, 3025, 2991, 1224, 3047, 615, 125, 1709]. zones [1381, 2205, 339]. ZooBank [2247]. zoogeography [969, 2380]. Zoogoneticus [411]. zooplanktivorous [2610]. zooplankton [342, 2610, 1782]. zosterae [2948]. Zubair [2836]. zygaena [2873]. Zygoparity [2307]. zygotic [1679].

References

Anonymous:2010:RPF

[1] Anonymous. Reproductive physiology of fishes. Journal of Fish Biology, 76(1):1–6, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takemura:2010:EIC

[2] A. Takemura, M. S. Rahman, and Y. J. Park. External and internal controls of lunar-related reproductive rhythms in fishes. Journal of Fish Biology, 76(1):7–26, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Migaud:2010:CKP

[3] H. Migaud, A. Davie, and J. F. Taylor. Current knowledge on the photoneuroendocrine regulation of reproduction in temperate fish species. Journal of Fish Biology, 76(1):27–68, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:2010:TSR

[4] N. W. Pankhurst and H. R. King. Temperature and salmonid reproduction: implications for aquaculture. Journal of Fish Biology, 76(1):69–85, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:2010:SGO

[5] J. F. Leatherland, M. Li, and S. Barkataki. Stressors, glucocorticoids and ovarian function in teleosts. Journal of Fish Biology, 76(1):86–111, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crossin:2010:TSE

[6] G. T. Crossin, S. G. Hinch, S. J. Cooke, D. A. Patterson, A. G. Lotto, G. Van Der Kraak, Y. Zohar, U. Klenke, and A. P. Farrell. Testing the synergistic effects of GnRH and testosterone on the reproductive physiology of pre-adult pink salmon Oncorhynchus gorbuscha. Journal of Fish Biology, 76(1):112–128, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dufour:2010:NCD

[7] S. Dufour, M.-E. Sebert, F.-A. Weltzien, K. Rousseau, and C. Pasqualini. Neuroendocrine control by dopamine of teleost reproduction. Journal of Fish Biology, 76(1):129–160, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akazome:2010:FEI

[8] Y. Akazome, S. Kanda, K. Okubo, and Y. Oka. Functional and evolutionary insights into vertebrate kisspeptin systems from studies of fish brain. Journal of Fish Biology, 76(1):161–182, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2010:BRR

[9] A. P. Scott, J. P. Sumpter, and N. Stacey. Book review: The role of the maturation-inducing steroid, 17,20β-dihydroxypregn-4-en-3-one, in male fishes: a review. Journal of Fish Biology, 76(1):183–224, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2010:RHN

[10] W.-T. Huang and C.-F. Weng. Roles of hepatocyte nuclear factors (HNF) in the regulation of reproduction in teleosts. Journal of Fish Biology, 76 (1):225–239, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cosson:2010:FAF

[11] J. Cosson. Frenetic activation of fish spermatozoa flagella entails shortterm motility, portending their precocious decadence. Journal of Fish Biology, 76(1):240–279, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gravel:2010:CEP

[12] M.-A. Gravel, P. Couture, and S. J. Cooke. Comparative energetics and physiology of parental care in smallmouth bass Micropterus dolomieu across a latitudinal gradient. Journal of Fish Biology, 76(1):280–300, January 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giordano:2010:MAH

[13] D. Giordano, R. Russo, D. Coppola, G. Di Prisco, and C. Verde. Molecular adaptations in haemoglobins of notothenioid fishes. Journal of Fish Biology, 76(2):301–318, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKeown:2010:PSB

[14] N. J. McKeown, R. A. Hynes, R. A. Duguid, A. Ferguson, and P. A. Prodöhl. Phylogeographic structure of brown trout Salmo trutta in Britain and Ireland: glacial refugia, postglacial colonization and origins of sympatric populations. Journal of Fish Biology, 76(2):319–347, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2010:PGI

[15] S.-H. Liu, W.-B. Yeh, and H.-K. Mok. Phylogeny of Glaucosomatidae inferred from molecular evidence. Journal of Fish Biology, 76(2):348–356, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sediva:2010:MPA

[16] A. Šedivá, A. Apostolou, J. Kohout, and J. Bohlen. Molecular phylogeographic analyses of the loach Oxynoemacheilus bureschi reveal postglacial range extensions across the Balkans. Journal of Fish Biology, 76 (2):357–368, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Itoi:2010:DSC

[17] S. Itoi, J. Odaka, K. Yuasa, S. Akeno, A. Nakajima, A. Suenaga, T. Noda, S. Akimoto, T. Myojin, Y. Ikeda, Y. Masuda, N. Takai, K. Yoshihara, and H. Sugita. Distribution and species composition of juvenile and adult scombropids (Teleostei, Scombropidae) in Japanese coastal waters. Journal of Fish Biology, 76(2):369–378, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohkubo:2010:MCB

[18] M. Ohkubo, F. Aranishi, and A. Shimizu. Molecular cloning and brain distribution of three types of gonadotropin-releasing hormone from mummichog Fundulus heteroclitus. Journal of Fish Biology, 76(2):379–394, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yasuike:2010:GTP

[19] M. Yasuike, S. Jantzen, G. A. Cooper, E. Leder, W. S. Davidson, and B. F. Koop. Grayling (Thymallinae) phylogeny within salmonids: complete mitochondrial DNA sequences of Thymallus arcticus and Thymallus thymallus. Journal of Fish Biology, 76(2):395–400, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Presti:2010:PRA

[20] R. Lo Presti, L. Gasco, C. Lisa, I. Zoccarato, and L. Di Stasio. PCR– RFLP analysis of mitochondrial DNA in tench Tinca tinca. Journal of Fish Biology, 76(2):401–407, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2010:NSS

[21] Z. X. Gao, H. P. Wang, H. Yao, L. Tiu, and W. M. Wang. No sexspecific markers detected in bluegill sunfish Lepomis macrochirus by AFLP. Journal of Fish Biology, 76(2):408–414, February 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:2010:MIH

[22] S. L. Edwards, J. C. Weakley, A. W. Diamanduros, and J. B. Claiborne. Molecular identification of Na^{+}–H^{+} exchanger isoforms (NHE2) in the gills of the euryhaline teleost Fundulus heteroclitus. Journal of Fish Biology, 76(2):415–426, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grunow:2010:VEA

[23] B. Grunow, P. Ciba, S. Rakers, M. Klinger, E. Anders, and C. Kruse. In vitro expansion of autonomously contracting, cardiomyogenic structures from rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 76 (2):427–434, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Genner:2010:PIN

[24] M. J. Genner and M. P. Haesler. Pliocene isolation of a north-west Saharan cichlid fish. Journal of Fish Biology, 76(2):435–441, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2010:MVN

[25] Felicity Huntingford. Margaret Varley, née Brown: Pioneering fish physiologist and inspirational teacher. Journal of Fish Biology, 76(3):443–445, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lapointe:2010:MAF

[26] N. W. R. Lapointe, L. D. Corkum, and N. E. Mandrak. Macrohabitat associations of fishes in shallow waters of the Detroit River. Journal of Fish Biology, 76(3):446–466, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Viana:2010:FFI

[27] A. P. Viana, F. Lucena Frédou, T. Frédou, M. F. Torres, and A. O. Bordalo. Fish fauna as an indicator of environmental quality in an urbanised region of the Amazon estuary. Journal of Fish Biology, 76(3):467–486, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Godinho:2010:LVM

[28] W. O. Godinho and T. M. C. Lotufo. Local v. microhabitat influences on the fish fauna of tidal pools in north-east Brazil. Journal of Fish Biology, 76(3):487–501, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cingi:2010:EWT

[29] S. Cingi, M. Keinänen, and P. J. Vuorinen. Elevated water temperature impairs fertilization and embryonic development of whitefish Coregonus lavaretus. Journal of Fish Biology, 76(3):502–521, February 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribeiro:2010:LHV

[30] F. Ribeiro and M. J. Collares-Pereira. Life-history variability of nonnative centrarchids in regulated river systems of the lower River Guadiana drainage (south-west Iberian Peninsula). Journal of Fish Biology, 76 (3):522–537, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Copeland:2010:EFH

[31] T. Copeland, B. R. Murphy, and J. J. Ney. The effects of feeding history and environment on condition, body composition and growth of bluegills Lepomis macrochirus. Journal of Fish Biology, 76(3):538–555, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiffany:2010:RSG

[32] B. N. Tiffany, L. A. Enzor, and W. A. Bennett. Responses of skilletfish Gobiesox strumosus to high temperature and low oxygen stress. Journal of Fish Biology, 76(3):556–563, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chizinski:2010:ISM

[33] C. J. Chizinski, K. L. Pope, G. R. Wilde, and R. E. Strauss. Implications of stunting on morphology of freshwater fishes. Journal of Fish Biology, 76(3):564–579, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee-Jenkins:2010:SFS

[34] S. S. Y. Lee-Jenkins and J.-G. J. Godin. Social familiarity and shoal formation in juvenile fishes. Journal of Fish Biology, 76(3):580–590, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luque:2010:MPB

[35] J. L. Luque, A. S. Cordeiro, and M. E. Oliva. Metazoan parasites as biological tags for stock discrimination of whitemouth croaker Micropogonias furnieri. Journal of Fish Biology, 76(3):591–600, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2010:MCS

[36] T. G. Pottinger. A multivariate comparison of the stress response in three salmonid and three cyprinid species: evidence for inter-family differences. Journal of Fish Biology, 76(3):601–621, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vainikka:2010:DTL

[37] A. Vainikka, I. Kallio-Nyberg, M. Heino, and M.-L. Koljonen. Divergent trends in life-history traits between Atlantic salmon Salmo salar of wild and hatchery origin in the Baltic Sea. Journal of Fish Biology, 76(3): 622–640, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lohmus:2010:ETG

[38] M. Lõhmus, M. Björklund, L. F. Sundström, and R. H. Devlin. Effects of temperature and growth hormone on individual growth trajectories of wild-type and transgenic Coho salmon Oncorhynchus kisutch. Journal of Fish Biology, 76(3):641–654, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jarial:2010:SKC

[39] M. S. Jarial and J. H. Wilkins. Structure of the kidney in the coelacanth Latimeria chalumnae with reference to osmoregulation. Journal of Fish Biology, 76(3):655–668, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azaza:2010:SDE

[40] M. S. Azaza, M. Legendre, M. M. Kraiem, and E. Baras. Size-dependent effects of daily thermal fluctuations on the growth and size heterogeneity of Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 76(3):

669–683, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turek:2010:PRG

[41] J. Turek, T. Randák, P. Horký, V. Žlábek, J. Velíšek, O. Slavík, and R. Hanák. Post-release growth and dispersal of pond and hatcheryreared European grayling Thymallus thymallus compared with their wild conspecifics in a small stream. Journal of Fish Biology, 76(3):684–693, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zelazowska:2010:FSE

[42] M. Żelazowska. Formation and structure of egg envelopes in Russian sturgeon Acipenser gueldenstaedtii (Acipenseriformes: Acipenseridae). Journal of Fish Biology, 76(3):694–706, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collares-Pereira:2010:RHO

[43] M. J. Collares-Pereira and M. M. Coelho. Reconfirming the hybrid origin and generic status of the Iberian cyprinid complex Squalius alburnoides. Journal of Fish Biology, 76(3):707–715, February 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kopecka-Pilarczyk:2010:EEP

[44] J. Kopecka-Pilarczyk. Effect of exposure to polycyclic aromatic hydrocarbons on Na^{+}-K^{+}-ATPase in juvenile gilthead seabream Sparus aurata. Journal of Fish Biology, 76(3):716–722, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vacchi:2010:FRD

[45] M. Vacchi, P. N. Psomadakis, N. Repetto, and M. Würtz. First record of the dog snapper Lutjanus jocu in the Mediterranean Sea. Journal of Fish Biology, 76(3):723–728, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tudorache:2010:FLP

[46] C. Tudorache, R. A. O’Keefe, and T. J. Benfey. Flume length and postexercise impingement affect anaerobic metabolism in brook charr Salvelinus fontinalis. Journal of Fish Biology, 76(3):729–733, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukunishi:2010:EES

[47] Y. Fukunishi, R. Masuda, and Y. Yamashita. Exposure of eggs to solar UV-B leads to reduced hatching rates in two sparid fishes, red sea bream Pagrus major and black sea bream Acanthopagrus schlegeli. Journal of Fish Biology, 76(3):734–741, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Ra

[48] Anonymous. Referees 2009. Journal of Fish Biology, 76(3):742–750, February 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozlan:2010:CKN

[49] R. E. Gozlan, J. R. Britton, I. Cowx, and G. H. Copp. Current knowledge on non-native freshwater fish introductions. Journal of Fish Biology, 76 (4):751–786, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kekalainen:2010:MVP

[50] J. Kekäläinen, J. Kähkönen, V. Kiviniemi, and H. Huuskonen. Morphological variation of perch Perca fluviatilis in humic lakes: the effect of predator density, competition and prey abundance. Journal of Fish Biology, 76(4):787–799, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaensson:2010:ECO

[51] A. Jaensson and K. H. Olsén. Effects of copper on olfactory-mediated endocrine responses and reproductive behaviour in mature male brown trout Salmo trutta parr to conspecific females. Journal of Fish Biology, 76(4):800–817, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cole:2010:CTW

[52] A. J. Cole, M. S. Pratchett, and G. P. Jones. Corallivory in tubelip wrasses: diet, feeding and trophic importance. Journal of Fish Biology, 76(4):818–835, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neely:2010:SRS

[53] B. C. Neely, M. A. Pegg, and G. E. Mestl. Seasonal resource selection by blue suckers Cycleptus elongatus. Journal of Fish Biology, 76(4):836– 851, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes:2010:LOM

[54] M. L. M. Gomes, S. L. P. Matta, V. A. Araújo, G. M. F. Silva, and A. A. Zacaro. Larval ontogeny and morphology of giant trahira Hoplias lacerdae. Journal of Fish Biology, 76(4):852–861, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:2010:RCD

[55] A. M. Davis, B. J. Pusey, D. C. Thorburn, J. L. Dowe, D. L. Morgan, and D. Burrows. Riparian contributions to the diet of terapontid grunters (Pisces: Terapontidae) in wet–dry tropical rivers. Journal of Fish Biology, 76(4):862–879, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gray:2010:VGM

[56] C. A. Gray, M. C. Ives, W. G. Macbeth, and B. W. Kendall. Variation in growth, mortality, length and age compositions of harvested populations of the herbivorous fish Girella tricuspidata. Journal of Fish Biology, 76 (4):880–899, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linlokken:2010:ETR

[57] A. N. Linløkken, E. Bergman, and L. Greenberg. Effect of temperature and roach Rutilus rutilus group size on swimming speed and prey capture rate of perch Perca fluviatilis and R. rutilus. Journal of Fish Biology, 76(4):900–912, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidson:2010:MIA

[58] D. Davidson and C. T. Marshall. Are morphometric indices accurate indicators of stored energy in herring Clupea harengus? Journal of Fish Biology, 76(4):913–929, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraboulet:2010:APE

[59] E. Fraboulet, Y. Lambert, R. Tremblay, and C. Audet. Assessment of paternal effect and physiological cost of metamorphosis on growth of young winter flounder Pseudopleuronectes americanus juveniles in a cold environment. Journal of Fish Biology, 76(4):930–948, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vrieze:2010:IOS

[60] L. A. Vrieze, R. Bjerselius, and P. W. Sorensen. Importance of the olfactory sense to migratory sea lampreys Petromyzon marinus seeking riverine spawning habitat. Journal of Fish Biology, 76(4):949–964, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gante:2010:PRN

[61] H. F. Gante, C. D. Santos, and M. J. Alves. Phylogenetic relationships of the newly described species Chondrostoma olisiponensis (Teleostei: Cyprinidae). Journal of Fish Biology, 76(4):965–974, March 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bohl:2010:EIM

[62] R. J. Bohl, T. B. Henry, and R. J. Strange. Electroshock-induced mortality in freshwater fish embryos increases with embryo diameter: a model based on results from 10 species. Journal of Fish Biology, 76(4):975– 986, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colin:2010:ASH

[63] P. L. Colin. Aggregation and spawning of the humphead wrasse Cheilinus undulatus (Pisces: Labridae): general aspects of spawning behaviour. Journal of Fish Biology, 76(4):987–1007, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lafon:2010:NRO

[64] A. Lafon, E. J. Niklitschek, J. Cornejo-Donoso, and K. González. New records of orange roughy Hoplostethus atlanticus juveniles in Chile. Journal of Fish Biology, 76(4):1008–1014, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolland:2010:RPA

[65] J. D. Bolland, I. G. Cowx, and M. C. Lucas. Retention of panjetapplied alcian blue by cyprinids. Journal of Fish Biology, 76(4):1015– 1018, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doupe:2010:RDF

[66] R. G. Doupé and M. J. Knott. Rapid digestion of fish prey by the highly invasive ‘detritivore’ Oreochromis mossambicus. Journal of Fish Biology, 76(4):1019–1024, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thuy:2010:COT

[67] N. H. Thuy, L. A. Tien, P. N. Tuyet, D. T. T. Huong, N. V. Cong, M. Bayley, T. Wang, and S. Lefevre. Critical oxygen tension increases during digestion in the perch Perca fluviatilis. Journal of Fish Biology, 76(4):1025–1031, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barcellos:2010:CZD

[68] L. J. G. Barcellos, F. Ritter, L. C. Kreutz, and L. Cericato. Can zebrafish Danio rerio learn about predation risk? The effect of a previous experience on the cortisol response in subsequent encounters with a predator. Journal of Fish Biology, 76(4):1032–1038, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cucherousset:2010:DES

[69] J. Cucherousset, J. R. Britton, W. R. C. Beaumont, M. Nyqvist, K. Sievers, and R. E. Gozlan. Determining the effects of species, environmental conditions and tracking method on the detection efficiency of portable PIT telemetry. Journal of Fish Biology, 76(4):1039–1045, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sandblom:2010:CAI

[70] E. Sandblom, A. Gräns, H. Seth, and M. Axelsson. Cholinergic and adrenergic influences on the heart of the African lungfish Protopterus annectens. Journal of Fish Biology, 76(4):1046–1054, March 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2010:OIM

[71] J. E. Smith, P. I. Macreadie, and S. E. Swearer. An osmotic induction method for externally marking saltwater fishes, Stigmatopora argus and Stigmatopora nigra, with calcein. Journal of Fish Biology, 76(4):1055– 1060, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Ea

[72] Anonymous. Erratum. Journal of Fish Biology, 76(4):1061–1063, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMeans:2010:E

[73] B. C. McMeans, J. A. Olin, and G. W. Benz. Erratum. Journal of Fish Biology, 76(4):1064–1065, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2010:E

[74] T. Tomiyama, K. Ebe, G. Kawata, and T. Fujii. Erratum. Journal of Fish Biology, 76(4):1066, March 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oleksiak:2010:GAN

[75] M. F. Oleksiak. Genomic approaches with natural fish populations. Journal of Fish Biology, 76(5):1067–1093, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vicari:2010:SDC

[76] M. R. Vicari, V. Nogaroto, R. B. Noleto, M. M. Cestari, M. B. Cioffi, M. C. Almeida, O. Moreira-Filho, L. A. C. Bertollo, and R. F. Artoni. Satellite DNA and chromosomes in Neotropical fishes: methods, applications and perspectives. Journal of Fish Biology, 76(5):1094–1116, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gross:2010:VRL

[77] M. C. Gross, C. H. Schneider, G. T. Valente, C. Martins, and E. Feldberg. Variability of 18S rDNA locus among Symphysodon fishes: chromosomal rearrangements. Journal of Fish Biology, 76(5):1117–1127, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2010:IHW

[78] J. B. Olsen, T. D. Beacham, M. Wetklo, L. W. Seeb, C. T. Smith, B. G. Flannery, and J. K. Wenburg. The influence of hydrology and waterway distance on population structure of Chinook salmon Oncorhynchus tshawytscha in a large river. Journal of Fish Biology, 76(5):1128–1148, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wei:2010:NHC

[79] Y.-B. Wei, T.-J. Fan, G.-J. Jiang, X.-H. Xu, and A. Sun. A novel heartcell line from brown-marbled grouper Epinephelus fuscoguttatus and its susceptibility to iridovirus. Journal of Fish Biology, 76(5):1149–1158, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergek:2010:PDP

[80] S. Bergek, G. Sundblad, and M. Björklund. Population differentiation in perch Perca fluviatilis: environmental effects on gene flow? Journal of Fish Biology, 76(5):1159–1172, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chiang:2010:MFS

[81] T.-Y. Chiang, H.-D. Lin, K.-T. Shao, and K.-C. Hsu. Multiple factors have shaped the phylogeography of Chinese spiny loach Cobitis sinensis in Taiwan as inferred from mitochondrial DNA variation. Journal of Fish Biology, 76(5):1173–1189, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sha:2010:GAE

[82] Z. Sha, S. Wang, Z. Zhuang, Q. Wang, Q. Wang, P. Li, H. Ding, N. Wang, Z. Liu, and S. Chen. Generation and analysis of 10 000 ESTs from the half-smooth tongue sole Cynoglossus semilaevis and identification of microsatellite and SNP markers. Journal of Fish Biology, 76(5):1190– 1204, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2010:ALN

[83] B. Wang, Y. Liu, X. Chen, and Z. Fan. Amitosis-like nuclear division in erythrocytes of triploid rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 76(5):1205–1211, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2010:EPC

[84] L. Wang, D. Chen, L. Yang, S. Huang, Y. Zhang, and H. Zhang. Expression patterns of the creatine metabolism-related molecules AGAT, GAMT and CT1 in adult zebrafish Danio rerio. Journal of Fish Biology, 76(5):1212–1219, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brutto:2010:GDB

[85] S. Lo Brutto, N. Hristovski, and M. Arculeo. Genetic divergence between morphological forms of brown trout Salmo trutta L. in the Balkan region of Macedonia. Journal of Fish Biology, 76(5):1220–1227, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgado-Santos:2010:SII

[86] M. Morgado-Santos, I. Matos, L. Vicente, and M. J. Collares-Pereira. Scaleprinting: individual identification based on scale patterns. Journal of Fish Biology, 76(5):1228–1232, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reinecke:2010:IEE

[87] M. Reinecke. Influences of the environment on the endocrine and paracrine fish growth hormone–insulin-like growth factor-i system. Journal of Fish Biology, 76(6):1233–1254, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Platell:2010:ISF

[88] M. E. Platell, S. A. Hesp, S. M. Cossington, E. Lek, S. E. Moore, and I. C. Potter. Influence of selected factors on the dietary compositions of three targeted and co-occurring temperate species of reef fishes: implications for food partitioning. Journal of Fish Biology, 76(6):1255–1276, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olin:2010:EFR

[89] M. Olin, M. Vinni, H. Lehtonen, M. Rask, J. Ruuhijärvi, K. Saulamo, and P. Ala-Opas. Environmental factors regulate the effects of roach Rutilus rutilus and pike Esox lucius on perch Perca fluviatilis populations in small boreal forest lakes. Journal of Fish Biology, 76(6):1277–1293, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedger:2010:REM

[90] R. D. Hedger, J. J. Dodson, D. Hatin, F. Caron, and D. Fournier. River and estuary movements of yellow-stage American eels Anguilla rostrata, using a hydrophone array. Journal of Fish Biology, 76(6):1294–1311, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Erisman:2010:RBP

[91] B. E. Erisman, M. T. Craig, and P. A. Hastings. Reproductive biology of the Panama graysby Cephalopholis panamensis (Teleostei: Epinephelidae). Journal of Fish Biology, 76(6):1312–1328, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monirian:2010:EHT

[92] J. Monirian, Z. Sutphin, and C. Myrick. Effects of holding temperature and handling stress on the upper thermal tolerance of threadfin shad Dorosoma petenense. Journal of Fish Biology, 76(6):1329–1342, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bushnell:2010:LSP

[93] M. E. Bushnell, J. T. Claisse, and C. W. Laidley. Lunar and seasonal patterns in fecundity of an indeterminate, multiple-spawning surgeonfish, the yellow tang Zebrasoma flavescens. Journal of Fish Biology, 76(6): 1343–1361, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2010:AMR

[94] W. T. White. Aspects of maturation and reproduction in hexanchiform and squaliform sharks. Journal of Fish Biology, 76(6):1362–1378, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2010:RBM

[95] R. M. Rideout and M. J. Morgan. Relationships between maternal body size, condition and potential fecundity of four north-west Atlantic demersal fishes. Journal of Fish Biology, 76(6):1379–1395, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marriott:2010:MSO

[96] R. J. Marriott, N. D. C. Jarvis, D. J. Adams, A. E. Gallash, J. Norriss, and S. J. Newman. Maturation and sexual ontogeny in the spangled emperor Lethrinus nebulosus. Journal of Fish Biology, 76(6):1396–1414, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghazali:2010:EMP

[97] I. El Ghazali, S. Saqrane, A. P. Carvalho, Y. Ouahid, F. F. Del Campo, V. Vasconcelos, and B. Oudra. Effects of the microcystin profile of a cyanobacterial bloom on growth and toxin accumulation in common carp Cyprinus carpio larvae. Journal of Fish Biology, 76(6):1415–1430, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Close:2010:RGT

[98] P. G. Close, P. M. Davies, and K. Trayler. Recruitment and growth of two small-bodied resident fish species (Gobiidae and Atherinidae) in oligohaline, seasonally open lagoons. Journal of Fish Biology, 76(6):1431– 1453, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeMiguel:2010:OAH

[99] R. De Miguel, E. Pino, A. Ramiro, F. Aranda, J. P. Peña, I. Doadrio, and C. Fernández-Delgado. On the occurrence of Anaecypris hispanica, an extremely endangered Iberian endemism, in the Guadalquivir River basin. Journal of Fish Biology, 76(6):1454–1465, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conway:2010:BEN

[100] K. W. Conway and R. L. Mayden. Balitora eddsi, a new species of hillstream loach (Ostariophysi: Balitoridae) from Nepal. Journal of Fish Biology, 76(6):1466–1473, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garner:2010:DCL

[101] S. R. Garner, B. D. Neff, and M. A. Bernards. Dietary carotenoid levels affect carotenoid and retinoid allocation in female Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 76(6):1474–1490, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garduno-Paz:2010:VSS

[102] M. V. Garduño-Paz, M. Demetriou, and C. E. Adams. Variation in scale shape among alternative sympatric phenotypes of Arctic charr Salvelinus alpinus from two lakes in Scotland. Journal of Fish Biology, 76(6):1491– 1497, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matamoros:2010:NSP

[103] W. A. Matamoros and J. F. Schaefer. A new species of Profundulus (Cyprinodontiformes: Profundulidae) from the Honduran central highlands. Journal of Fish Biology, 76(6):1498–1507, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karvonen:2010:PHR

[104] A. Karvonen, H. Halonen, and O. Seppälä. Priming of host resistance to protect cultured rainbow trout Oncorhynchus mykiss against eye flukes and parasite-induced cataracts. Journal of Fish Biology, 76(6):1508– 1515, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ozen:2010:DCA

[105] O. Ozen, A. Altin, and H. Ayyildiz. The diet of Clinitrachus argentatus (Blennioidei: Clinidae) in the northern Aegean and Marmara Seas. Journal of Fish Biology, 76(6):1516–1519, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2010:MDA

[106] E. L. Johnson, T. S. Clabough, C. C. Caudill, M. L. Keefer, C. A. Peery, and M. C. Richmond. Migration depths of adult steelhead Oncorhynchus mykiss in relation to dissolved gas supersaturation in a regulated river system. Journal of Fish Biology, 76(6):1520–1528, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nowak:2010:FRB

[107] M. Nowak, J. Koščo, W. Popek, and P. Epler. First record of the black bullhead Ameiurus melas (Teleostei: Ictaluridae) in Poland. Journal of Fish Biology, 76(6):1529–1532, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ewart:2010:BRA

[108] John W. Ewart. Book review: Aquaculture: an Introductory Text, 2nd edition — by R. R. Stickney. Journal of Fish Biology, 76(6):1533–1534, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morris:2010:BRF

[109] David J. Morris. Book review: Fish Defences Volume 1: Immunology — Edited by G. Zaccone, J. Meseguer, A. García-Ayala and B. G. Kapoor. Journal of Fish Biology, 76(6):1534–1535, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozlan:2010:BRF

[110] Rodolphe Gozlan. Book review: Fisheries Management: Pandemic Failure, Workable Solutions — by G. Pontecorvo, W. E. Schrank, M. Holliday and D. B. Olson. Journal of Fish Biology, 76(6):1535–1536, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2010:BRG

[111] Ian J. Winfield. Book review: Glimpses of Creatures in Their Physical Worlds — by S. Vogel. Journal of Fish Biology, 76(6):1536–1537, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britz:2010:BRM

[112] Ralf Britz. Book review: Mesozoic Fishes 4: Homology and Phylogeny — Edited by G. Arratia, H.-P. Schultze and M. V. H. Wilson. Journal of Fish Biology, 76(6):1537–1539, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlucci:2010:BRO

[113] Roberto Carlucci. Book review: Oceanic Anglerfishes: Extraordinary Diversity in the Deep Sea — by T. W. Pietsch. Journal of Fish Biology, 76(6):1539–1541, April 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hernandez-Miranda:2010:MFS

[114] E. Hernández-Miranda, R. A. Quiñones, G. Aedo, A. Valenzuela, N. Mermoud, C. Román, and F. Yañez. A major fish stranding caused by a natural hypoxic event in a shallow bay of the eastern South Pacific Ocean. Journal of Fish Biology, 76(7):1543–1564, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lamarre:2010:WMP

[115] S. G. Lamarre, P. U. Blier, W. R. Driedzic, and N. R. Le François. White muscle 20S proteasome activity is negatively correlated to growth rate at low temperature in the spotted wolffish Anarhichas minor. Journal of Fish Biology, 76(7):1565–1575, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2010:CSS

[116] F. A. Huntingford, G. Andrew, S. Mackenzie, D. Morera, S. M. Coyle, M. Pilarczyk, and S. Kadri. Coping strategies in a strongly schooling fish, the common carp Cyprinus carpio. Journal of Fish Biology, 76(7):1576– 1591, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Serrano:2010:SSP

[117] X. Serrano, M. Grosell, and J. E. Serafy. Salinity selection and preference of the grey snapper Lutjanus griseus: field and laboratory observations. Journal of Fish Biology, 76(7):1592–1608, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kvarnemo:2010:ITS

[118] C. Kvarnemo, O. Svensson, and W. Manson. Investment in testes, spermduct glands and lipid reserves differs between male morphs but not between early and late breeding season in Pomatoschistus minutus. Journal of Fish Biology, 76(7):1609–1625, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castriota:2010:TVD

[119] L. Castriota, M. Falautano, M. G. Finoia, D. Campo, M. P. Scarabello, and F. Andaloro. Temporal variations in the diet of pearly razorfish Xyrichtys novacula (Osteichthyes: Labridae). Journal of Fish Biology, 76(7):1626–1639, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hussy:2010:UOM

[120] K. Hüssy, H.-H. Hinrichsen, D. P. Fey, Y. Walther, and A. Velasco. The use of otolith microstructure to estimate age in adult Atlantic cod Gadus morhua. Journal of Fish Biology, 76(7):1640–1654, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2010:RBC

[121] P. Oliveira, F. H. V. Hazin, F. Carvalho, M. Rego, R. Coelho, A. Piercy, and G. Burgess. Reproductive biology of the crocodile shark Pseudocarcharias kamoharai. Journal of Fish Biology, 76(7):1655–1670, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sikkel:2010:SPE

[122] P. C. Sikkel and C. A. Fuller. Shoaling preference and evidence for maintenance of sibling groups by juvenile black perch Embiotoca jacksoni. Journal of Fish Biology, 76(7):1671–1681, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baremore:2010:RAA

[123] I. E. Baremore. Reproductive aspects of the Atlantic angel shark Squatina dumeril. Journal of Fish Biology, 76(7):1682–1695, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gallo:2010:PAS

[124] V. Gallo, M. J. Cavalcanti, R. F. L. Da Silva, H. M. A. Da Silva, and D. Pagnoncelli. Panbiogeographical analysis of the shark genus Rhizoprionodon (Chondrichthyes, Carcharhiniformes, Carcharhinidae). Journal of Fish Biology, 76(7):1696–1713, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoll:2010:EWE

[125] S. Stoll, H. Hofmann, and P. Fischer. Effect of wave exposure dynamics on gut content mass and growth of young-of-the-year fishes in the littoral zone of lakes. Journal of Fish Biology, 76(7):1714–1728, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wellington:2010:ETP

[126] C. G. Wellington, C. M. Mayer, J. M. Bossenbroek, and N. A. Stroh. Effects of turbidity and prey density on the foraging success of age 0 year yellow perch Perca flavescens. Journal of Fish Biology, 76(7):1729– 1741, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Han:2010:CRD

[127] Y. Han, M. Liu, L. Lan Zhang, B. Simpson, and G. Xue Zhang. Comparison of reproductive development in triploid and diploid female rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 76(7):1742–1750, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teichert:2010:EDL

[128] M. A. K. Teichert, E. Kvingedal, T. Forseth, O. Ugedal, and A. G. Finstad. Effects of discharge and local density on the growth of juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 76(7):1751–1769, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lymbery:2010:PNE

[129] A. J. Lymbery, M. Hassan, D. L. Morgan, S. J. Beatty, and R. G. Doupé. Parasites of native and exotic freshwater fishes in south-western Australia. Journal of Fish Biology, 76(7):1770–1785, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2010:ECM

[130] R. J. Kennedy and W. W. Crozier. Evidence of changing migratory patterns of wild Atlantic salmon Salmo salar smolts in the River Bush, Northern Ireland, and possible associations with climate change. Journal of Fish Biology, 76(7):1786–1805, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thieltges:2010:SGS

[131] D. W. Thieltges, T. Dolch, M. Krakau, and R. Poulin. Salinity gradient shapes distance decay of similarity among parasite communities in three marine fishes. Journal of Fish Biology, 76(7):1806–1814, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreira:2010:RRU

[132] C. R. Moreira, M. E. Bichuette, O. T. Oyakawa, M. C. C. De Pinna, and E. Trajano. Rediscovery and redescription of the unusual subterranean characiform Stygichthys typhlops, with notes on its life history. Journal of Fish Biology, 76(7):1815–1824, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kane:2010:EFH

[133] A. S. Kane, J. Song, M. B. Halvorsen, D. L. Miller, J. D. Salierno, L. E. Wysocki, D. Zeddies, and A. N. Popper. Exposure of fish to highintensity sonar does not induce acute pathology. Journal of Fish Biology, 76(7):1825–1840, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukuda:2010:OCS

[134] H. Fukuda, S. Torisawa, Y. Sawada, and T. Takagi. Ontogenetic changes in schooling behaviour during larval and early juvenile stages of Pacific bluefin tuna Thunnus orientalis. Journal of Fish Biology, 76(7):1841– 1847, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nurminen:2010:FEP

[135] L. Nurminen, Z. Pekcan-Hekim, and J. Horppila. Feeding efficiency of planktivorous perch Perca fluviatilis and roach Rutilus rutilus in varying turbidity: an individual-based approach. Journal of Fish Biology, 76 (7):1848–1855, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaber:2010:DCF

[136] M. Schaber, C. Petereit, and M. Paulsen. Diet composition and feeding of European anchovy Engraulis encrasicolus in Kiel Bight, western Baltic Sea. Journal of Fish Biology, 76(7):1856–1862, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aalbers:2010:FRC

[137] S. A. Aalbers, D. Bernal, and C. A. Sepulveda. The functional role of the caudal fin in the feeding ecology of the common thresher shark Alopias vulpinus. Journal of Fish Biology, 76(7):1863–1868, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thackeray:2010:BRP

[138] Stephen J. Thackeray. Book review: Plankton: a Guide to Their Ecology and Monitoring for Water Quality — Edited by I. M. Suthers and D. Rissik. Journal of Fish Biology, 76(7):1869–1870, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sweeting:2010:BRR

[139] Roger A. Sweeting. Book review: Reconciling Fisheries with Conservation: Proceedings of the Fourth World Fisheries Congress — Edited by J. Nielsen, J. J. Dodson, K. Friedland, T. R. Hamon, J. Musick and E. Verspoor. Journal of Fish Biology, 76(7):1870–1871, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:2010:BRB

[140] Robin N. Gibson. Book review: The Biology of Blennies — Edited by R. A. Patzner, E. J. Gonçalves, P. A. Hastings and B. G. Kapoor. Journal of Fish Biology, 76(7):1871–1872, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2010:BRW

[141] Laith A. Jawad. Book review: World Atlas of Marine Fishes, 2nd edition — By R. H. Kuiter and H. Debelius. Journal of Fish Biology, 76(7):1873– 1874, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filby:2010:BRZ

[142] Amy L. Filby. Book review: Zebrafish: Methods and Protocols. Methods in Molecular Biology — By G. J. Lieschke, A. C. Oates and K. Kawakami. Journal of Fish Biology, 76(7):1874–1876, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Ca

[143] Anonymous. Corrigendum. Journal of Fish Biology, 76(7):1877, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siegfried:2010:SDG

[144] K. R. Siegfried. In search of determinants: gene expression during gonadal sex differentiation. Journal of Fish Biology, 76(8):1879–1902, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2010:COO

[145] T. G. Evans. Co-ordination of osmotic stress responses through osmosensing and signal transduction events in fishes. Journal of Fish Biology, 76(8):1903–1925, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2010:MIE

[146] C. Yang and J. Su. Molecular identification and expression analysis of toll-like receptor 3 in common carp Cyprinus carpio. Journal of Fish Biology, 76(8):1926–1939, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2010:GSR

[147] H. J. Lee, S. Pittlik, J. C. Jones, W. Salzburger, M. Barluenga, and A. Meyer. Genetic support for random mating between left and rightmouth morphs in the dimorphic scale-eating cichlid fish Perissodus microlepis from Lake Tanganyika. Journal of Fish Biology, 76(8):1940– 1957, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2010:PSO

[148] W.-J. Kim, K.-K. Kim, H.-S. Han, B.-H. Nam, Y.-O. Kim, H.-J. Kong, J.-K. Noh, and M. Yoon. Population structure of the olive flounder (Paralichthys olivaceus) in Korea inferred from microsatellite marker analysis. Journal of Fish Biology, 76(8):1958–1971, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wallace:2010:INM

[149] E. M. Wallace and M. D. Tringali. Identification of a novel member in the family Albulidae (bonefishes). Journal of Fish Biology, 76(8):1972– 1983, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2010:EMC

[150] J.-X. Liu and B. Ely. Evolution of an MHC class Ia gene fragment in four North American Morone species. Journal of Fish Biology, 76(8):1984– 1994, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Inacio:2010:EMI

[151] A. Inácio, I. Matos, M. P. Machado, and M. M. Coelho. An easier method to identify the individual genomic composition of allopolyploid complexes. Journal of Fish Biology, 76(8):1995–2001, May 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:CF

[152] Anonymous. Conservation of fishes. Journal of Fish Biology, 76(9):2003– 2008, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chong:2010:DER

[153] V. C. Chong, P. K. Y. Lee, and C. M. Lau. Diversity, extinction risk and conservation of Malaysian fishes. Journal of Fish Biology, 76(9):2009– 2066, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2010:SFC

[154] A. K. Whitfield and P. D. Cowley. The status of fish conservation in South African estuaries. Journal of Fish Biology, 76(9):2067–2089, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2010:SCC

[155] W. T. White and P. M. Kyne. The status of chondrichthyan conservation in the Indo-Australasian region. Journal of Fish Biology, 76(9):2090– 2117, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta:2010:FAH

[156] M. Barletta, A. J. Jaureguizar, C. Baigun, N. F. Fontoura, A. A. Agostinho, V. M. F. Almeida-Val, A. L. Val, R. A. Torres, L. F. Jimenes-Segura, T. Giarrizzo, N. N. Fabré, V. S. Batista, C. Lasso, D. C. Taphorn, M. F. Costa, P. T. Chaves, J. P. Vieira, and M. F. M. Corrêa. Fish and aquatic habitat conservation in South America: a continental overview with emphasis on neotropical systems. Journal of Fish Biology, 76(9): 2118–2176, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2010:FGS

[157] P. G. Azevedo, F. O. Mesquita, and R. J. Young. Fishing for gaps in science: a bibliographic analysis of Brazilian freshwater ichthyology from 1986 to 2005. Journal of Fish Biology, 76(9):2177–2193, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:2010:HRF

[158] I. G. Cowx, R. Arlinghaus, and S. J. Cooke. Harmonizing recreational fisheries and conservation objectives for aquatic biodiversity in inland waters. Journal of Fish Biology, 76(9):2194–2215, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Veiga:2010:QRS

[159] P. Veiga, J. Ribeiro, J. M. S. Gonçalves, and K. Erzini. Quantifying recreational shore angling catch and harvest in southern Portugal (north-east Atlantic Ocean): implications for conservation and integrated fisheries management. Journal of Fish Biology, 76(9):2216–2237, June 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2010:IOL

[160] T. H. Ng and H. H. Tan. The introduction, origin and life-history attributes of the non-native cichlid Etroplus suratensis in the coastal waters of Singapore. Journal of Fish Biology, 76(9):2238–2260, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2010:TNQ

[161] S. J. Cooke, A. J. Danylchuk, M. J. Kaiser, and M. A. Rudd. Is there a need for a ‘100 questions exercise’ to enhance fisheries and aquatic conservation, policy, management and research? Lessons from a global 100 questions exercise on conservation of biodiversity. Journal of Fish Biology, 76(9):2261–2286, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:2010:BTT

[162] A. J. W. Ward, E. M. A. Hensor, M. M. Webster, and P. J. B. Hart. Behavioural thermoregulation in two freshwater fish species. Journal of Fish Biology, 76(10):2287–2298, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pimentel:2010:DFP

[163] C. R. Pimentel and J.-C. Joyeux. Diet and food partitioning between juveniles of mutton Lutjanus analis, dog Lutjanus jocu and lane Lutjanus synagris snappers (Perciformes: Lutjanidae) in a mangrove-fringed estuarine environment. Journal of Fish Biology, 76(10):2299–2317, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tveiten:2010:ESL

[164] H. Tveiten, P. A. Bjørn, H. K. Johnsen, B. Finstad, and R. S. McKinley. Effects of the sea louse Lepeophtheirus salmonis on temporal changes in cortisol, sex steroids, growth and reproductive investment in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 76(10):2318–2341, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xu:2010:SDS

[165] C. L. Xu, B. H. Letcher, and K. H. Nislow. Size-dependent survival of brook trout Salvelinus fontinalis in summer: effects of water temperature and stream flow. Journal of Fish Biology, 76(10):2342–2369, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Box:2010:DSI

[166] A. Box, S. Deudero, A. Blanco, A. M. Grau, and F. Riera. Differences in δ ^{13}C and δ ^{15}N stable isotopes in the pearly razorfish Xyrichtys novacula related to the sex, location and spawning period. Journal of Fish Biology, 76(10):2370–2381, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horn:2010:FHA

[167] P. L. Horn, J. Forman, and M. R. Dunn. Feeding habits of alfonsino Beryx splendens. Journal of Fish Biology, 76(10):2382–2400, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaramillo-Villa:2010:AVF

[168] U. Jaramillo-Villa, J. A. Maldonado-Ocampo, and F. Escobar. Altitudinal variation in fish assemblage diversity in streams of the central Andes of Colombia. Journal of Fish Biology, 76(10):2401–2417, June 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shimose:2010:HFP

[169] T. Shimose, K. Yokawa, and H. Saito. Habitat and food partitioning of billfishes (Xiphioidei). Journal of Fish Biology, 76(10):2418–2433, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foster:2010:ULH

[170] S. J. Foster and A. C. J. Vincent. Using life-history information to assess potential effects of shrimp trawling on small fishes. Journal of Fish Biology, 76(10):2434–2454, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walther:2010:IEF

[171] B. D. Walther, T. S. Elsdon, and B. M. Gillanders. Interactive effects of food quality, temperature and rearing time on condition of juvenile black bream Acanthopagrus butcheri. Journal of Fish Biology, 76(10): 2455–2468, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Louca:2010:HAR

[172] V. Louca, S. W. Lindsay, C. Piyapong, and M. C. Lucas. Habitat associations, reproduction and diet of the Guinean tilapia Tilapia guineensis of the Gambia River floodplains. Journal of Fish Biology, 76(10):2469– 2485, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VandeNieuwegiessen:2010:CSF

[173] P. G. Van de Nieuwegiessen, N. M. Ramli, B. P. F. J. M. Knegtel, J. A. J. Verreth, and J. W. Schrama. Coping strategies in farmed African catfish Clarias gariepinus. Does it affect their welfare? Journal of Fish Biology, 76(10):2486–2501, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogan:2010:RGA

[174] F. Hogan and K. D. Friedland. Retrospective growth analysis of Atlantic salmon Salmo salar and implications for abundance trends. Journal of Fish Biology, 76(10):2502–2520, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsieh:2010:IHF

[175] H. Y. Hsieh, W. T. Lo, D. C. Liu, and W. C. Su. Influence of hydrographic features on larval fish distribution during the south-westerly monsoon in the waters of Taiwan, western North Pacific Ocean. Journal of Fish Biology, 76(10):2521–2539, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dantas:2010:MPC

[176] D. V. Dantas, M. Barletta, M. F. Costa, S. C. T. Barbosa-Cintra, F. E. Possatto, J. A. A. Ramos, A. R. A. Lima, and U. Saint-Paul. Movement patterns of catfishes (Ariidae) in a tropical semi-arid estuary. Journal of Fish Biology, 76(10):2540–2557, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janhunen:2010:PEE

[177] M. Janhunen, J. Piironen, and N. Peuhkuri. Parental effects on embryonic viability and growth in Arctic charr Salvelinus alpinus at two incubation temperatures. Journal of Fish Biology, 76(10):2558–2570, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blessing:2010:HKF

[178] J. J. Blessing, J. C. Marshall, and S. R. Balcombe. Humane killing of fishes for scientific research: a comparison of two methods. Journal of Fish Biology, 76(10):2571–2577, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brooker:2010:DCD

[179] R. M. Brooker, P. L. Munday, and T. D. Ainsworth. Diets of coraldwelling fishes of the genus Gobiodon with evidence of corallivory. Journal of Fish Biology, 76(10):2578–2583, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pander:2010:SEF

[180] J. Pander and J. Geist. Salmonid-egg floating boxes as bioindication for riverine water quality and stocking success. Journal of Fish Biology, 76 (10):2584–2590, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salonen:2010:PCP

[181] M. Salonen and J. Engström-Öst. Prey capture of pike Esox lucius larvae in turbid water. Journal of Fish Biology, 76(10):2591–2596, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maclean:2010:BRM

[182] Norman Maclean. Book review: Molecular Research in Aquaculture — Edited by K. Overturf. Journal of Fish Biology, 76(10):2597–2598, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parenti:2010:BRR

[183] Lynne R. Parenti. Book review: Reproductive Biology and Phylogeny of Fishes (Agnathans and Bony Fishes) — Edited by B. G. M. Jamieson. Journal of Fish Biology, 76(10):2599–2600, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:A

[184] Anonymous. Announcement. Journal of Fish Biology, 76(10):2601, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:CVa

[185] Anonymous. Contents of volume 76. Journal of Fish Biology, 76(10): 2602–2612, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:IVa

[186] Anonymous. Index, volume 76. Journal of Fish Biology, 76(10):i–xxvii, June 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koedijk:2010:IFF

[187] R. M. Koedijk, A. Folkvord, A. Foss, K. Pittman, S. O. Stefansson, S. Handeland, and A. K. Imsland. The influence of first-feeding diet on the Atlantic cod Gadus morhua phenotype: survival, development and long-term consequences for growth. Journal of Fish Biology, 77(1): 1–19, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinyoles:2010:LHT

[188] D. Vinyoles, A. De Sostoa, C. Franch, A. Maceda-Veiga, F. Casals, and N. Caiola. Life-history traits of the stone loach Barbatula barbatula. Journal of Fish Biology, 77(1):20–32, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tveiten:2010:BRS

[189] H. Tveiten, M. Frantzen, A. M. Scott, and A. P. Scott. Book review: Synthesis of 17,20β, 21-trihydroxypregn-4-en-3-one by ovaries of reproductively mature Atlantic cod Gadus morhua. Journal of Fish Biology, 77(1):33–53, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2010:OHR

[190] C. Sun, C. Huang, X. Su, X. Zhao, and Q. Dong. Optimization of handling and refrigerated storage of guppy Poecilia reticulata sperm. Journal of Fish Biology, 77(1):54–66, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ripley:2010:MQC

[191] J. L. Ripley, P. S. Williams, and C. M. Foran. Morphological and quantitative changes in paternal brood-pouch vasculature during embryonic development in two Syngnathus pipefishes. Journal of Fish Biology, 77 (1):67–79, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eloranta:2010:SOS

[192] A. P. Eloranta, K. K. Kahilainen, and R. I. Jones. Seasonal and ontogenetic shifts in the diet of Arctic charr Salvelinus alpinus in a subarctic lake. Journal of Fish Biology, 77(1):80–97, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schuckel:2010:LPC

[193] S. Schückel, S. Ehrich, I. Kröncke, and H. Reiss. Linking prey composition of haddock Melanogrammus aeglefinus to benthic prey availability in three different areas of the northern North Sea. Journal of Fish Biology, 77(1):98–118, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2010:RBA

[194] K.-S. Chen, P. R. Crone, and C.-C. Hsu. Reproductive biology of albacore Thunnus alalunga. Journal of Fish Biology, 77(1):119–136, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brady:2010:CER

[195] D. C. Brady and T. E. Targett. Characterizing the escape response of juvenile summer flounder Paralichthys dentatus to diel-cycling hypoxia. Journal of Fish Biology, 77(1):137–152, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horka:2010:VSA

[196] P. Horká, A. Ibbotson, J. I. Jones, R. J. Cove, and L. J. Scott. Validation of scale-age determination in European grayling Thymallus thymallus using tag-recapture analysis. Journal of Fish Biology, 77(1):153–161, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:2010:FAA

[197] M. Watanabe and K. Maekawa. Factors affecting aggressive behaviour of spawning migratory males towards mature male parr in masu salmon Oncorhynchus masou. Journal of Fish Biology, 77(1):162–168, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2010:RER

[198] A. D. Marshall and M. B. Bennett. Reproductive ecology of the reef manta ray Manta alfredi in southern Mozambique. Journal of Fish Biology, 77(1):169–190, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wielgoss:2010:GSI

[199] S. Wielgoss, F. Hollandt, T. Wirth, and A. Meyer. Genetic signatures in an invasive parasite of Anguilla anguilla correlate with differential stock management. Journal of Fish Biology, 77(1):191–210, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gundersen:2010:SMC

[200] A. C. Gundersen, C. Stenberg, I. Fossen, B. Lyberth, J. Boje, and O. A. Jørgensen. Sexual maturity cycle and spawning of Greenland halibut Reinhardtius hippoglossoides in the Davis Strait. Journal of Fish Biology, 77(1):211–226, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Partridge:2010:DIE

[201] G. J. Partridge and R. J. Michael. Direct and indirect effects of simulated calcareous dredge material on eggs and larvae of pink snapper Pagrus auratus. Journal of Fish Biology, 77(1):227–240, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Estlander:2010:DSF

[202] S. Estlander, L. Nurminen, M. Olin, M. Vinni, S. Immonen, M. Rask, J. Ruuhijärvi, J. Horppila, and H. Lehtonen. Diet shifts and food selection of perch Perca fluviatilis and roach Rutilus rutilus in humic lakes of varying water colour. Journal of Fish Biology, 77(1):241–256, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teletchea:2010:CEL

[203] F. Teletchea and P. Fontaine. Comparison of early life-stage strategies in temperate freshwater fish species: trade-offs are directed towards first feeding of larvae in spring and early summer. Journal of Fish Biology, 77(1):257–278, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kopecka-Pilarczyk:2010:EEH

[204] J. Kopecka-Pilarczyk and J. Coimbra. The effect of elevated hydrostatic pressure upon selected biomarkers in juvenile blackspot seabream Pagellus bogaraveo in a 14 day-long experiment. Journal of Fish Biology, 77(1):279–284, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fischer:2010:FRS

[205] C. Fischer and I. Schlupp. Feeding rates in the sailfin molly Poecilia latipinna and its coexisting sexual parasite, the gynogenetic Amazon molly Poecilia formosa. Journal of Fish Biology, 77(1):285–291, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuroki:2010:VET

[206] M. Kuroki, R. M. Buckley, L. L. LeClair, and L. Hauser. Validation and efficacy of transgenerational mass marking of otoliths in viviparous fish larvae. Journal of Fish Biology, 77(1):292–298, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karino:2010:RIA

[207] K. Karino, Y. Shimada, H. Kudo, and A. Sato. Relative importance of the area and intensity of the orange spots of male guppies Poecilia reticulata as mating traits preferred by females. Journal of Fish Biology, 77(1):299–307, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahnesjo:2010:BRS

[208] Ingrid Ahnesjö. Book review: Seahorses and Their Relatives — By R. H. Kuiter. Journal of Fish Biology, 77(1):308–309, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Cb

[209] Anonymous. Corrigendum. Journal of Fish Biology, 77(1):310, July 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrett:2010:AEL

[210] R. D. H. Barrett. Adaptive evolution of lateral plates in three-spined stickleback Gasterosteus aculeatus: a case study in functional analysis of natural variation. Journal of Fish Biology, 77(2):311–328, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plank:2010:PGS

[211] S. M. Plank, C. G. Lowe, K. A. Feldheim, R. R. Wilson, Jr., and J. A. Brusslan. Population genetic structure of the round stingray Urobatis halleri (Elasmobranchii: Rajiformes) in southern California and the Gulf of California. Journal of Fish Biology, 77(2):329–340, August 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xu:2010:DPG

[212] D. Xu, J. Wei, H. Cui, J. Gong, Y. Yan, R. Lai, and Q. Qin. Differential profiles of gene expression in grouper Epinephelus coioides, infected with Singapore grouper iridovirus, revealed by suppression subtractive hybridization and DNA microarray. Journal of Fish Biology, 77(2):341– 360, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:2010:PAR

[213] A. M. Pereira, J. I. Robalo, J. Freyhof, C. Maia, J. P. Fonseca, A. Valente, and V. C. Almada. Phylogeographical analysis reveals multiple conservation units in brook lampreys Lampetra planeri of Portuguese streams. Journal of Fish Biology, 77(2):361–371, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohkubo:2010:MCL

[214] M. Ohkubo, S. Katayama, and A. Shimizu. Molecular cloning and localization of the luteinizing hormone β subunit and glycoprotein hormone α subunit from Japanese anchovy Engraulis japonicus. Journal of Fish Biology, 77(2):372–387, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Song:2010:PGS

[215] N. Song, X. M. Zhang, X. F. Sun, T. Yanagimoto, and T. X. Gao. Population genetic structure and larval dispersal potential of spottedtail goby Synechogobius ommaturus in the north-west Pacific. Journal of Fish Biology, 77(2):388–402, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2010:SIM

[216] J. J. Chen, Q. Y. Du, Y. Y. Yue, B. J. Dang, and Z. J. Chang. Screening and identification of male-specific DNA fragments in common carps Cyprinus carpio using suppression subtractive hybridization. Journal of Fish Biology, 77(2):403–413, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sriwattanarothai:2010:MME

[217] N. Sriwattanarothai, D. Steinke, P. Ruenwongsa, R. Hanner, and B. Panijpan. Molecular and morphological evidence supports the species status of the Mahachai fighter Betta sp. Mahachai and reveals new species of Betta from Thailand. Journal of Fish Biology, 77(2):414–424, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2010:GCB

[218] D. G. Roberts, C. A. Gray, R. J. West, and D. J. Ayre. Gamete compatibility between marine and estuarine Acanthopagrus spp. (Sparidae) and their hybrids. Journal of Fish Biology, 77(2):425–431, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Rb

[219] Anonymous. Retraction. Journal of Fish Biology, 77(2):432, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Eb

[220] Anonymous. Erratum. Journal of Fish Biology, 77(2):433, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadswell:2010:NAS

[221] M. J. Dadswell, A. D. Spares, J. M. Reader, and M. J. W. Stokesbury. The North Atlantic subpolar gyre and the marine migration of Atlantic salmon Salmo salar: the ‘Merry-Go-Round’ hypothesis. Journal of Fish Biology, 77(3):435–467, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:2010:KDC

[222] K. M. Ferreira and A. L. Netto-Ferreira. Knodus dorsomaculatus (Characiformes: Characidae), a new species from Teles Pires River, Tapajós River basin, Brazil. Journal of Fish Biology, 77(3):468–478, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weir:2010:MPV

[223] L. K. Weir, C. Breau, J. A. Hutchings, and R. A. Cunjak. Multiple paternity and variance in male fertilization success within Atlantic salmon Salmo salar redds in a naturally spawning population. Journal of Fish Biology, 77(3):479–493, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muncaster:2010:RCF

[224] S. Muncaster, E. Andersson, O. S. Kjesbu, G. L. Taranger, A. B. Skiftesvik, and B. Norberg. The reproductive cycle of female Ballan wrasse Labrus bergylta in high latitude, temperate waters. Journal of Fish Biology, 77(3):494–511, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tong:2010:CRD

[225] X. H. Tong, Q. H. Liu, S. H. Xu, J. Li, Z. Z. Xiao, and D. Y. Ma. Changes in RNA, DNA, protein contents and growth of turbot Scophthalmus maximus larvae and juveniles. Journal of Fish Biology, 77(3): 512–525, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellsworth:2010:RTD

[226] C. M. Ellsworth, M. C. Belk, and C. J. Keleher. Residence time and drift patterns of larval June sucker Chasmistes liorus in the lower Provo River as determined by otolith microstructure. Journal of Fish Biology, 77(3):526–537, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Behrens:2010:IMS

[227] J. W. Behrens, J. K. Petersen, G. Ærtebjerg, and J. F. Steffensen. Influence of moderate and severe hypoxia on the diurnal activity pattern of lesser sandeel Ammodytes tobianus. Journal of Fish Biology, 77(3):538– 551, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uusi-Heikkila:2010:SDR

[228] S. Uusi-Heikkilä, C. Wolter, T. Meinelt, and R. Arlinghaus. Sizedependent reproductive success of wild zebrafish Danio rerio in the laboratory. Journal of Fish Biology, 77(3):552–569, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liasko:2010:BTR

[229] R. Liasko, V. Liousia, P. Vrazeli, O. Papiggioti, R. Chortatou, Th. J. Abatzopoulos, and I. D. Leonardos. Biological traits of rare males in the population of Carassius gibelio (Actinopterygii: Cyprinidae) from Lake Pamvotis (north-west Greece). Journal of Fish Biology, 77(3):570– 584, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Szlinder-Richert:2010:VFC

[230] J. Szlinder-Richert, Z. Usydus, M. Wyszyński, and M. Adamczyk. Variation in fat content and fatty-acid composition of the Baltic herring Clupea harengus membras. Journal of Fish Biology, 77(3):585–599, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cossington:2010:GRB

[231] S. Cossington, S. A. Hesp, N. G. Hall, and I. C. Potter. Growth and reproductive biology of the foxfish Bodianus frenchii, a very long-lived and monandric protogynous hermaphroditic labrid. Journal of Fish Biology, 77(3):600–626, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawabata:2010:ETC

[232] Y. Kawabata, J. Okuyama, K. Asami, K. Okuzawa, K. Yoseda, and N. Arai. Effects of a tropical cyclone on the distribution of hatcheryreared black-spot tuskfish Choerodon schoenleinii determined by acoustic telemetry. Journal of Fish Biology, 77(3):627–642, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakabe:2010:ITC

[233] R. Sakabe and J. M. Lyle. The influence of tidal cycles and freshwater inflow on the distribution and movement of an estuarine resident fish Acanthopagrus butcheri. Journal of Fish Biology, 77(3):643–660, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2010:HUM

[234] J. K. Carlson, M. M. Ribera, C. L. Conrath, M. R. Heupel, and G. H. Burgess. Habitat use and movement patterns of bull sharks Carcharhinus leucas determined using pop-up satellite archival tags. Journal of Fish Biology, 77(3):661–675, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bubley:2010:LHS

[235] W. J. Bubley and O. Pashuk. Life history of a simultaneously hermaphroditic fish, Diplectrum formosum. Journal of Fish Biology, 77(3):676–691, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reed:2010:LSV

[236] T. E. Reed, G. Martinek, and T. P. Quinn. Lake-specific variation in growth, migration timing and survival of juvenile sockeye salmon Oncorhynchus nerka: separating environmental from genetic influences. Journal of Fish Biology, 77(3):692–705, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deMitcheson:2010:GDG

[237] Y. Sadovy de Mitcheson, M. Liu, and S. Suharti. Gonadal development in a giant threatened reef fish, the humphead wrasse Cheilinus undulatus, and its relationship to international trade. Journal of Fish Biology, 77 (3):706–718, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aragon:2010:QOF

[238] L. Aragón, G. Aranda, A. Santos, and A. Medina. Quantification of ovarian follicles in bluefin tuna Thunnus thynnus by two stereological methods. Journal of Fish Biology, 77(3):719–730, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pusey:2010:WOL

[239] B. J. Pusey, A. H. Arthington, B. Stewart-Koster, M. J. Kennard, and M. G. Read. Widespread omnivory and low temporal and spatial variation in the diet of fishes in a hydrologically variable northern Australian river. Journal of Fish Biology, 77(3):731–753, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polacik:2010:DOA

[240] M. Polačik and M. Reichard. Diet overlap among three sympatric African annual killifish species Nothobranchius spp. from Mozambique. Journal of Fish Biology, 77(3):754–768, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:2010:PCK

[241] S. C. Mills, J. Mourier, and R. Galzin. Plasma cortisol and 11ketotestosterone enzyme immunoassay (EIA) kit validation for three fish species: the orange clownfish Amphiprion percula, the orangefin anemonefish Amphiprion chrysopterus and the blacktip reef shark Carcharhinus melanopterus. Journal of Fish Biology, 77(3):769–777, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Rc

[242] Anonymous. Retraction. Journal of Fish Biology, 77(3):778, August 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mincarone:2010:RAH

[243] M. M. Mincarone and B. Fernholm. Review of the Australian hagfishes with description of two new species of Eptatretus (Myxinidae). Journal of Fish Biology, 77(4):779–801, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguirre:2010:SDH

[244] W. E. Aguirre and O. Akinpelu. Sexual dimorphism of head morphology in three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 77(4):802–821, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boomhower:2010:PVP

[245] J. Boomhower, M. Romero, J. Posada, S. Kobara, and W. Heyman. Prediction and verification of possible reef-fish spawning aggregation sites in Los Roques Archipelago National Park, Venezuela. Journal of Fish Biology, 77(4):822–840, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ho:2010:RHA

[246] H.-C. Ho, P. Chakrabarty, and J. S. Sparks. Review of the Halieutichthys aculeatus species complex (Lophiiformes: Ogcocephalidae), with descriptions of two new species. Journal of Fish Biology, 77(4):841–869, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rakotobe:2010:ATE

[247] L. Rakotobe, K. Mezhoud, M. Berkal, C. Djediat, V. Jeannoda, B. Bodo, S. Puiseux-Dao, L. Mambu, and M. Edery. Acute toxic effects of 8epidiosbulbin E, a 19-norclerodane diterpene from yam Dioscorea antaly, on medaka Oryzias latipes embryos. Journal of Fish Biology, 77(4):870– 878, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marinho:2010:MCN

[248] M. M. F. Marinho and F. Langeani. Moenkhausia celibela: a new species from the Amazon basin, Brazil (Characiformes: Characidae). Journal of Fish Biology, 77(4):879–889, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartlett:2010:NSF

[249] J. A. Bartlett, M. P. Ward, S. J. Landsman, and J. M. Epifanio. Nest-site fidelity in parental male bluegill Lepomis macrochirus: spatial patterns and the influence of prior mating success. Journal of Fish Biology, 77(4): 890–906, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honorio:2010:CSR

[250] P. P. F. Honório, R. T. C. Ramos, and B. M. Feitoza. Composition and structure of reef fish communities in Paraíba state, north-eastern Brazil. Journal of Fish Biology, 77(4):907–926, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2010:DAS

[251] K. L. Williams, S. W. Griffiths, S. McKelvey, and J. D. Armstrong. Deposition of Atlantic salmon Salmo salar carcasses in a Scottish upland catchment. Journal of Fish Biology, 77(4):927–934, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagarese:2010:IEP

[252] S. R. Sagarese and M. G. Frisk. An investigation on the effect of photoperiod and temperature on vertebral band deposition in little skate Leucoraja erinacea. Journal of Fish Biology, 77(4):935–946, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dallas:2010:CEA

[253] L. J. Dallas, A. D. Shultz, A. J. Moody, K. A. Sloman, and A. J. Danylchuk. Chemical excretions of angled bonefish Albula vulpes and their potential use as predation cues by juvenile lemon sharks Negaprion brevirostris. Journal of Fish Biology, 77(4):947–962, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Concha:2010:ECD

[254] F. Concha, C. Bustamante, M. C. Oddone, S. Hernández, and J. Lamilla. Egg capsules of the dusky catshark Bythaelurus canescens (Carcharhiniformes, Scyliorhinidae) from the south-eastern Pacific Ocean. Journal of Fish Biology, 77(4):963–971, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suen:2010:RRM

[255] J. P. Suen and W. C. Su. Reconstructing riverine mesohabitat unit composition using fish community data and an autecology matrix. Journal of Fish Biology, 77(4):972–984, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baras:2010:IGT

[256] E. Baras and M. C. Lucas. Individual growth trajectories of sibling Brycon moorei raised in isolation since egg stage, and their relationship with aggressive behaviour. Journal of Fish Biology, 77(4):985–997, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernholm:2010:NSH

[257] B. Fernholm and M. M. Mincarone. A new species of hagfish (Myxinidae: Eptatretus) from Papua New Guinea. Journal of Fish Biology, 77(4): 998–1005, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Denton:2010:UPS

[258] K. P. Denton, H. B. Rich, Jr., J. W. Moore, and T. P. Quinn. The utilization of a Pacific salmon Oncorhynchus nerka subsidy by three populations of charr Salvelinus spp. Journal of Fish Biology, 77(4):1006–1023, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2010:ART

[259] K. Morita, M. Fukuwaka, and N. Tanimata. Age-related thermal habitat use by Pacific salmon Oncorhynchus spp. Journal of Fish Biology, 77(4): 1024–1029, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gold:2010:SFB

[260] J. R. Gold, M. A. Renshaw, E. Saillant, and R. R. Vega. Spawning frequency of brood dams and sires in a marine fish stock-enhancement hatchery. Journal of Fish Biology, 77(4):1030–1040, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2010:BRF

[261] Laith A. Jawad. Book review: Freshwater Fishes of Iraq — By B. Coad. Journal of Fish Biology, 77(4):1041–1042, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Ec

[262] Anonymous. Erratum. Journal of Fish Biology, 77(4):1043, September 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cerda:2010:GRF

[263] J. Cerdà, S. Douglas, and M. Reith. Genomic resources for flatfish research and their applications. Journal of Fish Biology, 77(5):1045–1070, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2010:RLM

[264] M. Liu, Z. C. Lu, T. X. Gao, T. Yanagimoto, and Y. Sakurai. Remarkably low mtDNA control-region diversity and shallow population structure in Pacific cod Gadus macrocephalus. Journal of Fish Biology, 77(5):1071– 1082, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ou-yang:2010:ECN

[265] Z. L. Ou-yang, X. H. Huang, E. Y. Huang, Y. H. Huang, J. Gong, J. J. Sun, and Q. W. Qin. Establishment and characterization of a new marine fish cell line derived from red-spotted grouper Epinephelus akaara. Journal of Fish Biology, 77(5):1083–1095, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dash:2010:DCE

[266] C. Dash, P. Routray, S. Tripathy, D. K. Verma, B. C. Guru, P. K. Meher, S. Nandi, and A. E. Eknath. Derivation and characterization of embryonic stem-like cells of Indian major carp Catla catla. Journal of Fish Biology, 77(5):1096–1113, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berendzen:2010:PGE

[267] P. B. Berendzen, J. F. Dugan, and T. Gamble. Post-glacial expansion into the Paleozoic Plateau: evidence of an Ozarkian refugium for the Ozark minnow Notropis nubilus (Teleostei: Cypriniformes). Journal of Fish Biology, 77(5):1114–1136, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coscia:2010:HPS

[268] I. Coscia, V. Rountree, J. J. King, W. K. Roche, and S. Mariani. A highly permeable species boundary between two anadromous fishes. Journal of Fish Biology, 77(5):1137–1149, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haynes:2010:PGI

[269] G. D. Haynes, D. M. Gilligan, P. Grewe, C. Moran, and F. W. Nicholas. Population genetics of invasive common carp Cyprinus carpio L. in coastal drainages in eastern Australia. Journal of Fish Biology, 77(5): 1150–1157, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muenzel:2010:GSV

[270] F. M. Muenzel, W. Salzburger, M. Sanetra, B. Grabherr, and A. Meyer. Genetic structure of the vairone Telestes souffia in the eastern part of Lake Constance, central Europe. Journal of Fish Biology, 77(5):1158– 1164, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boomer:2010:GDS

[271] J. J. Boomer, V. Peddemors, and A. J. Stow. Genetic data show that Carcharhinus tilstoni is not confined to the tropics, highlighting the importance of a multifaceted approach to species identification. Journal of Fish Biology, 77(5):1165–1172, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamanoue:2010:PFT

[272] Y. Yamanoue, D. H. E. Setiamarga, and K. Matsuura. Pelvic fins in teleosts: structure, function and evolution. Journal of Fish Biology, 77 (6):1173–1208, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purcell:2010:ASL

[273] K. M. Purcell, P. L. Klerks, and P. L. Leberg. Adaptation to sea level rise: does local adaptation influence the demography of coastal fish populations? Journal of Fish Biology, 77(6):1209–1218, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hobbs:2010:IST

[274] J.-P. A. Hobbs and C. A. McDonald. Increased seawater temperature and decreased dissolved oxygen triggers fish kill at the Cocos (Keeling) Islands, Indian Ocean. Journal of Fish Biology, 77(6):1219–1229, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rennie:2010:PLH

[275] M. D. Rennie, C. F. Purchase, B. J. Shuter, N. C. Collins, P. A. Abrams, and G. E. Morgan. Prey life-history and bioenergetic responses across a predation gradient. Journal of Fish Biology, 77(6):1230–1251, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tobin:2010:TMT

[276] D. Tobin, P. J. Wright, and M. O’Sullivan. Timing of the maturation transition in haddock Melanogrammus aeglefinus. Journal of Fish Biology, 77(6):1252–1267, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kierl:2010:SPP

[277] N. C. Kierl and C. E. Johnston. Sound production in the pygmy sculpin Cottus paulus (Cottidae) during courtship and agonistic behaviours. Journal of Fish Biology, 77(6):1268–1281, October 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:2010:ELS

[278] A. M. Grant, M. Gardner, L. M. Hanson, A. P. Farrell, and C. J. Brauner. Early life stage salinity tolerance of wild and hatchery-reared juvenile pink salmon Oncorhynchus gorbuscha. Journal of Fish Biology, 77(6): 1282–1292, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niklitschek:2010:EFE

[279] E. J. Niklitschek and D. H. Secor. Experimental and field evidence of behavioural habitat selection by juvenile Atlantic Acipenser oxyrinchus oxyrinchus and shortnose Acipenser brevirostrum sturgeons. Journal of Fish Biology, 77(6):1293–1308, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roo:2010:OSD

[280] F. J. Roo, C.-M. Hernández-Cruz, J.-A. Socorro, H. Fernández-Palacios, and M.-S. Izquierdo. Occurrence of skeletal deformities and osteological development in red porgy Pagrus pagrus larvae cultured under different rearing techniques. Journal of Fish Biology, 77(6):1309–1324, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kanaji:2010:VOP

[281] Y. Kanaji, M. Kishida, Y. Watanabe, T. Kawamura, S. Xie, Y. Yamashita, C. Sassa, and Y. Tsukamoto. Variations in otolith patterns, sizes and body morphometrics of jack mackerel Trachurus japonicus juveniles. Journal of Fish Biology, 77(6):1325–1342, October 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brunnschweiler:2010:OAS

[282] J. M. Brunnschweiler, N. Queiroz, and D. W. Sims. Oceans apart? Shortterm movements and behaviour of adult bull sharks Carcharhinus leucas in Atlantic and Pacific Oceans determined from pop-off satellite archival tagging. Journal of Fish Biology, 77(6):1343–1358, October 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakefield:2010:ALD

[283] C. B. Wakefield. Annual, lunar and diel reproductive periodicity of a spawning aggregation of snapper Pagrus auratus (Sparidae) in a marine embayment on the lower west coast of Australia. Journal of Fish Biology, 77(6):1359–1378, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roth:2010:SFW

[284] B. M. Roth, T. R. Hrabik, C. T. Solomon, N. Mercado-Silva, and J. F. Kitchell. A simulation of food-web interactions leading to rainbow smelt Osmerus mordax dominance in Sparkling Lake, Wisconsin. Journal of Fish Biology, 77(6):1379–1405, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Webb:2010:EBC

[285] J. F. Webb, J. L. Herman, C. F. Woods, and D. R. Ketten. The ears of butterflyfishes (Chaetodontidae): ‘hearing generalists’ on noisy coral reefs? Journal of Fish Biology, 77(6):1406–1423, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cobb:2010:AAA

[286] C. S. Cobb, J. A. Brown, and J. C. Rankin. Antidiuretic action of angiotensin II in the river lamprey Lampetra fluviatilis: evidence for endocrine control of kidney function in cyclostomes. Journal of Fish Biology, 77(6):1424–1431, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:2010:SBU

[287] F. Ferreira, M. M. Santos, M. A. Reis-Henriques, N. M. Vieira, and N. M. Monteiro. Sexing blennies using genital papilla morphology or ano-genital distance. Journal of Fish Biology, 77(6):1432–1438, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendonca:2010:SCN

[288] F. Z. Mendonça, G. L. Volpato, R. S. Costa-Ferreira, and E. Gonçalves de Freitas. Substratum choice for nesting in male Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 77(6):1439–1445, October 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:2010:BRE

[289] Miran Aprahamian. Book review: Eels at the Edge: Science, Status and Conservation Concerns — Edited by J. M. Casselman and D. K. Cairns. Journal of Fish Biology, 77(6):1446–1447, October 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chassot:2010:CCR

[290] E. Chassot, D. M. Kaplan, V. Ortiz de Zarate, E. Romanov, and A. Fonteneau. Comment on ‘Clarification regarding the distribution of bigeye tuna Thunnus obesus in the Atlantic Ocean, including British waters’. Journal of Fish Biology, 77(7):1449–1454, November 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See reply

Powell:2010:RCA

[291] A. Powell, D. M. Herdson, G. S. Parry, and J. D. R. Houghton. Reply to Chassot et al.: Rare tuna species in British waters. Journal of Fish Biology, 77(7):1455–1458, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [290].

Lamatsch:2010:MTD

[292] D. K. Lamatsch, M. Stöck, R. Fuchs, M. Döbler, R. Wacker, J. Parzefall, I. Schlupp, and M. Schartl. Morphology, testes development and behaviour of unusual triploid males in microchromosome-carrying clones of Poecilia formosa. Journal of Fish Biology, 77(7):1459–1487, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belanger:2010:EST

[293] A. J. Belanger, I. Bobeica, and D. M. Higgs. The effect of stimulus type and background noise on hearing abilities of the round goby Neogobius melanostomus. Journal of Fish Biology, 77(7):1488–1504, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2010:RBS

[294] E. D. Farrell, S. Mariani, and M. W. Clarke. Reproductive biology of the starry smooth-hound shark Mustelus asterias: geographic variation and implications for sustainable exploitation. Journal of Fish Biology, 77(7): 1505–1525, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honda:2010:REC

[295] K. Honda, H. Kagiwada, N. Tojo, and K. Miyashita. Riverine environmental characteristics and seasonal habitat use by adult Sakhalin taimen Hucho perryi. Journal of Fish Biology, 77(7):1526–1541, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lai:2010:VRS

[296] X.-J. Lai and W.-S. Hong. Variation with reproductive status of PGE_{2} receptor immunoreactivities in the Bostrichthys sinensis olfactory system. Journal of Fish Biology, 77(7):1542–1551, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horn:2010:AVG

[297] P. L. Horn, H. L. Neil, L. J. Paul, and P. Marriott. Age validation and growth of bluenose Hyperoglyphe antarctica using the bomb chronometer method of radiocarbon ageing. Journal of Fish Biology, 77(7):1552–1563, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tovar-Avila:2010:URA

[298] J. Tovar-Ávila, R. W. Day, and T. I. Walker. Using rapid assessment and demographic methods to evaluate the effects of fishing on Heterodontus portusjacksoni off far-eastern Victoria, Australia. Journal of Fish Biology, 77(7):1564–1578, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kallert:2010:ART

[299] D. M. Kallert, S. Ponader, S. Adelt, P. Kaese, R. Geyer, W. Haas, and M. El-Matbouli. Analysis of rainbow trout Oncorhynchus mykiss epidermal mucus and evaluation of semiochemical activity for polar filament discharge in Myxobolus cerebralis actinospores. Journal of Fish Biology, 77(7):1579–1598, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medeiros:2010:HUB

[300] P. R. Medeiros, A. T. Souza, and M. I. Ilarri. Habitat use and behavioural ecology of the juveniles of two sympatric damselfishes (Actinopterygii: Pomacentridae) in the south-western Atlantic Ocean. Journal of Fish Biology, 77(7):1599–1615, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Easy:2010:CAS

[301] R. H. Easy and N. W. Ross. Changes in Atlantic salmon Salmo salar mucus components following short- and long-term handling stress. Journal of Fish Biology, 77(7):1616–1631, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elvidge:2010:SMR

[302] C. K. Elvidge, I. W. Ramnarine, J.-G. J. Godin, and G. E. Brown. Sizemediated response to public cues of predation risk in a tropical stream fish. Journal of Fish Biology, 77(7):1632–1644, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polgar:2010:DAR

[303] G. Polgar, A. Sacchetti, and P. Galli. Differentiation and adaptive radiation of amphibious gobies (Gobiidae: Oxudercinae) in semi-terrestrial habitats. Journal of Fish Biology, 77(7):1645–1664, November 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diaz:2010:GHG

[304] A. O. Díaz, A. M. García, A. H. Escalante, and A. L. Goldemberg. Glycoproteins histochemistry of the gills of Odontesthes bonariensis (Teleostei, Atherinopsidae). Journal of Fish Biology, 77(7):1665–1673, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wen:2010:EHM

[305] C. K.-C. Wen, M. S. Pratchett, K.-T. Shao, K.-P. Kan, and B. K. K. Chan. Effects of habitat modification on coastal fish assemblages. Journal of Fish Biology, 77(7):1674–1687, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnett:2010:SOP

[306] A. Barnett, J. D. Stevens, S. D. Frusher, and J. M. Semmens. Seasonal occurrence and population structure of the broadnose sevengill shark Notorynchus cepedianus in coastal habitats of south-east Tasmania. Journal of Fish Biology, 77(7):1688–1701, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulsen:2010:EFD

[307] S. B. Poulsen, J. C. Svendsen, L. F. Jensen, C. Schulz, T. Jäger-Kleinicke, and H. Schwarten. Effects of food deprivation on refuge use and dispersal in juvenile North Sea houting Coregonus oxyrinchus under experimental conditions. Journal of Fish Biology, 77(7):1702–1708, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takemura:2010:PRR

[308] A. Takemura, S. Tamotsu, T. Miwa, and H. Yamamoto. Preliminary results on the reproduction of a deep-sea snailfish Careproctus rhodomelas around the active hydrothermal vent on the Hatoma Knoll, Okinawa, Japan. Journal of Fish Biology, 77(7):1709–1715, November 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cook:2010:RTO

[309] C. L. Cook and D. J. Coughlin. Rainbow trout Oncorhynchus mykiss consume less energy when swimming near obstructions. Journal of Fish Biology, 77(7):1716–1723, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaudo:2010:MOI

[310] J. J. Vaudo, P. Matich, and M. R. Heithaus. Mother–offspring isotope fractionation in two species of placentatrophic sharks. Journal of Fish Biology, 77(7):1724–1727, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2010:BRS

[311] Ian J. Winfield. Book review: To Sea and Back: the Heroic Life of the Atlantic Salmon — Edited by R. Shelton. Journal of Fish Biology, 77(7): 1728–1729, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:Cc

[312] Anonymous. Corrigendum. Journal of Fish Biology, 77(7):1730, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnuson:2010:HHT

[313] J. J. Magnuson. History and heroes: the thermal niche of fishes and long-term lake ice dynamics. Journal of Fish Biology, 77(8):1731–1744, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Portner:2010:CCE

[314] H. O. Pörtner and M. A. Peck. Climate change effects on fishes and fisheries: towards a cause-and-effect understanding. Journal of Fish Biology, 77(8):1745–1779, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2010:TES

[315] A. D. Nunn, P. A. Frear, M. Lee, and I. G. Cowx. Is there evidence for a shift in fish growth and recruitment success linked to climate change? Journal of Fish Biology, 77(8):1780–1792, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2010:TRA

[316] J. M. Elliott and J. A. Elliott. Temperature requirements of Atlantic salmon Salmo salar, brown trout Salmo trutta and Arctic charr Salvelinus alpinus: predicting the effects of climate change. Journal of Fish Biology, 77(8):1793–1817, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strussmann:2010:ICC

[317] C. A. Strüssmann, D. O. Conover, G. M. Somoza, and L. A. Miranda. Implications of climate change for the reproductive capacity and survival of New World silversides (family Atherinopsidae). Journal of Fish Biology, 77(8):1818–1834, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orpwood:2010:IBR

[318] J. E. Orpwood, J. D. Armstrong, and S. W. Griffiths. Interactions between riparian shading and food supply: a seasonal comparison of effects on time budgets, space use and growth in Atlantic salmon Salmo salar. Journal of Fish Biology, 77(8):1835–1849, November 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zieba:2010:EET

[319] G. Zieba, M. G. Fox, and G. H. Copp. The effect of elevated temperature on spawning of introduced pumpkinseed Lepomis gibbosus in Europe. Journal of Fish Biology, 77(8):1850–1855, November 2010. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kopprio:2010:VPO

[320] G. A. Kopprio, R. H. Freije, C. A. Strüssmann, G. Kattner, M. S. Hoffmeyer, C. A. Popovich, and R. J. Lara. Vulnerability of pejerrey Odontesthes bonariensis populations to climate change in pampean lakes of Argentina. Journal of Fish Biology, 77(8):1856–1866, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lyons:2010:PEC

[321] J. Lyons, J. S. Stewart, and M. Mitro. Predicted effects of climate warming on the distribution of 50 stream fishes in Wisconsin, U.S.A. Journal of Fish Biology, 77(8):1867–1898, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brander:2010:CGM

[322] K. M. Brander. Cod Gadus morhua and climate change: processes, productivity and prediction. Journal of Fish Biology, 77(8):1899–1911, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:2010:CCG

[323] M. W. Aprahamian, C. D. Aprahamian, and A. M. Knights. Climate change and the green energy paradox: the consequences for twaite shad Alosa fallax from the River Severn, U.K. Journal of Fish Biology, 77(8): 1912–1930, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feary:2010:FCW

[324] D. A. Feary, J. A. Burt, A. G. Bauman, P. Usseglio, P. F. Sale, and G. H. Cavalcante. Fish communities on the world’s warmest reefs: what can they tell us about the effects of climate change in the future? Journal of Fish Biology, 77(8):1931–1947, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharma:2010:UHM

[325] R. Sharma and M. Liermann. Using hierarchical models to estimate effects of ocean anomalies on north-west Pacific Chinook salmon Oncorhynchus tshawytscha recruitment. Journal of Fish Biology, 77(8): 1948–1963, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lynch:2010:ICC

[326] A. J. Lynch, W. W. Taylor, and K. D. Smith. The influence of changing climate on the ecology and management of selected Laurentian Great Lakes fisheries. Journal of Fish Biology, 77(8):1964–1982, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ling:2010:SED

[327] N. Ling. Socio-economic drivers of freshwater fish declines in a changing climate: a New Zealand perspective. Journal of Fish Biology, 77(8): 1983–1992, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pardinas:2010:CCO

[328] A. F. Pardiñas, D. Campo, I. G. Pola, L. Miralles, F. Juanes, and E. Garcia-Vazquez. Climate change and oceanic barriers: genetic differentiation in Pomatomus saltatrix (Pisces: Pomatomidae) in the North Atlantic Ocean and the Mediterranean Sea. Journal of Fish Biology, 77(8):1993–1998, November 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hook:2010:PPE

[329] S. E. Hook. Promise and progress in environmental genomics: a status report on the applications of gene expression-based microarray studies in ecologically relevant fish species. Journal of Fish Biology, 77(9):1999– 2022, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eizaguirre:2010:MHC

[330] C. Eizaguirre and T. L. Lenz. Major histocompatibility complex polymorphism: dynamics and consequences of parasite-mediated local adaptation in fishes. Journal of Fish Biology, 77(9):2023–2047, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swatdipong:2010:HLP

[331] A. Swatdipong, A. Vasemägi, T. Niva, M.-L. Koljonen, and C. R. Primmer. High level of population genetic structuring in lake-run brown trout, Salmo trutta, of the Inari Basin, northern Finland. Journal of Fish Biology, 77(9):2048–2071, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OBryan:2010:PCG

[332] D. M. O’Bryan, Z. Xie, Y. Wang, J. Du, C. J. Brauner, J. G. Richards, C. M. Wood, X.-Q. Chen, and B. W. Murray. Phylogeography and conservation genetics of Lake Qinghai scaleless carp Gymnocypris przewalskii. Journal of Fish Biology, 77(9):2072–2092, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Page:2010:CPM

[333] T. J. Page and J. M. Hughes. Comparing the performance of multiple mitochondrial genes in the analysis of Australian freshwater fishes. Journal of Fish Biology, 77(9):2093–2122, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamasaki:2010:GEH

[334] K. Hamasaki, S. Toriya, H. Shishidou, T. Sugaya, and S. Kitada. Genetic effects of hatchery fish on wild populations in red sea bream Pagrus major (Perciformes, Sparidae) inferred from a partial sequence of mitochondrial DNA. Journal of Fish Biology, 77(9):2123–2136, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garvin:2010:ASN

[335] M. R. Garvin and A. J. Gharrett. Application of single nucleotide polymorphism markers to chum salmon Oncorhynchus keta: discovery, genotyping and linkage phase resolution. Journal of Fish Biology, 77(9): 2137–2162, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drozd:2010:SOT

[336] B. Drozd, M. Flajšhans, and P. Ráb. Sympatric occurrence of triploid, aneuploid and tetraploid weatherfish Misgurnus fossilis (Cypriniformes, Cobitidae). Journal of Fish Biology, 77(9):2163–2170, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halacka:2010:ESC

[337] K. Halačka, L. Vetešník, I. Papoušek, J. Mendel, and A. ŠImková. The epidermal structure of Carassius gibelio: a link with ploidy status in spawning and postspawning periods. Journal of Fish Biology, 77(9): 2171–2179, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciotti:2010:DVF

[338] B. J. Ciotti, T. E. Targett, R. D. M. Nash, R. S. Batty, M. T. Burrows, and A. J. Geffen. Development, validation and field application of an RNA-based growth index in juvenile plaice Pleuronectes platessa. Journal of Fish Biology, 77(10):2181–2209, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winker:2010:VCD

[339] H. Winker, O. L. F. Weyl, A. J. Booth, and B. R. Ellender. Validating and corroborating the deposition of two annual growth zones in asteriscus otoliths of common carp Cyprinus carpio from South Africa’s largest impoundment. Journal of Fish Biology, 77(10):2210–2228, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grandcourt:2010:RBI

[340] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Reproductive biology and implications for management of the spangled emperor Lethrinus nebulosus in the southern Arabian Gulf. Journal of Fish Biology, 77(10):2229–2247, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gigliotti:2010:SAE

[341] E. S. Gigliotti, D. F. M. Gherardi, E. T. Paes, R. B. Souza, and M. Katsuragawa. Spatial analysis of egg distribution and geographic changes in the spawning habitat of the Brazilian sardine Sardinella brasiliensis. Journal of Fish Biology, 77(10):2248–2267, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lankov:2010:FEP

[342] A. Lankov, H. Ojaveer, M. Simm, M. Põllupüü, and C. Möllmann. Feeding ecology of pelagic fish species in the Gulf of Riga (Baltic Sea): the importance of changes in the zooplankton community. Journal of Fish Biology, 77(10):2268–2284, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guerrero-Tortolero:2010:ECS

[343] D. A. Guerrero-Tortolero, R. Campos-Ramos, M. A. Burgos-Aceves, J. C. Pérez-Urbiola, and G. Colado-Durán. Effects of compressed seasonally changing day-length cycles on spawning performance, production of viable eggs and levels of vitellogenin in plasma in female yellowtail snapper Lutjanus argentiventris. Journal of Fish Biology, 77(10):2285–2297, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morbey:2010:PCS

[344] Y. E. Morbey, P. Couture, P. Busby, and B. J. Shuter. Physiological correlates of seasonal growth patterns in lake trout Salvelinus namaycush. Journal of Fish Biology, 77(10):2298–2314, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2010:OCS

[345] K. Takahashi, R. Masuda, and Y. Yamashita. Ontogenetic changes in the spatial learning capability of jack mackerel Trachurus japonicus. Journal of Fish Biology, 77(10):2315–2325, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mather:2010:DDR

[346] M. E. Mather, J. T. Finn, S. M. Pautzke, D. Fox, T. Savoy, H. M. Brundage III, L. A. Deegan, and R. M. Muth. Diversity in destinations, routes and timing of small adult and sub-adult striped bass Morone saxatilis on their southward autumn migration. Journal of Fish Biology, 77(10):2326–2337, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalogirou:2010:DSF

[347] S. Kalogirou, M. Corsini-Foka, A. Sioulas, H. Wennhage, and L. Pihl. Diversity, structure and function of fish assemblages associated with Posidonia oceanica beds in an area of the eastern Mediterranean Sea and the role of non-indigenous species. Journal of Fish Biology, 77(10):

2338–2357, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nika:2010:BTS

[348] N. Nika and T. Virbickas. Brown trout Salmo trutta redd superimposition by spawning Lampetra species in a lowland stream. Journal of Fish Biology, 77(10):2358–2372, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaefer:2010:RID

[349] S. A. Schaefer and J. Arroyave. Rivers as islands: determinants of the distribution of Andean astroblepid catfishes. Journal of Fish Biology, 77(10):2373–2390, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Etheridge:2010:HHP

[350] E. C. Etheridge, C. Harrod, C. W. Bean, and C. E. Adams. Has habitat heterogeneity promoted phenotypic and ecological sub-structuring among a Coregonus lavaretus population in a large Scottish lake? Journal of Fish Biology, 77(10):2391–2404, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2010:AGN

[351] I. P. Jacobsen and M. B. Bennett. Age and growth of Neotrygon picta, Neotrygon annotata and Neotrygon kuhlii from north-east Australia, with notes on their reproductive biology. Journal of Fish Biology, 77(10): 2405–2422, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz:2010:FRB

[352] M. Muñoz, C. Dimitriadis, M. Casadevall, S. Vila, E. Delgado, J. Lloret, and F. Saborido-Rey. Female reproductive biology of the bluemouth Helicolenus dactylopterus dactylopterus: spawning and fecundity. Journal of Fish Biology, 77(10):2423–2442, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harant:2010:ESL

[353] R. Harant and J. Bohlen. Enobarbus is a synonym of Lepidocephalichthys (Osteichthyes, Cobitidae). Journal of Fish Biology, 77(10):2443–2453, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2010:SCP

[354] J. R. Britton. Scale circuli patterns differentiate between hatchery-reared and wild Rutilus rutilus during evaluation of fish stocking. Journal of Fish Biology, 77(10):2454–2459, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grimardias:2010:EHI

[355] D. Grimardias, N. Merchermek, A. Manicki, J. Chebaux, P. Gaudin, M. Jarry, and E. Beall. Effects of habitat on individual reproductive success of mature male parr of Atlantic salmon Salmo salar. Journal of Fish Biology, 77(10):2460–2466, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2010:BRA

[356] Ian J. Winfield. Book review: Assessing the Conservation Value of Fresh Waters: an International Perspective — Edited by P. J. Boon and C. M. Pringle. Journal of Fish Biology, 77(10):2467, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2010:A

[357] John F. Craig. Announcement. Journal of Fish Biology, 77(10):2469, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:CVb

[358] Anonymous. Contents of volume 77. Journal of Fish Biology, 77(10): 2470–2480, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2010:IVb

[359] Anonymous. Index, volume 77. Journal of Fish Biology, 77(10):i–xxx, December 2010. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larsen:2011:GEA

[360] P. F. Larsen, P. M. Schulte, and E. E. Nielsen. Gene expression analysis for the identification of selection and local adaptation in fishes. Journal of Fish Biology, 78(1):1–22, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bilotta:2011:DEE

[361] G. S. Bilotta, P. Sibley, J. Hateley, and A. Don. The decline of the European eel Anguilla anguilla: quantifying and managing escapement to support conservation. Journal of Fish Biology, 78(1):23–38, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonBiela:2011:LTI

[362] V. R. von Biela, C. E. Zimmerman, and L. L. Moulton. Long-term increases in young-of-the-year growth of Arctic cisco Coregonus autumnalis and environmental influences. Journal of Fish Biology, 78(1):39–56, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharma:2011:GSP

[363] S. Sharma and J. M. Hughes. Genetic structure and phylogeography of two freshwater fishes, Rhadinocentrus ornatus and Hypseleotris compressa, in southern Queensland, Australia, inferred from allozymes and mitochondrial DNA. Journal of Fish Biology, 78(1):57–77, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bubic:2011:TSV

[364] T. Šegvić Bubić, L. Grubišić, V. Tičina, and I. Katavić. Temporal and spatial variability of pelagic wild fish assemblages around Atlantic bluefin tuna Thunnus thynnus farms in the eastern Adriatic Sea. Journal of Fish Biology, 78(1):78–97, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schmidt:2011:GAT

[365] D. J. Schmidt, D. A. Crook, J. P. O’Connor, and J. M. Hughes. Genetic analysis of threatened Australian grayling Prototroctes maraena suggests recruitment to coastal rivers from an unstructured marine larval source population. Journal of Fish Biology, 78(1):98–111, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kimura:2011:SSI

[366] M. R. Kimura and H. Munehara. Spawning substrata are important for breeding habitat selection but do not determine premating reproductive isolation in three sympatric Hexagrammos species. Journal of Fish Biology, 78(1):112–126, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jiang:2011:SMC

[367] P. Jiang, J. J. Bai, X. Ye, Q. Jian, M. Chen, and X. Q. Chen. Shoaling and mate choice of wild-type Tanichthys albonubes in the presence of the red fluorescent transgenic conspecifics. Journal of Fish Biology, 78 (1):127–137, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spieth:2011:MPF

[368] H. R. Spieth, T. Möller, Ch. Ptatscheck, A. Kazemi-Dinan, and W. Traunspurger. Meiobenthos provides a food resource for young cyprinids. Journal of Fish Biology, 78(1):138–149, January 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xiao:2011:MAD

[369] J. Xiao, T. M. Zou, L. Chen, S. J. Liu, J. Xiao, H. Zhang, Y. Long, J. P. Yan, R. R. Zhao, M. Tao, C. Zhang, C. P. You, and Y. Liu. Microsatellite analysis of different ploidy offspring of artificial gynogenesis in Cyprinus carpio. Journal of Fish Biology, 78(1):150–165, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutson:2011:SVP

[370] K. S. Hutson, E. L. Brock, and M. A. Steer. Spatial variation in parasite abundance: evidence of geographical population structuring in southern garfish Hyporhamphus melanochir. Journal of Fish Biology, 78(1):166– 182, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pansonato-Alves:2011:CDP

[371] J. C. Pansonato-Alves, M. R. Vicari, C. Oliveira, and F. Foresti. Chromosomal diversification in populations of Characidium cf. gomesi (Teleostei, Crenuchidae). Journal of Fish Biology, 78(1):183–194, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DiSanto:2011:PFT

[372] V. Di Santo and W. A. Bennett. Is post-feeding thermotaxis advantageous in elasmobranch fishes? Journal of Fish Biology, 78(1):195–207, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chassaing:2011:DSE

[373] O. Chassaing, C. Hänni, and P. Berrebi. Distinguishing species of European sturgeons Acipenser spp. using microsatellite allele sequences.

Journal of Fish Biology, 78(1):208–226, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konecna:2011:SDP

[374] M. Konečná and M. Reichard. Seasonal dynamics in population characteristics of European bitterling Rhodeus amarus in a small lowland river. Journal of Fish Biology, 78(1):227–239, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguilera:2011:NSH

[375] G. Aguilera, J. M. Mirande, and M. M. Azpelicueta. A new species of Heptapterus Bleeker 1858 (Siluriformes, Heptapteridae) from the Río Salí basin, north-western Argentina. Journal of Fish Biology, 78(1):240– 250, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2011:RAF

[376] A. K. Imsland, R. Koedijk, S. O. Stefansson, A. Foss, S. Hjörleifsdóttir, G. Ó. Hreggvidsson, E. Otterlei, and A. Folkvord. A retrospective approach to fractionize variation in body mass of Atlantic cod Gadus morhua. Journal of Fish Biology, 78(1):251–264, January 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:2011:TEN

[377] A. M. Davis, R. G. Pearson, B. J. Pusey, C. Perna, D. L. Morgan, and D. Burrows. Trophic ecology of northern Australia’s terapontids: ontogenetic dietary shifts and feeding classification. Journal of Fish Biology, 78(1):265–286, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colonello:2011:RBL

[378] J. C. Colonello, M. L. García, and R. C. Menni. Reproductive biology of the lesser guitarfish Zapteryx brevirostris from the south-western Atlantic Ocean. Journal of Fish Biology, 78(1):287–302, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nock:2011:PSE

[379] C. J. Nock, J. R. Ovenden, G. L. Butler, I. Wooden, A. Moore, and P. R. Baverstock. Population structure, effective population size and adverse effects of stocking in the endangered Australian eastern freshwater cod maccullochella ikei. Journal of Fish Biology, 78(1):303–321, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sha:2011:DCN

[380] Z.-X. Sha, X.-H. Luo, X.-L. Liao, S.-L. Wang, Q.-L. Wang, and S.-L. Chen. Development and characterization of 60 novel EST-SSR markers in half-smooth tongue sole Cynoglossus semilaevis. Journal of Fish Biology, 78(1):322–331, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liso:2011:PST

[381] S. Liso, K. Ø. Gjelland, Y. S. Reshetnikov, and P.-A. Amundsen. A planktivorous specialist turns rapacious: piscivory in invading vendace Coregonus albula. Journal of Fish Biology, 78(1):332–337, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nichols:2011:FSL

[382] O. C. Nichols and U. T. Tscherter. Feeding of sea lampreys Petromyzon marinus on minke whales Balaenoptera acutorostrata in the St Lawrence Estuary, Canada. Journal of Fish Biology, 78(1):338–343, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos:2011:IMP

[383] J. A. A. Ramos, M. Barletta, D. V. Dantas, A. R. A. Lima, and M. F. Costa. Influence of moon phase on fish assemblages in estuarine mangrove tidal creeks. Journal of Fish Biology, 78(1):344–354, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2011:DBH

[384] P. J. Smith, D. Steinke, P. J. McMillan, A. L. Stewart, S. M. McVeagh, J. M. Diaz de Astarloa, D. Welsford, and R. D. Ward. DNA barcoding highlights a cryptic species of grenadier Macrourus in the Southern Ocean. Journal of Fish Biology, 78(1):355–365, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caires:2011:NRB

[385] R. A. Caires and J. L. Figueiredo. The northernmost record of Bassanago albescens and comments on the occurrence of Rhynchoconger guppyi (Teleostei: Anguilliformes: Congridae) along the Brazilian coast. Journal of Fish Biology, 78(1):366–372, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsiridou:2011:DTP

[386] A. Imsiridou, G. Minos, A. Gakopoulou, V. Katsares, T. Karidas, and G. Katselis. Discrimination of two picarel species Spicara flexuosa and Spicara maena (Pisces: Centracanthidae) based on mitochondrial DNA sequences. Journal of Fish Biology, 78(1):373–377, January 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohner:2011:HLW

[387] C. A. Rohner, A. J. Richardson, A. D. Marshall, S. J. Weeks, and S. J. Pierce. How large is the world’s largest fish? Measuring whale sharks Rhincodon typus with laser photogrammetry. Journal of Fish Biology, 78(1):378–385, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibanez:2011:FRM

[388] A. L. Ibáñez, M. González-Castro, and E. Pacheco-Almanzar. First record of Mugil hospes in the Gulf of Mexico and its identification from Mugil curema using ctenii. Journal of Fish Biology, 78(1):386–390, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2011:BRT

[389] Ian J. Winfield. Book review: Tropical Fishes of the East Indies — Edited by T. W. Pietsch. Journal of Fish Biology, 78(1):391, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2011:A

[390] John F. Craig. Announcement. Journal of Fish Biology, 78(1):392, January 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2011:EJU

[391] J. D. Metcalfe and J. F. Craig. Ethical justification for the use and treatment of fishes in research: an update. Journal of Fish Biology, 78 (2):393–394, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conrad:2011:BSF

[392] J. L. Conrad, K. L. Weinersmith, T. Brodin, J. B. Saltz, and A. Sih. Behavioural syndromes in fishes: a review with implications for ecology and fisheries management. Journal of Fish Biology, 78(2):395–435, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imholt:2011:DDT

[393] C. Imholt, I. A. Malcolm, P. J. Bacon, C. N. Gibbins, C. Soulsby, M. Miles, and R. J. Fryer. Does diurnal temperature variability affect growth in juvenile Atlantic salmon Salmo salar? Journal of Fish Biology, 78(2):436–448, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2011:FBP

[394] O. P. Jensen, S. Hansson, T. Didrikas, J. D. Stockwell, T. R. Hrabik, T. Axenrot, and J. F. Kitchell. Foraging, bioenergetic and predation constraints on diel vertical migration: field observations and modelling of reverse migration by young-of-the-year herring Clupea harengus. Journal of Fish Biology, 78(2):449–465, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Friess:2011:GES

[395] C. Friess and G. R. Sedberry. Genetic evidence for a single stock of the deep-sea teleost Beryx decadactylus in the North Atlantic Ocean as inferred from mtDNA control region analysis. Journal of Fish Biology, 78 (2):466–478, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chargulaf:2011:CSS

[396] C. A. Chargulaf, K. A. Townsend, and I. R. Tibbetts. Community structure of soft sediment pool fishes in Moreton Bay, Australia. Journal of Fish Biology, 78(2):479–494, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Obaza:2011:UDE

[397] A. Obaza, D. L. DeAngelis, and J. C. Trexler. Using data from an encounter sampler to model fish dispersal. Journal of Fish Biology, 78 (2):495–513, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bundy:2011:YWY

[398] A. Bundy, J. S. Link, B. E. Smith, and A. M. Cook. You are what you eat, whenever or wherever you eat it: an integrative analysis of fish food habits in Canadian and U.S.A. waters. Journal of Fish Biology, 78 (2):514–539, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verliin:2011:FSC

[399] A. Verliin, J. Kotta, H. Orav-Kotta, L. Saks, and M. Vetemaa. Food selection of Coregonus lavaretus in a brackish water ecosystem. Journal of Fish Biology, 78(2):540–551, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anttila:2011:ETF

[400] K. Anttila, E. Jokikokko, J. Erkinaro, M. Järvilehto, and S. Mänttäri. Effects of training on functional variables of muscles in reared Atlantic salmon Salmo salar smolts: connection to downstream migration pattern. Journal of Fish Biology, 78(2):552–566, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MHetli:2011:AGR

[401] M. M’Hetli, I. Ben Khemis, N. Hamza, B. Turki, and O. Turki. Allometric growth and reproductive biology traits of pikeperch Sander lucioperca at the southern edge of its range. Journal of Fish Biology, 78(2):567–579, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gisbert:2011:PFD

[402] E. Gisbert, I. Fernández, and C. A. Alvarez-González. Prolonged feed deprivation does not permanently compromise digestive function in migrating European glass eels Anguilla anguilla. Journal of Fish Biology, 78 (2):580–592, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Archard:2011:IEP

[403] G. A. Archard and V. A. Braithwaite. Increased exposure to predators increases both exploration and activity level in Brachyrhaphis episcopi. Journal of Fish Biology, 78(2):593–601, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Easton:2011:GAE

[404] A. A. Easton, H. K. Moghadam, R. G. Danzmann, and M. M. Ferguson. The genetic architecture of embryonic developmental rate and genetic covariation with age at maturation in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 78(2):602–623, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

manel-la:2011:DRO

[405] N. Plantalech manel la, C. M. Chittenden, F. Økland, E. B. Thorstad, J. G. Davidsen, R. Sivertsgård, R. S. McKinley, and B. Finstad. Does river of origin influence the early marine migratory performance of Salmo salar? Journal of Fish Biology, 78(2):624–634, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robertsen:2011:TGV

[406] G. Robertsen, E. Kvingedal, and S. Einum. Is there genetic variation in the response to competition intensity in juvenile brown trout Salmo trutta? Journal of Fish Biology, 78(2):635–646, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Onal:2011:EID

[407] U. Önal, İ. Çelik, E. Tokşen, A. Sepil, and E. Çaydan. Early infection of discus Symphysodon aequifasciatus altricial larvae by Sciadicleithrum variabilum (Monogenea). Journal of Fish Biology, 78(2):647–650, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Patrick:2011:ESF

[408] S. M. Patrick, E. White, R. W. Brill, and H. A. Shiels. The effect of stimulation frequency on the transmural ventricular monophasic action potential in yellowfin tuna Thunnus albacares. Journal of Fish Biology, 78(2):651–658, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagliano:2011:UIL

[409] M. Gagliano, A. K. Lema, M. Depczynski, and S. Whalan. Use it and lose it: lipofuscin accumulation in the midbrain of a coral reef fish. Journal of Fish Biology, 78(2):659–666, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2011:AAM

[410] W. G. Anderson and P. J. Allen. An alternative approach for the measurement of trimethylamine oxide in body fluid samples of elasmobranchs. Journal of Fish Biology, 78(2):667–672, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boto:2011:EEM

[411] L. Boto, O. Domínguez-Domínguez, and I. Doadrio. Exploring the effect of microsatellite size homoplasy on reconstruction of phylogenetic relationships of picote splitfin Zoogoneticus quitzeoensis. Journal of Fish Biology, 78(2):673–679, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mellor:2011:FMZ

[412] D. T. Mellor, C. M. Tarsiewicz, and R. C. Jordan. Female Maylandia zebrao prefer victorious males. Journal of Fish Biology, 78(2):680–687, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:R

[413] Anonymous. Referees 2010. Journal of Fish Biology, 78(2):688–695, February 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:2011:RSR

[414] R. J. Wootton and J. F. Craig. Reporting statistical results. Journal of Fish Biology, 78(3):697–699, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ota:2011:SSD

[415] K. Ota and M. Kohda. Social status-dependent nest choice of territorial males under reproductive parasitism in a Lake Tanganyika cichlid Telmatochromis vittatus. Journal of Fish Biology, 78(3):700–712, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Youngson:2011:SAN

[416] A. F. Youngson, S. B. Piertney, J. L. Thorley, I. A. Malcolm, and C. Soulsby. Spatial association of nest construction by brown trout Salmo trutta. Journal of Fish Biology, 78(3):713–725, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreau:2011:GHT

[417] D. T. R. Moreau, I. A. Fleming, G. L. Fletcher, and J. A. Brown. Growth hormone transgenesis does not influence territorial dominance or growth and survival of first-feeding Atlantic salmon Salmo salar in food-limited stream microcosms. Journal of Fish Biology, 78(3):726–740, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trip:2011:RBO

[418] E. D. L. Trip, K. D. Clements, D. Raubenheimer, and J. H. Choat. Reproductive biology of an odacine labrid, Odax pullus. Journal of Fish Biology, 78(3):741–761, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morales-Nin:2011:OHV

[419] B. Morales-Nin, A. Grau, S. Pérez-Mayol, E. Pastor, and M. Palmer. Oxytetracycline hydrochloride vital labelling revisited: the case of Dicentrarchus labrax and Diplodus puntazzo. Journal of Fish Biology, 78 (3):762–782, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nielsen:2011:RBC

[420] J. G. Nielsen and P. R. Møller. Revision of the bathyal cusk-eels of the genus Bassogigas (Ophidiidae) with description of a new species from off Guam, west Pacific Ocean. Journal of Fish Biology, 78(3):783–795, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polacik:2011:ASA

[421] M. Polačik, M. T. Donner, and M. Reichard. Age structure of annual Nothobranchius fishes in Mozambique: is there a hatching synchrony? Journal of Fish Biology, 78(3):796–809, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiegand:2011:MIT

[422] M. D. Wiegand, T. A. Johnston, L. R. Brown, S. B. Brown, J. M. Casselman, and W. C. Leggett. Maternal influences on thiamine status of walleye Sander vitreus ova. Journal of Fish Biology, 78(3):810–824, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreno:2011:TPL

[423] P. Moreno, G. Claramunt, and L. R. Castro. Transition period from larva to juvenile in anchoveta Engraulis ringens. Length or age related? Journal of Fish Biology, 78(3):825–837, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zion:2011:RAC

[424] B. Zion, A. Barki, J. Grinshpon, L. Rosenfeld, and I. Karplus. Retention of acoustic conditioning in St Peter’s fish Sarotherodon galilaeus. Journal of Fish Biology, 78(3):838–847, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2011:DMC

[425] T. Yamamoto, K. Kawai, and S. Oshima. Distribution of mucous cells on the body surface of Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 78(3):848–859, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2011:MLH

[426] Y.-J. Lin, S. Yalçin-Özdilek, Y. Iizuka, A. Gümüş, and W.-N. Tzeng. Migratory life history of European eel Anguilla anguilla from freshwater regions of the River Asi, southern Turkey and their high otolith Sr:Ca ratios. Journal of Fish Biology, 78(3):860–868, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fanelli:2011:FPD

[427] E. Fanelli, F. Badalamenti, G. D’Anna, C. Pipitone, E. Riginella, and E. Azzurro. Food partitioning and diet temporal variation in two coexisting sparids, Pagellus erythrinus and Pagellus acarne. Journal of Fish Biology, 78(3):869–900, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vigliano:2011:ISG

[428] F. A. Vigliano, L. Muñoz, D. Hernández, P. Cerutti, R. Bermúdez, and M. I. Quiroga. An immunohistochemical study of the gut neuroendocrine system in juvenile pejerrey Odontesthes bonariensis (Valenciennes). Journal of Fish Biology, 78(3):901–911, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chakona:2011:AET

[429] A. Chakona, E. R. Swartz, and K. Magellan. Aerial exposure tolerance of a newly discovered galaxiid. Journal of Fish Biology, 78(3):912–922, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2011:SSB

[430] B. R. Moore, J. Stapley, Q. Allsop, S. J. Newman, A. Ballagh, D. J. Welch, and R. J. G. Lester. Stock structure of blue threadfin Eleutheronema tetradactylum across northern Australia, as indicated by parasites. Journal of Fish Biology, 78(3):923–936, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trape:2011:FRM

[431] S. Trape and J.-D. Durand. First record of Mugil capurrii (Mugilidae, Perciformes) in the Gulf of Guinea. Journal of Fish Biology, 78(3):937– 940, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delpiani:2011:FRD

[432] S. M. Delpiani, M. Y. Deli Antoni, S. A. Barbini, and D. E. Figueroa. First record of a dicephalic specimen of tope Galeorhinus galeus (Elasmobranchii: Triakidae). Journal of Fish Biology, 78(3):941–944, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miles:2011:UAS

[433] N. G. Miles and R. J. West. The use of an aeration system to prevent thermal stratification of a freshwater impoundment and its effect on downstream fish assemblages. Journal of Fish Biology, 78(3):945– 952, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vasek:2011:ONN

[434] M. Vašek, T. Jůza, M. Čech, M. Kratochvíl, M. Prchalová, J. Frouzová, M. Říha, M. Tušer, J. Seďa, and J. Kubečka. The occurrence of nonnative tubenose goby Proterorhinus semilunaris in the pelagic 0+ year fish assemblage of a central European reservoir. Journal of Fish Biology, 78(3):953–961, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karlsson:2011:CAO

[435] A. Karlsson, B. O. Rosseland, H. Thorarensen, and A. Kiessling. Changes in arterial oxygen tension and physiological status in resting, unrestrained Arctic charr Salvelinus alpinus (L.) exposed to mild hypoxia and hyperoxia. Journal of Fish Biology, 78(3):962–966, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sandkam:2011:EWD

[436] B. A. Sandkam and R. C. Fuller. The effects of water depth and light on oviposition and egg cannibalism in the bluefin killifish Lucania goodei. Journal of Fish Biology, 78(3):967–972, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Su:2011:MDF

[437] J. Su, J. Dong, T. Huang, R. Zhang, C. Yang, and J. Heng. Myeloid differentiation factor 88 gene is involved in antiviral immunity in grass carp Ctenopharyngodon idella. Journal of Fish Biology, 78(3):973–979, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2011:BRB

[438] Ian J. Winfield. Book review: Behaviour, Migrations, Distribution and Stocks of Sturgeons in the Volga–Caspian Basin — Edited by R. P. Khodorevskaya, G. I. Ruban and D. S. Pavlov. Journal of Fish Biology, 78(3):980–981, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodda:2011:BRD

[439] Kate Rodda. Book review: Development of Non-Teleost Fishes — Edited by Y. W. Kunz, C. A. Luer and B. G. Kapoor. Journal of Fish Biology, 78(3):981–982, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:Ca

[440] Anonymous. Corrigendum. Journal of Fish Biology, 78(3):983–984, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2011:CSE

[441] P. J. Smith, D. Steinke, P. J. McMillan, A. L. Stewart, S. M. McVeagh, J. M. Diaz de Astarloa, D. Welsford, and R. D. Ward. The current status of elasmobranchs: Biology, fisheries and conservation. Journal of Fish Biology, 78(3):985, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCa

[442] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 78(3):986, March 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lakra:2011:DCC

[443] W. S. Lakra and M. Goswami. Development and characterization of a continuous cell line PSCF from Puntius sophore. Journal of Fish Biology, 78(4):987–1001, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bal:2011:EWT

[444] G. Bal, E. Rivot, E. Prévost, C. Piou, and J. L. Baglinière. Effect of water temperature and density of juvenile salmonids on growth of youngof-the-year Atlantic salmon Salmo salar. Journal of Fish Biology, 78(4): 1002–1022, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gerry:2011:NAQ

[445] S. P. Gerry, J. Wang, and D. J. Ellerby. A new approach to quantifying morphological variation in bluegill Lepomis macrochirus. Journal of Fish Biology, 78(4):1023–1034, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cloutier:2011:EOA

[446] R. Cloutier, J. Lambrey de Souza, H. I. Browman, and A. B. Skiftesvik. Early ontogeny of the Atlantic halibut Hippoglossus hippoglossus head. Journal of Fish Biology, 78(4):1035–1053, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jovanovic:2011:ISC

[447] B. Jovanović, F. W. Goetz, G. W. Goetz, and D. Palić. Immunological stimuli change expression of genes and neutrophil function in fathead minnow Pimephales promelas Rafinesque. Journal of Fish Biology, 78 (4):1054–1072, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clevestam:2011:TSS

[448] P. D. Clevestam, M. Ogonowski, N. B. Sjöberg, and H. Wickström. Too short to spawn? Implications of small body size and swimming distance on successful migration and maturation of the European eel Anguilla anguilla. Journal of Fish Biology, 78(4):1073–1089, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamer:2011:CLE

[449] P. A. Hamer, S. Acevedo, G. P. Jenkins, and A. Newman. Connectivity of a large embayment and coastal fishery: spawning aggregations in one bay source local and broad-scale fishery replenishment. Journal of Fish Biology, 78(4):1090–1109, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjornsson:2011:SCM

[450] B. Björnsson, P. Reynisson, J. Solmundsson, and H. Valdimarsson. Seasonal changes in migratory and predatory activity of two species of gadoid preying on inshore northern shrimp Pandalus borealis. Journal of Fish Biology, 78(4):1110–1131, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaemingk:2011:ESD

[451] M. A. Kaemingk, J. C. Jolley, D. W. Willis, and B. D. S. Graeb. Exploring spatial distributions of larval yellow perch Perca flavescens, bluegill Lepomis macrochirus and their prey in relation to wind. Journal of Fish Biology, 78(4):1132–1151, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalogianni:2011:CRS

[452] E. Kalogianni, M. Alexis, C. Tsangaris, M. Abraham, S. E. Wendelaar Bonga, Y. Iger, E. H. Van Ham, and M. Th. Stoumboudi. Cellular responses in the skin of the gilthead sea bream Sparus aurata L. and the sea bass Dicentrarchus labrax (L.) exposed to high ammonia. Journal of Fish Biology, 78(4):1152–1169, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carreon-Martinez:2011:USC

[453] L. Carreon-Martinez, T. B. Johnson, S. A. Ludsin, and D. D. Heath. Utilization of stomach content DNA to determine diet diversity in piscivorous fishes. Journal of Fish Biology, 78(4):1170–1182, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2011:GNS

[454] J. A. Smith, L. J. Baumgartner, I. M. Suthers, and M. D. Taylor. Generalist niche, specialist strategy: the diet of an Australian percichthyid. Journal of Fish Biology, 78(4):1183–1199, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landsman:2011:ETC

[455] S. J. Landsman, A. J. Gingerich, D. P. Philipp, and C. D. Suski. The effects of temperature change on the hatching success and larval survival of largemouth bass Micropterus salmoides and smallmouth bass Micropterus dolomieu. Journal of Fish Biology, 78(4):1200–1212, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Navia:2011:SOT

[456] A. F. Navia, A. Torres, P. A. Mejía-Falla, and A. Giraldo. Sexual, ontogenetic, temporal and spatial effects on the diet of Urotrygon rogersi (Elasmobranchii: Myliobatiformes). Journal of Fish Biology, 78(4):1213– 1224, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chu:2011:COE

[457] W. Y. Chu, J. Chen, R. X. Zhou, F. L. Zhao, T. Meng, D. X. Chen, X. X. Nong, Z. Liu, S. Q. Lu, and J. S. Zhang. Characterization and ontogenetic expression analysis of the myosin light chains from the fast white muscle of mandarin fish Siniperca chuatsi. Journal of Fish Biology, 78(4):1225–1238, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vitorino:2011:MCC

[458] C. A. Vitorino, I. L. Souza, J. N. Rosa, G. T. Valente, C. Martins, and P. C. Venere. Molecular cytogenetics and its contribution to the understanding of the chromosomal diversification in Hoplias malabaricus (Characiformes). Journal of Fish Biology, 78(4):1239–1248, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2011:LHB

[459] I. P. Jacobsen and M. B. Bennett. Life history of the blackspotted whipray Himantura astra. Journal of Fish Biology, 78(4):1249–1268, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logez:2011:VBT

[460] M. Logez and D. Pont. Variation of brown trout Salmo trutta youngof-the-year growth along environmental gradients in Europe. Journal of Fish Biology, 78(4):1269–1276, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2011:BRG

[461] Lei Yang. Book review: Gonorynchiformes and Ostariophysan Relationships: a Comprehensive Review — Edited by T. Grande, F. J. Poyato-Ariza and R. Diogo. Journal of Fish Biology, 78(4):1277–1278, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:2011:BRE

[462] Audrey J. Geffen. Book review: Early Life History of Marine Fishes — Edited by B. S. Miller and A. W. Kendall. Journal of Fish Biology, 78 (4):1279–1280, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:CSEa

[463] Anonymous. The current status of elasmobranchs: Biology, fisheries and conservation. Journal of Fish Biology, 78(4):1281, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCb

[464] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 78(4):1282, April 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2011:CNR

[465] I. J. Harrison, P. Chakrabarty, J. Freyhof, and J. F. Craig. Correct nomenclature and recommendations for preserving and cataloguing voucher material and genetic sequences. Journal of Fish Biology, 78 (5):1283–1290, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeLaCruzAguero:2011:JLC

[466] J. De La Cruz Agüero. José Luis Castro Aguirre, mainstay of contemporary ichthyology in Mexico. Journal of Fish Biology, 78(5):1291–1293, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Litvin:2011:LCD

[467] S. Y. Litvin, V. G. Guida, and M. P. Weinstein. Lipid class dynamics and storage depots in juvenile weakfish Cynoscion regalis and their application to condition assessment. Journal of Fish Biology, 78(5):1294–1310, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bono:2011:ALH

[468] L. M. Bono, O. Rios-Cardenas, and M. R. Morris. Alternative life histories in Xiphophorus multilineatus: evidence for different ages at sexual maturity and growth responses in the wild. Journal of Fish Biology, 78 (5):1311–1322, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Masuda:2011:OSS

[469] R. Masuda. Ontogeny of swimming speed, schooling behaviour and jellyfish avoidance by Japanese anchovy Engraulis japonicus. Journal of Fish Biology, 78(5):1323–1335, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deAstarloa:2011:MMM

[470] J. M. Díaz de Astarloa, S. I. Bezzi, M. González Castro, E. Mabragaña, D. Hernández, S. M. Delpiani, D. E. Figueroa, M. B. Cousseau, M. Y. Deli Antoni, and L. Tringali. Morphological, morphometric, meristic and osteological evidence for two species of hake (Actinopterygii: Gadiformes: Merluccius) in Argentinean waters. Journal of Fish Biology, 78(5):1336–1358, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abadi:2011:UOF

[471] Z. Taghizadeh Rahmat Abadi, S. Khodabandeh, B. Abtahi, G. Charmantier, and M. Charmantier-Daures. Ultrastructure and osmoregulatory function of the kidney in larvae of the Persian sturgeon Acipenser persicus. Journal of Fish Biology, 78(5):1359–1374, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huveneers:2011:QME

[472] C. Huveneers, N. M. Otway, R. G. Harcourt, and M. Ellis. Quantification of the maternal–embryonal nutritional relationship of elasmobranchs: case study of wobbegong sharks (genus Orectolobus). Journal of Fish Biology, 78(5):1375–1389, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2011:EMM

[473] J. G. Davidsen, B. Finstad, F. Økland, E. B. Thorstad, T. A. Mo, and A. H. Rikardsen. Early marine migration of European silver eel Anguilla anguilla in northern Norway. Journal of Fish Biology, 78(5):1390–1404, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gutteridge:2011:AOB

[474] A. N. Gutteridge, M. B. Bennett, C. Huveneers, and I. R. Tibbetts. Assessing the overlap between the diet of a coastal shark and the surrounding prey communities in a sub-tropical embayment. Journal of Fish Biology, 78(5):1405–1422, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2011:PRF

[475] T. Tomiyama, M. Watanabe, G. Kawata, and K. Ebe. Post-release feeding and growth of hatchery-reared Japanese flounder Paralichthys olivaceus: relevance to stocking effectiveness. Journal of Fish Biology, 78 (5):1423–1436, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meager:2011:BRH

[476] J. J. Meager, P. Rodewald, P. Domenici, A. Fernö, T. Järvi, J. E. Skjæraasen, and G. K. Sverdrup. Behavioural responses of hatcheryreared and wild cod Gadus morhua to mechano-acoustic predator signals. Journal of Fish Biology, 78(5):1437–1450, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solem:2011:MDP

[477] Ø. Solem and O. K. Berg. Morphological differences in parr of Atlantic salmon Salmo salar from three regions in Norway. Journal of Fish Biology, 78(5):1451–1469, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salze:2011:MDL

[478] G. Salze, S. R. Craig, B. H. Smith, E. P. Smith, and E. McLean. Morphological development of larval cobia Rachycentron canadum and the influence of dietary taurine supplementation. Journal of Fish Biology, 78(5):1470–1491, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kielbassa:2011:IIA

[479] J. Kielbassa, S. Charles, and M. L. Delignette-Muller. The importance of incorporating age and sex when backcalculating length in bullhead Cottus gobio. Journal of Fish Biology, 78(5):1492–1507, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:2011:PAI

[480] L. Zheng, H. Tanaka, and S. Abe. Proteomic analysis of inviable salmonid hybrids between female masu salmon Oncorhynchus masou masou and male rainbow trout Oncorhynchus mykiss during early embryogenesis. Journal of Fish Biology, 78(5):1508–1528, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakabe:2011:IFI

[481] R. Sakabe, J. M. Lyle, and C. M. Crawford. The influence of freshwater inflows on spawning success and early growth of an estuarine resident fish species, Acanthopagrus butcheri. Journal of Fish Biology, 78(5):1529– 1544, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catalan:2011:TEA

[482] I. A. Catalán, A. Tejedor, F. Alemany, and P. Reglero. Trophic ecology of Atlantic bluefin tuna Thunnus thynnus larvae. Journal of Fish Biology, 78(5):1545–1560, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jenkins:2011:FEK

[483] G. P. Jenkins, A. Syme, and P. I. Macreadie. Feeding ecology of King George whiting Sillaginodes punctatus (Perciformes) recruits in seagrass and unvegetated habitats. Does diet reflect habitat utilization? Journal of Fish Biology, 78(5):1561–1573, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banon:2011:FRL

[484] R. Bañón and V. Santás. First record of Lagocephalus laevigatus (Tetraodontiformes, Tetraodontidae) from Galician waters (north-west Spain), a northernmost occurrence in the north-east Atlantic Ocean. Journal of Fish Biology, 78(5):1574–1578, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayden:2011:LCV

[485] B. Hayden, I. Coscia, and S. Mariani. Low cytochrome b variation in bream Abramis brama. Journal of Fish Biology, 78(5):1579–1587, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heule:2011:ODE

[486] C. Heule and W. Salzburger. The ontogenetic development of egg-spots in the haplochromine cichlid fish Astatotilapia burtoni. Journal of Fish Biology, 78(5):1588–1593, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:CSEb

[487] Anonymous. The current status of elasmobranchs: Biology, fisheries and conservation. Journal of Fish Biology, 78(5):1594, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCc

[488] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 78(5):1595, May 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahnesjo:2011:BSP

[489] I. Ahnesjö and J. F. Craig. The biology of Syngnathidae: pipefishes, seadragons and seahorses. Journal of Fish Biology, 78(6):1597–1602, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2011:EOS

[490] A. B. Wilson and J. W. Orr. The evolutionary origins of Syngnathidae: pipefishes and seahorses. Journal of Fish Biology, 78(6):1603–1623, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mobley:2011:GGS

[491] K. B. Mobley, C. M. Small, and A. G. Jones. The genetics and genomics of Syngnathidae: pipefishes, seahorses and seadragons. Journal of Fish Biology, 78(6):1624–1646, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosenqvist:2011:SSM

[492] G. Rosenqvist and A. Berglund. Sexual signals and mating patterns in Syngnathidae. Journal of Fish Biology, 78(6):1647–1661, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scobell:2011:RES

[493] S. K. Scobell and D. S. MacKenzie. Reproductive endocrinology of Syngnathidae. Journal of Fish Biology, 78(6):1662–1680, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vincent:2011:CMS

[494] A. C. J. Vincent, S. J. Foster, and H. J. Koldewey. Conservation and management of seahorses and other Syngnathidae. Journal of Fish Biology, 78(6):1681–1724, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kvarnemo:2011:EPN

[495] C. Kvarnemo, K. B. Mobley, C. Partridge, A. G. Jones, and I. Ahnesjö. Evidence of paternal nutrient provisioning to embryos in broad-nosed pipefish Syngnathus typhle. Journal of Fish Biology, 78(6):1725–1737, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodall:2011:HCP

[496] L. C. Woodall, H. J. Koldewey, and P. W. Shaw. Historical and contemporary population genetic connectivity of the European short-snouted seahorse Hippocampus hippocampus and implications for management. Journal of Fish Biology, 78(6):1738–1756, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin-Smith:2011:PII

[497] K. M. Martin-Smith. Photo-identification of individual weedy seadragons Phyllopteryx taeniolatus and its application in estimating population dynamics. Journal of Fish Biology, 78(6):1757–1768, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caldwell:2011:ATS

[498] I. R. Caldwell, M. Correia, J. Palma, and A. C. J. Vincent. Advances in tagging syngnathids, with the effects of dummy tags on behaviour of Hippocampus guttulatus. Journal of Fish Biology, 78(6):1769–1785, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanWassenbergh:2011:WLF

[499] S. Van Wassenbergh, G. Roos, P. Aerts, A. Herrel, and D. Adriaens. Why the long face? A comparative study of feeding kinematics of two pipefishes with different snout lengths. Journal of Fish Biology, 78(6): 1786–1798, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leysen:2011:MSF

[500] H. Leysen, J. Christiaens, B. De Kegel, M. N. Boone, L. Van Hoorebeke, and D. Adriaens. Musculoskeletal structure of the feeding system and implications of snout elongation in Hippocampus reidi and Dunckerocampus dactyliophorus. Journal of Fish Biology, 78(6):1799–1823, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2011:FSS

[501] T. M. Smith, J. S. Hindell, G. P. Jenkins, R. M. Connolly, and M. J. Keough. Fine-scale spatial and temporal variations in diets of the pipefish Stigmatopora nigra within seagrass patches. Journal of Fish Biology, 78 (6):1824–1832, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sogabe:2011:OSM

[502] A. Sogabe and I. Ahnesjö. The ovarian structure and mode of egg production in two polygamous pipefishes: a link to mating pattern. Journal of Fish Biology, 78(6):1833–1846, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2011:RBF

[503] I. Braga Goncalves, I. Ahnesjö, and C. Kvarnemo. The relationship between female body size and egg size in pipefishes. Journal of Fish Biology, 78(6):1847–1854, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundin:2011:SNP

[504] J. Sundin, Ö. Jacobsson, A. Berglund, and G. Rosenqvist. Straight-nosed pipefish Nerophis ophidion and broad-nosed pipefish Syngnathus typhle avoid eelgrass overgrown with filamentous algae. Journal of Fish Biology, 78(6):1855–1860, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindqvist:2011:MBN

[505] C. Lindqvist, J. Sundin, A. Berglund, and G. Rosenqvist. Male broadnosed pipefish Syngnathus typhle do not locate females by smell. Journal of Fish Biology, 78(6):1861–1867, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:CSEc

[506] Anonymous. The current status of elasmobranchs: Biology, fisheries and conservation. Journal of Fish Biology, 78(6):1868, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCd

[507] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 78(6):1869, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2011:TBB

[508] John F. Craig. Timothy bracegirdle bagenal 1925–2011 (fish ecologist). Journal of Fish Biology, 78(7):1871–1873, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baudron:2011:IWN

[509] A. R. Baudron, C. L. Needle, and C. T. Marshall. Implications of a warming North Sea for the growth of haddock Melanogrammus aeglefinus. Journal of Fish Biology, 78(7):1874–1889, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giannetto:2011:PSM

[510] D. Giannetto, G. La Porta, G. Maio, E. Pizzul, P. Turin, and M. Lorenzoni. Proposed standard mass equations for European chub Leuciscus cephalus in Italy. Journal of Fish Biology, 78(7):1890–1899, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feindel:2011:GDT

[511] N. J. Feindel, T. J. Benfey, and E. A. Trippel. Gonadal development of triploid Atlantic cod Gadus morhua. Journal of Fish Biology, 78(7): 1900–1912, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lek:2011:WED

[512] E. Lek, D. V. Fairclough, M. E. Platell, K. R. Clarke, J. R. Tweedley, and I. C. Potter. To what extent are the dietary compositions of three abundant, co-occurring labrid species different and related to latitude, habitat, body size and season? Journal of Fish Biology, 78(7):1913– 1943, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yavno:2011:RGN

[513] S. Yavno and L. D. Corkum. Round goby Neogobius melanostomus attraction to conspecific and heterospecific egg odours. Journal of Fish Biology, 78(7):1944–1953, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vollset:2011:III

[514] K. W. Vollset and K. M. Bailey. Interplay of individual interactions and turbidity affects the functional response of three-spined sticklebacks Gasterosteus aculeatus. Journal of Fish Biology, 78(7):1954–1964, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russon:2011:EQS

[515] I. J. Russon and P. S. Kemp. Experimental quantification of the swimming performance and behaviour of spawning run river lamprey Lampetra fluviatilis and European eel Anguilla anguilla. Journal of Fish Biology, 78(7):1965–1975, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dempson:2011:RTM

[516] J. B. Dempson, M. J. Robertson, C. J. Pennell, G. Furey, M. Bloom, M. Shears, L. M. N. Ollerhead, K. D. Clarke, R. Hinks, and G. J. Robertson. Residency time, migration route and survival of Atlantic salmon Salmo salar smolts in a Canadian fjord. Journal of Fish Biology, 78 (7):1976–1992, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sokolowska:2011:AGR

[517] E. Sokołowska and D. P. Fey. Age and growth of the round goby Neogobius melanostomus in the Gulf of Gdańsk several years after invasion. Is the Baltic Sea a new promised land? Journal of Fish Biology, 78 (7):1993–2009, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuji:2011:MEJ

[518] T. Fuji, A. Kasai, K. W. Suzuki, M. Ueno, and Y. Yamashita. Migration ecology of juvenile temperate seabass Lateolabrax japonicus: a carbon stable-isotope approach. Journal of Fish Biology, 78(7):2010–2025, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harry:2011:LHE

[519] A. V. Harry, W. G. Macbeth, A. N. Gutteridge, and C. A. Simpfendorfer. The life histories of endangered hammerhead sharks (Carcharhiniformes, Sphyrnidae) from the east coast of Australia. Journal of Fish Biology, 78(7):2026–2051, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cazorla:2011:DFS

[520] A. Lopez Cazorla, R. E. Pettigrosso, L. Tejera, and R. Camina. Diet and food selection by Ramnogaster arcuata (Osteichthyes, Clupeidae). Journal of Fish Biology, 78(7):2052–2066, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartulovic:2011:RCM

[521] V. Bartulović, J. Dulčić, S. Matić-Skoko, and B. Glamuzina. Reproductive cycles of Mugil cephalus, Liza ramada and Liza aurata (Teleostei: Mugilidae). Journal of Fish Biology, 78(7):2067–2073, June 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeBruyn:2011:SED

[522] M. De Bruyn, L. R. Parenti, and G. R. Carvalho. Successful extraction of DNA from archived alcohol-fixed white-eye fish specimens using an ancient DNA protocol. Journal of Fish Biology, 78(7):2074–2079, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2011:DTR

[523] B. P. Wang, J. Chen, J. D. Liu, N. Liu, and Q. X. Yu. Discriminating two races of Liobagrus marginatoides by cytogenetic analysis. Journal of Fish Biology, 78(7):2080–2084, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gotanda:2011:HMB

[524] K. M. Gotanda, E. E. Reardon, and L. J. Chapman. Hypoxia and male behaviour in an African cichlid Pseudocrenilabrus multicolor victoriae. Journal of Fish Biology, 78(7):2085–2092, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2011:BRD

[525] S. G. Jones and A. Douglas. Book review: Do Fish Feel Pain? Edited by V. Braithwaite. Journal of Fish Biology, 78(7):2093–2095, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:CSEd

[526] Anonymous. The current status of elasmobranchs: Biology, fisheries and conservation. Journal of Fish Biology, 78(7):2097, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCe

[527] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 78(7):2098, June 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2011:SA

[528] J. F. Craig. Submissions and authorship. Journal of Fish Biology, 79(1): 1–2, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Padros:2011:MHS

[529] F. Padrós, M. Villalta, E. Gisbert, and A. Estévez. Morphological and histological study of larval development of the Senegal sole Solea senegalensis: an integrative study. Journal of Fish Biology, 79(1):3–32, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valentincic:2011:ODC

[530] T. Valentincic, P. Miklavc, S. Kralj, and V. Zgonik. Olfactory discrimination of complex mixtures of amino acids by the black bullhead Ameiurus melas. Journal of Fish Biology, 79(1):33–52, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pietsch:2011:VTP

[531] C. Pietsch, N. Neumann, T. Preuer, and W. Kloas. In vivo treatment with progestogens causes immunosuppression of carp Cyprinus carpio leucocytes by affecting nitric oxide production and arginase activity. Journal of Fish Biology, 79(1):53–69, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebner:2011:EGP

[532] B. C. Ebner, B. Kroll, P. Godfrey, P. A. Thuesen, T. Vallance, B. Pusey, G. R. Allen, T. S. Rayner, and C. N. Perna. Is the elusive Gymnothorax polyuranodon really a freshwater moray? Journal of Fish Biology, 79(1): 70–79, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brian:2011:PDI

[533] J. V. Brian, N. Beresford, L. Margiotta-Casaluci, and J. P. Sumpter. Preliminary data on the influence of rearing temperature on the growth and reproductive status of fathead minnows Pimephales promelas. Journal of Fish Biology, 79(1):80–88, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Demain:2011:DFN

[534] D. K. Demain, A. Gallego, A. Jaworski, I. G. Priede, and E. G. Jones. Diet and feeding niches of juvenile Gadus morhua, Melanogrammus aeglefinus and Merlangius merlangus during the settlement transition in the northern North Sea. Journal of Fish Biology, 79(1):89–111, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Budarf:2011:ISG

[535] A. Chelsky Budarf, D. D. Burfeind, W. K. W. Loh, and I. R. Tibbetts. Identification of seagrasses in the gut of a marine herbivorous fish using DNA barcoding and visual inspection techniques. Journal of Fish Biology, 79(1):112–121, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsson:2011:GPS

[536] J. Olsson, K. Mo, A.-B. Florin, T. Aho, and N. Ryman. Genetic population structure of perch Perca fluviatilis along the Swedish coast of the Baltic Sea. Journal of Fish Biology, 79(1):122–137, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2011:DTS

[537] J. Roberts, J. C. Xavier, and D. J. Agnew. The diet of toothfish species Dissostichus eleginoides and Dissostichus mawsoni with overlapping distributions. Journal of Fish Biology, 79(1):138–154, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rourke:2011:VSE

[538] M. L. Rourke, H. C. McPartlan, B. A. Ingram, and A. C. Taylor. Variable stocking effect and endemic population genetic structure in Murray cod Maccullochella peelii. Journal of Fish Biology, 79(1):155–177, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2011:EPH

[539] X. Z. Zhang, L. P. Dai, Z. H. Wu, J. C. Jian, and Y. S. Lu. Expression pattern of heat shock protein 90 gene of humphead snapper Lutjanus sanguineus during pathogenic Vibrio harveyi stress. Journal of Fish Biology, 79(1):178–193, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daBarbiano:2011:ISC

[540] L. Alberici da Barbiano, A. S. Aspbury, C. C. Nice, and C. R. Gabor. The impact of social context on male mate preference in a unisexual– bisexual mating complex. Journal of Fish Biology, 79(1):194–204, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jamieson:2011:SIB

[541] A. J. Jamieson, T. Fujii, P. M. Bagley, and I. G. Priede. Scavenging interactions between the arrow tooth eel Synaphobranchus kaupii and the Portuguese dogfish Centroscymnus coelolepis. Journal of Fish Biology, 79(1):205–216, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuya:2011:RIL

[542] F. Tuya, T. Wernberg, and M. S. Thomsen. The relative influence of local to regional drivers of variation in reef fishes. Journal of Fish Biology, 79 (1):217–234, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horodysky:2011:CMR

[543] A. Z. Horodysky, R. W. Brill, P. G. Bushnell, J. A. Musick, and R. J. Latour. Comparative metabolic rates of common western North Atlantic Ocean sciaenid fishes. Journal of Fish Biology, 79(1):235–255, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2011:IST

[544] T. G. Pottinger, A. Cook, M. D. Jürgens, M. Sebire, P. A. Henrys, I. Katsiadaki, J. L. Balaam, A. J. Smith, and P. Matthiessen. Indices of stress in three-spined sticklebacks Gasterosteus aculeatus in relation to extreme weather events and exposure to wastewater effluent. Journal of Fish Biology, 79(1):256–279, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coxon:2011:SFV

[545] S. E. Coxon and W. Davison. Structure and function of the velar muscle in the New Zealand hagfish Eptatretus cirrhatus: response to temperature change and hypoxia. Journal of Fish Biology, 79(1):280–289, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2011:FSJ

[546] M. Jönsson, S. Hylander, L. Ranåker, P. A. Nilsson, and C. Brönmark. Foraging success of juvenile pike Esox lucius depends on visual conditions and prey pigmentation. Journal of Fish Biology, 79(1):290–297, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2011:RCR

[547] L. J. Williams, A. M. Cicia, G. B. Pellegrin, K. M. Smith, and J. A. Sulikowski. The reproductive cycle of the roundel skate Raja texana. Journal of Fish Biology, 79(1):298–305, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2011:BREa

[548] Ian J. Winfield. Book review: Ecosystem-based Fisheries Management: Confronting Trade-Offs — Edited by J. S. Link. Journal of Fish Biology, 79(1):306–307, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2011:BRS

[549] L. A. Jawad. Book review: Sharks and Their Relatives II: Biodiversity, Adaptive Physiology, and Conservation — Edited by J. C. Carrier, J. A. Musick and M. R. Heithaus. Journal of Fish Biology, 79(1):308–309, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frommen:2011:BRT

[550] J. G. Frommen. Book review: Tinbergen’s Legacy in Behaviour: 60 years of Landmark Stickleback Papers — Edited by F. A. von Hippel. Journal of Fish Biology, 79(1):310–311, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCf

[551] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 79(1):312, July 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Westphal:2011:MPS

[552] M. F. Westphal, S. R. Morey, J. C. Uyeda, and T. J. Morgan. Molecular phylogeny of the subfamily Amphistichinae (Teleostei: Embiotocidae) reveals parallel divergent evolution of red pigmentation in two rapidly evolving lineages of sand-dwelling surfperch. Journal of Fish Biology, 79 (2):313–330, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kirchhoff:2011:EIR

[553] N. T. Kirchhoff, T. D’Antignana, M. J. Leef, C. J. Hayward, R. J. Wilkinson, and B. F. Nowak. Effects of immunostimulants on ranched southern bluefin tuna Thunnus maccoyii: immune response, health and performance. Journal of Fish Biology, 79(2):331–355, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Renkawitz:2011:FEE

[554] M. D. Renkawitz and T. F. Sheehan. Feeding ecology of early marine phase Atlantic salmon Salmo salar post-smolts. Journal of Fish Biology, 79(2):356–373, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheehan:2011:STS

[555] T. F. Sheehan, M. D. Renkawitz, and R. W. Brown. Surface trawl survey for U.S. origin Atlantic salmon Salmo salar. Journal of Fish Biology, 79 (2):374–398, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luchetti:2011:COS

[556] E. A. Luchetti, S. P. Iglésias, and D. Y. Sellos. Chimaera opalescens n. sp., a new chimaeroid (Chondrichthyes: Holocephali) from the northeastern Atlantic Ocean. Journal of Fish Biology, 79(2):399–417, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcic:2011:NEC

[557] Z. Marčić, I. Buj, A. Duplić, M. Ćaleta, P. Mustafić, D. Zanella, P. Zupančič, and M. Mrakovčić. A new endemic cyprinid species from the Danube drainage. Journal of Fish Biology, 79(2):418–430, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linlokken:2011:IAB

[558] A. N. Linløkken and T. Hesthagen. The interactions of abiotic and biotic factors influencing perch Perca fluviatilis and roach Rutilus rutilus populations in small acidified boreal lakes. Journal of Fish Biology, 79 (2):431–448, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jeffries:2011:TCB

[559] K. M. Jeffries, S. G. Hinch, M. R. Donaldson, M. K. Gale, J. M. Burt, L. A. Thompson, A. P. Farrell, D. A. Patterson, and K. M. Miller. Temporal changes in blood variables during final maturation and senescence in male sockeye salmon Oncorhynchus nerka: reduced osmoregulatory ability can predict mortality. Journal of Fish Biology, 79(2):449–465, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lagrue:2011:FII

[560] C. Lagrue, D. W. Kelly, A. Hicks, and R. Poulin. Factors influencing infection patterns of trophically transmitted parasites among a fish community: host diet, host–parasite compatibility or both? Journal of Fish Biology, 79(2):466–485, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:2011:MCE

[561] M. Cui, Q. Z. Zhang, Z. J. Yao, and Z. H. Zhang. Molecular cloning and expression analysis of heat-shock protein 70 in orange-spotted grouper Epinephelus coioides following heat shock and Vibrio alginolyticus challenge. Journal of Fish Biology, 79(2):486–501, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liebich:2011:WCE

[562] T. Liebich, S. D. McCormick, D. Kircheis, K. Johnson, R. Regal, and T. Hrabik. Water chemistry and its effects on the physiology and survival of Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 79(2): 502–519, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bassett:2011:HRU

[563] D. Bassett and J. Montgomery. Home range use and movement patterns of the yellow moray eel Gymnothorax prasinus. Journal of Fish Biology, 79(2):520–525, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bassett:2011:FRP

[564] D. K. Bassett and J. C. Montgomery. The feeding response of Pseudophycas barbata to multisensory prey cues in a low light environment.

Journal of Fish Biology, 79(2):526–532, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dupuch:2011:BRP

[565] A. Dupuch, Y. Paradis, and P. Magnan. Behavioural responses of prey fishes to habitat complexity and predation risk induce bias in minnow trap catches. Journal of Fish Biology, 79(2):533–538, August 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowen:2011:EMI

[566] L. Bowen, J. D. Kiernan, K. Eder, I. Werner, C. M. Leutenegger, M. A. Adkison, R. P. Hedrick, and M. L. Johnson. Euthanization methods influence cytokine mRNA expression levels in age 0 year Oncorhynchus mykiss. Journal of Fish Biology, 79(2):539–545, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DiDario:2011:NRC

[567] F. Di Dario, A. C. Petry, M. M. Mincarone, M. M. S. Pereira, and R. M. Dos Santos. New records of coastal fishes in the northern Rio de Janeiro State, Brazil, with comments on the biogeography of the southwestern Atlantic Ocean. Journal of Fish Biology, 79(2):546–555, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lymbery:2011:BRF

[568] Alan J. Lymbery. Book review: Fish Defenses, Volume 2: Pathogens, Parasites and Predators — Edited by G. Zaccone, C. Perriere, A. Mathis and B. G. Kapoor. Journal of Fish Biology, 79(2):556–557, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2011:BRS

[569] Kyle A. Young. Book review: Salmonid Fisheries: Freshwater Habitat Management — Edited by P. Kemp. Journal of Fish Biology, 79(2):558– 559, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:FMCg

[570] Anonymous. Fish migration in the 21st Century: Opportunities and challenges. Journal of Fish Biology, 79(2):560, August 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moghadam:2011:WGD

[571] H. K. Moghadam, M. M. Ferguson, and R. G. Danzmann. Whole genome duplication: challenges and considerations associated with sequence orthology assignment in salmoninae. Journal of Fish Biology, 79(3):561– 574, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eriksen:2011:SMT

[572] M. S. Eriksen, G. Færevik, S. Kittilsen, M. I. McCormick, B. Damsgård, V. A. Braithwaite, B. O. Braastad, and M. Bakken. Stressed mothers — troubled offspring: a study of behavioural maternal effects in farmed Salmo salar. Journal of Fish Biology, 79(3):575–586, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoogenboom:2011:ICB

[573] M. O. Hoogenboom, J. D. Armstrong, M. S. Miles, T. Burton, T. G. G. Groothuis, and N. B. Metcalfe. Implantation of cocoa butter reduces egg and hatchling size in Salmo trutta. Journal of Fish Biology, 79(3): 587–596, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fridman:2011:OCL

[574] S. Fridman, J. E. Bron, and K. J. Rana. Ontogenetic changes in location and morphology of chloride cells during early life stages of the Nile tilapia Oreochromis niloticus adapted to fresh and brackish water. Journal of Fish Biology, 79(3):597–614, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grandcourt:2011:RBI

[575] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Reproductive biology and implications for management of the painted sweetlips Diagramma pictum in the southern Arabian Gulf. Journal of Fish Biology, 79(3):615–632, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broderick:2011:GPS

[576] D. Broderick, J. R. Ovenden, R. C. Buckworth, S. J. Newman, R. J. G. Lester, and D. J. Welch. Genetic population structure of grey mackerel Scomberomorus semifasciatus in northern Australia. Journal of Fish Biology, 79(3):633–661, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Potter:2011:IDU

[577] I. C. Potter, B. M. Chuwen, S. A. Hesp, N. G. Hall, S. D. Hoeksema, D. V. Fairclough, and T. M. Rodwell. Implications of the divergent use of a suite of estuaries by two exploited marine fish species. Journal of Fish Biology, 79(3):662–691, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaslet:2011:SIA

[578] A. Vaslet, C. France, D. L. Phillips, I. C. Feller, and C. C. Baldwin. Stable-isotope analyses reveal the importance of seagrass beds as feeding areas for juveniles of the speckled worm eel Myrophis punctatus (Teleostei: Ophichthidae) in Florida. Journal of Fish Biology, 79(3): 692–706, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krimmer:2011:BPR

[579] A. N. Krimmer, A. J. Paul, A. Hontela, and J. B. Rasmussen. Behavioural and physiological responses of brook trout Salvelinus fontinalis to midwinter flow reduction in a small ice-free mountain stream. Journal of Fish Biology, 79(3):707–725, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernies:2011:IRS

[580] D. Bernies, A. Brinker, and A. Daugschies. An invasion record for the swimbladder nematode Anguillicoloides crassus in European eel Anguilla anguilla in a deep cold-monomictic lake, from invasion to steady state. Journal of Fish Biology, 79(3):726–746, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2011:MCF

[581] J. F. Li, H. X. Ai, J. Zhang, M. X. Du, Z. Zhang, J. X. Zhang, and S. Q. Zhang. Molecular cloning, functional characterization and phylogenetic analysis of TRAIL in Japanese pufferfish Takifugu rubripes. Journal of Fish Biology, 79(3):747–760, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becker:2011:ASS

[582] A. Becker, A. K. Whitfield, P. D. Cowley, J. Järnegren, and T. F. Næsje. An assessment of the size structure, distribution and behaviour of fish populations within a temporarily closed estuary using dual frequency identification sonar (DIDSON). Journal of Fish Biology, 79(3):761–775, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dzieweczynski:2011:SBD

[583] T. L. Dzieweczynski and J. A. Crovo. Shyness and boldness differences across contexts in juvenile three-spined stickleback Gasterosteus aculeatus from an anadromous population. Journal of Fish Biology, 79(3): 776–788, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DaSilva:2011:TDC

[584] D. Da Silva, D. C. F. Costa, C. M. Alves, B. A. Block, and A. M. Landeira-Fernandez. Temperature dependence of cardiac sarcoplasmic reticulum Ca^{2+}-ATPase from rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 79(3):789–800, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mirimin:2011:QLI

[585] L. Mirimin, D. O’Keeffe, A. Ruggiero, M. Bolton-Warberg, S. Vartia, and R. FitzGerald. A quick, least-invasive, inexpensive and reliable method for sampling Gadus morhua postlarvae for genetic analysis. Journal of Fish Biology, 79(3):801–805, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mincarone:2011:FRF

[586] M. M. Mincarone, M. Mwale, and B. Fernholm. First record and further description of the Cape hagfish Myxine capensis (Myxinidae) off Mozambique, western Indian Ocean. Journal of Fish Biology, 79(3):806–811, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binning:2011:NLM

[587] S. A. Binning and C. J. Fulton. Non-lethal measurement of pectoral fin aspect ratio in coral-reef fishes. Journal of Fish Biology, 79(3):812–818, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Masuda:2011:BRF

[588] Reiji Masuda. Book review: Fish Locomotion: an Eco-Ethological Perspective — Edited by P. Domenici and B. G. Kapoor. Journal of Fish Biology, 79(3):819–820, September 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Handy:2011:EMN

[589] R. D. Handy, G. Al-Bairuty, A. Al-Jubory, C. S. Ramsden, D. Boyle, B. J. Shaw, and T. B. Henry. Effects of manufactured nanomaterials on fishes: a target organ and body systems physiology approach. Journal of Fish Biology, 79(4):821–853, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaneko:2011:CLB

[590] G. Kaneko, S. Furukawa, Y. Kurosu, T. Yamada, H. Takeshima, M. Nishida, T. Mitsuboshi, T. Otaka, K. Shirasu, T. Koda, Y. Takemasa, S. Aki, T. Mochizuki, H. Fukushima, Y. Fukuda, S. Kinoshita, S. Asakawa, and S. Watabe. Correlation with larval body size of mRNA levels of growth hormone, growth hormone receptor I and insulin-like growth factor I in larval torafugu Takifugu rubripes. Journal of Fish Biology, 79(4):854–874, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez:2011:IPR

[591] M. S. Rodriguez, H. Ortega, and R. Covain. Intergeneric phylogenetic relationships in catfishes of the Loricariinae (Siluriformes: Loricariidae), with the description of Fonchiiloricaria nanodon: a new genus and species from Peru. Journal of Fish Biology, 79(4):875–895, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laiz-Carrion:2011:CSD

[592] R. Laiz-Carrión, J. M. Quintanilla, J. M. Mercado, and A. García. Combined study of daily growth variability and nitrogen–carbon isotopic signature analysis of schooling Sardina pilchardus larvae. Journal of Fish Biology, 79(4):896–914, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trejo-Martinez:2011:RAY

[593] J. Trejo-Martínez, T. Brulé, A. Mena-Loría, T. Colás-Marrufo, and M. Sánchez-Crespo. Reproductive aspects of the yellowtail snapper Ocyurus chrysurus from the southern Gulf of Mexico. Journal of Fish Biology, 79(4):915–936, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKellar:2011:EFI

[594] A. E. McKellar and A. P. Hendry. Environmental factors influencing adult sex ratio in Poecilia reticulata: laboratory experiments. Journal of Fish Biology, 79(4):937–953, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2011:SFN

[595] W. T. White and S. P. Iglésias. Squalus formosus, a new species of spurdog shark (Squaliformes: Squalidae), from the western North Pacific Ocean. Journal of Fish Biology, 79(4):954–968, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulsen:2011:BRT

[596] S. B. Poulsen, L. F. Jensen, K. S. Nielsen, H. Malte, K. Aarestrup, and J. C. Svendsen. Behaviour of rainbow trout Oncorhynchus mykiss presented with a choice of normoxia and stepwise progressive hypoxia. Journal of Fish Biology, 79(4):969–979, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Semmens:2011:EIA

[597] D. Semmens and S. E. Swearer. Extended incubation affects larval morphology, hatching success and starvation resistance in a terrestrially spawning fish, Galaxias maculatus (Jenyns 1842). Journal of Fish Biology, 79(4):980–990, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bogutskaya:2011:NAS

[598] N. G. Bogutskaya, M. A. Zuykov, A. M. Naseka, and E. B. Anderson. Normal axial skeleton structure in common roach Rutilus rutilus (Actinopterygii: Cyprinidae) and malformations due to radiation contamination in the area of the Mayak (Chelyabinsk Province, Russia) nuclear plant. Journal of Fish Biology, 79(4):991–1016, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nannini:2011:ISV

[599] M. A. Nannini, D. H. Wahl, D. P. Philipp, and S. J. Cooke. The influence of selection for vulnerability to angling on foraging ecology in largemouth bass Micropterus salmoides. Journal of Fish Biology, 79(4):1017–1028, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zuniga-Vega:2011:MRV

[600] J. J. Zúñiga-Vega, M. Suárez-Rodríguez, H. Espinosa-Pérez, and J. B. Johnson. Morphological and reproductive variation among populations of the Pacific molly Poecilia butleri. Journal of Fish Biology, 79(4):

1029–1046, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huey:2011:EMH

[601] J. A. Huey, A. M. Baker, and J. M. Hughes. Evidence for multiple historical colonizations of an endoreic drainage basin by an Australian freshwater fish. Journal of Fish Biology, 79(4):1047–1067, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mesa:2011:SWH

[602] M. G. Mesa, R. J. Magie, E. S. Copeland, and H. E. Christiansen. Surgical wound healing in radio-tagged adult Pacific lamprey Entosphenus tridentatus held on different substrata. Journal of Fish Biology, 79(4): 1068–1075, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haugland:2011:AFL

[603] T. Haugland, G. Rudolfsen, L. Figenschou, and I. Folstad. Is the adipose fin and the lower jaw (kype) related to social dominance in male Arctic charr Salvelinus alpinus? Journal of Fish Biology, 79(4):1076–1083, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Link:2011:BRM

[604] Jason Link. Book review: Mismanagement of Marine Fisheries — Edited by A. Longhurst. Journal of Fish Biology, 79(4):1084–1085, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2011:BRE

[605] Mike Elliott. Book review: Ecology of Estuarine Fishes: Temperate Waters of the Western North Atlantic — Edited by K. W. Able and M. P. Fahay. Journal of Fish Biology, 79(4):1086–1088, October 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deLeaniz:2011:WCJ

[606] C. Garcia de Leaniz. Walter Crozier Jordan (Bill) 1962–2011. Journal of Fish Biology, 79(5):1089–1093, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milosevic:2011:GPE

[607] D. Milošević, K. A. Winkler, D. Marić, and S. Weiss. Genotypic and phenotypic evaluation of Rutilus spp. from Skadar, Ohrid and Prespa Lakes supports revision of endemic as well as taxonomic status of several taxa. Journal of Fish Biology, 79(5):1094–1110, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borrell:2011:HFC

[608] Y. J. Borrell, C. E. Carleos, J. A. Sánchez, E. Vázquez, V. Gallego, J. F. Asturiano, and G. Blanco. Heterozygosity–fitness correlations in the gilthead sea bream Sparus aurata using microsatellite loci from unknown and gene-rich genomic locations. Journal of Fish Biology, 79(5):1111– 1129, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roach:2011:EBM

[609] J. P. Roach, K. C. Hall, and M. K. Broadhurst. Effects of barotrauma and mitigation methods on released Australian bass Macquaria novemaculeata. Journal of Fish Biology, 79(5):1130–1145, November 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deslauriers:2011:IFL

[610] D. Deslauriers and J. D. Kieffer. The influence of flume length and group size on swimming performance in shortnose sturgeon Acipenser brevirostrum. Journal of Fish Biology, 79(5):1146–1155, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berg:2011:PWL

[611] O. K. Berg, G. Rød, Ø. Solem, and A. G. Finstad. Pre-winter lipid stores in brown trout Salmo trutta along altitudinal and latitudinal gradients. Journal of Fish Biology, 79(5):1156–1166, November 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vehanen:2011:BTS

[612] T. Vehanen and A. Huusko. Brown trout Salmo trutta express different morphometrics due to divergence in the rearing environment. Journal of Fish Biology, 79(5):1167–1181, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsu:2011:RAB

[613] T.-H. Hsu, A. G. Guillén Madrid, C. P. Burridge, H.-Y. Cheng, and J.-C. Gwo. Resolution of the Acanthopagrus black seabream complex based on mitochondrial and amplified fragment-length polymorphism analyses. Journal of Fish Biology, 79(5):1182–1192, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Last:2011:BPA

[614] P. R. Last and W. T. White. Biogeographic patterns in the Australian chondrichthyan fauna. Journal of Fish Biology, 79(5):1193–1213, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shaddick:2011:HZB

[615] K. Shaddick, C. P. Burridge, D. R. Jerry, T. S. Schwartz, K. Truong, D. M. Gilligan, and L. B. Beheregaray. A hybrid zone and bidirectional introgression between two catadromous species: Australian bass Macquaria novemaculeata and estuary perch Macquaria colonorum. Journal of Fish Biology, 79(5):1214–1235, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delling:2011:TRG

[616] B. Delling, M. Noren, S. O. Kullander, and J. A. González. Taxonomic review of the genus Trisopterus (Teleostei: Gadidae) with recognition of the capelan Trisopterus capelanus as a valid species. Journal of Fish Biology, 79(5):1236–1260, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabragana:2011:CEC

[617] E. Mabragaña, D. E. Figueroa, L. B. Scenna, J. M. Díaz de Astarloa, J. H. Colonello, and G. Delpiani. Chondrichthyan egg cases from the south-west Atlantic Ocean. Journal of Fish Biology, 79(5):1261–1290, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torisawa:2011:SBJ

[618] S. Torisawa, H. Fukuda, K. Suzuki, and T. Takagi. Schooling behaviour of juvenile Pacific bluefin tuna Thunnus orientalis depends on their vision development. Journal of Fish Biology, 79(5):1291–1303, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lord:2011:FLA

[619] C. Lord, H. Tabouret, F. Claverie, C. Pécheyran, and P. Keith. Femtosecond laser ablation ICP-MS measurement of otolith Sr:Ca and Ba:Ca composition reveal differential use of freshwater habitats for three amphidromous Sicyopterus (Teleostei: Gobioidei: Sicydiinae) species. Journal of Fish Biology, 79(5):1304–1321, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:2011:EFT

[620] T. D. Clark, M. R. Donaldson, S. M. Drenner, S. G. Hinch, D. A. Patterson, J. Hills, V. Ives, J. J. Carter, S. J. Cooke, and A. P. Farrell. The efficacy of field techniques for obtaining and storing blood samples from fishes. Journal of Fish Biology, 79(5):1322–1333, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roy:2011:IGB

[621] T. S. C. Roy and A. Gopalakrishnan. Identification of groupers based on pyloric caeca differentiation. Journal of Fish Biology, 79(5):1334–1339, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2011:GSD

[622] S. Yamamoto, S. Kitamura, H. Sakano, and K. Morita. Genetic structure and diversity of Japanese Kokanee Oncorhynchus nerka stocks as revealed by microsatellite and mitochondrial DNA markers. Journal of Fish Biology, 79(5):1340–1349, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsu:2011:TSC

[623] T.-H. Hsu, Y. T. Adiputra, C. P. Burridge, and J.-C. Gwo. Two spinefoot colour morphs: mottled spinefoot Siganus fuscescens and white-spotted spinefoot Siganus canaliculatus are synonyms. Journal of Fish Biology, 79(5):1350–1355, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hermann:2011:AOR

[624] C. M. Hermann, K. M. Sefc, and S. Koblmüller. Ancient origin and recent divergence of a haplochromine cichlid lineage from isolated water bodies in the East African Rift system. Journal of Fish Biology, 79(5): 1356–1369, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rueda:2011:ICE

[625] E. C. Rueda, P. Amavet, F. Brancolini, J. Sommer, and G. Ortí. Isolation and characterization of eight polymorphic microsatellite markers for the migratory characiform fish, Salminus brasiliensis. Journal of Fish Biology, 79(5):1370–1375, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:2011:ARP

[626] D. P. Robinson, W. Baverstock, A. Al-Jaru, K. Hyland, and K. A. Khazanehdari. Annually recurring parthenogenesis in a zebra shark Stegostoma fasciatum. Journal of Fish Biology, 79(5):1376–1382, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinagre:2011:DNO

[627] C. Vinagre, A. Silva, M. Lara, and H. N. Cabral. Diet and niche overlap of southern populations of brill Scophthalmus rhombus and turbot Scophthalmus maximus. Journal of Fish Biology, 79(5):1383–1391, November 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magurran:2011:BDF

[628] A. E. Magurran, S. Khachonpisitsak, and A. B. Ahmad. Biological diversity of fish communities: pattern and process. Journal of Fish Biology, 79(6):1393–1412, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skelton:2011:WTT

[629] P. H. Skelton and E. R. Swartz. Walking the tightrope: trends in African freshwater systematic ichthyology. Journal of Fish Biology, 79(6):1413– 1435, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulet:2011:TTF

[630] N. Poulet, L. Beaulaton, and S. Dembski. Time trends in fish populations in metropolitan France: insights from national monitoring data. Journal of Fish Biology, 79(6):1436–1452, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barausse:2011:LTC

[631] A. Barausse, A. Michieli, E. Riginella, L. Palmeri, and C. Mazzoldi. Long-term changes in community composition and life-history traits in a highly exploited basin (northern Adriatic Sea): the role of environment and anthropogenic pressures. Journal of Fish Biology, 79(6):1453–1486, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dudgeon:2011:ARF

[632] D. Dudgeon. Asian river fishes in the Anthropocene: threats and conservation challenges in an era of rapid environmental change. Journal of Fish Biology, 79(6):1487–1524, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart-Koster:2011:FRT

[633] B. Stewart-Koster, J. D. Olden, M. J. Kennard, B. J. Pusey, E. L. Boone, M. Douglas, and S. Jackson. Fish response to the temporal hierarchy of the natural flow regime in the Daly River, northern Australia. Journal of Fish Biology, 79(6):1525–1544, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2011:SMC

[634] Y.-H. Zhao, R. E. Gozlan, and C.-G. Zhang. Out of sight out of mind: current knowledge of Chinese cave fishes. Journal of Fish Biology, 79(6): 1545–1562, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pino-Del-Carpio:2011:CGK

[635] A. Pino-Del-Carpio, A. Villarroya, A. H. Ariño, J. Puig, and R. Miranda. Communication gaps in knowledge of freshwater fish biodiversity: implications for the management and conservation of Mexican biosphere reserves. Journal of Fish Biology, 79(6):1563–1591, December 2011. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fobert:2011:HCH

[636] E. Fobert, M. G. Fox, M. Ridgway, and G. H. Copp. Heated competition: how climate change will affect non-native pumpkinseed Lepomis gibbosus and native perch Perca fluviatilis interactions in the U.K. Journal of Fish Biology, 79(6):1592–1607, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sayer:2011:TCC

[637] C. D. Sayer, G. H. Copp, D. Emson, M. J. Godard, G. Zieba, and K. J. Wesley. Towards the conservation of crucian carp Carassius carassius: understanding the extent and causes of decline within part of its native English range. Journal of Fish Biology, 79(6):1608–1624, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roux:2011:SSA

[638] M. J. Roux, R. F. Tallman, and C. W. Lewis. Small-scale Arctic charr Salvelinus alpinus fisheries in Canada’s Nunavut: management challenges and options. Journal of Fish Biology, 79(6):1625–1647, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Railsback:2011:IFB

[639] S. F. Railsback and B. C. Harvey. Importance of fish behaviour in modelling conservation problems: food limitation as an example. Journal of Fish Biology, 79(6):1648–1662, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:2011:PSF

[640] I. G. Cowx and M. Portocarrero Aya. Paradigm shifts in fish conservation: moving to the ecosystem services concept. Journal of Fish Biology, 79(6):1663–1680, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe-McConnell:2011:EDC

[641] Ro Lowe-McConnell. Eric David Le Cren 1922–2011 (Freshwater Biological Association). Journal of Fish Biology, 79(7):1681–1684, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2011:MIU

[642] A. A. Santos, R. C. Gutierre, M. M. Antoniazzi, M. J. T. Ranzani-Paiva, M. R. R. Silva, C. T. F. Oshima, and M. I. Egami. Morphocytochemical, immunohistochemical and ultrastructural characterization of the head kidney of fat snook Centropomus parallelus. Journal of Fish Biology, 79(7):1685–1707, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arronte:2011:NDL

[643] J. C. Arronte, R. Bañón, D. T. G. Quigley, J. A. Pis-Millán, and J. Heredia. New data on Lepidion schmidti (Gadiformes: Moridae) from the north-east Atlantic Ocean. Journal of Fish Biology, 79(7):1708–1721, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ho:2011:RAL

[644] H.-C. Ho, B. Séret, and K.-T. Shao. Records of anglerfishes (Lophiiformes: Lophiidae) from the western South Pacific Ocean, with descriptions of two new species. Journal of Fish Biology, 79(7):1722–1745, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conway:2011:NSG

[645] K. W. Conway, D. R. Edds, J. Shrestha, and R. L. Mayden. A new species of gravel-dwelling loach (Ostariophysi: Nemacheilidae) from the Nepalese Himalayan foothills. Journal of Fish Biology, 79(7):1746–1759, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scharnweber:2011:DNO

[646] K. Scharnweber, M. Plath, K. O. Winemiller, and M. Tobler. Dietary niche overlap in sympatric asexual and sexual livebearing fishes Poecilia spp. Journal of Fish Biology, 79(7):1760–1773, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2011:BSP

[647] R. W. Blake and K. H. S. Chan. Biomechanics of swimming in the pufferfish Diodon holocanthus: propulsive momentum enhancement is an adaptation for thrust production in an undulatory median and pairedfin swimmer. Journal of Fish Biology, 79(7):1774–1794, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reinaldo:2011:FSC

[648] M. Ocampo Reinaldo, R. González, and M. A. Romero. Feeding strategy and cannibalism of the Argentine hake Merluccius hubbsi. Journal of Fish Biology, 79(7):1795–1814, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefebure:2011:TDG

[649] R. Lefébure, S. Larsson, and P. Byström. A temperature-dependent growth model for the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 79(7):1815–1827, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schrey:2011:HBP

[650] A. W. Schrey, R. Boley, and E. J. Heist. Hybridization between pallid sturgeon Scaphirhynchus albus and shovelnose sturgeon Scaphirhynchus platorynchus. Journal of Fish Biology, 79(7):1828–1850, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henriquez:2011:MPC

[651] V. P. Henriquez, M. T. Gonzalez, R. Licandeo, and J. Carvajal. Metazoan parasite communities of rock cod Eleginops maclovinus along southern Chilean coast and their use as biological tags at a local spatial scale. Journal of Fish Biology, 79(7):1851–1865, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haynes:2011:FEY

[652] P. S. Haynes, D. Brophy, F. De Raedemaecker, and D. McGrath. The feeding ecology of 0 year-group turbot Scophthalmus maximus and brill Scophthalmus rhombus on Irish west coast nursery grounds. Journal of Fish Biology, 79(7):1866–1882, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:2011:VAS

[653] E. Robinson, A. R. Jerrett, S. E. Black, and W. Davison. Visual acuity of snapper Pagrus auratus: effect of size and spectral composition. Journal of Fish Biology, 79(7):1883–1894, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Specziar:2011:SDP

[654] A. Specziár. Size-dependent prey selection in piscivorous pikeperch Sander lucioperca and Volga pikeperch Sander volgensis shaped by bimodal prey size distribution. Journal of Fish Biology, 79(7):1895–1917, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeRaedemaecker:2011:MPA

[655] F. De Raedemaecker, I. O’Connor, D. Brophy, and A. Black. Macrobenthic prey availability and the potential for food competition between 0 year group Pleuronectes platessa and Limanda limanda. Journal of Fish Biology, 79(7):1918–1939, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Troynikov:2011:PEP

[656] V. S. Troynikov, H. K. Gorfine, L. Ložys, Z. Pūtys, E. Jakubavičiūtė, and R. W. Day. Parameterization of European perch Perca fluviatilis length-at-age data using stochastic Gompertz growth models. Journal of Fish Biology, 79(7):1940–1949, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Apostolidis:2011:GDA

[657] A. P. Apostolidis, M. Th. Stoumboudi, E. Kalogianni, G. Cote, and L. Bernatchez. Genetic divergence among native trout Salmo trutta populations from southern Balkans based on mitochondrial DNA and microsatellite variation. Journal of Fish Biology, 79(7):1950–1960, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kocabas:2011:ESG

[658] M. Kocabas, N. Basçinar, S. Atasaral Sahin, and I. Okumus. Effect of salinity on growth of Abant trout Salmo trutta abanticus. Journal of Fish Biology, 79(7):1961–1967, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chargulaf:2011:DSA

[659] C. A. Chargulaf, N. C. Krück, and I. R. Tibbetts. Does sympatry affect trophic resource use in congeneric tidepool fishes? A tale of two gobies Favonigobius lentiginosus and Favonigobius exquisitus. Journal of Fish Biology, 79(7):1968–1983, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinheiro:2011:RFS

[660] H. T. Pinheiro, C. E. L. Ferreira, J.-C. Joyeux, R. G. Santos, and P. A. Horta. Reef fish structure and distribution in a south-western Atlantic Ocean tropical island. Journal of Fish Biology, 79(7):1984–2006, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohata:2011:OAP

[661] R. Ohata, R. Masuda, and Y. Yamashita. Ontogeny of antipredator performance in hatchery-reared Japanese anchovy Engraulis japonicus larvae exposed to visual or tactile predators in relation to turbidity. Journal of Fish Biology, 79(7):2007–2018, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crespel:2011:QGA

[662] A. Crespel, L. Bernatchez, D. Garant, and C. Audet. Quantitative genetic analysis of the physiological stress response in three strains of brook charr Salvelinus fontinalis and their hybrids. Journal of Fish Biology, 79(7):2019–2033, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garrido:2011:ISV

[663] S. Garrido and A. G. Murta. Interdecadal and spatial variations of diet composition in horse mackerel Trachurus trachurus. Journal of Fish Biology, 79(7):2034–2042, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamazaki:2011:FFN

[664] Y. Yamazaki, R. Yokoyama, T. Nagai, and A. Goto. Formation of a fluvial non-parasitic population of Lethenteron camtschaticum as the first step in petromyzontid speciation. Journal of Fish Biology, 79(7):2043– 2059, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reum:2011:LCM

[665] J. C. P. Reum. Lipid correction model of carbon stable isotopes for a cosmopolitan predator, spiny dogfish Squalus acanthias. Journal of Fish Biology, 79(7):2060–2066, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leknes:2011:CRC

[666] I. L. Leknes. Carbohydrates in rod and cone light absorbing segments in the firemouth cichlid Thorichthys meeki retina. Journal of Fish Biology, 79(7):2067–2073, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Charlier:2011:CGE

[667] J. Charlier, A. Palmé, L. Laikre, J. Andersson, and N. Ryman. Census (N_{C}) and genetically effective (n_{e}) population size in a lake-resident population of brown trout Salmo trutta. Journal of Fish Biology, 79(7): 2074–2082, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ji:2011:BAG

[668] C. Ji, A. H. Li, J. Xie, D. F. Zhang, and J. Y. Liu. Bactericidal activity of grass carp Ctenopharyngodon idella C9-deficient serum against Aeromonas hydrophila. Journal of Fish Biology, 79(7):2083–2086, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Willette:2011:FRT

[669] D. A. Willette, M. D. Santos, and M. A. Aragon. First report of the Taiwan sardinella Sardinella hualiensis (Clupeiformes: Clupeidae) in the Philippines. Journal of Fish Biology, 79(7):2087–2094, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatarenkov:2011:EHL

[670] A. Tatarenkov, S. M. Q. Lima, and J. C. Avise. Extreme homogeneity and low genetic diversity in Kryptolebias ocellatus from south-eastern Brazil suggest a recent foundation for this androdioecious fish population. Journal of Fish Biology, 79(7):2095–2105, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shields:2011:BRA

[671] Brian Shields. Book review: Atlantic Salmon Ecology — Edited by O. Aas, S. Einum, A. Klemetsen and J. Skurdal. Journal of Fish Biology, 79(7):2106–2107, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2011:BREb

[672] Ian J. Winfield. Book review: Ecology of Atlantic Salmon and Brown Trout: Habitat as a Template for Life Histories — Edited by B. Jonsson and N. Jonsson. Journal of Fish Biology, 79(7):2108–2109, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2011:Cb

[673] Anonymous. Corrigendum. Journal of Fish Biology, 79(7):2110, December 2011. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gjedrem:2012:GVT

[674] T. Gjedrem and B. O. Rosseland. Genetic variation for tolerance to acidic water in salmonids. Journal of Fish Biology, 80(1):1–14, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carballo:2012:SPC

[675] M. C. Carballo, F. Cremonte, G. T. Navone, and J. T. Timi. Similarity in parasite community structure may be used to trace latitudinal migrations of Odontesthes smitti along Argentinean coasts. Journal of Fish Biology, 80(1):15–28, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kramer:2012:DES

[676] A. Kramer, M. Kovačić, and R. A. Patzner. Dentition of eight species of Mediterranean Sea Gobiidae: do dentition characters of gobies reflect phylogenetic relationships? Journal of Fish Biology, 80(1):29–48, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:2012:PVA

[677] M. J. Weber, K. D. Rounds, and M. L. Brown. Phenotypic variation and associated predation risk of juvenile common carp Cyprinus carpio.

Journal of Fish Biology, 80(1):49–60, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeFaveri:2012:PIF

[678] J. DeFaveri, L. N. Zanella, D. Zanella, M. Mrakovčić, and J. Merilä. Phylogeography of isolated freshwater three-spined stickleback Gasterosteus aculeatus populations in the Adriatic Sea basin. Journal of Fish Biology, 80(1):61–85, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:2012:OSM

[679] A. J. Clark and A. P. Summers. Ontogenetic scaling of the morphology and biomechanics of the feeding apparatus in the Pacific hagfish Eptatretus stoutii. Journal of Fish Biology, 80(1):86–99, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2012:STE

[680] H. Zhao, J. Cui, Y. Wang, X. Liu, D. Zhao, and J. Duan. Spatialtemporal expression of pum1 and pum2 in medaka Oryzias latipes. Journal of Fish Biology, 80(1):100–109, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2012:PTS

[681] E. P. Williams, A. C. Peer, T. J. Miller, D. H. Secor, and A. R. Place. A phylogeny of the temperate seabasses (Moronidae) characterized by a translocation of the mt-nd 6 gene. Journal of Fish Biology, 80(1):110– 130, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitano:2012:RSD

[682] J. Kitano, S. Mori, and C. L. Peichel. Reduction of sexual dimorphism in stream-resident forms of three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 80(1):131–146, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stacey:2012:MPE

[683] N. E. Stacey, G. J. Van Der Kraak, and K. H. Olsén. Male primer endocrine responses to preovulatory female cyprinids under natural conditions in Sweden. Journal of Fish Biology, 80(1):147–165, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karlsen:2012:DCM

[684] J. D. Karlsen and N. G. Andersen. Does the cylinder model of gastric evacuation predict observed evacuation of mixed meals of prey of contrasting geometries in a piscivorous fish? Journal of Fish Biology, 80 (1):166–180, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andreou:2012:IGT

[685] D. Andreou, J. Vacquie-Garcia, J. Cucherousset, S. Blanchet, R. E. Gozlan, and G. Loot. Individual genetic tagging for teleosts: an empirical validation and a guideline for ecologists. Journal of Fish Biology, 80 (1):181–194, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen-Ankins:2012:ETP

[686] S. Allen-Ankins, R. J. Stoffels, P. A. Pridmore, and M. T. Vogel. The effects of turbidity, prey density and environmental complexity on the feeding of juvenile Murray cod Maccullochella peelii. Journal of Fish Biology, 80(1):195–206, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunton:2012:GMS

[687] K. J. Dunton, D. Chapman, A. Jordaan, K. Feldheim, S. J. O’Leary, K. A. McKown, and M. G. Frisk. Genetic mixed-stock analysis of Atlantic sturgeon Acipenser oxyrinchus oxyrinchus in a heavily exploited marine habitat indicates the need for routine genetic monitoring. Journal of Fish Biology, 80(1):207–217, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ning:2012:MPB

[688] N. S. P. Ning, S. Hladyz, B. Gawne, S. Maffei, A. Price, and D. L. Nielsen. Morphological, physiological and behavioural response patterns of carp gudgeon Hypseleotris spp. to food deprivation: implications for assessing health. Journal of Fish Biology, 80(1):218–224, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Syvaranta:2012:SIC

[689] J. Syväranta, C. Harrod, L. Kubicek, V. Cappanera, and J. D. R. Houghton. Stable isotopes challenge the perception of ocean sunfish Mola mola as obligate jellyfish predators. Journal of Fish Biology, 80 (1):225–231, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See comments [994, 993].

Jawad:2012:BRE

[690] L. A. Jawad. Book review: Ecology and Evolution of Poeciliid Fishes — Edited by J. P. Evans, A. Pilastro and I. Schlupp. Journal of Fish Biology, 80(1):232–233, January 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kornis:2012:TYI

[691] M. S. Kornis, N. Mercado-Silva, and M. J. Vander Zanden. Twenty years of invasion: a review of round goby Neogobius melanostomus biology, spread and ecological implications. Journal of Fish Biology, 80(2):235– 285, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deng:2012:EAF

[692] J. M. Deng, K. S. Mai, Q. H. Ai, W. B. Zhang, X. J. Wang, W. Xu, Z. G. Liufu, Y. H. Cai, and W. Chen. Effects of antinutritional factors on plasma lipoprotein levels in Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 80(2):286–300, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:2012:DPS

[693] S. F. Perry and L. Rivero-Lopez. Does the presence of a seawater gill morphology induced by dietary salt loading affect Cl^{−} uptake and acid–base regulation in freshwater rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 80(2):301–311, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babu:2012:DCN

[694] V. Sarath Babu, V. Chandra, K. S. N. Nambi, S. A. Majeed, G. Taju, M. S. Patole, and A. S. S. Hameed. Development and characterization of novel cell lines from Etroplus suratensis and their applications in virology, toxicology and gene expression. Journal of Fish Biology, 80(2):312–334, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McGrath:2012:SDL

[695] P. E. McGrath and E. J. Hilton. Sexual dimorphism in longnose gar Lepisosteus osseus. Journal of Fish Biology, 80(2):335–345, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soeparno:2012:RBP

[696] Soeparno, Y. Nakamura, T. Shibuno, and K. Yamaoka. Relationship between pelagic larval duration and abundance of tropical fishes on temperate coasts of Japan. Journal of Fish Biology, 80(2):346–357, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casselman:2012:UMH

[697] M. T. Casselman, K. Anttila, and A. P. Farrell. Using maximum heart rate as a rapid screening tool to determine optimum temperature for aerobic scope in Pacific salmon Oncorhynchus spp. Journal of Fish Biology, 80(2):358–377, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2012:VIM

[698] T. Yamamoto and S. Kitanishi. Variable incidences and morphological characteristics of female masu salmon Oncorhynchus masou with growth hormone pseudogene. Journal of Fish Biology, 80(2):378–386, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Etheridge:2012:PTU

[699] E. C. Etheridge, C. E. Adams, C. W. Bean, N. C. Durie, A. R. D. Gowans, C. Harrod, A. A. Lyle, P. S. Maitland, and I. J. Winfield. Are phenotypic traits useful for differentiating among a priori Coregonus taxa? Journal of Fish Biology, 80(2):387–407, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honda:2012:MPE

[700] K. Honda, T. Arai, S. Kobayashi, Y. Tsuda, and K. Miyashita. Migratory patterns of exotic brown trout Salmo trutta in south-western Hokkaido, Japan, on the basis of otolith Sr:Ca ratios and acoustic telemetry. Journal of Fish Biology, 80(2):408–426, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2012:PSB

[701] A. T. Correia, P. Gomes, J. M. S. Gonçalves, K. Erzini, and P. A. Hamer. Population structure of the black seabream Spondyliosoma cantharus along the south-west Portuguese coast inferred from otolith chemistry. Journal of Fish Biology, 80(2):427–443, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burt:2012:PII

[702] J. M. Burt, S. G. Hinch, and D. A. Patterson. Parental identity influences progeny responses to incubation thermal stress in sockeye salmon Onchorhynchus nerka. Journal of Fish Biology, 80(2):444–462, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:R

[703] Anonymous. Referees 2011. Journal of Fish Biology, 80(2):463–470, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:Ca

[704] Anonymous. Corrigendum. Journal of Fish Biology, 80(2):471, February 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sui:2012:GCF

[705] J. Sui, D. Y. Ma, Q. H. Liu, S. H. Xu, Z. Z. Xiao, F. Lin, Y. S. Xiao, and J. Li. Germ cells and fertilization differences among Japanese flounder Paralichthys olivaceus, summer flounder Paralichthys dentatus and their first and second generations. Journal of Fish Biology, 80(3):473– 485, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tong:2012:SDA

[706] X. H. Tong, Q. H. Liu, S. H. Xu, D. Y. Ma, Z. Z. Xiao, Y. S. Xiao, and J. Li. Skeletal development and abnormalities of the vertebral column and of the fins in hatchery-reared turbot Scophthalmus maximus. Journal of Fish Biology, 80(3):486–502, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edenbrow:2012:KFI

[707] M. Edenbrow and D. P. Croft. Kin and familiarity influence association preferences and aggression in the mangrove killifish Kryptolebias marmoratus. Journal of Fish Biology, 80(3):503–518, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2012:FTP

[708] S. M. Walsh, S. L. Hamilton, B. I. Ruttenberg, M. K. Donovan, and S. A. Sandin. Fishing top predators indirectly affects condition and reproduction in a reef-fish community. Journal of Fish Biology, 80(3):519–537, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2012:MFM

[709] N. S. Johnson, S.-S. Yun, T. J. Buchinger, and W. Li. Multiple functions of a multi-component mating pheromone in sea lamprey Petromyzon marinus. Journal of Fish Biology, 80(3):538–554, March 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2012:EDN

[710] S. A. Anderson, I. Salinas, S. P. Walker, Y. Gublin, S. Pether, Y. Y. Kohn, and J. E. Symonds. Early development of New Zealand hapuku Polyprion oxygeneios eggs and larvae. Journal of Fish Biology, 80(3): 555–571, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guerrero-Estevez:2012:GMS

[711] S. Guerrero-Estévez and N. Moreno-Mendoza. Gonadal morphogenesis and sex differentiation in the viviparous fish Chapalichthys encaustus (Teleostei, Cyprinodontiformes, Goodeidae). Journal of Fish Biology, 80 (3):572–594, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2012:IPP

[712] N. G. Andersen. Influences of potential predictor variables on gastric evacuation in Atlantic cod Gadus morhua feeding on fish prey: parameterization of a generic model. Journal of Fish Biology, 80(3):595–612, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fowler:2012:ESP

[713] A. M. Fowler and D. J. Booth. Evidence of sustained populations of a small reef fish on artificial structures. Does depth affect production on artificial reefs? Journal of Fish Biology, 80(3):613–629, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marreta:2012:CMW

[714] M. E. Marreta, F. L. C. Faldoni, and P. P. Parise-Maltempi. Cytogenetic mapping of the W chromosome in the genus Leporinus (Teleostei, Anostomidae) using a highly repetitive DNA sequence. Journal of Fish Biology, 80(3):630–637, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broadhurst:2012:RIJ

[715] M. K. Broadhurst, P. A. Butcher, K. C. Hall, B. R. Cullis, and S. P. McGrath. Resilience of inshore, juvenile snapper Pagrus auratus to angling and release. Journal of Fish Biology, 80(3):638–650, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nguyen:2012:SPG

[716] T. T. T. Nguyen and P. Sunnucks. Strong population genetic structure and its management implications in the mud carp Cirrhinus molitorella, an indigenous freshwater species subject to an aquaculture and culturebased fishery. Journal of Fish Biology, 80(3):651–668, March 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2012:MCC

[717] Y. J. Fu, F. G. Huang, T. Yuan, J. R. Gu, G. Q. Luo, and H. Xu. Molecular cloning, characterization and expression analysis of B cell translocation gene 1 in grass carp Ctenopharyngodon idella. Journal of Fish Biology, 80(3):669–678, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2012:IEA

[718] X. W. Chen, S. Jiang, and Z. Y. Shi. Identification and expression analysis of fabp2 gene from common carp Cyprinus carpio. Journal of Fish Biology, 80(3):679–691, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2012:NOT

[719] T. Arai, N. Chino, S. Z. Zulkifli, and A. Ismail. Notes on the occurrence of the tropical eel Anguilla bicolor bicolor in Peninsular Malaysia, Malaysia. Journal of Fish Biology, 80(3):692–697, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parkinson:2012:VOD

[720] K. L. Parkinson, D. J. Booth, and J. E. Lee. Validation of otolith daily increment formation for two temperate syngnathid fishes: the pipefishes Stigmatopora argus and Stigmatopora nigra. Journal of Fish Biology, 80 (3):698–704, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheppard:2012:EAE

[721] J. N. Sheppard, A. K. Whitfield, P. D. Cowley, and J. M. Hill. Effects of altered estuarine submerged macrophyte bed cover on the omnivorous Cape stumpnose Rhabdosargus holubi. Journal of Fish Biology, 80(3): 705–712, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernhardt:2012:SMP

[722] B. Bernhardt, K. P. Lampert, F. Leese, C. Mayer, and R. Tollrian. Are shoals of minnow Phoxinus phoxinus formed by close kin? Journal of Fish Biology, 80(3):713–721, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rhodes:2012:RSI

[723] S. B. Rhodes and I. Schlupp. Rapid and socially induced change of a badge of status. Journal of Fish Biology, 80(3):722–727, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larson:2012:WCP

[724] A. D. Larson, P. L. Scarborough, C. L. Sayre, and J. B. Johnson. What can phylogeographic breaks in the livebearer Poeciliopsis turrubarensis tell about fish community boundaries in western Costa Rica? Journal of Fish Biology, 80(3):728–736, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:E

[725] Anonymous. Erratum. Journal of Fish Biology, 80(3):737, March 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huebert:2012:CRS

[726] K. B. Huebert. Connecting recent studies on fish vertical navigation. Journal of Fish Biology, 80(4):739–740, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barkataki:2012:COF

[727] S. Barkataki, N. Aluru, M. Li, M. M. Vijayan, and J. F. Leatherland. Characteristics of ovarian follicle steroidogenesis during vitellogenesis in an asynchronously ovulating stock of rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 80(4):741–751, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertucci:2012:SPC

[728] F. Bertucci, J. Attia, M. Beauchaud, and N. Mathevon. Sounds produced by the cichlid fish Metriaclima zebra allow reliable estimation of size and provide information on individual identity. Journal of Fish Biology, 80 (4):752–766, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knip:2012:HUS

[729] D. M. Knip, M. R. Heupel, and C. A. Simpfendorfer. Habitat use and spatial segregation of adult spottail sharks Carcharhinus sorrah in tropical nearshore waters. Journal of Fish Biology, 80(4):767–784, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tine:2012:VGE

[730] M. Tine, B. Guinand, and J.-D. Durand. Variation in gene expression along a salinity gradient in wild populations of the euryhaline blackchinned tilapia Sarotherodon melanotheron. Journal of Fish Biology, 80(4):785–801, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milhomem:2012:GCN

[731] S. S. R. Milhomem, W. G. R. Crampton, J. C. Pieczarka, G. H. Shetka, D. S. Silva, and C. Y. Nagamachi. Gymnotus capanema, a new species of electric knife fish (Gymnotiformes, Gymnotidae) from eastern Amazonia, with comments on an unusual karyotype. Journal of Fish Biology, 80(4): 802–815, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herrmann:2012:MHH

[732] K. K. Herrmann and R. Poulin. The missing host hypothesis: do chemical cues from predators induce life cycle truncation of trematodes within their fish host? Journal of Fish Biology, 80(4):816–830, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:2012:NST

[733] L. P. Zheng, J. X. Yang, and X. Y. Chen. A new species of Triplophysa (Nemacheilidae: Cypriniformes), from Guangxi, southern China. Journal of Fish Biology, 80(4):831–841, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luo:2012:EBM

[734] Y. P. Luo and Q. Q. Wang. Effects of body mass and temperature on routine metabolic rate of juvenile largemouth bronze gudgeon Coreius guichenoti. Journal of Fish Biology, 80(4):842–851, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dalben:2012:CRF

[735] A. Dalben and S. R. Floeter. Cryptobenthic reef fishes: depth distribution and correlations with habitat complexity and sea urchins. Journal of Fish Biology, 80(4):852–865, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2012:CYG

[736] S. F. Wang, Y. Cai, Y. X. Qin, Y. C. Zhou, Y. Q. Su, and J. Wang. Characterization of yellow grouper Epinephelus awoara (Serranidae) karyotype by chromosome bandings and fluorescence in situ hybridization. Journal of Fish Biology, 80(4):866–875, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ballesta:2012:EAT

[737] S. Ballesta, L. M. Hanson, and A. P. Farrell. The effect of adrenaline on the temperature dependency of cardiac action potentials in pink salmon Oncorhynchus gorbuscha. Journal of Fish Biology, 80(4):876–885, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohtla:2012:ELM

[738] M. Rohtla, M. Vetemaa, K. Urtson, and A. Soesoo. Early life migration patterns of Baltic Sea pike Esox lucius. Journal of Fish Biology, 80(4): 886–893, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2012:BRE

[739] T. G. Pottinger. Book review: Encyclopedia of Fish Physiology — Edited by A. P. Farrell, J. J. Cech, J. G. Richards and E. D. Stevens. Journal of Fish Biology, 80(4):894–895, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:Aa

[740] Anonymous. Announcement. Journal of Fish Biology, 80(4):896, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2012:CSE

[741] W. T. White, S. J. M. Blaber, and J. F. Craig. The current status of elasmobranchs: biology, fisheries and conservation. Journal of Fish Biology, 80(5):897–900, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2012:RTC

[742] W. T. White and P. R. Last. A review of the taxonomy of chondrichthyan fishes: a modern perspective. Journal of Fish Biology, 80(5):901–917, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maisey:2012:WEI

[743] J. G. Maisey. What is an ‘elasmobranch’? The impact of palaeontology in understanding elasmobranch phylogeny and evolution. Journal of Fish Biology, 80(5):918–951, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spaet:2012:RER

[744] J. L. Y. Spaet, S. R. Thorrold, and M. L. Berumen. A review of elasmobranch research in the Red Sea. Journal of Fish Biology, 80(5):952–965, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnett:2012:ORH

[745] A. Barnett, J. M. Braccini, C. A. Awruch, and D. A. Ebert. An overview on the role of Hexanchiformes in marine ecosystems: biology, ecology and conservation status of a primitive order of modern sharks. Journal of Fish Biology, 80(5):966–990, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacNeil:2012:BGS

[746] M. A. MacNeil, B. C. McMeans, N. E. Hussey, P. Vecsei, J. Svavarsson, K. M. Kovacs, C. Lydersen, M. A. Treble, G. B. Skomal, M. Ramsey, and A. T. Fisk. Biology of the Greenland shark Somniosus microcephalus. Journal of Fish Biology, 80(5):991–1018, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowat:2012:RBF

[747] D. Rowat and K. S. Brooks. A review of the biology, fisheries and conservation of the whale shark Rhincodon typus. Journal of Fish Biology, 80(5):1019–1056, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2012:OBS

[748] J. R. Ellis, S. R. McCully, and M. J. Brown. An overview of the biology and status of undulate ray Raja undulata in the north-east Atlantic Ocean. Journal of Fish Biology, 80(5):1057–1074, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Couturier:2012:BEC

[749] L. I. E. Couturier, A. D. Marshall, F. R. A. Jaine, T. Kashiwagi, S. J. Pierce, K. A. Townsend, S. J. Weeks, M. B. Bennett, and A. J. Richardson. Biology, ecology and conservation of the mobulidae. Journal of Fish Biology, 80(5):1075–1119, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Portnoy:2012:MMP

[750] D. S. Portnoy and E. J. Heist. Molecular markers: progress and prospects for understanding reproductive ecology in elasmobranchs. Journal of Fish Biology, 80(5):1120–1140, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fitzpatrick:2012:APP

[751] J. L. Fitzpatrick, R. M. Kempster, T. S. Daly-Engel, S. P. Collin, and J. P. Evans. Assessing the potential for post-copulatory sexual selection in elasmobranchs. Journal of Fish Biology, 80(5):1141–1158, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tribuzio:2012:LHC

[752] C. A. Tribuzio and G. H. Kruse. Life history characteristics of a lightly exploited stock of Squalus suckleyi. Journal of Fish Biology, 80(5):1159– 1180, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irvine:2012:BVT

[753] S. B. Irvine, R. K. Daley, K. J. Graham, and J. D. Stevens. Biological vulnerability of two exploited sharks of the genus Deania (Centrophoridae). Journal of Fish Biology, 80(5):1181–1206, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dai:2012:BOC

[754] X. J. Dai, J. F. Zhu, X. J. Chen, L. X. Xu, and Y. Chen. Biological observations on the crocodile shark Pseudocarcharias kamoharai. Journal of Fish Biology, 80(5):1207–1212, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bustamante:2012:RBZ

[755] C. Bustamante, C. Vargas-Caro, M. C. Oddone, F. Concha, H. Flores, J. Lamilla, and M. B. Bennett. Reproductive biology of Zearaja chilensis (Chondrichthyes: Rajidae) in the south-east Pacific Ocean. Journal of Fish Biology, 80(5):1213–1226, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jirik:2012:EMW

[756] K. E. Jirik and C. G. Lowe. An elasmobranch maternity ward: female round stingrays Urobatis halleri use warm, restored estuarine habitat during gestation. Journal of Fish Biology, 80(5):1227–1245, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mejia-Falla:2012:RVU

[757] P. A. Mejía-Falla, A. F. Navia, and E. Cortés. Reproductive variables of Urotrygon rogersi (Batoidea: Urotrygonidae): a species with a triannual reproductive cycle in the eastern tropical Pacific Ocean. Journal of Fish Biology, 80(5):1246–1266, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trinnie:2012:BRC

[758] F. I. Trinnie, T. I. Walker, P. L. Jones, and L. J. Laurenson. Biennial reproductive cycle in an extensive matrotrophic viviparous batoid: the sandyback stingaree Urolophus bucculentus from south-eastern Australia. Journal of Fish Biology, 80(5):1267–1291, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geraghty:2012:MCT

[759] P. T. Geraghty, A. S. Jones, J. Stewart, and W. G. Macbeth. Microcomputed tomography: an alternative method for shark ageing. Journal of Fish Biology, 80(5):1292–1299, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bubley:2012:RSD

[760] W. J. Bubley, J. Kneebone, J. A. Sulikowski, and P. C. W. Tsang. Reassessment of spiny dogfish Squalus acanthias age and growth using vertebrae and dorsal-fin spines. Journal of Fish Biology, 80(5):1300–1319, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:2012:BSS

[761] N. G. Hall, C. Bartron, W. T. White, Dharmadi, and I. C. Potter. Biology of the silky shark Carcharhinus falciformis (Carcharhinidae) in the eastern Indian Ocean, including an approach to estimating age when timing of parturition is not well defined. Journal of Fish Biology, 80 (5):1320–1341, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papastamatiou:2012:AHD

[762] Y. P. Papastamatiou and C. G. Lowe. An analytical and hypothesisdriven approach to elasmobranch movement studies. Journal of Fish Biology, 80(5):1342–1360, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2012:UAP

[763] A. D. Marshall and S. J. Pierce. The use and abuse of photographic identification in sharks and rays. Journal of Fish Biology, 80(5):1361– 1379, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2012:SPB

[764] G. D. Williams, K. S. Andrews, S. L. Katz, M. L. Moser, N. Tolimieri, D. A. Farrer, and P. S. Levin. Scale and pattern of broadnose sevengill shark Notorynchus cepedianus movement in estuarine embayments. Journal of Fish Biology, 80(5):1380–1400, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramirez-Macias:2012:PCA

[765] D. Ramírez-Macías, M. Meekan, R. De La Parra-Venegas, F. Remolina-Suárez, M. Trigo-Mendoza, and R. Vázquez-Juárez. Patterns in composition, abundance and scarring of whale sharks Rhincodon typus near Holbox Island, Mexico. Journal of Fish Biology, 80(5):1401–1416, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Awruch:2012:MPD

[766] C. A. Awruch, S. D. Frusher, J. D. Stevens, and A. Barnett. Movement patterns of the draughtboard shark Cephaloscyllium laticeps (Scyliorhinidae) determined by passive tracking and conventional tagging. Journal of Fish Biology, 80(5):1417–1435, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DiGirolamo:2012:DTP

[767] A. L. DiGirolamo, S. H. Gruber, C. Pomory, and W. A. Bennett. Diel temperature patterns of juvenile lemon sharks Negaprion brevirostris, in a shallow-water nursery. Journal of Fish Biology, 80(5):1436–1448, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hussey:2012:SIE

[768] N. E. Hussey, M. A. MacNeil, J. A. Olin, B. C. McMeans, M. J. Kinney, D. D. Chapman, and A. T. Fisk. Stable isotopes and elasmobranchs: tissue types, methods, applications and assumptions. Journal of Fish Biology, 80(5):1449–1484, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boyle:2012:SIA

[769] M. D. Boyle, D. A. Ebert, and G. M. Cailliet. Stable-isotope analysis of a deep-sea benthic-fish assemblage: evidence of an enriched benthic food web. Journal of Fish Biology, 80(5):1485–1507, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Courtney:2012:PSS

[770] D. L. Courtney and R. Foy. Pacific sleeper shark Somniosus pacificus trophic ecology in the eastern North Pacific Ocean inferred from nitrogen and carbon stable-isotope ratios and diet. Journal of Fish Biology, 80 (5):1508–1545, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forman:2012:DSH

[771] J. S. Forman and M. R. Dunn. Diet and scavenging habits of the smooth skate Dipturus innominatus. Journal of Fish Biology, 80(5):1546–1562, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Garcia:2012:FHT

[772] J. López-García, A. F. Navia, P. A. Mejía-Falla, and E. A. Rubio. Feeding habits and trophic ecology of Dasyatis longa (Elasmobranchii: Myliobatiformes): sexual, temporal and ontogenetic effects. Journal of Fish Biology, 80(5):1563–1579, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2012:FED

[773] I. P. Jacobsen and M. B. Bennett. Feeding ecology and dietary comparisons among three sympatric Neotrygon (Myliobatoidei: Dasyatidae) species. Journal of Fish Biology, 80(5):1580–1594, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsu:2012:FMC

[774] H. H. Hsu, S. J. Joung, and K. M. Liu. Fisheries, management and conservation of the whale shark Rhincodon typus in Taiwan. Journal of Fish Biology, 80(5):1595–1607, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2012:PSW

[775] W. Li, Y. Wang, and B. Norman. A preliminary survey of whale shark Rhincodon typus catch and trade in China: an emerging crisis. Journal of Fish Biology, 80(5):1608–1618, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2012:SSS

[776] A. B. M. Moore, I. D. McCarthy, G. R. Carvalho, and R. Peirce. Species, sex, size and male maturity composition of previously unreported elasmobranch landings in Kuwait, Qatar and Abu Dhabi Emirate. Journal of Fish Biology, 80(5):1619–1642, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biery:2012:GRS

[777] L. Biery and D. Pauly. A global review of species-specific shark-fin-to-body-mass ratios and relevant legislation. Journal of Fish Biology, 80 (5):1643–1677, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2012:SCS

[778] J. F. Silva, J. R. Ellis, and T. L. Catchpole. Species composition of skates (Rajidae) in commercial fisheries around the British Isles and their discarding patterns. Journal of Fish Biology, 80(5):1678–1703, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2012:SAE

[779] A. Arkhipkin, P. Brickle, V. Laptikhovsky, J. Pompert, and A. Winter. Skate assemblage on the eastern Patagonian Shelf and Slope: structure, diversity and abundance. Journal of Fish Biology, 80(5):1704–1726, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gallagher:2012:ERA

[780] A. J. Gallagher, P. M. Kyne, and N. Hammerschlag. Ecological risk assessment and its application to elasmobranch conservation and management. Journal of Fish Biology, 80(5):1727–1748, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2012:RAS

[781] J. K. Carlson, L. F. Hale, A. Morgan, and G. Burgess. Relative abundance and size of coastal sharks derived from commercial shark longline catch and effort data. Journal of Fish Biology, 80(5):1749–1764, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnett:2012:ESC

[782] L. A. K. Barnett, D. A. Ebert, and G. M. Cailliet. Evidence of stability in a chondrichthyan population: case study of the spotted ratfish Hydrolagus colliei (Chondrichthyes: Chimaeridae). Journal of Fish Biology, 80(5):1765–1788, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dudgeon:2012:RAM

[783] C. L. Dudgeon, D. C. Blower, D. Broderick, J. L. Giles, B. J. Holmes, T. Kashiwagi, N. C. Krück, J. A. T. Morgan, B. J. Tillett, and J. R. Ovenden. A review of the application of molecular genetics for fisheries management and conservation of sharks and rays. Journal of Fish Biology, 80(5):1789–1843, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward-Paige:2012:RPC

[784] C. A. Ward-Paige, D. M. Keith, B. Worm, and H. K. Lotze. Recovery potential and conservation options for elasmobranchs. Journal of Fish Biology, 80(5):1844–1869, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromhead:2012:IFI

[785] D. Bromhead, S. Clarke, S. Hoyle, B. Muller, P. Sharples, and S. Harley. Identification of factors influencing shark catch and mortality in the Marshall Islands tuna longline fishery and management implications. Journal of Fish Biology, 80(5):1870–1894, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santana-Garcon:2012:BSP

[786] J. Santana-Garcon, S. Fordham, and S. Fowler. Blue shark Prionace glauca fin-to-carcass-mass ratios in Spain and implications for finning ban enforcement. Journal of Fish Biology, 80(5):1895–1903, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shadwick:2012:MFS

[787] R. E. Shadwick and J. A. Goldbogen. Muscle function and swimming in sharks. Journal of Fish Biology, 80(5):1904–1939, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:2012:BSS

[788] W. Meyer and U. Seegers. Basics of skin structure and function in elasmobranchs: a review. Journal of Fish Biology, 80(5):1940–1967, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yopak:2012:NCF

[789] K. E. Yopak. Neuroecology of cartilaginous fishes: the functional implications of brain scaling. Journal of Fish Biology, 80(5):1968–2023, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lisney:2012:VET

[790] T. J. Lisney, S. M. Theiss, S. P. Collin, and N. S. Hart. Vision in elasmobranchs and their relatives: 21st century advances. Journal of Fish Biology, 80(5):2024–2054, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kempster:2012:PEF

[791] R. M. Kempster, I. D. McCarthy, and S. P. Collin. Phylogenetic and ecological factors influencing the number and distribution of electroreceptors in elasmobranchs. Journal of Fish Biology, 80(5):2055–2088, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnes:2012:CMB

[792] D. W. Barnes. Cell and molecular biology of the spiny dogfish Squalus acanthias and little skate Leucoraja erinacea: insights from in vitro cultured cells. Journal of Fish Biology, 80(5):2089–2111, April 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giudice:2012:ENV

[793] G. Del Giudice, M. Prisco, M. Agnese, M. Verderame, L. Rosati, E. Limatola, and P. Andreuccetti. Effects of nonylphenol on vitellogenin synthesis in adult males of the spotted ray Torpedo marmorata. Journal of Fish Biology, 80(5):2112–2121, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:AF

[794] Anonymous. Airbreathing fishes. Journal of Fish Biology, 80(5):2123, April 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cioffi:2012:CRH

[795] M. B. Cioffi, O. Moreira-Filho, L. F. Almeida-Toledo, and L. A. C. Bertollo. The contrasting role of heterochromatin in the differentiation of sex chromosomes: an overview from Neotropical fishes. Journal of Fish Biology, 80(6):2125–2139, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tillett:2012:ERP

[796] B. J. Tillett, M. G. Meekan, I. C. Field, D. C. Thorburn, and J. R. Ovenden. Evidence for reproductive philopatry in the bull shark Carcharhinus leucas. Journal of Fish Biology, 80(6):2140–2158, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arechavala-Lopez:2012:DFG

[797] P. Arechavala-Lopez, P. Sanchez-Jerez, J. T. Bayle-Sempere, D. G. Sfakianakis, and S. Somarakis. Discriminating farmed gilthead sea bream Sparus aurata and European sea bass Dicentrarchus labrax from wild stocks through scales and otoliths. Journal of Fish Biology, 80(6):2159– 2175, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russell:2012:RST

[798] D. J. Russell, P. A. Thuesen, and F. E. Thomson. Reproductive strategies of two invasive tilapia species Oreochromis mossambicus and Tilapia mariae in northern Australia. Journal of Fish Biology, 80(6):2176–2197, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2012:PGS

[799] M. R. Menezes, G. Kumar, and S. P. Kunal. Population genetic structure of skipjack tuna Katsuwonus pelamis from the Indian coast using sequence analysis of the mitochondrial DNA D-loop region. Journal of Fish Biology, 80(6):2198–2212, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glowacki:2012:LDR

[800] Ł. Głowacki and T. Penczak. Large dam reservoirs are probably longperiod oscillators of fish diversity. Journal of Fish Biology, 80(6):2213– 2235, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uchiyama:2012:SIK

[801] M. Uchiyama, M. Komiyama, H. Yoshizawa, N. Shimizu, N. Konno, and K. Matsuda. Structures and immunolocalization of Na^{+}, K^{+}-ATPase, Na^{+} /H^{+} exchanger 3 and vacuolar-type H^{+}-ATPase in the gills of blennies (Teleostei: Blenniidae) inhabiting rocky intertidal areas. Journal of Fish Biology, 80(6):2236–2252, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domingos:2012:RFC

[802] J. A. Domingos, P. Fromm, C. Smith-Keune, and D. R. Jerry. A robust flow-cytometric protocol for assessing growth rate of hatchery-reared barramundi Lates calcarifer larvae. Journal of Fish Biology, 80(6):2253– 2266, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sykes:2012:EEO

[803] P. J. Sykes, H. Stryhn, C. A. McClure, C. L. Brooking, and K. L. Hammell. Evaluation of external operculum loop tags to individually identify cage-cultured Atlantic halibut Hippoglossus hippoglossus in commercial research trials. Journal of Fish Biology, 80(6):2267–2280, May 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2012:IFF

[804] J. Williams, J. S. Hindell, S. E. Swearer, and G. P. Jenkins. Influence of freshwater flows on the distribution of eggs and larvae of black bream Acanthopagrus butcheri within a drought-affected estuary. Journal of Fish Biology, 80(6):2281–2301, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergek:2012:PZV

[805] S. Bergek, F. Franzén, M. Quack, A. Hochkirch, T. Kinitz, T. Prestegaard, and M. Appelberg. Panmixia in Zoarces viviparus: implications for environmental monitoring studies. Journal of Fish Biology, 80(6): 2302–2316, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rasmussen:2012:ERB

[806] J. B. Rasmussen, A. N. Krimmer, A. J. Paul, and A. Hontela. Empirical relationships between body tissue composition and bioelectrical impedance of brook trout Salvelinus fontinalis from a Rocky Mountain stream. Journal of Fish Biology, 80(6):2317–2327, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2012:AGA

[807] K.-S. Chen, T. Shimose, T. Tanabe, C.-Y. Chen, and C.-C. Hsu. Age and growth of albacore Thunnus alalunga in the North Pacific Ocean. Journal of Fish Biology, 80(6):2328–2344, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makiguchi:2012:DFB

[808] Y. Makiguchi, Y. Sugie, T. Kojima, and Y. Naito. Detection of feeding behaviour in common carp Cyprinus carpio by using an acceleration data logger to identify mandibular movement. Journal of Fish Biology, 80(6): 2345–2356, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Contente:2012:FEA

[809] R. F. Contente, M. F. Stefanoni, and H. L. Spach. Feeding ecology of the American freshwater goby Ctenogobius shufeldti (Gobiidae, Perciformes) in a sub-tropical estuary. Journal of Fish Biology, 80(6):2357–2373, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sopinka:2012:DPA

[810] N. M. Sopinka, J. L. Fitzpatrick, J. E. Taves, M. G. Ikonomou, S. E. Marsh-Rollo, and S. Balshine. Does proximity to aquatic pollution affect reproductive traits in a wild-caught intertidal fish? Journal of Fish Biology, 80(6):2374–2383, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nilsson:2012:VCD

[811] P. A. Nilsson, H. Baktoft, M. Boel, K. Meier, L. Jacobsen, E. M. Rokkjær, T. Clausen, and C. Skov. Visibility conditions and diel period affect small-scale spatio-temporal behaviour of pike Esox lucius in the absence of prey and conspecifics. Journal of Fish Biology, 80(6):2384–2389, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2012:BRB

[812] L. A. Jawad. Book review: The Banggai Cardinalfish: Natural History, Conservation and Culture of Pterapogon kauderni — Edited by A. A. Vagelli. Journal of Fish Biology, 80(6):2390–2392, May 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2012:ITM

[813] K. B. Li, O. Q. Chang, F. Wang, C. Liu, Q. Wang, F. L. Liang, B. Y. Ma, and S. Q. Wu. Identification of a transparent mutant tiger barb Puntius tetrazona and its use for in vivo observation of a Pleistophora sp. (Microsporidia) infection. Journal of Fish Biology, 80(7):2393–2404, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colonello:2012:RBS

[814] J. C. Colonello, M. L. García, C. A. Lasta, and R. C. Menni. Reproductive biology of the spotback skate Atlantoraja castelnaui in the southwest Atlantic Ocean. Journal of Fish Biology, 80(7):2405–2419, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Underwood:2012:EAT

[815] Z. E. Underwood, C. A. Myrick, and K. B. Rogers. Effect of acclimation temperature on the upper thermal tolerance of Colorado River cutthroat trout Oncorhynchus clarkii pleuriticus: thermal limits of a North American salmonid. Journal of Fish Biology, 80(7):2420–2433, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Panarari-Antunes:2012:GDA

[816] R. S. Panarari-Antunes, A. J. Prioli, S. M. A. P. Prioli, V. N. Gomes, H. F. Júlio, Jr., C. S. Agostinho, J. P. Silva Filho, T. A. Boni, and L. M. Prioli. Genetic divergence among invasive and native populations of Plagioscion squamosissimus (Perciformes, Sciaenidae) in Neotropical regions. Journal of Fish Biology, 80(7):2434–2447, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2012:DON

[817] H. Jensen, M. Kiljunen, and P.-A. Amundsen. Dietary ontogeny and niche shift to piscivory in lacustrine brown trout Salmo trutta revealed by stomach content and stable isotope analyses. Journal of Fish Biology, 80(7):2448–2462, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2012:MPF

[818] S. Y. V. Liu and C. F. Dai. Molecular phylogeny of four morphologically similar Ostorhinchus taxa with specific comments on O. cf. properuptus dotted type. Journal of Fish Biology, 80(7):2463–2474, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Czesny:2012:DSV

[819] S. J. Czesny, J. Rinchard, B.-J. Lee, K. Dabrowski, J. M. Dettmers, and Y. Cao. Does spatial variation in egg thiamine and fatty-acid concentration of Lake Michigan lake trout Salvelinus namaycush lead to differential early mortality syndrome and yolk oedema mortality in offspring? Journal of Fish Biology, 80(7):2475–2493, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dickson:2012:MSS

[820] K. A. Dickson, J. M. Donley, M. W. Hansen, and J. A. Peters. Maximum sustainable speed, energetics and swimming kinematics of a tropical carangid fish, the green jack Caranx caballus. Journal of Fish Biology, 80(7):2494–2516, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Queirolo:2012:BFC

[821] D. Queirolo, E. Gaete, I. Montenegro, M. C. Soriguer, and K. Erzini. Behaviour of fish by-catch in the mouth of a crustacean trawl. Journal of Fish Biology, 80(7):2517–2527, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalogirou:2012:FEI

[822] S. Kalogirou, F. Mittermayer, L. Pihl, and H. Wennhage. Feeding ecology of indigenous and non-indigenous fish species within the family Sphyraenidae. Journal of Fish Biology, 80(7):2528–2548, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karinthanyakit:2012:MPR

[823] W. Karinthanyakit and A. Jondeung. Molecular phylogenetic relationships of pangasiid and schilbid catfishes in Thailand. Journal of Fish Biology, 80(7):2549–2570, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schuckel:2012:DOA

[824] S. Schückel, A. F. Sell, I. Kröncke, and H. Reiss. Diet overlap among flatfish species in the southern North Sea. Journal of Fish Biology, 80 (7):2571–2594, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wouters:2012:MIH

[825] J. Wouters, S. Janson, V. Lusková, and K. H. Olsén. Molecular identification of hybrids of the invasive gibel carp Carassius auratus gibelio and crucian carp Carassius carassius in Swedish waters. Journal of Fish Biology, 80(7):2595–2604, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2012:IEP

[826] C. Yang, J. Su, R. Zhang, L. Peng, and Q. Li. Identification and expression profiles of grass carp Ctenopharyngodon idella tlr 7 in responses to double-stranded RNA and virus infection. Journal of Fish Biology, 80 (7):2605–2622, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ponton:2012:DLE

[827] D. Ponton, N. Loiseau, and P. Chabanet. Does light explain damselfish Chromis viridis abundances observed over coral colonies? Journal of Fish Biology, 80(7):2623–2628, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2012:RBP

[828] C. D. Jones and T. J. Near. The reproductive behaviour of Pogonophryne scotti confirms widespread egg-guarding parental care among Antarctic notothenioids. Journal of Fish Biology, 80(7):2629–2635, June 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cannas:2012:EHV

[829] M. Cannas, P. Domenici, and C. Lefrançois. The effect of hypoxia on ventilation frequency in startled common sole Solea solea. Journal of Fish Biology, 80(7):2636–2642, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendonca:2012:ICP

[830] F. F. Mendonça, L. H. F. Ussami, D. T. Hashimoto, L. H. G. Pereira, F. Porto-Foresti, C. Oliveira, O. B. F. Gadig, and F. Foresti. Identification and characterization of polymorphic microsatellite loci in the blue shark Prionace glauca, and cross-amplification in other shark species. Journal of Fish Biology, 80(7):2643–2646, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2012:BRA

[831] Stephen J. Lockwood. Book review: All the Fish in the Sea: Maximum Sustainable Yield and the Failure of Fisheries Management — Edited by C. Finley. Journal of Fish Biology, 80(7):2647–2648, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:Cb

[832] Anonymous. Corrigendum. Journal of Fish Biology, 80(7):2649, June 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klinck:2012:VCC

[833] J. S. Klinck, A. Singh, and C. M. Wood. In vitro characterization of calcium transport along the gastrointestinal tract of freshwater rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 81(1):1–20, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wacker:2012:LMF

[834] S. Wacker, K. de Jong, E. Forsgren, and T. Amundsen. Large males fight and court more across a range of social environments: an experiment on the two spotted goby Gobiusculus flavescens. Journal of Fish Biology, 81(1):21–34, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forsythe:2012:ELC

[835] P. S. Forsythe, K. T. Scribner, J. A. Crossman, A. Ragavendran, E. A. Baker, C. Davis, and K. K. Smith. Environmental and lunar cues are predictive of the timing of river entry and spawning-site arrival in lake sturgeon Acipenser fulvescens. Journal of Fish Biology, 81(1):35–53, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Block:2012:DMH

[836] A. J. Block and P. M. Mabee. Development of the mandibular, hyoid arch and gill arch skeleton in the Chinese barb Puntius semifasciolatus: comparisons of ossification sequences among Cypriniformes. Journal of Fish Biology, 81(1):54–80, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newton:2012:GER

[837] J. R. Newton, C. De Santis, and D. R. Jerry. The gene expression response of the catadromous perciform barramundi Lates calcarifer to an acute heat stress. Journal of Fish Biology, 81(1):81–93, July 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawamura:2012:ESH

[838] K. Kawamura, M. Furukawa, M. Kubota, and Y. Harada. Effects of stocking hatchery fish on the phenotype of indigenous populations in the amago salmon Oncorhynchus masou ishikawae in Japan. Journal of Fish Biology, 81(1):94–109, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Terencio:2012:MSD

[839] M. L. Terencio, C. H. Schneider, and J. I. R. Porto. Molecular signature of the D-loop in the brown pencilfish Nannostomus eques (Characiformes, Lebiasinidae) reveals at least two evolutionary units in the Rio Negro basin, Brazil. Journal of Fish Biology, 81(1):110–124, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolbadinezhad:2012:EGS

[840] S. Malakpour Kolbadinezhad, A. Hajimoradloo, R. Ghorbani, H. Joshaghani, and J. M. Wilson. Effects of gradual salinity increase on osmoregulation in Caspian roach Rutilus caspicus. Journal of Fish Biology, 81(1):125– 134, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galster:2012:WSV

[841] B. J. Galster, M. R. Wuellner, and B. D. S. Graeb. Walleye Sander vitreus and smallmouth bass Micropterus dolomieu interactions: an historic stable-isotope analysis approach. Journal of Fish Biology, 81(1):135–147, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rees:2012:TDG

[842] B. B. Rees, T. E. Targett, B. J. Ciotti, C. A. Tolman, S. S. Akkina, and A. M. Gallaty. Temporal dynamics in growth and white skeletal muscle composition of the mummichog Fundulus heteroclitus during chronic hypoxia and hyperoxia. Journal of Fish Biology, 81(1):148–164, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lu:2012:IEA

[843] D. Q. Lu, S. B. Yi, M. Yao, Y. W. Li, X. C. Liu, Y. Zhang, and H. R. Lin. Identification and expression analysis of major histocompatibility complex IIB gene in orange-spotted grouper Epinephelus coioides. Journal of Fish Biology, 81(1):165–180, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pribyl:2012:IBI

[844] A. L. Pribyl, C. B. Schreck, S. J. Parker, and V. M. Weis. Identification of biomarkers indicative of barotrauma and recovery in black rockfish Sebastes melanops. Journal of Fish Biology, 81(1):181–196, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2012:ETS

[845] F. Lahnsteiner and N. Mansour. The effect of temperature on sperm motility and enzymatic activity in brown trout Salmo trutta, burbot Lota lota and grayling Thymallus thymallus. Journal of Fish Biology, 81(1):197–209, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oates:2012:RBR

[846] J. Oates, A. Manica, R. Bshary, and A. S. Grutter. Relationship between roving behaviour and the diet and client composition of the cleaner fish Labroides bicolor. Journal of Fish Biology, 81(1):210–219, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pappal:2012:RBB

[847] A. L. Pappal, R. A. Rountree, and D. G. MacDonald. Relationship between body size and habitat complexity preference in age-0 and -1 year winter flounder Pseudopleuronectes americanus. Journal of Fish Biology, 81(1):220–229, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2012:PIM

[848] B. R. Moore, D. J. Welch, S. J. Newman, and R. J. G. Lester. Parasites as indicators of movement and population connectivity of a non-diadromous, tropical estuarine teleost: king threadfin Polydactylus macrochir. Journal of Fish Biology, 81(1):230–252, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halvorsen:2012:MDD

[849] K. A. T. Halvorsen, E. Árnason, P. J. Smith, and J. Mork. Mitochondrial DNA differentiation between the antitropical blue whiting species Micromesistius poutassou and Micromesistius australis. Journal of Fish Biology, 81(1):253–269, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valdehita:2012:DIC

[850] A. Valdehita, M. L. Fernández-Cruz, F. Torrent, J. L. Sericano, and J. M. Navas. Differences in the induction of cyp1A and related genes in cultured rainbow trout Oncorhynchus mykiss. Additional considerations for the use of EROD activity as a biomarker. Journal of Fish Biology, 81(1):270–287, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2012:GFE

[851] L. N. Harris, E. B. Taylor, R. F. Tallman, and J. D. Reist. Gene flow and effective population size in two life-history types of broad whitefish Coregonus nasus from the Canadian Arctic. Journal of Fish Biology, 81(1):288–307, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macieira:2012:GEI

[852] R. M. Macieira, T. Giarrizzo, J. L. Gasparini, and I. Sazima. Geographic expansion of the invasive mud sleeper Butis koilomatodon (Perciformes: Eleotridae) in the western Atlantic Ocean. Journal of Fish Biology, 81 (1):308–313, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirt-Chabbert:2012:ESI

[853] J. A. Hirt-Chabbert and O. A. Young. Effects of surgically implanted PIT tags on growth, survival and tag retention of yellow shortfin eels Anguilla australis under laboratory conditions. Journal of Fish Biology, 81(1):314–319, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunter:2012:ETP

[854] E. Hunter, N. Taylor, C. J. Fox, M. Maillard, and M. I. Taylor. Effectiveness of TaqMan probes for detection of fish eggs and larvae in the stomach contents of a teleost predator. Journal of Fish Biology, 81(1): 320–328, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2012:EHP

[855] A. T. Correia, A. M. Coimbra, and A. Damasceno-Oliveira. Effect of the hydrostatic pressure on otolith growth of early juveniles of Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 81(1):329–334, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2012:CRA

[856] M. D. Powell, M. S. Burke, and D. Dahle. Cardiac remodelling of Atlantic halibut Hippoglossus hippoglossus induced by experimental anaemia with phenylhydrazine. Journal of Fish Biology, 81(1):335–344, July 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watz:2012:TDP

[857] J. Watz, J. J. Piccolo, L. Greenberg, and E. Bergman. Temperaturedependent prey capture efficiency and foraging modes of brown trout Salmo trutta. Journal of Fish Biology, 81(1):345–350, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanco:2012:NTO

[858] D. R. Blanco, L. A. C. Bertollo, R. L. Lui, M. R. Vicari, V. P. Margarido, R. F. Artoni, and O. Moreira-Filho. A new technique for obtaining mitotic chromosome spreads from fishes in the field. Journal of Fish Biology, 81(1):351–357, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mina:2012:BRA

[859] M. V. Mina. Book review: Atlas of Larvae and Juveniles of Freshwater Fishes of Russia — Edited by A. P. Makeyeva, D. S. Pavlov and D. A. Pavlov. Journal of Fish Biology, 81(1):358–359, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:Ab

[860] Anonymous. Announcement. Journal of Fish Biology, 81(1):360, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2012:FMC

[861] J. Metcalfe and J. F. Craig. Fish migration in the 21st Century: opportunities and challenges. Journal of Fish Biology, 81(2):361–364, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Righton:2012:ASM

[862] D. Righton, K. Aarestrup, D. Jellyman, P. Sébert, G. van den Thillart, and K. Tsukamoto. The Anguilla spp. migration problem: 40 million years of evolution and two millennia of speculation. Journal of Fish Biology, 81(2):365–386, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verreault:2012:SMA

[863] G. Verreault, M. Mingelbier, and P. Dumont. Spawning migration of American eel Anguilla rostrata from pristine (1843–1872) to contemporary (1963–1990) periods in the St Lawrence Estuary, Canada. Journal of Fish Biology, 81(2):387–407, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2012:MPC

[864] S. Hagihara, J. Aoyama, D. Limbong, and K. Tsukamoto. Morphological and physiological changes of female tropical eels, Anguilla celebesensis and Anguilla marmorata, in relation to downstream migration. Journal of Fish Biology, 81(2):408–426, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2012:RVO

[865] Y.-J. Lin, B. M. Jessop, O. L. F. Weyl, Y. Iizuka, S.-H. Lin, W.-N. Tzeng, and C.-L. Sun. Regional variation in otolith Sr:Ca ratios of African longfinned eel Anguilla mossambica and mottled eel Anguilla marmorata: a challenge to the classic tool for reconstructing migratory histories of fishes. Journal of Fish Biology, 81(2):427–441, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2012:DMS

[866] T. Arai and N. Chino. Diverse migration strategy between freshwater and seawater habitats in the freshwater eel genus Anguilla. Journal of Fish Biology, 81(2):442–455, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:2012:PMFa

[867] B. B. Chapman, K. Hulthén, J. Brodersen, P. A. Nilsson, C. Skov, L.-A. Hansson, and C. Brönmark. Partial migration in fishes: causes and consequences. Journal of Fish Biology, 81(2):456–478, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:2012:PMFb

[868] B. B. Chapman, C. Skov, K. Hulthén, J. Brodersen, P. A. Nilsson, L.- A. Hansson, and C. Brönmark. Partial migration in fishes: definitions, methodologies and taxonomic distribution. Journal of Fish Biology, 81 (2):479–499, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorstad:2012:CLS

[869] E. B. Thorstad, F. Whoriskey, I. Uglem, A. Moore, A. H. Rikardsen, and B. Finstad. A critical life stage of the Atlantic salmon Salmo salar: behaviour and survival during the smolt and initial post-smolt migration. Journal of Fish Biology, 81(2):500–542, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ueda:2012:PMI

[870] H. Ueda. Physiological mechanisms of imprinting and homing migration in Pacific salmon Oncorhynchus spp. Journal of Fish Biology, 81(2): 543–558, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnstone:2012:IOR

[871] K. A. Johnstone, K. P. Lubieniecki, B. F. Koop, and W. S. Davidson. Identification of olfactory receptor genes in Atlantic salmon Salmo salar. Journal of Fish Biology, 81(2):559–575, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinch:2012:DFS

[872] S. G. Hinch, S. J. Cooke, A. P. Farrell, K. M. Miller, M. Lapointe, and D. A. Patterson. Dead fish swimming: a review of research on the early migration and high premature mortality in adult Fraser River sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 81(2): 576–599, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2012:QMA

[873] J. E. Johnson, D. A. Patterson, E. G. Martins, S. J. Cooke, and S. G. Hinch. Quantitative methods for analysing cumulative effects on fish migration success: a review. Journal of Fish Biology, 81(2):600–631, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2012:CSO

[874] E. I. Hasegawa. Chum salmon Oncorhynchus keta respond to moonlight during homeward migrations. Journal of Fish Biology, 81(2):632–641, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fjeldstad:2012:CIA

[875] H. P. Fjeldstad, I. Uglem, O. H. Diserud, P. Fiske, T. Forseth, E. Kvingedal, N. A. Hvidsten, F. Økland, and J. Järnegren. A concept for improving Atlantic salmon Salmo salar smolt migration past hydro power intakes. Journal of Fish Biology, 81(2):642–663, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:2012:PIB

[876] A. B. Gill, M. Bartlett, and F. Thomsen. Potential interactions between diadromous fishes of U.K. conservation importance and the electromagnetic fields and subsea noise from marine renewable energy developments. Journal of Fish Biology, 81(2):664–695, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2012:MMA

[877] C. Y. Wang, Q. W. Wei, B. Kynard, H. Du, and H. Zhang. Migrations and movements of adult Chinese sturgeon Acipenser sinensis in the Yangtze River, China. Journal of Fish Biology, 81(2):696–713, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tamdrari:2012:HGC

[878] H. Tamdrari, J.-C. Brêthes, M. Castonguay, and D. E. Duplisea. Homing and group cohesion in Atlantic cod Gadus morhua revealed by tagging experiments. Journal of Fish Biology, 81(2):714–727, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DePontual:2012:DVM

[879] H. De Pontual, A. Jolivet, M. Bertignac, and R. Fablet. Diel vertical migration of European hake Merluccius merluccius and associated temperature histories: insights from a pilot data-storage tagging (DST) experiment. Journal of Fish Biology, 81(2):728–734, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2012:ATR

[880] J. L. A. Jensen and A. H. Rikardsen. Archival tags reveal that Arctic charr Salvelinus alpinus and brown trout Salmo trutta can use estuarine and marine waters during winter. Journal of Fish Biology, 81(2):735– 749, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anteneh:2012:SME

[881] W. Anteneh, A. Getahun, E. Dejen, F. A. Sibbing, L. A. J. Nagelkerke, M. De Graaf, T. Wudneh, J. Vijverberg, and A. P. Palstra. Spawning migrations of the endemic Labeobarbus (Cyprinidae, Teleostei) species of Lake Tana, Ethiopia: status and threats. Journal of Fish Biology, 81(2):750–765, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sturrock:2012:COE

[882] A. M. Sturrock, C. N. Trueman, A. M. Darnaude, and E. Hunter. Can otolith elemental chemistry retrospectively track migrations in fully marine fishes? Journal of Fish Biology, 81(2):766–795, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walther:2012:UOC

[883] B. D. Walther and K. E. Limburg. The use of otolith chemistry to characterize diadromous migrations. Journal of Fish Biology, 81(2):796– 825, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trueman:2012:IMM

[884] C. N. Trueman, K. M. MacKenzie, and M. R. Palmer. Identifying migrations in marine fishes through stable-isotope analysis. Journal of Fish Biology, 81(2):826–847, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Speranza:2012:MBD

[885] E. D. Speranza, N. Cappelletti, M. C. Migoya, L. M. Tatone, and J. C. Colombo. Migratory behaviour of a dominant detritivorous fish Prochilodus lineatus evaluated by multivariate biochemical and pollutant data. Journal of Fish Biology, 81(2):848–865, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makrakis:2012:DMP

[886] M. C. Makrakis, L. E. Miranda, S. Makrakis, H. M. Fontes Júnior, W. G. Morlis, J. H. P. Dias, and J. O. Garcia. Diversity in migratory patterns among Neotropical fishes in a highly regulated river basin. Journal of Fish Biology, 81(2):866–881, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2012:DHS

[887] A. Arkhipkin, P. Brickle, V. Laptikhovsky, and A. Winter. Dining hall at sea: feeding migrations of nektonic predators to the eastern Patagonian shelf. Journal of Fish Biology, 81(2):882–902, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leonard:2012:FPC

[888] G. Leonard, T. Maie, K. N. Moody, G. D. Schrank, R. W. Blob, and H. L. Schoenfuss. Finding paradise: cues directing the migration of the waterfall climbing Hawaiian gobioid Sicyopterus stimpsoni. Journal of Fish Biology, 81(2):903–920, July 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bacon:2012:OFS

[889] P. J. Bacon, J. C. MacLean, I. A. Malcolm, and W. S. C. Gurney. Ova fecundity in Scottish Atlantic salmon Salmo salar: predictions, selective forces and causal mechanisms. Journal of Fish Biology, 81(3):921–938, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yano:2012:FCS

[890] C. F. Yano and V. P. Margarido. First cytogenetic studies of the genus Heptapterus (Actinopterygii, Siluriformes): karyotype differentiation and review of cytogenetic data on the Heptapteridae family. Journal of Fish Biology, 81(3):939–953, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kai:2012:IHM

[891] Y. Kai, K.-D. Park, and T. Nakabo. The incomplete history of mitochondrial lineages between two rockfishes, Sebastes longispinis and Sebastes hubbsi (Scorpaeniformes: Scorpaenidae). Journal of Fish Biology, 81 (3):954–965, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruiz-Gomez:2012:BAE

[892] M. L. Ruiz-Gomez and F. A. Huntingford. Boldness and aggressiveness in early and late hatched three-spined sticklebacks Gasterosteus aculeatus. Journal of Fish Biology, 81(3):966–976, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2012:ETE

[893] F. Lahnsteiner, M. Kletzl, and T. Weismann. The effect of temperature on embryonic and yolk-sac larval development in the burbot Lota lota. Journal of Fish Biology, 81(3):977–986, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kotrschal:2012:IHD

[894] A. Kotrschal, L. F. Sundström, D. Brelin, R. H. Devlin, and N. Kolm. Inside the heads of David and Goliath: environmental effects on brain morphology among wild and growth-enhanced Coho salmon Oncorhynchus kisutch. Journal of Fish Biology, 81(3):987–1002, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2012:ALD

[895] G. M. Menezes, A. Rosa, O. Melo, and F. Porteiro. Annotated list of demersal fishes occurring at sedlo seamount, Azores north-east central Atlantic Ocean. Journal of Fish Biology, 81(3):1003–1018, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffin:2012:DPF

[896] R. Griffin, B. Pearce, and R. D. Handy. Dietary preference and feeding selectivity of common dragonet Callionymus lyra in U.K. Journal of Fish Biology, 81(3):1019–1031, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barros:2012:BAT

[897] B. Barros, Y. Sakai, H. Hashimoto, K. Gushima, and M. Vallinoto. ‘better off alone than in bad company’: agonistic colour display in mimetic juveniles of two ephippid species. Journal of Fish Biology, 81(3):1032– 1042, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Capossela:2012:MCR

[898] K. M. Capossela, R. W. Brill, M. C. Fabrizio, and P. G. Bushnell. Metabolic and cardiorespiratory responses of summer flounder Paralichthys dentatus to hypoxia at two temperatures. Journal of Fish Biology, 81(3):1043–1058, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foldvik:2012:SDC

[899] A. Foldvik, M. A. K. Teichert, S. Einum, A. G. Finstad, O. Ugedal, and T. Forseth. Spatial distribution correspondence of a juvenile Atlantic salmon Salmo salar cohort from age 0+ to 1+ years. Journal of Fish Biology, 81(3):1059–1069, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santic:2012:DFS

[900] M. Šantić, B. Rada, and A. Pallaoro. Diet and feeding strategy of thornback ray Raja clavata. Journal of Fish Biology, 81(3):1070–1084, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurenson:2012:ABG

[901] L. J. B. Laurenson, R. P. French, P. Jones, D. Ierodiaconou, S. Gray, V. L. Versace, A. Rattray, S. Brown, and J. Monk. Aspects of the biology of Galaxias maculatus. Journal of Fish Biology, 81(3):1085–1100, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DArcy:2012:EDB

[902] J. D’Arcy, E. Dunaevskaya, J. W. Treasurer, O. Ottesen, J. Maguire, N. Zhuravleva, A. Karlsen, C. Rebours, and R. D. FitzGerald. Embryonic development in ballan wrasse Labrus bergylta. Journal of Fish Biology, 81 (3):1101–1110, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grans:2012:BFB

[903] A. Gräns, M. Rosengren, L. Niklasson, and M. Axelsson. Behavioural fever boosts the inflammatory response in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 81(3):1111–1117, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stephenson:2012:CCM

[904] J. F. Stephenson. The chemical cues of male sea lice Lepeophtheirus salmonis encourage others to move between host Atlantic salmon Salmo salar. Journal of Fish Biology, 81(3):1118–1123, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verhille:2012:VBA

[905] C. Verhille and A. P. Farrell. The in vitro blood-O_{2} affinity of triploid rainbow trout Oncorhynchus mykiss at different temperatures and CO_{2} tensions. Journal of Fish Biology, 81(3):1124–1132, August 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vieira:2012:FRD

[906] R. P. Vieira, B. Christiansen, S. Christiansen, and J. M. S. Gonçalves. First record of the deep-water whalefish Cetichthys indagator (Actinopterygii: Cetomimidae) in the North Atlantic Ocean. Journal of Fish Biology, 81(3):1133–1137, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akhilesh:2012:ABP

[907] K. V. Akhilesh, K. K. Bineesh, W. T. White, and N. G. K. Pillai. Aspects of the biology of the pygmy ribbontail catshark Eridacnis radcliffei (Proscylliidae: Carcharhiniformes) from the south-west coast of India. Journal of Fish Biology, 81(3):1138–1144, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dinis:2012:BRP

[908] M. T. Dinis. Book review: Practical flatfish culture and stock enhancement — Edited by H. V. Daniels and W. O. Watanabe. Journal of Fish Biology, 81(3):1145–1146, August 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2012:CAL

[909] F. Wu, P. Su, L. Chen, M. Li, X. Liu, and Q. Li. Cloning of Arctic lamprey Lethenteron camtschaticum cd 9 with roles in the immune response. Journal of Fish Biology, 81(4):1147–1157, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conejeros:2012:GMH

[910] P. Conejeros, M. Power, S. Alekseyev, and B. Dixon. Global major histocompatibility Class II β (mh-IIβ)-polymorphism in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 81(4):1158–1174, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Locascio:2012:QTR

[911] J. V. Locascio, S. Burghart, and D. A. Mann. Quantitative and temporal relationships of egg production and sound production by black drum Pogonias cromis. Journal of Fish Biology, 81(4):1175–1191, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zare:2012:FRG

[912] P. Zare, H. K. Larson, and A. Toorang. First record of the gobiid fish Aulopareia ocellata (Teleostei: Gobiidae) from Qeshm Island, The Gulf, with discussion of its generic placement. Journal of Fish Biology, 81(4): 1192–1200, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomelsky:2012:EPS

[913] B. Gomelsky, K. J. Schneider, R. P. Glennon, and D. A. Plouffe. Effect of ploidy on scale-cover pattern in linear ornamental (koi) common carp Cyprinus carpio. Journal of Fish Biology, 81(4):1201–1209, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schneider:2012:CDM

[914] C. H. Schneider, M. C. Gross, M. L. Terencio, and J. I. R. Porto. Cryptic diversity in the mtDNA of the ornamental fish Carnegiella strigata. Journal of Fish Biology, 81(4):1210–1224, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harry:2012:CRE

[915] A. V. Harry, J. A. T. Morgan, J. R. Ovenden, A. J. Tobin, D. J. Welch, and C. A. Simpfendorfer. Comparison of the reproductive ecology of two sympatric blacktip sharks (Carcharhinus limbatus and Carcharhinus tilstoni) off north-eastern Australia with species identification inferred from vertebral counts. Journal of Fish Biology, 81(4):1225–1233, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warren:2012:MTF

[916] M. A. Warren and Y. E. Morbey. Migration timing of female Kokanee salmon Oncorhynchus nerka: diel patterns and effects of maturation state. Journal of Fish Biology, 81(4):1234–1247, September 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guenard:2012:ESM

[917] G. Guénard, D. Boisclair, O. Ugedal, T. Forseth, B. Jonsson, and I. A. Fleming. An experimental study of the multiple effects of brown trout Salmo trutta on the bioenergetics of two Arctic charr Salvelinus alpinus morphs. Journal of Fish Biology, 81(4):1248–1270, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellender:2012:AGM

[918] B. R. Ellender, O. L. F. Weyl, and H. Winker. Age and growth and maturity of southern Africa’s largest cyprinid fish, the largemouth yellowfish Labeobarbus kimberleyensis. Journal of Fish Biology, 81(4):1271–1284, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maia:2012:RBC

[919] C. Maia, K. Erzini, B. Serra-Pereira, and I. Figueiredo. Reproductive biology of cuckoo ray Leucoraja naevus. Journal of Fish Biology, 81(4): 1285–1296, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2012:PDL

[920] D. Harasti, K. Martin-Smith, and W. Gladstone. Population dynamics and life history of a geographically restricted seahorse, Hippocampus whitei. Journal of Fish Biology, 81(4):1297–1314, September 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yokota:2012:RBS

[921] L. Yokota, R. Goitein, M. D. Gianeti, and R. T. P. Lessa. Reproductive biology of the smooth butterfly ray Gymnura micrura. Journal of Fish Biology, 81(4):1315–1326, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gabor:2012:MWB

[922] C. R. Gabor and A. Contreras. Measuring water-borne cortisol in Poecilia latipinna: is the process stressful, can stress be minimized and is cortisol correlated with sex steroid release rates? Journal of Fish Biology, 81(4): 1327–1339, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunlap:2012:SIB

[923] P. V. Dunlap, A. L. Gould, M. L. Wittenrich, and M. Nakamura. Symbiosis initiation in the bacterially luminous sea urchin cardinalfish Siphamia versicolor. Journal of Fish Biology, 81(4):1340–1356, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OFarrell:2012:BSM

[924] B. O’Farrell, C. Dennis, J. A. Benzie, P. McGinnity, J. Carlsson, E. de Eyto, J. P. Coughlan, F. Igoe, R. Meehan, and T. F. Cross. Balancing selection on MHC class I in wild brown trout Salmo trutta. Journal of Fish Biology, 81(4):1357–1374, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flores:2012:RBA

[925] A. Flores, R. Wiff, P. Gálvez, and E. Díaz. Reproductive biology of alfonsino Beryx splendens. Journal of Fish Biology, 81(4):1375–1390, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bogevik:2012:EDL

[926] A. S. Bogevik, S. Natário, Ø. Karlsen, A. Thorsen, K. Hamre, G. Rosenlund, and B. Norberg. The effect of dietary lipid content and stress on egg quality in farmed Atlantic cod Gadus morhua. Journal of Fish Biology, 81(4):1391–1405, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mandelman:2012:BCS

[927] J. W. Mandelman, R. A. Morrison, J. M. Cavin, and M. A. Farrington. The blood chemical status of Atlantic cod Gadus morhua following capture by jig and demersal longline with differential hook removal methods. Journal of Fish Biology, 81(4):1406–1414, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almeida:2012:RKH

[928] J. S. Almeida, P. R. A. M. Affonso, and A. L. Dias. Remarkable karyotypic homogeneity in a widespread tropical fish species: Hoplosternum littorale (Siluriformes, Callichthyidae). Journal of Fish Biology, 81(4): 1415–1421, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Libungan:2012:FNM

[929] L. A. Libungan, G. Ólafsdóttir, S. Skírnisdóttir, S. Pálsson, C. Pampoulie, S. H. Björnsdóttir, K. Ólafsson, G. J. Óskarsson, and A. K. Daníelsdóttir. Fourteen new microsatellite markers for Atlantic herring Clupea harengus. Journal of Fish Biology, 81(4):1422–1426, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodrigues:2012:FRW

[930] N. V. Rodrigues, J. P. S. Correia, J. T. C. Graça, F. Rodrigues, R. Pinho, and M. Hirofumi. First record of a whale shark Rhincodon typus in continental Europe. Journal of Fish Biology, 81(4):1427–1429, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2012:SOD

[931] T. Tomiyama. Seasonal and ontogenetic diet shift in juvenile stone flounder Platichthys bicoloratus. Journal of Fish Biology, 81(4):1430–1435, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gressler:2012:IAT

[932] L. T. Gressler, T. V. Parodi, A. P. K. Riffel, S. T. DaCosta, and B. Baldisserotto. Immersion anaesthesia with tricaine methanesulphonate or propofol on different sizes and strains of silver catfish Rhamdia quelen. Journal of Fish Biology, 81(4):1436–1445, September 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McAdam:2012:ISC

[933] B. J. McAdam, T. B. Grabowski, and G. Marteinsdóttir. Identification of stock components using morphological markers. Journal of Fish Biology, 81(5):1447–1462, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2012:ETS

[934] D. C. Glover, D. R. DeVries, and R. A. Wright. Effects of temperature, salinity and body size on routine metabolism of coastal largemouth bass Micropterus salmoides. Journal of Fish Biology, 81(5):1463–1478, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Segherloo:2012:PSB

[935] I. Hashemzadeh Segherloo, H. Farahmand, A. Abdoli, L. Bernatchez, C. R. Primmer, A. Swatdipong, M. Karami, and B. Khalili. Phylogenetic status of brown trout Salmo trutta populations in five rivers from the southern Caspian Sea and two inland lake basins, Iran: a morphogenetic approach. Journal of Fish Biology, 81(5):1479–1500, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perrier:2012:ETF

[936] C. Perrier, J. C. Molinero, D. Gerdeaux, and O. Anneville. Effects of temperature and food supply on the growth of whitefish Coregonus lavaretus larvae in an oligotrophic peri-alpine lake. Journal of Fish Biology, 81(5): 1501–1513, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langerhans:2012:GQC

[937] R. B. Langerhans, M. E. Gifford, O. Domínguez-Domínguez, D. García-Bedoya, and T. J. DeWitt. Gambusia quadruncus (Cyprinodontiformes: Poeciliidae): a new species of mosquitofish from east-central México. Journal of Fish Biology, 81(5):1514–1539, October 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ochwada-Doyle:2012:CBT

[938] F. Ochwada-Doyle, D. Roberts, C. Gray, L. Barnes, J. Haddy, and J. Fearman. Characterizing the biological traits and life history of Acanthopagrus (Sparidae) hybrid complexes: implications for conservation and management. Journal of Fish Biology, 81(5):1540–1558, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraitsek:2012:NMP

[939] S. Kraitsek, E. Klossa-Kilia, and G. Kilias. Nuclear and mitochondrial phylogenetic analysis of big-scale sand smelt Atherina boyeri complex in Greece. Journal of Fish Biology, 81(5):1559–1577, October 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espinoza:2012:ODS

[940] M. Espinoza, T. M. Clarke, F. Villalobos-Rojas, and I. S. Wehrtmann. Ontogenetic dietary shifts and feeding ecology of the rasptail skate Raja velezi and the brown smoothhound shark Mustelus henlei along the Pacific coast of Costa Rica, Central America. Journal of Fish Biology, 81 (5):1578–1595, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacoby:2012:EFA

[941] D. M. P. Jacoby, D. W. Sims, and D. P. Croft. The effect of familiarity on aggregation and social behaviour in juvenile small spotted catsharks Scyliorhinus canicula. Journal of Fish Biology, 81(5):1596–1610, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Renchen:2012:IBR

[942] G. F. Renchen, S. J. Pittman, and M. E. Brandt. Investigating the behavioural responses of trapped fishes using underwater video surveillance. Journal of Fish Biology, 81(5):1611–1625, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halfyard:2012:ESM

[943] E. A. Halfyard, A. J. F. Gibson, D. E. Ruzzante, M. J. W. Stokesbury, and F. G. Whoriskey. Estuarine survival and migratory behaviour of Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 81(5): 1626–1645, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strobel:2012:DLF

[944] S. M. Strobel and T. A. Mooney. Detection of low-frequency tones and whale predator sounds by the American sand lance Ammodytes americanus. Journal of Fish Biology, 81(5):1646–1664, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lozano:2012:AGH

[945] C. Lozano, E. D. Houde, R. L. Wingate, and D. H. Secor. Age, growth and hatch dates of ingressing larvae and surviving juveniles of Atlantic menhaden Brevoortia tyrannus. Journal of Fish Biology, 81(5):1665– 1685, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riascos:2012:RHP

[946] J. M. Riascos, M. Vergara, J. Fajardo, V. Villegas, and A. S. Pacheco. The role of hyperiid parasites as a trophic link between jellyfish and fishes. Journal of Fish Biology, 81(5):1686–1695, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Natsopoulou:2012:PPD

[947] M. E. Natsopoulou, S. Pálsson, and G. Á. Ólafsdóttir. Parasites and parallel divergence of the number of individual MHC alleles between sympatric three-spined stickleback Gasterosteus aculeatus morphs in Iceland. Journal of Fish Biology, 81(5):1696–1714, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macdonald:2012:AVL

[948] J. I. Macdonald, D. G. McNeil, and D. A. Crook. Asteriscus v. lapillus: comparing the chemistry of two otolith types and their ability to delineate riverine populations of common carp Cyprinus carpio. Journal of Fish Biology, 81(5):1715–1729, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2012:ESY

[949] R. J. Kennedy, W. W. Crozier, and M. Allen. The effect of stocking with 0+ year age-class Atlantic salmon Salmo salar fry: a case study from the River Bush, Northern Ireland. Journal of Fish Biology, 81(5): 1730–1746, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Segherloo:2012:GDB

[950] I. Hashemzadeh Segherloo, L. Bernatchez, K. Golzarianpour, A. Abdoli, C. R. Primmer, and M. Bakhtiary. Genetic differentiation between two sympatric morphs of the blind Iran cave barb Iranocypris typhlops. Journal of Fish Biology, 81(5):1747–1753, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dupont-Nivet:2012:CIL

[951] M. Dupont-Nivet, C. Robert-Granié, S. Le Guillou, F. Tiquet, and E. Quillet. Comparison of isogenic lines provides evidence that phenotypic plasticity is under genetic control in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 81(5):1754–1762, October 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vidal:2012:PSC

[952] R. R. Vidal, E. W. Carson, and J. R. Gold. Population structure in Chilean hake Merluccius gayi as revealed by mitochondrial DNA sequences. Journal of Fish Biology, 81(5):1763–1772, October 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Longo:2012:CRV

[953] G. O. Longo and S. R. Floeter. Comparison of remote video and diver’s direct observations to quantify reef fishes feeding on benthos in coral and rocky reefs. Journal of Fish Biology, 81(5):1773–1780, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernard:2012:DCN

[954] A. M. Bernard, K. A. Feldheim, and M. S. Shivji. Development and characterization of 11 novel microsatellite loci for the roundscale spearfish Tetrapturus georgii and their cross-species amplification among other istiophorid species. Journal of Fish Biology, 81(5):1781–1786, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2012:BRA

[955] John F. Craig. Book review: Africa in Transition: the Journey from Traditional to Modern — Edited by G. Coulter. Journal of Fish Biology, 81(5):1787–1788, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madsen:2012:BRB

[956] Niels Madsen. Book review: Behaviour of Marine Fishes: Capture Processes and Conservation Challenges — Edited by P. He. Journal of Fish Biology, 81(5):1789–1790, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2012:Cc

[957] Anonymous. Corrigendum. Journal of Fish Biology, 81(5):1791, October 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2012:IBM

[958] M. Li and J. F. Leatherland. The interaction between maternal stress and the ontogeny of the innate immune system during teleost embryogenesis: implications for aquaculture practice. Journal of Fish Biology, 81(6): 1793–1814, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vuori:2012:SSV

[959] K. Vuori, M. Kiljunen, M. Kanerva, M.-L. Koljonen, and M. Nikinmaa. Stock-specific variation of trophic position, diet and environmental stress markers in Atlantic salmon Salmo salar during feeding migrations in the Baltic Sea. Journal of Fish Biology, 81(6):1815–1833, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kopf:2012:RBS

[960] R. K. Kopf, P. S. Davie, D. B. Bromhead, and J. W. Young. Reproductive biology and spatiotemporal patterns of spawning in striped marlin Kajikia audax. Journal of Fish Biology, 81(6):1834–1858, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aalbers:2012:ULT

[961] S. A. Aalbers and C. A. Sepulveda. The utility of a long-term acoustic recording system for detecting white seabass Atractoscion nobilis spawning sounds. Journal of Fish Biology, 81(6):1859–1870, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gabrielsson:2012:DTE

[962] R. M. Gabrielsson, J. Kim, M. R. Reid, C. H. Stirling, M. Numata, and G. P. Closs. Does the trace element composition of brown trout Salmo trutta eggs remain unchanged in spawning redds? Journal of Fish Biology, 81(6):1871–1879, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abolfathi:2012:CGJ

[963] M. Abolfathi, A. Hajimoradloo, R. Ghorbani, and A. Zamani. Compensatory growth in juvenile roach Rutilus caspicus: effect of starvation and re-feeding on growth and digestive surface area. Journal of Fish Biology, 81(6):1880–1890, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boguski:2012:GDE

[964] D. A. Boguski, S. B. Reid, D. H. Goodman, and M. F. Docker. Genetic diversity, endemism and phylogeny of lampreys within the genus Lampetra sensu stricto (Petromyzontiformes: Petromyzontidae) in western North America. Journal of Fish Biology, 81(6):1891–1914, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naspleda:2012:MVB

[965] J. Naspleda, A. Vila-Gispert, M. G. Fox, L. Zamora, and A. Ruiz-Navarro. Morphological variation between non-native lake- and streamdwelling pumpkinseed Lepomis gibbosus in the Iberian Peninsula. Journal of Fish Biology, 81(6):1915–1935, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lek:2012:DMS

[966] E. Lek, D. V. Fairclough, N. G. Hall, S. A. Hesp, and I. C. Potter. Do the maximum sizes, ages and patterns of growth of three reef-dwelling labrid species at two latitudes differ in a manner conforming to the metabolic theory of ecology? Journal of Fish Biology, 81(6):1936–1962, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ornelas-Garcia:2012:GDS

[967] C. P. Ornelas-García, F. Alda, E. Díaz-Pardo, A. Gutiérrez-Hernández, and I. Doadrio. Genetic diversity shaped by historical and recent factors in the live-bearing twoline skiffia Neotoca bilineata. Journal of Fish Biology, 81(6):1963–1984, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2012:DSV

[968] X. Zhang, K. Shimoda, K. Ura, S. Adachi, and Y. Takagi. Developmental structure of the vertebral column, fins, scutes and scales in bester sturgeon, a hybrid of beluga Huso huso and sterlet Acipenser ruthenus. Journal of Fish Biology, 81(6):1985–2004, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2012:FTS

[969] T. D. Harrison and A. K. Whitfield. Fish trophic structure in estuaries, with particular emphasis on estuarine typology and zoogeography. Journal of Fish Biology, 81(6):2005–2029, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boettiger:2012:PRE

[970] C. Boettiger, D. T. Lang, and P. C. Wainwright. rfishbase: exploring, manipulating and visualizing FishBase data from R. Journal of Fish Biology, 81(6):2030–2039, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peyrusse:2012:PAL

[971] K. Peyrusse, P. Sasal, and D. Lecchini. Parasite acquisition by larval coral-reef fishes. Journal of Fish Biology, 81(6):2040–2043, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruggeri:2012:INM

[972] P. Ruggeri, A. Splendiani, M. Giovannotti, P. Nisi Cerioni, and V. Caputo. Isolation of novel microsatellite loci in the black goby Gobius niger and cross-amplification in other gobiid species (Perciformes, Gobiidae). Journal of Fish Biology, 81(6):2044–2052, November 2012. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mourier:2012:MRM

[973] J. Mourier. Manta rays in the Marquesas Islands: first records of Manta birostris in French Polynesia and most easterly location of Manta alfredi in the Pacific Ocean, with notes on their distribution. Journal of Fish Biology, 81(6):2053–2058, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Easy:2012:ITA

[974] R. H. Easy, E. A. Trippel, M. D. B. Burt, and D. K. Cone. Identification of transferrin in Atlantic cod Gadus morhua epidermal mucus. Journal of Fish Biology, 81(6):2059–2063, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa:2012:TPB

[975] C. Correa. Tissue preservation biases in stable isotopes of fishes and molluscs from Patagonian lakes. Journal of Fish Biology, 81(6):2064– 2073, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bos:2012:LCD

[976] A. R. Bos and G. S. Gumanao. The lunar cycle determines availability of coral-reef fishes at fish markets. Journal of Fish Biology, 81(6):2074– 2079, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Afonso:2012:EGT

[977] A. S. Afonso, F. H. V. Hazin, R. R. Barreto, F. M. Santana, and R. P. Lessa. Extraordinary growth in tiger sharks Galeocerdo cuvier from the South Atlantic Ocean. Journal of Fish Biology, 81(6):2080–2085, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzoni:2012:ESM

[978] M. Lorenzoni, D. Giannetto, G. Maio, E. Pizzul, L. Pompei, P. Turin, S. Vincenzi, and A. Crivelli. Empirical standard mass equation for Salmo marmoratus. Journal of Fish Biology, 81(6):2086–2091, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrades:2012:NRW

[979] R. Andrades, H. T. Pinheiro, R. G. Santos, A. S. Martins, and P. A. S. Costa. A new record of whale shark Rhincodon typus in Brazilian waters: a report of association with Caranx crysos. Journal of Fish Biology, 81(6):2092–2094, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2012:BRF

[980] Ian J. Winfield. Book review: FAO Statistical Yearbook 2012: World Food and Agriculture — Edited by A. Prakash and M. Stigler. Journal of Fish Biology, 81(6):2095–2096, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2012:BRI

[981] Ian J. Winfield. Book review: Ikawai: Freshwater Fishes in Maori Culture and Ecomomy — Edited by R. M. McDowall. Journal of Fish Biology, 81(6):2097–2098, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2012:BRR

[982] Ian J. Winfield. Book review: Review of the State of the World Fishery Resources: Inland Fisheries — Edited by R. Welcomme. Journal of Fish Biology, 81(6):2099, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2012:BRS

[983] L. A. Jawad. Book review: Sparidae: Biology and Aquaculture of Gilthead Sea Bream and Other Species — Edited by M. A. Pavlidis and C. C. Mylonas. Journal of Fish Biology, 81(6):2100–2101, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2012:PFB

[984] F. A. Huntingford. The physiology of fish behaviour: a selective review of developments over the past 40 years. Journal of Fish Biology, 81(7): 2103–2126, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2012:SPF

[985] R. F. Oliveira. Social plasticity in fish: integrating mechanisms and function. Journal of Fish Biology, 81(7):2127–2150, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebbesson:2012:EEF

[986] L. O. E. Ebbesson and V. A. Braithwaite. Environmental effects on fish neural plasticity and cognition. Journal of Fish Biology, 81(7):2151–2174, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloman:2012:UPB

[987] K. A. Sloman and P. L. McNeil. Using physiology and behaviour to understand the responses of fish early life stages to toxicants. Journal of Fish Biology, 81(7):2175–2198, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:2012:TAB

[988] T. W. K. Fraser, P. G. Fjelldal, J. E. Skjæraasen, T. Hansen, and I. Mayer. Triploidy alters brain morphology in pre-smolt Atlantic salmon Salmo salar: possible implications for behaviour. Journal of Fish Biology, 81(7):2199–2212, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shaw:2012:BET

[989] J. C. Shaw and Ø. Øverli. Brain-encysting trematodes and altered monoamine activity in naturally infected killifish Fundulus parvipinnis. Journal of Fish Biology, 81(7):2213–2222, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slavik:2012:DDE

[990] O. Slavík and P. Horký. Diel dualism in the energy consumption of the European catfish Silurus glanis. Journal of Fish Biology, 81(7):2223– 2234, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gebhardt:2012:EBD

[991] K. Gebhardt, M. Böhme, and G. von der Emde. Electrocommunication behaviour during social interactions in two species of pulse-type weakly electric fishes (Mormyridae). Journal of Fish Biology, 81(7):2235–2254, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fechler:2012:MGM

[992] K. Fechler, D. Holtkamp, G. Neusel, J. I. Sanguinetti-Scheck, R. Budelli, and G. von der Emde. Mind the gap: the minimal detectable separation distance between two objects during active electrolocation. Journal of Fish Biology, 81(7):2255–2276, December 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logan:2013:CSI

[993] J. M. Logan and K. L. Dodge. Comment on ‘Stable isotopes challenge the perception of ocean sunfish Mola mola as obligate jellyfish predators’. Journal of Fish Biology, 82(1):1–9, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [689, 994].

Harrod:2013:RLD

[994] C. Harrod, J. Syväranta, L. Kubicek, V. Cappanera, and J. D. R. Houghton. Reply to Logan & Dodge: ‘Stable isotopes challenge the perception of ocean sunfish Mola mola as obligate jellyfish predators’. Journal of Fish Biology, 82(1):10–16, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [689, 993].

Rabitsch:2013:TCT

[995] W. Rabitsch, N. Milasowszky, S. Nehring, C. Wiesner, C. Wolter, and F. Essl. The times are changing: temporal shifts in patterns of fish invasions in central European fresh waters. Journal of Fish Biology, 82 (1):17–33, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2013:EGP

[996] X. D. Liu, G. T. Zhao, M. Y. Cai, and Z. Y. Wang. Estimated genetic parameters for growth-related traits in large yellow croaker Larimichthys crocea using microsatellites to assign parentage. Journal of Fish Biology, 82(1):34–41, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shirak:2013:DEA

[997] A. Shirak, Y. Palti, O. Bern, T. D. Kocher, E. Gootwine, E. Seroussi, G. Hulata, M. Ron, and R. R. Avtalion. A deleterious effect associated with UNH159 is attenuated in twin embryos of an inbred line of blue tilapia Oreochromis aureus. Journal of Fish Biology, 82(1):42–53, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arceo-Carranza:2013:DNT

[998] D. Arceo-Carranza, M. E. Vega-Cendejas, and M. Hernández de Santillana. Day and night trophic variations of dominant fish species in a lagoon influenced by freshwater seeps. Journal of Fish Biology, 82(1):54– 68, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walter:2013:CSO

[999] B. E. Walter. Cranial skeletogenesis and osteology of the redeye tetra Moenkhausia sanctaefilomenae. Journal of Fish Biology, 82(1):69–95, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wearmouth:2013:IRE

[1000] V. J. Wearmouth, E. J. Southall, D. Morritt, and D. W. Sims. Identifying reproductive events using archival tags: egg-laying behaviour of the small spotted catshark Scyliorhinus canicula. Journal of Fish Biology, 82(1): 96–110, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lynch:2013:TPG

[1001] M. P. Lynch and A. F. Mensinger. Temporal patterns in growth and survival of the round goby Neogobius melanostomus. Journal of Fish Biology, 82(1):111–124, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2013:PGS

[1002] Y. Zhang, S. Wang, J. Li, X. Zhang, L. Jiang, P. Xu, C. Lu, Y. Wan, and X. Sun. Primary genome scan for complex body shape-related traits in the common carp Cyprinus carpio. Journal of Fish Biology, 82(1):125– 140, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinkerton:2013:DTN

[1003] M. H. Pinkerton, J. Forman, S. J. Bury, J. Brown, P. Horn, and R. L. O’Driscoll. Diet and trophic niche of Antarctic silverfish Pleuragramma antarcticum in the Ross Sea, Antarctica. Journal of Fish Biology, 82 (1):141–164, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ngoma:2013:GMS

[1004] E. Ng’oma, S. Valdesalici, K. Reichwald, and A. Cellerino. Genetic and morphological studies of Nothobranchius (Cyprinodontiformes) from Malaŵi with description of Nothobranchius wattersi sp. nov. Journal of Fish Biology, 82(1):165–188, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guo:2013:CEC

[1005] Z. Guo, G. L. Wang, J. P. Fu, and P. Nie. Characterization and expression of Cd8 molecules in mandarin fish Siniperca chuatsi. Journal of Fish Biology, 82(1):189–205, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urke:2013:STP

[1006] H. A. Urke, T. Kristensen, J. V. Arnekleiv, T. O. Haugen, G. Kjærstad, S. O. Stefansson, L. O. E. Ebbesson, and T. O. Nilsen. Seawater tolerance and post-smolt migration of wild Atlantic salmon Salmo salar × brown trout S. trutta hybrid smolts. Journal of Fish Biology, 82(1):206–227, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2013:SSS

[1007] S. M. Taylor and M. B. Bennett. Size, sex and seasonal patterns in the assemblage of Carcharhiniformes in a sub-tropical bay. Journal of Fish Biology, 82(1):228–241, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardner:2013:CAS

[1008] J. P. A. Gardner and C. D. Struthers. Comparisons among survey methodologies to test for abundance and size of a highly targeted fish species. Journal of Fish Biology, 82(1):242–262, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shimose:2013:ODS

[1009] T. Shimose, H. Watanabe, T. Tanabe, and T. Kubodera. Ontogenetic diet shift of age-0 year Pacific bluefin tuna Thunnus orientalis. Journal of Fish Biology, 82(1):263–276, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rees:2013:LTI

[1010] D. Rees and I. Byrkjedal. Lack of taxonomic information from parietal spine size invalidates subspecies in the Atlantic hookear sculpin Artediellus atlanticus. Journal of Fish Biology, 82(1):277–285, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espinoza:2013:DCD

[1011] M. Espinoza, T. M. Clarke, F. Villalobos-Rojas, and I. S. Wehrtmann. Diet composition and diel feeding behaviour of the banded guitarfish Zapteryx xyster along the Pacific coast of Costa Rica, Central America.

Journal of Fish Biology, 82(1):286–305, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha:2013:RBG

[1012] F. Rocha and O. B. F. Gadig. Reproductive biology of the guitarfish Rhinobatos percellens (Chondrichthyes, Rhinobatidae) from the São Paulo Coast, Brazil, western South Atlantic Ocean. Journal of Fish Biology, 82(1):306–317, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aikio:2013:OGS

[1013] S. Aikio, G. Herczeg, A. Kuparinen, and J. Merilä. Optimal growth strategies under divergent predation pressure. Journal of Fish Biology, 82(1):318–331, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dulcic:2013:PFP

[1014] J. Dulčić and B. Dragičević. Paranthias furcifer (Perciformes: Serranidae), a new alien fish in the Mediterranean Sea. Journal of Fish Biology, 82(1):332–337, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gerry:2013:VDF

[1015] S. P. Gerry, M. Vogelzang, J. M. Ascher, and D. J. Ellerby. Variation in the diet and feeding morphology of polyphenic Lepomis macrochirus. Journal of Fish Biology, 82(1):338–346, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Myszkowski:2013:CGC

[1016] L. Myszkowski. Compensatory growth, condition and food utilization in barbel Barbus barbus juveniles reared at different feeding periodicities with a dry diet. Journal of Fish Biology, 82(1):347–353, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kagawa:2013:SRD

[1017] N. Kagawa. Social rank-dependent expression of arginine vasotocin in distinct preoptic regions in male Oryzias latipes. Journal of Fish Biology, 82(1):354–363, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everson:2013:BRL

[1018] Inigo Everson. Book review: Life in the World’s Oceans: Diversity, Distribution, and Abundance — Edited by A. D. McIntyre. Journal of Fish Biology, 82(1):364–365, January 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sequeira:2013:IGC

[1019] A. M. M. Sequeira, C. Mellin, M. G. Meekan, D. W. Sims, and C. J. A. Bradshaw. Inferred global connectivity of whale shark Rhincodon typus populations. Journal of Fish Biology, 82(2):367–389, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulcini:2013:DSM

[1020] D. Pulcini, P. A. Wheeler, S. Cataudella, T. Russo, and G. H. Thorgaard. Domestication shapes morphology in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 82(2):390–407, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kent:2013:TSP

[1021] J. Kent, G. Jenkins, and S. Acevedo. Temporal and spatial patterns in ichthyoplankton assemblages in bay and open coastal environments. Journal of Fish Biology, 82(2):408–429, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elgueta:2013:TIS

[1022] A. Elgueta, J. González, D. E. Ruzzante, S. J. Walde, and E. Habit. Trophic interference by Salmo trutta on Aplochiton zebra and Aplochiton taeniatus in southern Patagonian lakes. Journal of Fish Biology, 82(2): 430–443, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dhiyebi:2013:WCM

[1023] H. A. Dhiyebi, M. J. O’Donnell, and P. A. Wright. Water chemistry in the microenvironment of rainbow trout Oncorhynchus mykiss embryos is affected by development, the egg capsule and crowding. Journal of Fish Biology, 82(2):444–457, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Llompart:2013:BTG

[1024] F. M. Llompart, D. C. Colautti, T. Maiztegui, A. M. Cruz-Jiménez, and C. R. M. Baigún. Biological traits and growth patterns of pejerrey Odontesthes argentinensis. Journal of Fish Biology, 82(2):458–474, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takegaki:2013:TCD

[1025] T. Takegaki, T. Kaneko, and Y. Matsumoto. Tactic changes in dusky frillgoby Bathygobius fuscus sneaker males: effects of body size and nest availability. Journal of Fish Biology, 82(2):475–491, February 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2013:IMG

[1026] R. Wang, M. Chen, C. Li, L.-P. Li, X. Gan, J. Huang, A.-Y. Lei, Z.-H. Xu, and W.-W. Liang. Identification of multiple genes and their expression profiles in four strains of Oreochromis spp. in response to Streptococcus iniae. Journal of Fish Biology, 82(2):492–504, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weiss:2013:AND

[1027] S. J. Weiss, T. Kopun, and S. Sušnik Bajec. Assessing natural and disturbed population structure in European grayling Thymallus thymallus: melding phylogeographic, population genetic and jurisdictional perspectives for conservation planning. Journal of Fish Biology, 82(2):505–521, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Locke:2013:EDI

[1028] S. A. Locke, G. Bulté, M. R. Forbes, and D. J. Marcogliese. Estimating diet in individual pumpkinseed sunfish Lepomis gibbosus using stomach contents, stable isotopes and parasites. Journal of Fish Biology, 82(2): 522–537, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheaves:2013:RPS

[1029] M. Sheaves and B. Molony. Reproductive periodicity of the sparid, Acanthopagrus pacificus, on a hierarchy of temporal scales. Journal of Fish Biology, 82(2):538–554, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Francova:2013:OBC

[1030] K. Francová and M. Ondračková. Overwinter body condition, mortality and parasite infection in two size classes of 0+ year juvenile European bitterling Rhodeus amarus. Journal of Fish Biology, 82(2):555–568, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woods:2013:RPD

[1031] P. J. Woods, D. Young, S. Skúlason, S. S. Snorrason, and T. P. Quinn. Resource polymorphism and diversity of Arctic charr Salvelinus alpinus in a series of isolated lakes. Journal of Fish Biology, 82(2):569–587, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2013:CMM

[1032] Y. X. Liu, H. Z. Han, Q. L. Wang, L. Jiang, S. L. Wang, X. Y. Zhang, Y. Liu, Y. F. Wang, Y. J. Liu, and H. J. Liu. Choice of microsatellite markers for identifying homozygosity of mitotic gynogenetic diploids in Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 82 (2):588–599, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muir:2013:OSM

[1033] A. M. Muir, P. Vecsei, T. C. Pratt, C. C. Krueger, M. Power, and J. D. Reist. Ontogenetic shifts in morphology and resource use of cisco Coregonus artedi. Journal of Fish Biology, 82(2):600–617, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madison:2013:EPM

[1034] B. N. Madison, J. W. Heath, D. D. Heath, and N. J. Bernier. Effect of parental mate choice and semi-natural early rearing environment on the growth performance and seawater tolerance of Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 82(2):618–636, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gratton:2013:GDH

[1035] P. Gratton, G. Allegrucci, A. Gandolfi, and V. Sbordoni. Genetic differentiation and hybridization in two naturally occurring sympatric trout Salmo spp. forms from a small karstic lake. Journal of Fish Biology, 82 (2):637–657, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woolley:2013:OBD

[1036] L. D. Woolley and J. G. Qin. Ontogeny of body density and the swimbladder in yellowtail kingfish Seriola lalandi larvae. Journal of Fish Biology, 82(2):658–670, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poletto:2013:BSP

[1037] J. B. Poletto, D. E. Cocherell, A. P. Klimley, J. J. Cech, Jr., and N. A. Fangue. Behavioural salinity preferences of juvenile green sturgeon Acipenser medirostris acclimated to fresh water and full-strength salt water. Journal of Fish Biology, 82(2):671–685, February 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prigge:2013:TMS

[1038] E. Prigge, L. Marohn, and R. Hanel. Tracking the migratory success of stocked European eels Anguilla anguilla in the Baltic Sea. Journal of Fish Biology, 82(2):686–699, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soldo:2013:FMR

[1039] A. Soldo. First marine record of marble trout Salmo marmoratus. Journal of Fish Biology, 82(2):700–702, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalous:2013:NMD

[1040] L. Kalous, K. Rylková, J. Bohlen, R. Šanda, and M. Petrtýl. New mtDNA data reveal a wide distribution of the Japanese ginbuna Carassius langsdorfii in Europe. Journal of Fish Biology, 82(2):703–707, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kegler:2013:NES

[1041] P. Kegler, A. Kunzmann, S. Bröhl, and N. A. Herbert. No evidence of shelter providing a metabolic advantage to the false clown anemonefish Amphiprion ocellaris. Journal of Fish Biology, 82(2):708–713, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguirre-Villasenor:2013:NEP

[1042] H. Aguirre-Villaseñor, C. Salas-Singh, J. Madrid-Vera, J. Martínez-Ortiz, D. A. Didier, and D. A. Ebert. New eastern Pacific Ocean records of Hydrolagus melanophasma, with annotations of a juvenile female. Journal of Fish Biology, 82(2):714–724, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Utne-Palm:2013:IRB

[1043] A. C. Utne-Palm, L. Locatello, I. Mayer, M. J. Gibbons, and M. B. Rasotto. An insight into the reproductive biology of the bearded goby Sufflogobius bibarbatus. Journal of Fish Biology, 82(2):725–731, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2013:R

[1044] Anonymous. Referees 2012. Journal of Fish Biology, 82(2):732–739, February 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuenfeldt:2013:AMA

[1045] S. Neuenfeldt, D. Righton, F. Neat, P. J. Wright, H. Svedäng, K. Michalsen, S. Subbey, P. Steingrund, V. Thorsteinsson, C. Pampoulie, K. H. Andersen, M. W. Pedersen, and J. Metcalfe. Analysing migrations of Atlantic cod Gadus morhua in the north-east Atlantic Ocean: then, now and the future. Journal of Fish Biology, 82(3):741–763, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Currey:2013:CBT

[1046] L. M. Currey, A. J. Williams, B. D. Mapstone, C. R. Davies, G. Carlos, D. J. Welch, C. A. Simpfendorfer, A. C. Ballagh, A. L. Penny, E. M. Grandcourt, A. Mapleston, A. S. Wiebkin, and K. Bean. Comparative biology of tropical Lethrinus species (Lethrinidae): challenges for multispecies management. Journal of Fish Biology, 82(3):764–788, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Couturier:2013:PEI

[1047] C. S. Couturier, N. G. Andersen, C. Audet, and D. Chabot. Prey exoskeletons influence the course of gastric evacuation in Atlantic cod Gadus morhua. Journal of Fish Biology, 82(3):789–805, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Livi:2013:GSB

[1048] S. Livi, S. de Innocentiis, A. Longobardi, S. Cataudella, L. Tancioni, M. Rampacci, and G. Marino. Genetic structure of Barbus spp. populations in the Marches region of central Italy and its relevance to conservation actions. Journal of Fish Biology, 82(3):806–826, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OBrien:2013:GCG

[1049] C. O’Brien, L. Unruh, C. Zimmerman, W. E. Bradshaw, C. M. Holzapfel, and W. A. Cresko. Geography of the circadian gene clock and photoperiodic response in western North American populations of the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 82(3):827– 839, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Specziar:2013:WLD

[1050] A. Specziár, Á. I. György, and T. Erős. Within-lake distribution patterns of fish assemblages: the relative roles of spatial, temporal and random environmental factors in assessing fish assemblages using gillnets in a large and shallow temperate lake. Journal of Fish Biology, 82(3):840– 855, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanna:2013:MDV

[1051] D. Sanna, F. Biagi, H. B. Alaya, F. Maltagliati, A. Addis, A. Romero, J. De Juan, J.-P. Quignard, A. Castelli, P. Franzoi, P. Torricelli, M. Casu, M. Carcupino, and P. Francalacci. Mitochondrial DNA variability of the pipefish Syngnathus abaster. Journal of Fish Biology, 82(3):856– 876, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2013:ADH

[1052] J. T. Kelly and J. M. Hanson. Abundance, distribution and habitat characteristics of winter skate Leucoraja ocellata in the southern Gulf of St Lawrence: a population on the brink of extirpation? Journal of Fish Biology, 82(3):877–892, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landeira-Dabarca:2013:CFI

[1053] A. Landeira-Dabarca, C. Sieiro, and M. Álvarez. Change in food ingestion induces rapid shifts in the diversity of microbiota associated with cutaneous mucus of Atlantic salmon Salmo salar. Journal of Fish Biology, 82(3):893–906, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dell:2013:ETD

[1054] Q. Dell, S. P. Griffiths, M. L. Tonks, W. A. Rochester, M. J. Miller, M. A. Duggan, T. D. van der Velde, R. D. Pillans, G. J. Coman, R. H. Bustamante, and D. A. Milton. Effects of trawling on the diets of common demersal fish by-catch of a tropical prawn trawl fishery. Journal of Fish Biology, 82(3):907–926, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vila:2013:PPA

[1055] I. Vila, P. Morales, S. Scott, E. Poulin, D. Véliz, C. Harrod, and M. A. Méndez. Phylogenetic and phylogeographic analysis of the genus Orestias (Teleostei: Cyprinodontidae) in the southern Chilean Altiplano: the relevance of ancient and recent divergence processes in speciation. Journal of Fish Biology, 82(3):927–943, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlsson:2013:RES

[1056] J. Carlsson, D. T. Gauthier, J. E. L. Carlsson, J. P. Coughlan, E. Dillane, R. D. Fitzgerald, U. Keating, P. McGinnity, L. Mirimin, and T. F. Cross. Rapid, economical single-nucleotide polymorphism and microsatellite discovery based on de novo assembly of a reduced representation genome in a non-model organism: a case study of Atlantic cod Gadus morhua. Journal of Fish Biology, 82(3):944–958, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2013:MSA

[1057] J. T. Kelly and J. M. Hanson. Maturity, size at age and predator– prey relationships of winter skate Leucoraja ocellata in the southern Gulf of St Lawrence: potentially an undescribed endemic facing extirpation. Journal of Fish Biology, 82(3):959–978, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mourier:2013:PSS

[1058] J. Mourier, S. C. Mills, and S. Planes. Population structure, spatial distribution and life-history traits of blacktip reef sharks Carcharhinus melanopterus. Journal of Fish Biology, 82(3):979–993, March 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janac:2013:RRP

[1059] M. Janáč, P. Jurajda, L. Kružíková, K. Roche, and V. Prášek. Reservoir to river passage of age-0+ year fishes, indication of a dispersion pathway for a non-native species. Journal of Fish Biology, 82(3):994–1010, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanaka:2013:FET

[1060] H. Tanaka, C. Sassa, S. Ohshimo, and I. Aoki. Feeding ecology of two lanternfishes Diaphus garmani and Diaphus chrysorhynchus. Journal of Fish Biology, 82(3):1011–1031, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jackson:2013:FPJ

[1061] L. A. Jackson, C. F. Rakocinski, and R. B. Blaylock. Feeding performance of juvenile hatchery-reared spotted seatrout Cynoscion nebulosus. Journal of Fish Biology, 82(3):1032–1046, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2013:EAS

[1062] X. He, S. Lu, M. Liao, X. Zhu, M. Zhang, S. Li, X. You, and J. Chen. Effects of age and size on critical swimming speed of juvenile Chinese sturgeon Acipenser sinensis at seasonal temperatures. Journal of Fish Biology, 82(3):1047–1056, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vilizzi:2013:IME

[1063] L. Vilizzi and G. H. Copp. Interstitial movement and emergence of barbel Barbus barbus free embryos and larvae. Journal of Fish Biology, 82(3): 1057–1063, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rivas:2013:FRA

[1064] L. Peña Rivas, E. Azzurro, and D. Lloris. First record of the Atlantic bumper Chloroscombrus chrysurus (Teleostei: Carangidae) in the Mediterranean Sea. Journal of Fish Biology, 82(3):1064–1067, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kusterle:2013:GMS

[1065] S. Kusterle, E. Halttunen, E. B. Thorstad, T. F. Næsje, J. L. A. Jensen, A. Gallo-Bueno, E. Olague, and A. H. Rikardsen. The gill maggot Salmincola salmoneus as an indicator of repeat spawning in Atlantic salmon Salmo salar. Journal of Fish Biology, 82(3):1068–1073, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reinardy:2013:ADD

[1066] H. C. Reinardy, E. Skippins, T. B. Henry, and A. N. Jha. Assessment of DNA damage in sperm after repeated non-invasive sampling in zebrafish Danio rerio. Journal of Fish Biology, 82(3):1074–1081, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janse:2013:PTU

[1067] M. Janse, A. L. Kappe, and B. L. M. Van Kuijk. Paternity testing using the poisonous sting in captive white-spotted eagle rays Aetobatus narinari: a non-invasive tool for captive sustainability programmes. Journal of Fish Biology, 82(3):1082–1085, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clotfelter:2013:ETD

[1068] E. D. Clotfelter, S. J. H. Lapidus, and A. C. Brown. The effects of temperature and dissolved oxygen on antioxidant defences and oxidative damage in the fathead minnow Pimephales promelas. Journal of Fish Biology, 82(3):1086–1092, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hume:2013:SMM

[1069] J. B. Hume, C. E. Adams, B. Mable, and C. W. Bean. Sneak male mating tactics between lampreys (Petromyzontiformes) exhibiting alternative life-history strategies. Journal of Fish Biology, 82(3):1093–1100, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2013:BRO

[1070] Michel J. Kaiser. Book review: Overfishing: What Everyone Needs to Know — Edited by R. Hilborn and U. Hilborn. Journal of Fish Biology, 82(3):1101–1102, March 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Novomeska:2013:MVB

[1071] A. Novomeská, S. Katina, G. H. Copp, G. Pedicillo, M. Lorenzoni, L. Pompei, J. Cucherousset, and V. Kováč. Morphological variability of black bullhead Ameiurus melas in four non-native European populations. Journal of Fish Biology, 82(4):1103–1118, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warry:2013:UNE

[1072] F. Y. Warry, P. Reich, J. S. Hindell, J. McKenzie, and A. Pickworth. Using new electrofishing technology to amp-up fish sampling in estuarine habitats. Journal of Fish Biology, 82(4):1119–1137, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagrodski:2013:EES

[1073] A. Nagrodski, K. J. Murchie, K. M. Stamplecoskie, C. D. Suski, and S. J. Cooke. Effects of an experimental short-term cortisol challenge on the behaviour of wild creek chub Semotilus atromaculatus in mesocosm and stream environments. Journal of Fish Biology, 82(4):1138–1158, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitney:2013:PMT

[1074] C. K. Whitney, S. G. Hinch, and D. A. Patterson. Provenance matters: thermal reaction norms for embryo survival among sockeye salmon Oncorhynchus nerka populations. Journal of Fish Biology, 82(4):1159– 1176, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pratchett:2013:SCD

[1075] M. S. Pratchett, N. A. J. Graham, and A. J. Cole. Specialist corallivores dominate butterflyfish assemblages in coral-dominated reef habitats. Journal of Fish Biology, 82(4):1177–1191, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chin:2013:PSR

[1076] A. Chin, A. J. Tobin, M. R. Heupel, and C. A. Simpfendorfer. Population structure and residency patterns of the blacktip reef shark Carcharhinus melanopterus in turbid coastal environments. Journal of Fish Biology, 82(4):1192–1210, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russell:2013:AGT

[1077] D. J. Russell, F. E. Thomson, and P. A. Thuesen. Age and growth of two newly established invasive populations of Tilapia mariae in northern Australia. Journal of Fish Biology, 82(4):1211–1225, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:2013:ELP

[1078] P. H. C. Pereira and B. P. Ferreira. Effects of life phase and schooling patterns on the foraging behaviour of coral-reef fishes from the genus Haemulon. Journal of Fish Biology, 82(4):1226–1238, April 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okamura:2013:MIA

[1079] H. Okamura, A. E. Punt, Y. Semba, and M. Ichinokawa. Marginal increment analysis: a new statistical approach of testing for temporal periodicity in fish age verification. Journal of Fish Biology, 82(4):1239–1249, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Umberger:2013:EMI

[1080] C. M. Umberger, I. de Buron, W. A. Roumillat, and E. J. McElroy. Effects of a muscle-infecting parasitic nematode on the locomotor performance of their fish host. Journal of Fish Biology, 82(4):1250–1258, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langeani:2013:HIN

[1081] F. Langeani and C. R. Moreira. Hemiodus iratapuru, a new species of Hemiodontidae from the Rio Jari, Amazon Basin, Brazil (Teleostei, Characiformes). Journal of Fish Biology, 82(4):1259–1268, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Patterson:2013:EBM

[1082] J. T. Patterson, S. D. Mims, and R. A. Wright. Effects of body mass and water temperature on routine metabolism of American paddlefish Polyodon spathula. Journal of Fish Biology, 82(4):1269–1280, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ozer:2013:GER

[1083] F. Ozer and M. V. Ashley. Genetic evaluation of remnant and translocated shiners, Notropis heterodon and Notropis heterolepis. Journal of Fish Biology, 82(4):1281–1296, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lebreton:2013:TSY

[1084] B. Lebreton, P. Richard, G. Guillou, and G. F. Blanchard. Trophic shift in young-of-the-year Mugilidae during salt-marsh colonization. Journal of Fish Biology, 82(4):1297–1307, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaifu:2013:HSP

[1085] K. Kaifu, K. Yokouchi, M. J. Miller, J. Aoyama, and K. Tsukamoto. Head-shape polymorphism in Japanese eels Anguilla japonica in relation to differences of somatic growth in freshwater and brackish habitats. Journal of Fish Biology, 82(4):1308–1320, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bani:2013:CMS

[1086] A. Bani, S. Poursaeid, and V. M. Tuset. Comparative morphology of the sagittal otolith in three species of south Caspian gobies. Journal of Fish Biology, 82(4):1321–1332, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rhodes:2013:RBS

[1087] K. L. Rhodes, B. M. Taylor, C. B. Wichilmel, E. Joseph, R. J. Hamilton, and G. R. Almany. Reproductive biology of squaretail coralgrouper Plectropomus areolatus using age-based techniques. Journal of Fish Biology, 82(4):1333–1350, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Decru:2013:RLG

[1088] E. Decru, E. Vreven, and J. Snoeks. A revision of the Lower Guinean Hepsetus species (Characiformes; Hepsetidae) with the description of Hepsetus kingsleyae sp. nov. Journal of Fish Biology, 82(4):1351–1375, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horn:2013:MPI

[1089] P. L. Horn, J. S. Forman, and M. R. Dunn. Moon phase influences the diet of southern Ray’s bream Brama australis. Journal of Fish Biology, 82(4):1376–1389, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wasylenko:2013:CCA

[1090] B. A. Wasylenko, P. J. Blanchfield, and G. G. Pyle. Chemosensory cues attract lake trout Salvelinus namaycush and an egg predator to the spawning substratum. Journal of Fish Biology, 82(4):1390–1397, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kindsvater:2013:STD

[1091] H. K. Kindsvater and S. H. Alonzo. Short-term dynamics of nest occupancy in an allopaternal species, the tessellated darter Etheostoma olmstedi. Journal of Fish Biology, 82(4):1398–1402, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belo:2013:HRC

[1092] M. A. A. Belo, D. G. F. Souza, V. P. Faria, E. J. R. Prado, F. R. Moraes, and E. M. Onaka. Haematological response of curimbas Prochilodus lineatus, naturally infected with Neoechinorhynchus curemai. Journal of Fish Biology, 82(4):1403–1410, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2013:CST

[1093] J. G. Davidsen, E. B. Thorstad, H. Baktoft, S. Aune, F. Økland, and A. H. Rikardsen. Can sea trout Salmo trutta compromise successful eradication of Gyrodactylus salaris by hiding from CFT legumin (rotenone) treatments? Journal of Fish Biology, 82(4):1411–1418, April 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2013:FRD

[1094] C. M. Wagner, P. H. Rice, and A. P. Pease. First record of dicephalia in a bull shark Carcharhinus leucas (Chondrichthyes: Carcharhinidae) foetus from the Gulf of Mexico, U.S.A. Journal of Fish Biology, 82(4):1419– 1422, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liberoff:2013:SSF

[1095] A. L. Liberoff, C. M. Riva Rossi, M. L. Fogel, J. E. Ciancio, and M. A. Pascual. Shifts in δ ^{15}N signature following the onset of exogenous feeding in rainbow trout Oncorhynchus mykiss: importance of combining length and age data. Journal of Fish Biology, 82(4):1423–1432, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ball:2013:FCC

[1096] R. E. Ball, C. S. Jones, A. Lynghammar, L. R. Noble, and A. M. Griffiths. The first confirmed cases of full albinism in rajid species. Journal of Fish Biology, 82(4):1433–1440, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2013:NBS

[1097] X. L. Ma, L. W. Zheng, L. T. Mao, L. B. Zhou, and A. L. Wang. Neuropeptide Y in black seabream Acanthopagrus schlegelii: identification, distribution and mRNA expression responses to ghrelin. Journal of Fish Biology, 82(4):1441–1447, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2013:BRS

[1098] Ian J. Winfield. Book review: Scientific Communication for Natural Resource Professionals — Edited by C. A. Jennings, T. E. Lauer and B. Vondracek. Journal of Fish Biology, 82(4):1448–1449, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2013:BRR

[1099] Stephen J. Lockwood. Book review: Reef Fish Spawning Aggregations: Biology, Research and Management — Edited by Y. Savody de Mitcheson & P. L. Colin. Journal of Fish Biology, 82(4):1450–1451, April 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:2013:SPS

[1100] A. Pérez and N. N. Fabré. Spatial population structure of the Neotropical tiger catfish Pseudoplatystoma metaense: skull and otolith shape variation. Journal of Fish Biology, 82(5):1453–1468, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Furey:2013:FSM

[1101] N. B. Furey, M. A. Dance, and J. R. Rooker. Fine-scale movements and habitat use of juvenile southern flounder Paralichthys lethostigma in an estuarine seascape. Journal of Fish Biology, 82(5):1469–1483, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Utz:2013:HDC

[1102] R. M. Utz, C. F. Mesick, B. J. Cardinale, and T. Dunne. How does coarse gravel augmentation affect early-stage Chinook salmon Oncorhynchus tshawytscha embryonic survivorship? Journal of Fish Biology, 82(5): 1484–1496, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wuellner:2013:RCS

[1103] M. R. Wuellner, R. G. Bramblett, C. S. Guy, A. V. Zale, D. R. Roberts, and J. Johnson. Reach and catchment-scale characteristics are relatively uninfluential in explaining the occurrence of stream fish species. Journal of Fish Biology, 82(5):1497–1513, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meli:2013:KAB

[1104] A. Meli and D. J. Fraser. Kinship analysis of brook trout Salvelinus fontinalis during their breeding migration. Journal of Fish Biology, 82 (5):1514–1522, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farley:2013:AEV

[1105] J. H. Farley, A. J. Williams, N. P. Clear, C. R. Davies, and S. J. Nicol. Age estimation and validation for South Pacific albacore Thunnus alalunga. Journal of Fish Biology, 82(5):1523–1544, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zeng:2013:OSM

[1106] W. W. Zeng, Q. Wang, Y. Y. Wang, D. H. Xu, and S. Q. Wu. A one-step molecular biology method for simple and rapid detection of grass carp Ctenopharyngodon idella reovirus (GCRV) HZ08 strain. Journal of Fish Biology, 82(5):1545–1555, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2013:STV

[1107] J. Martin, G. Bareille, S. Berail, C. Pecheyran, F. Daverat, N. Bru, H. Tabouret, and O. Donard. Spatial and temporal variations in otolith chemistry and relationships with water chemistry: a useful tool to distinguish Atlantic salmon Salmo salar parr from different natal streams. Journal of Fish Biology, 82(5):1556–1581, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akhilesh:2013:BOB

[1108] K. V. Akhilesh, W. T. White, K. K. Bineesh, U. Ganga, and N. G. K. Pillai. Biological observations on the bristly catshark Bythaelurus hispidus from deep waters off the south-west coast of India. Journal of Fish Biology, 82(5):1582–1591, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebner:2013:URU

[1109] B. C. Ebner and D. L. Morgan. Using remote underwater video to estimate freshwater fish species richness. Journal of Fish Biology, 82(5): 1592–1612, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2013:MME

[1110] G. S. C. Silva, B. F. Melo, C. Oliveira, and R. C. Benine. Morphological and molecular evidence for two new species of Tetragonopterus (Characiformes: Characidae) from central Brazil. Journal of Fish Biology, 82(5):1613–1631, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Izzo:2013:ABS

[1111] C. Izzo and E. B. Drew. Analysis of body shape for differentiating among species of rajids. Journal of Fish Biology, 82(5):1632–1640, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DAnatro:2013:CPM

[1112] A. D’Anatro, D. E. Naya, E. P. Lessa, and O. Defeo. Contrasting patterns of morphological variation with dietary preferences in Micropogonias furnieri: insights from stable-isotope and digestive-trait analyses. Journal of Fish Biology, 82(5):1641–1658, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franssen:2013:UER

[1113] N. R. Franssen and M. Tobler. Upstream effects of a reservoir on fish assemblages 45 years following impoundment. Journal of Fish Biology, 82(5):1659–1670, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsu:2013:GCV

[1114] C. C. Hsu, W. Y. Pai, C. Y. Lai, M. W. Lu, and G. M. Her. Genetic characterization and in vivo image analysis of novel zebrafish Danio rerio pigment mutants. Journal of Fish Biology, 82(5):1671–1683, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2013:EMV

[1115] M. Liu, J.-L. Li, S.-X. Ding, and Z.-Q. Liu. Epinephelus moara: a valid species of the family Epinephelidae (Pisces: Perciformes). Journal of Fish Biology, 82(5):1684–1699, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Toobaie:2013:DAP

[1116] A. Toobaie, J.-W. Kim, I. J. Dolinsek, and J. W. A. Grant. Diel activity patterns of the fish community in a temperate stream. Journal of Fish Biology, 82(5):1700–1707, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hume:2013:ERD

[1117] J. B. Hume, C. E. Adams, C. W. Bean, and P. S. Maitland. Evidence of a recent decline in river lamprey Lampetra fluviatilis parasitism of a nationally rare whitefish Coregonus lavaretus: is there a diamond in the ruffe Gymnocephalus cernuus? Journal of Fish Biology, 82(5):1708– 1716, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraabol:2013:OTS

[1118] M. Kraabøl, G. Fjeld, S. I. Johnsen, J. G. Dokk, B. K. Dervo, and J. Skurdal. The old trout Salmo trutta in the pond. Journal of Fish Biology, 82(5):1717–1719, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sparkes:2013:RSN

[1119] T. C. Sparkes, V. Rush, D. A. Kopp, and S. A. Foster. Reproductive success in a natural population of male three-spined stickleback Gasterosteus aculeatus: effects of nuptial colour, parasites and body size.

Journal of Fish Biology, 82(5):1720–1727, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gentil:2013:PBF

[1120] J. Le Gentil, S. Launey, F. Marchand, D. Ombredane, and J.-L. Bagliniere. Prediction of biological form at the adult stage of brown trout Salmo trutta using morphological or colorimetric criteria in migrating juveniles. Journal of Fish Biology, 82(5):1728–1732, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allard:2013:ITE

[1121] J. B. Allard, H. Kamei, and C. Duan. Inducible transgenic expression in the short-lived fish Nothobranchius furzeri. Journal of Fish Biology, 82(5):1733–1738, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2013:HFN

[1122] S. Silva, M. J. Servia, R. Vieira-Lanero, D. J. Nachón, and F. Cobo. Haematophagous feeding of newly metamorphosed European sea lampreys Petromyzon marinus on strictly freshwater species. Journal of Fish Biology, 82(5):1739–1745, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuset:2013:AOM

[1123] V. M. Tuset, V. Parisi-Baradad, and A. Lombarte. Application of otolith mass and shape for discriminating scabbardfishes Aphanopus spp. in the north-eastern Atlantic Ocean. Journal of Fish Biology, 82(5):1746–1752, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz-Osorio:2013:FRB

[1124] L. A. Muñoz-Osorio, P. A. Mejía-Falla, and A. F. Navia. First record of a bicephalic embryo of smalltail shark Carcharhinus porosus. Journal of Fish Biology, 82(5):1753–1757, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavraro:2013:SDV

[1125] F. Cavraro, M. Zucchetta, P. Torricelli, and S. Malavasi. Sexual dimorphism of vertical bar patterning in the South European toothcarp Aphanius fasciatus. Journal of Fish Biology, 82(5):1758–1764, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2013:BRE

[1126] A. (Tony) P. Farrell. Book review: Ecological and Environmental Physiology of Fishes — Edited by F. B. Eddy and R. D. Handy. Journal of Fish Biology, 82(5):1765–1766, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morley:2013:BRF

[1127] Neil Morley. Book review: Fish Parasites: Pathobiology and Protection — Edited by P. T. K. Woo and K. Buchmann. Journal of Fish Biology, 82(5):1767–1768, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2013:Ca

[1128] Anonymous. Corrigendum. Journal of Fish Biology, 82(5):1769, May 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2013:EWE

[1129] N. G. Andersen and J. F. Craig. Editorial on writing equations and mathematical expressions: some basic principles. Journal of Fish Biology, 82(6):1771–1772, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Denil:2013:RFT

[1130] N. A. Denil, E. Yamashita, M. Kirino, and S. Kiyohara. Recurrent facial taste neurons of sea catfish Plotosus japonicus: morphology and organization in the ganglion. Journal of Fish Biology, 82(6):1773–1788, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keenan:2013:BTM

[1131] K. Keenan, C. R. Bradley, J. J. Magee, R. A. Hynes, R. J. Kennedy, W. W. Crozier, R. Poole, T. F. Cross, P. McGinnity, and P. A. Prodöhl. Beaufort trout MicroPlex: a high-throughput multiplex platform comprising 38 informative microsatellite loci for use in resident and anadromous (sea trout) brown trout Salmo trutta genetic studies. Journal of Fish Biology, 82(6):1789–1804, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OShea:2013:DPF

[1132] O. R. O’Shea, M. Thums, M. van Keulen, R. M. Kempster, and M. G. Meekan. Dietary partitioning by five sympatric species of stingray (Dasyatidae) on coral reefs. Journal of Fish Biology, 82(6):1805–1820, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinault:2013:MFC

[1133] M. Pinault, N. Loiseau, P. Chabanet, P. Durville, H. Magalon, J. P. Quod, and R. Galzin. Marine fish communities in shallow volcanic habitats. Journal of Fish Biology, 82(6):1821–1847, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2013:TTE

[1134] J. S. F. Lee, E. P. Tezak, and B. A. Berejikian. Telemetry tag effects on juvenile lingcod Ophiodon elongatus movement: a laboratory and field study. Journal of Fish Biology, 82(6):1848–1857, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muncaster:2013:NSC

[1135] S. Muncaster, B. Norberg, and E. Andersson. Natural sex change in the temperate protogynous Ballan wrasse Labrus bergylta. Journal of Fish Biology, 82(6):1858–1870, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magath:2013:MBT

[1136] V. Magath, L. Marohn, J. Fietzke, M. Frische, R. Thiel, and J. Dierking. Migration behaviour of twaite shad Alosa fallax assessed by otolith Sr:Ca and Ba:Ca profiles. Journal of Fish Biology, 82(6):1871–1887, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2013:CNC

[1137] S.-L. Lin, Y.-H. Cheng, C.-M. Wen, and S.-N. Chen. Characterization of a novel cell line from the caudal fin of koi carp Cyprinus carpio. Journal of Fish Biology, 82(6):1888–1903, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burress:2013:SSO

[1138] E. D. Burress, A. Duarte, W. S. Serra, M. M. Gangloff, and L. Siefferman. Species-specific ontogenetic diet shifts among Neotropical crenicichla: using stable isotopes and tissue stoichiometry. Journal of Fish Biology, 82(6):1904–1915, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tweedley:2013:SCE

[1139] J. R. Tweedley, D. J. Bird, I. C. Potter, H. S. Gill, P. J. Miller, G. O’Donovan, and A. H. Tjakrawidjaja. Species compositions and ecology of the riverine ichthyofaunas in two Sulawesian islands in the biodiversity hotspot of Wallacea. Journal of Fish Biology, 82(6):1916–1950, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrade:2013:LHT

[1140] H. Andrade, J. Santos, and R. Taylor. Life-history traits of the common snook Centropomus undecimalis in a Caribbean estuary and large-scale biogeographic patterns relevant to management. Journal of Fish Biology, 82(6):1951–1974, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dijkstra:2013:MDB

[1141] P. D. Dijkstra, O. Seehausen, and N. B. Metcalfe. Metabolic divergence between sibling species of cichlids Pundamilia nyererei and Pundamilia pundamilia. Journal of Fish Biology, 82(6):1975–1989, June 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lima:2013:EDM

[1142] A. R. A. Lima, M. Barletta, D. V. Dantas, J. A. A. Ramos, and M. F. Costa. Early development of marine catfishes (Ariidae): from mouth brooding to the release of juveniles in nursery habitats. Journal of Fish Biology, 82(6):1990–2014, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kocovsky:2013:GMD

[1143] P. M. Kocovsky, T. J. Sullivan, C. T. Knight, and C. A. Stepien. Genetic and morphometric differences demonstrate fine-scale population substructure of the yellow perch Perca flavescens: need for redefined management units. Journal of Fish Biology, 82(6):2015–2030, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brandl:2013:PFH

[1144] S. J. Brandl and D. R. Bellwood. Pair formation in the herbivorous rabbitfish Siganus doliatus. Journal of Fish Biology, 82(6):2031–2044, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mwale:2013:IUS

[1145] M. Mwale, H. Kaiser, N. P. Barker, A. B. Wilson, and P. R. Teske. Identification of a uniquely southern African clade of coastal pipefishes Syngnathus spp. Journal of Fish Biology, 82(6):2045–2062, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brandner:2013:EST

[1146] J. Brandner, J. Pander, M. Mueller, A. F. Cerwenka, and J. Geist. Effects of sampling techniques on population assessment of invasive round goby Neogobius melanostomus. Journal of Fish Biology, 82(6):2063–2079, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ensing:2013:AGS

[1147] D. Ensing, W. W. Crozier, P. Boylan, N. O’Maoiléidigh, and P. McGinnity. An analysis of genetic stock identification on a small geographical scale using microsatellite markers, and its application in the management of a mixed-stock fishery for Atlantic salmon Salmo salar in Ireland. Journal of Fish Biology, 82(6):2080–2094, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindstedt:2013:VSI

[1148] E. R. Lindstedt and I. M. Hamilton. Variation in social information use: the influences of information reliability and mass on decision making in a group-living fish Gambusia affinis. Journal of Fish Biology, 82(6):2095– 2103, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliason:2013:LCA

[1149] E. J. Eliason, S. M. Wilson, A. P. Farrell, S. J. Cooke, and S. G. Hinch. Low cardiac and aerobic scope in a coastal population of sockeye salmon Oncorhynchus nerka with a short upriver migration. Journal of Fish Biology, 82(6):2104–2112, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2013:SFT

[1150] T. Simon, R. M. Macieira, and J.-C. Joyeux. The shore fishes of the Trindade–Martin Vaz insular complex: an update. Journal of Fish Biology, 82(6):2113–2127, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindner:2013:FEI

[1151] K. Lindner, A. F. Cerwenka, J. Brandner, S. Gertzen, J. Borcherding, J. Geist, and U. K. Schliewen. First evidence for interspecific hybridization between invasive goby species Neogobius fluviatilis and Neogobius melanostomus (Teleostei: Gobiidae: Benthophilinae). Journal of Fish Biology, 82(6):2128–2134, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lecchini:2013:VHD

[1152] D. Lecchini, V. P. Waqalevu, S. Holles, M. Lerohellec, C. Brie, and S. D. Simpson. Vertical and horizontal distributions of coral-reef fish larvae in open water immediately prior to reef colonization. Journal of Fish Biology, 82(6):2135–2140, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Theodorou:2013:HCA

[1153] P. Theodorou, S. S. Snorrason, and G. Á. Ólafsdóttir. Habitat complexity affects how young of the year Atlantic cod Gadus morhua perceive predation threat from older conspecifics. Journal of Fish Biology, 82 (6):2141–2146, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spitz:2013:VED

[1154] J. Spitz and J. Jouma’a. Variability in energy density of forage fishes from the Bay of Biscay (north-east Atlantic Ocean): reliability of functional grouping based on prey quality. Journal of Fish Biology, 82(6):2147– 2152, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2013:Cb

[1155] Anonymous. Corrigendum. Journal of Fish Biology, 82(6):2153, June 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uji:2013:MDB

[1156] S. Uji, T. Seikai, T. Suzuki, and K. Okuzawa. Muscle development in the bamboo sole Heteromycteris japonicus with special reference to larval branchial levators. Journal of Fish Biology, 83(1):1–13, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ponton:2013:GMT

[1157] D. Ponton, L. Carassou, S. Raillard, and P. Borsa. Geometric morphometrics as a tool for identifying emperor fish (Lethrinidae) larvae and juveniles. Journal of Fish Biology, 83(1):14–27, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benfield:2013:FSO

[1158] M. C. Benfield, S. Cook, S. Sharuga, and M. M. Valentine. Five in situ observations of live oarfish Regalecus glesne (Regalecidae) by remotely operated vehicles in the oceanic waters of the northern Gulf of Mexico.

Journal of Fish Biology, 83(1):28–38, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tofoli:2013:DFS

[1159] R. M. Tófoli, G. H. Z. Alves, J. Higuti, A. M. Cunico, and N. S. Hahn. Diet and feeding selectivity of a benthivorous fish in streams: responses to the effects of urbanization. Journal of Fish Biology, 83(1):39–51, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vergara:2013:CMP

[1160] A. R. Vergara, T. Sigurdsson, and F. Saborido-Rey. Comparative morphology of pre-extrusion larvae, Sebastes mentella and Sebastes norvegicus (Pisces: Sebastidae) in Icelandic waters. Journal of Fish Biology, 83 (1):52–63, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Iwatsuki:2013:RAL

[1161] Y. Iwatsuki. Review of the Acanthopagrus latus complex (Perciformes: Sparidae) with descriptions of three new species from the Indo-West Pacific Ocean. Journal of Fish Biology, 83(1):64–95, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Starrs:2013:CBM

[1162] D. Starrs, B. C. Ebner, and C. J. Fulton. Can backcalculation models unravel complex larval growth histories in a tropical freshwater fish? Journal of Fish Biology, 83(1):96–110, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drevecky:2013:GDS

[1163] C. J. Drevecky, R. Falco, and W. E. Aguirre. Genetic divergence of a sympatric lake-resident–anadromous three-spined stickleback Gasterosteus aculeatus species pair. Journal of Fish Biology, 83(1):111–132, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bustamante:2013:CME

[1164] C. Bustamante, P. M. Kyne, and M. B. Bennett. Comparative morphology of the egg cases of Asymbolus analis, Asymbolus rubiginosus and Figaro boardmani (Carcharhiniformes: Scyliorhinidae) from southern Queensland, Australia. Journal of Fish Biology, 83(1):133–143, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Regus:2013:CLH

[1165] J. U. Regus, J. B. Johnson, S. A. Webb, and D. N. Reznick. Comparative life histories of fishes in the genus Phallichthys (Pisces: Poeciliidae). Journal of Fish Biology, 83(1):144–155, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbini:2013:EIE

[1166] S. A. Barbini, L. B. Scenna, D. E. Figueroa, and J. M. Díaz de Astarloa. Effects of intrinsic and extrinsic factors on the diet of Bathyraja macloviana, a benthophagous skate. Journal of Fish Biology, 83(1):156– 169, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nichol:2013:DVM

[1167] D. G. Nichol, S. Kotwicki, and M. Zimmermann. Diel vertical migration of adult Pacific cod Gadus macrocephalus in Alaska. Journal of Fish Biology, 83(1):170–189, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elshout:2013:LOT

[1168] P. M. F. Elshout, L. M. Dionisio Pires, R. S. E. W. Leuven, S. E. Wendelaar Bonga, and A. J. Hendriks. Low oxygen tolerance of different life stages of temperate freshwater fish species. Journal of Fish Biology, 83(1):190–206, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wong:2013:CBT

[1169] M. Y. L. Wong, A. Medina, C. Uppaluri, S. Arnold, J. R. Seymour, and P. M. Buston. Consistent behavioural traits and behavioural syndromes in pairs of the false clown anemonefish Amphiprion ocellaris. Journal of Fish Biology, 83(1):207–213, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersson:2013:RBE

[1170] M. Å. Andersson, D. C. Laursen, P. I. M. Silva, and E. Höglund. The relationship between emergence from spawning gravel and growth in farmed rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 83(1): 214–219, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whittington:2013:BCR

[1171] C. M. Whittington, K. Musolf, S. Sommer, and A. B. Wilson. Behavioural cues of reproductive status in seahorses Hippocampus abdominalis. Journal of Fish Biology, 83(1):220–226, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2013:BRC

[1172] Laith A. Jawad. Book review: Coastal Fishes of New Zealand — Fourth Edition. By M. Francis. Journal of Fish Biology, 83(1):227–228, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lancaster:2013:BRF

[1173] John E. Lancaster. Book review: Fish. By E. R. DeSombre and J. S. Barkin. Journal of Fish Biology, 83(1):229–230, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2013:BRS

[1174] Charles G. M. Paxton. Book review: Systematics, Biology, and Distribution of the Species of the Oceanic Oarfish Genus Regalecus (Teleostei, Lampridiformes, Regalecidae) by T. R. Roberts. Journal of Fish Biology, 83(1):231–232, July 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2013:SPD

[1175] A. P. Scott, T. Hill, N. Beresford, T. J. Runnalls, J. P. Sumpter, and T. Ellis. Out-of-season production of 17,20β-dihydroxypregn-4-en-3-one in the roach Rutilus rutilus. Journal of Fish Biology, 83(2):233–249, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2013:RBF

[1176] C. L. Sun, S. Z. Yeh, Y. J. Chang, H. Y. Chang, and S. L. Chu. Reproductive biology of female bigeye tuna Thunnus obesus in the western Pacific Ocean. Journal of Fish Biology, 83(2):250–271, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Afonso:2013:NRC

[1177] P. Afonso, F. M. Porteiro, J. Fontes, F. Tempera, T. Morato, F. Cardigos, and R. S. Santos. New and rare coastal fishes in the Azores Islands: occasional events or tropicalization process? Journal of Fish Biology, 83 (2):272–294, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cheung:2013:EPO

[1178] N. K. M. Cheung, A. C. K. Cheung, R. R. Ye, W. Ge, J. P. Giesy, and D. W. T. Au. Expression profile of oestrogen receptors and oestrogenrelated receptors is organ specific and sex dependent: the Japanese medaka Oryzias latipes model. Journal of Fish Biology, 83(2):295–310, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irving:2013:ELB

[1179] E. Irving and C. Brown. Examining the link between personality and laterality in a feral guppy Poecilia reticulata population. Journal of Fish Biology, 83(2):311–325, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanNoord:2013:FFE

[1180] J. E. Van Noord, E. A. Lewallen, and R. L. Pitman. Flyingfish feeding ecology in the eastern Pacific: prey partitioning within a speciose epipelagic community. Journal of Fish Biology, 83(2):326–342, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amcoff:2013:DFF

[1181] M. Amcoff and N. Kolm. Does female feeding motivation affect the response to a food-mimicking male ornament in the swordtail characin Corynopoma riisei? Journal of Fish Biology, 83(2):343–354, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reecht:2013:UMC

[1182] Y. Reecht, M.-J. Rochet, V. M. Trenkel, S. Jennings, and J. K. Pinnegar. Use of morphological characteristics to define functional groups of predatory fishes in the Celtic Sea. Journal of Fish Biology, 83(2):355– 377, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2013:DRA

[1183] D. Harasti and H. Malcolm. Distribution, relative abundance and size composition of the threatened serranid Epinephelus daemelii in New South Wales, Australia. Journal of Fish Biology, 83(2):378–395, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plambech:2013:EPH

[1184] M. Plambech, M. Van Deurs, J. F. Steffensen, B. Tirsgaard, and J. W. Behrens. Excess post-hypoxic oxygen consumption in Atlantic cod Gadus morhua. Journal of Fish Biology, 83(2):396–403, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pompini:2013:TIS

[1185] M. Pompini, A. M. Buser, M. R. Thali, B. A. Von Siebenthal, S. Nusslé, S. Guduff, and C. Wedekind. Temperature-induced sex reversal is not responsible for sex-ratio distortions in grayling Thymallus thymallus or brown trout Salmo trutta. Journal of Fish Biology, 83(2):404–411, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva:2013:ASB

[1186] M. E. Oliva. Is Anisakis simplex s.l. a biological marker for stock identification of Strangomera bentincki from Chile? Journal of Fish Biology, 83(2):412–416, August 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harborne:2013:EBP

[1187] A. R. Harborne. The ecology, behaviour and physiology of fishes on coral reef flats, and the potential impacts of climate change. Journal of Fish Biology, 83(3):417–447, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunha:2013:LJD

[1188] M. E. Cunha, P. Ré, H. Quental-Ferreira, P. J. Gavaia, and P. Pousão-Ferreira. Larval and juvenile development of dusky grouper Epinephelus marginatus reared in mesocosms. Journal of Fish Biology, 83(3):448–465, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leis:2013:PDT

[1189] J. M. Leis, A. C. Hay, P. Sasal, A. S. Hicks, and R. Galzin. Pelagic to demersal transition in a coral-reef fish, the orbicular batfish Platax orbicularis. Journal of Fish Biology, 83(3):466–479, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcalo:2013:BRS

[1190] A. Marçalo, J. Araújo, P. Pousão-Ferreira, G. J. Pierce, Y. Stratoudakis, and K. Erzini. Behavioural responses of sardines Sardina pilchardus to simulated purse-seine capture and slipping. Journal of Fish Biology, 83 (3):480–500, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Velez:2013:OTP

[1191] Z. Velez, P. C. Hubbard, E. N. Barata, and A. V. M. Canário. Olfactory transduction pathways in the Senegalese sole Solea senegalensis. Journal of Fish Biology, 83(3):501–514, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2013:UMB

[1192] R. J. Kennedy, I. Moffett, M. M. Allen, and S. M. Dawson. Upstream migratory behaviour of wild and ranched Atlantic salmon Salmo salar at a natural obstacle in a coastal spate river. Journal of Fish Biology, 83 (3):515–530, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alarcos:2013:SSM

[1193] A. J. Alarcos and J. T. Timi. Stocks and seasonal migrations of the flounder Xystreurys rasile as indicated by its parasites. Journal of Fish Biology, 83(3):531–541, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crichigno:2013:CMV

[1194] S. Crichigno, C. Conte-Grand, M. Battini, and V. Cussac. Cephalic morphological variation in freshwater silversides Odontesthes hatcheri and Odontesthes bonariensis in Patagonia: introgression and ecological relationships. Journal of Fish Biology, 83(3):542–559, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2013:ECV

[1195] J. Ma, S. Sun, L. Zeng, and Y. Lu. Establishment, characterization and viral susceptibility of two cell lines derived from leopard wrasse Macropharyngodon geoffroy. Journal of Fish Biology, 83(3):560–573, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fox:2013:PSR

[1196] S. Fox, I. Foisy, R. De La Parra Venegas, B. E. Galván Pastoriza, R. T. Graham, E. R. Hoffmayer, J. Holmberg, and S. J. Pierce. Population structure and residency of whale sharks Rhincodon typus at Utila, Bay Islands, Honduras. Journal of Fish Biology, 83(3):574–587, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoshikawa:2013:SDG

[1197] M. Yoshikawa. Sex differences in growth rates of early life stage Japanese eels Anguilla japonica under experimental conditions. Journal of Fish Biology, 83(3):588–597, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swatdipong:2013:GMS

[1198] A. Swatdipong, A. Vasemägi, T. Niva, M.-L. Koljonen, and C. R. Primmer. Genetic mixed-stock analysis of lake-run brown trout Salmo trutta fishery catches in the Inari Basin, northern Finland: implications for conservation and management. Journal of Fish Biology, 83(3):598–617, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boaden:2013:DHA

[1199] A. E. Boaden and M. J. Kingsford. Distributions and habitat associations of the bridled monocle bream Scolopsis bilineatus (Nemipteridae): a demographic approach. Journal of Fish Biology, 83(3):618–641, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lu:2013:SIR

[1200] D. Q. Lu, M. Yao, S. B. Yi, Y. W. Li, X. C. Liu, Y. Zhang, and H. R. Lin. Soluble interleukin-1 receptor, a potential negative regulator of orangespotted grouper Epinephelus coioides interleukin-1 system. Journal of Fish Biology, 83(3):642–658, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sekulic:2013:NDD

[1201] N. Sekulić, S. Marić, L. Galambos, D. Radošević, and J. Krpo-Ćetković. New distribution data and population structure of the European mudminnow Umbra krameri in Serbia and Bosnia and Herzegovina. Journal of Fish Biology, 83(3):659–666, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ho:2013:TAP

[1202] W. W. Ho, C. R. Turner, and G. T. Smith. Transition across polymorphic phenotypes observed in a male Sternarchogiton nattereri (Gymnotiformes: Apteronotidae). Journal of Fish Biology, 83(3):667–670, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

CardosodeCarvalho:2013:MIH

[1203] D. Cardoso de Carvalho, A. S. Seerig, B. Dos Santos Alves Figueiredo Brasil, D. Vieira Crepaldi, and D. Aparecida Andrade de Oliveria. Molecular identification of the hybrid between the catfish species Pseudoplatystoma corruscans and Pseudoplatystoma reticulatum using a set of eight microsatellite markers. Journal of Fish Biology, 83(3):671–676, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luzzatto:2013:RSP

[1204] D. C. Luzzatto, M. L. Estalles, and J. M. Díaz de Astarloa. Rafting seahorses: the presence of juvenile Hippocampus patagonicus in floating debris. Journal of Fish Biology, 83(3):677–681, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hurtado:2013:GDP

[1205] L. A. Hurtado, E. Carrera, A. Adite, and K. O. Winemiller. Genetic differentiation of a primitive teleost, the African bonytongue Heterotis niloticus, among river basins and within a floodplain river system in Benin, West Africa. Journal of Fish Biology, 83(3):682–690, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacDonald:2013:ROR

[1206] P. MacDonald. A rare occurrence of reversal in the common megrim Lepidorhombus whiffiagonis (Pleuronectiformes: Scophthalmidae) in the northern North Sea. Journal of Fish Biology, 83(3):691–694, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomita:2013:OCD

[1207] T. Tomita, T. Kawai, H. Matsubara, and R. Nagata. Occurrence of the Chilean devil ray Mobula tarapacana (Elasmobranchii: Batoidea: Myliobatiformes) in the Sea of Okhotsk: first record from cold temperate waters. Journal of Fish Biology, 83(3):695–698, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Selckmann:2013:PSA

[1208] G. M. Selckmann and J. Ramcharitar. Patterns of saccular afferent innervation in sciaenids. Journal of Fish Biology, 83(3):699–706, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2013:BRFa

[1209] Laith A. Jawad. Book review: Field Identification guide to the Living Resources of the Eastern and Southern Mediterranean Sea — By M. Bariche. Journal of Fish Biology, 83(3):707–708, September 2013. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2013:BRD

[1210] Culum Brown. Book review: Discovery of Australia’s Fishes: a History of Australian Ichthyology to 1930 — By B. Saunders. Journal of Fish Biology, 83(3):709–710, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lucas:2013:BRA

[1211] Martyn C. Lucas. Book review: Advances in Fish Tagging and Marking Technology — Edited by J. R. McKenzie, B. Parsons, A. C. Seitz, R. K. Kopf, M. Mesa and Q. Phelps. Journal of Fish Biology, 83(3):711–712, September 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2013:SPS

[1212] P. J. B. Hart, B. M. Gillanders, and J. F. Craig. Selected Papers from the Sixth World Fisheries Congress. Journal of Fish Biology, 83(4):713– 715, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2013:CPF

[1213] M. B. Santos, R. González-Quirós, I. Riveiro, M. Iglesias, M. Louzao, and G. J. Pierce. Characterization of the pelagic fish community of the northwestern and northern Spanish shelf waters. Journal of Fish Biology, 83 (4):716–738, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McIntyre:2013:VST

[1214] F. D. McIntyre, N. Collie, M. Stewart, L. Scala, and P. G. Fernandes. A visual survey technique for deep-water fishes: estimating anglerfish Lophius spp. abundance in closed areas. Journal of Fish Biology, 83 (4):739–753, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Righton:2013:ACG

[1215] D. Righton and A. M. Walker. Anguillids: conserving a global fishery. Journal of Fish Biology, 83(4):754–765, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gilman:2013:CME

[1216] E. Gilman, P. Suuronen, M. Hall, and S. Kennelly. Causes and methods to estimate cryptic sources of fishing mortality. Journal of Fish Biology, 83(4):766–803, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:2013:CDP

[1217] J. A. A. Perez, B. N. Pereira, D. A. Pereira, and R. Schroeder. Composition and diversity patterns of megafauna discards in the deep-water shrimp trawl fishery off Brazil. Journal of Fish Biology, 83(4):804–825, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eno:2013:ASH

[1218] N. C. Eno, C. L. J. Frid, K. Hall, K. Ramsay, R. A. M. Sharp, D. P. Brazier, S. Hearn, K. M. Dernie, K. A. Robinson, O. A. L. Paramor, and L. A. Robinson. Assessing the sensitivity of habitats to fishing: from seabed maps to sensitivity maps. Journal of Fish Biology, 83(4):826– 846, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pranovi:2013:VAF

[1219] F. Pranovi, A. Caccin, P. Franzoi, S. Malavasi, M. Zucchetta, and P. Torricelli. Vulnerability of artisanal fisheries to climate change in the Venice lagoon. Journal of Fish Biology, 83(4):847–864, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:2013:IPR

[1220] T. J. Pitcher, M. E. Lam, C. Ainsworth, A. Martindale, K. Nakamura, R. I. Perry, and T. Ward. Improvements to Rapfish: a rapid evaluation technique for fisheries integrating ecological and human dimensions. Journal of Fish Biology, 83(4):865–889, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houle:2013:EAI

[1221] J. E. Houle, K. H. Andersen, K. D. Farnsworth, and D. G. Reid. Emerging asymmetric interactions between forage and predator fisheries impose management trade-offs. Journal of Fish Biology, 83(4):890–904, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2013:CCB

[1222] A. Arkhipkin, J. Barton, S. Wallace, and A. Winter. Close cooperation between science, management and industry benefits sustainable exploitation of the Falkland Islands squid fisheries. Journal of Fish Biology, 83 (4):905–920, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ratz:2013:MTS

[1223] H.-J. Rätz, A. Charef, A. J. Abella, F. Colloca, A. Ligas, A. Mannini, and J. Lloret. A medium-term, stochastic forecast model to support sustainable, mixed fisheries management in the Mediterranean Sea. Journal of Fish Biology, 83(4):921–938, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Florin:2013:ELN

[1224] A.-B. Florin, U. Bergström, D. Ustups, K. Lundström, and P. R. Jonsson. Effects of a large northern European no-take zone on flatfish populations. Journal of Fish Biology, 83(4):939–962, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraan:2013:OPS

[1225] M. Kraan, S. Uhlmann, J. Steenbergen, A. T. M. Van Helmond, and L. Van Hoof. The optimal process of self-sampling in fisheries: lessons learned in The Netherlands. Journal of Fish Biology, 83(4):963–973, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2013:ISC

[1226] M. J. Armstrong, A. I. L. Payne, B. Deas, and T. L. Catchpole. Involving stakeholders in the commissioning and implementation of fishery science projects: experiences from the U.K. fisheries science partnership. Journal of Fish Biology, 83(4):974–996, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2013:FEP

[1227] S. J. Cooke, N. W. R. Lapointe, E. G. Martins, J. D. Thiem, G. D. Raby, M. K. Taylor, T. D. Beard, Jr., and I. G. Cowx. Failure to engage the public in issues related to inland fishes and fisheries: strategies for building public and political will to promote meaningful conservation. Journal of Fish Biology, 83(4):997–1018, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Field:2013:CBS

[1228] J. G. Field, C. G. Attwood, A. Jarre, K. Sink, L. J. Atkinson, and S. Petersen. Cooperation between scientists, NGOs and industry in support of sustainable fisheries: the South African hake Merluccius spp. trawl fishery experience. Journal of Fish Biology, 83(4):1019–1034, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Savolainen:2013:DEE

[1229] M. A. Savolainen, R. F. Kazmierczak, Jr., and R. H. Caffey. Determining the effect of environmental accidents on responses to a Gulf of Mexico recreational for-hire fishing industry survey. Journal of Fish Biology, 83 (4):1035–1045, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shepherd:2013:GFF

[1230] C. J. Shepherd and A. J. Jackson. Global fishmeal and fish-oil supply: inputs, outputs and markets. Journal of Fish Biology, 83(4):1046–1066, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beveridge:2013:MFN

[1231] M. C. M. Beveridge, S. H. Thilsted, M. J. Phillips, M. Metian, M. Troell, and S. J. Hall. Meeting the food and nutrition needs of the poor: the role of fish and the opportunities and challenges emerging from the rise of aquaculture. Journal of Fish Biology, 83(4):1067–1084, October 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2013:GJH

[1232] Ian Harrison, Malcolm Evans, Rosemary Lowe-McConnell, Lynne Parenti, Chris Sanford, Lex Snyder, Melanie Stiassny, and Rich Vari. Gordon John Howes 1938–2013 (fish systematist). Journal of Fish Biology, 83(5): 1085–1093, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hipfner:2013:STV

[1233] J. M. Hipfner and M. Galbraith. Spatial and temporal variation in the diet of the Pacific sand lance Ammodytes hexapterus in waters off the coast of British Columbia, Canada. Journal of Fish Biology, 83(5):1094– 1111, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akhilesh:2013:CCR

[1234] K. V. Akhilesh, K. K. Bineesh, W. T. White, C. P. R. Shanis, M. Hashim, U. Ganga, and N. G. K. Pillai. Catch composition, reproductive biology and diet of the bramble shark Echinorhinus brucus (Squaliformes: Echinorhinidae) from the south-eastern Arabian Sea. Journal of Fish Biology, 83(5):1112–1127, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birindelli:2013:TNS

[1235] J. L. O. Birindelli and H. A. Britski. Two new species of Leporinus (Characiformes: Anostomidae) from the Brazilian Amazon, and redescription of Leporinus striatus Kner 1858. Journal of Fish Biology, 83(5):1128–1160, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goetz:2013:DMI

[1236] F. A. Goetz, B. Baker, T. Buehrens, and T. P. Quinn. Diversity of movements by individual anadromous coastal cutthroat trout Oncorhynchus clarkii clarkii. Journal of Fish Biology, 83(5):1161–1182, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duan:2013:RTB

[1237] M. Duan, T. Zhang, W. Hu, S. Xie, L. F. Sundström, Z. Li, and Z. Zhu. Risk-taking behaviour may explain high predation mortality of GHtransgenic common carp Cyprinus carpio. Journal of Fish Biology, 83(5): 1183–1196, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Handeland:2013:LTE

[1238] S. O. Handeland, A. K. Imsland, B. Th. Björnsson, and S. O. Stefansson. Long-term effects of photoperiod, temperature and their interaction on growth, gill Na^{+}, K^{+}-ATPase activity, seawater tolerance and plasma growth-hormone levels in Atlantic salmon Salmo salar. Journal of Fish Biology, 83(5):1197–1209, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2013:GMF

[1239] A. Arkhipkin and V. Laptikhovsky. From gelatinous to muscle food chain: rock cod Patagonotothen ramsayi recycles coelenterate and tunicate resources on the Patagonian shelf. Journal of Fish Biology, 83(5): 1210–1220, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2013:AME

[1240] D. S. Murray, M. M. Bain, and C. E. Adams. Adhesion mechanisms in European whitefish Coregonus lavaretus eggs: is this a survival mechanism for high-energy spawning grounds? Journal of Fish Biology, 83(5): 1221–1233, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shen:2013:MCI

[1241] X. B. Shen, D. Xu, J. L. Li, and L. Q. Lu. Molecular cloning and immune responsive expression of a ribonuclease III orthologue involved in RNA interference, dicer, in grass carp Ctenopharyngodon idella. Journal of Fish Biology, 83(5):1234–1248, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meneghesso:2013:LHT

[1242] C. Meneghesso, E. Riginella, M. La Mesa, F. Donato, and C. Mazzoldi. Life-history traits and population decline of the Atlantic mackerel Scomber scombrus in the Adriatic Sea. Journal of Fish Biology, 83(5): 1249–1267, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deMestral:2013:RAR

[1243] L. G. de Mestral and C. M. Herbinger. Reduction in antipredator response detected between first and second generations of endangered juvenile Atlantic salmon Salmo salar in a captive breeding and rearing programme. Journal of Fish Biology, 83(5):1268–1286, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2013:CTC

[1244] Z.-X. Zhang, D.-Y. Geng, Q. Han, S.-D. Liang, and H.-R. Guo. The Cterminal cysteine of turbot Scophthalmus maximus translationally controlled tumour protein plays a key role in antioxidation and growthpromoting functions. Journal of Fish Biology, 83(5):1287–1301, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boglino:2013:HDA

[1245] A. Boglino, A. Wishkerman, M. J. Darias, K. B. Andree, P. de la Iglesia, A. Estévez, and E. Gisbert. High dietary arachidonic acid levels affect the process of eye migration and head shape in pseudoalbino Senegalese sole Solea senegalensis early juveniles. Journal of Fish Biology, 83(5):

1302–1320, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barriga:2013:VVN

[1246] J. P. Barriga, D. Milano, and V. E. Cussac. Variation in vertebral number and its morphological implication in Galaxias platei. Journal of Fish Biology, 83(5):1321–1333, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2013:SDS

[1247] T. Tomiyama. Sexual dimorphism in scales of marbled flounder Pseudopleuronectes yokohamae (Pleuronectiformes: Pleuronectidae), with comments on the relevance to their spawning behaviour. Journal of Fish Biology, 83(5):1334–1343, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2013:DUF

[1248] D. Harasti and W. Gladstone. Does underwater flash photography affect the behaviour, movement and site persistence of seahorses? Journal of Fish Biology, 83(5):1344–1353, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temperoni:2013:FSJ

[1249] B. Temperoni, M. D. Viñas, and C. C. Buratti. Feeding strategy of juvenile (age-0+ year) Argentine hake Merluccius hubbsi in the Patagonian nursery ground. Journal of Fish Biology, 83(5):1354–1370, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2013:IBI

[1250] M. D. Martin and T. C. Mendelson. Incomplete behavioural isolation and asymmetric female preference in darter sister species (Percidae: Etheostoma). Journal of Fish Biology, 83(5):1371–1380, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leris:2013:ZDR

[1251] I. Leris, D. G. Sfakianakis, and M. Kentouri. Are zebrafish Danio rerio males better swimmers than females? Journal of Fish Biology, 83(5): 1381–1386, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchio:2013:SSL

[1252] E. A. Marchio and K. R. Piller. Scale-stacking in the live-bearer Belonesox belizanus (Poeciliidae). Journal of Fish Biology, 83(5):1387–1390, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsche:2013:PEI

[1253] M. A. Matsche. A portable electro-immobilization and laparoscopy system for sex determination and gonadal biopsy in largemouth bass Micropterus salmoides. Journal of Fish Biology, 83(5):1391–1400, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Currey:2013:BLL

[1254] L. M. Currey, M. R. Heupel, C. A. Simpfendorfer, and T. D. Clark. Blood lactate loads of redthroat emperor Lethrinus miniatus associated with angling stress and exhaustive exercise. Journal of Fish Biology, 83(5):1401–1406, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulgar:2013:CUS

[1255] J. Pulgar, E. Poblete, M. Alvarez, J. P. Morales, B. Aranda, M. Aldana, and V. M. Pulgar. Can upwelling signals be detected in intertidal fishes of different trophic levels? Journal of Fish Biology, 83(5):1407–1415, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sopinka:2013:DAP

[1256] N. M. Sopinka, S. G. Hinch, A. G. Lotto, C. K. Whitney, and D. A. Patterson. Does among-population variation in burst swimming performance of sockeye salmon Oncorhynchus nerka fry reflect early life migrations? Journal of Fish Biology, 83(5):1416–1424, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

El-Regal:2013:FRS

[1257] M. A. Abu El-Regal and K. El-Moselhy. The first record of the slender sunfish Ranzania laevis from the Red Sea. Journal of Fish Biology, 83(5): 1425–1429, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:2013:DCM

[1258] D. G. Ferreira, B. A. Galindo, A. N. Alves, F. S. Almeida, C. F. Ruas, and S. H. Sofia. Development and characterization of 14 microsatellite loci in the Neotropical fish Geophagus brasiliensis (Perciformes, Cichlidae). Journal of Fish Biology, 83(5):1430–1438, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macri:2013:UMR

[1259] F. Macrì, L. Liotta, R. Bonfiglio, C. De Stefano, D. Ruscica, and G. Aiudi. Ultrasound measurement of reproductive organs in juvenile European sea bass Dicentrarchus labrax. Journal of Fish Biology, 83(5): 1439–1443, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Juza:2013:OAY

[1260] T. Jůza, T. Mrkvička, P. Blabolil, M. Čech, J. Peterka, M. Vašek, and J. Kubečka. Occurrence of age-0 year dwarf pikeperch Sander lucioperca in late summer — an overlooked phenomenon in reservoirs. Journal of Fish Biology, 83(5):1444–1452, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2013:IIC

[1261] F. A. Huntingford, F. L. Borçato, and F. O. Mesquita. Identifying individual common carp Cyprinus carpio using scale pattern. Journal of Fish Biology, 83(5):1453–1458, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nowak:2013:LCD

[1262] B. Nowak, K. Cadoret, S. W. Feist, and T. P. Bean. Laser-capture dissection and immunohistochemistry reveals chloride and mucous-cell specific gene expression in gills of seawater acclimated Atlantic salmon Salmo salar. Journal of Fish Biology, 83(5):1459–1467, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osorio:2013:CAC

[1263] R. Osorio and O. Schmachtenberg. Calcium-activated chloride channels do not contribute to the odorant transduction current in the marine teleost Isacia conceptionis. Journal of Fish Biology, 83(5):1468–1473, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2013:BRB

[1264] Ian J. Winfield. Book review: Britain’s Freshwater Fishes — By M. Everard. Journal of Fish Biology, 83(5):1474–1475, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartl:2013:BRH

[1265] Mark G. J. Hartl. Book review: Homeostasis and Toxicology of Non-Essential Metals — Edited by C. M. Wood, A. P. Farrell & C. J. Brauner.

Journal of Fish Biology, 83(5):1476–1477, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2013:BRF

[1266] Ian J. Winfield. Book review: Fisheries Techniques — Edited by A. V. Zale, D. L. Parrish & T. M. Sutton. Journal of Fish Biology, 83(5):1478, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2013:BRE

[1267] Culum Brown. Book review: Ecology of Australian Freshwater Fishes — Edited by P. Humphries and K. Walker. Journal of Fish Biology, 83(5): 1479–1482, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2013:BRR

[1268] Laith A. Jawad. Book review: Reef Fishes of the East Indies — By G. R. Allen and M. V. Erdmann. Journal of Fish Biology, 83(5):1483–1484, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2013:BRFb

[1269] Laith A. Jawad. Book review: Fish Physiology: Euryhaline Fishes — Edited by S. D. McCormick, A. P. Farrell & C. J. Brauner. Journal of Fish Biology, 83(5):1485–1486, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2013:BRS

[1270] Laith A. Jawad. Book review: Sharks of the World: a Fully Illustrated Guide — Edited by D. A. Ebert, S. Fowler & L. Compagno. Journal of Fish Biology, 83(5):1487–1488, November 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergstad:2013:NAD

[1271] O. A. Bergstad. North Atlantic demersal deep-water fish distribution and biology: present knowledge and challenges for the future. Journal of Fish Biology, 83(6):1489–1507, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sutton:2013:VEP

[1272] T. T. Sutton. Vertical ecology of the pelagic ocean: classical patterns and new perspectives. Journal of Fish Biology, 83(6):1508–1527, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:2013:CDS

[1273] I. G. Priede and R. Froese. Colonization of the deep sea by fishes. Journal of Fish Biology, 83(6):1528–1550, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mytilineou:2013:SDL

[1274] Ch. Mytilineou, K. Tsagarakis, P. Bekas, A. Anastasopoulou, S. Kavadas, A. Machias, J. Haralabous, C. J. Smith, G. Petrakis, J. Dokos, and A. Kapandagakis. Spatial distribution and life-history aspects of blackspot seabream Pagellus bogaraveo (Osteichthyes: Sparidae). Journal of Fish Biology, 83(6):1551–1575, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2013:PEF

[1275] F. C. Neat and N. Campbell. Proliferation of elongate fishes in the deep sea. Journal of Fish Biology, 83(6):1576–1591, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drazen:2013:RMP

[1276] J. C. Drazen, B. Dugan, and J. R. Friedman. Red muscle proportions and enzyme activities in deep-sea demersal fishes. Journal of Fish Biology, 83(6):1592–1612, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trueman:2013:TTM

[1277] C. N. Trueman, R. E. M. Rickaby, and S. Shephard. Thermal, trophic and metabolic life histories of inaccessible fishes revealed from stableisotope analyses: a case study using orange roughy Hoplostethus atlanticus. Journal of Fish Biology, 83(6):1613–1636, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anastasopoulou:2013:DFS

[1278] A. Anastasopoulou, Ch. Mytilineou, E. Lefkaditou, J. Dokos, C. J. Smith, A. Siapatis, P. Bekas, and K.-N. Papadopoulou. Diet and feeding strategy of blackmouth catshark Galeus melastomus. Journal of Fish Biology, 83(6):1637–1655, December 2013. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Backstrom:2014:BPC

[1279] T. Backström, E. Brännäs, J. Nilsson, and C. Magnhagen. Behaviour, physiology and carotenoid pigmentation in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 84(1):1–9, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gil:2014:AWC

[1280] M. M. Gil, M. Palmer, A. Grau, S. Deudero, J. I. Alconchel, and I. A. Catalán. Adapting to the wild: the case of aquaculture-produced and released meagres Argyrosomus regius. Journal of Fish Biology, 84(1):10– 30, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnan:2014:MVF

[1281] P. Magnan, R. Proulx, P. Berrebi, J. Blondel, P. Perret, and B. Roché. Morphological variation in the freshwater blenny Salaria fluviatilis from Corsican rivers: adaptive divergence, phenotypic plasticity or both? Journal of Fish Biology, 84(1):31–44, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wolkers:2014:TCA

[1282] C. P. B. Wolkers, M. Serra, R. E. Szawka, and E. C. Urbinati. The time course of aggressive behaviour in juvenile matrinxã Brycon amazonicus fed with dietary L-tryptophan supplementation. Journal of Fish Biology, 84(1):45–57, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beakes:2014:USC

[1283] M. P. Beakes, S. Sharron, R. Charish, J. W. Moore, W. H. Satterthwaite, E. Sturm, B. K. Wells, S. M. Sogard, and M. Mangel. Using scale characteristics and water temperature to reconstruct growth rates of juvenile steelhead Oncorhynchus mykiss. Journal of Fish Biology, 84(1):58–72, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richard:2014:EFR

[1284] A. Richard, J. O’Rourke, and J.-F. Rubin. External fluorescence retention of calcein-marked juvenile brown trout Salmo trutta raised in natural and artificial environments. Journal of Fish Biology, 84(1):73– 84, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bian:2014:TMS

[1285] X. Bian, X. Zhang, Y. Sakurai, X. Jin, T. Gao, R. Wan, and J. Yamamoto. Temperature-mediated survival, development and hatching variation of Pacific cod Gadus macrocephalus eggs. Journal of Fish Biology, 84(1):85–105, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

French:2014:BHF

[1286] B. French, I. C. Potter, S. A. Hesp, P. G. Coulson, and N. G. Hall. Biology of the harlequin fish Othos dentex (Serranidae), with particular emphasis on sexual pattern and other reproductive characteristics. Journal of Fish Biology, 84(1):106–132, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brannas:2014:TPL

[1287] E. Brännäs. Time–place learning and leader–follower relationships in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 84(1):133–144, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2014:GAS

[1288] J.-S. Moore, T. N. Loewen, L. N. Harris, and R. F. Tallman. Genetic analysis of sympatric migratory ecotypes of Arctic charr Salvelinus alpinus: alternative mating tactics or reproductively isolated strategies? Journal of Fish Biology, 84(1):145–162, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2014:SAA

[1289] D. T. Roberts, S. Mallett, N. C. Krück, W. Loh, and I. Tibbetts. Spawning activity of the Australian lungfish Neoceratodus forsteri in an impoundment. Journal of Fish Biology, 84(1):163–177, January 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urke:2014:DST

[1290] H. A. Urke, J. V. Arnekleiv, T. O. Nilsen, and K. J. Nilssen. Development of seawater tolerance and subsequent downstream migration in wild and stocked young-of-the-year derived Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 84(1):178–192, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wen:2014:CMG

[1291] A. Y. Wen, F. You, P. Sun, J. Li, D. D. Xu, Z. H. Wu, D. Y. Ma, and P. J. Zhang. CpG methylation of dmrt1 and cyp19a promoters in relation to their sexual dimorphic expression in the Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 84(1):193–205, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nikolakakis:2014:PQS

[1292] S. Nikolakakis, P. Bossier, G. Kanlis, K. Dierckens, and D. Adriaens. Protocol for quantitative shape analysis of deformities in early larval European seabass Dicentrarchus labrax. Journal of Fish Biology, 84(1): 206–224, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haynes:2014:TEL

[1293] P. S. Haynes, D. Brophy, and D. Mcgrath. The timing of early life events and growth rate estimates of age-0 year group brill Scophthalmus rhombus along the west coast of Ireland. Journal of Fish Biology, 84 (1):225–230, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Booth:2014:GTR

[1294] D. J. Booth, D. E. Poulos, J. Poole, and D. A. Feary. Growth and temperature relationships for juvenile fish species in seagrass beds: implications of climate change. Journal of Fish Biology, 84(1):231–236, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuparinen:2014:MHB

[1295] A. Kuparinen, M. Vinni, A. G. F. Teacher, K. Kähkönen, and J. Merilä. Mechanism of hybridization between bream Abramis brama and roach Rutilus rutilus in their native range. Journal of Fish Biology, 84(1):237– 242, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomita:2014:NRW

[1296] T. Tomita, T. Kawai, H. Matsubara, M. Kobayashi, and S. Katakura. Northernmost record of a whale shark Rhincodon typus from the Sea of Okhotsk. Journal of Fish Biology, 84(1):243–246, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pradella:2014:FAA

[1297] N. Pradella, A. M. Fowler, D. J. Booth, and P. I. Macreadie. Fish assemblages associated with oil industry structures on the continental shelf of north-western Australia. Journal of Fish Biology, 84(1):247–255, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2014:ROG

[1298] T. Yamamoto and S. Kitanishi. Reduced oceanic growth of growth hormone pseudogene-positive female masu salmon Oncorhynchus masou.

Journal of Fish Biology, 84(1):256–262, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OBrien:2014:FOA

[1299] G. C. O’Brien, F. Jacobs, S. W. Evans, and N. J. Smit. First observation of African tigerfish Hydrocynus vittatus predating on barn swallows Hirundo rustica in flight. Journal of Fish Biology, 84(1):263–266, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bremer:2014:SSN

[1300] J. R. Alvarado Bremer, B. L. Smith, D. L. Moulton, C.-P. Lu, and M. Cornic. Shake and stew: a non-destructive PCR-ready DNA isolation method from a single preserved fish larva. Journal of Fish Biology, 84(1):267–272, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2014:MPA

[1301] D. G. Roberts, C. A. Gray, and D. J. Ayre. Microsatellite primers for Australian recreationally and commercially important estuarine fishes. Journal of Fish Biology, 84(1):273–281, January 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tobler:2014:MVV

[1302] M. Tobler and N. Bertrand. Morphological variation in vanishing Mexican desert fishes of the genus Characodon (Goodeidae). Journal of Fish Biology, 84(2):283–296, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2014:MMF

[1303] A. C. Henderson, A. J. Reeve, and A. Ambu-Ali. Microanatomy of the male and female reproductive tracts in the long-tailed butterfly ray Gymnura poecilura, an elasmobranch with unusual characteristics. Journal of Fish Biology, 84(2):297–313, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Datovo:2014:NSI

[1304] A. Datovo and M. C. C. de Pinna. A new species of Ituglanis representing the southernmost record of the genus, with comments on phylogenetic relationships (Teleostei: Siluriformes: Trichomycteridae). Journal of Fish Biology, 84(2):314–327, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hobbs:2014:EET

[1305] J.-P. A. Hobbs, A. J. Frisch, S. Mutz, and B. M. Ford. Evaluating the effectiveness of teeth and dorsal fin spines for non-lethal age estimation of a tropical reef fish, coral trout Plectropomus leopardus. Journal of Fish Biology, 84(2):328–338, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gunnarsson:2014:AWF

[1306] Á. Gunnarsson. Atlantic wolf-fish Anarhichas lupus population diversity: growth and maturation. Journal of Fish Biology, 84(2):339–353, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsai:2014:TSS

[1307] C.-N. Tsai, W.-C. Chiang, C.-L. Sun, K.-T. Shao, S.-Y. Chen, and S.-Z. Yeh. Trophic size-structure of sailfish Istiophorus platypterus in eastern Taiwan estimated by stable isotope analysis. Journal of Fish Biology, 84 (2):354–371, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandez:2014:NSS

[1308] L. Fernández, E. A. Sanabria, L. B. Quiroga, and R. P. Vari. A new species of Silvinichthys (Siluriformes, Trichomycteridae) lacking pelvic fins from mid-elevation localities of the southern Andes, with comments on the genus. Journal of Fish Biology, 84(2):372–382, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2014:CDC

[1309] B. M. Taylor and J. H. Choat. Comparative demography of commercially important parrotfish species from Micronesia. Journal of Fish Biology, 84 (2):383–402, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bian:2014:ESU

[1310] X. Bian, X. Zhang, Y. Sakurai, X. Jin, T. Gao, R. Wan, and J. Yamamoto. Envelope surface ultrastructure and specific gravity of artificially fertilized Pacific cod Gadus macrocephalus eggs. Journal of Fish Biology, 84(2):403–421, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espinoza:2014:OSV

[1311] P. Espinoza and A. Bertrand. Ontogenetic and spatiotemporal variability in anchoveta Engraulis ringens diet off Peru. Journal of Fish Biology, 84 (2):422–435, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ji:2014:MCE

[1312] H. Ji, J. Wei, S. Wei, Y. Yan, Y. Huang, X. Huang, S. Zhou, Y. Zhou, and Q. Qin. Molecular cloning and expression of a C-type lectin-like protein from orange-spotted grouper Epinephelus coioides. Journal of Fish Biology, 84(2):436–447, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalamarz-Kubiak:2014:VEC

[1313] H. Kalamarz-Kubiak, I. Meiri-Ashkenazi, A. Kleszczyńska, and H. Rosenfeld. In vitro effect of cortisol and urotensin I on arginine vasotocin and isotocin secretion from pituitary cells of gilthead sea bream Sparus aurata. Journal of Fish Biology, 84(2):448–458, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez:2014:GEN

[1314] R. González, P. Dinghi, C. Corio, A. Medina, M. Maggioni, L. Storero, and A. Gosztonyi. Genetic evidence and new morphometric data as essential tools to identify the Patagonian seahorse Hippocampus patagonicus (Pisces, Syngnathidae). Journal of Fish Biology, 84(2):459–474, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camelier:2014:NSA

[1315] P. Camelier and A. M. Zanata. A new species of Astyanax Baird & Girard (Characiformes: Characidae) from the Rio Paraguaçu basin, Chapada Diamantina, Bahia, Brazil, with comments on bony hooks on all fins. Journal of Fish Biology, 84(2):475–490, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2014:ERL

[1316] C. F. Baker. Effect of ramp length and slope on the efficacy of a baffled fish pass. Journal of Fish Biology, 84(2):491–502, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulthen:2014:SIP

[1317] K. Hulthén, B. B. Chapman, P. A. Nilsson, L.-A. Hansson, C. Skov, H. Baktoft, J. Brodersen, and C. Brönmark. Sex identification and PITtagging: tools and prospects for studying intersexual differences in freshwater fishes. Journal of Fish Biology, 84(2):503–512, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bineesh:2014:RCC

[1318] K. K. Bineesh, K. V. Akhilesh, M. F. Gomon, E. M. Abdussamad, N. G. K. Pillai, and A. Gopalakrishnan. Redescription of Chlorophthalmus corniger, a senior synonym of Chlorophthalmus bicornis (family: Chlorophthalmidae). Journal of Fish Biology, 84(2):513–522, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:2014:MPB

[1319] M. A. Campbell and J. A. Lopéz. Mitochondrial phylogeography of a Beringian relict: the endemic freshwater genus of blackfish Dallia (Esociformes). Journal of Fish Biology, 84(2):523–538, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2014:C

[1320] Anonymous. Corrigendum. Journal of Fish Biology, 84(2):539, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temperoni:2014:R

[1321] B. Temperoni, M. D. Viñas, and C. C. Buratti. Referees 2013. Journal of Fish Biology, 84(2):540–546, February 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefevre:2014:ABFa

[1322] S. Lefevre, M. Bayley, D. J. Mckenzie, and J. F. Craig. Air-breathing fishes. Journal of Fish Biology, 84(3):547–553, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:2014:BWS

[1323] J. A. Nelson. Breaking wind to survive: fishes that breathe air with their gut. Journal of Fish Biology, 84(3):554–576, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2014:TVA

[1324] K. L. Martin. Theme and variations: amphibious air-breathing intertidal fishes. Journal of Fish Biology, 84(3):577–602, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chew:2014:ENM

[1325] S. F. Chew and Y. K. Ip. Excretory nitrogen metabolism and defence against ammonia toxicity in air-breathing fishes. Journal of Fish Biology, 84(3):603–638, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pace:2014:SPT

[1326] C. M. Pace and A. C. Gibb. Sustained periodic terrestrial locomotion in air-breathing fishes. Journal of Fish Biology, 84(3):639–660, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefevre:2014:SAB

[1327] S. Lefevre, P. Domenici, and D. J. McKenzie. Swimming in air-breathing fishes. Journal of Fish Biology, 84(3):661–681, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shartau:2014:ABI

[1328] R. B. Shartau and C. J. Brauner. Acid–base and ion balance in fishes with bimodal respiration. Journal of Fish Biology, 84(3):682–704, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefevre:2014:ABFb

[1329] S. Lefevre, T. Wang, A. Jensen, N. V. Cong, D. T. T. Huong, N. T. Phuong, and M. Bayley. Air-breathing fishes in aquaculture. What can we learn from physiology? Journal of Fish Biology, 84(3):705–731, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2014:MER

[1330] C. A. Brown and C. C. Green. Metabolic and embryonic responses to terrestrial incubation of Fundulus grandis embryos across a temperature gradient. Journal of Fish Biology, 84(3):732–747, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2014:SLT

[1331] K. Magellan, S. Pinchuck, and E. R. Swartz. Short and long-term strategies to facilitate aerial exposure in a galaxiid. Journal of Fish Biology, 84(3):748–758, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urbina:2014:ICG

[1332] M. A. Urbina, A. S. Meredith, C. N. Glover, and M. E. Forster. The importance of cutaneous gas exchange during aerial and aquatic respiration in galaxiids. Journal of Fish Biology, 84(3):759–773, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Toba:2014:RAS

[1333] A. Toba and A. Ishimatsu. Roles of air stored in burrows of the mudskipper Boleophthalmus pectinirostris for adult respiration and embryonic development. Journal of Fish Biology, 84(3):774–793, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendez-Sanchez:2014:EMO

[1334] J. F. Mendez-Sanchez and W. W. Burggren. Environmental modulation of the onset of air breathing and survival of Betta splendens and Trichopodus trichopterus. Journal of Fish Biology, 84(3):794–807, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blank:2014:HID

[1335] T. Blank and W. Burggren. Hypoxia-induced developmental plasticity of the gills and air-breathing organ of Trichopodus trichopterus. Journal of Fish Biology, 84(3):808–826, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Porteus:2014:ESH

[1336] C. S. Porteus, P. A. Wright, and W. K. Milsom. The effect of sustained hypoxia on the cardio-respiratory response of bowfin Amia calva: implications for changes in the oxygen transport system. Journal of Fish Biology, 84(3):827–843, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johannsson:2014:ABM

[1337] O. E. Johannsson, H. L. Bergman, C. M. Wood, P. Laurent, D. G. Kavembe, A. Bianchini, J. N. Maina, C. Chevalier, L. F. Bianchini, M. B. Papah, and R. O. Ojoo. Air breathing in Magadi tilapia Alcolapia grahami, under normoxic and hyperoxic conditions, and the association with sunlight and reactive oxygen species. Journal of Fish Biology, 84(3):844– 863, March 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2014:UNT

[1338] C. Azevedo, P. Matos, S. Rocha, E. Matos, E. Oliveira, S. Al-Quraishy, and G. Casal. Ultrastructure of novel thrombocytes in the dog snapper Lutjanus jocu. Journal of Fish Biology, 84(4):865–871, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2014:UNI

[1339] M. Correia, J. Palma, H. Koldewey, and J. P. Andrade. The use of a non-invasive tool for capture–recapture studies on a seahorse Hippocampus guttulatus population. Journal of Fish Biology, 84(4):872–884, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2014:CCE

[1340] J. H. Chen, C. H. Wang, Y. L. Li, H. M. Wang, X. J. Zhang, and B. L. Yan. cDNA cloning and expression characterization of serum transferrin gene from oriental weatherfish Misgurnus anguillicaudatus. Journal of Fish Biology, 84(4):885–896, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2014:CRM

[1341] W. E. Smith and T. J. Kwak. A capture–recapture model of amphidromous fish dispersal. Journal of Fish Biology, 84(4):897–912, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2014:OMT

[1342] W. E. Smith and T. J. Kwak. Otolith microchemistry of tropical diadromous fishes: spatial and migratory dynamics. Journal of Fish Biology, 84(4):913–928, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2014:GTL

[1343] T. Takahashi. Greenwoodochromini Takahashi from Lake Tanganyika is a junior synonym of Limnochromini Poll (Perciformes: Cichlidae). Journal of Fish Biology, 84(4):929–936, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borg:2014:SSD

[1344] J. P. G. Borg, M. Westerbom, and H. Lehtonen. Sex-specific distribution and diet of Platichthys flesus at the end of spawning in the northern Baltic Sea. Journal of Fish Biology, 84(4):937–951, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diogo:2014:APB

[1345] H. Diogo and J. G. Pereira. Assessing the potential biological implications of recreational inshore fisheries on sub-tidal fish communities of Azores (north-east Atlantic Ocean) using catch and effort data. Journal of Fish Biology, 84(4):952–970, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turek:2014:STE

[1346] K. C. Turek, M. A. Pegg, and K. L. Pope. Short-term evaluation of visible implant alpha tags in juveniles of three fish species under laboratory conditions. Journal of Fish Biology, 84(4):971–981, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sidhu:2014:UTT

[1347] R. Sidhu, K. Anttila, and A. P. Farrell. Upper thermal tolerance of closely related Danio species. Journal of Fish Biology, 84(4):982–995, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes:2014:OGL

[1348] E. A. P. Gomes, P. N. Campos, and A. C. T. Bonecker. Occurrence of Gobiidae larvae in a tropical Brazilian estuary, with particular emphasis on the use of size classes to categorize species guilds. Journal of Fish Biology, 84(4):996–1013, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2014:TVC

[1349] L. J. Wilson, M. T. Burrows, G. D. Hastie, and B. Wilson. Temporal variation and characterization of grunt sounds produced by Atlantic cod Gadus morhua and pollack Pollachius pollachius during the spawning season. Journal of Fish Biology, 84(4):1014–1030, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ashworth:2014:QDA

[1350] E. C. Ashworth, M. Depczynski, T. H. Holmes, and S. K. Wilson. Quantitative diet analysis of four mesopredators from a coral reef. Journal of Fish Biology, 84(4):1031–1045, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becerra-Jurado:2014:DPI

[1351] G. Becerra-Jurado, R. Cruikshanks, C. O’Leary, F. Kelly, R. Poole, and P. Gargan. Distribution, prevalence and intensity of Anguillicola crassus (Nematoda) infection in Anguilla anguilla in the Republic of Ireland.

Journal of Fish Biology, 84(4):1046–1062, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahe:2014:MVS

[1352] K. Mahe, M. C. Villanueva, S. Vaz, F. Coppin, P. Koubbi, and A. Carpentier. Morphological variability of the shape of striped red mullet Mullus surmuletus in relation to stock discrimination between the Bay of Biscay and the eastern English Channel. Journal of Fish Biology, 84 (4):1063–1073, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carter:2014:SVE

[1353] A. B. Carter, G. R. Russ, A. J. Tobin, A. J. Williams, C. R. Davies, and B. D. Mapstone. Spatial variation in the effects of size and age on reproductive dynamics of common coral trout Plectropomus leopardus. Journal of Fish Biology, 84(4):1074–1098, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayden:2014:TFR

[1354] B. Hayden, A. Massa-Gallucci, C. Harrod, M. O’Grady, J. Caffrey, and M. Kelly-Quinn. Trophic flexibility by roach Rutilus rutilus in novel habitats facilitates rapid growth and invasion success. Journal of Fish Biology, 84(4):1099–1116, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penny:2014:OCH

[1355] F. M. Penny and J. D. Kieffer. Oxygen consumption and haematology of juvenile shortnose sturgeon Acipenser brevirostrum during an acute 24 h saltwater challenge. Journal of Fish Biology, 84(4):1117–1135, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claydon:2014:MSS

[1356] J. A. B. Claydon, M. I. McCormick, and G. P. Jones. Multispecies spawning sites for fishes on a low-latitude coral reef: spatial and temporal patterns. Journal of Fish Biology, 84(4):1136–1163, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ficker:2014:FLE

[1357] H. Ficker, R. Mazzucco, H. Gassner, J. Wanzenböck, and U. Dieckmann. Fish length exclusively determines sexual maturation in the European whitefish Coregonus lavaretus species complex. Journal of Fish Biology, 84(4):1164–1170, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lucas:2014:ARB

[1358] J. Lucas, A. Schouman, L. Lyphout, X. Cousin, and C. Lefrancois. Allometric relationship between body mass and aerobic metabolism in zebrafish Danio rerio. Journal of Fish Biology, 84(4):1171–1178, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamidan:2014:RAH

[1359] N. Hamidan and A. Aloufi. Rediscovery of Acanthobrama hadiyahensis (Cyprinidae) in Saudi Arabia. Journal of Fish Biology, 84(4):1179–1184, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faria:2014:PDA

[1360] V. V. Faria, G. E. Chiaramonte, and M. J. P. van Oijen. Publication date and author spelling for the hidden angelshark Squatina occulta. Journal of Fish Biology, 84(4):1185–1187, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2014:SCA

[1361] R. L. Moran, C. N. von Ende, and B. H. King. Seasonal colour and antipredator behaviour in Etheostoma (Percidae). Journal of Fish Biology, 84(4):1188–1194, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nurminen:2014:FEP

[1362] L. Nurminen, S. Estlander, M. Olin, and H. Lehtonen. Feeding efficiency of planktivores under disturbance, the effect of water colour, predation threat and shoal composition. Journal of Fish Biology, 84(4):1195–1201, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langaard:2014:TSI

[1363] L. Langård, A. Slotte, G. Skaret, and A. Johannessen. Thermal stratification influences maturation and timing of spawning in a local Clupea harengus population. Journal of Fish Biology, 84(4):1202–1209, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ern:2014:ESO

[1364] R. Ern, D. T. T. Huong, N. V. Cong, M. Bayley, and T. Wang. Effect of salinity on oxygen consumption in fishes: a review. Journal of Fish Biology, 84(4):1210–1220, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaxion-Harm:2014:ETC

[1365] J. Jaxion-Harm and F. Ladich. Effects of temperature change on cortisol release by common carp Cyprinus carpio. Journal of Fish Biology, 84 (4):1221–1227, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kleinhappel:2014:MAL

[1366] T. K. Kleinhappel, A. Al-Zoubi, B. Al-Diri, O. Burman, P. Dickinson, L. John, A. Wilkinson, and T. W. Pike. A method for the automated long-term monitoring of three-spined stickleback Gasterosteus aculeatus shoal dynamics. Journal of Fish Biology, 84(4):1228–1233, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2014:VOD

[1367] M. G. Carvalho, A. S. Moreira, C. Moreira, H. Queiroga, P. T. Santos, and A. T. Correia. Validation of otolith daily increments in early juveniles of shanny Lipophrys pholis. Journal of Fish Biology, 84(4):1234– 1239, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2014:RMB

[1368] M. K. Taylor and S. J. Cooke. Repeatability of movement behaviour in a wild salmonid revealed by telemetry. Journal of Fish Biology, 84 (4):1240–1246, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moyano:2014:SMG

[1369] M. Moyano, S. Garrido, M. A. Teodósio, and M. A. Peck. Standard metabolism and growth dynamics of laboratory-reared larvae of Sardina pilchardus. Journal of Fish Biology, 84(4):1247–1255, April 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skrypzeck:2014:FRR

[1370] H. Skrypzeck, A. G. V. Salvanes, B. Currie, and A. Kotze. First records of reproductive behaviour and early development of the bearded goby Sufflogobius bibarbatus. Journal of Fish Biology, 84(4):1256–1261, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2014:BRK

[1371] Laith A. Jawad. Book review: Know your New Zealand Fishes — By J. Enderby and T.Enderby. Journal of Fish Biology, 84(4):1262–1263, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2014:BRO

[1372] Laith A. Jawad. Book review: Otolith Atlas of Taiwan Fishes — By C.-H. Lin and C.-W. Chang. Journal of Fish Biology, 84(4):1264–1265, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bourdon:2014:BRB

[1373] Robert J. Bourdon, Mauricio Carrasquilla-Henao, Brenna Collicutt, Cameron Freshwater, Olivia McMillan, Amber Messmer, James P. W. Robinson, Easton R. White, and Francis Juanes. Book review: The Biology of Sharks and Rays — By A. P. Klimley. Journal of Fish Biology, 84(4):1266–1267, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2014:BRS

[1374] Ian J. Winfield. Book review: Scientific Writing for Impact Factor Journals — By E. Lichtfouse. Journal of Fish Biology, 84(4):1268–1269, April 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taguchi:2014:USC

[1375] T. Taguchi, Y. Miura, D. Krueger, and S. Sugiura. Utilizing stomach content and faecal DNA analysis techniques to assess the feeding behaviour of largemouth bass Micropterus salmoides and bluegill Lepomis macrochirus. Journal of Fish Biology, 84(5):1271–1288, May 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakamura:2014:DSH

[1376] T. Nakamura and T. Doi. Do stocked hatchery-reared juveniles ecologically suppress wild juveniles in Salvelinus leucomaenis? Journal of Fish Biology, 84(5):1289–1299, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mckenzie:2014:CSL

[1377] R. W. Mckenzie, P. J. Motta, and J. R. Rohr. Comparative squamation of the lateral line canal pores in sharks. Journal of Fish Biology, 84 (5):1300–1311, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ilies:2014:DRA

[1378] I. Ilieş, I. M. Traniello, R. F. Sîrbulescu, and G. K. H. Zupanc. Determination of relative age using growth increments of scales as a minimally invasive method in the tropical freshwater Apteronotus leptorhynchus.

Journal of Fish Biology, 84(5):1312–1325, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulsen:2014:CDB

[1379] S. B. Poulsen, J. C. Svendsen, K. Aarestrup, and H. Malte. Calciumdependent behavioural responses to acute copper exposure in Oncorhynchus mykiss. Journal of Fish Biology, 84(5):1326–1339, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2014:AGP

[1380] J. White, C. A. Simpfendorfer, A. J. Tobin, and M. R. Heupel. Age and growth parameters of shark-like batoids. Journal of Fish Biology, 84(5):1340–1353, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pattrick:2014:EES

[1381] P. Pattrick and N. A. Strydom. The effects of exposure in sandy beach surf zones on larval fishes. Journal of Fish Biology, 84(5):1354–1376, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Michalkova:2014:EEP

[1382] V. Michálková and M. Ondračková. Experimental evidence for parasiteinduced over-winter mortality in juvenile Rhodeus amarus. Journal of Fish Biology, 84(5):1377–1388, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2014:RIB

[1383] M. Zhou and R. C. Fuller. Reproductive isolation between two darter species is enhanced and asymmetric in sympatry. Journal of Fish Biology, 84(5):1389–1400, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chicoli:2014:EFS

[1384] A. Chicoli, S. Butail, Y. Lun, J. Bak-Coleman, S. Coombs, and D. A. Paley. The effects of flow on schooling Devario aequipinnatus: school structure, startle response and information transmission. Journal of Fish Biology, 84(5):1401–1421, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Afeworki:2014:SLP

[1385] Y. Afeworki, J. J. Videler, Y. H. Berhane, and J. H. Bruggemann. Seasonal and life-phase related differences in growth in Scarus ferrugineus on a southern Red Sea fringing reef. Journal of Fish Biology, 84(5):1422– 1438, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anttila:2014:FWT

[1386] K. Anttila, E. J. Eliason, K. H. Kaukinen, K. M. Miller, and A. P. Farrell. Facing warm temperatures during migration: cardiac mRNA responses of two adult Oncorhynchus nerka populations to warming and swimming challenges. Journal of Fish Biology, 84(5):1439–1456, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2014:TSS

[1387] N. Jonsson and B. Jonsson. Time and size at seaward migration influence the sea survival of Salmo salar. Journal of Fish Biology, 84(5):1457– 1473, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldenberg:2014:SFC

[1388] S. U. Goldenberg and K. Erzini. Seagrass feeding choices and digestive strategies of the herbivorous fish Sarpa salpa. Journal of Fish Biology, 84(5):1474–1489, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imre:2014:BRA

[1389] I. Imre, R. T. Di Rocco, C. F. Belanger, G. E. Brown, and N. S. Johnson. The behavioural response of adult Petromyzon marinus to damagereleased alarm and predator cues. Journal of Fish Biology, 84(5):1490– 1502, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamashita:2014:EGD

[1390] Y. T. Yamashita, M. Aritaki, Y. Kurita, and M. Tanaka. Early growth and development of reciprocal hybrids of the starry flounder Platichthys stellatus and stone flounder Kareius bicoloratus. Journal of Fish Biology, 84(5):1503–1518, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stein:2014:PHN

[1391] D. L. Stein and J. C. Drazen. Paraliparis hawaiiensis, a new species of snailfish (Scorpaeniformes: Liparidae) and the first described from the Hawaiian Archipelago. Journal of Fish Biology, 84(5):1519–1526, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2014:RGI

[1392] L. Zhang, R. E. Gozlan, Z. Li, J. Liu, T. Zhang, W. Hu, and Z. Zhu. Rapid growth increases intrinsic predation risk in genetically modified Cyprinus carpio: implications for environmental risk. Journal of Fish Biology, 84(5):1527–1538, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Netto-Ferreira:2014:NSR

[1393] A. L. Netto-Ferreira, J. L. O. Birindelli, L. M. Sousa, and N. A. Menezes. A new species of Rhinopetitia Géry 1964 (Ostariophysi: Characiformes: Characidae) from the Rio Teles Pires, Rio Tapajós basin, Brazil. Journal of Fish Biology, 84(5):1539–1550, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2014:ECH

[1394] X. H. Wang, K. R. Wang, P. Nie, X. H. Chen, and J. Q. Ao. Establishment and characterization of a head kidney cell line from large yellow croaker Pseudosciaena crocea. Journal of Fish Biology, 84(5):1551–1561, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wasylenko:2014:SNS

[1395] B. A. Wasylenko, D. T. Callaghan, P. J. Blanchfield, and G. G. Pyle. Salvelinus namaycush spawning substratum attracts egg predators and opportunists through chemosensory cues. Journal of Fish Biology, 84 (5):1562–1573, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bourg:2014:MEI

[1396] B. Le Bourg, J. Kiszka, and P. Bustamante. Mother–embryo isotope (δ 15N, δ ^{13}C) fractionation and mercury (Hg) transfer in aplacental deepsea sharks. Journal of Fish Biology, 84(5):1574–1581, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baines:2014:TMF

[1397] D. C. Baines, M. A. Purnell, and P. J. B. Hart. Tooth microwear formation rate in Gasterosteus aculeatus. Journal of Fish Biology, 84(5): 1582–1589, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciccotto:2014:ABP

[1398] P. J. Ciccotto, D. J. Dresser, and T. C. Mendelson. Association between parasite load and orange, but not blue, male nuptial colouration in Etheostoma caeruleum. Journal of Fish Biology, 84(5):1590–1598, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Behrens:2014:PCL

[1399] J. W. Behrens, H. Seth, M. Axelsson, and K. Buchmann. The parasitic copepod Lernaeocera branchialis negatively affects cardiorespiratory function in Gadus morhua. Journal of Fish Biology, 84(5):1599–1606, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2014:RCM

[1400] S. Yamamoto, T. Yokoduka, K. Fujimoto, K. Takagi, and T. Ono. Radiocaesium concentrations in the muscle and eggs of salmonids from Lake Chuzenji, Japan, after the Fukushima fallout. Journal of Fish Biology, 84(5):1607–1613, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hume:2014:MAP

[1401] J. B. Hume, C. W. Bean, and C. E. Adams. Morphological abnormalities in a population of Lampetra planeri, with a short review of petromyzontid teratologies. Journal of Fish Biology, 84(5):1614–1619, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bologna:2014:MLL

[1402] P. A. X. Bologna. Mangrove loss leads to fish hyperutilization of seagrass beds in a UNESCO biosphere reserve. Journal of Fish Biology, 84(5): 1620–1625, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bardarson:2014:GGR

[1403] H. Bardarson, B. J. McAdam, G. Petursdottir, and G. Marteinsdottir. To glue or not to glue? Reassembling broken otoliths for population discrimination. Journal of Fish Biology, 84(5):1626–1633, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2014:BRB

[1404] Ian J. Winfield. Book review: Britain’s Game Fishes: Celebration and Conservation of Salmonids. By M. Everard and P. Knight. Journal of Fish Biology, 84(5):1634–1635, May 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2014:ORJ

[1405] Carl Smith. Obituary: Robert John Wootton (1942–2014). Journal of Fish Biology, 84(6):1637–1639, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2014:WTI

[1406] J. L. A. Jensen, A. H. Rikardsen, E. B. Thorstad, A. H. Suhr, J. G. Davidsen, and R. Primicerio. Water temperatures influence the marine area use of Salvelinus alpinus and Salmo trutta. Journal of Fish Biology, 84(6):1640–1653, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fanelli:2014:TEL

[1407] E. Fanelli, V. Papiol, J. E. Cartes, and O. Rodriguez-Romeu. Trophic ecology of Lampanyctus crocodilus on north-west Mediterranean Sea slopes in relation to reproductive cycle and environmental variables. Journal of Fish Biology, 84(6):1654–1688, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Higgins:2014:SAP

[1408] B. A. Higgins and M. H. Horn. Suction among pickers: jaw mechanics, dietary breadth and feeding behaviour in beach-spawning Leuresthes spp. compared with their relatives. Journal of Fish Biology, 84(6):1689– 1707, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eya:2014:EDL

[1409] J. C. Eya, R. Yossa, M. F. Ashame, C. F. Pomeroy, and A. L. Gannam. Effects of dietary lipid levels on mitochondrial gene expression in low and high-feed efficient families of rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 84(6):1708–1720, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gum:2014:GDU

[1410] B. Gum, J. Geist, S. Eckenfels, and A. Brinker. Genetic diversity of upper Lake Constance whitefish Coregonus spp. under the influence of fisheries: a DNA study based on archived scale samples from 1932, 1975 and 2006. Journal of Fish Biology, 84(6):1721–1739, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marohn:2014:IAE

[1411] L. Marohn, E. Prigge, and R. Hanel. Introduced American eels Anguilla rostrata in European waters: life-history traits in a non-native environment. Journal of Fish Biology, 84(6):1740–1747, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bai:2014:NAD

[1412] L.-R. Bai, A.-L. Wang, Z.-Y. Zhao, and Y.-T. Miao. Neurobehavioural analysis of developmental iron deficiency in Oreochromis aureus × Oreochromis niloticus. Journal of Fish Biology, 84(6):1748–1767, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2014:EBN

[1413] L. Jacobsen, H. Baktoft, N. Jepsen, K. Aarestrup, S. Berg, and C. Skov. Effect of boat noise and angling on lake fish behaviour. Journal of Fish Biology, 84(6):1768–1780, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drost:2014:UTL

[1414] H. E. Drost, E. C. Carmack, and A. P. Farrell. Upper thermal limits of cardiac function for Arctic cod Boreogadus saida, a key food web fish species in the Arctic Ocean. Journal of Fish Biology, 84(6):1781– 1792, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrews:2014:IGS

[1415] C. D. Andrews, J. F. Payne, and M. L. Rise. Identification of a gene set to evaluate the potential effects of loud sounds from seismic surveys on the ears of fishes: a study with Salmo salar. Journal of Fish Biology, 84(6):1793–1819, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kopf:2014:OCP

[1416] S. M. Kopf, P. Humphries, and R. J. Watts. Ontogeny of critical and prolonged swimming performance for the larvae of six Australian freshwater fish species. Journal of Fish Biology, 84(6):1820–1841, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leclercq:2014:GDS

[1417] E. Leclercq, B. Grant, A. Davie, and H. Migaud. Gender distribution, sexual size dimorphism and morphometric sexing in ballan wrasse Labrus bergylta. Journal of Fish Biology, 84(6):1842–1862, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zuckerman:2014:IBL

[1418] Z. C. Zuckerman, D. P. Philipp, and C. D. Suski. The influence of brood loss on nest abandonment decisions in largemouth bass Micropterus salmoides. Journal of Fish Biology, 84(6):1863–1875, June 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luque:2014:DAT

[1419] P. L. Luque, E. Rodriguez-Marin, J. Landa, M. Ruiz, P. Quelle, D. Macias, and J. M. Ortiz De Urbina. Direct ageing of Thunnus thynnus from the eastern Atlantic Ocean and western Mediterranean Sea using dorsal fin spines. Journal of Fish Biology, 84(6):1876–1903, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joensen:2014:DAN

[1420] H. Joensen and O. Grahl-Nielsen. Distinction among North Atlantic cod Gadus morhua stocks by tissue fatty acid profiles. Journal of Fish Biology, 84(6):1904–1925, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Busst:2014:PAL

[1421] G. M. Busst and J. R. Britton. Precision of the age–length increments of three cyprinids: effects of fish number and sub-sampling strategy. Journal of Fish Biology, 84(6):1926–1939, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villegas-Hernandez:2014:LHT

[1422] H. Villegas-Hernández, M. Muñoz, and J. Lloret. Life-history traits of temperate and thermophilic barracudas (Teleostei: Sphyraenidae) in the context of sea warming in the Mediterranean Sea. Journal of Fish Biology, 84(6):1940–1957, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kasper:2014:OGP

[1423] J. M. Kasper, B. Bertelsen, H. G. Ólafsson, J. C. Holst, J. Sturlaugsson, and S. P. Jónsson. Observations of growth and postspawning survival of lumpfish Cyclopterus lumpus from mark-recapture studies. Journal of Fish Biology, 84(6):1958–1963, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orsi:2014:LTC

[1424] M. L. Orsi and J. R. Britton. Long-term changes in the fish assemblage of a neotropical hydroelectric reservoir. Journal of Fish Biology, 84(6): 1964–1970, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dAuriac:2014:RQM

[1425] M. B. Anglès d’Auriac, H. A. Urke, and T. Kristensen. A rapid qPCR method for genetic sex identification of Salmo salar and Salmo trutta including simultaneous elucidation of interspecies hybrid paternity by high-resolution melt analysis. Journal of Fish Biology, 84(6):1971–1977, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Closs:2014:BRP

[1426] G. P. Closs. Book review: A Photographic Guide to Freshwater Fishes of New Zealand — By S. McQueen. Journal of Fish Biology, 84(6):1978– 1979, June 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fleming:2014:RAR

[1427] I. A. Fleming, D. L. Bottom, K. K. Jones, C. A. Simenstad, and J. F. Craig. Resilience of anadromous and resident salmonid populations. Journal of Fish Biology, 85(1):1–7, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vincenzi:2014:EED

[1428] S. Vincenzi, A. J. Crivelli, W. H. Satterthwaite, and M. Mangel. Ecoevolutionary dynamics induced by massive mortality events. Journal of Fish Biology, 85(1):8–30, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2014:RNR

[1429] B. E. Craig, C. A. Simenstad, and D. L. Bottom. Rearing in natural and recovering tidal wetlands enhances growth and life-history diversity of Columbia Estuary tributary Coho salmon Oncorhynchus kisutch population. Journal of Fish Biology, 85(1):31–51, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2014:CER

[1430] K. K. Jones, T. J. Cornwell, D. L. Bottom, L. A. Campbell, and S. Stein. The contribution of estuary-resident life histories to the return of adult Oncorhynchus kisutch. Journal of Fish Biology, 85(1):52–80, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hammar:2014:NRA

[1431] J. Hammar. Natural resilience in Arctic charr Salvelinus alpinus: life history, spatial and dietary alterations along gradients of interspecific interactions. Journal of Fish Biology, 85(1):81–118, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:2014:EHC

[1432] D. J. Fraser. Evolutionary hypotheses for a constraint to life-history resilience in depleted Salmo salar populations. Journal of Fish Biology, 85(1):119–131, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McPhee:2014:ERP

[1433] M. V. McPhee, D. C. Whited, K. V. Kuzishchin, and J. A. Stanford. The effects of riverine physical complexity on anadromy and genetic diversity in steelhead or rainbow trout Oncorhynchus mykiss around the Pacific rim. Journal of Fish Biology, 85(1):132–150, July 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2014:EEI

[1434] B. Jonsson and N. Jonsson. Early environment influences later performance in fishes. Journal of Fish Biology, 85(2):151–188, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kimura-Kawaguchi:2014:IHR

[1435] M. R. Kimura-Kawaguchi, M. Horita, S. Abe, K. Arai, M. Kawata, and H. Munehara. Identification of hemiclonal reproduction in three species of Hexagrammos marine reef fishes. Journal of Fish Biology, 85(2):189– 209, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osachoff:2014:ABS

[1436] H. L. Osachoff, K. N. Osachoff, A. E. Wickramaratne, E. K. Gunawardane, F. P. Venturini, and C. J. Kennedy. Altered burst swimming in rainbow trout Oncorhynchus mykiss exposed to natural and synthetic oestrogens. Journal of Fish Biology, 85(2):210–227, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beshera:2014:MDP

[1437] K. A. Beshera and P. M. Harris. Mitochondrial DNA phylogeography of the Labeobarbus intermedius complex (Pisces, Cyprinidae) from Ethiopia. Journal of Fish Biology, 85(2):228–245, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bornatowski:2014:FET

[1438] H. Bornatowski, R. R. Braga, V. Abilhoa, and M. F. M. Corrêa. Feeding ecology and trophic comparisons of six shark species in a coastal ecosystem off southern Brazil. Journal of Fish Biology, 85(2):246–263, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeager:2014:DVG

[1439] L. A. Yeager, C. A. Layman, and C. M. Hammerschlag-Peyer. Diet variation of a generalist fish predator, grey snapper Lutjanus griseus, across an estuarine gradient: trade-offs of quantity for quality? Journal of Fish Biology, 85(2):264–277, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wei:2014:CTD

[1440] R. Wei, C. Zhou, D. Yuan, T. Wang, F. Lin, H. Chen, H. Wu, Z. Xin, S. Yang, Y. Wang, D. Chen, J. Liu, Y. Gao, and Z. Li. Characterization, tissue distribution and regulation of neuropeptide Y in Schizothorax prenanti. Journal of Fish Biology, 85(2):278–291, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horreo:2014:LTE

[1441] J. L. Horreo, G. Machado-Schiaffino, A. M. Griffiths, D. Bright, J. R. Stevens, and E. Garcia-Vazquez. Long-term effects of stock transfers: synergistic introgression of allochthonous genomes in salmonids. Journal of Fish Biology, 85(2):292–306, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loh:2014:DBF

[1442] W. K. W. Loh, P. Bond, K. J. Ashton, D. T. Roberts, and I. R. Tibbetts. DNA barcoding of freshwater fishes and the development of a quantitative qPCR assay for the species-specific detection and quantification of fish larvae from plankton samples. Journal of Fish Biology, 85(2):307– 328, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feng:2014:CSZ

[1443] D. Feng, Q. Y. Huang, K. Liu, S. C. Zhang, and Z. H. Liu. Comparative studies of zebrafish Danio rerio lipoprotein lipase (lpl) and hepatic lipase (lipc) genes belonging to the lipase gene family: evolution and expression pattern. Journal of Fish Biology, 85(2):329–342, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2014:EDS

[1444] M. Lin, M. Chevalier, S. Lek, L. Zhang, R. E. Gozlan, J. Liu, T. Zhang, S. Ye, W. Li, and Z. Li. Eutrophication as a driver of r-selection traits in a freshwater fish. Journal of Fish Biology, 85(2):343–354, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DAguillo:2014:DCF

[1445] M. C. D’Aguillo, A. S. Harold, and T. L. Darden. Diet composition and feeding ecology of the naked goby Gobiosoma bosc (Gobiidae) from four western Atlantic estuaries. Journal of Fish Biology, 85(2):355–373, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ODonnell:2014:GPS

[1446] T. P. O’Donnell, M. R. Denson, and T. L. Darden. Genetic population structure of spotted seatrout Cynoscion nebulosus along the southeastern U.S.A. Journal of Fish Biology, 85(2):374–393, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stern:2014:DPS

[1447] N. Stern, Y. Levitt, B. S. Galil, A. Diamant, M. B. Yokeş, and M. Goren. Distribution and population structure of the alien Indo–Pacific Randall’s threadfin bream Nemipterus randalli in the eastern Mediterranean Sea. Journal of Fish Biology, 85(2):394–406, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2014:CSP

[1448] T. Wang, D. Yuan, C. Zhou, F. Lin, H. Chen, H. Wu, R. Wei, Z. Xin, J. Liu, Y. Gao, D. Chen, S. Yang, Y. Pu, and Z. Li. Characterization of Schizothorax prenanti cgnrhII gene: fasting affects cgnrhII expression. Journal of Fish Biology, 85(2):407–420, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Press:2014:TCO

[1449] Y. K. Press, R. S. McBride, and M. J. Wuenschel. Time course of oocyte development in winter flounder Pseudopleuronectes americanus and spawning seasonality for the Gulf of Maine, Georges Bank and southern New England stocks. Journal of Fish Biology, 85(2):421–445, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dietrich:2014:CLM

[1450] G. J. Dietrich, A. Pecio, J. Sikorska, J. Wolnicki, M. Słowińska, P. Hliwa, R. Kamiński, A. Andronowska, and A. Ciereszko. Characterization of lake minnow Eupallasella percnurus semen in relation to sperm morphology, regulation of sperm motility and interpopulation diversity. Journal of Fish Biology, 85(2):446–455, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanDerLingen:2014:SOI

[1451] C. D. Van Der Lingen and T. W. Miller. Spatial, ontogenetic and interspecific variability in stable isotope ratios of nitrogen and carbon of Merluccius capensis and Merluccius paradoxus off South Africa. Journal of Fish Biology, 85(2):456–472, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Francisco:2014:GGT

[1452] S. M. Francisco, J. I. Robalo, S. Stefanni, A. Levy, and V. C. Almada. Gaidropsarus (Gadidae, Teleostei) of the North Atlantic Ocean: a brief phylogenetic review. Journal of Fish Biology, 85(2):473–487, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marques:2014:GBN

[1453] P. S. Marques, M. F. Costa, C. C. D. Corrêa, M. M. Marinho, and R. Mazzoni. Grazing behaviour of a non-herbivorous characin: revisiting plasticity. Journal of Fish Biology, 85(2):488–493, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Escobar-Sanchez:2014:FCS

[1454] O. Escobar-Sánchez, X. G. Moreno-Sánchez, C. A. Aguilar-Cruz, and L. A. Abitia-Cárdenas. First case of synophthalmia and albinism in the Pacific angel shark Squatina californica. Journal of Fish Biology, 85 (2):494–501, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Portnoy:2014:PWR

[1455] D. S. Portnoy, C. M. Hollenbeck, J. S. Johnston, H. M. Casman, and J. R. Gold. Parthenogenesis in a whitetip reef shark Triaenodon obesus involves a reduction in ploidy. Journal of Fish Biology, 85(2):502–508, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaha:2014:PNP

[1456] M. Bláha, I. Šetlíková, J. Peterka, J. Musil, and T. Policar. Planktonic or non-planktonic food in young-of-the-year European perch Perca fluviatilis in ponds. Journal of Fish Biology, 85(2):509–515, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2014:MCE

[1457] X. W. Chen, S. Jiang, Y. F. Gu, and Z. Y. Shi. Molecular characterization and expression of cyp19a gene in Carassius auratus. Journal of Fish Biology, 85(2):516–522, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sparrevohn:2014:SPT

[1458] C. R. Sparrevohn, K. Aarestrup, C. Stenberg, and D. Righton. Scanning for PIT-tagged flatfish in a coastal area using a sledge equipped with an RFID antenna. Journal of Fish Biology, 85(2):523–529, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maric:2014:FMS

[1459] S. Marić, S. Alekseyev, A. Snoj, O. Askeyev, I. Askeyev, and S. Weiss. First mtDNA sequencing of Volga and Ob basin taimen Hucho taimen: European populations stem from a late Pleistocene expansion of H. taimen out of western Siberia and are not intermediate to Hucho hucho. Journal of Fish Biology, 85(2):530–539, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Devine:2014:FRB

[1460] B. M. Devine and J. A. D. Fisher. First records of the blue runner Caranx crysos (Perciformes: Carangidae) in Newfoundland waters. Journal of Fish Biology, 85(2):540–545, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez-Chavez:2014:PGE

[1461] C. C. Martínez-Chávez, A. Tello-Ballinas, J. Fonseca-Madrigal, L. G. Ross, and C. A. Martínez-Palacios. Photoperiodic growth enhancement in a tropical batch spawning atherinopsid, pike silverside Chirostoma estor. Journal of Fish Biology, 85(2):546–553, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2014:DFO

[1462] Laith A. Jawad. The diversity of fish otoliths past and present. Journal of Fish Biology, 85(2):554–555, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hospital:2014:BRB

[1463] Justin Hospital. Book review: Billion-dollar fish: the untold story of Alaska pollock. Journal of Fish Biology, 85(2):556–557, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2014:SAU

[1464] Stephen J. Lockwood. The sea around us. Journal of Fish Biology, 85 (2):558–559, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsikliras:2014:CHI

[1465] A. C. Tsikliras and K. I. Stergiou. Condition and health indicators of exploited marine fishes. Journal of Fish Biology, 85(2):560–561, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2014:E

[1466] Anonymous. Erratum. Journal of Fish Biology, 85(2):562, August 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2014:DNC

[1467] M.-Y. Lee, T. A. Munroe, and K.-T. Shao. Description of a new cryptic, shallow-water tonguefish (Pleuronectiformes: Cynoglossidae: Symphurus) from the western North Pacific Ocean. Journal of Fish Biology, 85 (3):563–585, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2014:RNC

[1468] J. G. Davidsen, M. Daverdin, J. V. Arnekleiv, L. Rønning, A. D. Sjursen, and J. I. Koksvik. Riverine and near coastal migration performance of hatchery brown trout Salmo trutta. Journal of Fish Biology, 85(3): 586–596, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holley:2014:PFS

[1469] E. D. Holley, K. J. Ludwig, D. C. McEwen, and B. D. Wisenden. Predictability of food supply, but not ration, increases exploratory behaviour in zebrafish Danio rerio. Journal of Fish Biology, 85(3):597–604, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birba:2014:PCC

[1470] A. Birba, M. R. Ramallo, L. Morandini, V. Villafañe, C. Tubert, R. Guimarães Moreira, and M. Pandolfi. The pineal complex in the cichlid Cichlasoma dimerus: effect of different photoperiods on its cell morphology. Journal of Fish Biology, 85(3):605–620, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spares:2014:FFA

[1471] A. D. Spares, M. J. Dadswell, J. MacMillan, R. Madden, R. K. O’Dor, and M. J. W. Stokesbury. To fast or feed: an alternative life history for anadromous brook trout Salvelinus fontinalis overwintering within a harbour. Journal of Fish Biology, 85(3):621–644, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2014:ESW

[1472] L. Persson and A. Alanärä. The effect of shelter on welfare of juvenile Atlantic salmon Salmo salar reared under a feed restriction regimen. Journal of Fish Biology, 85(3):645–656, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mitchell:2014:QIE

[1473] J. D. Mitchell, K. J. Collins, P. I. Miller, and L. A. Suberg. Quantifying the impact of environmental variables upon catch per unit effort of the blue shark Prionace glauca in the western English Channel. Journal of Fish Biology, 85(3):657–670, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lemos:2014:MRB

[1474] V. M. Lemos, A. S. Varela, Jr., P. R. Schwingel, J. H. Muelbert, and J. P. Vieira. Migration and reproductive biology of Mugil liza (Teleostei: Mugilidae) in south Brazil. Journal of Fish Biology, 85(3):671–687, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frazier:2014:VAG

[1475] B. S. Frazier, W. B. Driggers III, D. H. Adams, C. M. Jones, and J. K. Loefer. Validated age, growth and maturity of the bonnethead Sphyrna tiburo in the western North Atlantic Ocean. Journal of Fish Biology, 85 (3):688–712, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knickle:2014:EFS

[1476] D. C. Knickle and G. A. Rose. Examination of fine-scale spatial-temporal overlap and segregation between two closely related congeners Gadus morhua and Gadus ogac in coastal Newfoundland. Journal of Fish Biology, 85(3):713–735, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korwin-Kossakowski:2014:PCS

[1477] M. Korwin-Kossakowski. Physical changes in specimens of five species of Cyprinidae preserved in ethanol and frozen. Journal of Fish Biology, 85 (3):736–751, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christian:2014:EPI

[1478] J. M. Christian and G. L. Adams. Effects of pool isolation on trophic ecology of fishes in a highland stream. Journal of Fish Biology, 85(3): 752–772, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biswas:2014:IMP

[1479] S. Biswas, K. Jahir Hussain, N. P. I. Das, B. C. Russell, K. K. Satpathy, and S. S. Mishra. Imprint of monsoonal patterns on the fish assemblage in coastal waters of south-east India: a case study. Journal of Fish Biology, 85(3):773–799, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Modica:2014:FCF

[1480] L. Modica, J. E. Cartes, and M. Carrassón. Food consumption of five deep-sea fishes in the Balearic Basin (western Mediterranean Sea): are there daily feeding rhythms in fishes living below 1000 m? Journal of Fish Biology, 85(3):800–820, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrkjedal:2014:MMV

[1481] I. Byrkjedal, S. Hadler-Jacobsen, D. Rees, and A. Orlov. Molecular and morphological variation in saddled fathead sculpins (Psychrolutidae: Cottunculus) of the north-east Atlantic Ocean, with a synonymy of three species. Journal of Fish Biology, 85(3):821–837, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yanase:2014:UAT

[1482] K. Yanase, N. A. Herbert, and J. C. Montgomery. Unilateral ablation of trunk superficial neuromasts increases directional instability during steady swimming in the yellowtail kingfish Seriola lalandi. Journal of Fish Biology, 85(3):838–856, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pattrick:2014:LFV

[1483] P. Pattrick and N. A. Strydom. Larval fish variability in response to oceanographic features in a nearshore nursery area. Journal of Fish Biology, 85(3):857–881, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Preston:2014:SFA

[1484] A. C. Preston, J. F. Taylor, C. E. Adams, and H. Migaud. Surface feeding and aggressive behaviour of diploid and triploid brown trout Salmo trutta during allopatric pair-wise matchings. Journal of Fish Biology, 85(3): 882–900, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2014:ABD

[1485] B. M. Taylor, K. L. Rhodes, A. Marshell, and J. L. McIlwain. Age-based demographic and reproductive assessment of orangespine Naso lituratus and bluespine Naso unicornis unicornfishes. Journal of Fish Biology, 85 (3):901–916, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irschick:2014:NMM

[1486] D. J. Irschick and N. Hammerschlag. A new metric for measuring condition in large predatory sharks. Journal of Fish Biology, 85(3):917–926, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vieira:2014:SOC

[1487] R. P. Vieira and M. R. Cunha. In situ observation of chimaerid species in the Gorringe Bank: new distribution records for the north-east Atlantic Ocean. Journal of Fish Biology, 85(3):927–932, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bondarenko:2014:CIS

[1488] O. Bondarenko, B. Dzyuba, J. Cosson, M. Rodina, and O. Linhart. Calcium ion supplementation increases brook trout Salvelinus fontinalis spermatozoa activation at the end of the spawning season. Journal of Fish Biology, 85(3):933–937, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2014:MMO

[1489] C. G. B. Gonçalves, C. Medeiros, L. G. Abraçado, and D. Acosta-Avalos. Magnetic material in the ocellar spot and lateral line of tomtates Haemulon aurolineatum. Journal of Fish Biology, 85(3):938–943, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beckmann:2014:EHP

[1490] C. L. Beckmann, J. G. Mitchell, L. Seuront, D. A. J. Stone, and C. Huveneers. From egg to hatchling: preferential retention of fatty acid biomarkers in young-of-the-year Port Jackson sharks Heterodontus portusjacksoni. Journal of Fish Biology, 85(3):944–952, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faria:2014:CSS

[1491] A. M. Faria, R. Borges, and E. J. Gonçalves. Critical swimming speeds of wild-caught sand-smelt Atherina presbyter larvae. Journal of Fish Biology, 85(3):953–959, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schedina:2014:ICE

[1492] I. M. Schedina, S. Pfautsch, S. Hartmann, N. Dolgener, A. Polgar, P. G. Bianco, R. Tiedemann, and V. Ketmaier. Isolation and characterization of eight microsatellite loci in the brook lamprey Lampetra planeri (Petromyzontiformes) using 454 sequence data. Journal of Fish Biology, 85(3):960–964, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henrich:2014:EVT

[1493] T. Henrich, N. Hafer, and K. B. Mobley. Effects of VIE tagging and partial tissue sampling on the immune response of three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 85(3):965–971, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2014:ICP

[1494] Y. Li, J. Wang, K. Zhao, B. Xie, and Z. Peng. Isolation and characterization of 16 polymorphic microsatellite loci for Euchiloglanis kishinouyei. Journal of Fish Biology, 85(3):972–977, September 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melia:2014:SME

[1495] P. Melià, A. J. Crivelli, C. Durif, R. Poole, and D. Bevacqua. A simplified method to estimate body growth parameters of the European eel Anguilla anguilla. Journal of Fish Biology, 85(3):978–984, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2014:MCS

[1496] Stephen J. Lockwood. Marine conservation: science, policy, and management. Journal of Fish Biology, 85(3):985–986, September 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shrimpton:2014:FMP

[1497] J. M. Shrimpton, K. D. Warren, N. L. Todd, C. J. McRae, G. J. Glova, K. H. Telmer, and A. D. Clarke. Freshwater movement patterns by juvenile Pacific salmon Oncorhynchus spp. before they migrate to the ocean: Oh the places you’ll go! Journal of Fish Biology, 85(4):987–1004, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCormick:2014:MPH

[1498] S. D. McCormick, A. Haro, D. T. Lerner, M. F. O’Dea, and A. M. Regish. Migratory patterns of hatchery and stream-reared Atlantic salmon Salmo salar smolts in the Connecticut River, U.S.A. Journal of Fish Biology, 85 (4):1005–1022, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zydlewski:2014:PCD

[1499] G. B. Zydlewski, D. S. Stich, and S. D. McCormick. Photoperiod control of downstream movements of Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 85(4):1023–1041, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riley:2014:DSB

[1500] W. D. Riley, A. T. Ibbotson, D. L. Maxwell, P. I. Davison, W. R. C. Beaumont, and M. J. Ives. Development of schooling behaviour during the downstream migration of Atlantic salmon Salmo salar smolts in a chalk stream. Journal of Fish Biology, 85(4):1042–1059, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norrgaard:2014:EFR

[1501] J. R. Norrgård, E. Bergman, M. Schmitz, and L. A. Greenberg. Effects of feeding regimes and early maturation on migratory behaviour of landlocked hatchery-reared Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 85(4):1060–1073, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stich:2014:SAS

[1502] D. S. Stich, M. M. Bailey, and J. D. Zydlewski. Survival of Atlantic salmon Salmo salar smolts through a hydropower complex. Journal of Fish Biology, 85(4):1074–1096, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Billman:2014:BMD

[1503] E. J. Billman, L. D. Whitman, R. K. Schroeder, C. S. Sharpe, D. L. G. Noakes, and C. B. Schreck. Body morphology differs in wild juvenile Chinook salmon Oncorhynchus tshawytscha that express different migratory phenotypes in the Willamette River, Oregon, U.S.A. Journal of Fish Biology, 85(4):1097–1110, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zydlewski:2014:SCC

[1504] J. Zydlewski, G. Zydlewski, B. Kennedy, and W. Gale. Smolting in coastal cutthroat trout Onchorhynchus clarkii clarkii. Journal of Fish Biology, 85(4):1111–1130, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urke:2014:LTH

[1505] H. A. Urke, J. V. Arnekleiv, T. O. Nilsen, K. J. Nilssen, L. Rønning, J. B. Ulvund, and T. Kristensen. Long-term hypo-osmoregulatory capacity in downstream migrating Atlantic salmon Salmo salar L. smolts. Journal of Fish Biology, 85(4):1131–1144, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnason:2014:LTR

[1506] T. Árnason, S. Gunnarsson, A. K. Imsland, H. Thorarensen, H. Smáradóttir, A. Steinarsson, A. Gústavsson, M. Johansson, and B. Th. Björnsson. Long-term rearing of Arctic charr Salvelinus alpinus under different salinity regimes at constant temperature. Journal of Fish Biology, 85(4): 1145–1162, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Handeland:2014:OAS

[1507] S. O. Handeland, A. K. Imsland, T. O. Nilsen, L. O. E. Ebbesson, C. D. Hosfeld, C. Pedrosa, H. Toften, and S. O. Stefansson. Osmoregulation in Atlantic salmon Salmo salar smolts transferred to seawater at different temperatures. Journal of Fish Biology, 85(4):1163–1176, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hammenstig:2014:ERD

[1508] D. Hammenstig, E. Sandblom, M. Axelsson, and J. I. Johnsson. Effects of rearing density and dietary fat content on burst-swim performance and oxygen transport capacity in juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 85(4):1177–1191, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norrgaard:2014:EFQ

[1509] J. R. Norrgård, E. Bergman, L. A. Greenberg, and M. Schmitz. Effects of feed quality and quantity on growth, early maturation and smolt development in hatchery-reared landlocked Atlantic salmon Salmo salar. Journal of Fish Biology, 85(4):1192–1210, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gunnarsson:2014:ESD

[1510] S. Gunnarsson, M. Johansson, A. Gústavsson, T. Árnason, J. Árnason, H. Smáradóttir, B. Th. Björnsson, H. Thorarensen, and A. K. Imsland. Effects of short-day treatment on long-term growth performance and maturation of farmed Arctic charr Salvelinus alpinus reared in brackish water. Journal of Fish Biology, 85(4):1211–1226, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundh:2014:DII

[1511] H. Sundh, T. O. Nilsen, J. Lindström, L. Hasselberg-Frank, S. O. Stefansson, S. D. McCormick, and K. Sundell. Development of intestinal ion-transporting mechanisms during smoltification and seawater acclimation in Atlantic salmon Salmo salar. Journal of Fish Biology, 85(4):

1227–1252, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munakata:2014:ESD

[1512] A. Munakata, G. Miura, and H. Matsuda. Evaluation of seasonal and daily changes of plasma thyroxine and cortisol levels in wild masu salmon Oncorhynchus masou, sampled by a Japanese fishing method. Journal of Fish Biology, 85(4):1253–1262, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yada:2014:SCG

[1513] T. Yada, K. Miyamoto, G. Miura, and A. Munakata. Seasonal changes in gene expression of corticoid receptors in anadromous and nonanadromous strains of rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 85(4):1263–1278, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Einarsdottir:2014:PGH

[1514] I. E. Einarsdóttir, N. Gong, E. Jönsson, H. Sundh, L. Hasselberg-Frank, T. O. Nilsen, S. O. Stefansson, K. Sundell, and B. Th. Björnsson. Plasma growth hormone-binding protein levels in Atlantic salmon Salmo salar during smoltification and seawater transfer. Journal of Fish Biology, 85 (4):1279–1296, October 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luzzatto:2014:CGE

[1515] D. C. Luzzatto, M. L. Estalles, and J. M. Díaz De Astarloa. Comment on ‘Genetic evidence and new morphometric data as essential tools to identify the Patagonian seahorse Hippocampus patagonicus (Pisces, Syngnathidae) by González et al. (2014)’. Journal of Fish Biology, 85(5): 1297–1299, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez:2014:RLA

[1516] R. González, P. Dinghi, C. Corio, A. Medina, M. Maggioni, L. Storero, and A. Gosztonyi. Reply to Luzzatto et al. (2014): “Comment on ‘Genetic evidence and new morphometric data as essential tools to identify the Patagonian seahorse Hippocampus patagonicus (Pisces, Syngnathidae), González et al. (2014)”’. Journal of Fish Biology, 85(5):1300–1302, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fricke:2014:CON

[1517] R. Fricke, L. A. Jawad, and J. M. Al-Mamry. Callionymus omanensis, a new species of dragonet from Oman, north-western Indian Ocean (Teleostei: Callionymidae). Journal of Fish Biology, 85(5):1303–1319, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Veale:2014:CIT

[1518] L. Veale, J. R. Tweedley, K. R. Clarke, C. S. Hallett, and I. C. Potter. Characteristics of the ichthyofauna of a temperate microtidal estuary with a reverse salinity gradient, including inter-decadal comparisons. Journal of Fish Biology, 85(5):1320–1354, November 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Esbaugh:2014:ESA

[1519] A. J. Esbaugh, T. Kristensen, H. Takle, and M. Grosell. The effects of sustained aerobic swimming on osmoregulatory pathways in Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 85(5):1355–1368, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2014:CPP

[1520] X. X. Yang, X. N. Hou, B. Xu, X. Hao, G. J. Jiang, and T. J. Fan. Cell-penetrating peptide delivery of biologically active oct4 protein into cultured Takifugu rubripes spermary cells. Journal of Fish Biology, 85(5): 1369–1380, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2014:DMO

[1521] Z. He, Y. Li, Y. Wu, S. Shi, C. Sun, Q. Deng, J. Xie, T. Wang, W. Zhang, and L. Zhang. Differentiation and morphogenesis of the ovary and expression of gonadal development-related genes in the protogynous hermaphroditic ricefield eel Monopterus albus. Journal of Fish Biology, 85(5):1381–1394, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solovyev:2014:FHO

[1522] M. M. Solovyev, E. N. Kashinskaya, G. I. Izvekova, E. Gisbert, and V. V. Glupov. Feeding habits and ontogenic changes in digestive enzyme patterns in five freshwater teleosts. Journal of Fish Biology, 85(5):1395– 1412, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harasti:2014:OSB

[1523] D. Harasti, K. Martin-Smith, and W. Gladstone. Ontogenetic and sexbased differences in habitat preferences and site fidelity of White’s seahorse Hippocampus whitei. Journal of Fish Biology, 85(5):1413–1428, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bond:2014:DGE

[1524] M. H. Bond, B. R. Beckman, L. Rohrbach, and T. P. Quinn. Differential growth in estuarine and freshwater habitats indicated by plasma IGF1 concentrations and otolith chemistry in Dolly Varden Salvelinus malma. Journal of Fish Biology, 85(5):1429–1445, November 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos:2014:TNH

[1525] J. A. A. Ramos, M. Barletta, D. V. Dantas, A. R. A. Lima, and M. F. Costa. Trophic niche and habitat shifts of sympatric Gerreidae. Journal of Fish Biology, 85(5):1446–1469, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wall:2014:TSM

[1526] C. C. Wall, P. Simard, M. Lindemuth, C. Lembke, D. F. Naar, C. Hu, B. B. Barnes, F. E. Muller-Karger, and D. A. Mann. Temporal and spatial mapping of red grouper Epinephelus morio sound production. Journal of Fish Biology, 85(5):1470–1488, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vasconcelos:2014:FRG

[1527] L. P. Vasconcelos, D. C. Alves, and L. C. Gomes. Fish reproductive guilds downstream of dams. Journal of Fish Biology, 85(5):1489–1506, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bourret:2014:UOC

[1528] S. L. Bourret, B. P. Kennedy, C. C. Caudill, and P. M. Chittaro. Using otolith chemical and structural analysis to investigate reservoir habitat use by juvenile Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 85(5):1507–1525, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saint-Jean:2014:ECN

[1529] S. Rodriguez Saint-Jean, C. González, M. Monrás, A. Romero, N. Ballesteros, R. Enríquez, and S. Perez-Prieto. Establishment and characterization of a new cell line (SSP-9) derived from Atlantic salmon Salmo salar that expresses type I ifn. Journal of Fish Biology, 85(5):1526–1545, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mincarone:2014:DSB

[1530] M. M. Mincarone, F. Di Dario, and P. A. S. Costa. Deepsea bigscales, pricklefishes, gibberfishes and whalefishes (Teleostei: Stephanoberycoidei) off Brazil: new records, range extensions for the south-western Atlantic Ocean and remarks on the taxonomy of Poromitra. Journal of Fish Biology, 85(5):1546–1570, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacerda:2014:TPI

[1531] C. H. F. Lacerda, M. Barletta, and D. V. Dantas. Temporal patterns in the intertidal faunal community at the mouth of a tropical estuary. Journal of Fish Biology, 85(5):1571–1602, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kekalainen:2014:IAB

[1532] J. Kekäläinen, T. Podgorniak, T. Puolakka, P. Hyvärinen, and A. Vainikka. Individually assessed boldness predicts Perca fluviatilis behaviour in shoals, but is not associated with the capture order or angling method. Journal of Fish Biology, 85(5):1603–1616, November 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bustamante:2014:AFE

[1533] C. Bustamante, C. Vargas-Caro, and M. B. Bennett. Not all fish are equal: functional biodiversity of cartilaginous fishes (Elasmobranchii and Holocephali) in Chile. Journal of Fish Biology, 85(5):1617–1633, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watson:2014:DCN

[1534] A. C. Watson, L. A. Siemann, and R. T. Hanlon. Dynamic camouflage by Nassau groupers Epinephelus striatus on a Caribbean coral reef. Journal of Fish Biology, 85(5):1634–1649, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rieucau:2014:SDA

[1535] G. Rieucau, A. De Robertis, K. M. Boswell, and N. O. Handegard. School density affects the strength of collective avoidance responses in wildcaught Atlantic herring Clupea harengus: a simulated predator encounter experiment. Journal of Fish Biology, 85(5):1650–1664, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Losee:2014:GCJ

[1536] J. P. Losee, J. Fisher, D. J. Teel, R. E. Baldwin, D. J. Marcogliese, and K. C. Jacobson. Growth and condition of juvenile Coho salmon Oncorhynchus kisutch relate positively to species richness of trophically transmitted parasites. Journal of Fish Biology, 85(5):1665–1681, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Utsunomia:2014:GDA

[1537] R. Utsunomia, J. C. Pansonato Alves, L. R. S. Paiva, G. J. Costa Silva, C. Oliveira, L. A. C. Bertollo, and F. Foresti. Genetic differentiation among distinct karyomorphs of the wolf fish Hoplias malabaricus species complex (Characiformes, Erythrinidae) and report of unusual hybridization with natural triploidy. Journal of Fish Biology, 85(5):1682–1692, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monroy:2014:TIB

[1538] M. Monroy, A. Maceda-Veiga, N. Caiola, and A. De Sostoa. Trophic interactions between native and introduced fish species in a littoral fish community. Journal of Fish Biology, 85(5):1693–1706, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hopperdietzel:2014:GMH

[1539] C. Hopperdietzel, R. M. Hirschberg, H. Hünigen, J. Wolter, K. Richardson, and J. Plendl. Gross morphology and histology of the alimentary tract of the convict cichlid Amatitlania nigrofasciata. Journal of Fish Biology, 85(5):1707–1725, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feldheim:2014:ICM

[1540] K. A. Feldheim, B. R. Kreiser, B. Schmidt, D. D. Duvernell, and J. F. Schaefer. Isolation and characterization of microsatellite loci for the blackstripe topminnow Fundulus notatus and their variability in two closely related species. Journal of Fish Biology, 85(5):1726–1732, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanSteenberge:2014:LRC

[1541] M. Van Steenberge, L. Gajdzik, A. Chilala, J. Snoeks, and E. Vreven. Labeo rosae (Cypriniformes: Cyprinidae) in the Congo basin: a relict distribution or a historical introduction? Journal of Fish Biology, 85(5): 1733–1738, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hernandez:2014:EMP

[1542] S. Hernández, C. Duffy, M. P. Francis, and P. A. Ritchie. Evidence for multiple paternity in the school shark Galeorhinus galeus found in New Zealand waters. Journal of Fish Biology, 85(5):1739–1745, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2014:VSA

[1543] P. G. Jellyman and J. S. Harding. Variable survival across low pH gradients in freshwater fish species. Journal of Fish Biology, 85(5):1746–1752, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hsu:2014:FRT

[1544] H. H. Hsu, C. Y. Lin, and S. J. Joung. The first record, tagging and release of a neonatal whale shark Rhincodon typus in Taiwan. Journal of Fish Biology, 85(5):1753–1756, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawabata:2014:ESP

[1545] Y. Kawabata, G. N. Nishihara, T. Yamaguchi, T. Takebe, K. Teruya, T. Sato, and K. Soyano. The effect of spine postures on the hydrodynamic drag in Epinephelus ongus larvae. Journal of Fish Biology, 85(5):1757– 1765, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lavretsky:2014:DMH

[1546] P. Lavretsky, R. S. Schwartz, M. R. Baerwald, and B. May. Developing major histocompatibility markers in a species of concern: the Sacramento perch Archoplites interruptus. Journal of Fish Biology, 85(5):1766–1776, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walleser:2014:CMU

[1547] L. R. Walleser, D. R. Howard, M. B. Sandheinrich, M. P. Gaikowski, and J. J. Amberg. Confocal microscopy as a useful approach to describe gill rakers of Asian species of carp and native filter-feeding fishes of the upper Mississippi River system. Journal of Fish Biology, 85(5):1777– 1784, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCallum:2014:SPT

[1548] E. S. McCallum, P. M. Capelle, and S. Balshine. Seasonal plasticity in telencephalon mass of a benthic fish. Journal of Fish Biology, 85(5): 1785–1792, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:2014:DSB

[1549] A. M. Pereira, I. Doadrio, J. I. Robalo, and V. C. Almada. Different stocks of brook lamprey in Spain and their origin from Lampetra fluviatilis at two distinct times and places. Journal of Fish Biology, 85(5): 1793–1798, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bonar:2014:BRE

[1550] Scott A. Bonar. Book review: Ecology of North American freshwater fishes. Journal of Fish Biology, 85(5):1799–1800, November 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2014:CC

[1551] Stephen J. Lockwood. Coastal conservation. Journal of Fish Biology, 85(5):1801–1802, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2014:ELS

[1552] Ian J. Winfield. The extreme life of the sea. Journal of Fish Biology, 85(5):1803, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2014:BRR

[1553] Laith A. Jawad. The biology of reefs and reef organisms. Journal of Fish Biology, 85(5):1804–1805, November 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzen:2014:UME

[1554] K. Lorenzen. Understanding and managing enhancements: why fisheries scientists should care. Journal of Fish Biology, 85(6):1807–1829, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garlock:2014:EHR

[1555] T. M. Garlock, C. T. Monk, K. Lorenzen, M. D. Matthews, and C. M. St Mary. Effects of hatchery rearing on Florida largemouth bass Micropterus floridanus resource allocation and performance under semi-natural conditions. Journal of Fish Biology, 85(6):1830–1842, December 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arlinghaus:2014:SSP

[1556] R. Arlinghaus, B. Beardmore, C. Riepe, J. Meyerhoff, and T. Pagel. Species-specific preferences of German recreational anglers for freshwater fishing experiences, with emphasis on the intrinsic utilities of fish stocking and wild fishes. Journal of Fish Biology, 85(6):1843–1867, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camp:2014:SEA

[1557] E. V. Camp, K. Lorenzen, R. N. M. Ahrens, and M. S. Allen. Stock enhancement to address multiple recreational fisheries objectives: an integrated model applied to red drum Sciaenops ocellatus in Florida. Journal of Fish Biology, 85(6):1868–1889, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellender:2014:MCA

[1558] B. R. Ellender, D. J. Woodford, O. L. F. Weyl, and I. G. Cowx. Managing conflicts arising from fisheries enhancements based on non-native fishes in southern Africa. Journal of Fish Biology, 85(6):1890–1906, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchings:2014:USU

[1559] J. A. Hutchings. Unintentional selection, unanticipated insights: introductions, stocking and the evolutionary ecology of fishes. Journal of Fish Biology, 85(6):1907–1926, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stringwell:2014:MPM

[1560] R. Stringwell, A. Lock, C. J. Stutchbury, E. Baggett, J. Taylor, P. J. Gough, and C. Garcia de Leaniz. Maladaptation and phenotypic mismatch in hatchery-reared Atlantic salmon Salmo salar released in the wild. Journal of Fish Biology, 85(6):1927–1945, December 2014. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsson:2014:EEB

[1561] J. I. Johnsson, S. Brockmark, and J. Näslund. Environmental effects on behavioural development consequences for fitness of captive-reared fishes in the wild. Journal of Fish Biology, 85(6):1946–1971, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:2014:SSR

[1562] L. J. Roberts, J. Taylor, P. J. Gough, D. W. Forman, and C. Garcia de Leaniz. Silver spoons in the rough: can environmental enrichment improve survival of hatchery Atlantic salmon Salmo salar in the wild? Journal of Fish Biology, 85(6):1972–1991, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2014:DRF

[1563] J. G. Davidsen, M. Daverdin, A. D. Sjursen, L. Rønning, J. V. Arnekleiv, and J. I. Koksvik. Does reduced feeding prior to release improve the marine migration of hatchery brown trout Salmo trutta smolts? Journal of Fish Biology, 85(6):1992–2002, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibanez:2014:IDF

[1564] A. L. Ibáñez, M. Pérez-Ramírez, and J. L. García-Calderón. Institutional development of freshwater fish stocking in Mexico. Journal of Fish Biology, 85(6):2003–2009, December 2014. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawson:2015:NLH

[1565] J. M. Lawson, S. J. Foster, A. C. O. Lim, V. C. Chong, and A. C. J. Vincent. Novel life-history data for threatened seahorses provide insight into fishery effects. Journal of Fish Biology, 86(1):1–15, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ryan:2015:ERC

[1566] D. K. Ryan and M. Kelly-Quinn. Effects of riparian canopy cover on salmonid diet and prey selectivity in low nutrient streams. Journal of Fish Biology, 86(1):16–31, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2015:NFC

[1567] X. Fu, N. Li, Y. Lai, X. Luo, Y. Wang, C. Shi, Z. Huang, S. Wu, and J. Su. A novel fish cell line derived from the brain of Chinese perch Siniperca chuatsi: development and characterization. Journal of Fish Biology, 86(1):32–45, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2015:ECO

[1568] A. Sun, T. Z. Wang, N. Wang, X. F. Liu, Z. X. Sha, and S. L. Chen. Establishment and characterization of an ovarian cell line from half-smooth tongue sole Cynoglossus semilaevis. Journal of Fish Biology, 86(1):46– 59, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berrebi:2015:TBT

[1569] P. Berrebi. Three brown trout Salmo trutta lineages in Corsica described through allozyme variation. Journal of Fish Biology, 86(1):60–73, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cao:2015:IEL

[1570] X. L. Cao, J. J. Chen, Y. Cao, G. X. Nie, Q. Y. Wan, L. F. Wang, and J. G. Su. Identification and expression of the laboratory of genetics and physiology 2 gene in common carp Cyprinus carpio. Journal of Fish Biology, 86(1):74–91, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:2015:WDA

[1571] N. A. Hvidsten, O. H. Diserud, A. J. Jensen, J. G. Jensås, B. O. Johnsen, and O. Ugedal. Water discharge affects Atlantic salmon Salmo salar smolt production: a 27 year study in the River Orkla, Norway. Journal of Fish Biology, 86(1):92–104, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penney:2015:FAP

[1572] Z. L. Penney and C. M. Moffitt. Fatty-acid profiles of white muscle and liver in stream-maturing steelhead trout Oncorhynchus mykiss from early migration to kelt emigration. Journal of Fish Biology, 86(1):105– 120, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulas:2015:DFB

[1573] A. Mulas, A. Bellodi, R. Cannas, Al. Cau, D. Cuccu, M. F. Marongiu, C. Porcu, and M. C. Follesa. Diet and feeding behaviour of longnosed skate Dipturus oxyrinchus. Journal of Fish Biology, 86(1):121–138, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yokozawa:2015:OCF

[1574] N. Yokozawa, O. Nakamura, E. Saito, and S. Tsutsui. Ovarian cavity fluid of the viviparous surfperch Neoditrema ransonnetii suppresses the spontaneous cytotoxic activity of head–kidney leucocytes against xenogeneic targets. Journal of Fish Biology, 86(1):139–147, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romany:2015:MIS

[1575] J. S. Romany, A. E. McElroy, and M. D. Fast. Measures of immune system status in young-of-the-year winter flounder Pseudopleuronectes americanus. Journal of Fish Biology, 86(1):148–161, January 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuz-Sulub:2015:SAT

[1576] A. Tuz-Sulub and T. Brulé. Spawning aggregations of three protogynous groupers in the southern Gulf of Mexico. Journal of Fish Biology, 86 (1):162–185, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2015:PSS

[1577] G. Garcia, S. Pereyra, V. Gutierrez, S. Oviedo, P. Miller, and A. Domingo. Population structure of Squatina guggenheim (Squatiniformes, Squatinidae) from the south-western Atlantic Ocean. Journal of Fish Biology, 86(1):186–202, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2015:VDM

[1578] G. Plaza and F. Cerna. Validation of daily microincrement deposition in otoliths of juvenile and adult Peruvian anchovy Engraulis ringens. Journal of Fish Biology, 86(1):203–216, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welsh:2015:ISH

[1579] D. P. Welsh and R. C. Fuller. Influence of sex and habitat on the size and shape of anal and dorsal fins of the blackstripe topminnow Fundulus notatus. Journal of Fish Biology, 86(1):217–227, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clausen:2015:ESD

[1580] L. A. W. Clausen, K.-J. Stæhr, A. Rindorf, and H. Mosegaard. Effect of spatial differences in growth on distribution of seasonally co-occurring herring Clupea harengus stocks. Journal of Fish Biology, 86(1):228–247, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2015:ETF

[1581] G.-L. Li, M.-Z. Zhang, S.-P. Deng, H.-P. Chen, and C.-H. Zhu. Effects of temperature and fish oil supplementation on ovarian development and fgeneoxl 2 mRNA expression in spotted scat Scatophagus argus. Journal of Fish Biology, 86(1):248–260, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keyl:2015:SSD

[1582] F. Keyl, A. J. Kempf, and A. F. Sell. Sexual size dimorphism in three North Sea gadoids. Journal of Fish Biology, 86(1):261–275, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katinic:2015:TRP

[1583] P. J. Katinic, D. A. Patterson, and R. C. Ydenberg. Thermal regime, predation danger and the early marine exit of sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 86(1):276–287, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedges:2015:HRP

[1584] K. J. Hedges and M. V. Abrahams. Hypoxic refuges, predator–prey interactions and habitat selection by fishes. Journal of Fish Biology, 86 (1):288–303, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hafs:2015:DTC

[1585] A. W. Hafs and K. J. Hartman. Development of temperature correction equations for bioelectrical impedance analysis models for brook trout Salvelinus fontinalis. Journal of Fish Biology, 86(1):304–316, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaston:2015:MVB

[1586] K. A. Gaston and T. E. Lauer. Morphometric variation in bluegill Lepomis macrochirus and green sunfish Lepomis cyanellus in lentic and lotic systems. Journal of Fish Biology, 86(1):317–332, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drew:2015:AGM

[1587] M. Drew, W. T. White, Dharmadi, A. V. Harry, and C. Huveneers. Age, growth and maturity of the pelagic thresher Alopias pelagicus and the scalloped hammerhead Sphyrna lewini. Journal of Fish Biology, 86 (1):333–354, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fey:2015:SGR

[1588] D. P. Fey. Size and growth rate differences of larval Baltic sprat Sprattus sprattus collected with bongo and MIK nets. Journal of Fish Biology, 86 (1):355–359, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrkjedal:2015:ASA

[1589] I. Byrkjedal, J. S. Christiansen, O. V. Karamushko, G. Langhelle, and A. Lynghammar. Arctic skate Amblyraja hyperborea preys on remarkably large glacial eelpouts Lycodes frigidus. Journal of Fish Biology, 86(1): 360–364, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Butler:2015:ABT

[1590] C. M. Butler, J. M. Logan, J. M. Provaznik, E. R. Hoffmayer, M. D. Staudinger, J. M. Quattro, M. A. Roberts, G. W. Ingram, Jr., A. G. Pollack, and M. E. Lutcavage. Atlantic bluefin tuna Thunnus thynnus feeding ecology in the northern Gulf of Mexico: a preliminary description of diet from the western Atlantic spawning grounds. Journal of Fish Biology, 86(1):365–374, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:2015:DAG

[1591] A. M. Ferguson, E. S. Harvey, M. J. Rees, and N. A. Knott. Does the abundance of girellids and kyphosids correlate with cover of the palatable green algae, Ulva spp.? A test on temperate rocky intertidal reefs. Journal of Fish Biology, 86(1):375–384, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delorenzo:2015:ADT

[1592] D. M. Delorenzo, D. M. Bethea, and J. K. Carlson. An assessment of the diet and trophic level of Atlantic sharpnose shark Rhizoprionodon terraenovae. Journal of Fish Biology, 86(1):385–391, January 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Furey:2015:EFS

[1593] N. B. Furey, S. G. Hinch, A. G. Lotto, and D. A. Beauchamp. Extensive feeding on sockeye salmon Oncorhynchus nerka smolts by bull trout Salvelinus confluentus during initial outmigration into a small, unregulated and inland British Columbia river. Journal of Fish Biology, 86 (1):392–401, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhivotovsky:2015:EIC

[1594] L. A. Zhivotovsky, S. Yu. Kordicheva, E. G. Shaikhaev, G. A. Rubtsova, K. I. Afanasiev, M. V. Shitova, S. A. Fuller, G. O. Shaikhaev, and A. J. Gharrett. Efficiency of the inbreeding coefficient f and other estimators in detecting null alleles, as revealed by empirical data of locus oke3 across 65 populations of chum salmon Oncorhynchus keta. Journal of Fish Biology, 86(1):402–408, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kikko:2015:DPL

[1595] T. Kikko, D. Ishizaki, K. Ninomiya, Y. Kai, and Y. Fujioka. Diel patterns of larval drift of honmoroko Gnathopogon caerulescens in an inlet of Ibanaiko Lagoon, Lake Biwa, Japan. Journal of Fish Biology, 86(1):409– 415, January 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2015:OSG

[1596] Y. W. Li, X. Li, Z. Wang, Z. Q. Mo, X. M. Dan, X. C. Luo, and A. X. Li. Orange-spotted grouper Epinephelus coioides Tak1: molecular identification, expression analysis and functional study. Journal of Fish Biology, 86(2):417–430, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hu:2015:TLR

[1597] G.-B. Hu, X.-P. Li, D.-H. Liu, Q.-M. Liu, and S.-C. Zhang. A toll-like receptor 3 homologue that is up-regulated by poly I:C and DNA virus in turbot Scophthalmus maximus. Journal of Fish Biology, 86(2):431–447, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joh:2015:COM

[1598] M. Joh, T. Matsuda, and A. Miyazono. Common otolith microstructure related to key early life-history events in flatfishes identified in the larvae and juveniles of cresthead flounder Pseudopleuronectes schrenki. Journal of Fish Biology, 86(2):448–462, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stocks:2015:IPT

[1599] J. R. Stocks, C. A. Gray, and M. D. Taylor. Intra-population trends in the maturation and reproduction of a temperate marine herbivore Girella elevata across latitudinal clines. Journal of Fish Biology, 86(2):463–483, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dharmadi:2015:SCA

[1600] Dharmadi, Fahmi, and W. T. White. Species composition and aspects of the biology of Orectolobiformes from Indonesian waters. Journal of Fish Biology, 86(2):484–492, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egge:2015:MMV

[1601] J. J. D. Egge, P. W. Nicholson, and A. W. Stark. Morphological and molecular variation in the least madtom Noturus hildebrandi (Siluriformes: Ictaluridae), a Mississippi Embayment endemic: evidence for a cryptic lineage in the Hatchie River. Journal of Fish Biology, 86(2): 493–526, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2015:PCH

[1602] T. L. A. Costa, J. A. Thayer, and L. F. Mendes. Population characteristics, habitat and diet of a recently discovered stingray Dasyatis marianae: implications for conservation. Journal of Fish Biology, 86(2): 527–543, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horka:2015:RTS

[1603] P. Horká, P. Horký, T. Randák, J. Turek, K. Rylková, and O. Slavík. Radio-telemetry shows differences in the behaviour of wild and hatcheryreared European grayling Thymallus thymallus in response to environmental variables. Journal of Fish Biology, 86(2):544–557, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:2015:AGR

[1604] H. A. Thompson and T. P. Simon. Age and growth of round goby Neogobius melanostomus associated with depth and habitat in the western basin of Lake Erie. Journal of Fish Biology, 86(2):558–574, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajasilta:2015:FBH

[1605] M. Rajasilta, J. Eklund, J. Hänninen, I. Vuorinen, and P. Laine. Female Baltic herring Clupea harengus allocate resources from growth to reproduction in poor feeding conditions. Journal of Fish Biology, 86 (2):575–591, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feng:2015:MCP

[1606] J. B. Feng, S. K. Liu, R. J. Wang, J. R. Zhang, X. L. Wang, L. Kaltenboeck, J. L. Li, and Z. J. Liu. Molecular characterization, phylogenetic analysis and expression profiling of myoglobin and cytoglobin genes in response to heat stress in channel catfish Ictalurus punctatus. Journal of Fish Biology, 86(2):592–604, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yan:2015:EFF

[1607] G. J. Yan, X. K. He, Z. D. Cao, and S. J. Fu. Effects of fasting and feeding on the fast-start swimming performance of southern catfish Silurus meridionalis. Journal of Fish Biology, 86(2):605–614, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2015:PAT

[1608] L. Li, Y. Shi, R. Wang, T. Huang, W. Liang, H. Luo, X. Gan, W. Huang, J. Li, A. Lei, and M. Chen. Proteomic analysis of tilapia Oreochromis niloticus Streptococcus agalactiae strains with different genotypes and serotypes. Journal of Fish Biology, 86(2):615–636, February 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rickborn:2015:LHT

[1609] A. J. Rickborn and P. M. Buston. Life-history transitions of the coral-reef fish Elacatinus lori. Journal of Fish Biology, 86(2):637–650, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karjalainen:2015:PSC

[1610] J. Karjalainen, T. J. Ruokonen, T. J. MarjomäKi, A. Martikainen, M. Pursiainen, J. Sarvala, M. Tarvainen, and A.-M. Ventelä. Predation by signal crayfish Pacifastacus leniusculus on fish eggs and its consequences for coregonid recruitment. Journal of Fish Biology, 86(2): 651–667, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cardenas-Palomo:2015:DFH

[1611] N. Cárdenas-Palomo, J. Herrera-Silveira, I. Velázquez-Abunader, O. Reyes, and U. Ordoñez. Distribution and feeding habitat characterization of whale sharks Rhincodon typus in a protected area in the north Caribbean Sea. Journal of Fish Biology, 86(2):668–686, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Isbert:2015:MPC

[1612] W. Isbert, C. Rodríguez-Cabello, I. Frutos, I. Preciado, F. E. Montero, and A. Pérez del Olmo. Metazoan parasite communities and diet of the velvet belly lantern shark Etmopterus spinax (Squaliformes: Etmopteridae): a comparison of two deep-sea ecosystems. Journal of Fish Biology, 86(2):687–706, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dantas:2015:FES

[1613] D. V. Dantas, M. Barletta, and M. F. Costa. Feeding ecology and seasonal diet overlap between Stellifer brasiliensis and Stellifer stellifer in a tropical estuarine ecocline. Journal of Fish Biology, 86(2):707–733, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rochowski:2015:RBG

[1614] B. E. A. Rochowski, K. J. Graham, R. W. Day, and T. I. Walker. Reproductive biology of the greeneye spurdog Squalus chloroculus (Squaliformes, Squalidae). Journal of Fish Biology, 86(2):734–754, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mesa:2015:AMI

[1615] M. G. Mesa and B. P. Rose. An assessment of morphometric indices, blood chemistry variables and an energy meter as indicators of the whole body lipid content in Micropterus dolomieu, Sander vitreus and Ictalurus punctatus. Journal of Fish Biology, 86(2):755–764, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boivin:2015:HDS

[1616] B. Boivin, M. Castonguay, C. Audet, S. A. Pavey, M. Dionne, and L. Bernatchez. How does salinity influence habitat selection and growth in juvenile American eels Anguilla rostrata? Journal of Fish Biology, 86 (2):765–784, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nobrega:2015:STV

[1617] M. F. Nóbrega, P. G. Kinas, R. Lessa, and E. Ferrandis. Spatial and temporal variation in artisanal catches of dolphinfish Coryphaena hippurus off north-eastern Brazil. Journal of Fish Biology, 86(2):785–804, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baran:2015:ACP

[1618] R. Baran, J. Kubecka, M. Kubín, B. Lojkásek, T. Mrkvička, D. Ricard, and M. Rulík. Abundance of Cottus poecilopus is influenced by O_{2} saturation, food density and Salmo trutta in three tributaries of the Rožnovská Bečva River, Czech Republic. Journal of Fish Biology, 86(2):805–811, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulgar:2015:EUR

[1619] J. Pulgar, P. Lagos, D. Maturana, M. Valdés, M. Aldana, and V. M. Pulgar. Effect of UV radiation on habitat selection by Girella laevifrons and Graus nigra (Kyphosidae). Journal of Fish Biology, 86(2):812–821, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Favero:2015:UNT

[1620] J. M. Favero, M. Katsuragawa, M. L. Zani-Teixeira, and J. T. Turner. Using new tools to identify eggs of Engraulis anchoita (Clupeiformes, Engraulidae). Journal of Fish Biology, 86(2):822–826, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yetsko:2015:ESS

[1621] K. Yetsko and G. Sancho. The effects of salinity on swimming performance of two estuarine fishes, Fundulus heteroclitus and Fundulus majalis. Journal of Fish Biology, 86(2):827–833, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wienerroither:2015:NRL

[1622] R. M. Wienerroither, O. Bjelland, L. Bachmann, and C. Junge. Northernmost record of the little gulper shark Centrophorus uyato in the northeastern Atlantic Ocean, with taxonomical notes on Centrophorus zeehaani. Journal of Fish Biology, 86(2):834–844, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:2015:VHS

[1623] N. N. Chang, E. Y. Liu, Y. C. Liao, and J. C. Shiao. Vertical habitat shift of viviparous and oviparous deep-sea cusk eels revealed by otolith microstructure and stable-isotope composition. Journal of Fish Biology, 86(2):845–853, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roch:2015:GDC

[1624] S. Roch, J. Behrmann-Godel, and A. Brinker. Genetically distinct colour morphs of European perch Perca fluviatilis in Lake Constance differ in susceptibility to macroparasites. Journal of Fish Biology, 86(2):854–863, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2015:R

[1625] Anonymous. Referees 2014. Journal of Fish Biology, 86(2):864–870, February 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OMalley:2015:IIR

[1626] K. G. O’Malley, B. Perales, and A. C. Whitcomb. Influence of immunerelevant genotype on the reproductive success of a salmonid alternative mating strategy. Journal of Fish Biology, 86(3):871–881, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gholami:2015:GCP

[1627] Z. Gholami, H. R. Esmaeili, D. Erpenbeck, and B. Reichenbacher. Genetic connectivity and phenotypic plasticity in the cyprinodont Aphanius farsicus from the Maharlu Basin, south-western Iran. Journal of Fish Biology, 86(3):882–906, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pattoura:2015:AEJ

[1628] P. Pattoura, E. Lefkaditou, and P. Megalofonou. Age estimation of juvenile European hake Merluccius merluccius based on otolith microstructure analysis: a slow or fast growth pattern? Journal of Fish Biology, 86(3):907–923, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beentjes:2015:GPA

[1629] M. P. Beentjes and D. J. Jellyman. Growth patterns and age validation from otolith ring deposition in New Zealand longfin eels Anguilla dieffenbachii recaptured after 10 years at large. Journal of Fish Biology, 86(3):924–939, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ho:2015:NAE

[1630] H.-C. Ho and M. McGrouther. A new anglerfish from eastern Australia and New Caledonia (Lophiiformes: Chaunacidae: Chaunacops), with new data and submersible observation of Chaunacops melanostomus. Journal of Fish Biology, 86(3):940–951, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fost:2015:PPA

[1631] B. A. Fost and C. P. Ferreri. pH preference and avoidance responses of adult brook trout Salvelinus fontinalis and brown trout Salmo trutta. Journal of Fish Biology, 86(3):952–966, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchheister:2015:DTG

[1632] A. Buchheister and R. J. Latour. Diets and trophic-guild structure of a diverse fish assemblage in Chesapeake Bay, U.S.A. Journal of Fish Biology, 86(3):967–992, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez-Rodriguez:2015:CHB

[1633] R. Pérez-Rodríguez, O. Domínguez-Domínguez, I. Doadrio, E. Cuevas-García, and G. Pérez-Ponce de León. Comparative historical biogeography of three groups of Nearctic freshwater fishes across central Mexico. Journal of Fish Biology, 86(3):993–1015, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2015:BSL

[1634] L. Jacobsen, S. Berg, H. Baktoft, and C. Skov. Behavioural strategy of large perch Perca fluviatilis varies between a mesotrophic and a hypereutrophic lake. Journal of Fish Biology, 86(3):1016–1029, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauli:2015:OSB

[1635] B. Diaz Pauli, M. Wiech, M. Heino, and A. C. Utne-Palm. Opposite selection on behavioural types by active and passive fishing gears in a simulated guppy Poecilia reticulata fishery. Journal of Fish Biology, 86 (3):1030–1045, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulson:2015:DVW

[1636] P. G. Coulson, M. E. Platell, K. R. Clarke, and I. C. Potter. Dietary variations within a family of ambush predators (Platycephalidae) occupying different habitats and environments in the same geographical region. Journal of Fish Biology, 86(3):1046–1077, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rey:2015:CEA

[1637] J. Rey, L. Fernandez-Peralta, L. F. Quintanilla, M. Hidalgo, C. Presas, F. Salmeron, and M. A. Puerto. Contrasting energy allocation strategies of two sympatric Merluccius species in an upwelling system. Journal of Fish Biology, 86(3):1078–1097, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pei:2015:CFT

[1638] Y. Y. Pei, R. Huang, Y. M. Li, L. J. Liao, Z. Y. Zhu, and Y. P. Wang. Characterizations of four toll-like receptor 4s in grass carp Ctenopharyngodon idellus and their response to grass carp reovirus infection and lipopolysaccharide stimulation. Journal of Fish Biology, 86(3):1098– 1108, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silveira:2015:DDS

[1639] M. M. Silveira, J. J. Oliveira, and A. C. Luchiari. Dusky damselfish Stegastes fuscus relational learning: evidences from associative and spatial tasks. Journal of Fish Biology, 86(3):1109–1120, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hatano:2015:OME

[1640] M. Hatano, T. Abe, T. Wada, and H. Munehara. Ontogenetic metamorphosis and extreme sexual dimorphism in lumpsuckers: Eumicrotremus asperrimus, Cyclopteropsis bergi and Cyclopteropsis lindbergi, may be synonymous. Journal of Fish Biology, 86(3):1121–1128, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2015:GAM

[1641] Y. X. Liu, Z. H. Sun, Y. F. Wang, and R. Q. Yang. Genetic analysis for main length ratio associated with morphological traits in Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 86(3):1129– 1138, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jeuthe:2015:TSA

[1642] H. Jeuthe, E. Brännäs, and J. Nilsson. Thermal stress in Arctic charr Salvelinus alpinus broodstock: a 28 year case study. Journal of Fish Biology, 86(3):1139–1152, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dominguez-Lopez:2015:PGP

[1643] M. Domínguez-López, P. Díaz-Jaimes, M. Uribe-Alcocer, and C. Quiñonez-Velázquez. Post-glacial population expansion of the Monterey Spanish mackerel Scomberomorus concolor in the Gulf of California. Journal of Fish Biology, 86(3):1153–1162, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingley:2015:SPS

[1644] S. J. Ingley. Sympatric populations of sister species of brachyrhaphis. Journal of Fish Biology, 86(3):1163–1170, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernal:2015:LTS

[1645] M. A. Bernal, N. L. Sinai, C. Rocha, M. R. Gaither, F. Dunker, and L. A. Rocha. Long-term sperm storage in the brownbanded bamboo shark Chiloscyllium punctatum. Journal of Fish Biology, 86(3):1171– 1176, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kroon:2015:PAN

[1646] F. Kroon, S. Phillips, D. Burrows, and A. Hogan. Presence and absence of non-native fish species in the Wet Tropics region, Australia. Journal of Fish Biology, 86(3):1177–1185, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seemann:2015:EOE

[1647] F. Seemann, T. Knigge, S. Olivier, and T. Monsinjon. Exogenous 17 βoestradiol (E2) modifies thymus growth and regionalization in European sea bass Dicentrarchus labrax. Journal of Fish Biology, 86(3):1186–1198, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2015:ICT

[1648] S. Yamamoto and M. Sekino. Isolation and characterization of triand tetra-repeat microsatellite loci in the white-spotted charr Salvelinus leucomaenis (Salmonidae). Journal of Fish Biology, 86(3):1199–1202, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domingues:2015:FNR

[1649] R. R. Domingues, S. Montealegre-Quijano, J. M. R. Soto, and A. F. Amorim. Five new records of Luvarus imperialis (Acanthuroidei: Luvaridae) in the south-west Atlantic Ocean. Journal of Fish Biology, 86 (3):1203–1208, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villanova:2015:ICP

[1650] G. V. Villanova, M. Vera, J. Díaz, P. Martinez, N. B. Calcaterra, and S. E. Arranz. Isolation and characterization of 20 polymorphic microsatellite loci in the migratory freshwater fish Leporinus obtusidens (Characiformes: Anostomidae) using 454 shotgun pyrosequencing. Journal of Fish Biology, 86(3):1209–1217, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mirimin:2015:TVM

[1651] L. Mirimin and R. Roodt-Wilding. Testing and validating a modified CTAB DNA extraction method to enable molecular parentage analysis of fertilized eggs and larvae of an emerging South African aquaculture species, the dusky kob Argyrosomus japonicus. Journal of Fish Biology, 86(3):1218–1223, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunlap:2015:SFB

[1652] Paul Dunlap. Symbiosis in fishes: the biology of interspecific partnerships. Journal of Fish Biology, 86(3):1224–1225, March 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2015:WTF

[1653] A. K. Whitfield. Why are there so few freshwater fish species in most estuaries? Journal of Fish Biology, 86(4):1227–1250, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wirgin:2015:POA

[1654] I. Wirgin, L. Maceda, C. Grunwald, and T. L. King. Population origin of Atlantic sturgeon Acipenser oxyrinchus oxyrinchus by-catch in U.S. Atlantic coast fisheries. Journal of Fish Biology, 86(4):1251–1270, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lika:2015:DEU

[1655] K. Lika, M. Pavlidis, N. Mitrizakis, A. Samaras, and N. Papandroulakis. Do experimental units of different scale affect the biological performance of European sea bass Dicentrarchus labrax larvae? Journal of Fish Biology, 86(4):1271–1285, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:2015:MDL

[1656] K. O. Perez and L. A. Fuiman. Maternal diet and larval diet influence survival skills of larval red drum Sciaenops ocellatus. Journal of Fish Biology, 86(4):1286–1304, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2015:MUA

[1657] G. E. White and C. Brown. Microhabitat use affects goby (Gobiidae) cue choice in spatial learning task. Journal of Fish Biology, 86(4):1305– 1318, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Preciado:2015:STV

[1658] I. Preciado, A. Punzón, and F. Velasco. Spatio-temporal variability in the cannibalistic behaviour of European hake Merluccius merluccius: the influence of recruit abundance and prey availability. Journal of Fish Biology, 86(4):1319–1334, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Namiki:2015:GML

[1659] C. Namiki, M. Katsuragawa, and M. L. Zani-Teixeira. Growth and mortality of larval Myctophum affine (Myctophidae, Teleostei). Journal of Fish Biology, 86(4):1335–1347, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costalago:2015:CFA

[1660] D. Costalago, S. Garrido, and I. Palomera. Comparison of the feeding apparatus and diet of European sardines Sardina pilchardus of Atlantic and Mediterranean waters: ecological implications. Journal of Fish Biology, 86(4):1348–1362, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braicovich:2015:HPA

[1661] P. E. Braicovich and J. T. Timi. Homogeneity of parasite assemblages of Dules auriga (Serranidae) in hydrographically heterogeneous sites. Journal of Fish Biology, 86(4):1363–1376, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Libungan:2015:OSP

[1662] L. A. Libungan, G. J. Óskarsson, A. Slotte, J. A. Jacobsen, and S. Pálsson. Otolith shape: a population marker for Atlantic herring Clupea harengus. Journal of Fish Biology, 86(4):1377–1395, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cogliati:2015:DCP

[1663] K. M. Cogliati, C. Danukarjanto, A. C. Pereira, M. J. Lau, A. Hassan, A. F. Mistakidis, B. M. Bolker, B. D. Neff, and S. Balshine. Diet and cannibalism in plainfin midshipman Porichthys notatus. Journal of Fish Biology, 86(4):1396–1415, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2015:IPT

[1664] F. Zhao, P. Zhuang, T. Zhang, Y. Wang, J. Hou, J. Liu, and L. Zhang. Isosmotic points and their ecological significance for juvenile Chinese sturgeon Acipenser sinensis. Journal of Fish Biology, 86(4):1416–1420, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartels:2015:VPC

[1665] H. Bartels, M. F. Docker, M. Krappe, M. M. White, C. Wrede, and I. C. Potter. Variations in the presence of chloride cells in the gills of lampreys (Petromyzontiformes) and their evolutionary implications. Journal of Fish Biology, 86(4):1421–1428, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cella-Ribeiro:2015:TSD

[1666] A. Cella-Ribeiro, L. F. Assakawa, G. Torrente-Vilara, J. Zuanon, R. G. Leite, C. Doria, and F. Duponchelle. Temporal and spatial distribution of young Brachyplatystoma spp. (Siluriformes: Pimelodidae) along the rapids stretch of the Madeira River (Brazil) before the construction of two hydroelectric dams. Journal of Fish Biology, 86(4):1429–1437, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houde:2015:PMC

[1667] A. L. S. Houde, C. C. Wilson, and B. D. Neff. Predictability of multispecies competitive interactions in three populations of Atlantic salmon Salmo salar. Journal of Fish Biology, 86(4):1438–1443, April 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LeLoch:2015:STI

[1668] F. Le Loc’h, J.-D. Durand, K. Diop, and J. Panfili. Spatio-temporal isotopic signatures (δ ^{13}C and δ ^{15}N) reveal that two sympatric West African mullet species do not feed on the same basal production sources. Journal of Fish Biology, 86(4):1444–1453, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2015:OO

[1669] Ian J. Winfield. Opulent oceans. Journal of Fish Biology, 86(4):1454– 1455, April 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holley:2015:MCG

[1670] L. L. Holley, M. K. Heidman, R. M. Chambers, and S. L. Sanderson. Mucous contribution to gut nutrient content in American gizzard shad Dorosoma cepedianum. Journal of Fish Biology, 86(5):1457–1470, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Serra-Pereira:2015:UBV

[1671] B. Serra-Pereira, K. Erzini, and I. Figueiredo. Using biological variables and reproductive strategy of the undulate ray Raja undulata to evaluate productivity and susceptibility to exploitation. Journal of Fish Biology, 86(5):1471–1490, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faustino:2015:MMA

[1672] F. Faustino, L. C. Makino, E. Neumann, and L. S. O. Nakaghi. Morphological and morphometric aspects of early life stages of piabanha Brycon gouldingi (Characidae). Journal of Fish Biology, 86(5):1491–1506, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmentier:2015:IVR

[1673] E. Parmentier, L. Berten, P. Rigo, F. Aubrun, S. L. Nedelec, S. D. Simpson, and D. Lecchini. The influence of various reef sounds on coral-fish larvae behaviour. Journal of Fish Biology, 86(5):1507–1518, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newbold:2015:RBE

[1674] L. R. Newbold, F. A. Hockley, C. F. Williams, J. Cable, A. J. Reading, N. Auchterlonie, and P. S. Kemp. Relationship between European eel Anguilla anguilla infection with non-native parasites and swimming behaviour on encountering accelerating flow. Journal of Fish Biology, 86 (5):1519–1533, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maric:2015:GMV

[1675] S. Marić, A. Snoj, N. Sekulić, J. Krpo-Ćetković, R. Šanda, and V. Jojić. Genetic and morphological variability of the European mudminnow Umbra krameri (Teleostei, Umbridae) in Serbia and in Bosnia and Herzegovina, a basis for future conservation activities. Journal of Fish Biology, 86(5):1534–1548, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribeiro:2015:PSS

[1676] F. F. Ribeiro and J. G. Qin. Prey size selection and cannibalistic behaviour of juvenile barramundi Lates calcarifer. Journal of Fish Biology, 86(5):1549–1566, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munroe:2015:RMP

[1677] S. E. M. Munroe, M. R. Heupel, A. T. Fisk, M. Logan, and C. A. Simpfendorfer. Regional movement patterns of a small-bodied shark revealed by stable-isotope analysis. Journal of Fish Biology, 86(5):1567– 1586, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piddocke:2015:AGM

[1678] T. P. Piddocke, G. L. Butler, P. A. Butcher, J. Stewart, D. J. Bucher, and L. Christidis. Age and growth of mangrove red snapper Lutjanus argentimaculatus at its cool-water-range limits. Journal of Fish Biology, 86(5):1587–1600, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eckmann:2015:AIP

[1679] R. Eckmann. Absence of intrinsic post-zygotic incompatibilities in artificial crosses between sympatric coregonid species from upper Lake Constance. Journal of Fish Biology, 86(5):1601–1611, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jia:2015:BCO

[1680] Y. D. Jia, H. X. Niu, Z. Meng, X. F. Liu, and J. L. Lei. Biochemical composition of the ovarian fluid and its effects on the fertilization capacity of turbot Scophthalmus maximus during the spawning season. Journal of Fish Biology, 86(5):1612–1620, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harmon:2015:ICP

[1681] B. S. Harmon, R. Hilborn, and T. P. Quinn. Infection by the cestode parasite Schistocephalus sp. and effects on diet, body condition and survival of sculpins Cottus aleuticus and Cottus cognatus. Journal of Fish Biology, 86(5):1621–1629, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2015:HDA

[1682] R. M. Rideout, S. A. Smith, and M. J. Morgan. High-density aggregations of rodlet cells in the gonads of Greenland halibut Reinhardtius hippoglossoides, a deep-water marine flatfish. Journal of Fish Biology, 86(5):1630–1637, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2015:TIE

[1683] M. M. Rahman, C. Gasparini, G. M. Turchini, and J. P. Evans. Testing the interactive effects of carotenoids and polyunsaturated fatty acids on ejaculate traits in the guppy Poecilia reticulata (Pisces: Poeciliidae). Journal of Fish Biology, 86(5):1638–1643, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Furtbauer:2015:VIE

[1684] I. Fürtbauer, A. J. King, and M. Heistermann. Visible implant elastomer (VIE) tagging and simulated predation risk elicit similar physiological stress responses in three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 86(5):1644–1649, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoareau:2015:ERP

[1685] T. B. Hoareau, A. W. Klopper, S. M. R. Dos Santos, C. J. Oosthuizen, and P. Bloomer. Evaluating the resolution power of new microsatellites for species identification and stock delimitation in the Cape hakes Merluccius paradoxus and Merluccius capensis (Teleostei: Merlucciidae). Journal of Fish Biology, 86(5):1650–1657, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2015:DPS

[1686] D. L. Morgan, M. G. Allen, B. C. Ebner, J. M. Whitty, and S. J. Beatty. Discovery of a pupping site and nursery for critically endangered green sawfish Pristis zijsron. Journal of Fish Biology, 86(5):1658–1663, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langford:2015:IGI

[1687] Terry E. L. Langford. Identification guide to the inshore fish of the British Isles. Journal of Fish Biology, 86(5):1664–1665, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2015:E

[1688] Anonymous. Erratum. Journal of Fish Biology, 86(5):1666, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stringwell:2015:E

[1689] R. Stringwell, A. Lock, C. J. Stutchbury, E. Baggett, J. Taylor, P. J. Gough, and C. Garcia de Leaniz. Erratum. Journal of Fish Biology, 86 (5):1667–1668, May 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zelazowska:2015:ONC

[1690] M. Żelazowska, W. Jankowska, E. Plewniak, and U. Rajek. Ovarian nests in cultured Russian sturgeon Acipenser gueldenstaedtii and North American paddlefish Polyodon spathula comprised of previtellogenic oocytes. Journal of Fish Biology, 86(6):1669–1679, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Izzo:2015:CWC

[1691] C. Izzo, Z. A. Doubleday, A. G. Schultz, S. H. Woodcock, and B. M. Gillanders. Contribution of water chemistry and fish condition to otolith chemistry: comparisons across salinity environments. Journal of Fish Biology, 86(6):1680–1698, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundt-Hansen:2015:FAS

[1692] L. Sundt-Hansen, J. Huisman, H. Skoglund, and K. Hindar. Farmed Atlantic salmon Salmo salar L. parr may reduce early survival of wild fish. Journal of Fish Biology, 86(6):1699–1712, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leurs:2015:RAU

[1693] G. Leurs, C. P. O’Connell, S. Andreotti, M. Rutzen, and H. Vonk Noordegraaf. Risks and advantages of using surface laser photogrammetry on free-ranging marine organisms: a case study on white sharks Carcharodon carcharias. Journal of Fish Biology, 86(6):1713–1728, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lauerburg:2015:SSF

[1694] R. A. M. Lauerburg, F. Keyl, P. Kotterba, J. Floeter, and A. Temming. Sex-specific food intake in whiting Merlangius merlangus. Journal of Fish Biology, 86(6):1729–1753, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spares:2015:RMP

[1695] A. D. Spares, M. J. W. Stokesbury, M. J. Dadswell, R. K. O’Dor, and T. A. Dick. Residency and movement patterns of Arctic charr Salvelinus alpinus relative to major estuaries. Journal of Fish Biology, 86(6):1754– 1780, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almeida-Silva:2015:TEF

[1696] P. H. Almeida-Silva, R. A. Tubino, L. C. Zambrano, D. A. Hunder, S. R. Garritano, and C. Monteiro-Neto. Trophic ecology and food consumption of fishes in a hypersaline tropical lagoon. Journal of Fish Biology, 86(6): 1781–1795, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cardeira:2015:MAC

[1697] J. Cardeira, A. C. Mendes, P. Pousão-Ferreira, M. L. Cancela, and P. J. Gavaia. Micro-anatomical characterization of vertebral curvatures in Senegalese sole Solea senegalensis. Journal of Fish Biology, 86(6):1796– 1810, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McDermott:2015:QAP

[1698] S. P. McDermott, N. C. Bransome, S. E. Sutton, B. E. Smith, J. S. Link, and T. J. Miller. Quantifying alosine prey in the diets of marine piscivores in the Gulf of Maine. Journal of Fish Biology, 86(6):1811– 1829, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medeiros:2015:OUE

[1699] A. M. Medeiros, O. J. Luiz, and C. Domit. Occurrence and use of an estuarine habitat by giant manta ray Manta birostris. Journal of Fish Biology, 86(6):1830–1838, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2015:VEW

[1700] K. Magellan. Voluntary emergence and water detection in a newly recognized amphibious fish. Journal of Fish Biology, 86(6):1839–1844, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Calich:2015:MSR

[1701] H. J. Calich and S. E. Campana. Mating scars reveal mate size in immature female blue shark Prionace glauca. Journal of Fish Biology, 86 (6):1845–1851, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ariyomo:2015:EHL

[1702] T. O. Ariyomo and P. J. Watt. Effect of hunger level and time of day on boldness and aggression in the zebrafish Danio rerio. Journal of Fish Biology, 86(6):1852–1859, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galindo:2015:ICP

[1703] B. A. Galindo, D. G. Ferreira, F. S. Almeida, J. Carlsson, and S. H. Sofia. Isolation and characterization of 13 polymorphic microsatellite loci in Hypostomus ancistroides (Teleostei, Loricariidae) and cross-amplification in related species. Journal of Fish Biology, 86(6):1860–1866, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amsler:2015:SOP

[1704] M. O. Amsler, K. E. Smith, J. B. McClintock, H. Singh, S. Thatje, S. C. Vos, C. J. Brothers, A. Brown, D. Ellis, J. Anderson, and R. B. Aronson. In situ observations of a possible skate nursery off the western Antarctic Peninsula. Journal of Fish Biology, 86(6):1867–1872, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barry:2015:LSC

[1705] J. Barry, K. Mcharg, J. A. Dodd, and C. E. Adams. Local scale, coastal currents influence recruitment to freshwater populations in the European eel Anguilla anguilla: a case study from the Isle of Man. Journal of Fish Biology, 86(6):1873–1880, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colmenero:2015:CUO

[1706] A. I. Colmenero, V. M. Tuset, J.-M. Fortuño, and P. Sánchez. The chorion ultrastructure of ova of Lophius spp. Journal of Fish Biology, 86(6):1881–1886, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oh:2015:MMC

[1707] J. Oh and S. Kim. Morphological and molecular characterization of separated pelagic eggs from Lophius litulon (Lophiiformes; Lophiidae). Journal of Fish Biology, 86(6):1887–1891, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ball:2015:FFP

[1708] Rachel E. Ball. Future of fisheries: perspectives for emerging professionals. Journal of Fish Biology, 86(6):1892–1893, June 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilhelm:2015:BOZ

[1709] M. R. Wilhelm, J.-P. Roux, C. L. Moloney, and A. Jarre. Biannual otolith-zone formation of young shallow-water hake Merluccius capensis in the northern Benguela: age verification using otoliths sampled by a top predator. Journal of Fish Biology, 87(1):1–16, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cid:2015:MSV

[1710] P. Cid, M. J. Doldán, and E. De Miguel Villegas. Morphogenesis of the saccus vasculosus of turbot Scophthalmus maximus: assessment of cell proliferation and distribution of parvalbumin and calretinin during ontogeny. Journal of Fish Biology, 87(1):17–27, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2015:EAV

[1711] L. Li, Y. Jia, P. Li, S. Yin, G. Zhang, X. Wang, Y. Wang, X. Wang, X. Zang, and Y. Ding. Expression and activity of V-H^{+}-ATPase in gill and kidney of marbled eel Anguilla marmorata in response to salinity challenge. Journal of Fish Biology, 87(1):28–42, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanuzzo:2015:STO

[1712] F. S. Zanuzzo, E. C. Urbinati, G. W. Nash, and A. K. Gamperl. Steelhead trout Oncorhynchus mykiss metabolic rate is affected by dietary Aloe vera inclusion but not by mounting an immune response against formalinkilled Aeromonas salmonicida. Journal of Fish Biology, 87(1):43–53, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beirao:2015:IPO

[1713] J. Beirão, C. F. Purchase, B. F. Wringe, and I. A. Fleming. Interpopulation ovarian fluid variation differentially modulates sperm motility in Atlantic cod Gadus morhua. Journal of Fish Biology, 87(1):54–68, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pappalardo:2015:CPG

[1714] A. M. Pappalardo, E. G. Gonzalez, C. Tigano, I. Doadrio, and V. Ferrito. Comparative pattern of genetic structure in two Mediterranean killifishes Aphanius fasciatus and Aphanius iberus inferred from both mitochondrial and nuclear data. Journal of Fish Biology, 87(1):69–87, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnhagen:2015:BRN

[1715] C. Magnhagen, T. Backström, Ø. Øverli, S. Winberg, J. Nilsson, M. A. Vindas, and E. Brännäs. Behavioural responses in a net restraint test predict interrenal reactivity in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 87(1):88–99, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cohen:2015:CLH

[1716] S. N. Cohen, J. U. Regus, Y. Reynoso, T. Mastro, and D. N. Reznick. Comparative life histories of fishes in the subgenus Limia (Pisces: Poeciliidae). Journal of Fish Biology, 87(1):100–114, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:2015:ECF

[1717] Y. Zheng, L. M. Peng, F. You, Y. X. Zou, P. J. Zhang, and S. L. Chen. Establishment and characterization of a fish-cell line from the brain of Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 87(1): 115–122, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diez:2015:BCS

[1718] G. Díez, M. Soto, and J. M. Blanco. Biological characterization of the skin of shortfin mako shark Isurus oxyrinchus and preliminary study of the hydrodynamic behaviour through computational fluid dynamics. Journal of Fish Biology, 87(1):123–137, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonHerbing:2015:CSR

[1719] I. Hunt von Herbing, T.-C. F. Pan, F. Méndez-Sánchez, M. Garduño-Paz, O. Hernández-Gallegos, M. L. Ruiz-Gómez, and G. Rodríguez-Vargas. Chronic stress of rainbow trout Oncorhynchus mykiss at high altitude: a field study. Journal of Fish Biology, 87(1):138–158, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becker:2015:DBM

[1720] R. A. Becker, N. G. Sales, G. M. Santos, G. B. Santos, and D. C. Carvalho. DNA barcoding and morphological identification of neotropical ichthyoplankton from the Upper Paraná and São Francisco. Journal of Fish Biology, 87(1):159–168, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freshwater:2015:VDI

[1721] C. Freshwater, M. Trudel, T. D. Beacham, C.-E. Neville, S. Tucker, and F. Juanes. Validation of daily increments and a marine-entry check in the otoliths of sockeye salmon Oncorhynchus nerka post-smolts. Journal of Fish Biology, 87(1):169–178, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salinas-de-Leon:2015:FRS

[1722] P. Salinas de León, E. Rastoin, and D. Acuña-Marrero. First record of a spawning aggregation for the tropical eastern Pacific endemic grouper Mycteroperca olfax in the Galapagos Marine Reserve. Journal of Fish Biology, 87(1):179–186, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2015:SSD

[1723] B. Jonsson and N. Jonsson. Sexual size dimorphism in anadromous brown trout Salmo trutta. Journal of Fish Biology, 87(1):187–193, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polacik:2015:CFC

[1724] M. Polačik, P. Jurajda, R. Blažek, and M. Janáč. Carcass feeding as a cryptic foraging mode in round goby Neogobius melanostomus. Journal of Fish Biology, 87(1):194–199, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lyons:2015:ILH

[1725] K. Lyons, A. Preti, D. J. Madigan, R. J. D. Wells, M. E. Blasius, O. E. Snodgrass, D. Kacev, J. D. Harris, H. Dewar, S. Kohin, K. MacKenzie, and C. G. Lowe. Insights into the life history and ecology of a large shortfin mako shark Isurus oxyrinchus captured in southern California. Journal of Fish Biology, 87(1):200–211, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2015:FA

[1726] Laith A. Jawad. Fishes of Afghanistan. Journal of Fish Biology, 87(1): 212–213, July 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nandi:2015:SSF

[1727] S. Nandi and S. K. Saikia. Size-selective feeding on phytoplankton by two morpho-groups of the small freshwater fish Amblypharyngodon mola. Journal of Fish Biology, 87(2):215–230, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2015:EPE

[1728] H. B. Zhao, X. Y. Zhang, G. Q. Feng, M. M. Guo, P. Chang, C. Qi, X. P. Zhong, Q. C. Zhou, and J. L. Wang. Expression of plzfa in embryo and adult of medaka Oryzias latipes. Journal of Fish Biology, 87(2):231– 240, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shan:2015:IEA

[1729] S. J. Shan, D. Z. Liu, L. Wang, Y. Y. Zhu, F. M. Zhang, T. Li, L. G. An, and G. W. Yang. Identification and expression analysis of irak1 gene in common carp Cyprinus carpio L.: indications for a role of antibacterial and antiviral immunity. Journal of Fish Biology, 87(2):241–255, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaslet:2015:TBJ

[1730] A. Vaslet, D. L. Phillips, C. A. M. France, I. C. Feller, and C. C. Baldwin. Trophic behaviour of juvenile reef fishes inhabiting interlinked mangrove– seagrass habitats in offshore mangrove islets. Journal of Fish Biology, 87(2):256–273, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katselis:2015:SSD

[1731] G. Katselis, K. Koukou, A. Ramfos, and D. K. Moutopoulos. Sex-specific daily spawning seaward migration of striped mullet Mugil cephalus in a coastal lagoon. Journal of Fish Biology, 87(2):274–285, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorval:2015:CGM

[1732] E. Dorval, J. D. McDaniel, B. J. Macewicz, and D. L. Porzio. Changes in growth and maturation parameters of Pacific sardine Sardinops sagax collected off California during a period of stock recovery from 1994 to 2010. Journal of Fish Biology, 87(2):286–310, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lara-Mendoza:2015:FHS

[1733] R. E. Lara-Mendoza, J. F. Márquez-Farías, and J. C. Román-Reyes. Feeding habits of the speckled guitarfish Rhinobatos glaucostigma (Elasmobranchii: Rhinobatidae). Journal of Fish Biology, 87(2):311–322, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Escalle:2015:RMM

[1734] L. Escalle, C. W. Speed, M. G. Meekan, W. T. White, R. C. Babcock, R. D. Pillans, and C. Huveneers. Restricted movements and mangrove dependency of the nervous shark Carcharhinus cautus in nearshore coastal waters. Journal of Fish Biology, 87(2):323–341, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Havn:2015:ECR

[1735] T. B. Havn, I. Uglem, Ø. Solem, S. J. Cooke, F. G. Whoriskey, and E. B. Thorstad. The effect of catch-and-release angling at high water temperatures on behaviour and survival of Atlantic salmon Salmo salar during spawning migration. Journal of Fish Biology, 87(2):342–359, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2015:CEM

[1736] D. L. Zhang, D. H. Yu, J. Chen, C. Chen, and Z. Y. Wang. Co-expression of march5b and tlr7 in large yellow croaker Larimichthys crocea in response to Cryptocaryon irritans infection. Journal of Fish Biology, 87 (2):360–370, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chabot:2015:MLC

[1737] C. L. Chabot. Microsatellite loci confirm a lack of population connectivity among globally distributed populations of the tope shark Galeorhinus galeus (Triakidae). Journal of Fish Biology, 87(2):371–385, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2015:ICL

[1738] Z.-Y. Chen, W.-Y. Hsieh, and Y.-S. Lai. Identification and characterization of a late gene of grouper iridovirus 61l and antibody production against the protein encoded by it. Journal of Fish Biology, 87(2):386– 399, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertucci:2015:SPD

[1739] F. Bertucci, P. Lejeune, J. Payrot, and E. Parmentier. Sound production by dusky grouper Epinephelus marginatus at spawning aggregation sites. Journal of Fish Biology, 87(2):400–421, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmes:2015:AGT

[1740] B. J. Holmes, V. M. Peddemors, A. N. Gutteridge, P. T. Geraghty, R. W. K. Chan, I. R. Tibbetts, and M. B. Bennett. Age and growth of the tiger shark Galeocerdo cuvier off the east coast of Australia. Journal of Fish Biology, 87(2):422–448, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aalvik:2015:SEC

[1741] I. M. Aalvik, E. Moland, E. M. Olsen, and N. C. Stenseth. Spatial ecology of coastal Atlantic cod Gadus morhua associated with parasite load. Journal of Fish Biology, 87(2):449–464, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2015:MAE

[1742] A. B. M. Moore. Morphological abnormalities in elasmobranchs. Journal of Fish Biology, 87(2):465–471, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forsman:2015:PEL

[1743] A. Forsman, P. Tibblin, H. Berggren, O. Nordahl, P. Koch-Schmidt, and P. Larsson. Pike Esox lucius as an emerging model organism for studies in ecology and evolutionary biology: a review. Journal of Fish Biology, 87(2):472–479, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olafsdottir:2015:MRS

[1744] G. Á. Ólafsdóttir, G. S. Gunnarsson, and H. Karlsson. More rapid shift to a benthic niche in larger Gadus morhua juveniles. Journal of Fish Biology, 87(2):480–486, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hundt:2015:FEN

[1745] M. Hundt, A. Scharbert, U. Weibel, G. Kuhn, K. Metzner, P. Jatteau, A. Pies, R. Schulz, and R. Gergs. First evidence of natural reproduction of the Allis shad Alosa alosa in the River Rhine following re-introduction measures. Journal of Fish Biology, 87(2):487–493, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fost:2015:PAP

[1746] B. A. Fost and C. P. Ferreri. Preference and avoidance pH of brook trout Salvelinus fontinalis and brown trout Salmo trutta exposed to different holding pH. Journal of Fish Biology, 87(2):494–501, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kever:2015:EST

[1747] L. Kéver, K. S. Boyle, and E. Parmentier. Effects of seawater temperature on sound characteristics in Ophidion rochei (Ophidiidae). Journal of Fish Biology, 87(2):502–509, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:2015:HTV

[1748] J. A. Nelson and G. K. Lipkey. Hypoxia tolerance variance between swimming and resting striped bass Morone saxatilis. Journal of Fish Biology, 87(2):510–518, August 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Avise:2015:PGE

[1749] J. C. Avise and A. Tatarenkov. Population genetics and evolution of the mangrove rivulus Kryptolebias marmoratus, the world’s only selffertilizing hermaphroditic vertebrate. Journal of Fish Biology, 87(3): 519–538, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sreetharan:2015:EDL

[1750] S. Sreetharan, C. Thome, C. Mitz, J. Eme, C. A. Mueller, E. N. Hulley, R. G. Manzon, C. M. Somers, D. R. Boreham, and J. Y. Wilson. Embryonic development of lake whitefish Coregonus clupeaformis: a staging series, analysis of growth and effects of fixation. Journal of Fish Biology, 87(3):539–558, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2015:IVF

[1751] D. R. Carvalho, D. Castro, M. Callisto, M. Z. Moreira, and P. S. Pompeu. Isotopic variation in five species of stream fishes under the influence of different land uses. Journal of Fish Biology, 87(3):559–578, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casaretto:2015:SPI

[1752] L. Casaretto, M. Picciulin, and A. D. Hawkins. Seasonal patterns and individual differences in the calls of male haddock Melanogrammus aeglefinus. Journal of Fish Biology, 87(3):579–603, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leiser:2015:MPF

[1753] J. K. Leiser, J. L. Gagliardi-Seeley, B. D. Wisenden, and M. Itzkowitz. Mating patterns of female Leon Springs pupfish Cyprinodon bovinus. Journal of Fish Biology, 87(3):604–615, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nishio:2015:LHR

[1754] M. Nishio, T. Kawamoto, R. Kawakami, K. Edo, and Y. Yamazaki. Life history and reproductive ecology of the endangered itasenpara bitterling Acheilognathus longipinnis (Cyprinidae) in the Himi region, central Japan. Journal of Fish Biology, 87(3):616–633, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coluccia:2015:CCS

[1755] E. Coluccia, F. Deidda, R. Cannas, C. Lobina, D. Cuccu, A. M. Deiana, and S. Salvadori. Comparative cytogenetics of six Indo–Pacific moray eels (Anguilliformes: Muraenidae) by chromosomal banding and fluorescence in situ hybridization. Journal of Fish Biology, 87(3):634–645, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mille:2015:SOM

[1756] T. Mille, K. Mahe, M. C. Villanueva, H. De Pontual, and B. Ernande. Sagittal otolith morphogenesis asymmetry in marine fishes. Journal of Fish Biology, 87(3):646–663, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matta:2015:RBA

[1757] M. E. Matta. Reproductive biology of the Alaska skate Bathyraja parmifera, with comments on an intersexual individual. Journal of Fish Biology, 87(3):664–678, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2015:SHP

[1758] M. Correia, I. R. Caldwell, H. J. Koldewey, J. P. Andrade, and J. Palma. Seahorse (Hippocampinae) population fluctuations in the Ria Formosa Lagoon, south Portugal. Journal of Fish Biology, 87(3):679–690, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bower:2015:SGM

[1759] L. M. Bower and K. R. Piller. Shaping up: a geometric morphometric approach to assemblage ecomorphology. Journal of Fish Biology, 87(3): 691–714, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaither:2015:TDE

[1760] M. R. Gaither, M. A. Bernal, I. Fernandez-Silva, M. Mwale, S. A. Jones, C. Rocha, and L. A. Rocha. Two deep evolutionary lineages in the circumtropical glasseye Heteropriacanthus cruentatus (Teleostei, Priacanthidae) with admixture in the south-western Indian Ocean. Journal of Fish Biology, 87(3):715–727, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rochowski:2015:GVL

[1761] B. E. A. Rochowski, T. I. Walker, and R. W. Day. Geographical variability in life-history traits of a midslope dogfish: the brier shark Deania calcea. Journal of Fish Biology, 87(3):728–747, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herrera:2015:DPM

[1762] M. Herrera, P. Saenz-Agudelo, G. B. Nanninga, and M. L. Berumen. Development of polymorphic microsatellite loci for conservation genetic studies of the coral reef fish Centropyge bicolor. Journal of Fish Biology, 87(3):748–753, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2015:DGF

[1763] K. V. Reis, P. C. Venere, I. Sampaio, P. S. Rêgo, M. Vallinoto, and I. L. Souza. Downstairs gene flow: the effects of a linear sequence of waterfalls on the only population of the endangered minnow Astyanax xavante. Journal of Fish Biology, 87(3):754–762, September 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2015:UTT

[1764] Z. Chen, R. H. Devlin, and A. P. Farrell. Upper thermal tolerance of wild-type, domesticated and growth hormone-transgenic Coho salmon Oncorhynchus kisutch. Journal of Fish Biology, 87(3):763–773, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Battaglia:2015:RBO

[1765] P. Battaglia, D. Malara, G. Ammendolia, T. Romeo, and F. Andaloro. Relationships between otolith size and fish length in some mesopelagic teleosts (Myctophidae, Paralepididae, Phosichthyidae and Stomiidae). Journal of Fish Biology, 87(3):774–782, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Antoni:2015:NTD

[1766] M. Y. Deli Antoni, M. González-Castro, and J. M. Díaz de Astarloa. New tools (DNA barcoding), old hypothesis: the case of the taxonomic identity of the Argentine hakes (Actinopterygii: Merluccius). Journal of Fish Biology, 87(3):783–793, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corrigan:2015:CGP

[1767] S. Corrigan, D. Kacev, and J. Werry. A case of genetic polyandry in the shortfin mako Isurus oxyrinchus. Journal of Fish Biology, 87(3):794–798, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stanczak:2015:MML

[1768] K. Stańczak, S. Krejszeff, M. Debowska, K. Mierzejewska, M. Woźniak, and P. Hliwa. Mass marking of Leuciscus idus larvae using Artemia salina as a vector of fluorescent dyes. Journal of Fish Biology, 87(3):

799–804, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fogelson:2015:GHU

[1769] S. B. Fogelson, R. P. E. Yanong, A. Kane, C. N. Teal, I. K. Berzins, S. A. Smith, C. Brown, and A. Camus. Gross, histological and ultrastructural morphology of the aglomerular kidney in the lined seahorse Hippocampus erectus. Journal of Fish Biology, 87(3):805–813, September 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turko:2015:EEP

[1770] A. J. Turko and P. A. Wright. Evolution, ecology and physiology of amphibious killifishes (Cyprinodontiformes). Journal of Fish Biology, 87 (4):815–835, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lagrue:2015:MFB

[1771] C. Lagrue and R. Poulin. Measuring fish body condition with or without parasites: does it matter? Journal of Fish Biology, 87(4):836–847, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gold:2015:CRH

[1772] D. K. Gold, T. Loirat, and A. P. Farrell. Cardiorespiratory responses to haemolytic anaemia in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 87(4):848–859, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birindelli:2015:NSC

[1773] J. L. O. Birindelli, L. M. Sarmento-Soares, and F. C. T. Lima. A new species of Centromochlus (Siluriformes, Auchenipteridae, Centromochlinae) from the middle Rio Tocantins basin, Brazil. Journal of Fish Biology, 87(4):860–875, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedreschi:2015:TFO

[1774] D. Pedreschi, S. Mariani, J. Coughlan, C. C. Voigt, M. O’Grady, J. Caffrey, and M. Kelly-Quinn. Trophic flexibility and opportunism in pike Esox lucius. Journal of Fish Biology, 87(4):876–894, October 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smylie:2015:PCO

[1775] M. Smylie, V. Shervette, and C. McDonough. Prey composition and ontogenetic shift in coastal populations of longnose gar Lepisosteus osseus. Journal of Fish Biology, 87(4):895–911, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ba:2015:SGO

[1776] A. Ba, K. Diouf, F. Guilhaumon, and J. Panfili. Slow growth of the overexploited milk shark Rhizoprionodon acutus affects its sustainability in West Africa. Journal of Fish Biology, 87(4):912–929, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klibansky:2015:SFA

[1777] N. Klibansky and F. S. Scharf. Success and failure assessing gonad maturity in sequentially hermaphroditic fishes: comparisons between macroscopic and microscopic methods. Journal of Fish Biology, 87(4):930–957, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fujioka:2015:RSR

[1778] Y. Fujioka, T. Kikko, J. Saegusa, T. Usuki, and T. Ohtsuka. Response of sex ratio to timing of breeding in the small cyprinid Gnathopogon caerulescens. Journal of Fish Biology, 87(4):958–966, October 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2015:TSS

[1779] R. S. A. White, P. A. McHugh, C. N. Glover, and A. R. McIntosh. Trapshyness subsidence is a threshold function of mark–recapture interval in brown mudfish Neochanna apoda populations. Journal of Fish Biology, 87(4):967–980, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vanesyan:2015:TSR

[1780] A. Vanesyan, F. H. Rodd, and W. S. Ryu. Tracking the startle response of guppies Poecilia reticulata in three dimensions. Journal of Fish Biology, 87(4):981–999, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espino:2015:MAF

[1781] F. Espino, A. Brito, R. Haroun, and F. Tuya. Macroecological analysis of the fish fauna inhabiting Cymodocea nodosa seagrass meadows. Journal of Fish Biology, 87(4):1000–1018, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trenkel:2015:EDD

[1782] V. M. Trenkel, P. Lorance, S. M. M. Fässler, and Å. S. Høines. Effects of density dependence, zooplankton and temperature on blue whiting Micromesistius poutassou growth. Journal of Fish Biology, 87(4):1019– 1030, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saunders:2015:TPM

[1783] R. A. Saunders, M. A. Collins, P. Ward, G. Stowasser, R. Shreeve, and G. A. Tarling. Trophodynamics of Protomyctophum (Myctophidae) in the Scotia Sea (Southern Ocean). Journal of Fish Biology, 87(4):1031– 1058, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gorski:2015:VHU

[1784] K. Górski, C. De Gruijter, and R. Tana. Variation in habitat use along the freshwater–marine continuum by grey mullet Mugil cephalus at the southern limits of its distribution. Journal of Fish Biology, 87(4):1059– 1071, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guallart:2015:ARO

[1785] J. Guallart, P. García-Salinas, A. E. Ahuir-Baraja, M. Guimerans, J. R. Ellis, and M. Roche. Angular roughshark Oxynotus centrina (Squaliformes: Oxynotidae) in captivity feeding exclusively on elasmobranch eggs: an overlooked feeding niche or a matter of individual taste? Journal of Fish Biology, 87(4):1072–1079, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otero-Ferrer:2015:FRH

[1786] F. Otero-Ferrer, R. Herrera, A. López, J. Socorro, L. Molina, and C. Bouza. First records of Hippocampus algiricus in the Canary Islands (north-east Atlantic Ocean) with an observation of hybridization with Hippocampus hippocampus. Journal of Fish Biology, 87(4):1080–1089, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciccotto:2015:EPF

[1787] P. J. Ciccotto and T. C. Mendelson. Evolution of the premaxillary fraenum and substratum in snubnose darters and allies (Percidae:

Etheostoma). Journal of Fish Biology, 87(4):1090–1098, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorscheider:2015:KDA

[1788] C. A. Lorscheider, C. H. Zawadzki, M. R. Vicari, I. C. Martins-Santos, and R. F. Artoni. Karyotypic diversity of the armoured catfish genus Hypostomus (Siluriformes: Loricariidae) in the context of its occurrence and distribution. Journal of Fish Biology, 87(4):1099–1110, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2015:TSD

[1789] K. Morita and A. Nakashima. Temperature seasonality during fry outmigration influences the survival of hatchery-reared chum salmon Oncorhynchus keta. Journal of Fish Biology, 87(4):1111–1117, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2015:SSD

[1790] N. D. Phillips, C. Harrod, A. R. Gates, T. M. Thys, and J. D. R. Houghton. Seeking the sun in deep, dark places: mesopelagic sightings of ocean sunfishes (Molidae). Journal of Fish Biology, 87(4):1118–1126, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2015:BES

[1791] Laith A. Jawad. Biology and ecology of sardines and anchovies. Journal of Fish Biology, 87(4):1127–1128, October 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2015:EEA

[1792] J. T. Kelly, D. T. Lerner, M. F. O’Dea, A. M. Regish, M. Y. Monette, J. P. Hawkes, K. H. Nislow, and S. D. McCormick. Evidence for episodic acidification effects on migrating Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 87(5):1129–1146, November 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torres-Velarde:2015:EMS

[1793] J. Torres-Velarde, L. Ibarra-Castro, E. Rodríguez-Ibarra, I. Sifuentes-Romero, R. Hernández-Cornejo, and A. García-Gasca. Expression of myostatin in the spotted rose snapper Lutjanus guttatus during larval and juvenile development under cultured conditions. Journal of Fish Biology, 87(5):1147–1164, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:2015:EAH

[1794] R. A. Weber, J. J. Pérez Maceira, M. J. Aldegunde, J. B. Peleteiro, L. O. García Martín, and M. Aldegunde. Effects of acute handling stress on cerebral monoaminergic neurotransmitters in juvenile Senegalese sole Solea senegalensis. Journal of Fish Biology, 87(5):1165–1175, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Folkvord:2015:DGS

[1795] A. Folkvord, K. W. Vollset, and I. A. Catalán. Differences in growth and survival between cod Gadus morhua and herring Clupea harengus early stages co-reared at variable prey concentrations. Journal of Fish Biology, 87(5):1176–1190, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagelkerke:2015:SGD

[1796] L. A. J. Nagelkerke, K. M. Leon-Kloosterziel, H.-J. Megens, M. De Graaf, O. E. Diekmann, and F. A. Sibbing. Shallow genetic divergence and species delineations in the endemic Labeobarbus species flock of Lake Tana, Ethiopia. Journal of Fish Biology, 87(5):1191–1208, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahiri:2015:EDP

[1797] S. Lahiri and M. L. Fine. Explosive development of pectoral muscle fibres in large juvenile blue catfish Ictalurus furcatus. Journal of Fish Biology, 87(5):1209–1218, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conradsen:2015:SDM

[1798] C. Conradsen and K. McGuigan. Sexually dimorphic morphology and swimming performance relationships in wild-type zebrafish Danio rerio. Journal of Fish Biology, 87(5):1219–1233, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowinski:2015:WAB

[1799] P. K. Rowiński, F. Mateos-Gonzalez, E. Sandblom, F. Jutfelt, A. Ekström, and L. F. Sundström. Warming alters the body shape of European perch Perca fluviatilis. Journal of Fish Biology, 87(5):1234–1247, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rheault:2015:DMC

[1800] G. Rheault, M. Langevin, G. Cabana, and H. Glémet. Dermal melanin concentration of yellow perch Perca flavescens in relation to water transparency. Journal of Fish Biology, 87(5):1248–1253, November 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gustavson:2015:EAI

[1801] M. S. Gustavson, P. C. Collins, J. A. Finarelli, D. Egan, R. Ó. Conchúir, G. D. Wightman, J. J. King, D. T. Gauthier, K. Whelan, J. E. L. Carlsson, and J. Carlsson. An eDNA assay for Irish Petromyzon marinus and Salmo trutta and field validation in running water. Journal of Fish Biology, 87(5):1254–1262, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:2015:CMP

[1802] Stephen J. Lockwood. The controversy over marine protected areas: science meets policy. Journal of Fish Biology, 87(5):1263–1264, November 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sims:2015:BEC

[1803] D. W. Sims. The biology, ecology and conservation of elasmobranchs: recent advances and new frontiers. Journal of Fish Biology, 87(6):1265– 1270, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cailliet:2015:PEL

[1804] G. M. Cailliet. Perspectives on elasmobranch life-history studies: a focus on age validation and relevance to fishery management. Journal of Fish Biology, 87(6):1271–1292, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skomal:2015:SOW

[1805] G. B. Skomal, E. M. Hoyos-Padilla, A. Kukulya, and R. Stokey. Subsurface observations of white shark Carcharodon carcharias predatory behaviour using an autonomous underwater vehicle. Journal of Fish Biology, 87(6):1293–1312, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lea:2015:MBL

[1806] J. S. E. Lea, N. E. Humphries, C. R. Clarke, and D. W. Sims. To Madagascar and back: long-distance, return migration across open ocean by a pregnant female bull shark Carcharhinus leucas. Journal of Fish Biology, 87(6):1313–1321, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guttridge:2015:OHU

[1807] T. L. Guttridge, S. J. B. Gulak, B. R. Franks, J. K. Carlson, S. H. Gruber, K. S. Gledhill, M. E. Bond, G. Johnson, and R. D. Grubbs. Occurrence and habitat use of the critically endangered smalltooth sawfish Pristis pectinata in the Bahamas. Journal of Fish Biology, 87(6):1322–1341, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bullock:2015:BRJ

[1808] R. W. Bullock, T. L. Guttridge, I. G. Cowx, M. Elliott, and S. H. Gruber. The behaviour and recovery of juvenile lemon sharks Negaprion brevirostris in response to external accelerometer tag attachment. Journal of Fish Biology, 87(6):1342–1354, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruce:2015:SAS

[1809] B. Bruce and R. Bradford. Segregation or aggregation? Sex-specific patterns in the seasonal occurrence of white sharks Carcharodon carcharias at the Neptune Islands, South Australia. Journal of Fish Biology, 87(6): 1355–1370, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gledhill:2015:GSP

[1810] K. S. Gledhill, S. T. Kessel, T. L. Guttridge, A. C. Hansell, A. E. Bester van der Merwe, K. A. Feldheim, S. H. Gruber, and D. D. Chapman. Genetic structure, population demography and seasonal occurrence of blacktip shark Carcharhinus limbatus in Bimini, The Bahamas. Journal of Fish Biology, 87(6):1371–1388, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dureuil:2015:EGT

[1811] M. Dureuil and B. Worm. Estimating growth from tagging data: an application to north-east Atlantic tope shark Galeorhinus galeus. Journal of Fish Biology, 87(6):1389–1410, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2015:RCL

[1812] S. R. McCully Phillips and J. R. Ellis. Reproductive characteristics and life-history relationships of starry smooth-hound Mustelus asterias in British waters. Journal of Fish Biology, 87(6):1411–1433, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romero-Caicedo:2015:RWG

[1813] A. F. Romero-Caicedo and M. Carrera-Fernández. Reproduction of the whitesnout guitarfish Rhinobatos leucorhynchus in the Ecuadorian Pacific Ocean. Journal of Fish Biology, 87(6):1434–1448, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winter:2015:IVS

[1814] A. Winter, J. Pompert, A. Arkhipkin, and P. E. Brewin. Interannual variability in the skate assemblage on the South Patagonian shelf and slope. Journal of Fish Biology, 87(6):1449–1468, December 2015. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2015:DDS

[1815] F. C. Neat, F. Burns, E. Jones, and T. Blasdale. The diversity, distribution and status of deep-water elasmobranchs in the Rockall Trough, north-east Atlantic Ocean. Journal of Fish Biology, 87(6):1469–1488, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osgood:2015:RSR

[1816] G. J. Osgood and J. K. Baum. Reef sharks: recent advances in ecological understanding to inform conservation. Journal of Fish Biology, 87(6): 1489–1523, December 2015. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chabot:2016:MRF

[1817] D. Chabot, D. J. McKenzie, and J. F. Craig. Metabolic rate in fishes: definitions, methods and significance for conservation physiology. Journal of Fish Biology, 88(1):1–9, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:2016:OCR

[1818] J. A. Nelson. Oxygen consumption rate v. rate of energy utilization of fishes: a comparison and brief history of the two measurements. Journal of Fish Biology, 88(1):10–25, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2016:DSI

[1819] M. B. S. Svendsen, P. G. Bushnell, and J. F. Steffensen. Design and setup of intermittent-flow respirometry system for aquatic organisms. Journal of Fish Biology, 88(1):26–50, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2016:SVO

[1820] M. B. S. Svendsen, P. G. Bushnell, E. A. F. Christensen, and J. F. Steffensen. Sources of variation in oxygen consumption of aquatic animals demonstrated by simulated constant oxygen consumption and respirometers of different sizes. Journal of Fish Biology, 88(1):51–64, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodgers:2016:EMA

[1821] G. G. Rodgers, P. Tenzing, and T. D. Clark. Experimental methods in aquatic respirometry: the importance of mixing devices and accounting for background respiration. Journal of Fish Biology, 88(1):65–80, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chabot:2016:DSM

[1822] D. Chabot, J. F. Steffensen, and A. P. Farrell. The determination of standard metabolic rate in fishes. Journal of Fish Biology, 88(1):81–121, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norin:2016:MRM

[1823] T. Norin and T. D. Clark. Measurement and relevance of maximum metabolic rate in fishes. Journal of Fish Biology, 88(1):122–151, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chabot:2016:MSD

[1824] D. Chabot, R. Koenker, and A. P. Farrell. The measurement of specific dynamic action in fishes. Journal of Fish Biology, 88(1):152–172, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peck:2016:MRR

[1825] M. A. Peck and M. Moyano. Measuring respiration rates in marine fish larvae: challenges and advances. Journal of Fish Biology, 88(1):173–205, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefevre:2016:MOU

[1826] S. Lefevre, M. Bayley, and D. J. McKenzie. Measuring oxygen uptake in fishes with bimodal respiration. Journal of Fish Biology, 88(1):206– 231, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claireaux:2016:RFE

[1827] G. Claireaux and D. Chabot. Responses by fishes to environmental hypoxia: integration through Fry’s concept of aerobic metabolic scope. Journal of Fish Biology, 88(1):232–251, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Snyder:2016:ECV

[1828] S. Snyder, L. E. Nadler, J. S. Bayley, M. B. S. Svendsen, J. L. Johansen, P. Domenici, and J. F. Steffensen. Effect of closed v. intermittent-flow respirometry on hypoxia tolerance in the shiner perch Cymatogaster aggregata. Journal of Fish Biology, 88(1):252–264, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosewarne:2016:MMS

[1829] P. J. Rosewarne, J. M. Wilson, and J. C. Svendsen. Measuring maximum and standard metabolic rates using intermittent-flow respirometry: a student laboratory investigation of aerobic metabolic scope and environmental hypoxia in aquatic breathers. Journal of Fish Biology, 88 (1):265–283, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2016:RAT

[1830] J. D. Metcalfe, S. Wright, C. Tudorache, and R. P. Wilson. Recent advances in telemetry for estimating the energy metabolism of wild fishes. Journal of Fish Biology, 88(1):284–297, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2016:DIV

[1831] N. B. Metcalfe, T. E. Van Leeuwen, and S. S. Killen. Does individual variation in metabolic phenotype predict fish behaviour and performance? Journal of Fish Biology, 88(1):298–321, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2016:PPA

[1832] A. P. Farrell. Pragmatic perspective on aerobic scope: peaking, plummeting, pejus and apportioning. Journal of Fish Biology, 88(1):322–343, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Enders:2016:EEF

[1833] E. C. Enders and D. Boisclair. Effects of environmental fluctuations on fish metabolism: Atlantic salmon Salmo salar as a case study. Journal of Fish Biology, 88(1):344–358, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliason:2016:OUP

[1834] E. J. Eliason and A. P. Farrell. Oxygen uptake in Pacific salmon Oncorhynchus spp.: when ecology and physiology meet. Journal of Fish Biology, 88(1):359–388, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jorgensen:2016:MIF

[1835] C. Jørgensen, K. Enberg, and M. Mangel. Modelling and interpreting fish bioenergetics: a role for behaviour, life-history traits and survival tradeoffs. Journal of Fish Biology, 88(1):389–402, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lea:2016:KES

[1836] J. M. D. Lea, A. N. Keen, R. L. Nudds, and H. A. Shiels. Kinematics and energetics of swimming performance during acute warming in brown trout Salmo trutta. Journal of Fish Biology, 88(1):403–417, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefevre:2016:ITT

[1837] S. Lefevre, I. Findorf, M. Bayley, D. T. T. Huong, and T. Wang. Increased temperature tolerance of the air-breathing Asian swamp eel Monopterus albus after high-temperature acclimation is not explained by improved cardiorespiratory performance. Journal of Fish Biology, 88 (1):418–432, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lucas:2016:TCP

[1838] J. Lucas, A. Bonnieux, L. Lyphout, X. Cousin, P. Miramand, and C. Lefrançois. Trophic contamination by pyrolytic polycyclic aromatic hydrocarbons does not affect aerobic metabolic scope in zebrafish Danio rerio. Journal of Fish Biology, 88(1):433–442, January 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2016:RLM

[1839] Gordon McGregor Reid. Rosemary Lowe-McConnell 1921–2014 (Freshwater Biological Association and British Museum of Natural History).

Journal of Fish Biology, 88(2):443–447, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorn:2016:ESS

[1840] M. W. Thorn and Y. E. Morbey. Evidence for the secondary sexual development of the anal fin in female Kokanee salmon Oncorhynchus nerka. Journal of Fish Biology, 88(2):448–458, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teixeira:2016:TNS

[1841] T. F. Teixeira, A. L. Netto-Ferreira, J. L. O. Birindelli, and L. M. Sousa. Two new species of Hyphessobrycon (Characiformes: Characidae) from the headwaters of the Tapajós and Xingu River basins, Pará, Brazil. Journal of Fish Biology, 88(2):459–476, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jazdzewski:2016:HPU

[1842] M. Jażdżewski, L. Marszał, and M. Przybylski. Habitat preferences of Ukrainian brook lamprey Eudontomyzon mariae ammocoetes in the lowland rivers of Central Europe. Journal of Fish Biology, 88(2):477–491, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pavon-Munoz:2016:RDG

[1843] T. Pavón-Muñoz, R. Bejarano-Escobar, M. Blasco, G. Martín-Partido, and J. Francisco-Morcillo. Retinal development in the gilthead seabream Sparus aurata. Journal of Fish Biology, 88(2):492–507, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weigel:2016:NEA

[1844] E. G. Weigel, R. M. Tinghitella, and J. W. Boughman. No evidence for adjustment of maternal investment under alternative mate availability regimes. Journal of Fish Biology, 88(2):508–522, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeMartini:2016:CBS

[1845] E. E. DeMartini and K. G. Howard. Comparisons of body sizes at sexual maturity and at sex change in the parrotfishes of Hawaii: input needed for management regulations and stock assessments. Journal of Fish Biology, 88(2):523–541, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deary:2016:DOF

[1846] A. L. Deary, P. Pattrick, and N. A. Strydom. Development and ossification of the feeding apparatus in the larvae of two co-occurring species of kob (Sciaenidae), Argyrosomus japonicus and Argyrosomus inodorus, in South Africa. Journal of Fish Biology, 88(2):542–556, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medeiros:2016:STC

[1847] L. R. Medeiros, M. Elliott, and J. J. Nagler. Stressor timing, not cortisol, is an important embryo viability determinant in female rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 88(2):557–566, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leng:2016:MCE

[1848] X. Q. Leng, H. J. Du, C. J. Li, and H. Cao. Molecular characterization and expression pattern of dmrt1 in the immature Chinese sturgeon Acipenser sinensis. Journal of Fish Biology, 88(2):567–579, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2016:IIP

[1849] C. E. Adams, C. W. Bean, J. A. Dodd, A. Down, E. C. Etheridge, A. R. D. Gowans, O. Hooker, R. Knudsen, A. A. Lyle, I. J. Winfield, and K. Præbel. Inter and intra-population phenotypic and genotypic structuring in the European whitefish Coregonus lavaretus, a rare freshwater fish in Scotland. Journal of Fish Biology, 88(2):580–594, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stich:2016:PPP

[1850] D. S. Stich, G. B. Zydlewski, and J. D. Zydlewski. Physiological preparedness and performance of Atlantic salmon Salmo salar smolts in relation to behavioural salinity preferences and thresholds. Journal of Fish Biology, 88(2):595–617, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2016:ECI

[1851] B. Jonsson, N. Jonsson, and J. Albretsen. Environmental change influences the life history of salmon Salmo salar in the North Atlantic Ocean. Journal of Fish Biology, 88(2):618–637, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2016:USI

[1852] T. M. Evans and J. E. Bauer. Using stable isotopes and C:N ratios to examine the life-history strategies and nutritional sources of larval lampreys. Journal of Fish Biology, 88(2):638–654, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nguyen:2016:CFR

[1853] P. L. Nguyen, Z. J. Jackson, and D. L. Peterson. Comparison of fin ray sampling methods on white sturgeon Acipenser transmontanus growth and swimming performance. Journal of Fish Biology, 88(2):655–667, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Straube:2016:FRS

[1854] N. Straube, K. P. Lampert, M. F. Geiger, J. D. Weiß, and J. X. Kirchhauser. First record of second-generation facultative parthenogenesis in a vertebrate species, the whitespotted bambooshark Chiloscyllium plagiosum. Journal of Fish Biology, 88(2):668–675, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barry:2016:FCM

[1855] J. Barry, M. Newton, J. A. Dodd, M. C. Lucas, P. Boylan, and C. E. Adams. Freshwater and coastal migration patterns in the silver-stage eel Anguilla anguilla. Journal of Fish Biology, 88(2):676–689, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Witten:2016:PPD

[1856] P. E. Witten, M. A. G. Owen, R. Fontanillas, M. Soenens, C. McGurk, and A. Obach. A primary phosphorus-deficient skeletal phenotype in juvenile Atlantic salmon Salmo salar: the uncoupling of bone formation and mineralization. Journal of Fish Biology, 88(2):690–708, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camarillo-Sepulveda:2016:STF

[1857] N. Camarillo-Sepulveda, D. Hamoutene, L. Lush, K. Burt, H. Volkoff, and I. A. Fleming. Sperm traits in farmed and wild Atlantic salmon Salmo salar. Journal of Fish Biology, 88(2):709–717, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drost:2016:UTL

[1858] H. E. Drost, J. Fisher, F. Randall, D. Kent, E. C. Carmack, and A. P. Farrell. Upper thermal limits of the hearts of Arctic cod Boreogadus saida: adults compared with larvae. Journal of Fish Biology, 88(2): 718–726, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Day:2016:BBA

[1859] R. D. Day, F. Mueller, L. Carseldine, N. Meyers-Cherry, and I. R. Tibbetts. Ballistic beloniformes attacking through Snell’s Window. Journal of Fish Biology, 88(2):727–734, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakefield:2016:ECR

[1860] C. B. Wakefield, G. I. Moore, A. E. Bertram, M. Snow, and S. J. Newman. Extraordinary capture of a Randall’s snapper Randallichthys filamentosus in the temperate south-eastern Indian Ocean and its molecular phylogenetic relationship within the etelinae. Journal of Fish Biology, 88 (2):735–740, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harmon:2016:CPB

[1861] T. S. Harmon, T. Y. Kamerman, A. L. Corwin, and A. B. Sellas. Consecutive parthenogenetic births in a spotted eagle ray Aetobatus narinari. Journal of Fish Biology, 88(2):741–745, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2016:CFG

[1862] T. Yamamoto and S. Kitanishi. Comparison of the frequency of the growth hormone pseudogene between juvenile and adult female masu salmon Oncorhynchus masou. Journal of Fish Biology, 88(2):746–750, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cisneros-Montemayor:2016:AWA

[1863] A. M. Cisneros-Montemayor, K. West, I. S. Boiro, and A. C. J. Vincent. An assessment of West African seahorses in fisheries catch and trade. Journal of Fish Biology, 88(2):751–759, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baudouin:2016:COS

[1864] M. Baudouin, M. Marengo, A. Pere, J.-M. Culioli, M.-C. Santoni, B. Marchand, and E. D. H. Durieux. Comparison of otolith and scale readings for age and growth estimation of common dentex Dentex dentex. Journal of Fish Biology, 88(2):760–766, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez-Hernandez:2016:AAR

[1865] J. Sánchez-Hernández, H.-M. Gabler, and P.-A. Amundsen. From autochthonous to allochthonous resources: seasonal shifts in food use by stream-dwelling YOY Arctic charr Salvelinus alpinus through the ice-free season. Journal of Fish Biology, 88(2):767–773, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2016:CRP

[1866] E. L. Johnson, K. L. Weinersmith, and R. L. Earley. Changes in reproductive physiology of mangrove rivulus Kryptolebias marmoratus following exposure to environmentally relevant doses of ethinyl oestradiol. Journal of Fish Biology, 88(2):774–786, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ju:2016:DET

[1867] C. S. Ju, L. B. He, Y. Y. Pei, Y. Jiang, R. Huang, Y. M. Li, L. J. Liao, S. H. Jang, and Y. P. Wang. Differential expression of two Ctype lectins in grass carp Ctenopharyngodon idella and their response to grass carp reovirus. Journal of Fish Biology, 88(2):787–793, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Womble:2016:OAA

[1868] M. R. Womble and S. A. Bullard. Ocular anomaly in Atlantic midshipman Porichthys plectrodon (Batrachoidiformes: Batrachoididae) from the Mississippi Canyon, north-central Gulf of Mexico. Journal of Fish Biology, 88(2):794–799, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chaves:2016:PIS

[1869] L. C. T. Chaves, J. Hall, J. L. L. Feitosa, and I. M. Côté. Photoidentification as a simple tool for studying invasive lionfish Pterois volitans populations. Journal of Fish Biology, 88(2):800–804, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunn:2016:TUT

[1870] R. P. Dunn. Tool use by a temperate wrasse, California sheephead Semicossyphus pulcher. Journal of Fish Biology, 88(2):805–810, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dias:2016:RTP

[1871] P. J. Dias, C. B. Wakefield, D. V. Fairclough, G. Jackson, M. J. Travers, and M. Snow. Real-time PCR validation of visually identified snapper Chrysophrys auratus (Sparidae) eggs. Journal of Fish Biology, 88(2): 811–819, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pawson:2016:BES

[1872] Mike Pawson. Biology of European sea bass. Journal of Fish Biology, 88 (2):820–821, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:BRP

[1873] Ian J. Winfield. Book review: Plankton: wonders of the drifting world. Journal of Fish Biology, 88(2):822, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:BRF

[1874] Ian J. Winfield. Book review: Few and far between: on the trail of Britain’s rarest animals. Journal of Fish Biology, 88(2):823–824, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:BEC

[1875] Ian J. Winfield. Biology and ecology of carp. Journal of Fish Biology, 88 (2):825–826, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:FGT

[1876] Ian J. Winfield. Fishes: a guide to their diversity. Journal of Fish Biology, 88(2):827–828, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2016:E

[1877] Anonymous. Erratum. Journal of Fish Biology, 88(2):829, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dAuriac:2016:R

[1878] M. B. Anglès d’Auriac, H. A. Urke, and T. Kristensen. Referees 2015. Journal of Fish Biology, 88(2):830–836, February 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weigmann:2016:ACL

[1879] S. Weigmann. Annotated checklist of the living sharks, batoids and chimaeras (Chondrichthyes) of the world, with a focus on biogeographical diversity. Journal of Fish Biology, 88(3):837–1037, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See comment [2248] and reply [2249].

Geiger:2016:CGM

[1880] M. F. Geiger, C. Schreiner, G. B. Delmastro, and F. Herder. Combining geometric morphometrics with molecular genetics to investigate a putative hybrid complex: a case study with barbels Barbus spp. (Teleostei: Cyprinidae). Journal of Fish Biology, 88(3):1038–1055, March 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rey:2016:DRE

[1881] S. Rey, L. Ribas, D. Morera Capdevila, A. Callol, F. A. Huntingford, M. Pilarczyk, S. Kadri, and S. MacKenzie. Differential responses to environmental challenge by common carp Cyprinus carpio highlight the importance of coping style in integrative physiology. Journal of Fish Biology, 88(3):1056–1069, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belleggia:2016:TEY

[1882] M. Belleggia, N. Andrada, S. Paglieri, F. Cortés, A. M. Massa, D. E. Figueroa, and C. Bremec. Trophic ecology of yellownose skate Zearaja chilensis, a top predator in the south-western Atlantic Ocean. Journal of Fish Biology, 88(3):1070–1087, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cox:2016:PDA

[1883] C. A. Cox, J. W. Quinn, L. C. Lewis, S. R. Adams, and G. L. Adams. Population demographics of American eels Anguilla rostrata in two Arkansas, U.S.A., catchments that drain into the Gulf of Mexico. Journal of Fish Biology, 88(3):1088–1103, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovacic:2016:NSG

[1884] M. Kovačić and R. Šanda. A new species of Gobius (Perciformes: Gobiidae) from the Mediterranean Sea and the redescription of Gobius bucchichi. Journal of Fish Biology, 88(3):1104–1124, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weinersmith:2016:RBH

[1885] K. L. Weinersmith, A. F. Hanninen, A. Sih, R. McElreath, and R. L. Earley. The relationship between handling time and cortisol release rates changes as a function of brain parasite densities in California killifish Fundulus parvipinnis. Journal of Fish Biology, 88(3):1125–1142, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva:2016:MPB

[1886] M. E. Oliva, J. F. Espinola, and L. A. Ñacari. Metazoan parasites of Brama australis from southern Chile: a tool for stock discrimination? Journal of Fish Biology, 88(3):1143–1148, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rigby:2016:LHT

[1887] C. L. Rigby, W. T. White, J. J. Smart, and C. A. Simpfendorfer. Life histories of two deep-water Australian endemic elasmobranchs: Argus skate Dipturus polyommata and eastern spotted gummy shark Mustelus walkeri. Journal of Fish Biology, 88(3):1149–1174, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pujolar:2016:TCA

[1888] J. M. Pujolar, S. Vincenzi, L. Zane, and A. J. Crivelli. Temporal changes in allele frequencies in a small marble trout Salmo marmoratus population threatened by extreme flood events. Journal of Fish Biology, 88 (3):1175–1190, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laporte:2016:SPR

[1889] M. Laporte, J. Claude, P. Berrebi, P. Perret, and P. Magnan. Shape plasticity in response to water velocity in the freshwater blenny Salaria fluviatilis. Journal of Fish Biology, 88(3):1191–1203, March 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramirez:2016:TMF

[1890] J. L. Ramirez, L. F. Carvalho-Costa, P. C. Venere, D. C. Carvalho, W. P. Troy, and P. M. Galetti, Jr. Testing monophyly of the freshwater fish Leporinus (Characiformes, Anostomidae) through molecular analysis. Journal of Fish Biology, 88(3):1204–1214, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKelvey:2016:SLG

[1891] K. S. McKelvey, M. K. Young, W. L. Knotek, K. J. Carim, T. M. Wilcox, T. M. Padgett-Stewart, and M. K. Schwartz. Sampling large geographic areas for rare species using environmental DNA: a study of bull trout Salvelinus confluentus occupancy in western Montana. Journal of Fish Biology, 88(3):1215–1222, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domingo:2016:TSG

[1892] A. Domingo, R. Coelho, E. Cortes, B. Garcia-Cortes, F. Mas, J. Mejuto, P. Miller, A. Ramos-Cartelle, M. N. Santos, and K. Yokawa. Is the tiger shark Galeocerdo cuvier a coastal species? Expanding its distribution range in the Atlantic Ocean using at-sea observer data. Journal of Fish Biology, 88(3):1223–1228, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otteraa:2016:IDT

[1893] H. Otterå and O. T. Skilbrei. Influence of depth, time and human activity on detection rate of acoustic tags: a case study on two fish farms. Journal of Fish Biology, 88(3):1229–1235, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Calvino:2016:FRG

[1894] P. Calviño and F. Alonso. First record of the genus Jenynsia from marine water on the coast of Punta del Este, Maldonado, Uruguay (Cyprinodontiformes: Anablepidae). Journal of Fish Biology, 88(3):1236–1240, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kent:2016:LRW

[1895] D. Kent, H. E. Drost, J. Fisher, T. Oyama, and A. P. Farrell. Laboratory rearing of wild Arctic cod Boreogadus saida from egg to adulthood. Journal of Fish Biology, 88(3):1241–1248, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2016:CGH

[1896] B.-M. Kim, B.-Y. Lee, J.-H. Lee, J.-S. Rhee, and J.-S. Lee. Conservation of Hox gene clusters in the self-fertilizing fish Kryptolebias marmoratus (Cyprinodontiformes; Rivulidae). Journal of Fish Biology, 88(3):1249– 1256, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnell:2016:MIT

[1897] C. P. O’Connell and G. Leurs. A minimally invasive technique to assess several life-history characteristics of the endangered great hammerhead shark Sphyrna mokarran. Journal of Fish Biology, 88(3):1257–1264, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shiao:2016:SRD

[1898] J. C. Shiao, E. Y. Liu, and T. D. Sui. Up-and-down shift in residence depth of slickheads (Alepocephalidae) revealed by otolith stable oxygen isotopic composition. Journal of Fish Biology, 88(3):1265–1272, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brager:2016:SMA

[1899] Zs. Bräger, T. Moritz, A. C. Tsikliras, J. Gonzalvo, M. Radulović, and Á. Staszny. Scale morphometry allows discrimination of European sardine Sardina pilchardus and round sardinella Sardinella aurita and among their local populations. Journal of Fish Biology, 88(3):1273–1281, March 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krylov:2016:DCE

[1900] V. V. Krylov, Yu. V. Chebotareva, and Yu. G. Izyumov. Delayed consequences of extremely low-frequency magnetic fields and the influence of adverse environmental conditions on roach Rutilus rutilus embryos. Journal of Fish Biology, 88(4):1283–1300, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nissling:2016:FRA

[1901] A. Nissling, A. Thorsen, and F. F. G. da Silva. Fecundity regulation by atresia in turbot Scophthalmus maximus in the Baltic Sea. Journal of Fish Biology, 88(4):1301–1320, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellmann:2016:ISS

[1902] J. K. Hellmann, I. Y. Ligocki, C. M. O’Connor, A. R. Reddon, T. M. Farmer, S. E. Marsh-Rollo, S. Balshine, and I. M. Hamilton. The influence of status and the social environment on energy stores in a social fish. Journal of Fish Biology, 88(4):1321–1334, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Starks:2016:ELH

[1903] T. A. Starks, M. L. Miller, and J. M. Long. Early life history of three pelagic-spawning minnows Macrhybopsis spp. in the lower Missouri River. Journal of Fish Biology, 88(4):1335–1349, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koehn:2016:CMF

[1904] J. D. Koehn and S. J. Nicol. Comparative movements of four large fish species in a lowland river. Journal of Fish Biology, 88(4):1350–1368, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kide:2016:CBA

[1905] N. G. Kide, A. Dunz, J.-F. Agnèse, J. Dilyte, A. Pariselle, C. Carneiro, E. Correia, J. C. Brito, L. O. Yarba, Y. Kone, and J.-D. Durand. Cichlids of the Banc d’Arguin National Park, Mauritania: insight into the diversity of the genus Coptodon. Journal of Fish Biology, 88(4):1369– 1393, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stiassny:2016:RSC

[1906] M. L. J. Stiassny and H. Sakharova. Review of the smiliogastrin cyprinids of the Kwilu River (Kasai Basin, central Africa), revised diagnosis for Clypeobarbus (Cyprinidae: Cyprininae: Smiliogastrini) and description of a new species. Journal of Fish Biology, 88(4):1394–1412, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delpiani:2016:RBS

[1907] G. Delpiani. Reproductive biology of the southern thorny skate Amblyraja doellojuradoi (Chondrichthyes, Rajidae). Journal of Fish Biology, 88(4):1413–1429, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romero-Caicedo:2016:RPP

[1908] A. F. Romero-Caicedo, F. Galván-Magaña, A. Hernández-Herrera, and M. Carrera-Fernández. Reproductive parameters of the Pacific angel shark Squatina californica (Selachii: Squatinidae). Journal of Fish Biology, 88(4):1430–1440, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lteif:2016:PBE

[1909] M. Lteif, R. Mouawad, G. Khalaf, P. Lenfant, and M. Verdoit-Jarraya. Population biology of an endangered species: the common guitarfish Rhinobatos rhinobatos in Lebanese marine waters of the eastern Mediterranean Sea. Journal of Fish Biology, 88(4):1441–1459, April 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Overdyk:2016:RTP

[1910] L. M. Overdyk, H. E. Braid, A. M. Naaum, S. S. Crawford, and R. H. Hanner. Real-time PCR identification of lake whitefish Coregonus clupeaformis in the Laurentian Great Lakes. Journal of Fish Biology, 88 (4):1460–1474, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ligon:2016:USS

[1911] F. K. Ligon, R. J. Nakamoto, B. C. Harvey, and P. F. Baker. Use of streambed substrate as refuge by steelhead or rainbow trout Oncorhynchus mykiss during simulated freshets. Journal of Fish Biology, 88(4):1475–1485, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tewson:2016:DVA

[1912] L. H. Tewson, I. G. Cowx, and A. D. Nunn. Diel variations in the assemblage structure and foraging ecology of larval and 0+ year juvenile fishes in a man-made floodplain waterbody. Journal of Fish Biology, 88 (4):1486–1500, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oyafuso:2016:TST

[1913] Z. S. Oyafuso, R. J. Toonen, and E. C. Franklin. Temporal and spatial trends in prey composition of wahoo Acanthocybium solandri: a diet analysis from the central North Pacific Ocean using visual and DNA bar-coding techniques. Journal of Fish Biology, 88(4):1501–1523, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feiner:2016:DOT

[1914] Z. S. Feiner, D. P. Coulter, S. C. Guffey, and T. O. Höök. Does overwinter temperature affect maternal body composition and egg traits in yellow perch Perca flavescens? Journal of Fish Biology, 88(4):1524–1543, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laskowski:2016:BSA

[1915] K. L. Laskowski, C. T. Monk, G. Polverino, J. Alós, S. Nakayama, G. Staaks, T. Mehner, and R. Arlinghaus. Behaviour in a standardized assay, but not metabolic or growth rate, predicts behavioural variation in an adult aquatic top predator Esox lucius in the wild. Journal of Fish Biology, 88(4):1544–1563, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faulks:2016:AMH

[1916] L. K. Faulks and Ö. Östman. Adaptive major histocompatibility complex (MHC) and neutral genetic variation in two native Baltic Sea fishes (perch Perca fluviatilis and zander Sander lucioperca) with comparisons to an introduced and disease susceptible population in Australia (P. fluviatilis): assessing the risk of disease epidemics. Journal of Fish Biology, 88(4):1564–1583, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luo:2016:STE

[1917] Q. Luo, C. Wu, S. Sun, F. Lu, L. Xie, H. Zhao, X. Zhong, and Q. Zhou. The spatial-temporal expression and functional divergence of bach homologs in zebrafish Danio rerio. Journal of Fish Biology, 88(4):1584– 1597, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2016:PMD

[1918] M. Liu, Y.-Y. Wang, X.-J. Shan, B. Kang, and S.-X. Ding. Primary male development of two sequentially hermaphroditic groupers, Epinephelus akaara and Epinephelus awoara (Perciformes: Epinephelidae). Journal of Fish Biology, 88(4):1598–1613, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paczolt:2016:LRM

[1919] K. A. Paczolt, W. E. Martin, N. L. Ratterman, and A. G. Jones. A low rate of multiple maternity for pregnant male northern pipefish Syngnathus fuscus. Journal of Fish Biology, 88(4):1614–1619, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aylesworth:2016:NRJ

[1920] L. Aylesworth, J. M. Lawson, P. Laksanawimol, P. Ferber, and T.-L. Loh. New records of the Japanese seahorse Hippocampus mohnikei in Southeast Asia lead to updates in range, habitat and threats. Journal of Fish Biology, 88(4):1620–1630, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vanderpham:2016:HRS

[1921] J. P. Vanderpham, S. Nakagawa, A. M. Senior, and G. P. Closs. Habitatrelated specialization of lateral-line system morphology in a habitatgeneralist and a habitat-specialist New Zealand eleotrid. Journal of Fish Biology, 88(4):1631–1641, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naylor:2016:TTB

[1922] P. Naylor and D. M. P. Jacoby. Territoriality in the tompot blenny Parablennius gattorugine from photographic records. Journal of Fish Biology, 88(4):1642–1647, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:2016:LBS

[1923] M. R. Hughes, J. A. Dodd, P. S. Maitland, and C. E. Adams. Lake bathymetry and species occurrence predict the distribution of a lacustrine apex predator. Journal of Fish Biology, 88(4):1648–1654, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Picciulin:2016:SPS

[1924] M. Picciulin, M. Bolgan, A. B. Corò, G. Calcagno, and S. Malavasi. Sound production by the Shi drum Umbrina cirrosa and comparison with the brown meagre Sciaena umbra: a passive acoustic monitoring perspective. Journal of Fish Biology, 88(4):1655–1660, April 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolles:2016:FIR

[1925] J. W. Jolles, A. Manica, and N. J. Boogert. Food intake rates of inactive fish are positively linked to boldness in three-spined sticklebacks Gasterosteus aculeatus. Journal of Fish Biology, 88(4):1661–1668, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morris:2016:MEM

[1926] Corey Morris. Migration ecology of marine fishes. Journal of Fish Biology, 88(4):1669–1670, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2016:EFCa

[1927] Laith A. Jawad. Endemic fishes of Croatia. Journal of Fish Biology, 88 (4):1671–1672, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2016:ACD

[1928] Charles G. M. Paxton. Animal cannibalism: the dark side of evolution. Journal of Fish Biology, 88(4):1673–1674, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bean:2016:FFE

[1929] Colin W. Bean. Freshwater fisheries ecology. Journal of Fish Biology, 88(4):1675–1676, April 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boys:2016:PRA

[1930] C. A. Boys, W. Robinson, B. Miller, B. Pflugrath, L. J. Baumgartner, A. Navarro, R. Brown, and Z. Deng. A piecewise regression approach for determining biologically relevant hydraulic thresholds for the protection of fishes at river infrastructure. Journal of Fish Biology, 88(5):1677– 1692, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2016:MCE

[1931] W. T. Zhang, K. Liu, J. S. Xiang, L. Y. Zhang, W. J. Liu, Z. D. Dong, Y. Z. Li, H. L. Li, S. L. Chen, and N. Wang. Molecular cloning, expression of, and regulation by thyroid-hormone receptor α A in the half-smooth tongue sole Cynoglossus semilaevis during metamorphosis. Journal of Fish Biology, 88(5):1693–1707, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piercy:2016:HMA

[1932] A. N. Piercy, D. J. Murie, and J. J. Gelsleichter. Histological and morphological aspects of reproduction in the sandbar shark Carcharhinus plumbeus in the U.S. south-eastern Atlantic Ocean and Gulf of Mexico. Journal of Fish Biology, 88(5):1708–1730, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghazilou:2016:AVP

[1933] A. Ghazilou, M. R. Shokri, and W. Gladstone. Animal v. plant-based bait: does the bait type affect census of fish assemblages and trophic groups by baited remote underwater video (BRUV) systems? Journal of Fish Biology, 88(5):1731–1745, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoff:2016:IMN

[1934] G. R. Hoff. Identification of multiple nursery habitats of skates in the eastern Bering Sea. Journal of Fish Biology, 88(5):1746–1757, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colonello:2016:RPP

[1935] J. H. Colonello, F. Cortés, M. Belleggia, and A. M. Massa. Reproductive and population parameters of spiny dogfish Squalus acanthias in the south-western Atlantic Ocean. Journal of Fish Biology, 88(5):1758–1775, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Montie:2016:LTM

[1936] E. W. Montie, C. Kehrer, J. Yost, K. Brenkert, T. O’Donnell, and M. R. Denson. Long-term monitoring of captive red drum Sciaenops ocellatus reveals that calling incidence and structure correlate with egg deposition. Journal of Fish Biology, 88(5):1776–1795, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daros:2016:HRM

[1937] F. A. Daros, H. L. Spach, and A. T. Correia. Habitat residency and movement patterns of Centropomus parallelus juveniles in a subtropical estuarine complex. Journal of Fish Biology, 88(5):1796–1810, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tornabene:2016:MLS

[1938] L. Tornabene, S. Valdez, M. V. Erdmann, and F. L. Pezold. Multi-locus sequence data reveal a new species of coral reef goby (Teleostei: Gobiidae: Eviota), and evidence of Pliocene vicariance across the Coral Triangle. Journal of Fish Biology, 88(5):1811–1834, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahjoub:2016:MAS

[1939] M. S. Mahjoub, G. Dur, S. Souissi, F. G. Schmitt, and J. S. Hwang. Multifractal anisotropic swimming: the optimal foraging behaviour of grouper larvae. Journal of Fish Biology, 88(5):1835–1846, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parkinson:2016:RGS

[1940] K. L. Parkinson and D. J. Booth. Rapid growth and short life spans characterize pipefish populations in vulnerable seagrass beds. Journal of Fish Biology, 88(5):1847–1855, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rhodes:2016:GRH

[1941] K. L. Rhodes, B. M. Taylor, D. Hernandez-Ortiz, and J. Cuetos-Bueno. Growth and reproduction of the highfin grouper Epinephelus maculatus.

Journal of Fish Biology, 88(5):1856–1869, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Botta:2016:GHR

[1942] P. E. Botta, I. Simó, A. A. Sciara, and S. E. Arranz. Growth hormone receptors in the atherinid Odontesthes bonariensis: characterization and expression profile after fasting–refeeding and growth hormone administration. Journal of Fish Biology, 88(5):1870–1885, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2016:MGE

[1943] N. G. Andersen, D. Chabot, and C. S. Couturier. Modelling gastric evacuation in gadoids feeding on crustaceans. Journal of Fish Biology, 88(5):1886–1903, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cejko:2016:SSP

[1944] B. I. Cejko, M. Słowińska, S. Judycka, and R. K. Kowalski. Substrate specificity of proteolytic activity in the testes fluid and seminal plasma of the common carp Cyprinus carpio. Journal of Fish Biology, 88(5):1904– 1917, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tietze:2016:TOB

[1945] S. M. Tietze and G. W. Gerald. Trade-offs between salinity preference and antipredator behaviour in the euryhaline sailfin molly Poecilia latipinna. Journal of Fish Biology, 88(5):1918–1931, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tan:2016:GMM

[1946] M. P. Tan, A. F. J. Jamsari, and M. N. Siti Azizah. Genotyping of microsatellite markers to study genetic structure of the wild striped snakehead Channa striata in Malaysia. Journal of Fish Biology, 88(5):1932– 1948, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cao:2016:IEA

[1947] X. L. Cao, J. J. Chen, Y. Cao, G. X. Nie, and J. G. Su. Identification and expression analysis of the sting gene, a sensor of viral DNA, in common carp Cyprinus carpio. Journal of Fish Biology, 88(5):1949–1964, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:2016:DWM

[1948] S. D. Ross, H. Gislason, N. G. Andersen, P. Lewy, and J. R. Nielsen. The diet of whiting Merlangius merlangus in the western Baltic Sea. Journal of Fish Biology, 88(5):1965–1988, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa:2016:NDS

[1949] S. B. Correa and J. T. Anderson. A non-destructive sampling protocol for field studies of seed dispersal by fishes. Journal of Fish Biology, 88 (5):1989–2003, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lavergne:2016:DCE

[1950] E. Lavergne, U. Zajonz, F. Krupp, F. Naseeb, and M. S. Aideed. Diversity and composition of estuarine and lagoonal fish assemblages of Socotra Island, Yemen. Journal of Fish Biology, 88(5):2004–2026, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katwate:2016:PSN

[1951] U. Katwate, S. Jadhav, P. Kumkar, R. Raghavan, and N. Dahanukar. Pethia sanjaymoluri, a new species of barb (Teleostei: Cyprinidae) from the northern Western Ghats, India. Journal of Fish Biology, 88(5):2027– 2050, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ajemian:2016:ADF

[1952] M. J. Ajemian, M. D. Kenworthy, J. L. Sánchez-Lizaso, and J. Cebrian. Aggregation dynamics and foraging behaviour of striped red mullet Mullus surmuletus in the western Mediterranean Sea. Journal of Fish Biology, 88(5):2051–2059, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mavraki:2016:UNA

[1953] N. Mavraki, M. Georgiadis, C. Koutsikopoulos, and E. Tzanatos. Unravelling the nocturnal appearance of bogue Boops boops shoals in the anthropogenically modified shallow littoral. Journal of Fish Biology, 88 (5):2060–2066, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brooks:2016:UGI

[1954] J. L. Brooks, T. L. Guttridge, B. R. Franks, R. D. Grubbs, D. D. Chapman, S. H. Gruber, J. D. Dibattista, and K. A. Feldheim. Using genetic inference to re-evaluate the minimum longevity of the lemon shark Negaprion brevirostris. Journal of Fish Biology, 88(5):2067–2074, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mapp:2016:TDR

[1955] J. J. I. Mapp, M. H. Fisher, R. C. Atwood, G. D. Bell, M. K. Greco, S. Songer, and E. Hunter. Three-dimensional rendering of otolith growth using phase contrast synchrotron tomography. Journal of Fish Biology, 88(5):2075–2080, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goulding:2016:QVA

[1956] A. T. Goulding and A. P. Farrell. Quantification of ventricular β_{2}- adrenoceptor density and ligand binding affinity in wild sockeye salmon Oncorhynchus nerka smolts using a novel modification to the tritiated ligand technique. Journal of Fish Biology, 88(5):2081–2087, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houde:2016:GAM

[1957] A. L. S. Houde, C. C. Wilson, and T. E. Pitcher. Genetic architecture and maternal contributions of early-life survival in lake trout Salvelinus namaycush. Journal of Fish Biology, 88(5):2088–2094, May 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xu:2016:GWI

[1958] H. Xu, C. Li, Q. Zeng, I. Agrawal, X. Zhu, and Z. Gong. Genomewide identification of suitable zebrafish Danio rerio reference genes for normalization of gene expression data by RT-qPCR. Journal of Fish Biology, 88(6):2095–2110, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2016:TPT

[1959] P. M. Harrison, L. F. G. Gutowsky, E. G. Martins, D. A. Patterson, S. J. Cooke, and M. Power. Temporal plasticity in thermal-habitat selection of burbot Lota lota a diel-migrating winter-specialist. Journal of Fish Biology, 88(6):2111–2129, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pohls:2016:CCS

[1960] P. Pohls, L. González-Dávalos, O. Mora, A. Shimada, A. Varela-Echavarria, E. M. Toledo-Cuevas, and C. A. Martínez-Palacios. A complete chitinolytic system in the atherinopsid pike silverside Chirostoma estor: gene expression and activities. Journal of Fish Biology, 88(6): 2130–2143, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kessel:2016:TDL

[1961] S. T. Kessel, A. C. Hansell, S. H. Gruber, T. L. Guttridge, N. E. Hussey, and R. G. Perkins. Three decades of longlining in Bimini, Bahamas, reveals long-term trends in lemon shark Negaprion brevirostris (Carcharhinidae) catch per unit effort. Journal of Fish Biology, 88(6):2144– 2156, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2016:PSG

[1962] D. J. Jellyman and S. K. Crow. Population size, growth and movements of Anguilla australis in a small lake. Journal of Fish Biology, 88(6):2157– 2174, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manuel:2016:EAS

[1963] R. Manuel, J. G. J. Boerrigter, M. Cloosterman, M. Gorissen, G. Flik, R. van den Bos, and H. van de Vis. Effects of acute stress on aggression and the cortisol response in the African sharptooth catfish Clarias gariepinus: differences between day and night. Journal of Fish Biology, 88(6):2175–2187, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhu:2016:CIR

[1964] Y. Y. Zhu, W. X. Xing, S. J. Shan, S. Q. Zhang, Y. Q. Li, T. Li, L. An, and G. W. Yang. Characterization and immune response expression of the Rig-I-like receptor mda5 in common carp Cyprinus carpio. Journal of Fish Biology, 88(6):2188–2202, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tam:2016:TAE

[1965] J. C. Tam, J. S. Link, S. I. Large, B. Bogstad, A. Bundy, A. M. Cook, G. E. Dingsør, A. V. Dolgov, D. Howell, A. Kempf, J. K. Pinnegar, A. Rindorf, S. Schückel, A. F. Sell, and B. E. Smith. A trans-Atlantic examination of haddock Melanogrammus aeglefinus food habits. Journal of Fish Biology, 88(6):2203–2218, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2016:PSC

[1966] A. J. Jensen, M. Berg, G. Bremset, B. Finstad, N. A. Hvidsten, J. G. Jensås, B. O. Johnsen, and E. Lund. Passing a seawater challenge test is not indicative of hatchery-reared Atlantic salmon Salmo salar smolts performing as well at sea as their naturally produced conspecifics. Journal of Fish Biology, 88(6):2219–2235, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vollset:2016:EHS

[1967] K. W. Vollset, H. Skoglund, T. Wiers, and B. T. Barlaup. Effects of hydropeaking on the spawning behaviour of Atlantic salmon Salmo salar and brown trout Salmo trutta. Journal of Fish Biology, 88(6):2236– 2250, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yu:2016:ENC

[1968] Y. Yu, S. Wei, Z. Wang, X. Huang, Y. Huang, J. Cai, C. Li, and Q. Qin. Establishment of a new cell line from the snout tissue of golden pompano Trachinotus ovatus, and its application in virus susceptibility. Journal of Fish Biology, 88(6):2251–2262, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langkau:2016:SBA

[1969] M. C. Langkau, D. Clavé, M. B. Schmidt, and J. Borcherding. Spawning behaviour of Allis shad Alosa alosa: new insights based on imaging sonar data. Journal of Fish Biology, 88(6):2263–2274, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forman:2016:DDO

[1970] J. S. Forman, P. L. Horn, and D. W. Stevens. Diets of deepwater oreos (Oreosomatidae) and orange roughy Hoplostethus atlanticus. Journal of Fish Biology, 88(6):2275–2302, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2016:BDD

[1971] A. B. M. Moore, A. C. Henderson, E. D. Farrell, and L. B. Weekes. Biological data from a data-deficient shark: the Arabian smoothhound Mustelus mosis (Carcharhiniformes: Triakidae). Journal of Fish Biology, 88(6):2303–2307, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2016:Ca

[1972] Anonymous. Corrigendum. Journal of Fish Biology, 88(6):2308, June 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta:2016:FAH

[1973] M. Barletta, S. J. M. Blaber, and J. F. Craig. Fish and aquatic habitat conservation in South America. Journal of Fish Biology, 89(1):1–3, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2016:SCC

[1974] M. F. Costa and M. Barletta. Special challenges in the conservation of fishes and aquatic environments of South America. Journal of Fish Biology, 89(1):4–11, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2016:FBC

[1975] R. E. Reis, J. S. Albert, F. Di Dario, M. M. Mincarone, P. Petry, and L. A. Rocha. Fish biodiversity and conservation in South America. Journal of Fish Biology, 89(1):12–47, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casimiro:2016:BTL

[1976] A. C. R. Casimiro, D. A. Z. Garcia, A. P. Vidotto-Magnoni, J. R. S. Vitule, and M. L. Orsi. Biodiversity: is there light for native fish assemblages at the end of the Anthropocene tunnel? Journal of Fish Biology, 89(1): 48–49, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lobon-Cervia:2016:ERF

[1977] J. Lobón-Cerviá, R. Mazzoni, and C. F. Rezende. Effects of riparian forest removal on the trophic dynamics of a Neotropical stream fish assemblage. Journal of Fish Biology, 89(1):50–64, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jimenez-Segura:2016:FFF

[1978] L. F. Jiménez-Segura, G. Galvis-Vergara, P. Cala-Cala, C. A. García-Alzate, S. López-Casas, M. I. Ríos-Pulgarín, G. A. Arango, N. J. Mancera-Rodríguez, F. Gutiérrez-Bonilla, and R. Álvarez-León. Freshwater fish faunas, habitats and conservation challenges in the Caribbean river basins of north-western South America. Journal of Fish Biology, 89(1):65–101, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rios-Pulgarin:2016:RHC

[1979] M. I. Ríos-Pulgarín, M. Barletta, and N. J. Mancera-Rodríguez. The role of the hydrological cycle on the distribution patterns of fish assemblages in an Andean stream. Journal of Fish Biology, 89(1):102–130, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rios-Pulgarin:2016:HCE

[1980] M. I. Ríos-Pulgarín, M. Barletta, and N. J. Mancera-Rodriguez. Hydrological cycle effects on the aquatic community in a Neotropical stream of the Andean Piedmont during the 2007–2010 ENSO events. Journal of Fish Biology, 89(1):131–156, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Casas:2016:PMM

[1981] S. López-Casas, L. F. Jiménez-Segura, A. A. Agostinho, and C. M. Pérez. Potamodromous migrations in the Magdalena River basin: bimodal reproductive patterns in neotropical rivers. Journal of Fish Biology, 89(1): 157–171, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Machado-Allison:2016:CAE

[1982] A. Machado-Allison. The conservation of aquatic ecosystems of the Orinoco River basin. Journal of Fish Biology, 89(1):172–173, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lasso:2016:FAH

[1983] C. A. Lasso, A. Machado-Allison, and D. C. Taphorn. Fishes and aquatic habitats of the Orinoco River Basin: diversity and conservation. Journal of Fish Biology, 89(1):174–191, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Val:2016:EDF

[1984] A. L. Val, P. M. Fearnside, and V. M. F. Almeida-Val. Environmental disturbances and fishes in the Amazon. Journal of Fish Biology, 89(1): 192–193, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ropke:2016:SDF

[1985] C. P. Röpke, S. A. Amadio, K. O. Winemiller, and J. Zuanon. Seasonal dynamics of the fish assemblage in a floodplain lake at the confluence of the Negro and Amazon Rivers. Journal of Fish Biology, 89(1):194– 212, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sacramento:2016:DTD

[1986] P. A. Sacramento, G. I. Manetta, and E. Benedito. Diet-tissue discrimination factors (Δ ^{13}C and Δ ^{15}N) and turnover rate in somatic tissues of a neotropical detritivorous fish on C_{3} and C_{4} diets. Journal of Fish Biology, 89(1):213–219, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2016:EMB

[1987] B. G. Marshall, B. R. Forsberg, M. Thomé-Souza, R. Peleja, M. Z. Moreira, and C. E. C. Freitas. Evidence of mercury biomagnification in the food chain of the cardinal tetra Paracheirodon axelrodi (Osteichthyes: Characidae) in the Rio Negro, central Amazon, Brazil. Journal of Fish Biology, 89(1):220–240, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersen:2016:RAS

[1988] T. A. Petersen, S. M. Brum, F. Rossoni, G. F. V. Silveira, and L. Castello. Recovery of Arapaima sp. populations by community-based management in floodplains of the Purus River, Amazon. Journal of Fish Biology, 89 (1):241–248, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Souza-Araujo:2016:MMM

[1989] J. Souza-Araujo, T. Giarrizzo, M. O. Lima, and M. B. G. Souza. Mercury and methyl mercury in fishes from Bacajá River (Brazilian Amazon): evidence for bioaccumulation and biomagnification. Journal of Fish Biology, 89(1):249–263, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Val:2016:VEI

[1990] A. L. Val, M. de N. Paula-Silva, V. M. F. Almeida-Val, and C. M. Wood. In vitro effects of increased temperature and decreased pH on blood oxygen affinity of 10 fish species of the Amazon. Journal of Fish Biology, 89(1):264–279, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petry:2016:FCS

[1991] A. C. Petry, T. F. R. Guimarães, F. M. Vasconcellos, S. M. Hartz, F. G. Becker, R. S. Rosa, G. Goyenola, E. P. Caramaschi, J. M. Díaz de Astarloa, L. M. Sarmento-Soares, J. P. Vieira, A. M. Garcia, F. Teixeira de Mello, F. A. G. de Melo, M. Meerhoff, J. L. Attayde, R. F. Menezes, N. Mazzeo, and F. Di Dario. Fish composition and species richness in eastern South American coastal lagoons: additional support for the freshwater ecoregions of the world. Journal of Fish Biology, 89(1):280–314, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fontoura:2016:AFC

[1992] N. F. Fontoura, J. P. Vieira, F. G. Becker, L. R. Rodrigues, L. R. Malabarba, U. H. Schulz, O. O. Möller, A. M. Garcia, and F. S. Vilella. Aspects of fish conservation in the upper Patos Lagoon basin. Journal of Fish Biology, 89(1):315–336, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ceni:2016:FAF

[1993] G. Ceni, N. F. Fontoura, and H. N. Cabral. The freshwater artisanal fishery of patos lagoon. Journal of Fish Biology, 89(1):337–354, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gogola:2016:RST

[1994] T. M. Gogola, P. S. Daga, É. A. Gubiani, P. L. R. da Silva, and P. V. Sanches. The role of submerged trees in the early development of fishes in a Neotropical reservoir. Journal of Fish Biology, 89(1):355–368, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Habit:2016:CFF

[1995] E. Habit and V. Cussac. Conservation of the freshwater fauna of Patagonia: an alert to the urgent need for integrative management and sustainable development. Journal of Fish Biology, 89(1):369–370, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lessa:2016:AGM

[1996] R. Lessa, H. A. Andrade, K. L. De Lima, and F. M. Santana. Age and growth of the midwater crocodile shark Pseudocarcharias kamoharai. Journal of Fish Biology, 89(1):371–385, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weyl:2016:RCG

[1997] O. L. F. Weyl, V. S. Daga, B. R. Ellender, and J. R. S. Vitule. A review of Clarias gariepinus invasions in Brazil and South Africa. Journal of Fish Biology, 89(1):386–402, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crichigno:2016:DIC

[1998] S. Crichigno, P. Cordero, G. Blasetti, and V. Cussac. Dispersion of the invasive common carp Cyprinus carpio in southern South America: changes and expectations, westward and southward. Journal of Fish Biology, 89(1):403–416, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maiztegui:2016:ISC

[1999] T. Maiztegui, C. R. M. Baigún, J. R. Garcia de Souza, P. Minotti, and D. C. Colautti. Invasion status of the common carp Cyprinus carpio in inland waters of Argentina. Journal of Fish Biology, 89(1):417–430, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2016:RST

[2000] M. A. Azevedo, C. B. Fialho, and L. R. Malabarba. Reproductive strategies in two inseminating species of Glandulocaudini, Mimagoniates microlepis and Mimagoniates rheocharis (Characiformes: Characidae: Stevardiinae). Journal of Fish Biology, 89(1):431–444, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valentin:2016:ODD

[2001] F. N. Valentin, S. R. Batlouni, N. F. Nascimento, R. C. Silva, B. Manzini, C. C. Hilbig, M. Pereira-Santos, and L. S. O. Nakaghi. Ovarian differentiation and development in cachara Pseudoplatystoma fasciatum. Journal of Fish Biology, 89(1):445–466, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrade:2016:TKT

[2002] M. C. Andrade, M. Jégu, and T. Giarrizzo. Tometes kranponhah and Tometes ancylorhynchus (Characiformes: Serrasalmidae), two new phytophagous serrasalmids, and the first Tometes species described from the Brazilian Shield. Journal of Fish Biology, 89(1):467–494, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2016:RHM

[2003] N. A. Menezes, S. H. Weitzman, and T. F. Teixeira. Redescription of Hysteronotus megalostomus (Characiformes: Characidae: Stevardiinae), a poorly known characid from tributaries of the Rio São Fancisco, Brazil with comments on the conservation of the species. Journal of Fish Biology, 89(1):495–509, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marinho:2016:NBT

[2004] M. M. F. Marinho, F. C. P. Dagosta, P. Camelier, and O. T. Oyakawa. A name for the ‘blueberry tetra’, an aquarium trade popular species of Hyphessobrycon Durbin (Characiformes, Characidae), with comments on fish species descriptions lacking accurate type locality. Journal of Fish Biology, 89(1):510–521, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Restrepo-Escobar:2016:VBG

[2005] N. Restrepo-Escobar, J. C. Hurtado-Alarcón, N. J. Mancera-Rodríguez, and E. J. Márquez. Variations of body geometry in Brycon henni (Teleostei: Characiformes, Bryconidae) in different rivers and streams. Journal of Fish Biology, 89(1):522–528, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Restrepo-Escobar:2016:MMC

[2006] N. Restrepo-Escobar, J. D. Rangel-Medrano, N. J. Mancera-Rodríguez, and E. J. Márquez. Molecular and morphometric characterization of two dental morphs of Saccodon dariensis (Parodontidae). Journal of Fish Biology, 89(1):529–536, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:2016:REF

[2007] S. J. M. Blaber and M. Barletta. A review of estuarine fish research in South America: what has been achieved and what is the future for sustainability and conservation? Journal of Fish Biology, 89(1):537– 568, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaureguizar:2016:FDR

[2008] A. J. Jaureguizar, A. Solari, F. Cortés, A. C. Milessi, M. I. Militelli, M. D. Camiolo, M. Luz Clara, and M. García. Fish diversity in the Río de la Plata and adjacent waters: an overview of environmental influences on its spatial and temporal structure. Journal of Fish Biology, 89(1):569– 600, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saint-Paul:2016:NHA

[2009] U. Saint-Paul and H. Schneider. The need for a holistic approach in mangrove-related fisheries research: a specific review of the German and Brazilian research project MADAM. Journal of Fish Biology, 89(1):601– 618, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lima:2016:CCI

[2010] A. R. A. Lima, M. Barletta, M. F. Costa, J. A. A. Ramos, D. V. Dantas, P. A. M. C. Melo, A. K. S. Justino, and G. V. B. Ferreira. Changes in the composition of ichthyoplankton assemblage and plastic debris in mangrove creeks relative to moon phases. Journal of Fish Biology, 89 (1):619–640, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa:2016:STD

[2011] B. Corrêa and M. Vianna. Spatial and temporal distribution patterns of the silver mojarra Eucinostomus argenteus (Perciformes: Gerreidae) in a tropical semi-enclosed bay. Journal of Fish Biology, 89(1):641–660, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sales:2016:DJR

[2012] N. S. Sales, T. L. P. Dias, A. Baeta, and A. L. M. Pessanha. Dependence of juvenile reef fishes on semi-arid hypersaline estuary microhabitats as nurseries. Journal of Fish Biology, 89(1):661–679, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figueiredo:2016:CST

[2013] G. G. A. A. Figueiredo and A. L. M. Pessanha. Comparative study of trophic organization of juvenile fish assemblages of three tidal creeks in a tropical semi-arid estuary. Journal of Fish Biology, 89(1):680–695, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos:2016:SSO

[2014] J. A. A. Ramos, M. Barletta, D. V. Dantas, and M. F. Costa. Seasonal and spatial ontogenetic movements of Gerreidae in a Brazilian tropical estuarine ecocline and its application for nursery habitat conservation. Journal of Fish Biology, 89(1):696–712, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis-Filho:2016:MMP

[2015] J. A. Reis-Filho and T. Giarrizzo. Microgobius meeki as a potential bio-indicator of habitat disturbance in shallow estuarine areas: a useful tool for the assessment of estuarine quality. Journal of Fish Biology, 89(1):713–734, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prestrelo:2016:BAE

[2016] L. Prestrelo and C. Monteiro-Neto. Before–after environmental impact assessment of an artificial channel opening on a south-western Atlantic choked lagoon system. Journal of Fish Biology, 89(1):735–752, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aschenbrenner:2016:STV

[2017] A. Aschenbrenner, B. P. Ferreira, and J. R. Rooker. Spatial and temporal variability in the otolith chemistry of the Brazilian snapper Lutjanus alexandrei from estuarine and coastal environments. Journal of Fish Biology, 89(1):753–769, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MontAlverne:2016:TSF

[2018] R. Mont’Alverne, P. E. R. Pereyra, and A. M. Garcia. Trophic segregation of a fish assemblage along lateral depth gradients in a subtropical coastal lagoon revealed by stable isotope analyses. Journal of Fish Biology, 89(1):770–792, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dantas:2016:HUC

[2019] D. V. Dantas and M. Barletta. Habitat use by Centropomus undecimalis in a rocky area of estuarine beach in north-east Brazil. Journal of Fish Biology, 89(1):793–803, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loureiro:2016:EHD

[2020] S. N. Loureiro, J. A. Reis-Filho, and T. Giarrizzo. Evidence for habitatdriven segregation of an estuarine fish assemblage. Journal of Fish Biology, 89(1):804–820, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2016:SPD

[2021] D. R. da Silva, Jr., R. Paranhos, and M. Vianna. Spatial patterns of distribution and the influence of seasonal and abiotic factors on demersal ichthyofauna in an estuarine tropical bay. Journal of Fish Biology, 89 (1):821–846, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Passos:2016:EIF

[2022] C. V. B. Passos, N. N. Fabré, A. C. M. Malhado, V. S. Batista, and R. J. Ladle. Estuarization increases functional diversity of demersal fish assemblages in tropical coastal ecosystems. Journal of Fish Biology, 89 (1):847–862, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa-Herrera:2016:FSM

[2023] T. Correa-Herrera, L. F. Jiménez-Segura, and M. Barletta. Fish species from a micro-tidal delta in the Caribbean Sea. Journal of Fish Biology, 89(1):863–875, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bueno:2016:ESA

[2024] L. S. Bueno, A. A. Bertoncini, C. C. Koenig, F. C. Coleman, M. O. Freitas, J. R. Leite, T. F. De Souza, and M. Hostim-Silva. Evidence for spawning aggregations of the endangered Atlantic goliath grouper Epinephelus itajara in southern Brazil. Journal of Fish Biology, 89(1): 876–889, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta:2016:EDO

[2025] M. Barletta, F. J. A. Cysneiros, and A. R. A. Lima. Effects of dredging operations on the demersal fish fauna of a South American tropical– subtropical transition estuary. Journal of Fish Biology, 89(1):890–920, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franco:2016:RFA

[2026] A. C. S. Franco, M. C. N. Ramos Chaves, M. P. B. Castel-Branco, and L. Neves Dos Santos. Responses of fish assemblages of sandy beaches to different anthropogenic and hydrodynamic influences. Journal of Fish Biology, 89(1):921–938, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cordeiro:2016:SDN

[2027] C. A. M. M. Cordeiro, T. C. Mendes, A. R. Harborne, and C. E. L. Ferreira. Spatial distribution of nominally herbivorous fishes across environmental gradients on Brazilian rocky reefs. Journal of Fish Biology, 89(1):939–958, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2016:OSF

[2028] R. R. de S. Oliveira, R. M. Macieira, and T. Giarrizzo. Ontogenetic shifts in fishes between vegetated and unvegetated tidepools: assessing the effect of physical structure on fish habitat selection. Journal of Fish Biology, 89(1):959–976, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lessa:2016:GDC

[2029] R. Lessa and F. M. Santana. Growth of the dolphinfish Coryphaena hippurus from north-eastern Brazil with an appraisal of the efficacy of scales and otoliths for ageing. Journal of Fish Biology, 89(1):977–989, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2016:MFA

[2030] T. Simon, H. T. Pinheiro, R. L. Moura, A. Carvalho-Filho, L. A. Rocha, A. S. Martins, E. Mazzei, R. B. Francini-Filho, G. M. Amado-Filho, and J.-C. Joyeux. Mesophotic fishes of the abrolhos shelf, the largest reef ecosystem in the South Atlantic. Journal of Fish Biology, 89(1):990– 1001, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cardoso:2016:DDC

[2031] L. G. Cardoso and M. Haimovici. Density-dependent changes in the feeding behaviour of Macrodon atricauda of southern Brazil. Journal of Fish Biology, 89(1):1002–1008, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marques:2016:AGM

[2032] S. Marques and B. P. Ferreira. Age and growth of the mutton hamlet Alphestes afer, with a review of the size and age of sex change among epinephelids. Journal of Fish Biology, 89(1):1009–1025, July 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santander-Neto:2016:RBU

[2033] J. Santander-Neto, M. L. G. Araújo, and R. P. Lessa. Reproductive biology of Urotrygon microphthalmum (Batoidea: Urotrygonidae) from north-eastern Brazil, tropical west Atlantic Ocean. Journal of Fish Biology, 89(1):1026–1042, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araujo:2016:RBS

[2034] P. R. V. Araújo, M. C. Oddone, and G. Velasco. Reproductive biology of the stingrays, Myliobatis goodei and Myliobatis ridens (Chondrichthyes:

Myliobatidae), in southern Brazil. Journal of Fish Biology, 89(1):1043– 1067, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cussac:2016:FFP

[2035] V. E. Cussac, E. Habit, J. Ciancio, M. A. Battini, C. Riva Rossi, J. P. Barriga, C. Baigún, and S. Crichigno. Freshwater fishes of Patagonia: conservation and fisheries. Journal of Fish Biology, 89(1):1068–1097, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Possatto:2016:MST

[2036] F. E. Possatto, M. K. Broadhurst, H. L. Spach, K. O. Winemiller, R. B. Millar, K. M. Santos, and M. R. Lamour. Mapping the spatio-temporal distribution of threatened batoids to improve conservation in a subtropical estuary. Journal of Fish Biology, 89(1):1098–1104, July 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rolim:2016:NRB

[2037] F. A. Rolim, M. M. Rotundo, and T. Vaske-Júnior. Notes on the reproductive biology of the Brazilian electric ray Narcine brasiliensis (Elasmobranchii: Narcinidae). Journal of Fish Biology, 89(1):1105–1111, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunes:2016:NTE

[2038] J. L. S. Nunes, G. Rincon, N. M. Piorski, and A. P. B. Martins. Nearterm embryos in a Pristis pristis (Elasmobranchii: Pristidae) from Brazil. Journal of Fish Biology, 89(1):1112–1120, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinheiro:2016:TS

[2039] H. T. Pinheiro. Thiony Simon 1985–2016. Journal of Fish Biology, 89 (1):1121–1123, July 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossouw:2016:AMP

[2040] C. Rossouw, S. P. Wintner, and A. E. Bester-Van Der Merwe. Assessing multiple paternity in three commercially exploited shark species: Mustelus mustelus, Carcharhinus obscurus and Sphyrna lewini. Journal of Fish Biology, 89(2):1125–1141, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrnes:2016:IPD

[2041] E. E. Byrnes and C. Brown. Individual personality differences in Port Jackson sharks Heterodontus portusjacksoni. Journal of Fish Biology, 89 (2):1142–1157, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breve:2016:SMS

[2042] N. W. P. Brevé, H. V. Winter, H. M. J. Van Overzee, E. D. Farrell, and P. A. Walker. Seasonal migration of the starry smooth-hound shark Mustelus asterias as revealed from tag-recapture data of an angler-led tagging programme. Journal of Fish Biology, 89(2):1158–1177, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2016:EBN

[2043] C. A. Blake and C. R. Gabor. Exploratory behaviour and novel predator recognition: behavioural correlations across contexts. Journal of Fish Biology, 89(2):1178–1189, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2016:AJA

[2044] S. A. M. Elliott, P. A. Ahti, M. R. Heath, W. R. Turrell, and D. M. Bailey. An assessment of juvenile Atlantic cod Gadus morhua distribution and growth using diver operated stereo-video surveys. Journal of Fish Biology, 89(2):1190–1207, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kyne:2016:RPR

[2045] P. M. Kyne, A. J. Courtney, I. P. Jacobsen, and M. B. Bennett. Reproductive parameters of rhinobatid and urolophid batoids taken as by-catch in the Queensland (Australia) east coast otter-trawl fishery. Journal of Fish Biology, 89(2):1208–1226, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khan:2016:SME

[2046] U. Khan, K. Seyhan, N. Başçinar, and N. S. Başçinar. Satiation meal and the effects of meal and body sizes on gastric evacuation rate in brook trout Salvelinus fontinalis fed commercial pellets. Journal of Fish Biology, 89(2):1227–1238, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alves:2016:DDD

[2047] V. E. N. Alves, J. Patrício, M. Dolbeth, A. Pessanha, A. R. T. Palma, E. W. Dantas, and A. L. Vendel. Do different degrees of human activity affect the diet of Brazilian silverside Atherinella brasiliensis? Journal of Fish Biology, 89(2):1239–1257, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galvan:2016:ETS

[2048] D. E. Galván, J. Jañez, and A. J. Irigoyen. Estimating tissue-specific discrimination factors and turnover rates of stable isotopes of nitrogen and carbon in the smallnose fanskate Sympterygia bonapartii (Rajidae). Journal of Fish Biology, 89(2):1258–1270, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schemmel:2016:RLH

[2049] E. M. Schemmel, M. K. Donovan, C. Wiggins, M. Anzivino, and A. M. Friedlander. Reproductive life history of the introduced peacock grouper Cephalopholis argus in Hawaii. Journal of Fish Biology, 89(2):1271– 1284, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohara:2016:BFH

[2050] W. M. Ohara, I. D. Da Costa, and M. L. Fonseca. Behaviour, feeding habits and ecology of the blind catfish Phreatobius sanguijuela (Ostariophysi: Siluriformes). Journal of Fish Biology, 89(2):1285–1301, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

F:2016:NLC

[2051] A. Polanco F., A. Acero P., and R. Betancur-R. No longer a circumtropical species: revision of the lizardfishes in the Trachinocephalus myops species complex, with description of a new species from the Marquesas Islands. Journal of Fish Biology, 89(2):1302–1323, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wong:2016:AOI

[2052] J. Y. Wong, C. Chu, V. C. Chong, S. K. Dhillon, and K. H. Loh. Automated otolith image classification with multiple views: an evaluation on Sciaenidae. Journal of Fish Biology, 89(2):1324–1344, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:2016:HDA

[2053] W. A. Thompson, T. M. Rodela, and J. G. Richards. Hardness does not affect the physiological responses of wild and domestic strains of diploid and triploid rainbow trout Oncorhynchus mykiss to short-term exposure to pH 9.5. Journal of Fish Biology, 89(2):1345–1358, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gould:2016:LHS

[2054] A. L. Gould, K. E. Dougan, S. T. Koenigbauer, and P. V. Dunlap. Life history of the symbiotically luminous cardinalfish Siphamia tubifer (Perciformes: Apogonidae). Journal of Fish Biology, 89(2):1359–1377, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langille:2016:MPS

[2055] B. L. Langille, R. Perry, D. Keefe, O. Barker, and H. D. Marshall. Mitochondrial population structure and post-glacial dispersal of longnose sucker Catostomus catostomus in Labrador, Canada: evidence for multiple refugial origins and limited ongoing gene flow. Journal of Fish Biology, 89(2):1378–1392, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulson:2016:BCT

[2056] P. G. Coulson, N. G. Hall, and I. C. Potter. Biological characteristics of three co-occurring species of armorhead from different genera vary markedly from previous results for the pentacerotidae. Journal of Fish Biology, 89(2):1393–1418, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bezerra:2016:VLF

[2057] N. P. A. Bezerra, P. Travassos, and F. H. V. Hazin. Vulnerability to longline fisheries of three hammerhead shark Sphyrna species in the southwestern and equatorial Atlantic Ocean. Journal of Fish Biology, 89(2): 1419–1433, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2016:CEA

[2058] L. B. He, H. Wang, L. F. Luo, S. H. Jiang, L. Y. Liu, Y. M. Li, R. Huang, L. J. Liao, Z. Y. Zhu, and Y. P. Wang. Characterization, expression analysis and localization pattern of toll-like receptor 1 (tlr1) and toll-like receptor 2 (tlr2) genes in grass carp Ctenopharyngodon idella. Journal of Fish Biology, 89(2):1434–1440, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fajkowska:2016:EDV

[2059] M. Fajkowska, M. Rzepkowska, D. Adamek, T. Ostaszewska, and M. Szczepkowski. Expression of dmrt1 and vtg genes during gonad formation, differentiation and early maturation in cultured Russian sturgeon Acipenser gueldenstaedtii. Journal of Fish Biology, 89(2):1441–1449, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa-Pereira:2016:FMT

[2060] R. Costa-Pereira, M. S. Araújo, F. Paiva, and L. E. R. Tavares. Functional morphology of the tetra fish Astyanax lacustris differs between divergent habitats in the Pantanal wetlands. Journal of Fish Biology, 89 (2):1450–1458, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira-Rodriguez:2016:APE

[2061] N. Ferreira-Rodriguez, J. Gessner, and I. Pardo. Assessing the potential of the European Atlantic sturgeon Acipenser sturio to control bivalve invasions in Europe. Journal of Fish Biology, 89(2):1459–1465, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramee:2016:FIE

[2062] S. W. Ramee and P. J. Allen. Freshwater influences on embryos, hatching and larval survival of euryhaline Gulf killifish Fundulus grandis and potential constraints on habitat distribution. Journal of Fish Biology, 89 (2):1466–1472, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henry:2016:SEL

[2063] L.-A. Henry, M. F. W. Stehmann, L. De Clippele, H. S. Findlay, N. Golding, and J. M. Roberts. Seamount egg-laying grounds of the deep-water skate Bathyraja richardsoni. Journal of Fish Biology, 89(2):1473–1481, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2016:TGM

[2064] B. Jonsson and N. Jonsson. Trans-generational maternal effect: temperature influences egg size of the offspring in Atlantic salmon Salmo salar. Journal of Fish Biology, 89(2):1482–1487, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pan:2016:HTD

[2065] Y. K. Pan, R. Ern, and A. J. Esbaugh. Hypoxia tolerance decreases with body size in red drum Sciaenops ocellatus. Journal of Fish Biology, 89 (2):1488–1493, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konstantinidis:2016:NRL

[2066] P. Konstantinidis, E. J. Hilton, and A. C. Matarese. New records of larval stages of the eel cod genus Muraenolepis Günther 1880 (Gadiformes: Muraenolepididae) from the western Antarctic Peninsula. Journal of Fish Biology, 89(2):1494–1500, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Booth:2016:AHS

[2067] D. J. Booth. Ability to home in small site-attached coral reef fishes. Journal of Fish Biology, 89(2):1501–1506, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2016:FBE

[2068] Michael Elliott. Flatfishes: biology and exploitation. second edition. Journal of Fish Biology, 89(2):1507–1508, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2016:EFCb

[2069] Laith A. Jawad. Ecology of fishes on coral reefs. Journal of Fish Biology, 89(2):1509–1510, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:2016:BRF

[2070] Nigel Milner. Book review: Fish atlas of the Celtic Sea, North Sea, and Baltic Sea. Journal of Fish Biology, 89(2):1511–1512, August 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malek:2016:TSC

[2071] A. J. Malek, J. S. Collie, and D. L. Taylor. Trophic structure of a coastal fish community determined with diet and stable isotope analyses. Journal of Fish Biology, 89(3):1513–1536, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figueroa:2016:CAS

[2072] E. Figueroa, I. Valdebenito, O. Merino, A. Ubilla, J. Risopatrón, and J. G. Farias. Cryopreservation of Atlantic salmon Salmo salar sperm: effects on sperm physiology. Journal of Fish Biology, 89(3):1537–1550, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Varella:2016:TPN

[2073] H. R. Varella, J. Zuanon, S. O. Kullander, and H. López-Fernández. Teleocichla preta, a new species of cichlid from the Rio Xingu Basin in Brazil (Teleostei: Cichlidae). Journal of Fish Biology, 89(3):1551–1569, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cochran:2016:PSW

[2074] J. E. M. Cochran, R. S. Hardenstine, C. D. Braun, G. B. Skomal, S. R. Thorrold, K. Xu, M. G. Genton, and M. L. Berumen. Population structure of a whale shark Rhincodon typus aggregation in the Red Sea. Journal of Fish Biology, 89(3):1570–1582, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacob:2016:FDN

[2075] H. Jacob, M. Metian, R. M. Brooker, E. Duran, N. Nakamura, N. Roux, P. Masanet, O. Soulat, and D. Lecchini. First description of the neuroanatomy of a larval coral reef fish Amphiprion ocellaris. Journal of Fish Biology, 89(3):1583–1591, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ito:2016:SSF

[2076] S. Ito, H. Iwao, J. Sakata, M. Inoue, K. Omori, and Y. Yanagisawa. Simultaneous spawning by female stream goby Rhinogobius sp. and the association with brood cannibalism by nesting males. Journal of Fish Biology, 89(3):1592–1602, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siegenthaler:2016:ARC

[2077] A. Siegenthaler, P. R. W. Niemantsverdriet, M. Laterveer, and I. M. A. Heitkönig. Aversive responses of captive sandbar sharks Carcharhinus plumbeus to strong magnetic fields. Journal of Fish Biology, 89(3):1603– 1611, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abudayah:2016:PRL

[2078] W. H. Abudayah and A. Mathis. Predator recognition learning in rainbow darters Etheostoma caeruleum: specific learning and neophobia.

Journal of Fish Biology, 89(3):1612–1623, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flaten:2016:FMS

[2079] A. C. Flaten, J. G. Davidsen, E. B. Thorstad, F. Whoriskey, L. Rønning, A. D. Sjursen, A. H. Rikardsen, and J. V. Arnekleiv. The first months at sea: marine migration and habitat use of sea trout Salmo trutta postsmolts. Journal of Fish Biology, 89(3):1624–1640, September 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newton:2016:DSM

[2080] M. Newton, J. Barry, J. A. Dodd, M. C. Lucas, P. Boylan, and C. E. Adams. Does size matter? A test of size-specific mortality in Atlantic salmon Salmo salar smolts tagged with acoustic transmitters. Journal of Fish Biology, 89(3):1641–1650, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2016:PSM

[2081] R. J. Kennedy and M. Allen. The pre-spawning migratory behaviour of Atlantic salmon Salmo salar in a large lacustrine catchment. Journal of Fish Biology, 89(3):1651–1665, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sweetser:2016:OCG

[2082] P. W. Sweetser and E. J. Hilton. Osteology of the crescent gunnel Pholis laeta and the phylogeny of the family Pholidae (Cottiformes, Zoarcoidei). Journal of Fish Biology, 89(3):1666–1691, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ke:2016:MCC

[2083] F. Ke, H. B. Zhang, Y. Wang, L. F. Hou, H. J. Dong, Z. F. Wang, G. W. Pan, and X. Y. Cao. Molecular cloning and characterization of a C-type lectin in yellow catfish Tachysurus fulvidraco. Journal of Fish Biology, 89(3):1692–1703, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ryan:2016:TSR

[2084] D. Ryan, S. Shephard, and F. L. Kelly. Temporal stability and rates of post-depositional change in geochemical signatures of brown trout Salmo trutta scales. Journal of Fish Biology, 89(3):1704–1719, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:2016:GJS

[2085] J. N. Thompson and D. A. Beauchamp. Growth of juvenile steelhead Oncorhynchus mykiss under size-selective pressure limited by seasonal bioenergetic and environmental constraints. Journal of Fish Biology, 89(3):1720–1739, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DePasquale:2016:DAS

[2086] J. A. DePasquale. Dynamic apical surface rings in superficial layer cells of koi Cyprinus carpio scale epidermis. Journal of Fish Biology, 89(3): 1740–1753, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harvey:2016:PRF

[2087] A. C. Harvey, M. F. Solberg, K. A. Glover, M. I. Taylor, S. Creer, and G. R. Carvalho. Plasticity in response to feed availability: Does feeding regime influence the relative growth performance of domesticated, wild and hybrid Atlantic salmon Salmo salar parr? Journal of Fish Biology, 89(3):1754–1768, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swaminathan:2016:NFC

[2088] T. R. Swaminathan, Raj Kumar, P. M. E. Jency, R. Charan, M. U. Syamkrishnan, V. S. Basheer, N. Sood, and J. K. Jena. A new fish cell line derived from the caudal fin of freshwater angelfish Pterophyllum scalare: development and characterization. Journal of Fish Biology, 89(3):1769– 1781, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davison:2016:LFV

[2089] P. I. Davison, V. Créach, W.-J. Liang, D. Andreou, J. R. Britton, and G. H. Copp. Laboratory and field validation of a simple method for detecting four species of non-native freshwater fish using eDNA. Journal of Fish Biology, 89(3):1782–1793, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahrenstorff:2016:SCP

[2090] T. D. Ahrenstorff and T. R. Hrabik. Seasonal changes in partial, reverse diel vertical migrations of cisco Coregonus artedi. Journal of Fish Biology, 89(3):1794–1809, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chiarello-Sosa:2016:RSP

[2091] J. M. Chiarello-Sosa, M. A. Battini, and J. P. Barriga. Reproductive strategy of the Patagonian catfish Hatcheria macraei. Journal of Fish Biology, 89(3):1810–1827, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kinney:2016:OAV

[2092] M. J. Kinney, R. J. D. Wells, and S. Kohin. Oxytetracycline age validation of an adult shortfin mako shark Isurus oxyrinchus after 6 years at liberty. Journal of Fish Biology, 89(3):1828–1833, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delrieu-Trottin:2016:WEC

[2093] E. Delrieu-Trottin, S. Planes, and J. T. Williams. When endemic coralreef fish species serve as models: endemic mimicry patterns in the Marquesas Islands. Journal of Fish Biology, 89(3):1834–1838, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schemmel:2016:GMS

[2094] E. Schemmel and K. Cole. Gonad morphology of Susan’s dwarfgoby Eviota susanae. Journal of Fish Biology, 89(3):1839–1844, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Locatello:2016:UFP

[2095] L. Locatello, C. Mazzoldi, M. Santon, S. Sparaciari, and M. B. Rasotto. Unexpected female preference for smaller males in the marbled goby Pomatoschistus marmoratus. Journal of Fish Biology, 89(3):1845–1850, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becker:2016:TAF

[2096] A. Becker, A. K. Whitfield, P. D. Cowley, V. J. Cole, and M. D. Taylor. Tidal amplitude and fish abundance in the mouth region of a small estuary. Journal of Fish Biology, 89(3):1851–1856, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolgan:2016:DAC

[2097] M. Bolgan, J. O’Brien, R. A. Rountree, and M. Gammell. Does the Arctic charr Salvelinus alpinus produce sounds in a captive setting? Journal of Fish Biology, 89(3):1857–1865, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scherer:2016:LAE

[2098] U. Scherer, M. Buck, and W. Schuett. Lateralisation in agonistic encounters: do mirror tests reflect aggressive behaviour? A study on a West African cichlid. Journal of Fish Biology, 89(3):1866–1872, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sampaio:2016:HLM

[2099] C. L. S. Sampaio, J. Santander-neto, and T. L. A. Costa. Hogfish Lachnolaimus maximus (Labridae) confirmed in the south-western Atlantic Ocean. Journal of Fish Biology, 89(3):1873–1879, September 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kousteni:2016:OBT

[2100] V. Kousteni and P. Megalofonou. Observations on the biological traits of the rare shark Oxynotus centrina (Chondrichthyes: Oxynotidae) in the Hellenic Seas. Journal of Fish Biology, 89(3):1880–1888, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeLaCruz-Aguero:2016:URD

[2101] J. De La Cruz-Agüero, R. Moncayo-Estrada, V. M. Cota-Gómez, H. Villalobos-Ortiz, and A. Valdez-Pelayo. Unusual records of deepwater teleosteans trawled off the western coast of Mexico. Journal of Fish Biology, 89(3):1889–1896, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2016:EPA

[2102] C. M. Wagner, K. E. Kierczynski, J. B. Hume, and T. M. Luhring. Exposure to a putative alarm cue reduces downstream drift in larval sea lamprey Petromyzon marinus in the laboratory. Journal of Fish Biology, 89(3):1897–1904, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gilby:2016:QIS

[2103] B. L. Gilby, C. J. Henderson, I. R. Tibbetts, and D. D. Burfeind. Quantifying the influence of small omnivorous fishes on seagrass epiphyte load. Journal of Fish Biology, 89(3):1905–1912, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proudlove:2016:BEM

[2104] Graham Proudlove. Biology and evolution of the Mexican cavefish. Journal of Fish Biology, 89(3):1913–1914, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2016:Cb

[2105] Anonymous. Corrigendum. Journal of Fish Biology, 89(3):1915, September 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2016:BPF

[2106] A. K. Whitfield. Biomass and productivity of fishes in estuaries: a South African case study. Journal of Fish Biology, 89(4):1917–1930, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2016:EIW

[2107] J. C. B. Stevens, M. J. H. Hickford, and D. R. Schiel. Evidence of iteroparity in the widely distributed diadromous fish inanga Galaxias maculatus and potential implications for reproductive output. Journal of Fish Biology, 89(4):1931–1946, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nirchio:2016:CCS

[2108] M. Nirchio, C. Oliveira, Z. R. Siccha-Ramirez, V. F. Sene, O. R. Sánchez-Romero, N. R. Ehemann, V. Milana, A. R. Rossi, and L. Sola. Cryptic Caribbean species of Scorpaena (Actinopterygii: Scorpaeniformes) suggested by cytogenetic and molecular data. Journal of Fish Biology, 89 (4):1947–1957, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thieren:2016:HOA

[2109] E. Thieren, A. Ervynck, D. Brinkhuizen, A. Locker, and W. Van Neer. The Holocene occurrence of Acipenser spp. in the southern North Sea: the archaeological record. Journal of Fish Biology, 89(4):1958–1973, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vollset:2016:IBM

[2110] K. W. Vollset, S. Mahlum, J. G. Davidsen, H. Skoglund, and B. T. Barlaup. Interaction between migration behaviour and estuarine mortality in cultivated Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 89(4):1974–1990, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buckland-Nicks:2016:NDN

[2111] J. A. Buckland-Nicks. New details of the neural architecture of the salmonid adipose fin. Journal of Fish Biology, 89(4):1991–2003, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohan:2016:FAB

[2112] S. D. Mohan, J. A. Mohan, T. L. Connelly, B. D. Walther, and J. W. Mc-Clelland. Fatty-acid biomarkers and tissue-specific turnover: validation from a controlled feeding study in juvenile Atlantic croaker Micropogonias undulatus. Journal of Fish Biology, 89(4):2004–2023, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Speranza:2016:FAA

[2113] E. D. Speranza, M. Colombo, L. M. Tatone, N. Cappelletti, M. C. Migoya, and J. C. Colombo. Fatty acid alterations in the detritivorous Prochilodus lineatus promoted by opportunistic feeding on sewage discharges in the Río de la Plata estuary. Journal of Fish Biology, 89 (4):2024–2037, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sarma:2016:MHU

[2114] D. Sarma, R. Das, M. S. Akhtar, A. Ciji, N. K. Sharma, and A. K. Singh. Morpho-histological and ultra architectural changes during early development of endangered golden mahseer Tor putitora. Journal of Fish Biology, 89(4):2038–2054, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzeo:2016:CTS

[2115] I. Mazzeo, E. Giorgini, G. Gioacchini, F. Maradonna, M. C. Vílchez, S. Baloche, S. Dufour, L. Pérez, O. Carnevali, and J. F. Asturiano. A comparison of techniques for studying oogenesis in the European eel Anguilla anguilla. Journal of Fish Biology, 89(4):2055–2069, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:2016:IIC

[2116] W. Clark, E. Leclercq, H. Migaud, J. Nairn, and A. Davie. Isolation, identification and characterisation of ballan wrasse Labrus bergylta plasma pigment. Journal of Fish Biology, 89(4):2070–2084, October 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dessen:2016:DFA

[2117] J.-E. Dessen, R. Arge, M. S. Thomassen, and K.-A. Rørvik. Differences in fat accumulation between immature male and female Atlantic salmon Salmo salar after dietary administration of tetradecylthioacetic acid. Journal of Fish Biology, 89(4):2085–2097, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhivotovsky:2016:HBC

[2118] L. A. Zhivotovsky, T. G. Tochilina, E. G. Shaikhaev, V. P. Pogodin, T. V. Malinina, and A. J. Gharrett. Hybrids between chum Oncorhynchus keta and pink Oncorhynchus gorbuscha salmon: age, growth and morphology and effects on salmon production. Journal of Fish Biology, 89(4):2098– 2106, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadswell:2016:AMF

[2119] M. J. Dadswell, S. A. Wehrell, A. D. Spares, M. F. Mclean, J. W. Beardsall, L. M. Logan-Chesney, G. S. Nau, C. Ceapa, A. M. Redden, and M. J. W. Stokesbury. The annual marine feeding aggregation of Atlantic sturgeon Acipenser oxyrinchus in the inner Bay of Fundy: population characteristics and movement. Journal of Fish Biology, 89(4): 2107–2132, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lima:2016:LIP

[2120] A. R. A. Lima and M. Barletta. Lunar influence on prey availability, diet shifts and niche overlap between Engraulidae larvae in tropical mangrove creeks. Journal of Fish Biology, 89(4):2133–2152, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bacha:2016:ISS

[2121] M. Bacha, A. M. Jeyid, S. Jaafour, A. Yahyaoui, M. Diop, and R. Amara. Insights on stock structure of round sardinella Sardinella aurita off northwest Africa based on otolith shape analysis. Journal of Fish Biology, 89 (4):2153–2166, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banon:2016:ONB

[2122] R. Bañón, T. Maño, and G. Mucientes. Observations of newborn blue sharks Prionace glauca in shallow inshore waters of the north-east Atlantic Ocean. Journal of Fish Biology, 89(4):2167–2177, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Branco-Nunes:2016:FRP

[2123] I. S. L. Branco-Nunes, F. V. Albuquerque, D. M. Nunes, P. G. V. Oliveira, and F. H. V. Hazin. First record of predation between Dasyatis species. Journal of Fish Biology, 89(4):2178–2181, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rueger:2016:HED

[2124] T. Rueger, N. M. Gardiner, and G. P. Jones. Homing is not for everyone: displaced cardinalfish find a new place to live. Journal of Fish Biology, 89 (4):2182–2188, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krumme:2016:TMV

[2125] U. Krumme and F. Bingel. Tetracycline marks visible in Baltic cod Gadus morhua otoliths stored for 40 years. Journal of Fish Biology, 89 (4):2189–2194, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gimenez:2016:RBO

[2126] J. Giménez, A. Manjabacas, V. M. Tuset, and A. Lombarte. Relationships between otolith and fish size from Mediterranean and north-eastern Atlantic species to be used in predator–prey studies. Journal of Fish Biology, 89(4):2195–2202, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chow:2016:GAL

[2127] S. Chow, T. Yanagimoto, H. Kurogi, S. A. Appleyard, and J. J. Pogonoski. A giant anguilliform leptocephalus Thalassenchelys foliaceus Castle & Raju 1975 is a junior synonym of Congriscus maldivensis (Norman 1939). Journal of Fish Biology, 89(4):2203–2211, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horn:2016:AAA

[2128] T. S. Horn, R. Machado, C. V. C. Weiss, J. Giora, and L. R. Malabarba. Analysis of an astounding aggregation of Ranzania laevis (Molidae: Actinopterygii) in the tropical south-western Atlantic. Journal of Fish Biology, 89(4):2212–2218, October 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fields:2016:PSC

[2129] A. T. Fields, K. A. Feldheim, J. Gelsleichter, C. Pfoertner, and D. D. Chapman. Population structure and cryptic speciation in bonnethead sharks Sphyrna tiburo in the south-eastern U.S.A. and Caribbean. Journal of Fish Biology, 89(5):2219–2233, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Homola:2016:CPR

[2130] J. J. Homola, C. R. Ruetz III, S. L. Kohler, and R. A. Thum. Complex postglacial recolonization inferred from population genetic structure of mottled sculpin Cottus bairdii in tributaries of eastern Lake Michigan, U.S.A. Journal of Fish Biology, 89(5):2234–2250, November 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polverino:2016:BLR

[2131] G. Polverino, D. Bierbach, S. S. Killen, S. Uusi-Heikkilä, and R. Arlinghaus. Body length rather than routine metabolic rate and body condition correlates with activity and risk-taking in juvenile zebrafish Danio rerio. Journal of Fish Biology, 89(5):2251–2267, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaefer:2016:IGM

[2132] F. J. Schaefer, B. Hermelink, P. Husmann, W. Meeus, J. Adriaen, and S. Wuertz. Induction of gonadal maturation at different temperatures in burbot Lota lota. Journal of Fish Biology, 89(5):2268–2281, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stern:2016:IAR

[2133] N. Stern, B. Rinkevich, and M. Goren. Integrative approach revises the frequently misidentified species of Sardinella (Clupeidae) of the Indo-West Pacific Ocean. Journal of Fish Biology, 89(5):2282–2305, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogers:2016:DTA

[2134] R. Rogers, S. Rowe, and M. J. Morgan. Depth and temperature associations of haddock Melanogrammus aeglefinus off southern Newfoundland. Journal of Fish Biology, 89(5):2306–2325, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finucci:2016:RBF

[2135] B. Finucci, C. Bustamante, E. G. Jones, and M. R. Dunn. Reproductive biology and feeding habits of the prickly dogfish Oxynotus bruniensis. Journal of Fish Biology, 89(5):2326–2344, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanes:2016:SDC

[2136] L. E. K. Lanés, R. S. Godoy, L. Maltchik, M. Polačik, R. Blažek, M. Vrtílek, and M. Reichard. Seasonal dynamics in community structure, abundance, body size and sex ratio in two species of Neotropical annual fishes. Journal of Fish Biology, 89(5):2345–2364, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebel:2016:WBE

[2137] J. D. Ebel, S. J. Leroux, M. J. Robertson, and J. B. Dempson. Whole body-element composition of Atlantic salmon Salmo salar influenced by migration direction and life stage in three distinct populations. Journal of Fish Biology, 89(5):2365–2374, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamanoue:2016:RGV

[2138] Y. Yamanoue. Revision of the genus Verilus (Perciformes: Acropomatidae) with a description of a new species. Journal of Fish Biology, 89(5): 2375–2398, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2016:LEL

[2139] S. M. Taylor, A. V. Harry, and M. B. Bennett. Living on the edge: latitudinal variations in the reproductive biology of two coastal species of sharks. Journal of Fish Biology, 89(5):2399–2418, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braicovich:2016:ARH

[2140] P. E. Braicovich, E. N. Ieno, M. Sáez, J. Despos, and J. T. Timi. Assessing the role of host traits as drivers of the abundance of long-lived parasites in fish-stock assessment studies. Journal of Fish Biology, 89(5): 2419–2433, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grande:2016:EAS

[2141] M. Grande, H. Murua, I. Zudaire, E. J. Arsenault-Pernet, F. Pernet, and N. Bodin. Energy allocation strategy of skipjack tuna Katsuwonus pelamis during their reproductive cycle. Journal of Fish Biology, 89(5): 2434–2448, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

French:2016:WDB

[2142] W. E. French, B. Vondracek, L. C. Ferrington, Jr., J. C. Finlay, and D. J. Dieterman. Winter diet of brown trout Salmo trutta in groundwaterdominated streams: influence of environmental factors on spatial and temporal variation. Journal of Fish Biology, 89(5):2449–2464, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:CFF

[2143] Ian J. Winfield. Conservation of freshwater fishes. Journal of Fish Biology, 89(5):2465–2466, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:SED

[2144] Ian J. Winfield. Sharks: an eponym dictionary. Journal of Fish Biology, 89(5):2467, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2016:IFA

[2145] Ian J. Winfield. Introductory fisheries analyses with r. Journal of Fish Biology, 89(5):2468–2469, November 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2016:FGG

[2146] G. R. Carvalho, L. Hauser, J. Martinsohn, and K. Naish. Fish, genes and genomes: contributions to ecology, evolution and management. Journal of Fish Biology, 89(6):2471–2478, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waples:2016:TER

[2147] R. S. Waples. Tiny estimates of the N_{e} /N ratio in marine fishes: Are they real? Journal of Fish Biology, 89(6):2479–2504, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ovenden:2016:CEG

[2148] J. R. Ovenden, G. M. Leigh, D. C. Blower, A. T. Jones, A. Moore, C. Bustamante, R. C. Buckworth, M. B. Bennett, and C. L. Dudgeon. Can estimates of genetic effective population size contribute to fisheries stock assessments? Journal of Fish Biology, 89(6):2505–2518, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernatchez:2016:MGV

[2149] L. Bernatchez. On the maintenance of genetic variation and adaptation to environmental change: considerations from population genomics in fishes. Journal of Fish Biology, 89(6):2519–2556, December 2016. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barasa:2016:HGD

[2150] J. E. Barasa, R. Abila, J. P. Grobler, M. Agaba, E. J. Chemoiwa, and B. Kaunda-Arara. High genetic diversity and population differentiation in Clarias gariepinus of Yala Swamp: evidence from mitochondrial DNA sequences. Journal of Fish Biology, 89(6):2557–2570, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drinan:2016:SGP

[2151] D. P. Drinan, H. M. Galindo, T. Loher, and L. Hauser. Subtle genetic population structure in Pacific halibut Hippoglossus stenolepis. Journal of Fish Biology, 89(6):2571–2594, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marques:2016:GDA

[2152] A. C. P. B. Marques, A. C. S. Franco, F. Salgueiro, E. García-Berthou, and L. N. Santos. Genetic divergence among invasive and native populations of the yellow peacock cichlid Cichla kelberi. Journal of Fish Biology, 89(6):2595–2606, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yu:2016:GDP

[2153] H. J. Yu, Y. Kai, and J.-K. Kim. Genetic diversity and population structure of Hyporhamphus sajori (Beloniformes: Hemiramphidae) inferred from mtDNA control region and msDNA markers. Journal of Fish Biology, 89(6):2607–2624, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Mayoral:2016:GSW

[2154] E. Garcia-Mayoral, M. Olsen, R. Hedeholm, S. Post, E. E. Nielsen, and D. Bekkevold. Genetic structure of West Greenland populations of lumpfish Cyclopterus lumpus. Journal of Fish Biology, 89(6):2625–2642, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mascolino:2016:AMA

[2155] S. Mascolino, C. Benvenuto, C. Gubili, C. Sacchi, B. Boufana, and S. Mariani. The ART of mating: alternative reproductive tactics and mating success in a nest-guarding fish. Journal of Fish Biology, 89(6): 2643–2657, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aykanat:2016:PGC

[2156] T. Aykanat, M. Lindqvist, V. L. Pritchard, and C. R. Primmer. From population genomics to conservation and management: a workflow for targeted analysis of markers identified using genome-wide approaches in Atlantic salmon Salmo salar. Journal of Fish Biology, 89(6):2658–2679, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2016:AII

[2157] E. Verspoor, D. Knox, and S. Marshall. Assessment of interbreeding and introgression of farm genes into a small Scottish Atlantic salmon Salmo salar stock: ad hoc samples — ad hoc results? Journal of Fish Biology, 89(6):2680–2696, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pearse:2016:SSA

[2158] D. E. Pearse. Saving the spandrels? Adaptive genomic variation in conservation and fisheries management. Journal of Fish Biology, 89(6): 2697–2716, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leitwein:2016:GWN

[2159] M. Leitwein, P.-A. Gagnaire, E. Desmarais, S. Guendouz, M. Rohmer, P. Berrebi, and B. Guinand. Genome-wide nucleotide diversity of hatchery-reared Atlantic and Mediterranean strains of brown trout Salmo trutta compared to wild Mediterranean populations. Journal of Fish Biology, 89(6):2717–2734, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piccolo:2016:CGC

[2160] J. J. Piccolo. Conservation genomics: coming to a salmonid near you. Journal of Fish Biology, 89(6):2735–2740, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2016:AAC

[2161] L. M. Oliveira, T. Knebelsberger, M. Landi, P. Soares, M. J. Raupach, and F. O. Costa. Assembling and auditing a comprehensive DNA barcode reference library for European marine fishes. Journal of Fish Biology, 89(6):2741–2754, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casey:2016:RGF

[2162] J. Casey, E. Jardim, and J. Th. Martinsohn. The role of genetics in fisheries management under the E.U. common fisheries policy. Journal of Fish Biology, 89(6):2755–2767, December 2016. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIa

[2163] Anonymous. Issue information. Journal of Fish Biology, 90(1):1–2, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lozano-Vilano:2017:DSM

[2164] M. L. Lozano-Vilano and M. De La Maza-Benignos. Diversity and status of Mexican killifishes. Journal of Fish Biology, 90(1):3–38, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2017:NFM

[2165] G. J. Johnson, R. C. Buckworth, H. Lee, J. A. T. Morgan, J. R. Ovenden, and C. R. McMahon. A novel field method to distinguish between cryptic carcharhinid sharks, Australian blacktip shark Carcharhinus tilstoni and common blacktip shark C. limbatus, despite the presence of hybrids. Journal of Fish Biology, 90(1):39–60, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suda:2017:GVL

[2166] A. Suda, N. Nagata, A. Sato, Y. Narimatsu, H. H. Nadiatul, and M. Kawata. Genetic variation and local differences in Pacific cod Gadus macrocephalus around Japan. Journal of Fish Biology, 90(1):61–79, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delomas:2017:SPG

[2167] T. A. Delomas, B. Gomelsky, A. Anil, K. J. Schneider, and J. L. Warner. Spontaneous polyploidy, gynogenesis and androgenesis in second generation (F_{2}) koi Cyprinus carpio × goldfish Carassius auratus hybrids.

Journal of Fish Biology, 90(1):80–92, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koster:2017:ISS

[2168] W. M. Koster, D. R. Dawson, C. Liu, P. D. Moloney, D. A. Crook, and J. R. Thomson. Influence of streamflow on spawning-related movements of golden perch Macquaria ambigua in south-eastern Australia. Journal of Fish Biology, 90(1):93–108, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skjaeraasen:2017:TOD

[2169] J. E. Skjæraasen, J. A. Devine, J. A. Godiksen, M. Fonn, H. Otterå, O. S. Kjesbu, B. Norberg, Ø. Langangen, and Ø. Karlsen. Timecourse of oocyte development in saithe Pollachius virens. Journal of Fish Biology, 90(1):109–128, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Portella:2017:EMA

[2170] T. Portella, J. Lobón-Cerviá, L. R. Manna, H. G. Bergallo, and R. Mazzoni. Eco-morphological attributes and feeding habits in coexisting characins. Journal of Fish Biology, 90(1):129–146, January 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stell:2017:AFR

[2171] S. K. Stell and P. Moller. Anal-fin ray morphology indicates sexual maturity in Brevimyrus niger (Teleostei, Mormyridae). Journal of Fish Biology, 90(1):147–155, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2017:HSH

[2172] J. Martins, F. Almada, A. Gonçalves, P. Duarte-Coelho, and P. E. Jorge. Home sweet home: evidence for nest-fidelity in the rocky intertidal fish, the shanny L ipophrys pholis. Journal of Fish Biology, 90(1):156–166, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gracan:2017:FET

[2173] R. Gračan, D. Zavodnik, P. Krstinić, B. Dragičević, and B. Lazar. Feeding ecology and trophic segregation of two sympatric mesopredatory sharks in the heavily exploited coastal ecosystem of the Adriatic Sea. Journal of Fish Biology, 90(1):167–184, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

May:2017:EST

[2174] L. E. May and J. D. Kieffer. The effect of substratum type on aspects of swimming performance and behaviour in shortnose sturgeon Acipenser brevirostrum. Journal of Fish Biology, 90(1):185–200, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wen:2017:ICN

[2175] C. M. Wen, M. M. Chen, F. H. Nan, and C. S. Wang. Immunocytochemical characterisation of neural stem-progenitor cells from green terror cichlid Aequidens rivulatus. Journal of Fish Biology, 90(1):201– 221, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2017:EAB

[2176] H. H. Lee, L. R. Thomas, K. R. Piner, and M. N. Maunder. Effects of age-based movement on the estimation of growth assuming random-at-age or random-at-length data. Journal of Fish Biology, 90(1):222–235, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanaka:2017:CPF

[2177] H. Tanaka, T. Hamatsu, and K. Mori. Comparison of potential fecundity models for walleye pollock Gadus chalcogrammus in the Pacific waters off Hokkaido, Japan. Journal of Fish Biology, 90(1):236–248, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sepulveda:2017:STP

[2178] F. A. Sepúlveda and M. T. González. Spatio-temporal patterns of genetic variations in populations of yellowtail kingfish Seriola lalandi from the south-eastern Pacific Ocean and potential implications for its fishery management. Journal of Fish Biology, 90(1):249–264, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harborne:2017:EDD

[2179] A. R. Harborne, J. D. Selwyn, J. M. Lawson, and M. Gallo. Environmental drivers of diurnal visits by transient predatory fishes to Caribbean patch reefs. Journal of Fish Biology, 90(1):265–282, January 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sans-Coma:2017:DVD

[2180] V. Sans-Coma, C. Rodríguez, M. A. López-Unzu, M. Lorenzale, B. Fernández, L. Vida, and A. C. Durán. Dicephalous v. diprosopus sharks: record of a two-headed embryo of Galeus atlanticus and review of the literature. Journal of Fish Biology, 90(1):283–293, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedger:2017:PSA

[2181] R. D. Hedger, A. H. Rikardsen, and E. B. Thorstad. Pop-up satellite archival tag effects on the diving behaviour, growth and survival of adult Atlantic salmon Salmo salar at sea. Journal of Fish Biology, 90(1):294– 310, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bose:2017:OMV

[2182] A. P. H. Bose, J. B. Adragna, and S. Balshine. Otolith morphology varies between populations, sexes and male alternative reproductive tactics in a vocal toadfish Porichthys notatus. Journal of Fish Biology, 90(1):311– 325, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barros:2017:ESC

[2183] G. Barros, J. Zuanon, and C. Deus. Effects of species co-occurrence on the trophic-niche breadth of characids in Amazon forest streams. Journal of Fish Biology, 90(1):326–340, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knope:2017:SFH

[2184] M. L. Knope, K. A. Tice, and D. C. Rypkema. Site fidelity and homing behaviour of intertidal sculpins revisited. Journal of Fish Biology, 90 (1):341–355, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chamon:2017:NSL

[2185] C. C. Chamon and L. M. Sousa. A new species of the leopard pleco genus Pseudacanthicus (Siluriformes: Loricariidae) from the Rio Xingu, Brazil. Journal of Fish Biology, 90(1):356–369, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2017:DDE

[2186] K. Hasegawa, T. Mori, and C. Yamazaki. Density-dependent effects of non-native brown trout Salmo trutta on the species–area relationship in stream fish assemblages. Journal of Fish Biology, 90(1):370–383, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galbraith:2017:DLS

[2187] H. S. Galbraith, C. J. Blakeslee, A. K. Schmucker, N. S. Johnson, M. J. Hansen, and W. Li. Donor life stage influences juvenile American eel Anguilla rostrata attraction to conspecific chemical cues. Journal of Fish Biology, 90(1):384–395, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pichler:2017:SEI

[2188] H. A. Pichler, C. A. Gray, M. K. Broadhurst, H. L. Spach, and I. Nagelkerken. Seasonal and environmental influences on recruitment patterns and habitat usage among resident and transient fishes in a World Heritage Site subtropical estuary. Journal of Fish Biology, 90(1):396–416, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burgess:2017:EES

[2189] K. B. Burgess and M. B. Bennett. Effects of ethanol storage and lipid and urea extraction on δ ^{15}N and δ ^{13}C isotope ratios in a benthic elasmobranch, the bluespotted maskray Neotrygon kuhlii. Journal of Fish Biology, 90(1):417–423, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:2017:IBP

[2190] M. J. Hansen, P. M. O’Leary, and A. J. W. Ward. Interactions between Plagiotremus spp., Labroides dimidiatus and their clients: evidence for behavioural niche partitioning. Journal of Fish Biology, 90(1):424–434, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2017:RMP

[2191] E. Kelly and S. Gibson-Kueh. Renal mucous production in Australian freshwater plotosid and ariid catfishes. Journal of Fish Biology, 90(1): 435–442, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pearish:2017:SMQ

[2192] S. Pearish, L. Hostert, and A. M. Bell. A standardized method for quantifying consistent individual differences in schooling behaviour. Journal of Fish Biology, 90(1):443–450, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britz:2017:FWF

[2193] R. Britz. Fishes of the world. fifth edition. Journal of Fish Biology, 90 (1):451–459, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIb

[2194] Anonymous. Issue information. Journal of Fish Biology, 90(1):460–469, January 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIc

[2195] Anonymous. Issue information. Journal of Fish Biology, 90(2):471, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walther:2017:FOC

[2196] B. D. Walther, K. E. Limburg, C. M. Jones, and J. J. Schaffler. Frontiers in otolith chemistry: insights, advances and applications. Journal of Fish Biology, 90(2):473–479, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Limburg:2017:ITD

[2197] K. E. Limburg and M. Elfman. Insights from two-dimensional mapping of otolith chemistry. Journal of Fish Biology, 90(2):480–491, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2017:BZU

[2198] C. M. Jones, M. Palmer, and J. J. Schaffler. Beyond zar: the use and abuse of classification statistics for otolith chemistry. Journal of Fish Biology, 90(2):492–504, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2017:CCU

[2199] A. K. Carlson, Q. E. Phelps, and B. D. S. Graeb. Chemistry to conservation: using otoliths to advance recreational and commercial fisheries management. Journal of Fish Biology, 90(2):505–527, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Humston:2017:IGR

[2200] R. Humston, S. S. Doss, C. Wass, C. Hollenbeck, S. R. Thorrold, S. Smith, and C. P. Bataille. Isotope geochemistry reveals ontogeny of dispersal and exchange between main-river and tributary habitats in smallmouth bass Micropterus dolomieu. Journal of Fish Biology, 90(2): 528–548, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pracheil:2017:SPA

[2201] B. M. Pracheil, B. C. Chakoumakos, M. Feygenson, G. W. Whitledge, R. P. Koenigs, and R. M. Bruch. Sturgeon and paddlefish (Acipenseridae) sagittal otoliths are composed of the calcium carbonate polymorphs vaterite and calcite. Journal of Fish Biology, 90(2):549–558, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMillan:2017:EEH

[2202] M. N. McMillan, C. Izzo, B. Wade, and B. M. Gillanders. Elements and elasmobranchs: hypotheses, assumptions and limitations of elemental analysis. Journal of Fish Biology, 90(2):559–594, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bock:2017:RBW

[2203] L. R. Bock, G. W. Whitledge, B. Pracheil, and P. Bailey. Relationships between water and paddlefish Polyodon spathula dentary elemental and stable-isotopic signatures: potential application for reconstructing environmental history. Journal of Fish Biology, 90(2):595–610, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tzadik:2017:LHS

[2204] O. E. Tzadik, E. B. Peebles, and C. D. Stallings. Life-history studies by non-lethal sampling: using microchemical constituents of fin rays as chronological recorders. Journal of Fish Biology, 90(2):611–625, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phelps:2017:LSA

[2205] Q. E. Phelps, R. N. Hupfeld, and G. W. Whitledge. Lake sturgeon Acipenser fulvescens and shovelnose sturgeon Scaphirhynchus platorynchus environmental life history revealed using pectoral fin-ray microchemistry: implications for interjurisdictional conservation through fishery closure zones. Journal of Fish Biology, 90(2):626–639, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IId

[2206] Anonymous. Issue information. Journal of Fish Biology, 90(2):640–649, February 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIe

[2207] Anonymous. Issue information. Journal of Fish Biology, 90(3):651–652, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2017:RCP

[2208] J. R. Ellis, S. R. McCully Phillips, and F. Poisson. A review of capture and post-release mortality of elasmobranchs. Journal of Fish Biology, 90(3):653–722, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townhill:2017:FLO

[2209] B. L. Townhill, J. K. Pinnegar, D. A. Righton, and J. D. Metcalfe. Fisheries, low oxygen and climate change: how much do we really know? Journal of Fish Biology, 90(3):723–750, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tix:2017:ECD

[2210] J. A. Tix, C. T. Hasler, C. Sullivan, J. D. Jeffrey, and C. D. Suski. Elevated carbon dioxide has limited acute effects on Lepomis macrochirus behaviour. Journal of Fish Biology, 90(3):751–772, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aoyama:2017:GDE

[2211] J. Aoyama, T. Yoshinaga, C. Tanaka, and K. Ishii. Geographic distribution and environmental control of vertebral count in Ammodytes spp. along the northern Pacific coast of Japan. Journal of Fish Biology, 90 (3):773–785, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melo:2017:RGD

[2212] M. R. S. Melo. A review of the genus Dysalotus (Percomorphacea: Chiasmodontidae), with the description of Dysalotus pouliulii sp. nov. Journal of Fish Biology, 90(3):786–802, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lai:2017:CEM

[2213] R. F. Lai, I. Jakovlić, H. Liu, J. Wei, F. B. Zhan, P. H. Yang, and W. M. Wang. Characterization and expression of Megalobrama amblycephala toll-like receptor 22 involved in the response to Aeromonas hydrophila. Journal of Fish Biology, 90(3):803–818, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christensen:2017:POO

[2214] E. A. F. Christensen, M. B. S. Svendsen, and J. F. Steffensen. Plasma osmolality and oxygen consumption of perch Perca fluviatilis in response to different salinities and temperatures. Journal of Fish Biology, 90(3): 819–833, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2017:NCD

[2215] S. W. Liu, X. F. Pan, J. X. Yang, and X. Y. Chen. A new cave-dwelling loach, Triplophysa xichouensis sp. nov. (Teleostei Nemacheilidae) from Yunnan, China. Journal of Fish Biology, 90(3):834–846, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pattrick:2017:VST

[2216] P. Pattrick and N. A. Strydom. Variability in spatial and temporal occurrence of presettlement and settlement-stage fishes associated with shallow reefs. Journal of Fish Biology, 90(3):847–866, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bellodi:2017:LHT

[2217] A. Bellodi, C. Porcu, R. Cannas, Al. Cau, M. F. Marongiu, A. Mulas, S. Vittori, and M. C. Follesa. Life-history traits of the long-nosed skate Dipturus oxyrinchus. Journal of Fish Biology, 90(3):867–888, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romo-Curiel:2017:RCM

[2218] A. E. Romo-Curiel, O. Sosa-Nishizaki, J. C. Pérez-Jiménez, and M. C. Rodríguez-Medrano. Reproductive cycle and maternal–embryonic nutritional relationship of shovelnose guitarfish Pseudobatos productus in the Gulf of California. Journal of Fish Biology, 90(3):889–905, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guest:2017:ERP

[2219] T. W. Guest, C. F. Rakocinski, A. N. Evans, and R. B. Blaylock. Effects of release procedures on the primary stress response and post-release survival and growth of hatchery-reared spotted seatrout Cynoscion nebulosus. Journal of Fish Biology, 90(3):906–921, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2017:EEC

[2220] D. C. Costa, C. C. Mattioli, W. S. Silva, R. Takata, F. O. P. Leme, A. L. Oliveira, and R. K. Luz. The effect of environmental colour on the growth, metabolism, physiology and skin pigmentation of the carnivorous freshwater catfish Lophiosilurus alexandri. Journal of Fish Biology, 90 (3):922–935, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atsumi:2017:MFE

[2221] K. Atsumi, H. Y. Song, H. Senou, K. Inoue, and K. Mabuchi. Morphological features of an endangered Japanese strain of Cyprinus carpio: reconstruction based on seven SNP markers. Journal of Fish Biology, 90(3):936–953, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miyazaki:2017:SLU

[2222] T. Miyazaki, S. L. Kondrashev, S. Kasagi, K. Mizusawa, and A. Takahashi. Sequence and localization of an ultraviolet (sws1) opsin in the retina of the Japanese sardine Sardinops melanostictus (Teleostei: Clupeiformes). Journal of Fish Biology, 90(3):954–967, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tscha:2017:SLV

[2223] M. K. Tschá, R. A. Baggio, F. M. Marteleto, V. Abilhoa, L. Bachmann, and W. A. Boeger. Sea-level variations have influenced the demographic history of estuarine and freshwater fishes of the coastal plain of Paraná, Brazil. Journal of Fish Biology, 90(3):968–979, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dockery:2017:ESP

[2224] D. R. Dockery, T. E. McMahon, K. M. Kappenman, and M. Blank. Evaluation of swimming performance for fish passage of longnose dace Rhinichthys cataractae using an experimental flume. Journal of Fish Biology, 90(3):980–1000, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Padula:2017:MDP

[2225] V. M. Padula, D. Causey, and J. A. López. Mitochondrial DNA phylogeography of least cisco Coregonus sardinella in Alaska. Journal of Fish Biology, 90(3):1001–1020, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Ramos:2017:VPI

[2226] M. S. González-Ramos, A. Santos-Moreno, E. F. Rosas-Alquicira, and G. Fuentes-Mascorro. Validation of photo-identification as a mark– recapture method in the spotted eagle ray Aetobatus narinari. Journal of Fish Biology, 90(3):1021–1030, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunbar:2017:FCP

[2227] M. Dunbar, C. Onuora, S. Morgan, F. E. Stone, T. M. Huckaba, and I. R. Davenport. Follicle cell processes: a shark thing? Journal of Fish Biology, 90(3):1031–1036, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Killen:2017:DMS

[2228] S. S. Killen, T. Norin, and L. G. Halsey. Do method and species lifestyle affect measures of maximum metabolic rate in fishes? Journal of Fish Biology, 90(3):1037–1046, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feldheim:2017:MBC

[2229] K. A. Feldheim, A. Clews, A. Henningsen, L. Todorov, C. McDermott, M. Meyers, J. Bradley, A. Pulver, E. Anderson, and A. Marshall. Multiple births by a captive swellshark Cephaloscyllium ventriosum via facultative parthenogenesis. Journal of Fish Biology, 90(3):1047–1053, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakinan:2017:IDB

[2230] S. Sakinan, A. Karahan, and M. Ok. Integration of DNA barcoding for the initial recordings of Lessepsian fishes: a case study of the Indo– Pacific slender ponyfish Equulites elongatus. Journal of Fish Biology, 90 (3):1054–1061, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2017:RDC

[2231] V. A. Oliveira, M. A. P. Ferreira, R. M. Rocha, and L. F. A. Montag. Reproduction of the duck catfish Ageneiosus ucayalensis in a ria river system. Journal of Fish Biology, 90(3):1062–1069, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2017:GII

[2232] J. L. McKenzie and J. R. Alvarado Bremer. Genetic identification of istiophorid larvae from the Gulf of Mexico based on the analysis of mitochondrial DNA control region sequences. Journal of Fish Biology, 90 (3):1070–1079, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Musschoot:2017:REG

[2233] T. Musschoot and J. Snoeks. Re-establishment of the genus Monostichodus Vaillant 1886 (Characiformes, Distichodontidae). Journal of Fish Biology, 90(3):1080–1082, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schnetz:2017:VRS

[2234] L. Schnetz, J. Kriwet, and C. Pfaff. Virtual reconstruction of the skeletal labyrinth of two lamnid sharks (Elasmobranchii, Lamniformes). Journal of Fish Biology, 90(3):1083–1089, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beatty:2017:RPE

[2235] S. J. Beatty and D. L. Morgan. Rapid proliferation of an endemic galaxiid following eradication of an alien piscivore (Perca fluviatilis) from a reservoir. Journal of Fish Biology, 90(3):1090–1097, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:2017:LJD

[2236] L. H. A. Pereira, A. Bialetzki, and A. C. T. Bonecker. Larval and juvenile development of Tatia intermedia (Siluriformes: Auchenipteridae). Journal of Fish Biology, 90(3):1098–1103, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yarmohammadi:2017:DEG

[2237] M. Yarmohammadi, M. Pourkazemi, and R. Kazemi. Differential expression of foxl2 and cyp19a1a mRNA during gonad developmental stages in great sturgeon Huso huso. Journal of Fish Biology, 90(3):1104–1111, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergendahl:2017:BBS

[2238] I. Ahlbeck Bergendahl, S. Miller, C. Depasquale, L. Giralico, and V. A. Braithwaite. Becoming a better swimmer: structural complexity enhances agility in a captive-reared fish. Journal of Fish Biology, 90(3): 1112–1117, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carniatto:2017:HST

[2239] N. Carniatto, R. Fugi, and S. M. Thomaz. Highly segregated trophic niche of two congeneric fish species in Neotropical floodplain lakes. Journal of Fish Biology, 90(3):1118–1125, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palsson:2017:NHR

[2240] J. Palsson and O. S. Astthorsson. New and historical records of the ocean sunfish Mola mola in Icelandic waters. Journal of Fish Biology, 90 (3):1126–1132, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahuir-Baraja:2017:FCR

[2241] A. E. Ahuir-Baraja, Y. Yamanoue, and L. Kubicek. First confirmed record of Mola sp. A in the western Mediterranean Sea: morphological, molecular and parasitological findings. Journal of Fish Biology, 90(3): 1133–1141, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehemann:2017:LDR

[2242] N. R. Ehemann, L. V. González-González, and A. W. Trites. Lesser devil rays Mobula cf. hypostoma from Venezuela are almost twice their previously reported maximum size and may be a new sub-species. Journal of Fish Biology, 90(3):1142–1148, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Longenecker:2017:HDD

[2243] K. Longenecker, R. Langston, and M. Crane. Hermaphroditism in the dash-and-dot goatfish Parupeneus barberinus. Journal of Fish Biology, 90(3):1149–1152, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2017:RW

[2244] Laith A. Jawad. Rays of the world. Journal of Fish Biology, 90(3):1153– 1154, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIf

[2245] Anonymous. Issue information. Journal of Fish Biology, 90(3):1155– 1164, March 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIg

[2246] Anonymous. Issue information. Journal of Fish Biology, 90(4):1165– 1166, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2017:CPU

[2247] I. J. Harrison, J. Arroyave, N. Lujan, and J. F. Craig. Correct procedure for uploading information on new taxonomic names to ZooBank. Journal of Fish Biology, 90(4):1167–1169, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borsa:2017:CAC

[2248] P. Borsa. Comment on ‘Annotated checklist of the living sharks, batoids and chimaeras (Chondrichthyes) of the world, with a focus on biogeographical diversity by Weigmann (2016)’. Journal of Fish Biology, 90 (4):1170–1175, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [1879, 2249].

Weigmann:2017:RBC

[2249] S. Weigmann. Reply to Borsa (2017): Comment on ‘Annotated checklist of the living sharks, batoids and chimaeras (Chondrichthyes) of the world, with a focus on biogeographical diversity by Weigmann (2016)’. Journal of Fish Biology, 90(4):1176–1181, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [1879, 2248].

Vitule:2017:CFB

[2250] J. R. S. Vitule, A. P. L. da Costa, F. A. Frehse, L. A. V. Bezerra, T. V. T. Occhi, V. S. Daga, and A. A. Padial. Comment on ‘Fish biodiversity and conservation in South America by Reis et al. (2016)’. Journal of Fish Biology, 90(4):1182–1190, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2017:RVA

[2251] R. E. Reis, J. S. Albert, F. Di Dario, M. M. Mincarone, P. Petry, and L. A. Rocha. Reply to Vitule et al. (2017): Comment on ‘Fish biodiversity and conservation in South America by Reis et al. (2016)’. Journal of Fish Biology, 90(4):1191–1195, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sutton:2017:DEL

[2252] T. M. Sutton. Distribution and ecology of lampreys Lethenteron spp. in interior Alaskan rivers. Journal of Fish Biology, 90(4):1196–1213, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeoh:2017:DSS

[2253] D. E. Yeoh, F. J. Valesini, C. S. Hallett, D. A. Abdo, and J. Williams. Diel shifts in the structure and function of nearshore estuarine fish communities. Journal of Fish Biology, 90(4):1214–1243, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:2017:TSB

[2254] J. A. Lewis and T. E. Pitcher. Tactic-specific benefits of polyandry in Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 90 (4):1244–1256, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2017:BHR

[2255] E. Y. Chen, J. B. K. Leonard, and H. Ueda. The behavioural homing response of adult chum salmon Oncorhynchus keta to amino-acid profiles. Journal of Fish Biology, 90(4):1257–1264, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:2017:SDF

[2256] N. C. Ferreira, R. M. Guereschi, C. Machado, C. A. Lopes, and A. P. O. Nuñer. Structure and diversity of fishes in a freshwater and coastal subtropical lagoon. Journal of Fish Biology, 90(4):1265–1282, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laptikhovsky:2017:SCB

[2257] V. Laptikhovsky, B. Reid, and P. Brickle. Spawning in a cold bath: reproduction of polar and deep-sea fish family Psychrolutidae. Journal of Fish Biology, 90(4):1283–1296, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imre:2017:MSS

[2258] I. Imre, R. T. Di Rocco, H. McClure, N. S. Johnson, and G. E. Brown. Migratory-stage sea lamprey Petromyzon marinus stop responding to conspecific damage-released alarm cues after 4 h of continuous exposure in laboratory conditions. Journal of Fish Biology, 90(4):1297–1304, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schultz:2017:GAS

[2259] L. D. Schultz, B. E. Chasco, S. L. Whitlock, M. H. Meeuwig, and C. B. Schreck. Growth and annual survival estimates to examine the ecology of larval lamprey and the implications of ageing error in fitting models. Journal of Fish Biology, 90(4):1305–1320, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filmalter:2017:FES

[2260] J. D. Filmalter, P. D. Cowley, M. Potier, F. Ménard, M. J. Smale, Y. Cherel, and L. Dagorn. Feeding ecology of silky sharks Carcharhinus falciformis associated with floating objects in the western Indian Ocean. Journal of Fish Biology, 90(4):1321–1337, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogan:2017:RLG

[2261] J. D. Hogan, R. Kozdon, M. J. Blum, J. F. Gilliam, J. W. Valley, and P. B. McIntyre. Reconstructing larval growth and habitat use in an amphidromous goby using otolith increments and microchemistry. Journal of Fish Biology, 90(4):1338–1355, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa-Herrera:2017:SDS

[2262] T. Correa-Herrera, M. Barletta, A. R. A. Lima, L. F. Jiménez-Segura, and L. B. Arango-Sánchez. Spatial distribution and seasonality of ichthyoplankton and anthropogenic debris in a river delta in the Caribbean Sea. Journal of Fish Biology, 90(4):1356–1387, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribeiro:2017:TRS

[2263] F. R. V. Ribeiro, L. H. Rapp Py-Daniel, and S. J. Walsh. Taxonomic revision of the South American catfish genus Ageneiosus (Siluriformes: Auchenipteridae) with the description of four new species. Journal of Fish Biology, 90(4):1388–1478, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolgan:2017:BAC

[2264] M. Bolgan, J. O’Brien, M. Picciulin, L. Manning, and M. Gammell. Behaviour of Arctic charr Salvelinus alpinus during an induced mating season in captivity: how male relative size influences male behavioural investment and female preference over time. Journal of Fish Biology, 90 (4):1479–1505, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parodi:2017:EST

[2265] J. Parodi, G. Guerra, M. Cuevas, A. Ramírez-Reveco, and F. Romero. Effects of storage time on the motility, mortality and calcium levels of Atlantic salmon Salmo salar spermatozoa. Journal of Fish Biology, 90 (4):1506–1516, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2017:VDI

[2266] Y. Huang, F. Chen, W. Tang, Z. Lai, and X. Li. Validation of daily increment deposition and early growth of mud carp Cirrhinus molitorella. Journal of Fish Biology, 90(4):1517–1532, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawakami:2017:PST

[2267] T. Kawakami, Y. Yamada, S. Tanaka, and K. Tsukamoto. Prolongation of somitogenesis in two anguilliform species, the Japanese eel Anguilla japonica and pike eel Muraenesox cinereus, with refined descriptions of their early development. Journal of Fish Biology, 90(4):1533–1547, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loose:2017:CMS

[2268] E. L. Loose, E. J. Hilton, and J. E. Graves. Comparative morphology of the scales of roundscale spearfish Tetrapturus georgii and white marlin Kajikia albida. Journal of Fish Biology, 90(4):1548–1569, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Higgins:2017:NNE

[2269] B. A. Higgins, D. Pearson, and R. S. Mehta. El Niño episodes coincide with California moray Gymnothorax mordax settlement around Santa Catalina Island, California. Journal of Fish Biology, 90(4):1570–1583, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradney:2017:SSR

[2270] D. R. Bradney, A. Davidson, S. P. Evans, B. E. Wueringer, D. L. Morgan, and P. D. Clausen. Sawfishes stealth revealed using computational fluid dynamics. Journal of Fish Biology, 90(4):1584–1596, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schreiber:2017:TRG

[2271] B. Schreiber, J. Monka, B. Drozd, M. Hundt, M. Weiss, T. Oswald, R. Gergs, and R. Schulz. Thermal requirements for growth, survival and aerobic performance of weatherfish larvae Misgurnus fossilis. Journal of Fish Biology, 90(4):1597–1608, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lik:2017:SCE

[2272] J. Lik, M. Dukowska, M. Grzybkowska, and J. Leszczyńska. Summer coexistence of small-sized cyprinid and percid individuals in natural and impounded stretches of a lowland river: food niche partitioning among fishes. Journal of Fish Biology, 90(4):1609–1630, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tellechea:2017:PAM

[2273] J. S. Tellechea, M. L. Fine, and W. Norbis. Passive acoustic monitoring, development of disturbance calls and differentiation of disturbance and advertisement calls in the Argentine croaker Umbrina canosai (Sciaenidae). Journal of Fish Biology, 90(4):1631–1643, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tentelier:2017:TMC

[2274] C. Tentelier, O. Barroso-Gomila, O. Lepais, A. Manicki, I. Romero-Garmendia, and B. M. Jugo. Testing mate choice and overdominance at MH in natural families of Atlantic salmon Salmo salar. Journal of Fish Biology, 90(4):1644–1659, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hollo:2017:BGH

[2275] T. Hollo, B. M. Watson, S. V. Johnston, and R. H. Devlin. Behaviour of growth hormone transgenic Coho salmon Oncorhynchus kisutch in marine mesocosms assessed by acoustic tag telemetry. Journal of Fish Biology, 90(4):1660–1667, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2017:DAM

[2276] W. Zhang, X. Q. Tang, X. Z. Sheng, J. Xing, and W. B. Zhan. Development and application of monoclonal antibodies against IgM of black rockfish Sebastes schlegeli. Journal of Fish Biology, 90(4):1668–1675, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:R

[2277] Anonymous. Referees 2016. Journal of Fish Biology, 90(4):1676–1681, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIh

[2278] Anonymous. Issue information. Journal of Fish Biology, 90(4):1682– 1691, April 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIi

[2279] Anonymous. Issue information. Journal of Fish Biology, 90(5):1693– 1694, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2017:RPV

[2280] G. McGregor Reid. Richard Peter Vari 1949–2016. Journal of Fish Biology, 90(5):1695–1699, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nikolsky:2017:GS

[2281] V. Nikolsky and J. Lloret. Georgiy shulman 1929–2016. Journal of Fish Biology, 90(5):1700–1701, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Calegari:2017:NSM

[2282] B. B. Calegari and R. E. Reis. New species of the miniature genus Gelanoglanis (Siluriformes: Auchenipteridae) from the Tocantins River basin (Brazil) and osteological description of G. nanonocticolus. Journal of Fish Biology, 90(5):1702–1716, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torniainen:2017:IMF

[2283] J. Torniainen, M. J. Kainz, R. I. Jones, M. Keinänen, P. J. Vuorinen, and M. Kiljunen. Influence of the marine feeding area on the muscle and egg fatty-acid composition of Atlantic salmon Salmo salar spawners estimated from the scale stable isotopes. Journal of Fish Biology, 90 (5):1717–1733, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valesini:2017:ICC

[2284] F. J. Valesini, A. Cottingham, C. S. Hallett, and K. R. Clarke. Interdecadal changes in the community, population and individual levels of the fish fauna of an extensively modified estuary. Journal of Fish Biology, 90(5):1734–1767, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciccotto:2017:EMN

[2285] P. J. Ciccotto and T. C. Mendelson. The evolution of male nuptial colour in a sexually dimorphic group of fishes (Percidae: Etheostomatinae). Journal of Fish Biology, 90(5):1768–1784, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salvanes:2017:ABH

[2286] A. G. V. Salvanes. Are antipredator behaviours of hatchery Salmo salar juveniles similar to wild juveniles? Journal of Fish Biology, 90(5):1785– 1796, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grandi:2017:IIO

[2287] G. Grandi, M. Pezzi, M. G. Marchetti, and M. Chicca. Immunocytochemical identification and ontogeny of adenohypophyseal cells in a cave fish, Phreatichthys andruzzii (Cypriniformes: Cyprinidae). Journal of Fish Biology, 90(5):1797–1822, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

French:2017:OFD

[2288] B. French, M. E. Platell, K. R. Clarke, and I. C. Potter. Optimization of foraging and diet by the piscivorous Othos dentex (Serranidae). Journal of Fish Biology, 90(5):1823–1841, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perelman:2017:FDC

[2289] J. N. Perelman, K. N. Schmidt, I. Haro, I. R. Tibbetts, and M. T. Zischke. Feeding dynamics, consumption rates and daily ration of wahoo Acanthocybium solandri in Indo–Pacific waters. Journal of Fish Biology, 90(5):1842–1860, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McBride:2017:GLA

[2290] R. S. McBride, A. K. Johnson, E. K. Lindsay, H. J. Walsh, and R. A. Richards. Goosefish Lophius americanus fecundity and spawning frequency, with implications for population reproductive potential. Journal of Fish Biology, 90(5):1861–1882, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fricke:2017:SEN

[2291] R. Fricke and M. Abu El-Regal. Schindleria elongata, a new species of paedomorphic gobioid from the Red Sea (Teleostei: Schindleriidae). Journal of Fish Biology, 90(5):1883–1890, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lidster:2017:ISU

[2292] K. Lidster, G. D. Readman, M. J. Prescott, and S. F. Owen. International survey on the use and welfare of zebrafish Danio rerio in research.

Journal of Fish Biology, 90(5):1891–1905, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nevatte:2017:MMB

[2293] R. J. Nevatte, J. E. Williamson, N. G. F. Vella, V. Raoult, and B. E. Wueringer. Morphometry and microanatomy of the barbels of the common sawshark Pristiophorus cirratus (Pristiophoridae): implications for pristiophorid behaviour. Journal of Fish Biology, 90(5):1906–1925, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polgar:2017:HSC

[2294] G. Polgar, S. Zaccara, M. Babbucci, F. Fonzi, C. M. Antognazza, N. Ishak, Z. Sulaiman, and G. Crosa. Habitat segregation and cryptic adaptation of species of Periophthalmus (Gobioidei: Gobiidae). Journal of Fish Biology, 90(5):1926–1943, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2017:ECC

[2295] P. Li, L. Zhou, S. Wei, M. Yang, S. Ni, Y. Yu, J. Cai, and Q. Qin. Establishment and characterization of a cell line from the head kidney of golden pompano Trachinotus ovatus and its application in toxicology and virus susceptibility. Journal of Fish Biology, 90(5):1944–1959, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowerman:2017:REE

[2296] T. E. Bowerman, A. Pinson-Dumm, C. A. Peery, and C. C. Caudill. Reproductive energy expenditure and changes in body morphology for a population of Chinook salmon Oncorhynchus tshawytscha with a long distance migration. Journal of Fish Biology, 90(5):1960–1979, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:2017:SEA

[2297] K. C. Hall, M. K. Broadhurst, P. A. Butcher, L. Cameron, S. J. Rowland, and R. B. Millar. Sublethal effects of angling and release on golden perch Macquaria ambigua: implications for reproduction and fish health. Journal of Fish Biology, 90(5):1980–1998, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jakubcinova:2017:WCM

[2298] K. Jakubčinová, P. Simonović, B. Števove, J. Čanak Atlagić, and V. Kováč. What can morphology tell us about ecology of four invasive goby species? Journal of Fish Biology, 90(5):1999–2019, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finucci:2017:SOB

[2299] B. Finucci, D. W. Stevens, E. G. Jones, and M. R. Dunn. Some observations on the biology of two rarely seen deep-sea chimaerids, Chimaera carophila and Hydrolagus homonycteris. Journal of Fish Biology, 90(5): 2020–2040, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kingsford:2017:SLN

[2300] M. J. Kingsford, M. D. O’Callaghan, L. Liggins, and G. Gerlach. The short-lived neon damsel Pomacentrus coelestis: implications for population dynamics. Journal of Fish Biology, 90(5):2041–2059, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergendahl:2017:FRA

[2301] I. Ahlbeck Bergendahl, P. B. Holliland, S. Hansson, and O. Karlöf. Feeding range of age 1+ year Eurasian perch Perca fluviatilis in the Baltic Sea. Journal of Fish Biology, 90(5):2060–2072, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalish-Achrai:2017:SCM

[2302] N. Kalish-Achrai, E. Monsonego-Ornan, and R. Shahar. Structure, composition, mechanics and growth of spines of the dorsal fin of blue tilapia Oreochromis aureus and common carp Cyprinus carpio. Journal of Fish Biology, 90(5):2073–2096, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouyoucos:2017:EWF

[2303] I. A. Bouyoucos, C. D. Suski, J. W. Mandelman, and E. J. Brooks. Effect of weight and frontal area of external telemetry packages on the kinematics, activity levels and swimming performance of small-bodied sharks. Journal of Fish Biology, 90(5):2097–2110, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmann:2017:RBH

[2304] G. C. S. Hoffmann, M. O. Freitas, R. L. Moura, M. Previero, and V. Abilhoa. Reproductive biology of Haemulon plumierii in the south-western Atlantic Ocean’s most extensive reefs: implications for fisheries management. Journal of Fish Biology, 90(5):2111–2124, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stepien:2017:PGS

[2305] C. A. Stepien, S. I. Karsiotis, T. J. Sullivan, and K. E. Klymus. Population genetic structure and comparative diversity of smallmouth bass Micropterus dolomieu: congruent patterns from two genomes. Journal of Fish Biology, 90(5):2125–2147, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubbard:2017:HOS

[2306] P. C. Hubbard, F. Baduy, J. L. Saraiva, P. M. Guerreiro, and A. V. M. Canário. High olfactory sensitivity to conspecific intestinal fluid in the chameleon cichlid Australoheros facetus: could faeces signal dominance? Journal of Fish Biology, 90(5):2148–2156, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sequeira:2017:ZSS

[2307] V. Sequeira, E. Couto, A. Neves, A. R. Vieira, R. B. Paiva, A. V. M. Canario, and L. S. Gordo. Zygoparity and sex steroid hormone profiles in bluemouth Helicolenus dactylopterus. Journal of Fish Biology, 90(5): 2157–2169, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crespel:2017:DPF

[2308] A. Crespel, A. Dupont-Prinet, L. Bernatchez, G. Claireaux, R. Tremblay, and C. Audet. Divergence in physiological factors affecting swimming performance between anadromous and resident populations of brook charr Salvelinus fontinalis. Journal of Fish Biology, 90(5):2170–2193, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2017:IEA

[2309] Y. Huang, J. Cai, J. F. Tang, H. Y. Zhang, Z. W. Wang, J. C. Jian, Z. H. Wu, and Y. S. Lu. Identification and expression analysis of terminal deoxynucleotidyl transferase in humphead snapper Lutjanus sanguineus. Journal of Fish Biology, 90(5):2194–2199, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazloumi:2017:MMP

[2310] N. Mazloumi, J. L. Johansen, Z. A. Doubleday, and B. M. Gillanders. Q_{10} measures of metabolic performance and critical swimming speed in King George whiting Sillaginodes punctatus. Journal of Fish Biology, 90(5):2200–2205, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2017:RIC

[2311] L. Murray, M. D. Rennie, J. C. Svendsen, and E. C. Enders. Respirometry increases cortisol levels in rainbow trout Oncorhynchus mykiss: implications for measurements of metabolic rate. Journal of Fish Biology, 90(5):2206–2213, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guillerault:2017:ADM

[2312] N. Guillerault, S. Bouletreau, A. Iribar, A. Valentini, and F. Santoul. Application of DNA metabarcoding on faeces to identify European catfish Silurus glanis diet. Journal of Fish Biology, 90(5):2214–2219, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2017:GFF

[2313] L. A. Jawad. A guide to the freshwater fauna of Nee Soon swamp forest. Journal of Fish Biology, 90(5):2220–2221, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottewell:2017:BRF

[2314] D. Ottewell. Book review: Freshwater fishes of Britain. Journal of Fish Biology, 90(5):2222–2223, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:2017:ESE

[2315] Nigel Milner. Ecology of salmonids in estuaries around the world. Journal of Fish Biology, 90(5):2224–2225, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIj

[2316] Anonymous. Issue information. Journal of Fish Biology, 90(5):2226– 2235, May 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIk

[2317] Anonymous. Issue information. Journal of Fish Biology, 90(6):2237– 2238, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:2017:LDA

[2318] F. N. Solomon, D. Rodrigues, E. J. Gonçalves, E. A. Serrão, and R. Borges. Larval development and allometric growth of the black-faced blenny Tripterygion delaisi. Journal of Fish Biology, 90(6):2239–2254, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delpiani:2017:DQA

[2319] G. Delpiani, M. C. Spath, M. Deli Antoni, and M. Delpiani. Description and quantitative analysis of the dentition of the southern thorny skate Amblyraja doellojuradoi. Journal of Fish Biology, 90(6):2255–2270, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamoutene:2017:ELT

[2320] D. Hamoutene, J. Perez-Casanova, K. Burt, L. Lush, J. Caines, C. Collier, and R. Hinks. Early life traits of farm and wild Atlantic salmon Salmo salar and first generation hybrids in the south coast of Newfoundland. Journal of Fish Biology, 90(6):2271–2288, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Acharjee:2017:EAP

[2321] A. Acharjee, R. Chaube, and K. P. Joy. Effects of altered photoperiod and temperature on expression levels of gonadotrophin subunit mRNAs in the female stinging catfish Heteropneustes fossilis. Journal of Fish Biology, 90(6):2289–2311, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mao:2017:ISR

[2322] H. Mao, K. Chen, X. Zhu, Q. Luo, J. Zhao, W. Li, X. Wu, and H. Xu. Identification of suitable reference genes for quantitative real-time PCR normalization in blotched snakehead Channa maculata. Journal of Fish Biology, 90(6):2312–2322, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larter:2017:IEM

[2323] K. F. Larter and B. B. Rees. Influence of euthanasia method on blood and gill variables in normoxic and hypoxic Gulf killifish Fundulus grandis. Journal of Fish Biology, 90(6):2323–2343, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vanegas-Rios:2017:NSC

[2324] J. A. Vanegas-Ríos and A. Urbano-Bonilla. A new species of Chrysobrycon (Characiformes, Characidae, Stevardiinae) from the Amazon River basin in Colombia, with a new diagnostic characteristic for the genus. Journal of Fish Biology, 90(6):2344–2362, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nishizawa:2017:VBR

[2325] H. Nishizawa, Y. Kono, N. Arai, J. Shoji, and H. Mitamura. Ventilatory and behavioural responses of the marbled sole Pseudopleuronectes yokohamae to progressive hypoxia. Journal of Fish Biology, 90(6):2363– 2374, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bal:2017:ELT

[2326] G. Bal, L. Montorio, E. Rivot, E. Prévost, J.-L. Baglinière, and M. Nevoux. Evidence for long-term change in length, mass and migration phenology of anadromous spawners in French Atlantic salmon Salmo salar. Journal of Fish Biology, 90(6):2375–2393, June 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Machado:2017:RRE

[2327] F. S. Machado, A. Carvalho-Filho, and T. Giarrizzo. Redescription and range extension of the endangered Paiva’s blenny Lupinoblennius paivai (Perciformes: Blenniidae). Journal of Fish Biology, 90(6):2394–2401, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shihab:2017:HPG

[2328] I. Shihab, A. Gopalakrishnan, N. Vineesh, M. Muktha, K. V. Akhilesh, and P. Vijayagopal. Histological profiling of gonads depicting protandrous hermaphroditism in Eleutheronema tetradactylum. Journal of Fish Biology, 90(6):2402–2411, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanLeeuwen:2017:SLL

[2329] C. H. A. Van Leeuwen, T. Dokk, T. O. Haugen, P. M. Kiffney, and J. Museth. Small larvae in large rivers: observations on downstream movement of European grayling Thymallus thymallus during early life stages. Journal of Fish Biology, 90(6):2412–2424, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coker:2017:ICC

[2330] D. J. Coker, J. P. Nowicki, and N. A. J. Graham. Influence of coral cover and structural complexity on the accuracy of visual surveys of coral-reef fish communities. Journal of Fish Biology, 90(6):2425–2433, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shi:2017:APR

[2331] X. Shi, J. S. Møller, J. Højgaard, J. L. Johansen, J. F. Steffensen, D. Liu, and P. Domenici. The angular position of a refuge affects escape responses in staghorn sculpin Leptocottus armatus. Journal of Fish Biology, 90(6): 2434–2442, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Splendiani:2017:MSC

[2332] A. Splendiani, T. Fioravanti, M. Giovannotti, L. Olivieri, P. Ruggeri, P. Nisi Cerioni, S. Vanni, F. Enrichetti, and V. Caputo Barucchi. Museum samples could help to reconstruct the original distribution of Salmo trutta complex in Italy. Journal of Fish Biology, 90(6):2443–2451, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2017:GMR

[2333] B. M. Taylor and C. Pardee. Growth and maturation of the redlip parrotfish Scarus rubroviolaceus. Journal of Fish Biology, 90(6):2452–2461, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanDeurs:2017:UAQ

[2334] M. Van Deurs, A. Andersson, J. Vinterstare, A. Didenko, A. Persson, C. Brönmark, and P. A. Nilsson. Using accelerometry to quantify prey attack and handling behaviours in piscivorous pike Esox lucius. Journal of Fish Biology, 90(6):2462–2469, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes:2017:FRC

[2335] P. Gomes, A. R. Vieira, R. Oliveira, H. Silva, R. Martins, and M. Carneiro. First record of Cynoscion regalis (Pisces, Sciaenidae) in Portuguese continental waters. Journal of Fish Biology, 90(6):2470–2474, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2017:DNW

[2336] K. Hasegawa. Displacement of native white-spotted charr Salvelinus leucomaenis by non-native brown trout Salmo trutta after resolution of habitat fragmentation by a migration barrier. Journal of Fish Biology, 90(6):2475–2479, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyegaard:2017:SPS

[2337] M. Nyegaard, N. Loneragan, and M. B. Santos. Squid predation by slender sunfish Ranzania laevis (Molidae). Journal of Fish Biology, 90 (6):2480–2487, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beattie:2017:RLP

[2338] M. Beattie, D. P. Nowacek, A. K. Bogdanoff, L. Akins, and J. A. Morris, Jr. The roar of the lionfishes Pterois volitans and Pterois miles. Journal of Fish Biology, 90(6):2488–2495, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catelani:2017:FRP

[2339] P. A. Catelani, A. B. Bauer, F. Di Dario, F. M. Pelicice, and A. C. Petry. First record of pughead deformity in Cichla kelberi (Teleostei: Cichlidae), an invasive species in an estuarine system in south-eastern Brazil. Journal of Fish Biology, 90(6):2496–2503, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diez:2017:DAS

[2340] G. Díez, G. Moreno, T. Galaz, L. Dagorn, and J. Murua. Digestive activity and stomach temperature in farmed bluefin tuna Thunnus thynnus measured by acoustic tag. Journal of Fish Biology, 90(6):2504–2511, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proudlove:2017:LDI

[2341] Graham Proudlove. Life in the dark: illuminating biodiversity in the shadowy haunts of planet earth. Journal of Fish Biology, 90(6):2512– 2513, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIl

[2342] Anonymous. Issue information. Journal of Fish Biology, 90(6):2514– 2523, June 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIm

[2343] Anonymous. Issue information. Journal of Fish Biology, 91(1):1–2, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vilizzi:2017:GPC

[2344] L. Vilizzi and G. H. Copp. Global patterns and clines in the growth of common carp Cyprinus carpio. Journal of Fish Biology, 91(1):3–40, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Godwin:2017:RGW

[2345] S. C. Godwin, L. M. Dill, M. Krkošek, M. H. H. Price, and J. D. Reynolds. Reduced growth in wild juvenile sockeye salmon Oncorhynchus nerka infected with sea lice. Journal of Fish Biology, 91(1):41–57, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albers:2017:RSS

[2346] J. L. Albers and M. L. Wildhaber. Reproductive strategy, spawning induction, spawning temperatures and early life history of captive sicklefin chub Macrhybopsis meeki. Journal of Fish Biology, 91(1):58–79, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Alonso:2017:MFA

[2347] R. Rodríguez-Alonso, M. Megías, M. A. Pombal, and P. Molist. Morphological and functional aspects of the epidermis of the sea lamprey Petromyzon marinus throughout development. Journal of Fish Biology, 91(1):80–100, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Law:2017:RBB

[2348] C. S. W. Law and Y. Sadovy de Mitcheson. Reproductive biology of black seabream Acanthopagrus schlegelii, threadfin porgy Evynnis cardinalis and red pargo Pagrus major in the northern South China Sea with consideration of fishery status and management needs. Journal of Fish Biology, 91(1):101–125, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:2017:MDZ

[2349] A. W. Thompson, A. I. Furness, C. Stone, C. M. Rade, and G. Ortí. Microanatomical diversification of the zona pellucida in aplochelioid killifishes. Journal of Fish Biology, 91(1):126–143, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mueller:2017:CND

[2350] M. Mueller, J. Pander, J. Knott, and J. Geist. Comparison of nine different methods to assess fish communities in lentic flood-plain habitats.

Journal of Fish Biology, 91(1):144–174, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pujolar:2017:NPD

[2351] J. M. Pujolar, A. L. Ferchaud, D. Bekkevold, and M. M. Hansen. Nonparallel divergence across freshwater and marine three-spined stickleback Gasterosteus aculeatus populations. Journal of Fish Biology, 91(1):175– 194, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scribner:2017:GAE

[2352] K. T. Scribner, C. Soiseth, J. McGuire, G. K. Sage, L. Thorsteinson, J. L. Nielsen, and E. Knudsen. Genetic assessment of the effects of streamscape succession on Coho salmon Oncorhynchus kisutch colonization in recently deglaciated streams. Journal of Fish Biology, 91(1):195–218, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2017:DSM

[2353] D. J. Jellyman and M. J. Unwin. Diel and seasonal movements of silver eels, Anguilla dieffenbachii, emigrating from a lake subject to hydroelectric control. Journal of Fish Biology, 91(1):219–241, July 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Quiroga:2017:DEE

[2354] J. J. Rodríguez-Quiroga, C. Otero-Rodiño, P. Suárez, T. P. Nieto, J. M. García Estévez, F. San Juan, and J. L. Soengas. Differential effects of exposure to parasites and bacteria on stress response in turbot Scophthalmus maximus simultaneously stressed by low water depth. Journal of Fish Biology, 91(1):242–259, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiens:2017:BHW

[2355] J. Wiens, R. Ho, A. K. Brassinga, C. A. Deck, P. J. Walsh, R. N. Ben, K. Mcclymont, T. Charlton, A. N. Evans, and W. G. Anderson. Biosynthesis of 1α-hydroxycorticosterone in the winter skate Leucoraja ocellata: evidence to suggest a novel steroidogenic route. Journal of Fish Biology, 91(1):260–277, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brigden:2017:IVR

[2356] K. E. Brigden, C. T. Marshall, B. E. Scott, E. F. Young, and P. Brickle. Interannual variability in reproductive traits of the Patagonian toothfish Dissostichus eleginoides around the sub-Antarctic island of South Georgia. Journal of Fish Biology, 91(1):278–301, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2017:CMS

[2357] C. J. Moran, M. W. O’Neill, J. W. Armbruster, and A. C. Gibb. Can members of the south-western Gila robusta species complex be distinguished by morphological features? Journal of Fish Biology, 91(1):302– 316, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zawadzki:2017:HSS

[2358] C. H. Zawadzki, T. P. A. Ramos, and M. Sabaj. Hypostomus sertanejo (Siluriformes: Loricariidae), new armoured catfish species from northeastern Brazil. Journal of Fish Biology, 91(1):317–330, July 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egan:2017:DTG

[2359] J. P. Egan, U.-S. Chew, C.-H. Kuo, V. Villarroel-Diaz, P. J. Hundt, N. G. Iwinski, M. P. Hammer, and A. M. Simons. Diets and trophic guilds of small fishes from coastal marine habitats in western Taiwan. Journal of Fish Biology, 91(1):331–345, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendes:2017:RCP

[2360] Y. A. Mendes, M. A. P. Ferreira, C. M. C. Lobato, G. M. F. Silva, L. F. A. Montag, and R. M. Rocha. Reproductive characteristics of pike-characids Boulengerella cuvieri (Ctenoluciidae) in the middle Xingu River, Eastern Amazon. Journal of Fish Biology, 91(1):346–353, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roff:2017:MSA

[2361] G. Roff, C. Doropoulos, G. Mereb, and P. J. Mumby. Mass spawning aggregation of the giant bumphead parrotfish Bolbometopon muricatum. Journal of Fish Biology, 91(1):354–361, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boulcott:2017:ESS

[2362] P. Boulcott, J. Clarke, and P. J. Wright. Effect of size on spawning time in the lesser sandeel Ammodytes marinus. Journal of Fish Biology, 91(1):362–367, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shi:2017:CVH

[2363] X. Shi, S. Zhao, C. Ding, W. Jiang, B. Kynard, L. Liu, S. Niu, and D. Liu. Comparison of vertical and horizontal swimming behaviour of the weather loach Misgurnus anguillicaudatus. Journal of Fish Biology, 91(1):368–374, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diamante:2017:GRP

[2364] N. A. Diamante, S. M. A. P. Prioli, A. V. Oliveira, T. M. C. Fabrin, L. M. Prioli, and A. J. Prioli. Genetic relationships of Plagioscion squamosissimus (Perciformes, Sciaenidae) from five Neotropical river basins evaluated using mitochondrial atpase6/8 gene sequences. Journal of Fish Biology, 91(1):375–384, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIn

[2365] Anonymous. Issue information. Journal of Fish Biology, 91(1):385–394, July 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIo

[2366] Anonymous. Issue information. Journal of Fish Biology, 91(2):395–396, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2017:CMR

[2367] Y. Zhang and M. J. H. Gilbert. Comment on ‘Measurement and relevance of maximum metabolic rate in fishes by Norin & Clark (2016)’. Journal of Fish Biology, 91(2):397–402, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norin:2017:RZG

[2368] T. Norin and T. D. Clark. Reply to Zhang & Gilbert (2017): Comment on ‘Measurement and relevance of maximum metabolic rate in fishes by Norin & Clark (2016)’. Journal of Fish Biology, 91(2):403–408, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2017:SWO

[2369] C. H. Stevens, D. P. Croft, G. C. Paull, and C. R. Tyler. Stress and welfare in ornamental fishes: what can be learned from aquaculture? Journal of Fish Biology, 91(2):409–428, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elston:2017:DCP

[2370] C. Elston, R. G. von Brandis, and P. D. Cowley. Dietary composition and prey selectivity of juvenile porcupine rays Urogymnus asperrimus. Journal of Fish Biology, 91(2):429–442, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2017:RBP

[2371] M. F. Martins and M. C. Oddone. Reproductive biology of Psammobatis rutrum (Chondrichthyes: Arhynchobatidae) in south Brazil, south-west Atlantic. Journal of Fish Biology, 91(2):443–459, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergman:2017:TPF

[2372] J. N. Bergman, M. J. Lajeunesse, and P. J. Motta. Teeth penetration force of the tiger shark Galeocerdo cuvier and sandbar shark Carcharhinus plumbeus. Journal of Fish Biology, 91(2):460–472, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Labak:2017:PNE

[2373] I. Labak, V. Pavić, M. Zjalić, S. Blažetić, B. Viljetić, E. Merdić, and M. Heffer. PSA-NCAM expression in the teleost optic tectum is related to ecological niche and use of vision in finding food. Journal of Fish Biology, 91(2):473–489, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matich:2017:TRA

[2374] P. Matich, J. J. Kiszka, K. R. Gastrich, and M. R. Heithaus. Trophic redundancy among fishes in an East African nearshore seagrass community inferred from stable-isotope analysis. Journal of Fish Biology, 91 (2):490–509, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Felix:2017:EES

[2375] M. L. González-Félix, M. Perez-Velazquez, and H. Cañedo-Orihuela. The effects of environmental salinity on the growth and physiology of totoaba Totoaba macdonaldi and shortfin corvina Cynoscion parvipinnis. Journal of Fish Biology, 91(2):510–527, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melo:2017:TNS

[2376] B. F. Melo and C. Oliveira. Three new species of Curimatopsis (Characiformes: Curimatidae) from the Amazon basin. Journal of Fish Biology, 91(2):528–544, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schumacher:2017:NSH

[2377] E. L. Schumacher, B. D. Owens, T. A. Uyeno, A. J. Clark, and J. S. Reece. No support for Heincke’s law in hagfish (Myxinidae): lack of an association between body size and the depth of species occurrence. Journal of Fish Biology, 91(2):545–557, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2017:RCS

[2378] A. B. Anderson, E. M. Salas, L. A. Rocha, and S. R. Floeter. The recent colonization of south Brazil by the Azores chromis Chromis limbata. Journal of Fish Biology, 91(2):558–573, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayden:2017:VSI

[2379] B. Hayden, S. Tongnunui, F. W. H. Beamish, P. Nithirojpakdee, and R. A. Cunjak. Variation in stable-isotope ratios between fin and muscle tissues can alter assessment of resource use in tropical river fishes. Journal of Fish Biology, 91(2):574–586, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinu:2017:RSD

[2380] J. Vinu, M. P. Rajeeshkumar, U. V. Parmeswaran, K. S. Sumod, K. V. Akhilesh, H. Manjebrayakath, and V. N. Sanjeevan. Redescription and sexual dimorphism of Andaman leg-skate Cruriraja andamanica (Chondrichthyes: Rajiformes) with comments on the zoogeography of the genus Cruriraja. Journal of Fish Biology, 91(2):587–602, August 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2017:EHL

[2381] Y. Liu, Q. H. Zhang, Y. W. Dong, C. H. You, S. Q. Wang, Y. Q. Li, and Y. Y. Li. Establishment of a hepatocyte line for studying biosynthesis of long-chain polyunsaturated fatty acids from a marine teleost, the whitespotted spinefoot Siganus canaliculatus. Journal of Fish Biology, 91(2): 603–616, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Decru:2017:SRA

[2382] E. Decru, E. Vreven, K. De Gelas, E. Verheyen, and J. Snoeks. Species richness in the African pike genus Hepsetus: a perfect match between genetics and morphology. Journal of Fish Biology, 91(2):617–627, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Esin:2017:OSC

[2383] E. V. Esin, E. S. Bocharova, N. S. Mugue, and G. N. Markevich. Occurrence of sympatric charr groups, Salvelinus, Salmonidae, in the lakes of Kamchatka: a legacy of the last glaciations. Journal of Fish Biology, 91(2):628–644, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morris:2017:EJR

[2384] M. R. J. Morris, E. Petrovitch, E. Bowles, H. A. Jamniczky, and S. M. Rogers. Exploring Jordan’s rule in Pacific three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 91(2):645–663, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barker:2017:IYY

[2385] A. M. Barker, B. S. Frazier, D. M. Bethea, J. R. Gold, and D. S. Portnoy. Identification of young-of-the-year great hammerhead shark Sphyrna mokarran in northern Florida and South Carolina. Journal of Fish Biology, 91(2):664–668, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duffy:2017:SRE

[2386] H. J. Duffy, T. B. Letessier, and R. A. Irving. Significant range extensions for two fish species at Pitcairn Island, South Pacific. Journal of Fish Biology, 91(2):669–672, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez:2017:CEF

[2387] C. M. Martinez and M. L. J. Stiassny. Can an eel be a flatfish? Observations on enigmatic asymmetrical heterenchelyids from the Guinea coast of West Africa. Journal of Fish Biology, 91(2):673–678, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Afeworki:2017:AMM

[2388] Y. Afeworki, J. J. Videler, F. Y. Tekle, H. E. Kibrom, and J. H. Bruggemann. Alternative male mating tactics of the endemic Red Sea parrotfish Scarus ferrugineus. Journal of Fish Biology, 91(2):679–685, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freitas:2017:RBT

[2389] M. O. Freitas and V. Abilhoa. Reproductive biology of the tarpon snook Centropomus pectinatus (Perciformes, Centropomidae) in estuarine waters in the south-western Atlantic. Journal of Fish Biology, 91(2):686– 694, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taal:2017:EAS

[2390] I. Taal, M. Rohtla, L. Saks, R. Svirgsden, M. Kesler, L. Matetski, and M. Vetemaa. Evidence of Atlantic salmon Salmo salar fry movement between fresh water and a brackish environment. Journal of Fish Biology, 91(2):695–703, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cabrera:2017:RAI

[2391] M. B. Cabrera, S. Bogan, P. Posadas, G. M. Somoza, J. I. Montoya-Burgos, and Y. P. Cardoso. Risks associated with introduction of poeciliids for control of mosquito larvae: first record of the non-native Gambusia holbrooki in Argentina. Journal of Fish Biology, 91(2):704–710, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIp

[2392] Anonymous. Issue information. Journal of Fish Biology, 91(2):711–720, August 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIq

[2393] Anonymous. Issue information. Journal of Fish Biology, 91(3):721–722, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tate:2017:LBR

[2394] M. Tate, R. E. McGoran, C. R. White, and S. J. Portugal. Life in a bubble: the role of the labyrinth organ in determining territory, mating and aggressive behaviours in anabantoids. Journal of Fish Biology, 91 (3):723–749, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2017:NSH

[2395] F. R. Carvalho, F. G. Cabeceira, and L. N. Carvalho. New species of Hyphessobrycon from the Rio Teles Pires, Rio Tapajós basin, Brazil (Ostariophysi, Characiformes). Journal of Fish Biology, 91(3):750–763, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neofytou:2017:SOV

[2396] M. Neofytou, D. G. Sfakianakis, G. Koumoundouros, C. C. Mylonas, and M. Kentouri. Skeletal ontogeny of the vertebral column and of the fins in shi drum Umbrina cirrosa (Teleostei: Perciformes: Sciaenidae), a new candidate species for aquaculture. Journal of Fish Biology, 91(3): 764–788, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2017:NDG

[2397] J. J. Wright. A new diminutive genus and species of catfish from Lake Tanganyika (Siluriformes: Clariidae). Journal of Fish Biology, 91(3): 789–805, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broach:2017:PVS

[2398] J. S. Broach, C. L. Ohs, and N. E. Breen. Protracted volitional spawning of pinfish Lagodon rhomboides and changes in egg quality and fatty-acid composition throughout the spawning season. Journal of Fish Biology, 91 (3):806–817, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chalde:2017:PTA

[2399] T. Chalde and L. A. Miranda. Pituitary–thyroid axis development during the larval–juvenile transition in the pejerrey Odontesthes bonariensis. Journal of Fish Biology, 91(3):818–834, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohara:2017:HPT

[2400] W. M. Ohara, V. P. Abrahão, and V. C. Espíndola. Hyphessobrycon platyodus (Teleostei: Characiformes), a new species from the Rio Madeira basin, Brazil, with comments on how multicuspid teeth relate to feeding habits in Characidae. Journal of Fish Biology, 91(3):835–850, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marinho:2017:CDM

[2401] M. M. Ferreira Marinho. Comparative development in Moenkhausia pittieri and Paracheirodon innesi (Ostariophysi: Characiformes) with comments on heterochrony and miniaturization in the Characidae. Journal of Fish Biology, 91(3):851–865, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pan:2017:GMS

[2402] Z. J. Pan, C. K. Zhu, H. Wang, F. J. Zhou, and X. G. Qiang. Gonadal morphogenesis and sex differentiation in cultured Ussuri catfish Tachysurus ussuriensis. Journal of Fish Biology, 91(3):866–879, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gregoir:2017:LSD

[2403] A. F. Grégoir, C. Philippe, T. Pinceel, J. Reniers, E. S. J. Thoré, B. Vanschoenwinkel, and L. Brendonck. Life stage dependent responses to desiccation risk in the annual killifish Nothobranchius wattersi. Journal of Fish Biology, 91(3):880–895, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rovani:2017:LHI

[2404] A. T. Rovani and L. G. Cardoso. Life history and initial assessment of fishing impacts on the by-catch species Dules auriga (Teleostei: Serranidae) in southern Brazil. Journal of Fish Biology, 91(3):896–911, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borgstrom:2017:PCD

[2405] R. Borgstrøm, J. Trømborg, T. O. Haugen, and B. O. Rosseland. Plerocercoids of the cestode Diphyllobothrium ditremum in brown trout Salmo trutta: substantial increase in infection after establishment of European minnow Phoxinus phoxinus. Journal of Fish Biology, 91(3):912–927, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seeley:2017:CEI

[2406] M. E. Seeley, W. K. Logan, and B. D. Walther. Consistency of elemental and isotope-ratio patterns across multiple scales from individual fish. Journal of Fish Biology, 91(3):928–946, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torquato:2017:VZF

[2407] F. Torquato, H. M. Jensen, P. Range, S. S. Bach, R. Ben-Hamadou, E. E. Sigsgaard, P. F. Thomsen, P. R. Møller, and R. Riera. Vertical zonation and functional diversity of fish assemblages revealed by ROV videos at oil platforms in The Gulf. Journal of Fish Biology, 91(3): 947–967, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabragana:2017:ECG

[2408] E. Mabragaña, D. M. Vazquez, V. Gabbanelli, D. Sabadin, S. A. Barbini, and L. O. Lucifora. Egg cases of the graytail skate Bathyraja griseocauda and the cuphead skate Bathyraja scaphiops from the south-west Atlantic Ocean. Journal of Fish Biology, 91(3):968–974, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2017:ABS

[2409] F. C. Neat. Aggregating behaviour, social interactions and possible spawning in the deep-water fish Coryphaenoides rupestris. Journal of Fish Biology, 91(3):975–980, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rizzari:2017:OMW

[2410] J. R. Rizzari, J. M. Semmens, A. Fox, and C. Huveneers. Observations of marine wildlife tourism effects on a non-focal species. Journal of Fish Biology, 91(3):981–988, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okada:2017:MMA

[2411] R. Okada, T. Inui, Y. Iguchi, T. Kitagawa, K. Takata, and T. Kitagawa. Molecular and morphological analyses revealed a cryptic species of dojo loach Misgurnus anguillicaudatus (Cypriniformes: Cobitidae) in Japan. Journal of Fish Biology, 91(3):989–996, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIr

[2412] Anonymous. Issue information. Journal of Fish Biology, 91(3):997–1006, September 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIs

[2413] Anonymous. Issue information. Journal of Fish Biology, 91(4):1007– 1008, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2017:DTL

[2414] J. Li, J. H. Lan, X. Y. Chen, and L. N. Du. Description of Triplophysa luochengensis sp. nov. (Teleostei: Nemacheilidae) from a karst cave in Guangxi, China. Journal of Fish Biology, 91(4):1009–1017, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jin:2017:ECL

[2415] Y. L. Jin, L. M. Chen, Y. Le, Y. L. Li, Y. H. Hong, K. T. Jia, and M. S. Yi. Establishment of a cell line with high transfection efficiency from zebrafish Danio rerio embryos and its susceptibility to fish viruses. Journal of Fish Biology, 91(4):1018–1031, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

French:2017:TWT

[2416] G. C. A. French, M. Stürup, S. Rizzuto, J. H. van Wyk, D. Edwards, R. W. Dolan, S. P. Wintner, A. V. Towner, and W. O. H. Hughes. The tooth, the whole tooth and nothing but the tooth: tooth shape and ontogenetic shift dynamics in the white shark Carcharodon carcharias. Journal of Fish Biology, 91(4):1032–1047, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojaveer:2017:GAR

[2417] H. Ojaveer, R. Gross, K. Laur, T. Arula, and R. Klais. Genetic analysis reveals the diversity of larval Gobiidae in a temperate estuary. Journal of Fish Biology, 91(4):1048–1061, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tagliafico:2017:RAA

[2418] A. Tagliafico, S. Rangel, and M. K. Broadhurst. Reproductive aspects of the Atlantic angel shark Squatina dumeril in the southern Caribbean Sea. Journal of Fish Biology, 91(4):1062–1071, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sehr:2017:HUE

[2419] M. Sehr and H. Keckeis. Habitat use of the European mudminnow Umbra krameri and association with other fish species in a disconnected Danube side arm. Journal of Fish Biology, 91(4):1072–1093, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burns:2017:RMA

[2420] M. D. Burns and K. S. Cole. Reproductive morphology and its application in testing molecular systematic hypotheses in the family Gobiidae (Teleostei, Gobiiformes). Journal of Fish Biology, 91(4):1094–1108, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catarino:2017:IGS

[2421] D. Catarino, D. Stanković, G. Menezes, and S. Stefanni. Insights into the genetic structure of the rabbitfish Chimaera monstrosa (Holocephali) across the Atlantic–Mediterranean transition zone. Journal of Fish Biology, 91(4):1109–1122, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daGraca:2017:NSA

[2422] W. J. da Graça, C. A. M. Oliveira, F. C. T. Lima, H. P. da Silva, and I. M. Fernandes. A new species of Astyanax (Characiformes: Characidae) from Dolina Água Milagrosa, Rio Paraguai basin, Mato Grosso, Brazil. Journal of Fish Biology, 91(4):1123–1138, October 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshak:2017:IBR

[2423] A. R. Marshak and K. L. Heck, Jr. Interactions between range-expanding tropical fishes and the northern Gulf of Mexico red snapper Lutjanus campechanus. Journal of Fish Biology, 91(4):1139–1165, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spath:2017:DAR

[2424] M. C. Spath, M. Deli Antoni, and G. Delpiani. Dentition of the apron ray Discopyge tschudii (Elasmobranchii: Narcinidae). Journal of Fish Biology, 91(4):1166–1177, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Viana:2017:SBS

[2425] S. T. de. F. L. Viana, M. R. de Carvalho, and D. A. Ebert. Squalus bassi sp. nov., a new long-snouted spurdog (Chondrichthyes: Squaliformes: Squalidae) from the Agulhas Bank. Journal of Fish Biology, 91(4):1178– 1207, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Engin:2017:NSP

[2426] S. Engin and D. Seyhan. A new species of Pomatoschistus (Teleostei, Gobiidae): the Mediterranean’s smallest marine fish. Journal of Fish Biology, 91(4):1208–1223, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bovcon:2017:PTF

[2427] N. D. Bovcon, P. D. Cochia, J. Ruibal Núñez, M. Vucica, and D. E. Figueroa. Presence of the telescope fish Mendosoma lineatum in Patagonian waters, a new species in the ichthyological fauna from the south-west Atlantic Ocean. Journal of Fish Biology, 91(4):1224–1227, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehemann:2017:BNR

[2428] N. R. Ehemann, X. A. Pérez-Palafox, K. Pabón-Aldana, P. A. Mejía-Falla, A. F. Navia, and V. H. Cruz-Escalona. Biological notes on the reef stingray, Urobatis concentricus, an endemic species of Mexico. Journal of Fish Biology, 91(4):1228–1235, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domit:2017:IBM

[2429] C. Domit, M. K. Broadhurst, and H. Bornatowski. Interactions between Manta birostris and Sotalia guianensis in a World Heritage listed Brazilian estuary. Journal of Fish Biology, 91(4):1236–1240, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2017:ETR

[2430] M. C. L. M. Oliveira, R. Mont’alverne, L. A. Sampaio, M. B. Tesser, L. R. V. Ramos, and A. M. Garcia. Elemental turnover rates and trophic discrimination in juvenile Lebranche mullet Mugil liza under experimental conditions. Journal of Fish Biology, 91(4):1241–1249, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2017:MOS

[2431] Ian J. Winfield. Marine ornamental species aquaculture. Journal of Fish Biology, 91(4):1250–1251, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIt

[2432] Anonymous. Issue information. Journal of Fish Biology, 91(4):1252– 1261, October 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIu

[2433] Anonymous. Issue information. Journal of Fish Biology, 91(5):1263– 1264, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodgers:2017:PFC

[2434] T. Rodgers. Proper fin-clip sample collection for molecular analyses in the age of eDNA. Journal of Fish Biology, 91(5):1265–1267, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eldoy:2017:MDU

[2435] S. H. Eldøy, J. G. Davidsen, E. B. Thorstad, F. G. Whoriskey, K. Aarestrup, T. F. Næsje, L. Rønning, A. D. Sjursen, A. H. Rikardsen, and J. V. Arnekleiv. Marine depth use of sea trout Salmo trutta in fjord areas of central Norway. Journal of Fish Biology, 91(5):1268–1283, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:2017:NVE

[2436] J. P. Lewis, W. F. Patterson III, and J. K. Carlson. Natural variability and effects of cleaning and storage procedures on vertebral chemistry of the blacktip shark Carcharhinus limbatus. Journal of Fish Biology, 91(5): 1284–1300, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pastana:2017:NSD

[2437] M. N. L. Pastana, F. C. P. Dagosta, and A. L. H. Esguícero. A new sexually dichromatic miniature Hyphessobrycon (Teleostei: Characiformes: Characidae) from the Rio Formiga, upper Rio Juruena basin, Mato Grosso, Brazil, with a review of sexual dichromatism in Characiformes. Journal of Fish Biology, 91(5):1301–1318, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vera-Duarte:2017:DBS

[2438] J. Vera-Duarte, C. A. Bustos, and M. F. Landaeta. Diet and body shape changes of pāroko Kelloggella disalvoi (Gobiidae) from intertidal pools of Easter Island. Journal of Fish Biology, 91(5):1319–1336, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flowers:2017:SAW

[2439] K. I. Flowers, A. C. Henderson, J. L. Lupton, and D. D. Chapman. Site affinity of whitespotted eagle rays Aetobatus narinari assessed using photographic identification. Journal of Fish Biology, 91(5):1337–1349, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:2017:MAS

[2440] T. Hesthagen, B. M. Larsen, G. Bolstad, P. Fiske, and B. Jonsson. Mitigation of acidified salmon rivers — effects of liming on young brown trout Salmo trutta. Journal of Fish Biology, 91(5):1350–1364, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frafjord:2017:ROS

[2441] K. Frafjord, T. Bakken, L. Kubicek, A.-H. Rønning, and P. O. Syvertsen. Records of ocean sunfish Mola mola along the Norwegian coast spanning two centuries, 1801–2015. Journal of Fish Biology, 91(5):1365–1377, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Steffensen:2017:URD

[2442] K. D. Steffensen, L. A. Powell, and M. A. Pegg. Using the robust design framework and relative abundance to predict the population size of pallid sturgeon Scaphirhynchus albus in the lower Missouri River. Journal of Fish Biology, 91(5):1378–1391, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rhodes:2017:GSS

[2443] K. Rhodes, D. Hernandez-Ortiz, M. Ioanis, W. Washington, S. Maxim, K. Olpet, and S. Malakai. Goldspotted spinefoot Siganus punctatus (Siganidae) age-based reproductive life history and fisheries vulnerability. Journal of Fish Biology, 91(5):1392–1406, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2017:TCH

[2444] Z. P. Wang, D. Wang, C. L. Wang, W. J. Xie, Y. F. Zhu, and X. W. Chen. Transcriptome characterization of HPG axis from Chinese sea perch Lateolabrax maculatus. Journal of Fish Biology, 91(5):1407–1418, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thuong:2017:OOD

[2445] N. P. Thuong, B. Verstraeten, B. D. Kegel, J. Christiaens, T. D. Wolf, P. Sorgeloos, D. Bonte, and D. Adriaens. Ontogenesis of opercular deformities in gilthead sea bream Sparus aurata: a histological description. Journal of Fish Biology, 91(5):1419–1434, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitada:2017:PPD

[2446] S. Kitada, K. Nakajima, and K. Hamasaki. Population panmixia and demographic expansion of a highly piscivorous marine fish Scomberomorus niphonius. Journal of Fish Biology, 91(5):1435–1448, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rault:2017:DTN

[2447] J. Rault, H. Le Bris, M. Robert, L. Pawlowski, M. Denamiel, and D. Kopp. Diets and trophic niches of the main commercial fish species from the Celtic Sea. Journal of Fish Biology, 91(5):1449–1474, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hench:2017:TCH

[2448] K. Hench, W. O. Mcmillan, R. Betancur-R., and O. Puebla. Temporal changes in hamlet communities (Hypoplectrus spp., Serranidae) over 17 years. Journal of Fish Biology, 91(5):1475–1490, November 2017. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lteif:2017:PBL

[2449] M. Lteif, R. Mouawad, G. Khalaf, P. Lenfant, B. Seret, and M. Verdoit-Jarraya. Population biology of the little gulper shark Centrophorus uyato in Lebanese waters. Journal of Fish Biology, 91(5):1491–1509, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernal-Duran:2017:SME

[2450] V. Bernal-Durán, N. Jahnsen-Guzmán, and M. F. Landaeta. Sharing morphospaces: early ontogenetic shape changes in two clingfish larvae (Pisces: Gobiesocidae) from the south-east Pacific Ocean. Journal of Fish Biology, 91(5):1510–1516, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chow:2017:RIB

[2451] S. Chow, H. Kurogi, T. Yamamoto, T. Tomoda, N. Mochioka, F. Shirotori, T. Yoshinaga, D. Ambe, M. Okazaki, S. Nagai, and T. Yanagimoto. Reproductive isolation between sympatric Anguilla japonica and Anguilla marmorata. Journal of Fish Biology, 91(5):1517–1525, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2017:TFN

[2452] Ian J. Winfield. Transforming the fisheries: neoliberalism, nature, and the commons. Journal of Fish Biology, 91(5):1526–1527, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2017:BRM

[2453] Ian J. Winfield. Book review: Marine fishes of Florida. Journal of Fish Biology, 91(5):1528–1529, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2017:TIC

[2454] Ian. J. Winfield. Tilapia in intensive co-culture. Journal of Fish Biology, 91(5):1530–1531, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIv

[2455] Anonymous. Issue information. Journal of Fish Biology, 91(5):1532– 1541, November 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIw

[2456] Anonymous. Issue information. Journal of Fish Biology, 91(6):1543– 1544, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2017:EHJ

[2457] J. F. Craig. Editorial on the history of the Journal of Fish Biology, 1969 to 2017. Journal of Fish Biology, 91(6):1545–1548, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nolan:2017:LHS

[2458] E. T. Nolan, K. J. Downes, A. Richardson, A. Arkhipkin, P. Brickle, J. Brown, R. J. Mrowicki, Z. Shcherbich, N. Weber, and S. B. Weber. Life-history strategies of the rock hind grouper Epinephelus adscensionis at Ascension Island. Journal of Fish Biology, 91(6):1549–1568, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anders:2017:SDS

[2459] N. Anders, A. Fernö, O.-B. Humborstad, S. Løkkeborg, G. Rieucau, and A. C. Utne-Palm. Size-dependent social attraction and repulsion explains the decision of Atlantic cod Gadus morhua to enter baited pots. Journal of Fish Biology, 91(6):1569–1581, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nevatte:2017:FIF

[2460] R. J. Nevatte, B. E. Wueringer, D. E. Jacob, J. M. Park, and J. E. Williamson. First insights into the function of the sawshark rostrum through examination of rostral tooth microwear. Journal of Fish Biology, 91(6):1582–1602, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curtis:2017:LHE

[2461] J. M. R. Curtis, S. V. Santos, J. L. Nadeau, B. Gunn, K. Bigney Wilner, H. Balasubramanian, S. Overington, C.-M. Lesage, J. D’entremont, and K. Wieckowski. Life history and ecology of the elusive European shortsnouted seahorse Hippocampus hippocampus. Journal of Fish Biology, 91(6):1603–1622, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raby:2017:FTU

[2462] G. D. Raby, T. B. Johnson, S. T. Kessel, T. J. Stewart, and A. T. Fisk. A field test of the use of pop-off data storage tags in freshwater fishes. Journal of Fish Biology, 91(6):1623–1641, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morais:2017:SPF

[2463] R. A. Morais, C. E. L. Ferreira, and S. R. Floeter. Spatial patterns of fish standing biomass across Brazilian reefs. Journal of Fish Biology, 91(6): 1642–1667, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castro:2017:LFC

[2464] M. S. Castro and A. C. T. Bonecker. Larval fish collected from soundscattering layers in an offshore tropical area. Journal of Fish Biology, 91(6):1668–1682, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2017:UVI

[2465] T. M. Evans. Using visible implant elastomer to study ammocoete populations with Cormack–Jolly–Seber models. Journal of Fish Biology, 91(6):1683–1698, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stien:2017:CMS

[2466] L. H. Stien, J. Nilsson, S. Bui, J. E. Fosseidengen, T. S. Kristiansen, Ø. Øverli, and O. Folkedal. Consistent melanophore spot patterns allow long-term individual recognition of Atlantic salmon Salmo salar. Journal of Fish Biology, 91(6):1699–1712, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norman:2017:DTS

[2467] B. M. Norman, J. M. Whitty, S. J. Beatty, S. D. Reynolds, and D. L. Morgan. Do they stay or do they go? Acoustic monitoring of whale sharks at Ningaloo Marine Park, Western Australia. Journal of Fish Biology, 91(6):1713–1720, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akhilesh:2017:BOB

[2468] K. V. Akhilesh, G. B. Purushottama, Thakurdas, and S. J. Kizhakudan. Biological observations on the broadfin shark Lamiopsis temminckii (Carcharhiniformes: Carcharhinidae). Journal of Fish Biology, 91(6): 1721–1729, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barry:2017:CRP

[2469] T. P. Barry, G. K. Dehnert, P. D. Hoppe, and P. W. Sorensen. Chemicals released by predation increase the growth rate of yellow perch, Perca flavescens. Journal of Fish Biology, 91(6):1730–1736, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Musseau:2017:CNN

[2470] C. Musseau, S. Vincenzi, D. Jesenšek, and A. J. Crivelli. Cannibalism in non-native brown trout Salmo trutta and rainbow trout Oncorhynchus mykiss stream-dwelling populations. Journal of Fish Biology, 91(6): 1737–1744, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quiroga:2017:IRT

[2471] S. Quiroga, F. Kacoliris, I. García, H. Povedano, M. Velasco, and S. Zalba. Invasive rainbow trout Oncorhynchus mykiss preying on the endangered naked characin Gymnocharacinus bergii at its thermal limits. Journal of Fish Biology, 91(6):1745–1749, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2017:GAM

[2472] Paul J. B. Hart. Global atlas of marine fisheries: a critical appraisal of catches and ecosystem impacts. Journal of Fish Biology, 91(6):1750– 1751, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proudlove:2017:EDD

[2473] G. Proudlove. Evolution in the dark: Darwin’s loss without selection. Journal of Fish Biology, 91(6):1752–1753, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2017:IIx

[2474] Anonymous. Issue information. Journal of Fish Biology, 91(6):1754– 1763, December 2017. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIa

[2475] Anonymous. Issue information. Journal of Fish Biology, 92(1):1–2, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2018:HDF

[2476] Y. Z. Wang, J. F. Sun, A. J. Lv, S. L. Zhang, Y. Y. Sung, H. Y. Shi, X. C. Hu, S. J. Chen, and K. Z. Xing. Histochemical distribution of four types of enzymes and mucous cells in the gastrointestinal tract of reared half-smooth tongue sole Cynoglossus semilaevis. Journal of Fish Biology, 92(1):3–16, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2018:GPF

[2477] Y. Li, Z. Jia, X. Liang, D. Matulic, M. Hussein, and J. Gao. Growth performance, fatty-acid composition, lipid deposition and hepatic-lipid metabolism-related gene expression in juvenile pond loach Misgurnus anguillicaudatus fed diets with different dietary soybean oil levels. Journal of Fish Biology, 92(1):17–33, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yates:2018:STD

[2478] P. Yates, P. Ziegler, D. Welsford, J. McIvor, B. Farmer, and E. Woodcock. Spatio-temporal dynamics in maturation and spawning of Patagonian toothfish Dissostichus eleginoides on the sub-Antarctic Kerguelen Plateau. Journal of Fish Biology, 92(1):34–54, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prado:2018:GSE

[2479] F. D. Do Prado, R. Fernandez-Cebrián, F. Foresti, C. Oliveira, P. Martínez, and F. Porto-Foresti. Genetic structure and evidence of anthropogenic effects on wild populations of two Neotropical catfishes: baselines for conservation. Journal of Fish Biology, 92(1):55–72, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OShea:2018:CFE

[2480] O. R. O’Shea, B. E. Wueringer, M. M. Winchester, and E. J. Brooks. Comparative feeding ecology of the yellow ray Urobatis jamaicensis (Urotrygonidae) from the Bahamas. Journal of Fish Biology, 92(1):73– 84, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meng:2018:EPD

[2481] F. X. Meng, M. Li, M. Z. Song, L. X. Yuan, Y. F. Gong, Y. X. Qian, G. Shi, and R. X. Wang. Di-2-ethylhexyl phthalate (DEHP) exposure disturbs lipid metabolism in juvenile yellow catfish Tachysurus fulvidraco. Journal of Fish Biology, 92(1):85–93, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dulski:2018:CGM

[2482] T. Dulski, Z. Zakeś, and S. Ciesielski. Characterization of the gut microbiota in early life stages of pikeperch Sander lucioperca. Journal of Fish Biology, 92(1):94–104, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:2018:RBF

[2483] H.-Y. Chang, C.-L. Sun, S.-Z. Yeh, Y.-J. Chang, N.-J. Su, and G. Di-Nardo. Reproductive biology of female striped marlin Kajikia audax in the western Pacific Ocean. Journal of Fish Biology, 92(1):105–130, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Longoni:2018:DSS

[2484] L. Schönhofen Longoni, J. Giora, and C. Bernhardt Fialho. Development of secondary sexual characters and their relationship to ontogeny and seasonal reproductive period in Hyphessobrycon igneus (Ostariophysi: Characiformes). Journal of Fish Biology, 92(1):131–149, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

South:2018:ITD

[2485] J. South, D. Welsh, A. Anton, J. D. Sigwart, and J. T. A. Dick. Increasing temperature decreases the predatory effect of the intertidal shanny Lipophrys pholis on an amphipod prey. Journal of Fish Biology, 92(1): 150–164, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daFonseca:2018:IMA

[2486] A. P. da Fonseca, M. V. Volcan, and R. B. Robaldo. Incubation media affect the survival, pathway and time of embryo development in Neotropical annual fish Austrolebias nigrofasciatus (Rivulidae). Journal of Fish Biology, 92(1):165–176, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Avey:2018:HEG

[2487] S. R. Avey, M. Ojehomon, J. F. Dawson, and T. E. Gillis. How the expression of green fluorescent protein and human cardiac actin in the heart influences cardiac function and aerobic performance in zebrafish Danio rerio. Journal of Fish Biology, 92(1):177–189, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Damerau:2018:MGP

[2488] M. Damerau, M. Freese, and R. Hanel. Multi-gene phylogeny of jacks and pompanos (Carangidae), including placement of monotypic vadigo Campogramma glaycos. Journal of Fish Biology, 92(1):190–202, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kullmann:2018:CME

[2489] B. Kullmann, M. Hempel, and R. Thiel. Chemical marking of European glass eels Anguilla anguilla with alizarin red S and in combination with strontium: in situ evaluation of short-term salinity effects on survival and efficient mass-marking. Journal of Fish Biology, 92(1):203–213, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2018:LMV

[2490] P. Sun and B. J. Tang. Low mtDNA variation and shallow population structure of the Chinese pomfret Pampus chinensis along the China coast. Journal of Fish Biology, 92(1):214–228, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birnie-Gauvin:2018:CVS

[2491] K. Birnie-Gauvin, K. S. Peiman, M. H. Larsen, K. Aarestrup, K. M. Gilmour, and S. J. Cooke. Comparison of vegetable shortening and cocoa butter as vehicles for cortisol manipulation in Salmo trutta. Journal of Fish Biology, 92(1):229–236, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kern:2018:MUE

[2492] P. Kern, R. L. Cramp, M. A. Gordos, J. R. Watson, and C. E. Franklin. Measuring U_{crit} and endurance: equipment choice influences estimates of fish swimming performance. Journal of Fish Biology, 92(1):237–247, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vazquez:2018:ABD

[2493] D. M. Vazquez, M. Belleggia, L. Schejter, and E. Mabragaña. Avoiding being dragged away: finding egg cases of Schroederichthys bivius (Chondrichthyes: Scyliorhinidae) associated with benthic invertebrates. Journal of Fish Biology, 92(1):248–253, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duchi:2018:FSN

[2494] A. Duchi. Flank spot number and its significance for systematics, taxonomy and conservation of the near-threatened Mediterranean trout Salmo cettii: evidence from a genetically pure population. Journal of Fish Biology, 92(1):254–260, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yabu:2018:FRN

[2495] M. H. S. Yabu, A. P. Vidotto-Magnoni, A. C. R. Casimiro, D. A. Z. Garcia, A. D. A. Costa, F. D. D. Prado, F. Porto-Foresti, and M. L. Orsi. First record of non-native hybrid catfish Pseudoplatystoma reticulatum × Leiarius marmoratus in the Upper Paraná River basin, Brazil. Journal of Fish Biology, 92(1):261–267, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeMeyer:2018:BHS

[2496] J. De Meyer, A. Herrel, C. Belpaire, G. Goemans, C. Ide, B. De Kegel, J. Christiaens, and D. Adriaens. Broader head, stronger bite: In vivo bite forces in European eel Anguilla anguilla. Journal of Fish Biology, 92(1):268–273, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reynisson:2018:PAL

[2497] H. Reynisson and G. Á. Ólafsdóttir. Plasticity in activity and latency to explore differs between juvenile Atlantic cod Gadus morhua across a temperature gradient. Journal of Fish Biology, 92(1):274–280, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIb

[2498] Anonymous. Issue information. Journal of Fish Biology, 92(1):281–290, January 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIc

[2499] Anonymous. Issue information. Journal of Fish Biology, 92(2):291–292, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Day:2018:LAY

[2500] J. W. Day, P. Sánchez Gil, and A. Sepúlveda Lozada. Luis alejandro Yáñez–Arancibia 1944–2016. Journal of Fish Biology, 92(2):293–297, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2018:ECB

[2501] X. F. Liu, Y. H. Wu, S. N. Wei, N. Wang, Y. Z. Li, N. W. Zhang, P. F. Li, Q. W. Qin, and S. L. Chen. Establishment and characterization of a brain-cell line from kelp grouper Epinephelus moara. Journal of Fish Biology, 92(2):298–307, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Compaire:2018:RRS

[2502] J. C. Compaire, P. Casademont, C. Gómez-Cama, and M. C. Soriguer. Reproduction and recruitment of sympatric fish species on an intertidal rocky shore. Journal of Fish Biology, 92(2):308–329, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jahnsen-Guzman:2018:PCA

[2503] N. Jahnsen-Guzmán, V. Bernal-Durán, and M. F. Landaeta. Parasitic copepods affect morphospace and diet of larvae of a temperate reef fish. Journal of Fish Biology, 92(2):330–346, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tang:2018:SPS

[2504] Q. Y. Tang, X. B. Li, D. Yu, Y. R. Zhu, B. Q. Ding, H. Z. Liu, and P. D. Danley. Saurogobio punctatus sp. nov., a new cyprinid gudgeon (Teleostei: Cypriniformes) from the Yangtze River, based on both morphological and molecular data. Journal of Fish Biology, 92(2):347–364, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konopinski:2018:GCP

[2505] M. K. Konopiński and A. Amirowicz. Genetic composition of a population of natural common bream Abramis brama × roach Rutilus rutilus hybrids and their morphological characteristics in comparison with parent species. Journal of Fish Biology, 92(2):365–385, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allman:2018:FAE

[2506] R. J. Allman, W. F. Patterson III, C. L. Fioramonti, and A. E. Pacicco. Factors affecting estimates of size at age and growth in grey triggerfish Balistes capriscus from the northern Gulf of Mexico. Journal of Fish Biology, 92(2):386–398, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Samways:2018:AFW

[2507] K. M. Samways, D. X. Soto, and R. A. Cunjak. Aquatic food-web dynamics following incorporation of nutrients derived from Atlantic anadromous fishes. Journal of Fish Biology, 92(2):399–419, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bas:2018:ESE

[2508] M. Bas and L. Cardona. Effects of skeletal element identity, delipidation and demineralization on the analysis of stable isotope ratios of C and N in fish bone. Journal of Fish Biology, 92(2):420–437, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chamon:2018:NSG

[2509] C. C. Chamon, T. N. A. Pereira, M. B. Mendonça, and A. Akama. New species of the genus Spectracanthicus (Loricariidae, Hypostominae, Ancistrini) from the Rio Javaés (Rio Araguaia basin), with a description of gross brain morphology. Journal of Fish Biology, 92(2):438–451, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eshaghzadeh:2018:SPS

[2510] H. Eshaghzadeh, A. Akbarzadeh, M. Yarmohammadi, and E. Gisbert. Skeletogenesis in the Persian sturgeon Acipenser persicus and its correlation with gene expression of vitamin K-dependent proteins during larval development. Journal of Fish Biology, 92(2):452–469, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Visconti:2018:LHT

[2511] V. Visconti, E. D. L. Trip, M. H. Griffiths, and K. D. Clements. Life-history traits of the leatherjacket Meuschenia scaber, a long-lived monacanthid. Journal of Fish Biology, 92(2):470–486, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2018:RTO

[2512] Y. Zhang, T. M. Healy, W. Vandersteen, P. M. Schulte, and A. P. Farrell. A rainbow trout Oncorhynchus mykiss strain with higher aerobic scope in normoxia also has superior tolerance of hypoxia. Journal of Fish Biology, 92(2):487–503, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gauthier:2018:MCA

[2513] A. R. G. Gauthier, D. L. Whitehead, I. R. Tibbetts, B. W. Cribb, and M. B. Bennett. Morphological comparison of the ampullae of Lorenzini of three sympatric benthic rays. Journal of Fish Biology, 92(2):504–514, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dutra:2018:NSC

[2514] G. M. Dutra, B. F. Melo, and A. L. Netto-Ferreira. A new species of Curimatopsis steindachner (Characiformes: Curimatidae) from the Rio Nhamundá, Amazon basin. Journal of Fish Biology, 92(2):515–522, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kent:2018:LLG

[2515] J. Kent, G. Jenkins, and C. D. H. Sherman. Low levels of genetic structuring in King George whiting Sillaginodes punctatus across two geographic regions. Journal of Fish Biology, 92(2):523–531, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2018:EVD

[2516] R. L. Moran, R. M. Soukup, M. Zhou, and R. C. Fuller. Egg viability decreases rapidly with time since ovulation in the rainbow darter Etheostoma caeruleum: implications for the costs of choosiness. Journal of Fish Biology, 92(2):532–536, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bornatowski:2018:HIN

[2517] H. Bornatowski, R. Loose, C. L. S. Sampaio, O. B. F. Gadig, A. Carvalho-Filho, and R. R. Domingues. Human introduction or natural dispersion? Atlantic Ocean occurrence of the Indo–Pacific whitetip reef shark Triaenodon obesus. Journal of Fish Biology, 92(2):537–542, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeWysiecki:2018:HCV

[2518] A. M. De Wysiecki, A. C. Milessi, R. Wiff, and A. J. Jaureguizar. Highest catch of the vulnerable broadnose sevengill shark Notorynchus cepedianus in the south-west Atlantic. Journal of Fish Biology, 92(2):543–548, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:Ca

[2519] Anonymous. Corrigendum. Journal of Fish Biology, 92(2):549, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tagliafico:2018:II

[2520] A. Tagliafico, S. Rangel, and M. K. Broadhurst. Issue information. Journal of Fish Biology, 92(2):550–559, February 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IId

[2521] Anonymous. Issue information. Journal of Fish Biology, 92(3):561, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simpson:2018:UFP

[2522] S. D. Simpson, I. Barber, and I. J. Winfield. Understanding fish populations. Journal of Fish Biology, 92(3):563–568, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2018:EIV

[2523] J. D. Armstrong, S. McKelvey, G. W. Smith, P. Rycroft, and R. J. Fryer. Effects of individual variation in length, condition and run-time on return rates of wild-reared Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 92(3):569–578, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gregory:2018:BRB

[2524] S. D. Gregory, J. D. Armstrong, and J. R. Britton. Is bigger really better? Towards improved models for testing how Atlantic salmon Salmo salar smolt size affects marine survival. Journal of Fish Biology, 92(3):579– 592, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomson:2018:CAB

[2525] M. Thomson and A. R. Lyndon. Comparing anadromous brown trout Salmo trutta in small, neighbouring catchments across contrasting landscapes: What is the role of environment in determining life-history characteristics? Journal of Fish Biology, 92(3):593–606, March 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cauwelier:2018:ART

[2526] E. Cauwelier, D. C. Stewart, C. P. Millar, J. Gilbey, and S. J. Middlemas. Across rather than between river genetic structure in Atlantic salmon Salmo salar in north-east Scotland, UK: potential causes and management implications. Journal of Fish Biology, 92(3):607–620, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ikediashi:2018:ASS

[2527] C. Ikediashi, J. R. Paris, R. A. King, W. R. C. Beaumont, A. Ibbotson, and J. R. Stevens. Atlantic salmon Salmo salar in the chalk streams of England are genetically unique. Journal of Fish Biology, 92(3):621– 641, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arevalo:2018:EFS

[2528] E. Arevalo, S. Panserat, I. Seiliez, A. Larrañaga, and A. Bardonnet. Effect of food shortage and temperature on age 0+ salmonids: a contribution to predict the effects of climate change. Journal of Fish Biology, 92(3):642–652, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2018:CST

[2529] C. J. Lee, C. R. Tyler, and G. C. Paull. Can simple tank changes benefit the welfare of laboratory zebrafish Danio rerio? Journal of Fish Biology, 92(3):653–659, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2018:LPS

[2530] I. M. Côté and N. S. Smith. The lionfish Pterois sp. invasion: Has the worst-case scenario come to pass? Journal of Fish Biology, 92(3):660– 689, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobs:2018:COF

[2531] A. Jacobs, C. Doran, D. S. Murray, J. Duffill Telsnig, K. L. Laskowski, N. A. R. Jones, S. K. Auer, and K. Præbel. On the challenges and opportunities facing fish biology: a discussion of five key knowledge gaps. Journal of Fish Biology, 92(3):690–698, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bacles:2018:EEN

[2532] C. F. E. Bacles, C. Bouchard, F. Lange, A. Manicki, C. Tentelier, and O. Lepais. Estimating the effective number of breeders from single parr samples for conservation monitoring of wild populations of Atlantic salmon Salmo salar. Journal of Fish Biology, 92(3):699–726, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paris:2018:UMF

[2533] J. R. Paris, K. D. Sherman, E. Bell, C. Boulenger, C. Delord, M. B. M. El-Mahdi, E. A. Fairfield, A. M. Griffiths, C. Gutmann Roberts, R. D. Hedger, L. E. Holman, L. H. Hooper, N. E. Humphries, I. Katsiadaki, R. A. King, A. Lemopoulos, C. J. Payne, G. Peirson, K. K. Richter, M. I. Taylor, C. N. Trueman, B. Hayden, and J. R. Stevens. Understanding and managing fish populations: keeping the toolbox fit for purpose. Journal of Fish Biology, 92(3):727–751, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salvanes:2018:AHE

[2534] A. G. V. Salvanes and M. J. Gibbons. Adaptation to hypoxic environments; bearded gobies Sufflogobius bibarbatus in the Benguela upwelling ecosystem. Journal of Fish Biology, 92(3):752–772, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:2018:NEP

[2535] S. C. Mills, J. L. O’Donnell, G. Bernardi, and R. Beldade. Natural endocrine profiles of the group-living skunk anemonefish Amphiprion akallopisos in relation to their size-based dominance hierarchy. Journal of Fish Biology, 92(3):773–789, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cheung:2018:FFF

[2536] W. W. L. Cheung. The future of fishes and fisheries in the changing oceans. Journal of Fish Biology, 92(3):790–803, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:2018:FCW

[2537] T. A. C. Gordon, H. R. Harding, F. K. Clever, I. K. Davidson, W. Davison, D. W. Montgomery, R. C. Weatherhead, F. M. Windsor, J. D. Armstrong, A. Bardonnet, E. Bergman, J. R. Britton, I. M. Côté, D. D’agostino, L. A. Greenberg, A. R. Harborne, K. K. Kahilainen, N. B. Metcalfe, S. C. Mills, N. J. Milner, F. H. Mittermayer, L. Montorio, S. L. Nedelec, J. M. Prokkola, L. A. Rutterford, A. G. V. Salvanes, S. D. Simpson, A. Vainikka, J. K. Pinnegar, and E. M. Santos. Fishes in a changing world: learning from the past to promote sustainability of fish populations. Journal of Fish Biology, 92(3):804–827, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Obregon:2018:VUF

[2538] C. Obregón, A. R. Lyndon, J. Barker, H. Christiansen, B. J. Godley, S. Kurland, J. J. Piccolo, R. Potts, R. Short, A. Tebb, and S. Mariani. Valuing and understanding fish populations in the Anthropocene: key questions to address. Journal of Fish Biology, 92(3):828–845, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lyndon:2018:IFS

[2539] A. R. Lyndon. Influence of the FSBI Small Research Grants scheme: an analysis and appraisal. Journal of Fish Biology, 92(3):846–850, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinnegar:2018:WDC

[2540] J. K. Pinnegar. Why the damselfish Chromis chromis is a key species in the Mediterranean rocky littoral — a quantitative perspective. Journal of Fish Biology, 92(3):851–872, March 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIe

[2541] Anonymous. Issue information. Journal of Fish Biology, 92(4):873–874, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2018:NSP

[2542] D. Turan, G. Kalayci, C. Kaya, Y. Bektaş, and F. Küçük. A new species of Petroleuciscus (Teleostei: Cyprinidae) from the Büyük Menderes River, southwestern Anatolia, Turkey. Journal of Fish Biology, 92(4): 875–887, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chatigny:2018:ETC

[2543] F. Chatigny, D. B. Groman, S. A. Martinson, and E. D. Stevens. Evaluation of tissue changes following intramuscular infiltration of lidocaine in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 92(4):888– 900, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lek:2018:DRD

[2544] E. Lek, M. E. Platell, D. V. Fairclough, N. G. Hall, and I. C. Potter. Diets of reef-dwelling labrids (choerodon species) vary with body size, season and habitat: influence of foraging ability, specialization and opportunism. Journal of Fish Biology, 92(4):901–928, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logan-Chesney:2018:ASA

[2545] L. M. Logan-Chesney, M. J. Dadswell, R. H. Karsten, I. Wirgin, and M. J. W. Stokesbury. Atlantic sturgeon Acipenser oxyrinchus surfacing behaviour. Journal of Fish Biology, 92(4):929–943, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Montana:2018:CEI

[2546] C. G. Montaña and C. M. Schalk. Correlates of elemental-isotopic composition of stream fishes: the importance of land-use, species identity and body size. Journal of Fish Biology, 92(4):944–960, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2018:MES

[2547] L. Persson, K. Leonardsson, and A. Alanärä. Manipulation of the energetic state of Atlantic salmon Salmo salar juveniles and the effect on migration speed. Journal of Fish Biology, 92(4):961–978, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2018:MCE

[2548] X. T. Zhang, K. J. Wei, Y. Y. Chen, Z. C. Shi, L. K. Liu, J. Li, G. R. Zhang, and W. Ji. Molecular cloning and expression analysis of tyr and tyrp1 genes in normal and albino yellow catfish Tachysurus fulvidraco. Journal of Fish Biology, 92(4):979–998, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Londono-Burbano:2018:NSS

[2549] A. Londoño-Burbano. A new species of Sturisoma (Loricariidae: Loricariinae) from the Madre de Dios River basin, Peru, with a key to all congeners and comments on the type series of Sturisoma rostratum. Journal of Fish Biology, 92(4):999–1015, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mion:2018:EFL

[2550] M. Mion, A. Thorsen, F. Vitale, J. Dierking, J. P. Herrmann, B. Huwer, B. von Dewitz, and M. Casini. Effect of fish length and nutritional condition on the fecundity of distressed Atlantic cod Gadus morhua from the Baltic Sea. Journal of Fish Biology, 92(4):1016–1034, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellender:2018:PFE

[2551] B. R. Ellender, O. L. F. Weyl, M. E. Alexander, A. M. Luger, L. A. J. Nagelkerke, and D. J. Woodford. Out of the pot and into the fire: Explaining the vulnerability of an endangered small headwater stream fish to black-bass Micropterus spp. invasion. Journal of Fish Biology, 92 (4):1035–1050, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruivo:2018:CGR

[2552] R. Ruivo, A. Capitão, L. F. C. Castro, and M. M. Santos. The cycling gonad: retinoic-acid synthesis and degradation patterns during adult zebrafish Danio rerio oogenesis. Journal of Fish Biology, 92(4):1051– 1064, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2018:CFI

[2553] J. B. Olsen, R. J. Brown, O. L. Russ, K. Harper, and J. K. Wenburg. Contemporary factors influencing genetic diversity in the Alaska humpback whitefish Coregonus clupeaformis complex. Journal of Fish Biology, 92(4):1065–1081, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niella:2018:CCS

[2554] Y. V. Niella, L. A. G. Duarte, V. R. Bandeira, O. Crespo, D. Beare, and F. H. V. Hazin. Cookie-cutter shark Isistius spp. predation upon different tuna species from the south-western Atlantic Ocean. Journal of Fish Biology, 92(4):1082–1089, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schmidt:2018:CGI

[2555] B. V. Schmidt and J. F. Schaefer. Comparative genetic isolation patterns for multiple headwater fishes in three geographic regions. Journal of Fish Biology, 92(4):1090–1109, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quimbayo:2018:CIG

[2556] J. P. Quimbayo and F. A. Zapata. Cleaning interactions by gobies on a tropical eastern Pacific coral reef. Journal of Fish Biology, 92(4):1110– 1125, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mello:2018:ELD

[2557] P. H. Mello, B. C. Araújo, M. F. Campos, J. A. Rodrigues-Filho, C. E. O. Garcia, and R. G. Moreira. Embryonic and larval development and fattyacid profile of Epinephelus marginatus spawned in captivity: tools applied to captive rearing. Journal of Fish Biology, 92(4):1126–1148, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Camelier:2018:NRS

[2558] P. Camelier, F. C. P. Dagosta, and M. M. F. Marinho. New remarkable sexually dimorphic miniature species of Hyphessobrycon (Characiformes: Characidae) from the upper Rio Tapajós basin. Journal of Fish Biology, 92(4):1149–1162, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gwilliam:2018:IGM

[2559] M. P. Gwilliam, A. C. Winkler, W. M. Potts, C. V. Santos, W. H. H. Sauer, P. W. Shaw, and N. J. McKeown. Integrated genetic and morphological data support eco-evolutionary divergence of Angolan and South African populations of Diplodus hottentotus. Journal of Fish Biology, 92 (4):1163–1176, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finucci:2018:VPO

[2560] B. Finucci, L. Gallus, A. Amaroli, S. Candiani, M. Rottigni, M. A. Masini, and S. Ferrando. Vacchi’s palatal organ: a widespread trait in holocephali. Journal of Fish Biology, 92(4):1177–1182, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maduna:2018:ESS

[2561] S. N. Maduna, J. H. Van Wyk, C. Da Silva, E. Gennari, and A. E. Bester-Van Der Merwe. Evidence for sperm storage in common smoothhound shark Mustelus mustelus and paternity assessment in a single litter from South Africa. Journal of Fish Biology, 92(4):1183–1191, April 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yuan:2018:PRS

[2562] X. Yuan, Y. P. Huang, L. Cai, D. Johnson, Z. Y. Tu, and Y. H. Zhou. Physiological responses to swimming fatigue in juvenile largemouth bronze gudgeon Coreius guichenoti. Journal of Fish Biology, 92 (4):1192–1197, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nobile:2018:DNN

[2563] A. B. Nobile, F. P. Lima, D. Freitas-Souza, J. Queiroz, D. A. Z. Garcia, M. L. Orsi, and A. P. Vidotto-Magnoni. Distribution of non-native suckermouth armoured catfish Pterygoplichthys ambrosettii in the upper Paraná River basin, Brazil. Journal of Fish Biology, 92(4):1198–1206, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becerril-Garcia:2018:FRP

[2564] E. E. Becerril-Garcia, D. Petatan-Ramirez, I. Ortiz-Aguirre, and A. Ayala-Bocos. First record of the Pacific seahorse Hippocampus ingens in Guadalupe Island, Mexico. Journal of Fish Biology, 92(4):1207– 1210, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Violi:2018:AEM

[2565] B. Violi, M. R. Gaither, F. Burns, A. Rus Hoelzel, and F. Neat. Assessing ecological and molecular divergence between the closely related species Hydrolagus pallidus and H. affinis (Chimaeridae). Journal of Fish Biology, 92(4):1211–1217, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2018:LCG

[2566] B. B. Gonçalves, N. F. Nascimento, M. P. Santos, R. M. Bertolini, G. S. Yasui, and P. C. Giaquinto. Low concentrations of glyphosate-based herbicide cause complete loss of sperm motility of yellowtail tetra fish Astyanax lacustris. Journal of Fish Biology, 92(4):1218–1224, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castellini:2018:AFS

[2567] D. L. Castellini, M. González-Castro, A. Lopez-Cazorla, and J. M. Díaz De Astarloa. Abnormalities in flatfishes of the south-west Atlantic Ocean. Journal of Fish Biology, 92(4):1225–1234, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohata:2018:OTH

[2568] R. Ohata, R. Masuda, and Y. Yamashita. Ontogeny of turbiditaxis in hatchery-reared Japanese anchovy Engraulis japonicus. Journal of Fish Biology, 92(4):1235–1240, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2018:IAS

[2569] R. Britton. Invasive alien species in aquatic environments: practical information and management insights. vol. 1. practical information & vol. 2 management insights. Journal of Fish Biology, 92(4):1241, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawrence:2018:ZMP

[2570] Christian Lawrence. Zebrafish: methods and protocols. second edition. Journal of Fish Biology, 92(4):1242–1243, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:R

[2571] Anonymous. Referees 2017. Journal of Fish Biology, 92(4):1244–1249, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:Aa

[2572] Anonymous. Announcement. Journal of Fish Biology, 92(4):1250–1252, April 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIf

[2573] Anonymous. Issue information. Journal of Fish Biology, 92(5):1253– 1254, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Middlemiss:2018:EGS

[2574] K. L. Middlemiss, D. G. Cook, A. R. Jerrett, and W. Davison. Effects of group size on school structure and behaviour in yellow-eyed mullet Aldrichetta forsteri. Journal of Fish Biology, 92(5):1255–1272, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ryu:2018:DCC

[2575] J. H. Ryu, M. S. Kim, J. H. Kang, D. H. Kim, Y. K. Nam, and S. P. Gong. Derivation of the clonal-cell lines from Siberian sturgeon (Acipenser baerii) head-kidney cell lines and its applicability to foreign gene expression and virus culture. Journal of Fish Biology, 92(5):1273– 1289, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavole:2018:UGT

[2576] L. M. Cavole, L. G. Cardoso, M. S. Almeida, and M. Haimovici. Unravelling growth trajectories from complicated otoliths — the case of Brazilian codling Urophycis brasiliensis. Journal of Fish Biology, 92(5):1290–1311, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2018:JLS

[2577] P. F. Amorim. Jenynsia lineata species complex, revision and new species description (Cyprinodontiformes: Anablepidae). Journal of Fish Biology, 92(5):1312–1332, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosales-Velazquez:2018:ACC

[2578] M. O. Rosales-Velázquez, A. F. Muttray, and R. H. Devlin. Assessing the capacity for compensatory growth in growth-hormone transgenic Coho salmon Oncorhynchus kisutch. Journal of Fish Biology, 92(5):1333–1341, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukuda:2018:LSA

[2579] N. Fukuda, H. Kurogi, D. Ambe, S. Chow, T. Yamamoto, K. Yokouchi, A. Shinoda, Y. Masuda, M. Sekino, K. Saitoh, M. Masujima, T. Watanabe, N. Mochioka, and H. Kuwada. Location, size and age at onset of metamorphosis in the Japanese eel Anguilla japonica. Journal of Fish Biology, 92(5):1342–1358, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polacik:2018:LVE

[2580] M. Polačik, M. Reichard, and M. Vrtílek. Local variation in embryo development rate in annual fish. Journal of Fish Biology, 92(5):1359– 1370, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crowe:2018:CLB

[2581] L. M. Crowe, O. O’Brien, T. H. Curtis, S. M. Leiter, R. D. Kenney, P. Duley, and S. D. Kraus. Characterization of large basking shark Cetorhinus maximus aggregations in the western North Atlantic Ocean. Journal of Fish Biology, 92(5):1371–1384, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2018:TPE

[2582] Y. Zhang, G. Claireaux, H. Takle, S. M. Jørgensen, and A. P. Farrell. A three-phase excess post-exercise oxygen consumption in Atlantic salmon Salmo salar and its response to exercise training. Journal of Fish Biology, 92(5):1385–1403, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:2018:NUP

[2583] R. O’Neill, N. Ó Maoiléidigh, P. McGinnity, N. Bond, and S. Culloty. The novel use of pop-off satellite tags (PSATs) to investigate the migratory behaviour of European sea bass Dicentrarchus labrax. Journal of Fish Biology, 92(5):1404–1421, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2018:IGC

[2584] X. L. Wu, Z. K. Tang, W. Li, Z. J. Chu, X. Y. Hong, X. P. Zhu, and H. Y. Xu. Identifying the germ cells during embryogenesis and gametogenesis by germ-line gene vasa in an anadromous fish, American shad Alosa sapidissima. Journal of Fish Biology, 92(5):1422–1434, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schulze:2018:HMD

[2585] M. J. Schulze, R. Henriques, K. A. Feldheim, R. C. K. Bowie, and S. von der Heyden. How many daddies: microsatellite genotyping reveals polyandry in a live-bearing clinid fish Muraenoclinus dorsalis. Journal of Fish Biology, 92(5):1435–1445, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2018:RDP

[2586] R. R. Silva and L. H. Rapp Py-Daniel. Redescription of Dekeyseria picta (Siluriformes; Loricariidae), a poorly known ancistrin from the Amazon and Orinoco Basins. Journal of Fish Biology, 92(5):1446–1462, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espindola:2018:INS

[2587] V. C. Espíndola, L. F. C. Tencatt, F. M. Pupo, L. Villa-Verde, and M. R. Britto. From the inside out: a new species of armoured catfish Corydoras with the description of poorly-explored character sources (Teleostei, Siluriformes, Callichthyidae). Journal of Fish Biology, 92(5):1463–1486, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arunrugstichai:2018:CCA

[2588] S. Arunrugstichai, J. D. True, and W. T. White. Catch composition and aspects of the biology of sharks caught by Thai commercial fisheries in the Andaman Sea. Journal of Fish Biology, 92(5):1487–1504, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Su:2018:SBS

[2589] J. J. Su, X. Y. Shang, Q. Y. Wan, and J. G. Su. SNP-based susceptibility– resistance association and mRNA expression regulation analyses of tlr7 to grass carp Ctenopharyngodon idella reovirus. Journal of Fish Biology, 92(5):1505–1525, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2018:AGM

[2590] S. Hagihara, J. Aoyama, D. Limbong, and K. Tsukamoto. Age and growth of migrating tropical eels, Anguilla celebesensis and Anguilla marmorata. Journal of Fish Biology, 92(5):1526–1544, May 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blabolil:2018:ABL

[2591] P. Blabolil, J. Duras, T. Jůza, L. Kočvara, J. Matěna, M. Muška, M. Říha, L. Vejřík, M. Holubová, and J. Peterka. Assessment of burbot Lota lota (L. 1758) population sustainability in central European reservoirs. Journal of Fish Biology, 92(5):1545–1559, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faustino:2018:UAB

[2592] F. Faustino, L. C. Makino, E. Neumann, and L. S. O. Nakaghi. Ultrastructure aspects of Brycon gouldingi (Teleostei, Characidae) related to swimming ability and feeding during larval development. Journal of Fish Biology, 92(5):1560–1573, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaffney:2018:CFS

[2593] K. A. Gaffney and M. M. Webster. Consistency of fish-shoal social network structure under laboratory conditions. Journal of Fish Biology, 92 (5):1574–1589, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See comment [2736].

Kikko:2018:JME

[2594] T. Kikko, D. Ishizaki, K. Kuwamura, H. Okamoto, M. Ujiie, A. Ide, J. Saegusa, Y. Kai, K. Nakayama, and Y. Fujioka. Juvenile migration of the exclusively pelagic cyprinid, Gnathopogon caerulescens (Honmoroko) in Lake Biwa, Central Japan. Journal of Fish Biology, 92(5):1590–1603, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murawala:2018:VEA

[2595] H. Murawala, S. Patel, I. Ranadive, I. Desai, and S. Balakrishnan. Variation in expression and activity pattern of mmp2 and mmp9 on different time scales in the regenerating caudal fin of Poecilia latipinna. Journal of Fish Biology, 92(5):1604–1619, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christensen:2018:BSP

[2596] E. A. F. Christensen and M. Grosell. Behavioural salinity preference of juvenile yellow perch Perca flavescens. Journal of Fish Biology, 92 (5):1620–1626, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colla:2018:UPA

[2597] S. Colla, F. Pranovi, R. Fiorin, S. Malavasi, and M. Picciulin. Using passive acoustics to assess habitat selection by the brown meagre Sciaena umbra in a northern Adriatic Sea mussel farm. Journal of Fish Biology, 92(5):1627–1634, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brackley:2018:CDL

[2598] R. Brackley, M. C. Lucas, R. Thomas, C. E. Adams, and C. W. Bean. Comparison of damage to live v. euthanized Atlantic salmon Salmo salar smolts from passage through an Archimedean screw turbine. Journal of Fish Biology, 92(5):1635–1644, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goatley:2018:EDT

[2599] C. H. R. Goatley, S. Wroe, S. B. Tebbett, and D. R. Bellwood. An evaluation of a double-tailed deformity in a coral-reef surgeonfish Acanthurus nigrofuscus (Acanthuridae) using micro-computed tomography. Journal of Fish Biology, 92(5):1645–1650, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McLennan:2018:TAS

[2600] D. McLennan, E. Rush, S. McKelvey, and N. B. Metcalfe. Timing of Atlantic salmon Salmo salar smolt migration predicts successful passage through a reservoir. Journal of Fish Biology, 92(5):1651–1656, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Navarro:2018:PIT

[2601] J. Navarro, A. Perezgrueso, C. Barría, and M. Coll. Photo-identification as a tool to study small-spotted catshark Scyliorhinus canicula. Journal of Fish Biology, 92(5):1657–1662, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jabado:2018:RCR

[2602] R. W. Jabado, P. M. Kyne, E. Nazareth, and D. N. Sutaria. A rare contemporary record of the Critically Endangered Ganges shark Glyphis gangeticus. Journal of Fish Biology, 92(5):1663–1669, May 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:Ab

[2603] Anonymous. Announcement. Journal of Fish Biology, 92(5):1670–1672, May 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIg

[2604] Anonymous. Issue information. Journal of Fish Biology, 92(6):1673– 1674, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wei:2018:ENF

[2605] S. Wei, Y. Yu, and Q. Qin. Establishment of a new fish cell line from the caudal fin of golden pompano Trachinotus ovatus and its susceptibility to iridovirus. Journal of Fish Biology, 92(6):1675–1686, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roxo:2018:HDN

[2606] F. F. Roxo, G. S. C. Silva, and B. F. Melo. Hisonotus devidei, a new species from the São Francisco basin, Brazil (Siluriformes: Loricariidae). Journal of Fish Biology, 92(6):1687–1701, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2018:VMS

[2607] J. R. Ellis, G. J. Burt, G. Grilli, S. R. McCully Phillips, T. L. Catchpole, and D. L. Maxwell. At-vessel mortality of skates (Rajidae) taken in coastal fisheries and evidence of longer-term survival. Journal of Fish Biology, 92(6):1702–1719, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loeb:2018:NSA

[2608] M. V. Loeb, H. R. Varella, and N. A. Menezes. A new species of Anchoviella (Clupeiformes: Engraulidae) from the western Amazon River in Peru, with comments on congeners in the Peruvian Amazon River. Journal of Fish Biology, 92(6):1720–1730, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boucher:2018:ESR

[2609] M. A. Boucher, D. W. Baker, C. J. Brauner, and J. M. Shrimpton. The effect of substrate rearing on growth, aerobic scope and physiology of larval white sturgeon Acipenser transmontanus. Journal of Fish Biology, 92(6):1731–1746, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mbandzi:2018:RFZ

[2610] N. Mbandzi, R. J. Wasserman, S. H. P. Deyzel, N. G. Vine, and A. K. Whitfield. River flow, zooplankton and dominant zooplanktivorous fish dynamics in a warm-temperate South African estuary. Journal of Fish Biology, 92(6):1747–1767, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lombarte:2018:ISO

[2611] A. Lombarte, M. Miletić, M. Kovačić, J. L. Otero-Ferrer, and V. M. Tuset. Identifying sagittal otoliths of Mediterranean Sea gobies: variability among phylogenetic lineages. Journal of Fish Biology, 92(6):1768– 1787, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franco-Belussi:2018:REJ

[2612] L. Franco-Belussi, C. De Oliveira, and H. N. Sköld. Regulation of eye and jaw colouration in three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 92(6):1788–1804, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teulier:2018:CEI

[2613] L. Teulier, L. Guillard, C. Leon, C. Romestaing, and Y. Voituron. Consequences of electroshock-induced narcosis in fish muscle: from mitochondria to swim performance. Journal of Fish Biology, 92(6):1805–1818, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lebeda:2018:FCA

[2614] I. Lebeda, C. Steinbach, and M. Flajšhans. Flow cytometry for assessing the efficacy of interspecific gynogenesis induction in sturgeon. Journal of Fish Biology, 92(6):1819–1831, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bayse:2018:PCL

[2615] S. M. Bayse, A. M. Regish, and S. D. McCormick. Proximate composition, lipid utilization and validation of a non-lethal method to determine lipid content in migrating American shad Alosa sapidissima. Journal of Fish Biology, 92(6):1832–1848, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2018:IAU

[2616] I. Moore, J. A. Dodd, M. Newton, C. W. Bean, I. Lindsay, P. Jarosz, and C. E. Adams. The influence of aquaculture unit proximity on the pattern of Lepeophtheirus salmonis infection of anadromous Salmo trutta populations on the Isle of Skye, Scotland. Journal of Fish Biology, 92 (6):1849–1865, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Angulo:2018:NSA

[2617] A. Angulo, A. C. Santos, M. López, F. Langeani, and C. D. Mcmahan. A new species of Astyanax (Characiformes: Characidae) from Costa Rica and Panama, with a key to the lower Central American species of the genus. Journal of Fish Biology, 92(6):1866–1887, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saemi-Komsari:2018:EDA

[2618] M. Saemi-Komsari, H. Mousavi-Sabet, C. F. Kratochwil, M. Sattari, S. Eagderi, and A. Meyer. Early developmental and allometric patterns in the electric yellow cichlid Labidochromis caeruleus. Journal of Fish Biology, 92(6):1888–1901, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrade:2018:NMS

[2619] M. C. Andrade, M. Jégu, P. A. Buckup, and A. L. Netto-Ferreira. A new Myleus species (Characiformes: Serrasalmidae) from the Rio Tapajós basin, Brazil. Journal of Fish Biology, 92(6):1902–1914, June 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faccioli:2018:CFD

[2620] C. K. Faccioli, R. A. Chedid, R. H. Mori, A. C. do Amaral, I. B. F. Vicentini, and C. A. Vicentini. Cytochemical features of the digestive tract mucosa of Hemisorubim platyrhynchos (Siluriformes: Pimelodidae). Journal of Fish Biology, 92(6):1915–1928, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Melo:2018:NSH

[2621] J. E. García-Melo, J. G. Albornoz-Garzón, L. J. García-Melo, F. A. Villa-Navarro, and J. A. Maldonado-Ocampo. A new species of Hemibrycon (Characiformes, Characidae, Stevardiinae) from the upper Magdalena River basin in Colombia. Journal of Fish Biology, 92(6):1929–1955, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cabrera-Castro:2018:MDF

[2622] R. Cabrera-Castro, C. Zabala, M. C. Soriguer, A. Domezain, and J. A. Hernando. Morphological development in the first life phase of Adriatic sturgeon Acipenser naccarii under controlled conditions. Journal of Fish Biology, 92(6):1956–1974, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Novelli:2018:EDL

[2623] B. Novelli, F. Otero Ferrer, J. A. Socorro, and L. Molina Domínguez. Early development of the longsnout seahorse Hippocampus reidi (Syngnathidae) within the male brood pouch. Journal of Fish Biology, 92 (6):1975–1984, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Charvet:2018:AGE

[2624] P. Charvet, F. M. Santana, K. L. De Lima, and R. Lessa. Age and growth of the endemic Xingu River stingray Potamotrygon leopoldi validated using fluorescent dyes. Journal of Fish Biology, 92(6):1985–1999, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claassens:2018:MPD

[2625] L. Claassens and A. N. Hodgson. Monthly population density and structure patterns of an endangered seahorse Hippocampus capensis: a comparison between natural and artificial habitats. Journal of Fish Biology, 92(6):2000–2015, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2018:TRM

[2626] R. J. Kennedy, M. Allen, and R. Wilson. Tag retention and mortality of adult Atlantic salmon Salmo salar gastrically tagged with different sized telemetry transmitters. Journal of Fish Biology, 92(6):2016–2021, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hou:2018:SPS

[2627] Y. Hou, L. Cai, X. Wang, X. Chen, D. Zhu, D. Johnson, and X. Shi. Swimming performance of 12 Schizothoracinae species from five rivers. Journal of Fish Biology, 92(6):2022–2028, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bett:2018:OGE

[2628] N. N. Bett, S. G. Hinch, K. H. Kaukinen, S. Li, and K. M. Miller. Olfactory gene expression in migrating adult sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 92(6):2029–2038, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koehler:2018:NRA

[2629] L. Koehler. New records of angular rough sharks Oxynotus centrina in the coastal waters of Malta, with observations on post-capture resilience and release behaviour. Journal of Fish Biology, 92(6):2039–2044, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Falconer:2018:CCI

[2630] L. Falconer. Climate change impacts on fisheries and aquaculture: a global analysis. Journal of Fish Biology, 92(6):2045–2046, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arantes:2018:BRF

[2631] C. Arantes. Book review: Field guide to the fishes of the Amazon, Orinoco & Guianas. Journal of Fish Biology, 92(6):2047–2048, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:Cb

[2632] Anonymous. Corrigendum. Journal of Fish Biology, 92(6):2049, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paris:2018:E

[2633] J. R. Paris, K. D. Sherman, E. Bell, C. Boulenger, C. Delord, M. B. M. El-Mahdi, E. A. Fairfield, A. M. Griffiths, C. Gutmann Roberts, R. D. Hedger, L. E. Holman, L. H. Hooper, N. E. Humphries, I. Katsiadaki, R. A. King, A. Lemopoulos, C. J. Payne, G. Peirson, K. K. Richter, M. I. Taylor, C. N. Trueman, B. Hayden, and J. R. Stevens. Erratum. Journal of Fish Biology, 92(6):2050, June 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIh

[2634] Anonymous. Issue information. Journal of Fish Biology, 93(1):1–2, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2018:EJ

[2635] A. P. Farrell. Editorial July 2018. Journal of Fish Biology, 93(1):3–4, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mo:2018:PSI

[2636] Tor A. Mo, Eva B. Thorstad, Odd T. Sandlund, Henrik H. Berntsen, Peder Fiske, and Ingebrigt Uglem. The pink salmon invasion: a Norwegian perspective. Journal of Fish Biology, 93(1):5–7, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See comment [2711].

Armstrong:2018:SIP

[2637] John D. Armstrong, Colin W. Bean, and Alan Wells. The Scottish invasion of pink salmon in 2017. Journal of Fish Biology, 93(1):8–11, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradley:2018:RBS

[2638] Daniel J. Bradley, William Gladstone, and Paul E. Gribben. Relationships between the spread of Caulerpa filiformis and fish communities on temperate rocky reefs. Journal of Fish Biology, 93(1):12–20, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2018:ISD

[2639] Seishi Hagihara, Jun Aoyama, Daniel Limbong, and Katsumi Tsukamoto. Interspecific and sexual differences in riverine distribution of tropical eels Anguilla spp. Journal of Fish Biology, 93(1):21–29, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbosa:2018:EFP

[2640] Thiago A. P. Barbosa, Daniela C. O. Rosa, Bruno E. Soares, Christophe H. A. Costa, Maria C. Esposito, and Luciano F. A. Montag. Effect of flood pulses on the trophic ecology of four piscivorous fishes from the eastern Amazon. Journal of Fish Biology, 93(1):30–39, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morbey:2018:FBD

[2641] Yolanda E. Morbey. Female-biased dimorphism in size and age at maturity is reduced at higher latitudes in lake whitefish Coregonus clupeaformis. Journal of Fish Biology, 93(1):40–46, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nissling:2018:PSS

[2642] Anders Nissling and Roger Larsson. Population specific sperm production in European flounder Platichthys flesus: Adaptation to salinity at spawning. Journal of Fish Biology, 93(1):47–52, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Basak:2018:SAT

[2643] Reetuparna Basak, Alivia Roy, and Umesh Rai. In silico analysis, temporal expression and gonadotropic regulation of receptors for folliclestimulating hormone and luteinizing hormone in testis of spotted snakehead Channa punctata. Journal of Fish Biology, 93(1):53–71, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ofelio:2018:HDL

[2644] Claudia Ofelio, Alcira Ofelia Díaz, Giuseppe Radaelli, and Miquel Planas. Histological development of the long-snouted seahorse Hippocampus guttulatus during ontogeny. Journal of Fish Biology, 93(1):72–87, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2018:DNS

[2645] Wei-Jun Wu, An-You He, Jun-Xing Yang, and Li-Na Du. Description of a new species of Triplophysa (Teleostei: Nemacheilidae) from Guizhou Province, China. Journal of Fish Biology, 93(1):88–94, July 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greenwell:2018:RDF

[2646] Claire N. Greenwell, Peter G. Coulson, James R. Tweedley, and Neil R. Loneragan. Regional differences in the feeding of the ambush predator Neosebastes pandus and comparisons of diets in the Scorpaenidae, Triglidae and Platycephalidae. Journal of Fish Biology, 93(1):95–109, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abrahao:2018:CVN

[2647] Vitor Pimenta Abrahão and Mario C. C. de Pinna. Cetopsis varii, a new species of whale catfish (Siluriformes: Cetopsidae) from the Meta Basin, Colombia. Journal of Fish Biology, 93(1):110–118, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rieucau:2018:UUA

[2648] Guillaume Rieucau, Jeremy J. Kiszka, Jose Carlos Castillo, Johann Mourier, Kevin M. Boswell, and Michael R. Heithaus. Using unmanned aerial vehicle (UAV) surveys and image analysis in the study of large surface-associated marine species: a case study on reef sharks Carcharhinus melanopterus shoaling behaviour. Journal of Fish Biology, 93(1): 119–127, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marta:2018:CSA

[2649] Anatol Marta, Natalia Kamenshchik, and Ion Toderas. Change of strategy or always overlooked? Third recent record of the beardless tadpole goby Benthophiloides brauneri from an inland location far from the coast. Journal of Fish Biology, 93(1):128–131, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verba:2018:MMO

[2650] Julia Tovar Verba, Manoela Lima de Oliveira Borges, Maria Nazareth Ferreira da Silva, Lorena Costa Pinto, and José Gurgel Rabello Neto. Mice on menu: opportunistic feeding behaviour of the Amazonian silver arowana Osteoglossum bicirrhosum. Journal of Fish Biology, 93(1):132– 133, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2018:MSA

[2651] Richard J. Kennedy, Robert Rosell, Michael Millane, Dennis Doherty, and Michelle Allen. Migration and survival of Atlantic salmon Salmo salar smolts in a large natural lake. Journal of Fish Biology, 93(1):134– 137, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raby:2018:EEC

[2652] Graham D. Raby, Josefin Sundin, Fredrik Jutfelt, Steven J. Cooke, and Timothy D. Clark. Exposure to elevated carbon dioxide does not impair short-term swimming behaviour or shelter-seeking in a predatory coralreef fish. Journal of Fish Biology, 93(1):138–142, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez-Hernandez:2018:ELB

[2653] Javier Sánchez-Hernández and Fernando Cobo. Examining the link between dietary specialization and foraging modes of stream-dwelling brown trout Salmo trutta. Journal of Fish Biology, 93(1):143–146, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zak:2018:IRG

[2654] Jakub Žák, Tomáš Jůza, Petr Blabolil, Roman Baran, Daniel Bartoň, Vladislav Draštík, Jaroslava Frouzová, Michaela Holubová, Henk A. M. Ketelaars, Luboš Kočvara, Jan Kubečka, Tomáš Mrkvička, Milan Muška, Milan Říha, Zuzana Sajdlová, Marek Šmejkal, Michal Tušer, Mojmír Vašek, Lukáš Vejřík, Ivana Vejříková, and Arco J. Wagenvoort. Invasive round goby Neogobius melanostomus has sex-dependent locomotor activity and is under-represented in catches from passive fishing gear compared with seine catches. Journal of Fish Biology, 93(1):147–152, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chase:2018:DFW

[2655] Tory John Chase, Jessica P. Nowicki, and Darren J. Coker. Diurnal foraging of a wild coral-reef fish Parapercis australis in relation to latesummer temperatures. Journal of Fish Biology, 93(1):153–158, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honkanen:2018:CMP

[2656] Hannele M. Honkanen, Jessica R. Rodger, Alastair Stephen, Keith Adams, James Freeman, and Colin E. Adams. Counterintuitive migration patterns by Atlantic salmon Salmo salar smolts in a large lake.

Journal of Fish Biology, 93(1):159–162, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mollen:2018:CAC

[2657] Frederik H. Mollen. Checklist anxiety: the case for sharks, skates and rays. Journal of Fish Biology, 93(1):163–164, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weyl:2018:JFB

[2658] Olaf Weyl, Albert Chakona, and Ian Harrison. Journal of Fish Biology special issue on biology and ecology of African freshwater fishes. Journal of Fish Biology, 93(1):165, July 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIi

[2659] Anonymous. Issue information. Journal of Fish Biology, 93(2):167–168, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2018:GBU

[2660] Tony Farrell. The good, the bad, but not the ugly. Journal of Fish Biology, 93(2):169, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soares:2018:ISM

[2661] Marta C. Soares, Robert Gerlai, and Caio Maximino. The integration of sociality, monoamines and stress neuroendocrinology in fish models: Applications in the neurosciences. Journal of Fish Biology, 93(2):170– 191, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadswell:2018:LTE

[2662] Michael J. Dadswell, Aaron D. Spares, Montana F. Mclean, Patrick J. Harris, and Roger A. Rulifson. Long-term effect of a tidal, hydroelectric propeller turbine on the populations of three anadromous fish species. Journal of Fish Biology, 93(2):192–206, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brownscombe:2018:EFS

[2663] Jacob W. Brownscombe, Robert J. Lennox, Andy J. Danylchuk, and Steven J. Cooke. Estimating fish swimming metrics and metabolic rates with accelerometers: the influence of sampling frequency. Journal of Fish Biology, 93(2):207–214, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2018:RSN

[2664] Xiu L. Chen, Eei Y. Lui, Yuen K. Ip, and Siew H. Lam. RNA sequencing, de novo assembly and differential analysis of the gill transcriptome of freshwater climbing perch Anabas testudineus after 6 days of seawater exposure. Journal of Fish Biology, 93(2):215–228, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corona-Herrera:2018:EEM

[2665] Guillermo A. Corona-Herrera, Silvia E. Arranz, Carlos A. Martínez-Palacios, Pamela Navarrete-Ramírez, Elva M. Toledo-Cuevas, Juan J. Valdez-Alarcón, and Carlos C. Martínez-Chávez. Experimental evidence of masculinization by continuous illumination in a temperature sex determination teleost (Atherinopsidae) model: is oxidative stress involved? Journal of Fish Biology, 93(2):229–237, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neves:2018:RPK

[2666] Mayara P. Neves, Jislaine C. Da Silva, Dirceu Baumgartner, Gilmar Baumgartner, and Rosilene L. Delariva. Is resource partitioning the key? The role of intra-interspecific variation in coexistence among five small endemic fish species (Characidae) in subtropical rivers. Journal of Fish Biology, 93(2):238–249, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rubio:2018:DCB

[2667] Kathryn S. Rubio, Matthew Ajemian, Gregory W. Stunz, Terence A. Palmer, Benoit Lebreton, and Jennifer Beseres Pollack. Dietary composition of black drum Pogonias cromis in a hypersaline estuary reflects water quality and prey availability. Journal of Fish Biology, 93(2):250– 262, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atkinson:2018:OLL

[2668] Emma M. Atkinson, Andrew W. Bateman, Lawrence M. Dill, Martin Krkošek, John D. Reynolds, and Sean C. Godwin. Oust the louse: leaping behaviour removes sea lice from wild juvenile sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 93(2):263–271, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singkam:2018:ODW

[2669] Abdul R. Singkam and Andrew D. C. MacColl. Otolith development in wild populations of stickleback: Jones & Hynes method does not apply to most populations. Journal of Fish Biology, 93(2):272–281, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2018:BSC

[2670] Allison J. Santos, Alyssa R. Frederick, Benjamin A. Higgins, Andres Carrillo, Ariel L. Carter, Kathryn A. Dickson, Donovan P. German, and Michael H. Horn. The beach-spawning California grunion Leuresthes tenuis eats and digests conspecific eggs. Journal of Fish Biology, 93 (2):282–289, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lei:2018:EPM

[2671] Cai X. Lei, Jing J. Tian, Hong Ji, and Yang Li. EPA plays multiple roles in regulating lipid accumulation of grass carp Ctenopharyngodon idella adipose tissue in vitro and in vivo. Journal of Fish Biology, 93(2):290– 301, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tougard:2018:SMT

[2672] Christelle Tougard, Fabienne Justy, Bruno Guinand, Emmanuel J. P. Douzery, and Patrick Berrebi. Salmo macrostigma (Teleostei, Salmonidae): Nothing more than a brown trout (S. trutta) lineage? Journal of Fish Biology, 93(2):302–310, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berbel-Filho:2018:UCD

[2673] Waldir M. Berbel-Filho, Telton P. A. Ramos, Uedson P. Jacobina, Danielle J. G. Maia, Rodrigo A. Torres, and Sergio M. Q. Lima. Updated checklist and DNA barcode-based species delimitations reveal taxonomic uncertainties among freshwater fishes from the mid-north-eastern Caatinga ecoregion, north-eastern Brazil. Journal of Fish Biology, 93(2): 311–323, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stiver:2018:IMM

[2674] Kelly A. Stiver, Holly K. Kindsvater, Natascia Tamburello, Kellie L. Heckman, Joanne Klein, and Suzanne H. Alonzo. Intentional multiple mating by females in a species where sneak fertilization circumvents female choice for parental males. Journal of Fish Biology, 93(2):324–333, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:2018:GFC

[2675] Stine D. Ross, J. Rasmus Nielsen, Henrik Gislason, Anders Nielsen, and Niels G. Andersen. Growth and food consumption of whiting Merlangius merlangus. Journal of Fish Biology, 93(2):334–343, August 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2018:CMB

[2676] Guy M. W. Stevens, Julie P. Hawkins, and Callum M. Roberts. Courtship and mating behaviour of manta rays Mobula alfredi and M. birostris in the Maldives. Journal of Fish Biology, 93(2):344–359, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fenkes:2018:BSR

[2677] Miriam Fenkes, Holly A. Shiels, and Robert L. Nudds. Body shape and robustness response to water flow during development of brown trout Salmo trutta parr. Journal of Fish Biology, 93(2):360–369, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2018:DGP

[2678] Guido Plaza, Francisco Cerna, Mauricio F. Landaeta, Alejandra Hernández, and Jorge E. Contreras. Daily growth patterns and age-at-recruitment of the anchoveta Engraulis ringens as indicated by a multiannual analysis of otolith microstructure across developmental stages. Journal of Fish Biology, 93(2):370–381, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Law:2018:AGB

[2679] Calton S. W. Law and Yvonne Sadovy de Mitcheson. Age and growth of black seabream Acanthopagrus schlegelii (Sparidae) in Hong Kong and adjacent waters of the northern South China Sea. Journal of Fish Biology, 93(2):382–390, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lear:2018:MRE

[2680] Karissa O. Lear, Adrian C. Gleiss, and Nicholas M. Whitney. Metabolic rates and the energetic cost of external tag attachment in juvenile blacktip sharks Carcharhinus limbatus. Journal of Fish Biology, 93(2):391– 395, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:2018:VDS

[2681] Chia-Hao Chang, Jessica E. Schult, Jonathan Sanders, Shih-Hui Liu, and Rebecca C. Fuller. The vertical distributions and spawning site choices of red and yellow bluefin killifish Lucania goodei colour morphs. Journal of Fish Biology, 93(2):396–400, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Martinez:2018:AWS

[2682] Juana López-Martínez, Marcos Porchas-Quijada, Francisco J. Álvarez-Tello, and Marco A. Porchas-Cornejo. Association of the whale shark Rhincodon typus with the cannonball jellyfish Stomolophus meleagris. Journal of Fish Biology, 93(2):401–404, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diedericks:2018:SEI

[2683] Genevieve Diedericks, Romina Henriques, Sophie von der Heyden, Olaf L. F. Weyl, and Cang Hui. Sleeping with the enemy: introgressive hybridization in two invasive centrarchids. Journal of Fish Biology, 93(2): 405–410, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lastrucci:2018:UPR

[2684] Naomi S. Lastrucci, Lucas T. Nunes, Alberto Lindner, and Sergio R. Floeter. An updated phylogeny of the redlip blenny genus Ophioblennius. Journal of Fish Biology, 93(2):411–414, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2018:MIB

[2685] Samuel M. Williams, Julian G. Pepperell, Michael Bennett, and Jennifer R. Ovenden. Misidentification of istiophorid billfishes by fisheries observers raises uncertainty over stock status. Journal of Fish Biology, 93(2):415–419, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DaSilva:2018:RPG

[2686] João Paulo C. B. Da Silva and G. David Johnson. Reconsidering pectoral girdle and fin morphology in Anguillidae (Elopomorpha: Anguilliformes). Journal of Fish Biology, 93(2):420–423, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DiazAndrade:2018:OFC

[2687] María C. Díaz Andrade, Ana C. Moya, Anahí Wehitt, Edgardo E. Di Giácomo, and Elena J. Galíndez. Observations of follicle cell processes in a holocephalan. Journal of Fish Biology, 93(2):424–427, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hassell:2018:FRI

[2688] Kathryn L. Hassell, Christopher A. Rawson, and Marthe M. Gagnon. First report of intersex in a lutjanid: the goldband snapper Pristipomoides multidens. Journal of Fish Biology, 93(2):428–431, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2018:MEF

[2689] Laith A. Jawad. Marine ecological field methods: a guide for marine biologists and fisheries scientists. Journal of Fish Biology, 93(2):432, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2018:FEP

[2690] Ian J. Winfield. Freshwater ecosystems in protected areas: conservation and management. Journal of Fish Biology, 93(2):433, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2018:SBS

[2691] Ian J. Winfield. Sea bass and sea bream: a practical approach to disease control and health management. Journal of Fish Biology, 93(2): 434, August 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIj

[2692] Anonymous. Issue information. Journal of Fish Biology, 93(3):435, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCormick:2018:AOU

[2693] S. D. McCormick, J. M. Shrimpton, T. O. Nilsen, and L. O. Ebbesson. Advances in our understanding of the parr–smolt transformation of juvenile salmon: a summary of the 10th International Workshop on Salmon Smoltification. Journal of Fish Biology, 93(3):437–439, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strand:2018:PRT

[2694] Jo E. T. Strand, David Hazlerigg, and Even H. Jørgensen. Photoperiod revisited: is there a critical day length for triggering a complete parr– smolt transformation in Atlantic salmon Salmo salar? Journal of Fish Biology, 93(3):440–448, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Self:2018:ERS

[2695] Katharine E. Self, Carl B. Schreck, Karen M. Cogliati, Eric J. Billman, and David L. G. Noakes. The effect of rearing structures on behaviour and movement of juvenile steelhead Oncorhynchus mykiss. Journal of Fish Biology, 93(3):449–454, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barlaup:2018:SMT

[2696] Bjørn T. Barlaup, Hans Rund, Eirik S. Normann, Sebastian Stranzl, Shad Mahlum, and Knut W. Vollset. Out of sync: monitoring the time of sea entry of wild and hatchery salmon Salmo salar smolt using floating passive-integrated transponder antennae. Journal of Fish Biology, 93(3): 455–464, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Self:2018:ESG

[2697] Katharine E. Self, Carl B. Schreck, Karen M. Cogliati, Eric J. Billman, and David L. G. Noakes. Egg size and growth in steelhead Oncorhynchus mykiss. Journal of Fish Biology, 93(3):465–468, September 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2018:ESH

[2698] Andy Moore, Lucia Privitera, Mark J. Ives, Jane Uzyczak, and William R. C. Beaumont. The effects of a small hydropower scheme on the migratory behaviour of Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 93(3):469–476, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riley:2018:UWC

[2699] William D. Riley, Anton T. Ibbotson, Stephen D. Gregory, Ian C. Russell, Rasmus B. Lauridsen, William R. C. Beaumont, Alastair C. Cook, and David L. Maxwell. Under what circumstances does the capture and tagging of wild Atlantic salmon Salmo salar smolts affect probability of return as adults? Journal of Fish Biology, 93(3):477–489, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Inatani:2018:ATD

[2700] Yu Inatani, Toshinao Ineno, Shiori Sone, Naoto Matsumoto, Katsuhisa Uchida, and Munetaka Shimizu. Assessment of the timing and degree of smolt development in southern populations of masu salmon Oncorhynchus masou. Journal of Fish Biology, 93(3):490–500, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elsner:2018:DSW

[2701] Rick A. Elsner and J. Mark Shrimpton. Is the duration of the smolt window related to migration distance in Coho salmon Oncorhynchus kisutch? Journal of Fish Biology, 93(3):501–509, September 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bassett:2018:TSD

[2702] Marley C. Bassett, David A. Patterson, and J. Mark Shrimpton. Temporal and spatial differences in smolting among Oncorhynchus nerka populations throughout fresh and seawater migration. Journal of Fish Biology, 93(3):510–518, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundh:2018:ECB

[2703] Henrik Sundh, Albin Gräns, Jeroen Brijs, Erik Sandblom, Michael Axelsson, Charlotte Berg, and Kristina Sundell. Effects of coeliacomesenteric blood flow reduction on intestinal barrier function in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 93(3):519–527, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mizuno:2018:EDS

[2704] Shinya Mizuno, Shigehiko Urawa, Mahito Miyamoto, Makoto Hatakeyama, Yoshitaka Sasaki, Nobuhisa Koide, Shoichi Tada, and Hiroshi Ueda. Effects of dietary supplementation with oregano essential oil on prevention of the ectoparasitic protozoans Ichthyobodo salmonis and Trichodina truttae in juvenile chum salmon Oncorhynchus keta. Journal of Fish Biology, 93(3):528–539, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2018:PSA

[2705] Morgan S. Brown, Paul L. Jones, Jared J. Tromp, Christian A. van Rijn, Robert A. Collins, and Luis O. B. Afonso. The physiology of saltwater acclimation in large juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 93(3):540–549, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vargas-Chacoff:2018:EET

[2706] Luis Vargas-Chacoff, Amy M. Regish, Andrew Weinstock, and Stephen D. McCormick. Effects of elevated temperature on osmoregulation and stress responses in Atlantic salmon Salmo salar smolts in fresh water and seawater. Journal of Fish Biology, 93(3):550–559, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCormick:2018:EOA

[2707] Stephen D. McCormick and Amy M. Regish. Effects of ocean acidification on salinity tolerance and seawater growth of Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 93(3):560–566, September 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balseiro:2018:CBA

[2708] Pablo Balseiro, Øyvind Moe, Ingrid Gamlem, Munetaka Shimizu, Harald Sveier, Tom O. Nilsen, Nobuto Kaneko, Lars Ebbesson, Cindy Pedrosa, Valentina Tronci, Are Nylund, and Sigurd O. Handeland. Comparison between Atlantic salmon Salmo salar post-smolts reared in open sea cages and in the Preline raceway semi-closed containment aquaculture system. Journal of Fish Biology, 93(3):567–579, September 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIk

[2709] Anonymous. Issue information. Journal of Fish Biology, 93(4):581–582, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2018:PGE

[2710] David McKenzie. Patterns of gene expression in larvae of the giant grouper. Journal of Fish Biology, 93(4):583, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2018:CPS

[2711] Nina Jonsson and Bror Jonsson. Comment on “The pink salmon invasion: a Norwegian perspective” (Mo et al., 2018). Journal of Fish Biology, 93 (4):584–585, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [2636].

Isbert:2018:PCW

[2712] Wolf Isbert, Francisco E. Montero, Ana Pérez del Olmo, Àngel López-Sanz, Olga Reñones, and Covadonga Orejas. Parasite communities of the white seabream Diplodus sargus sargus in the marine protected area of Medes Islands, north-west Mediterranean Sea. Journal of Fish Biology, 93(4):586–596, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otwoma:2018:GPS

[2713] Levy M. Otwoma, Valeska Diemel, Hauke Reuter, Marc Kochzius, and Achim Meyer. Genetic population structure of the convict surgeonfish Acanthurus triostegus: a phylogeographic reassessment across its range. Journal of Fish Biology, 93(4):597–608, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tang:2018:CTD

[2714] Ni Tang, Jin Hao, Xin Zhang, Yuan B. Wu, Shu Y. Wang, Jin W. Qi, Zhen Z. Tian, Bin Wang, Hu Chen, De F. Chen, and Zhi Q. Li. Characterization, tissue distribution of apela and periprandial, fasting and refeeding changes of apela mRNA in Siberian sturgeon Acipenser baerii. Journal of Fish Biology, 93(4):609–615, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waraniak:2018:MDA

[2715] Justin M. Waraniak, Edward A. Baker, and Kim T. Scribner. Molecular diet analysis reveals predator–prey community dynamics and environmental factors affecting predation of larval lake sturgeon Acipenser fulvescens in a natural system. Journal of Fish Biology, 93(4):616–629, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tonkin:2018:SSC

[2716] Zeb Tonkin, Jarod P. Lyon, Paul Moloney, Stephen R. Balcombe, and Graeme Hackett. Spawning-stock characteristics and migration of a lakebound population of the endangered Macquarie perch Macquaria australasica. Journal of Fish Biology, 93(4):630–640, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bone:2018:ECC

[2717] Bryden Bone and Gail K. Davoren. Egg cannibalism in capelin Mallotus villosus at beach and deep-water spawning habitats in the north-west Atlantic Ocean. Journal of Fish Biology, 93(4):641–648, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manning:2018:HIE

[2718] Clayton G. Manning, Sarah J. Foster, David Harasti, and Amanda C. J. Vincent. A holistic investigation of the ecological correlates of abundance and body size for the endangered White’s seahorse Hippocampus whitei. Journal of Fish Biology, 93(4):649–663, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gossieaux:2018:IHB

[2719] Philippine Gossieaux, Pascal Sirois, Louis Bernatchez, and Dany Garant. Introgressive hybridization between wild and domestic individuals and its relationship with parasitism in brook charr Salvelinus fontinalis. Journal of Fish Biology, 93(4):664–673, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bose:2018:GOM

[2720] Aneesh P. H. Bose, Erin S. McCallum, Kate Raymond, Julie R. Marentette, and Sigal Balshine. Growth and otolith morphology vary with alternative reproductive tactics and contaminant exposure in the round goby Neogobius melanostomus. Journal of Fish Biology, 93(4):674–684, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wieczorek:2018:SIA

[2721] Alina M. Wieczorek, Anne Marie Power, Patricia Browne, and Conor T. Graham. Stable-isotope analysis reveals the importance of soft-bodied prey in the diet of lesser spotted dogfish Scyliorhinus canicula. Journal of Fish Biology, 93(4):685–693, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2018:TIA

[2722] Kelli C. Anderson, Richard Knuckey, Maximilano Cánepa, and Abigail Elizur. A transcriptomic investigation of appetite-regulation and digestive processes in giant grouper Epinephelus lanceolatus during early larval development. Journal of Fish Biology, 93(4):694–710, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendes:2018:RBT

[2723] Yanne A. Mendes, James T. Lee, Ivana K. S. Viana, Rossineide M. Rocha, and Maria A. P. Ferreira. Reproductive biology of the tiger pleco Panaqolus tankei (Loricariidae) in a lentic system of the Amazon Basin. Journal of Fish Biology, 93(4):711–714, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:2018:SRB

[2724] Arthur B. Bauer, Luciano G. Fischer, Acácio R. G. Tomás, Roberta A. Santos, and Michael M. Mincarone. The southernmost records of Brotula barbata (Bloch & Schneider 1801) (Ophidiiformes: Ophidiidae) in the western Atlantic Ocean. Journal of Fish Biology, 93(4):715–718, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bramburger:2018:PID

[2725] Andrew J. Bramburger, Katherine E. Moir, and M. Brian C. Hickey. Preferential incorporation of dark, coloured materials into nests by a mound-nesting stream cyprinid. Journal of Fish Biology, 93(4):719–722, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waltham:2018:TAE

[2726] Nathan J. Waltham and Jason R. Schaffer. Thermal and asphyxia exposure risk to freshwater fish in feral-pig-damaged tropical wetlands. Journal of Fish Biology, 93(4):723–728, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagihara:2018:IDD

[2727] Seishi Hagihara, Jun Aoyama, Daniel Limbong, and Katsumi Tsukamoto. Interspecific difference in downstream migratory season between two tropical eels, Anguilla celebesensis and Anguilla marmorata. Journal of Fish Biology, 93(4):729–732, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banon:2018:GRE

[2728] Rafael Bañón, Alexandre Alonso-Fernández, David Barros-García, María-Berta Rios, and Alejandro de Carlos. Geographic range expansion of Ephippion guttifer (Tetraodontidae) in the north-eastern Atlantic. Journal of Fish Biology, 93(4):733–737, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Concha:2018:FOC

[2729] Francisco Concha, Naití Morales, and Sebastián Hernández. First observations on captive hatching and incubation period of the yellow-nose skate Dipturus chilensis (Rajiformes: Rajidae), from the South-Eastern Pacific Ocean. Journal of Fish Biology, 93(4):738–740, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richu:2018:PDP

[2730] Avinipully Richu, Neelesh Dahanukar, Anvar Ali, Kutty Ranjeet, and Rajeev Raghavan. Population dynamics of a poorly known serranid, the duskytail grouper Epinephelus bleekeri in the Arabian Sea. Journal of Fish Biology, 93(4):741–744, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castro:2018:OAC

[2731] Kathleen M. da Silva Alcântara Castro, Maiara P. do Santos, Marcelo F. G. Brito, Claudio J. Bidau, and Pablo A. Martinez. Ontogenetic allometry conservatism across five teleost orders. Journal of Fish Biology, 93(4):745–749, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2018:ISR

[2732] Kentaro Morita, Jun ichi Tsuboi, Genki Sahashi, Takeshi Kikko, Daisuke Ishizaki, Daisuke Kishi, Shinsuke Endo, and Yusuke Koseki. Iteroparity of stream resident masu salmon Oncorhynchus masou. Journal of Fish Biology, 93(4):750–754, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bachman:2018:GRE

[2733] Beverly A. Bachman, Richard Kraus, Cheston T. Peterson, Ralph Dean Grubbs, and Esther C. Peters. Growth and reproduction of Echeneis naucrates from the eastern Gulf of Mexico. Journal of Fish Biology, 93 (4):755–758, October 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIl

[2734] Anonymous. Issue information. Journal of Fish Biology, 93(5):759–760, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Graham:2018:LTS

[2735] Nicholas A. J. Graham. Long-term studies in the Philippines illuminate the relative role of marine reserves versus benthic degradation in driving coral reef fish densities. Journal of Fish Biology, 93(5):761, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolles:2018:CCF

[2736] Jolle W. Jolles. Comment on “Consistency of fish-shoal social network structure under laboratory conditions (Gaffney & Webster, 2018)”. Journal of Fish Biology, 93(5):762–763, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [2593].

Gaffney:2018:RJC

[2737] Katie A. Gaffney and Mike M. Webster. Reply to Jolles (2018): Comment on “Consistency of fish-shoal social network structure under laboratory conditions (Gaffney & Webster, 2018)”. Journal of Fish Biology, 93(5): 764–765, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McDowell:2018:HLG

[2738] Jan R. McDowell and Heid L. Brightman. High level of genetic connectivity in a deep-water reef fish, Caulolatilus microps. Journal of Fish Biology, 93(5):766–777, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schonhuth:2018:PWC

[2739] Susana Schönhuth, Roderick B. Gagne, Fernando Alda, David A. Neely, Richard L. Mayden, and Michael J. Blum. Phylogeography of the widespread creek chub Semotilus atromaculatus (Cypriniformes: Leuciscidae). Journal of Fish Biology, 93(5):778–791, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalayci:2018:GME

[2740] Gokhan Kalayci, Rafet C. Ozturk, Erol Capkin, and Ilhan Altinok. Genetic and molecular evidence that brown trout Salmo trutta belonging to the Danubian lineage are a single biological species. Journal of Fish Biology, 93(5):792–804, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Itakura:2018:DWS

[2741] Hikaru Itakura, Kohma Arai, Kenzo Kaifu, Kotaro Shirai, Akihito Yoneta, Yoichi Miyake, David H. Secor, and Shingo Kimura. Distribution of wild and stocked Japanese eels in the lower reaches of the Tone River catchment revealed by otolith stable-isotope ratios. Journal of Fish Biology, 93(5):805–813, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjornsson:2018:FSB

[2742] Björn Björnsson, Hjalti Karlsson, and Andreas Macrander. Food searching behaviour in adult Atlantic cod Gadus morhua during acoustic training: social learning and leadership within a school. Journal of Fish Biology, 93(5):814–829, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vanegas-Rios:2018:NSD

[2743] James A. Vanegas-Ríos, María de las M. Azpelicueta, and Luiz R. Malabarba. A new species of Diapoma (Characiformes, Characidae, Stevardiinae) from the Rio Paraná basin, with an identification key to the species of the genus. Journal of Fish Biology, 93(5):830–841, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logan:2018:EBC

[2744] Ryan K. Logan, Connor F. White, Chuck Winkler, Salvador J. Jorgensen, John B. O’Sullivan, Christopher G. Lowe, and Kady Lyons. An evaluation of body condition and morphometric relationships within southern California juvenile white sharks Carcharodon carcharias. Journal of Fish Biology, 93(5):842–849, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulmanns:2018:TDP

[2745] Nathalie Pülmanns, Gustavo Adolfo Castellanos-Galindo, and Uwe Krumme. Tidal-diel patterns in feeding and abundance of armed snook Centropomus armatus from macrotidal mangrove creeks of the tropical eastern Pacific Ocean. Journal of Fish Biology, 93(5):850–859, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatom-Naecker:2018:BFK

[2746] Theresa-Anne M. Tatom-Naecker and Mark W. Westneat. Burrowing fishes: Kinematics, morphology and phylogeny of sand-diving wrasses (Labridae). Journal of Fish Biology, 93(5):860–873, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janowicz:2018:AGR

[2747] Mariola E. Janowicz, Włodzimierz Załachowski, Agnieszka Rybczyk, Sheri Dalton, Eurielle Fernandes, and Nelson F. Fontoura. Age, growth and reproductive biology of threatened westslope cutthroat trout Oncorhynchus clarkii lewisi inhabiting small mountain streams. Journal of Fish Biology, 93(5):874–886, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russ:2018:DSR

[2748] Garry R. Russ, Cody S. Payne, Brock J. Bergseth, Justin R. Rizzari, Rene A. Abesamis, and Angel C. Alcala. Decadal-scale response of detritivorous surgeonfishes (family Acanthuridae) to no-take marine reserve protection and changes in benthic habitat. Journal of Fish Biology, 93 (5):887–900, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slobodian:2018:DNP

[2749] Veronica Slobodian and Murilo N. L. Pastana. Description of a new Pimelodella (Siluriformes: Heptapteridae) species with a discussion on the upper pectoral girdle homology of Siluriformes. Journal of Fish Biology, 93(5):901–916, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:2018:FHR

[2750] Peter J. Allen, Wes Baumgartner, Erin Brinkman, Robert J. DeVries, Heather A. Stewart, Daniel L. Aboagye, Shane W. Ramee, Michael A. Ciaramella, Charlie M. Culpepper III, and Lora Petrie-Hanson. Fin healing and regeneration in sturgeon. Journal of Fish Biology, 93(5):917–930, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jenkins:2018:DTS

[2751] Gregory P. Jenkins, Julia A. Kent, Ryan J. Woodland, Fiona Warry, Stephen E. Swearer, and Perran L. M. Cook. Delayed timing of successful spawning of an estuarine dependent fish, black bream Acanthopagrus butcheri. Journal of Fish Biology, 93(5):931–941, November 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitaker:2018:VDM

[2752] Justine M. Whitaker, Amy B. Welsh, Darryl W. Hondorp, James C. Boase, George T. Merovich, Stuart Welsh, and Charles Krueger. Variation in DNA methylation is associated with migratory phenotypes of lake sturgeon Acipenser fulvescens in the St. Clair River, MI, USA. Journal of Fish Biology, 93(5):942–951, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poletto:2018:AMS

[2753] Jamilynn B. Poletto, Benjamin Martin, Eric Danner, Sarah E. Baird, Dennis E. Cocherell, Natnael Hamda, Joseph J. Cech, Jr., and Nann A. Fangue. Assessment of multiple stressors on the growth of larval green sturgeon Acipenser medirostris: implications for recruitment of early lifehistory stages. Journal of Fish Biology, 93(5):952–960, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abecia:2018:IMR

[2754] Janine E. D. Abecia, Osmar J. Luiz, and Alison J. King. Intraspecific morphological and reproductive trait variation in mouth almighty Glossamia aprion (Apogonidae) across different flow environments. Journal of Fish Biology, 93(5):961–971, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sherwood:2018:PRR

[2755] Joshua L. Sherwood, Andrew J. Stites, Michael J. Dreslik, and Jeremy S. Tiemann. Predicting the range of a regionally threatened, benthic fish using species distribution models and field surveys. Journal of Fish Biology, 93(5):972–977, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Losee:2018:SAG

[2756] James P. Losee, Andrew M. Claiborne, Phillip E. Dionne, Hannah S. Faulkner, and Todd R. Seamons. Size, age, growth and site fidelity of anadromous cutthroat trout Oncorhynchus clarkii clarkii in the Salish Sea. Journal of Fish Biology, 93(5):978–987, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soeth:2018:FRP

[2757] Marcelo Soeth, Johnatas Adelir-Alves, Robin Loose, Felippe A. Daros, and Henry L. Spach. First record of Pomacanthus maculosus (Perciformes, Pomacanthidae) in the south-western Atlantic Ocean. Journal of Fish Biology, 93(5):988–991, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eduardo:2018:FRI

[2758] Leandro N. Eduardo, Bárbara T. Villarins, Flávia Lucena-Frédou, Thierry Frédou, Alex S. Lira, Arnaud Bertrand, and Michael M. Mincarone. First record of the intermediate scabbardfish Aphanopus intermedius (Scombriformes: Trichiuridae) in the western South Atlantic Ocean. Journal of Fish Biology, 93(5):992–995, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christoffersen:2018:BHS

[2759] Mads Christoffersen, Jon C. Svendsen, Jesper A. Kuhn, Anders Nielsen, Anastasija Martjanova, and Josianne G. Støttrup. Benthic habitat selection in juvenile European eel Anguilla anguilla: implications for coastal habitat management and restoration. Journal of Fish Biology, 93(5):

996–999, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baird:2018:NCS

[2760] Sarah E. Baird, Anna E. Steel, Dennis E. Cocherell, Joseph J. Cech, Jr., and Nann A. Fangue. Native Chinook salmon Oncorhynchus tshawytscha and non-native brook trout Salvelinus fontinalis prefer similar water temperatures. Journal of Fish Biology, 93(5):1000–1004, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McLoone:2018:CPE

[2761] Paul McLoone, Samuel Shephard, Karen Delanty, Kieran Rocks, Rory Feeney, and Fiona Kelly. Coexistence of pike Esox lucius and brown trout Salmo trutta in Irish lakes. Journal of Fish Biology, 93(5):1005–1011, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colaco-Fernandes:2018:AHP

[2762] Talles Colaço-Fernandes and Sidineia Amadio. Abnormal hermaphroditism in Pellona flavipinnis (Clupeiformes: Pristigasteridae). Journal of Fish Biology, 93(5):1012–1015, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2018:EIT

[2763] Bror Jonsson and Nina Jonsson. Egg incubation temperature affects the timing of the Atlantic salmon Salmo salar homing migration. Journal of Fish Biology, 93(5):1016–1020, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burcea:2018:GEI

[2764] Alexandru Burcea, Gina-Oana Popa, Iulia E. Florescu (Gune), Sami Gharbia, Gisela Gaina, Andreea Dudu, Marilena Maereanu, Anca Hermenean, Sergiu E. Georgescu, and Marieta Costache. Gene expression involved in the sexual development of Best Beluga hybrid sturgeons. Journal of Fish Biology, 93(5):1021–1026, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2018:CFB

[2765] Ian J. Winfield. Cleaner fish biology and aquaculture applications. Journal of Fish Biology, 93(5):1027, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everson:2018:ASK

[2766] Inigo Everson. The Antarctic silverfish: a keystone species in a changing ecosystem. Journal of Fish Biology, 93(5):1028, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2018:ITO

[2767] Ian J. Winfield. Impacts of tannery operations on guppy, Poecilia reticulata. Journal of Fish Biology, 93(5):1029, November 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2018:IIm

[2768] Anonymous. Issue information. Journal of Fish Biology, 93(6):1031– 1032, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wosnick:2018:UPM

[2769] Natascha Wosnick, Kye R. Adams, and Carolina A. Freire. Ultrasonography as a promising methodology to indicate captured-induced abortion in viviparous elasmobranchs. Journal of Fish Biology, 93(6):1033–1037, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corsso:2018:CAB

[2770] Jessica T. Corsso, Otto B. F. Gadig, Rodrigo R. P. Barreto, and Fabio S. Motta. Condition analysis of the Brazilian sharpnose shark Rhizoprionodon lalandii: evidence of maternal investment for initial post-natal life. Journal of Fish Biology, 93(6):1038–1045, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2018:CMI

[2771] Laith A. Jawad. A comparative morphological investigation of otoliths of six parrotfish species (Scaridae) from the Solomon Islands. Journal of Fish Biology, 93(6):1046–1058, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giora:2018:AHA

[2772] Júlia Giora and Tiago P. Carvalho. Anatomy and homology of the accessory electric organs of the toothless knifefishes (Rhamphichthyoidea: Gymnotiformes). Journal of Fish Biology, 93(6):1059–1068, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chance:2018:ERA

[2773] Rachel J. Chance, Gary A. Cameron, Mark Fordyce, Patricia Noguera, Tiehui Wang, Catherine Collins, Christopher J. Secombes, and Bertrand Collet. Effects of repeated anaesthesia on gill and general health of Atlantic salmon, Salmo salar. Journal of Fish Biology, 93(6):1069–1081, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qin:2018:DEH

[2774] Qin B. Qin, Qi W. Liu, Yu W. Zhou, Chong Q. Wang, Huan Qin, Chun Zhao, and Shao J. Liu. Differential expression of HPG-axis genes in autotetraploids derived from red crucian carp Carassius auratus red var., [female sign] × blunt snout bream Megalobrama amblycephala, [male sign]. Journal of Fish Biology, 93(6):1082–1089, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temperoni:2018:ENG

[2775] Brenda Temperoni, Agueda E. Massa, Carla Derisio, Patricia Martos, Carla Berghoff, and María D. Viñas. Effect of nursery ground variability on condition of age 0+ year Merluccius hubbsi. Journal of Fish Biology, 93(6):1090–1101, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vincenzi:2018:SEH

[2776] Simone Vincenzi, Dušan Jesenšek, John C. Garza, and Alain J. Crivelli. Stronger effects of heterozygosity on survival in harsher environments. Journal of Fish Biology, 93(6):1102–1106, December 2018. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guillerault:2018:CRG

[2777] Nicolas Guillerault, Géraldine Loot, Simon Blanchet, and Frederic Santoul. Catch-related and genetic outcome of adult northern pike Esox lucius stocking in a large river system. Journal of Fish Biology, 93(6): 1107–1112, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2018:CGG

[2778] LiB. He, Lv Xiong, AiD. Zhang, YongM. Li, Rong Huang, LanJ. Liao, ZuoY. Zhu, and YaP. Wang. Changes in gene and genotype frequencies during the development of the grass carp Ctenopharyngodon idella. Journal of Fish Biology, 93(6):1113–1120, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendonca:2018:SAD

[2779] Sibele A. Mendonça, Bruno C. L. Macena, André S. Afonso, and Fabio H. V. Hazin. Seasonal aggregation and diel activity by the sicklefin devil ray Mobula tarapacana off a small, equatorial outcrop of the mid-Atlantic ridge. Journal of Fish Biology, 93(6):1121–1129, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shink:2018:GVP

[2780] Katie G. Shink, Trent M. Sutton, James M. Murphy, and J. Andrés López. Genetic variation and population structure among larval Lethenteron spp. within the Yukon River drainage, Alaska. Journal of Fish Biology, 93(6):1130–1140, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neumann:2018:LDB

[2781] Erika Neumann, Maria Carmo F. Paes, José Mário R. Mendes, Francisco Manoel S. Braga, and Laura Satiko O. Nakaghi. Larval development of Brycon amazonicus (Teleostei, Bryconidae) with a focus on locomotory, respiratory and feeding structures. Journal of Fish Biology, 93(6):1141– 1150, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barreto:2018:IAR

[2782] Silvia B. Barreto, André T. Silva, Henrique Batalha-Filho, Paulo Roberto A. M. Affonso, and Angela M. Zanata. Integrative approach reveals a new species of Nematocharax (Teleostei: Characidae). Journal of Fish Biology, 93(6):1151–1162, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2018:EUP

[2783] Brenda Anderson, Carolyn Belcher, JoAnn Slack, and Jim Gelsleichter. Evaluation of the use of portable ultrasonography to determine pregnancy status and fecundity in bonnethead shark Sphyrna tiburo. Journal of Fish Biology, 93(6):1163–1170, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ralliere:2018:FIC

[2784] Cécile Rallière, Adele Branthonne, and Pierre-Yves Rescan. Formation of intramuscular connective tissue network in fish: first insight from the rainbow trout (Oncorhynchus mykiss ). Journal of Fish Biology, 93(6): 1171–1177, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delomas:2018:EHS

[2785] Thomas A. Delomas and Konrad Dabrowski. Effects of homozygosity on sex determination in zebrafish Danio rerio. Journal of Fish Biology, 93(6):1178–1187, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjornson:2018:BCR

[2786] Forrest Bjornson and Gary W. Anderson. Body condition, rather than size, predicts risk-taking and resource holding potential in hatchery reared juvenile lake sturgeon Acipenser fulvescens. Journal of Fish Biology, 93(6):1188–1196, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robles:2018:ECT

[2787] Laurence E. Robles, Patrick C. Cabaitan, and Maria Eleanor B. Aurellado. Effects of competition on the territorial behaviour of a farmer damselfish, Plectroglyphidodon lacrymatus (Perciformes: Pomacentridae). Journal of Fish Biology, 93(6):1197–1206, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gao:2018:MCI

[2788] Feng-Ying Gao, Dong Zhang, Mai-Xin Lu, Jian-Meng Cao, Zhi-Gang Liu, Xiao-Li Ke, Miao Wang, De-Feng Zhang, and Meng-Meng Yi. MHC class IIA polymorphisms and their association with resistance– susceptibility to Streptococcus agalactiae in Nile tilapia, Oreochromis niloticus. Journal of Fish Biology, 93(6):1207–1215, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vernygora:2018:DAS

[2789] Oksana V. Vernygora, Corey S. Davis, Alison M. Murray, and Felix A. H. Sperling. Delimitation of Alosa species (Teleostei: Clupeiformes) from the Sea of Azov: integrating morphological and molecular approaches. Journal of Fish Biology, 93(6):1216–1228, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bovcon:2018:FRP

[2790] Nelson D. Bovcon, Pablo D. Cochia, Ximena Navoa, Pablo Ledesma, Guillermo M. Caille, and Claudio R. M. Baigun. First report on a pupping area of the tope shark Galeorhinus galeus (Carcharhiniformes, Triakidae) in the south-west Atlantic. Journal of Fish Biology, 93(6): 1229–1232, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ashida:2018:ESH

[2791] Hiroshi Ashida, Shinpei Ohashi, Masachika Masujima, Masaaki Nakano, and Tamaki Shimose. Evidence of spawning by hermaphroditic skipjack tuna Katsuwonus pelamis. Journal of Fish Biology, 93(6):1233–1237, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wanzenbock:2018:SFF

[2792] Josef Wanzenböck. Success factors for fish larval production. Journal of Fish Biology, 93(6):1238, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2018:ACL

[2793] Ian J. Winfield. Arctic charr in the lochs of Scotland: an assessment of distribution and status. Journal of Fish Biology, 93(6):1239, December 2018. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIa

[2794] Anonymous. Issue information. Journal of Fish Biology, 94(1):1–2, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacoby:2019:ESR

[2795] David M. P. Jacoby. An eye in the sky reveals the collective dynamics of free-ranging shark aggregations. Journal of Fish Biology, 94(1): 3, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lujan:2019:AYS

[2796] Nathan Lujan. After 175 years, a species proposed by Darwin and Jenyns is tested. Journal of Fish Biology, 94(1):4, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrissey-McCaffrey:2019:NNS

[2797] Emma Morrissey-McCaffrey, Samuel Shephard, Fiona L. Kelly, and Mary Kelly-Quinn. Non-native species and lake warming negatively affect Arctic char Salvelinus alpinus abundance; deep thermal refugia facilitate coexistence. Journal of Fish Biology, 94(1):5–16, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quimpo:2019:SVR

[2798] Timothy J. R. Quimpo, Patrick C. Cabaitan, Ronald D. D. Olavides, Edwin E. Dumalagan, Jr., Jeffrey Munar, and Fernando P. Siringan. Spatial variability in reef-fish assemblages in shallow and upper mesophotic coral ecosystems in the Philippines. Journal of Fish Biology, 94(1):17– 28, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedberg:2019:FLD

[2799] Per Hedberg, Fanny F. Rybak, Martin Gullström, Narriman S. Jiddawi, and Monika Winder. Fish larvae distribution among different habitats in coastal East Africa. Journal of Fish Biology, 94(1):29–39, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2019:SSD

[2800] Brianne Kelly, Hugues P. Benoît, Gerald Chaput, Ross A. Jones, and Michael Power. Spawning-strategy-dependent diets in two North American populations of Atlantic salmon Salmo salar. Journal of Fish Biology, 94(1):40–52, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morat:2019:DUC

[2801] Fabien Morat, Marine J. Briand, Christophe Pécheyran, and Yves Letourneur. Differential uses of coral reef habitats by a poorly-known cryptic fish predator. Journal of Fish Biology, 94(1):53–61, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulsen:2019:NOO

[2802] Jan Y. Poulsen. New observations and ontogenetic transformation of photogenic tissues in the tubeshoulder Sagamichthys schnakenbecki (Platytroctidae, Alepocephaliformes). Journal of Fish Biology, 94(1): 62–76, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mangel:2019:EPS

[2803] Jeffrey C. Mangel, Mariela Pajuelo, Andrea Pasara-Polack, German Vela, Eduardo Segura-Cobeña, and Joanna Alfaro-Shigueto. The effect of Peruvian small-scale fisheries on sunfishes (Molidae). Journal of Fish Biology, 94(1):77–85, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2019:EEE

[2804] Carole J. Lee, Gregory C. Paull, and Charles R. Tyler. Effects of environmental enrichment on survivorship, growth, sex ratio and behaviour in laboratory maintained zebrafish Danio rerio. Journal of Fish Biology, 94(1):86–95, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moyes:2019:CDS

[2805] Faye Moyes and Anne E. Magurran. Change in the dominance structure of two marine-fish assemblages over three decades. Journal of Fish Biology, 94(1):96–102, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reizenberg:2019:VTT

[2806] Jody-Lee Reizenberg, Lesley E. Bloy, Olaf L. F. Weyl, Jeremy M. Shelton, and Helen F. Dallas. Variation in thermal tolerances of native freshwater fishes in South Africa’s Cape Fold Ecoregion: examining the east– west gradient in species’ sensitivity to climate warming. Journal of Fish Biology, 94(1):103–112, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macnaughton:2019:SMR

[2807] Camille J. Macnaughton, Colin Kovachik, and Eva C. Enders. Standard metabolic rate models for Carmine Shiner (Notropis percobromus) and Common Shiner (Luxilus cornutus) across different temperature regimes. Journal of Fish Biology, 94(1):113–121, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlucci:2019:CHA

[2808] Roberto Carlucci, Donatella Mentino, Daniela Semeraro, Pasquale Ricci, Letizia Sion, and Giovanni Scillitani. Comparative histochemical analysis of intestinal glycoconjugates in the blunthead pufferfish Sphoeroides pachygaster and grey triggerfish Balistes capriscus (Teleostei: Tetraodontiformes). Journal of Fish Biology, 94(1):122–131, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2019:DSS

[2809] Liangming Chen, Rong Huang, Denghui Zhu, Cheng Yang, Libo He, Yongming Li, Lanjie Liao, Zuoyan Zhu, and Yaping Wang. Deep sequencing of small RNAs from 11 tissues of grass carp Ctenopharyngodon idella and discovery of sex-related microRNAs. Journal of Fish Biology, 94(1):132–141, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maxfield:2019:PSC

[2810] Jessica M. Maxfield and Kathleen S. Cole. Patterns of structural change in gonads of the divine dwarfgoby Eviota epiphanes as they sexually transition. Journal of Fish Biology, 94(1):142–153, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnekleiv:2019:DGD

[2811] Jo V. Arnekleiv, Jan G. Davidsen, Timothy F. Sheehan, Sarah J. Lehnert, Ian R. Bradbury, Lars Rønning, Aslak D. Sjursen, Gaute Kjærstad, Barbara A. Lubinski, and Kjell J. Nilssen. Demographic and genetic description of Greenland’s only indigenous Atlantic salmon Salmo salar population. Journal of Fish Biology, 94(1):154–164, January 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:2019:WSR

[2812] Darren A. Whitehead, Edgar E. Becerril-García, David Petatán-Ramírez, Abraham Vázquez-Haikin, Rogelio González-Armas, and Felipe Galván-Magaña. Whale shark Rhincodon typus strandings in the Gulf of California, Mexico. Journal of Fish Biology, 94(1):165–167, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breves:2019:PCB

[2813] Jason P. Breves. Prolactin controls branchial clcn2c but not atp1a1a.2 in zebrafish Danio rerio. Journal of Fish Biology, 94(1):168–172, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smukall:2019:NAN

[2814] Matthew J. Smukall, Steven T. Kessel, Bryan R. Franks, Kevin A. Feldheim, Tristan L. Guttridge, and Samuel H. Gruber. No apparent negative tagging effects after 13 years at liberty for lemon shark, Negaprion brevirostris implanted with acoustic transmitter. Journal of Fish Biology, 94 (1):173–177, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2019:ABF

[2815] Morten Bo Søndergaard Svendsen, Jacob L. Johansen, Peter G. Bushnell, Peter V. Skov, Tommy Norin, Paolo Domenici, John F. Steffensen, and Augusto Abe. Are all bony fishes oxygen regulators? Evidence for oxygen regulation in a putative oxygen conformer, the swamp eel Synbranchus marmoratus. Journal of Fish Biology, 94(1):178–182, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanson:2019:PAD

[2816] Nora N. Hanson, Gordon W. Smith, Stuart J. Middlemas, and Christopher D. Todd. Precision and accuracy of dahl-lea back-calculated smolt lengths from adult scales of Atlantic salmon Salmo salar. Journal of Fish Biology, 94(1):183–186, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anju:2019:DLI

[2817] Chakkalathara Venu Anju, Neelesh Dahanukar, Arya Sidharthan, Kutty Ranjeet, and Rajeev Raghavan. Demographics of Lagocephalus inermis in the Arabian Sea unveils complex conservation challenges. Journal of Fish Biology, 94(1):187–190, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Austin:2019:SEP

[2818] Catherine S. Austin, Timothy E. Essington, and Thomas P. Quinn. Spawning and emergence phenology of bull trout Salvelinus confluentus under differing thermal regimes. Journal of Fish Biology, 94(1):191–195, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janse:2019:CWE

[2819] Max Janse and Henk A. Luten. Contraception within an elasmobranch captive breeding programme: a case report on the eagle ray Aetobatus ocellatus using gonadotrophin-releasing hormone (GnRH) agonist Deslorelin (Suprelorin). Journal of Fish Biology, 94(1):196–199, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haque:2019:FCR

[2820] Alifa B. Haque and Sudipta A. Das. First confirmed record of the Critically Endangered green sawfish Pristis zijsron from Bangladeshi waters. Journal of Fish Biology, 94(1):200–203, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2019:EFB

[2821] Andy D. Nunn. Essential fish biology: diversity, structure and function. Journal of Fish Biology, 94(1):204, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proudlove:2019:BRA

[2822] Graham Proudlove. Book review: The Astyanax Caves of Mexico: Cavefishes of Tamaulipas, San Luis Potosí, and Guerrero. Journal of Fish Biology, 94(1):205, January 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIb

[2823] Anonymous. Issue information. Journal of Fish Biology, 94(2):207–208, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lujan:2019:SLL

[2824] Nathan K. Lujan. Shining light on larval fish diversity on reefs. Journal of Fish Biology, 94(2):209, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grande:2019:DSS

[2825] Henrique Grande, Sergio de M. Rezende, Thiony E. Simon, Fabiana C. Félix-Hackradt, José A. García-Charton, Mauro Maida, Ana L. B. Gaspar, Ronaldo B. Francini-Filho, Thierry Fredou, and Beatrice P. Ferreira. Diversity of settlement-stage reef fishes captured by light-trap in a tropical south-west Atlantic Ocean coastal reef system. Journal of Fish Biology, 94(2):210–222, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2019:SDC

[2826] Qian Ma, Shu F. Liu, Xin X. Wang, Zi L. Xiang, and Zhi M. Zhuang. Skeletal development of the chondrocranium in the tongue sole Cynoglossus semilaevis (Pleuronectiformes: Cynoglossidae). Journal of Fish Biology, 94(2):223–230, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaneko:2019:MRC

[2827] Sanshiro Kaneko, Reiji Masuda, and Yoh Yamashita. Memory retention capacity using two different training methods, appetitive and aversive learning, in juvenile red sea bream Chrysophrys major. Journal of Fish Biology, 94(2):231–240, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doring:2019:SEO

[2828] Julian Döring, Carola Wagner, Maik Tiedemann, Patrice Brehmer, and Werner Ekau. Spawning energetics and otolith microchemistry provide insights into the stock structure of bonga shad Ethmalosa fimbriata. Journal of Fish Biology, 94(2):241–250, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Song:2019:TBT

[2829] Wei Song, Lingzhi Li, Hongliang Huang, Keji Jiang, Fengying Zhang, Luming Wang, Ming Zhao, and Lingbo Ma. Tissue-based transcriptomics of Chionodraco hamatus: Sequencing, de novo assembly, annotation and marker discovery. Journal of Fish Biology, 94(2):251–260, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eom:2019:VMP

[2830] Junho Eom and Chris M. Wood. The ventilation mechanism of the Pacific hagfish Eptatretus stoutii. Journal of Fish Biology, 94(2):261– 276, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowell:2019:RBC

[2831] Timothy J. Rowell, Octavio Aburto-Oropeza, Juan J. Cota-Nieto, Mark A. Steele, and Brad E. Erisman. Reproductive behaviour and concurrent sound production of Gulf grouper Mycteroperca jordani (Epinephelidae) at a spawning aggregation site. Journal of Fish Biology, 94(2):277–296, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahmad:2019:NCC

[2832] Raahyma Ahmad, Helene Paradis, Danny Boyce, James McDonald, and Robert L. Gendron. Novel characteristics of the cultured lumpfish Cyclopterus lumpus eye during post-hatch larval and juvenile developmental stages. Journal of Fish Biology, 94(2):297–312, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gauthier:2019:CME

[2833] Arnault R. G. Gauthier, Darryl L. Whitehead, Ian R. Tibbetts, and Michael B. Bennett. Comparative morphology of the electrosensory system of the epaulette shark Hemiscyllium ocellatum and brown-banded bamboo shark Chiloscyllium punctatum. Journal of Fish Biology, 94 (2):313–319, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almeida:2019:LDM

[2834] Mara C. Almeida, Hugo J. Message, Paulo V. Sanches, Dirceu Baumgartner, Robie A. Bombardelli, Gilmar Baumgartner, and Langeani Francisco. Larval development of mandi-pintado Pimelodus britskii (Siluriformes: Pimelodidae). Journal of Fish Biology, 94(2):320–324, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2019:GPS

[2835] Shoichiro Yamamoto. Genetic population structure of Japanese river sculpin Cottus pollux (Cottidae) large-egg type, inferred from mitochondrial DNA sequences. Journal of Fish Biology, 94(2):325–329, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2019:OSS

[2836] Laith A. Jawad, Ahmed M. H. Al-Dirawi, Hayder I. Al-Hilali, and Uday T. S. Al-Asadi. Observations of stranded and swimming whale sharks Rhincodon typus in Khor Al-Zubair, NW Arabian Gulf and Shatt Al-Arab Estuary, Iraq. Journal of Fish Biology, 94(2):330–334, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattei:2019:BSA

[2837] Xavier Mattei, Bernard Marchand, and Yann Quilichini. A biflagellate spermatozoon in the African bonytongue Heterotis niloticus (Teleostei, Osteoglossidae). Journal of Fish Biology, 94(2):335–338, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vecchioni:2019:DBM

[2838] Luca Vecchioni, Federico Marrone, Eya Belaiba, Francesco Tiralongo, Lilia Bahri-Sfar, and Marco Arculeo. The DNA barcoding of Mediterranean combtooth blennies suggests the paraphyly of some taxa (Perciformes, Blenniidae). Journal of Fish Biology, 94(2):339–344, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2019:RLS

[2839] Wilson J. E. M. Costa, José L. O. Mattos, and Pedro F. Amorim. Rediscovery of Leptopanchax splendens (Cyprinodontiformes: Aplocheilidae): a seasonal killifish from the Atlantic Forest of south-eastern Brazil that was recently considered extinct. Journal of Fish Biology, 94(2):345–347, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:Ea

[2840] Anonymous. Erratum. Journal of Fish Biology, 94(2):348, February 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIc

[2841] Anonymous. Issue information. Journal of Fish Biology, 94(3):349–350, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubbard:2019:CSF

[2842] Peter Hubbard. Changing sex: it’s a fishy thing. Journal of Fish Biology, 94(3):351, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faustino-Fuster:2019:TNS

[2843] Dario R. Faustino-Fuster, Flávio A. Bockmann, and Luiz R. Malabarba. Two new species of Heptapterus (Siluriformes: Heptapteridae) from the Uruguay River basin, Brazil. Journal of Fish Biology, 94(3):352–373, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulson:2019:LHC

[2844] Peter G. Coulson. The life-history of Cheilodactylus rubrolabiatus from south-western Australia and comparison of biological characteristics of the Cheilodactylidae and Latridae: support for an amalgamation of families. Journal of Fish Biology, 94(3):374–390, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balde:2019:VKB

[2845] Bocar S. Baldé, Fambaye N. Sow, Kamarel Ba, Werner Ekau, Patrice Brehmer, Justin Kantoussan, Massal Fall, and Malick Diouf. Variability of key biological parameters of round sardinella Sardinella aurita and the effects of environmental changes. Journal of Fish Biology, 94(3):391– 401, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibarra-Zatarain:2019:RSM

[2846] Zohar Ibarra-Zatarain, Katia Parati, Silvia Cenadelli, and Neil Duncan. Reproductive success of a marine teleost was correlated with proactive and reactive stress-coping styles. Journal of Fish Biology, 94(3):402– 413, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2019:SCM

[2847] Jiaxing Chen, Ling Xiao, Cheng Peng, Zhifeng Ye, Dengdong Wang, Yuqing Yang, Haifa Zhang, Mi Zhao, Shuisheng Li, Haoran Lin, and Yong Zhang. Socially controlled male-to-female sex reversal in the protogynous orange-spotted grouper, Epinephelus coioides. Journal of Fish Biology, 94(3):414–421, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reemeyer:2019:EPI

[2848] Jessica E. Reemeyer, Jasmine C. Harris, Ariel M. Hernandez, and Bernard B. Rees. Effects of passive integrated transponder tagging on cortisol release, aerobic metabolism and growth of the Gulf killifish Fundulus grandis. Journal of Fish Biology, 94(3):422–433, March 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2019:SMM

[2849] Jessica S. Miller, Aneesh P. H. Bose, John L. Fitzpatrick, and Sigal Balshine. Sperm maturation and male tactic-specific differences in ejaculates in the plainfin midshipman fish Porichthys notatus. Journal of Fish Biology, 94(3):434–445, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roufidou:2019:OFT

[2850] Chrysoula Roufidou and Bertil Borg. Ovarian fluid in the three-spined stickleback Gasterosteus aculeatus: effects of egg overripening and sex steroid treatment. Journal of Fish Biology, 94(3):446–457, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2019:OCN

[2851] Davut Turan, Cüneyt Kaya, Gökhan Kalayci, Esra Bayçelebi, and İsmail Aksu. Oxynoemacheilus cemali, a new species of stone loach (Teleostei: Nemacheilidae) from the Çoruh River drainage, Turkey. Journal of Fish Biology, 94(3):458–468, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2019:BSC

[2852] Kelsey L. Martin, Dan C. Abel, Derek P. Crane, Neil Hammerschlag, and Erin J. Burge. Blacktip shark Carcharhinus limbatus presence at fishing piers in South Carolina: association and environmental drivers. Journal of Fish Biology, 94(3):469–480, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gianeti:2019:ASM

[2853] Michel D. Gianeti, Francisco M. Santana, Leandro Yokota, Jonas E. Vasconcelos, June F. Dias, and Rosangela P. Lessa. Age structure and multi-model growth estimation of longnose stingray Hypanus guttatus (Dasyatidae: Myliobatoidei) from north-east Brazil. Journal of Fish Biology, 94(3):481–488, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Branconi:2019:CED

[2854] Rebecca Branconi, Marian Y. L. Wong, and Peter M. Buston. Comparison of efficiency of direct observations by scuba diver and indirect observations via video camera for measuring reef-fish behaviour. Journal of Fish Biology, 94(3):489–497, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vorsatz:2019:OBS

[2855] Lyle D. Vorsatz, Carl D. van der Lingen, and Mark J. Gibbons. Observations on the biology and seasonal variation in feeding of the east coast round herring Etrumeus wongratanai (Clupeiformes), off Scottburgh, KwaZulu-Natal, South Africa. Journal of Fish Biology, 94(3):498–511, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domingues:2019:UMN

[2856] Rodrigo R. Domingues, Domingos Garrone-Neto, Alexandre W. S. Hilsdorf, and Otto B. F. Gadig. Use of mucus as a non-invasive sampling method for DNA barcoding of stingrays and skates (batoid elasmobranchs). Journal of Fish Biology, 94(3):512–516, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mancuso:2019:FRM

[2857] Monique Mancuso, Serena Savoca, and Teresa Bottari. First record of microplastics ingestion by European hake Merluccius merluccius from the Tyrrhenian Sicilian coast (Central Mediterranean Sea). Journal of Fish Biology, 94(3):517–519, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stuckert:2019:EAC

[2858] Adam M. M. Stuckert, Sara Drury, Christopher M. Anderson, Tyler B. T. Bowling, and Jeffrey S. Mckinnon. Evolution and assessment of colour patterns in stream-resident and anadromous male threespine stickleback Gasterosteus aculeatus from three regions. Journal of Fish Biology, 94 (3):520–525, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guallart:2019:NRS

[2859] Javier Guallart, Gabriel Morey, and Àlex Bartolí. New record of a sharpnose sevengill shark Heptranchias perlo (Elasmobranchii, Hexanchidae) from the Balearic Sea, western Mediterranean Sea. Journal of Fish Biology, 94(3):526–531, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tandler:2019:DCE

[2860] Talia Tandler, Emma Gellman, Dayna De La Cruz, and David J. Ellerby. Drag coefficient estimates from coasting bluegill sunfish Lepomis macrochirus. Journal of Fish Biology, 94(3):532–534, March 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IId

[2861] Anonymous. Issue information. Journal of Fish Biology, 94(4):537–538, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2019:MCE

[2862] David J. McKenzie. Measuring cortisol to evaluate stress in fishes. Journal of Fish Biology, 94(4):539, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadoul:2019:MCM

[2863] Bastien Sadoul and Benjamin Geffroy. Measuring cortisol, the major stress hormone in fishes. Journal of Fish Biology, 94(4):540–555, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloman:2019:ECF

[2864] Katherine A. Sloman, Ian A. Bouyoucos, Edward J. Brooks, and Lynne U. Sneddon. Ethical considerations in fish research. Journal of Fish Biology, 94(4):556–577, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bemis:2019:LHD

[2865] Katherine E. Bemis, James C. Tyler, and Dahiana Arcila. Life history, distribution and molecular phylogenetics of the upward-mouth spikefish Atrophacanthus japonicus (Teleostei: Tetraodontiformes: Triacanthodidae). Journal of Fish Biology, 94(4):578–584, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skinner:2019:CSM

[2866] Christina Skinner, Steven P. Newman, Stephen Box, Andrzej Narozanski, and Nicholas V. C. Polunin. Chronic spearfishing may indirectly affect reef health through reductions in parrotfish bite rates. Journal of Fish Biology, 94(4):585–594, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadauskas-Henrique:2019:DDO

[2867] Helen Sadauskas-Henrique, Chris M. Wood, Luciana R. Souza-Bastos, Rafael M. Duarte, Donald S. Smith, and Adalberto L. Val. Does dissolved organic carbon from Amazon black water (Brazil) help a native species, the tambaqui Colossoma macropomum to maintain ionic homeostasis in acidic water? Journal of Fish Biology, 94(4):595–605, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gkagkavouzis:2019:GPS

[2868] Konstantinos Gkagkavouzis, Nikoleta Karaiskou, Theodora Katopodi, Ioannis Leonardos, Theodore J. Abatzopoulos, and Alexander Triantafyllidis. The genetic population structure and temporal genetic stability of gilthead sea bream Sparus aurata populations in the Aegean and Ionian Seas, using microsatellite DNA markers. Journal of Fish Biology, 94(4):606–613, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2019:IFR

[2869] Øivind Andersen, Vera Vieira, Jens-Erik Dessen, and Ian A. Johnston. Influence of feed ration size on somatic and muscle growth in landlocked dwarf and farmed Atlantic salmon Salmo salar. Journal of Fish Biology, 94(4):614–620, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quattrini:2019:ASV

[2870] Andrea M. Quattrini, Jennifer McClain-Counts, Stephen J. Artabane, Adela Roa-Varón, Tara C. McIver, Michael Rhode, and Steve W. Ross. Assemblage structure, vertical distributions and stable-isotope compositions of anguilliform leptocephali in the Gulf of Mexico. Journal of Fish Biology, 94(4):621–647, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deOliveira:2019:DTP

[2871] Maria I. B. de Oliveira, Lorena V. de Matos, Lídia A. da Silva, Edsandra C. Chagas, Grazyelle S. da Silva, and Ana L. S. Gomes. The digestive tube of Piaractus brachypomus: gross morphology, histology/ histochemistry of the mucosal layer and the effects of parasitism by Neoechinorhynchus sp. Journal of Fish Biology, 94(4):648–659, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsutsui:2019:MGG

[2872] Shigeyuki Tsutsui, Tatsuki Yoshinaga, Shun Watanabe, Katsumi Tsukamoto, and Osamu Nakamura. Mucosal galectin genes in all freshwater eels of the genus Anguilla. Journal of Fish Biology, 94(4):660–670, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Felix-Lopez:2019:PFP

[2873] Daniela G. Félix-López, Nataly Bolaño-Martinez, Pindaro Díaz-Jaimes, Erick C. Oñate-González, Jorge S. Ramírez-Pérez, Emiliano García-Rodríguez, David Corro-Espinosa, Jesus E. Osuna-Soto, and Nancy C. Saavedra-Sotelo. Possible female philopatry of the smooth hammerhead shark Sphyrna zygaena revealed by genetic structure patterns. Journal of Fish Biology, 94(4):671–679, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curnick:2019:EDR

[2874] David J. Curnick, Aaron B. Carlisle, Matthew J. Gollock, Robert J. Schallert, and Nigel E. Hussey. Evidence for dynamic resource partitioning between two sympatric reef shark species within the British Indian Ocean Territory. Journal of Fish Biology, 94(4):680–685, April 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIe

[2875] Anonymous. Issue information. Journal of Fish Biology, 94(5):687–688, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2019:NW

[2876] David J. McKenzie. Noisy waters. Journal of Fish Biology, 94(5):691, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Popper:2019:OFB

[2877] Arthur N. Popper and Anthony D. Hawkins. An overview of fish bioacoustics and the impacts of anthropogenic sounds on fishes. Journal of Fish Biology, 94(5):692–713, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2019:GSR

[2878] Jianhua Wang, Jia Cheng, Yulong Li, Huiling Yan, Ping Wu, Xin Zhu, Li Liu, Lin Chen, Wuying Chu, and Jianshe Zhang. Gene structure, recombinant expression and function characterization of Siniperca chuatsi Fsrp-3. Journal of Fish Biology, 94(5):714–724, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreira:2019:GDP

[2879] Cláudia Moreira, Alberto T. Correia, Paulo Vaz-Pires, and Elsa Froufe. Genetic diversity and population structure of the blue jack mackerel Trachurus picturatus across its western distribution. Journal of Fish Biology, 94(5):725–731, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burggren:2019:DCO

[2880] Warren W. Burggren, Jose F. Mendez-Sanchez, Gil Martínez Bautista, Emyr Peña, Rafael Martínez García, and Carlos A. Alvarez González. Developmental changes in oxygen consumption and hypoxia tolerance in the heat and hypoxia-adapted tabasco line of the Nile tilapia Oreochromis niloticus, with a survey of the metabolic literature for the genus Oreochromis. Journal of Fish Biology, 94(5):732–744, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schwinn:2019:ALD

[2881] Michael Schwinn, Henrik Baktoft, Kim Aarestrup, and Anders Koed. Artificial lakes delay the migration of brown trout Salmo trutta smolts: a comparison of migratory behaviour in a stream and through an artificial lake. Journal of Fish Biology, 94(5):745–751, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2019:VMH

[2882] Takaomi Arai and Naoko Chino. Variations in the migratory history of the tropical catadromous eels Anguilla bicolor bicolor and A. bicolor pacifica in south-east Asian waters. Journal of Fish Biology, 94(5):752– 758, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuss:2019:EBV

[2883] Theodora Fuss, Sabine Nöbel, and Klaudia Witte. It’s in the eye of the beholder: visual lateralisation in response to the social environment in poeciliids. Journal of Fish Biology, 94(5):759–771, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2019:ECG

[2884] Yu Li, Wei Song, Yei Fei Zhu, Tian Yu Zhu, Ling Bo Ma, and Ming You Li. Evolutionarily conserved vasa identifies embryonic and gonadal germ cells in spinyhead croaker Collichthys lucidus. Journal of Fish Biology, 94(5):772–780, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rangel:2019:RPB

[2885] Bianca de Sousa Rangel, Nigel E. Hussey, Aline D. Gomes, Alexandre Rodrigues, Luiz A. Martinelli, and Renata Guimarães Moreira. Resource partitioning between two young-of-year cownose rays Rhinoptera bonasus and R. brasiliensis within a communal nursery inferred by trophic biomarkers. Journal of Fish Biology, 94(5):781–788, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weideli:2019:SFD

[2886] Ornella C. Weideli, Yannis P. Papastamatiou, and Serge Planes. Size frequency, dispersal distances and variable growth rates of young sharks in a multi-species aggregation. Journal of Fish Biology, 94(5):789–797, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nogueira:2019:HBN

[2887] Acácio F. Nogueira, Francisco Langeani, and André L. Netto-Ferreira. Hemiodus bimaculatus, a new species of Hemiodontidae from the Rio Tapajós drainage, Brazil (Ostariophysi: Characiformes). Journal of Fish Biology, 94(5):798–803, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2019:ENT

[2888] Ai Sun, Hua Zhu, Ying Dong, Wei Wang, and Hong Xia Hu. Establishment of a novel testicular cell line from sterlet Acipenser ruthenus and evaluation of its applications. Journal of Fish Biology, 94(5):804– 809, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loeppky:2019:OSA

[2889] Alison R. Loeppky, Bryan C. Chakoumakos, Brenda M. Pracheil, and W. Gary Anderson. Otoliths of sub-adult Lake Sturgeon Acipenser fulvescens contain aragonite and vaterite calcium carbonate polymorphs. Journal of Fish Biology, 94(5):810–814, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zarri:2019:KTM

[2890] Liam J. Zarri, Sarah B. Mehl, Eric P. Palkovacs, and Nann A. Fangue. Key transitions in morphological development improve age estimates in white sturgeon Acipenser transmontanus. Journal of Fish Biology, 94(5): 815–819, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mumby:2019:SGR

[2891] Peter J. Mumby. Survival of a grey reef shark Carcharhinus amblyrhynchos without a dorsal fin. Journal of Fish Biology, 94(5):820–822, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crivelaro:2019:BPE

[2892] Roberta M. Crivelaro, Roberto Thiesen, Marcela Aldrovani, Paloma E. S. Silva, Alexandre A. F. Barros Sobrinho, and Paola C. Moraes. Behavioural and physiological effects of methadone in the perioperative period on the Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 94(5):823–827, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Risto:2019:BRM

[2893] Melissa Risto. Book review: Marine fishes of Arctic Canada. Journal of Fish Biology, 94(5):828, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2019:BET

[2894] Laith A. Jawad. Biology and ecology of toxic pufferfish. Journal of Fish Biology, 94(5):829, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:Ca

[2895] Anonymous. Corrigendum. Journal of Fish Biology, 94(5):830, May 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIf

[2896] Anonymous. Issue information. Journal of Fish Biology, 94(6):831, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Engelhard:2019:SUE

[2897] Georg H. Engelhard, John K. Pinnegar, Martin I. Taylor, and Audrey J. Geffen. The sustainable use and exploitation of fishes. Journal of Fish Biology, 94(6):833–836, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stead:2019:UST

[2898] Selina Marguerite Stead. Using systems thinking and open innovation to strengthen aquaculture policy for the United Nations Sustainable Development Goals. Journal of Fish Biology, 94(6):837–844, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2019:SRS

[2899] Steven J. Cooke, William M. Twardek, Andrea J. Reid, Robert J. Lennox, Sascha C. Danylchuk, Jacob W. Brownscombe, Shannon D. Bower, Robert Arlinghaus, Kieran Hyder, and Andy J. Danylchuk. Searching for responsible and sustainable recreational fisheries in the Anthropocene. Journal of Fish Biology, 94(6):845–856, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jimenez-Alvarado:2019:HPC

[2900] David Jiménez-Alvarado, Airam Sarmiento-Lezcano, Airam Guerra-Marrero, Fernando Tuya, Ángelo Santana Del Pino, Michael J. Sealey, and José J. Castro. Historical photographs of captures of recreational fishers indicate overexploitation of nearshore resources at an oceanic island. Journal of Fish Biology, 94(6):857–864, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matern:2019:ERF

[2901] Sven Matern, Matthias Emmrich, Thomas Klefoth, Christian Wolter, Robert Nikolaus, Nicola Wegener, and Robert Arlinghaus. Effect of recreational-fisheries management on fish biodiversity in gravel pit lakes, with contrasts to unmanaged lakes. Journal of Fish Biology, 94(6):865– 881, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Montero:2019:TSP

[2902] Lorena Couce Montero, Villy Christensen, Alberto Bilbao Sieyro, Yeray Pérez González, David Jiménez-Alvarado, and José J. Castro. Temporal and spatial predictions of effect of alternative fishing policies for the Gran Canaria marine ecosystem. Journal of Fish Biology, 94(6): 882–895, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:2019:EIE

[2903] Katie Thomas, Tom Hansen, Deirdre Brophy, Niall Ó Maoiléidigh, and Per Gunnar Fjelldal. Experimental investigation of the effects of temperature and feeding regime on scale growth in Atlantic salmon Salmo salar post-smolts. Journal of Fish Biology, 94(6):896–908, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evers:2019:WTA

[2904] Hans-Georg Evers, John K. Pinnegar, and Martin I. Taylor. Where are they all from? — sources and sustainability in the ornamental freshwater fish trade. Journal of Fish Biology, 94(6):909–916, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinnegar:2019:UUK

[2905] John K. Pinnegar and Joanna M. Murray. Understanding the United Kingdom marine aquarium trade — a mystery shopper study of species on sale. Journal of Fish Biology, 94(6):917–924, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2019:WCO

[2906] Tracey A. King. Wild caught ornamental fish: a perspective from the UK ornamental aquatic industry on the sustainability of aquatic organisms and livelihoods. Journal of Fish Biology, 94(6):925–936, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nguyen:2019:HCF

[2907] Hai-Thanh T. Nguyen, A-Nga T. Tran, Le Thi L. Ha, Dang N. Ngo, Binh T. Dang, and Audrey J. Geffen. Host choice and fitness of anemonefish Amphiprion ocellaris (Perciformes: Pomacentridae) living with host anemones (Anthozoa: Actiniaria) in captive conditions. Journal of Fish Biology, 94(6):937–947, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cocks:2019:FRR

[2908] Lewis N. Cocks, Jonathan K. L. Mee, and Alex P. Shepherd. First record of recurring reproduction of captive tawny nurse sharks Nebrius ferrugineus. Journal of Fish Biology, 94(6):948–951, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agnew:2019:WDS

[2909] David J. Agnew. Who determines sustainability? Journal of Fish Biology, 94(6):952–957, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:2019:SIA

[2910] Ching-Tsun Chang, Wei-Chuan Chiang, Yun-Chen Chang, Michael K. Musyl, Chi-Lu Sun, Daniel J. Madigan, Aaron B. Carlisle, Hung-Hung Hsu, Qi-Xuan Chang, Nan-Jay Su, Yuan-Shing Ho, and Chen-Te Tseng. Stable isotope analysis reveals ontogenetic feeding shifts in Pacific blue marlin (Makaira nigricans) off eastern Taiwan. Journal of Fish Biology, 94(6):958–965, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2019:BDP

[2911] Joana F. Silva and Jim R. Ellis. Bycatch and discarding patterns of dogfish and sharks taken in English and Welsh commercial fisheries. Journal of Fish Biology, 94(6):966–980, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vieira:2019:DWF

[2912] Rui P. Vieira, Clive N. Trueman, Lisa Readdy, Andrew Kenny, and John K. Pinnegar. Deep-water fisheries along the British Isles continental slopes: status, ecosystem effects and future perspectives. Journal of Fish Biology, 94(6):981–992, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2019:CWC

[2913] Régis V. S. Santos, Ana M. Novoa-Pabon, Hélder M. Silva, and Mário R. Pinho. Can we consider the stocks of alfonsinos Beryx splendens and Beryx decadactylus from the Azores a discrete fishery management unit? Journal of Fish Biology, 94(6):993–1000, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sendek:2019:ESO

[2914] Dmitry S. Sendek and Dmitry V. Bogdanov. European smelt Osmerus eperlanus in the eastern Gulf of Finland, Baltic Sea: Stock status and fishery. Journal of Fish Biology, 94(6):1001–1010, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:2019:WMM

[2915] Robert B. Thorpe. What is multispecies MSY? A worked example from the North Sea. Journal of Fish Biology, 94(6):1011–1018, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2019:MTF

[2916] Vittoria L. Elliott, Phen Chheng, Sophorn Uy, and Gordon W. Holtgrieve. Monitoring of tropical freshwater fish resources for sustainable use. Journal of Fish Biology, 94(6):1019–1025, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lamb:2019:CDF

[2917] Philip D. Lamb, Ewan Hunter, John K. Pinnegar, Jeroen van der Kooij, Simon Creer, and Martin I. Taylor. Cryptic diets of forage fish: jellyfish consumption observed in the Celtic Sea and western English Channel. Journal of Fish Biology, 94(6):1026–1032, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrett:2019:EPH

[2918] James H. Barrett. An environmental (pre)history of European fishing: past and future archaeological contributions to sustainable fisheries. Journal of Fish Biology, 94(6):1033–1044, June 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIg

[2919] Anonymous. Issue information. Journal of Fish Biology, 95(1):1–2, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:2019:SEF

[2920] Andrew B. Gill. The sensory ecology of fishes. Journal of Fish Biology, 95 (1):3–4, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2019:CCV

[2921] N. Justin Marshall, Fabio Cortesi, Fanny de Busserolles, Uli E. Siebeck, and Karen L. Cheney. Colours and colour vision in reef fishes: Past, present and future research directions. Journal of Fish Biology, 95(1): 5–38, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Putland:2019:EFH

[2922] Rosalyn L. Putland, John C. Montgomery, and Craig A. Radford. Ecology of fish hearing. Journal of Fish Biology, 95(1):39–52, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mogdans:2019:SEF

[2923] Joachim Mogdans. Sensory ecology of the fish lateral-line system: Morphological and physiological adaptations for the perception of hydrodynamic stimuli. Journal of Fish Biology, 95(1):53–72, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Formicki:2019:MF

[2924] Krzysztof Formicki, Agata Korzelecka-Orkisz, and Adam Tański. Magnetoreception in fish. Journal of Fish Biology, 95(1):73–91, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crampton:2019:EEE

[2925] William G. R. Crampton. Electroreception, electrogenesis and electric signal evolution. Journal of Fish Biology, 95(1):92–134, July 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newton:2019:EMF

[2926] Kyle C. Newton, Andrew B. Gill, and Stephen M. Kajiura. Electroreception in marine fishes: chondrichthyans. Journal of Fish Biology, 95 (1):135–154, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kasumyan:2019:TSF

[2927] Alexander O. Kasumyan. The taste system in fishes and the effects of environmental variables. Journal of Fish Biology, 95(1):155–178, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schweikert:2019:VRS

[2928] Lorian E. Schweikert, Eleanor M. Caves, Sarah E. Solie, Tracey T. Sutton, and Sönke Johnsen. Variation in rod spectral sensitivity of fishes is best predicted by habitat and depth. Journal of Fish Biology, 95(1): 179–185, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nieman:2019:VPI

[2929] Chelsey L. Nieman and Suzanne M. Gray. Visual performance impaired by elevated sedimentary and algal turbidity in walleye Sander vitreus and emerald shiner Notropis atherinoides. Journal of Fish Biology, 95 (1):186–199, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jokela-Maatta:2019:PEP

[2930] Mirka Jokela-Määttä, Martta Viljanen, Noora Nevala, Kristian Donner, and Christer Brönmark. Photoreceptors and eyes of pikeperch Sander lucioperca, pike Esox lucius, perch Perca fluviatilis and roach Rutilus rutilus from a clear and a brown lake. Journal of Fish Biology, 95(1): 200–213, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flamarique:2019:LED

[2931] Inigo Novales Flamarique. Light exposure during embryonic and yolksac alevin development of Chinook salmon Oncorhynchus tshawytscha does not alter the spectral phenotype of photoreceptors. Journal of Fish Biology, 95(1):214–221, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saez:2019:MSL

[2932] Sylvia Sáez, German Pequeño, Roberto Jaramillo, and Luis Vargas-Chacoff. Mechanosensory system of the lateral line in the subantarctic Patagonian blenny Eleginops maclovinus. Journal of Fish Biology, 95(1):222–227, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kerr:2019:MFD

[2933] James R. Kerr and Paul S. Kemp. Masking a fish’s detection of environmental stimuli: application to improving downstream migration at river infrastructure. Journal of Fish Biology, 95(1):228–237, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mussen:2019:AUV

[2934] Timothy D. Mussen and Joseph J. Cech, Jr. Assessing the use of vibrations and strobe lights at fish screens as enhanced deterrents for two estuarine fishes. Journal of Fish Biology, 95(1):238–246, July 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jesus:2019:ABA

[2935] Joaquim Jesus, Maria Clara P. Amorim, Paulo J. Fonseca, Amílcar Teixeira, Silvestre Natário, João Carrola, Simone Varandas, Luís Torres Pereira, and Rui M. V. Cortes. Acoustic barriers as an acoustic deterrent for native potamodromous migratory fish species. Journal of Fish Biology, 95(1):247–255, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vetter:2019:ELS

[2936] Brooke J. Vetter, Loranzie S. Rogers, and Allen F. Mensinger. The effect of light stimuli on dark-adapted visual sensitivity in invasive silver carp Hypophthalmichthys molitrix and bighead carp H. nobilis. Journal of Fish Biology, 95(1):256–262, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Znotinas:2019:AAV

[2937] Katherine R. Znotinas and Emily M. Standen. Aerial and aquatic visual acuity of the grey bichir Polypterus senegalus, as estimated by optokinetic response. Journal of Fish Biology, 95(1):263–273, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maximino:2019:SEO

[2938] Caio Maximino, Rhayra X. do Carmo Silva, Kimberly dos Santos Campos, Jeisiane S. de Oliveira, Sueslene P. Rocha, Maryana P. Pyterson, Dainara P. dos Santos Souza, Leonardo M. Feitosa, Saulo R. Ikeda, Ana F. N. Pimentel, Pâmila N. F. Ramos, Bruna P. D. Costa, Anderson M. Herculano, Denis B. Rosemberg, Diógenes H. Siqueira-Silva, and Monica Lima-Maximino. Sensory ecology of ostariophysan alarm substances. Journal of Fish Biology, 95(1):274–286, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barkhymer:2019:OMC

[2939] Alison J. Barkhymer, Shawn G. Garrett, and Brian D. Wisenden. Olfactorily-mediated cortisol response to chemical alarm cues in zebrafish Danio rerio. Journal of Fish Biology, 95(1):287–292, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ueda:2019:SMN

[2940] Hiroshi Ueda. Sensory mechanisms of natal stream imprinting and homing in Oncorhynchus spp. Journal of Fish Biology, 95(1):293–303, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horreo:2019:NAR

[2941] Jose L. Horreo, Andrew M. Griffiths, Gonzalo Machado-Schiaffino, Jamie R. Stevens, and Eva Garcia-Vazquez. Northern areas as refugia for temperate species under current climate warming: Atlantic salmon (Salmo salar L.) as a model in northern Europe. Journal of Fish Biology, 95(1):304–310, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hu:2019:PRS

[2942] Yinan Hu, John E. Majoris, Peter M. Buston, and Jacqueline F. Webb. Potential roles of smell and taste in the orientation behaviour of coralreef fish larvae: insights from morphology. Journal of Fish Biology, 95 (1):311–323, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chatterjee:2019:EEO

[2943] Subhendu K. Chatterjee, Chandan Malick, Samir Bhattacharya, Vettath R. Suresh, Rakesh Kundu, and Surjya K. Saikia. Ectopic expression of olfactory receptors and associated G-protein subunits in the head integument of the amphihaline migratory fish hilsa Tenualosa ilisha. Journal of Fish Biology, 95(1):324–334, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bressman:2019:WDF

[2944] Noah R. Bressman, Mark Simms, Benjamin M. Perlman, and Miriam A. Ashley-Ross. Where do fish go when stranded on land? Terrestrial orientation of the mangrove rivulus Kryptolebias marmoratus. Journal of Fish Biology, 95(1):335–344, July 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIh

[2945] Anonymous. Issue information. Journal of Fish Biology, 95(2):345–346, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lujan:2019:FHT

[2946] Nathan K. Lujan. Fishes that hold their breath? Journal of Fish Biology, 95(2):347, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossi:2019:RBS

[2947] Amélie Rossi, Marc Levaray, Christelle Paillon, Eric D. H. Durieux, Vanina Pasqualini, and Sylvia Agostini. Relationship between swimming capacities and morphological traits of fish larvae at settlement stage: a study of several coastal Mediterranean species. Journal of Fish Biology, 95(2):348–356, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rose:2019:SSV

[2948] Emily Rose, Mackenzie Simmonds, Alana L. Hayashida-Boyles, and Heather D. Masonjones. Seasonal and spatial variation in the reproductive biology of the dwarf seahorse Hippocampus zosterae. Journal of Fish Biology, 95(2):357–366, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oskarsson:2019:VEM

[2949] Gudmundur J. Óskarsson, Christopher T. Taggart, and Robert L. Stephenson. Variation in egg mass within two Atlantic herring Clupea harengus stocks. Journal of Fish Biology, 95(2):367–378, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arroyave:2019:DBU

[2950] Jairo Arroyave, Christopher M. Martinez, and Melanie L. J. Stiassny. DNA barcoding uncovers extensive cryptic diversity in the African longfin tetra Bryconalestes longipinnis (Alestidae: Characiformes). Journal of Fish Biology, 95(2):379–392, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2019:SRG

[2951] Dongmei Ma, Jiajia Fan, Yuanyuan Tian, Peng Jiang, Junjie Wang, Huaping Zhu, and Junjie Bai. Selection of reference genes for quantitative real-time PCR normalisation in largemouth bass Micropterus salmoides fed on alternative diets. Journal of Fish Biology, 95(2):393–400, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frater:2019:DGA

[2952] Paul N. Frater, Birgir Hrafnkelsson, Bjarki Th. Elvarsson, and Gunnar Stefansson. Drivers of growth for Atlantic cod (Gadus morhua L.) in Icelandic waters — a Bayesian approach to determine spatiotemporal variation and its causes. Journal of Fish Biology, 95(2):401–410, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

L:2019:NSP

[2953] Maria D. Escobar L., Rafaela P. Ota, Antonio Machado-Allison, Juana Andrade-López, Izeni P. Farias, and Tomas Hrbek. A new species of Piaractus (Characiformes: Serrasalmidae) from the Orinoco Basin with a redescription of Piaractus brachypomus. Journal of Fish Biology, 95 (2):411–427, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Serrat:2019:TOB

[2954] Alba Serrat, Josep Lloret, Xenia Frigola-Tepe, and Marta Muñoz. Tradeoffs between life-history traits in a coldwater fish in the Mediterranean Sea: the case of blue whiting Micromesistius poutassou. Journal of Fish Biology, 95(2):428–443, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bounket:2019:OSA

[2955] Bernadette Bounket, Pierre Gibert, Vincent Gennotte, Christine Argillier, Georges Carrel, Anthony Maire, Maxime Logez, and Fabien Morat. Otolith shape analysis and daily increment validation during ontogeny of larval and juvenile European chub Squalius cephalus. Journal of Fish Biology, 95(2):444–452, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tencatt:2019:DOM

[2956] Luiz F. C. Tencatt, Flávio C. T. Lima, and Marcelo R. Britto. Deconstructing an octogenarian misconception reveals the true Corydoras arcuatus Elwin 1938 (Siluriformes: Callichthyidae) and a new Corydoras species from the Amazon basin. Journal of Fish Biology, 95(2):453–471, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Satterfield:2019:ESS

[2957] Darien R. Satterfield and Mark A. Steele. Effects of size and sex on the courting success and foraging behaviour of Embiotoca jacksoni. Journal of Fish Biology, 95(2):472–479, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daroux:2019:MFL

[2958] Aurélie Daroux, François Martignac, Marie Nevoux, Jean L. Baglinière, Dominique Ombredane, and Jean Guillard. Manual fish length measurement accuracy for adult river fish using an acoustic camera (DIDSON). Journal of Fish Biology, 95(2):480–489, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burgos-Vazquez:2019:RCR

[2959] Maria I. Burgos-Vázquez, Felipe Galván-Magaña, Maribel Carrera-Fernández, Rosa I. Ochoa-Báez, and Maria C. Oddone. Reproductive characteristics and reproductive tract anatomy of the California butterfly ray Gymnura marmorata (Myliobatiformes: Gymnuridae). Journal of Fish Biology, 95(2):490–501, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Long:2019:EGC

[2960] Nicholas P. Long and Stacy C. Farina. Enormous gill chambers of deepsea coffinfishes (Lophiiformes: Chaunacidae) support unique ventilatory specialisations such as breath holding and extreme inflation. Journal of Fish Biology, 95(2):502–509, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhu:2019:MCF

[2961] Yaoyao Zhu, Shijuan Shan, Hanxiao Feng, Lei Jiang, Liguo An, Guiwen Yang, and Hua Li. Molecular characterization and functional analysis of interferon regulatory factor 9 (irf9) in common carp Cyprinus carpio: a pivotal molecule in the Ifn response against pathogens. Journal of Fish Biology, 95(2):510–519, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pouca:2019:LSP

[2962] Catarina Vila Pouca and Culum Brown. Lack of social preference between unfamiliar and familiar juvenile Port Jackson sharks Heterodontus portusjacksoni. Journal of Fish Biology, 95(2):520–526, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gray:2019:SWG

[2963] Benjamin P. Gray, Mary Anne Bishop, and Sean Powers. Structure of winter groundfish feeding guilds in Pacific herring Clupea pallasii and walleye pollock Gadus chalcogrammus nursery fjords. Journal of Fish Biology, 95(2):527–539, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2019:ERG

[2964] Fang Ma, Zhe Liu, Jinqiang Huang, Yujun Kang, and Jianfu Wang. Evaluation of reference genes for quantitative real-time PCR analysis of messenger RNAs and microRNAs in rainbow trout Oncorhynchus mykiss under heat stress. Journal of Fish Biology, 95(2):540–554, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simpson:2019:EWS

[2965] Michael Simpson, Rebecca L. Morris, David Harasti, and Ross A. Coleman. The endangered White’s seahorse Hippocampus whitei chooses artificial over natural habitats. Journal of Fish Biology, 95(2):555–561, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filous:2019:LHC

[2966] Alexander Filous, Robert J. Lennox, Richard R. Coleman, Alan M. Friedlander, Eric E. G. Clua, and Andy J. Danylchuk. Life-history characteristics of an exploited bonefish Albula glossodonta population in a remote South Pacific atoll. Journal of Fish Biology, 95(2):562–574, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solaas:2019:CSE

[2967] Martine R. Solås, Helge Skoglund, and Anne G. V. Salvanes. Can structural enrichment reduce predation mortality and increase recaptures of hatchery-reared Atlantic salmon Salmo salar L. fry released into the wild? Journal of Fish Biology, 95(2):575–588, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broadhurst:2019:MEI

[2968] Matt K. Broadhurst, Camila Domit, Tailisi H. Trevizani, Vincent Raoult, and Russell B. Millar. Mother–embryo isotope fractionation in the pygmy devilray Mobula kuhlii cf. eregoodootenkee. Journal of Fish Biology, 95 (2):589–593, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daSilva:2019:PMN

[2969] João Paulo Capretz Batista da Silva and Thiago Silva Loboda. Potamotrygon marquesi, a new species of neotropical freshwater stingray (Potamotrygonidae) from the Brazilian Amazon Basin. Journal of Fish Biology, 95(2):594–612, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz-Cordovez:2019:ELT

[2970] Rodrigo Muñoz-Cordovez, Lucas De La Maza, Alejandro Pérez-Matus, and Sergio A. Carrasco. Embryonic and larval traits of the temperate damselfish Chromis crusma reveal important similarities with other Pomacentridae throughout the family’s thermal range. Journal of Fish Biology, 95(2):613–623, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luzzatto:2019:LVN

[2971] Diego C. Luzzatto and María L. Estalles. Leptonotus vincentae, a new pipefish species (Syngnathidae: Syngnathinae) from the south-west Atlantic Ocean near northern Patagonia. Journal of Fish Biology, 95(2):

624–632, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hahn:2019:GTI

[2972] Lisiane Hahn, Priscilla C. Silva, Maria C. Malabarba, Luiz R. Malabarba, Luís F. Da Câmara, Leonardo D. Nunes, Leonardo S. Machado, Eduardo G. Martins, and Ronaldo B. Barthem. Genetics and telemetry indicate unexpected movements among structured populations for Brachyplatystoma platynemum in the Amazon. Journal of Fish Biology, 95(2):633–637, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watz:2019:SCH

[2973] Johan Watz. Structural complexity in the hatchery rearing environment affects activity, resting metabolic rate and post-release behaviour in brown trout Salmo trutta. Journal of Fish Biology, 95(2):638–641, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morales:2019:FST

[2974] Naiti Morales, Amy Rose Coghlan, Gonzalo Hayden, and Paula Guajardo. First sighting of a tropical benthic reef shark species at Rapa Nui: chance dispersal or a sign of things to come? Journal of Fish Biology, 95(2):642–646, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kroska:2019:ESC

[2975] Anita C. Kroska, Nathan Wolf, Roman Dial, and Bradley P. Harris. Exploring sample cross-contamination in fish epidermal mucus. Journal of Fish Biology, 95(2):647–650, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Millane:2019:UWO

[2976] Michael Millane, Laura Walsh, William K. Roche, and Patrick G. Gargan. Unprecedented widespread occurrence of pink salmon Oncorhynchus gorbuscha in Ireland in 2017. Journal of Fish Biology, 95(2):651–654, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hume:2019:HTI

[2977] John B. Hume. Higher temperatures increase developmental rate & reduce body size at hatching in the small-eyed skate Raja microocellata: implications for exploitation of an elasmobranch in warming seas. Journal of Fish Biology, 95(2):655–658, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez-Hernandez:2019:SGA

[2978] Javier Sánchez-Hernández and Fernando Cobo. Suboptimal growth among individuals of brown trout (Salmo trutta) in a temperate river. Journal of Fish Biology, 95(2):659–662, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pryor:2019:TUG

[2979] Kimberley J. Pryor and Ashley M. Milton. Tool use by the graphic tuskfish Choerodon graphicus. Journal of Fish Biology, 95(2):663–667, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kikko:2019:DGI

[2980] Takeshi Kikko, Daisuke Ishizaki, Taiga Yodo, Shizuo Aino, Kunihiko Kuwamura, Haruo Okamoto, Morihito Nemoto, Kazuki Yoneda, Nobuyuki Oue, Akihisa Sakai, Yasuhiro Fujioka, Yoshiaki Kai, Takuya Sato, and Kouji Nakayama. Daily growth increments in otoliths of wildcaught honmoroko Gnathopogon caerulescens. Journal of Fish Biology, 95(2):668–672, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vrtilek:2019:RGL

[2981] Milan Vrtílek, Jakub Žák, Matej Polačik, Radim Blažek, and Martin Reichard. Rapid growth and large body size in annual fish populations are compromised by density-dependent regulation. Journal of Fish Biology, 95(2):673–678, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McDevitt:2019:EDM

[2982] Allan D. McDevitt, Naiara Guimarães Sales, Samuel S. Browett, Abbie O. Sparnenn, Stefano Mariani, Owen S. Wangensteen, Ilaria Coscia, and Chiara Benvenuto. Environmental DNA metabarcoding as an effective and rapid tool for fish monitoring in canals. Journal of Fish Biology, 95(2):679–682, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soares:2019:SDB

[2983] Karla D. A. Soares. Sexually dimorphic body proportions in the catshark genus Scyliorhinus (Chondrichthyes: Carcharhiniformes:

Scyliorhinidae). Journal of Fish Biology, 95(2):683–685, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2019:PSA

[2984] Ian J. Winfield. Plastic soup: an atlas of ocean pollution. Journal of Fish Biology, 95(2):686, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2019:BEP

[2985] Neil B. Metcalfe. Behavior and ecology of Pacific salmon and trout. Journal of Fish Biology, 95(2):687–688, August 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIi

[2986] Anonymous. Issue information. Journal of Fish Biology, 95(3):689–690, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:AFT

[2987] Anonymous. Ancestry.fish: Tracing fish lineages with genomic DNA. Journal of Fish Biology, 95(3):691, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:2019:APR

[2988] Andrew Ferguson, Thomas E. Reed, Tom F. Cross, Philip McGinnity, and Paulo A. Prodöhl. Anadromy, potamodromy and residency in brown trout Salmo trutta: the role of genes and the environment. Journal of Fish Biology, 95(3):692–718, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prodohl:2019:IAB

[2989] Paulo A. Prodöhl, Andrew Ferguson, Caroline R. Bradley, Robin Ade, Colin Roberts, E. J. Keay, Artur R. Costa, and Rosaleen Hynes. Impacts of acidification on brown trout Salmo trutta populations and the contribution of stocking to population recovery and genetic diversity. Journal of Fish Biology, 95(3):719–742, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Francis:2019:GBC

[2990] Malcolm P. Francis and Caoimhghin Ó Maolagáin. Growth-band counts from elephantfish Callorhinchus milii fin spines do not correspond with independently estimated ages. Journal of Fish Biology, 95(3):743–752, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehemann:2019:UTL

[2991] Nicolás R. Ehemann, Lorem D. V. González-González, Alejandro Tagliafico, and Simon Weigmann. Updated taxonomic list and conservation status of chondrichthyans from the exclusive economic zone of Venezuela, with first generic and specific records. Journal of Fish Biology, 95(3): 753–771, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silveira:2019:MRA

[2992] Mayara M. Silveira, Jéssica Ferreira de Souza, Ana L. Pires Moreira, Priscila F. Silva, and Ana C. Luchiari. Memory retention of appetitive and aversive conditioning in the damselfish Stegastes fuscus. Journal of Fish Biology, 95(3):772–780, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayden:2019:FFL

[2993] Brian Hayden, Myriam Ferron, Richard A. Cunjak, and Kurt Samways. Fruit of the forest — larval sea lamprey Petromyzon marinus are fuelled by allochthonous resources. Journal of Fish Biology, 95(3):781–792, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elsner:2019:BCD

[2994] Rick A. Elsner and J. Mark Shrimpton. Behavioural changes during the parr–smolt transformation in Coho salmon Oncorhynchus kisutch: is it better to be cool? Journal of Fish Biology, 95(3):793–801, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yurtseva:2019:ATS

[2995] Anastasia Yurtseva, Kristina Noreikiene, Dmitry Lajus, Zitong Li, Tarja Alapassi, Tatiana Ivanova, Mikhail Ivanov, Pavel Golovin, Sami Vesala, and Juha Merilä. Aging three-spined sticklebacks Gasterosteus aculeatus: comparison of estimates from three structures. Journal of Fish Biology, 95(3):802–811, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendes:2019:OTM

[2996] Thiago C. Mendes, Juan P. Quimbayo, Helena F. Bouth, Luana P. S. Silva, and Carlos E. L. Ferreira. The omnivorous triggerfish Melichthys niger is a functional herbivore on an isolated Atlantic oceanic island. Journal of Fish Biology, 95(3):812–819, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borie:2019:DCF

[2997] Alfredo Borie, Diogo B. Hungria, Heba Ali, Carolina R. Doria, Michael L. Fine, and Paulo E. Travassos. Disturbance calls of five migratory Characiformes species and advertisement choruses in Amazon spawning sites. Journal of Fish Biology, 95(3):820–832, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeLosRios-Mendez:2019:LID

[2998] Juanita De Los Ríos-Méndez, Gabriela A. Kanamori, and Mauricio F. Landaeta. Lunar influences in the diet habits and selectivity of larval clingfish Gobiesox marmoratus. Journal of Fish Biology, 95(3):833–846, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haxton:2019:ISS

[2999] Tim Haxton and Mike Friday. Interaction of sauger Sander canadensis and walleye Sander vitreus in a large, shallow northern river. Journal of Fish Biology, 95(3):847–854, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peel:2019:SSD

[3000] Richard A. Peel, Jaclyn M. Hill, Geraldine C. Taylor, Denis Tweddle, and Olaf L. F. Weyl. Species succession and the development of a lacustrine fish community in an ephemeral lake. Journal of Fish Biology, 95(3): 855–869, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cogliati:2019:REA

[3001] Karen M. Cogliati, Julia R. Unrein, Carl B. Schreck, and David L. G. Noakes. Rearing environment affects spatial learning in juvenile Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 95(3):870– 880, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moyer:2019:BPF

[3002] Joshua K. Moyer, Stephanie F. Shannon, and Duncan J. Irschick. Bite performance and feeding behaviour of the sand tiger shark Carcharias taurus. Journal of Fish Biology, 95(3):881–892, September 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvas:2019:IES

[3003] Malthe Hvas and Frode Oppedal. Influence of experimental set-up and methodology for measurements of metabolic rates and critical swimming speed in Atlantic salmon Salmo salar. Journal of Fish Biology, 95(3): 893–902, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2019:RDM

[3004] Ruiqi Zhang, Jinliang Zhao, YueYue Hao, Xiaoyin Cao, Yan Zhao, and Shoujie Tang. Retinal development in mandarinfish Siniperca chuatsi and morphological analysis of the photoreceptor layer. Journal of Fish Biology, 95(3):903–917, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2019:NST

[3005] Vinícius J. C. Reis, Mário C. C. de Pinna, and Tiago C. Pessali. A new species of Trichomycterus Valenciennes 1832 (Trichomycteridae: Siluriformes) from the Rio Doce drainage with remarkable similarities with Bullockia and a CT-scan survey. Journal of Fish Biology, 95(3):918–931, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albornoz-Garzon:2019:NSH

[3006] Juan G. Albornoz-Garzón, Alejandro Méndez-López, Carlos DoNascimiento, and Flávio C. T. Lima. A new species of Hemigrammus (Characiformes: Characidae) from western Amazon River basin, Colombia. Journal of Fish Biology, 95(3):932–939, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dickson:2019:OCA

[3007] Juleen M. Dickson and Kathryn A. Dickson. Ontogenetic change in the amount and position of slow-oxidative myotomal muscle in relationship to regional endothermy in juvenile yellowfin tuna Thunnus albacares. Journal of Fish Biology, 95(3):940–951, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitmarsh:2019:LLC

[3008] Sasha K. Whitmarsh, Peter G. Fairweather, and Charlie Huveneers. Lack of light colour effects when sampling fish at night in low visibility environments. Journal of Fish Biology, 95(3):952–955, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winter:2019:NLS

[3009] Emily R. Winter, Marina Nyqvist, and J. Robert Britton. Non-lethal sampling for stable isotope analysis of pike Esox lucius: how mucus, scale and fin tissue compare to muscle. Journal of Fish Biology, 95(3): 956–958, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liberoff:2019:MIE

[3010] Ana L. Liberoff, Alejandra A. Leal, María E. Díaz De Vivar, Carla M. Riva-Rossi, and Miguel A. Pascual. Maternal influence on early lipid content in an introduced partially anadromous population of Oncorhynchus mykiss. Journal of Fish Biology, 95(3):959–964, September 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedeholm:2019:TCS

[3011] Rasmus Hedeholm, Trine Qvist, Julius Nielsen, and Peter Grønkjær. Temporal changes in size-at-maturity of black dogfish Centroscyllium fabricii. Journal of Fish Biology, 95(3):965–968, September 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gallagher:2019:EPE

[3012] Austin J. Gallagher, Michael J. Lawrence, Sofia M. R. Jain-Schlaepfer, Kathleen M. Gilmour, Alexander D. M. Wilson, and Steven J. Cooke. Effects of predator exposure on baseline and stress-induced glucocorticoid hormone concentrations in pumpkinseed Lepomis gibbosus. Journal of Fish Biology, 95(3):969–973, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebner:2019:JSG

[3013] Brendan C. Ebner, James A. Donaldson, and Danswell Starrs. Juvenile silver grunter Mesopristes argenteus shift benthic objects to access food. Journal of Fish Biology, 95(3):974–978, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horiuchi:2019:BDE

[3014] Tomoya Horiuchi, Reiji Masuda, Hiroaki Murakami, Satoshi Yamamoto, and Toshifumi Minamoto. Biomass-dependent emission of environmental DNA in jack mackerel Trachurus japonicus juveniles. Journal of Fish Biology, 95(3):979–981, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:2019:STW

[3015] Darren Andrew Whitehead, David Petatán Ramírez, Damien Olivier, Rogelio Gonzalez Armas, Francesca Pancaldi, and Felipe Galvan-Magana. Seasonal trends in whale shark Rhincodon typus sightings in an established tourism site in the Gulf of California, Mexico. Journal of Fish Biology, 95(3):982–984, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2019:BRM

[3016] Laith A. Jawad. Book review: Management of Marine Protected Areas: a Network Perspective. By P. D. Goriup (Ed.). 312 pp. Published by John Wiley & Sons, Ltd, Oxford, U.K., 2017. Price £89.95. ISBN: 978-1-119-07577-6. Journal of Fish Biology, 95(3):985, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McConville:2019:BRD

[3017] James McConville. Book review: Diagnosis and Control of Diseases of Fish and Shellfish. By B. Austin & A. Newaj-Fyzul (Eds). 320 pp. Published by John Wiley & Sons, Ltd., Oxford, U.K., 2017. Price £130.00. ISBN: 978-1-119-15210-1. Journal of Fish Biology, 95(3):986, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:Eb

[3018] Anonymous. Erratum. Journal of Fish Biology, 95(3):987–988, September 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIj

[3019] Anonymous. Issue information. Journal of Fish Biology, 95(4):989–990, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2019:SET

[3020] Michel J. Kaiser. Sheltering from the effects of thermal stress. Journal of Fish Biology, 95(4):991, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:2019:DBS

[3021] Yuuki Y. Watanabe and Yannis P. Papastamatiou. Distribution, body size and biology of the megamouth shark Megachasma pelagios. Journal of Fish Biology, 95(4):992–998, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2019:SRS

[3022] Eric Garcia, Cristy A. Rice, Douglas J. Eernisse, Kristy L. Forsgren, Juan P. Quimbayo, and Greg W. Rouse. Systematic relationships of sympatric pipefishes (Syngnathus spp.): a mismatch between morphological and molecular variation. Journal of Fish Biology, 95(4):999–1012, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yin-Liao:2019:FAT

[3023] Irene Yin-Liao, Patricia A. Wright, and Frédéric Laberge. Factors affecting terrestrial movement in the amphibious mangrove rivulus (Kryptolebias marmoratus). Journal of Fish Biology, 95(4):1013–1021, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shao:2019:EED

[3024] Tianyun Shao, Xiaoyan Chen, Dingxuan Zhai, Tao Wang, Xiaohua Long, and Zhaopu Liu. Evaluation of the effects of different stocking densities on growth and stress responses of juvenile hybrid grouper [female sign] Epinephelus fuscoguttatus × [male sign] Epinephelus lanceolatus in recirculating aquaculture systems. Journal of Fish Biology, 95(4):1022–1029, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chong:2019:RBP

[3025] Liu Chong, Andrés Flores, Juan Carlos Quiroz, Francisco Contreras, and Rodrigo Wiff. Reproductive biology and population structure of white warehou Seriolella caerulea and silver warehou Seriolella punctata in the austral zone off Chile. Journal of Fish Biology, 95(4):1030–1039, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsuda:2019:WST

[3026] Keishi Matsuda, Masaaki Kamoshida, and Yoshitsugu Masuda. Wavelength-specific thresholds of artificially reared Japanese eel Anguilla japonica larvae determined from negative-phototactic behaviours. Journal of Fish Biology, 95(4):1040–1045, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagley:2019:DBM

[3027] Justin C. Bagley, Pedro De Podestà Uchôa de Aquino, María Florencia Breitman, Francisco Langeani, and Guarino R. Colli. DNA barcode and minibarcode identification of freshwater fishes from Cerrado headwater streams in Central Brazil. Journal of Fish Biology, 95(4):1046–1060, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2019:IBR

[3028] Shannon L. White, Benjamen C. Kline, Nathaniel P. Hitt, and Tyler Wagner. Individual behaviour and resource use of thermally stressed brook trout Salvelinus fontinalis portend the conservation potential of thermal refugia. Journal of Fish Biology, 95(4):1061–1071, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coiraton:2019:PGB

[3029] Claire Coiraton, Javier Tovar-Ávila, Karla C. Garcés-García, José A. Rodríguez-Madrigal, Rodney Gallegos-Camacho, Dario A. Chávez-Arrenquín, and Felipe Amezcua. Periodicity of the growth-band formation in vertebrae of juvenile scalloped hammerhead shark Sphyrna lewini from the Mexican Pacific Ocean. Journal of Fish Biology, 95(4): 1072–1085, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kamikawa:2019:RDF

[3030] Keith T. Kamikawa, Robert L. Humphreys, Jr., Brian W. Bowen, and Alan M. Friedlander. Recruitment dynamics and fishery characteristics of juvenile goatfishes Mulloidichthys spp. in Hawai’i. Journal of Fish Biology, 95(4):1086–1093, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Darcy:2019:EIT

[3031] Andrew P. Darcy, Graham D. Raby, Timothy B. Johnson, Trevor E. Pitcher, and Aaron T. Fisk. Effects of intracoelomic transmitter implantation on metabolic rate, swimming performance, growth and survival in juveniles of two salmonids. Journal of Fish Biology, 95(4):1094–1106, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:2019:SAG

[3032] Jessica S. Thompson. Salinity affects growth but not thermal preference of adult mummichogs Fundulus heteroclitus. Journal of Fish Biology, 95 (4):1107–1115, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palance:2019:LCM

[3033] Danial G. Palance, Beverly J. Macewicz, Kevin Stierhoff, David A. Demer, and Juan P. Zwolinski. Length conversions and mass–length relationships of five forage-fish species in the California current ecosystem. Journal of Fish Biology, 95(4):1116–1124, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Queiroz-Sousa:2019:SRF

[3034] Jamile Queiroz-Sousa, Sally A. Keith, Gianmarco S. David, Heleno Brandão, André B. Nobile, Jaciara V. K. Paes, Ana C. Souto, Felipe P. Lima, Reinaldo J. Silva, Raoul Henry, and Katherine Richardson. Species richness and functional structure of fish assemblages in three freshwater habitats: effects of environmental factors and management. Journal of Fish Biology, 95(4):1125–1136, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadras:2019:ETS

[3035] Hadiseh Dadras, Sergey Boryshpolets, Amin Golpour, Tomas Policar, Miroslav Blecha, and Borys Dzyuba. Effects of temperature on sperm motility of burbot Lota lota: spontaneous activation and calcium dependency. Journal of Fish Biology, 95(4):1137–1144, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jakobsdottir:2019:OTD

[3036] Klara B. Jakobsdóttir, Jónas P. Jónasson, Julian M. Burgos, and Jónbjörn Pálsson. Observations of two deep-sea chimaeroids in Icelandic waters. Journal of Fish Biology, 95(4):1145–1150, October 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mucientes:2019:IFS

[3037] Gonzalo Mucientes, José Irisarri, and David Villegas-Ríos. Interannual fine-scale site fidelity of male ballan wrasse Labrus bergylta revealed by photo-identification and tagging. Journal of Fish Biology, 95(4):1151– 1155, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adelir-Alves:2019:PMA

[3038] J. Adelir-Alves, D. Spier, H. L. N. Gerum, L. F. Machado, H. L. Spach, B. R. Boza, and C. Oliveira. Plectorhinchus macrolepis (Actinopterygii: Haemulidae) in the western Atlantic Ocean. Journal of Fish Biology, 95 (4):1156–1160, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2019:SET

[3039] Nick A. R. Jones, Rowena Spence, Faith A. M. Jones, and Helen C. Spence-Jones. Shade as enrichment: testing preferences for shelter in two model fish species. Journal of Fish Biology, 95(4):1161–1165, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2019:SWV

[3040] Ian J. Winfield. Social wellbeing and the values of small-scale fisheries. Journal of Fish Biology, 95(4):1166, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:Cb

[3041] Anonymous. Corrigendum. Journal of Fish Biology, 95(4):1167, October 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIk

[3042] Anonymous. Issue information. Journal of Fish Biology, 95(5):1169– 1170, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2019:SLM

[3043] Michel J. Kaiser. Shining a light on mitigation measures to reduce undesirable mortality in fish. Journal of Fish Biology, 95(5):1171, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heisey:2019:PSE

[3044] Paul G. Heisey, Dilip Mathur, Joanne L. Phipps, Jerry Christopher Avalos, Cory E. Hoffman, Steven W. Adams, and Eric De-Oliveira. Passage survival of European and American eels at Francis and propeller turbines. Journal of Fish Biology, 95(5):1172–1183, November 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2019:LHT

[3045] Yu-Jia Lin, Lotfi Rabaoui, Rommel H. Maneja, Mohammad A. Qurban, Khaled Abdulkader, Hussain Al-Nazry, and Ruben H. Roa-Ureta. Life history traits and temporal trends of abundance of the orange-spotted trevally (Carangoides bajad) from Saudi waters of the Gulf. Journal of Fish Biology, 95(5):1184–1194, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pirog:2019:RBM

[3046] Agathe Pirog, Hélène Magalon, Thomas Poirout, and Sébastien Jaquemet. Reproductive biology, multiple paternity and polyandry of the bull shark Carcharhinus leucas. Journal of Fish Biology, 95(5):1195–1206, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mai:2019:CZU

[3047] Ana C. G. Mai, Cristiano Q. Albuquerque, Valéria M. Lemos, Paulo R. Schwingel, Gian-Franco Ceni, Tatiana D. Saint’Pierre, and João P. Vieira. Coastal zone use and migratory behaviour of the southern population of Mugil liza in Brazil. Journal of Fish Biology, 95(5):1207–1214, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dermond:2019:HMD

[3048] Philip Dermond, Nicola Sperlich, and Jakob Brodersen. Heritable morphological differentiation in salmonids from two distinct stream types. Journal of Fish Biology, 95(5):1215–1222, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2019:TYL

[3049] Ross S. Glover, Robert J. Fryer, Chris Soulsby, and Iain A. Malcolm. These are not the trends you are looking for: poorly calibrated single-pass electrofishing data can bias estimates of trends in fish abundance. Journal of Fish Biology, 95(5):1223–1235, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caltabellotta:2019:AGT

[3050] Fabio P. Caltabellotta, Zachary A. Siders, Debra J. Murie, Fabio S. Motta, Gregor M. Cailliet, and Otto B. F. Gadig. Age and growth of three endemic threatened guitarfishes Pseudobatos horkelii, P. percellens and Zapteryx brevirostris in the western South Atlantic Ocean. Journal of Fish Biology, 95(5):1236–1248, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ford:2019:CLL

[3051] Matthew I. Ford, Chris K. Elvidge, Paul H. Patrick, Michael Sills, and Steven J. Cooke. Coloured LED light as a potential behavioural guidance tool for age 0 and 2 year walleye Sander vitreus. Journal of Fish Biology, 95(5):1249–1256, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2019:ASS

[3052] Richard James Kennedy, William O’Connor, and Michelle Allen. Atlantic salmon Salmo salar passing a natural barrier before and after construction of a hydroelectric station. Journal of Fish Biology, 95(5):1257–1264, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchant:2019:MCC

[3053] James L. Marchant and Anthony P. Farrell. Membrane and calcium clock mechanisms contribute variably as a function of temperature to setting cardiac pacemaker rate in zebrafish Danio rerio. Journal of Fish Biology, 95(5):1265–1274, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zavala-Munoz:2019:NPM

[3054] Francisca Zavala-Muñoz, Javier Vera-Duarte, Claudia A. Bustos, Jorge Angulo-Aros, and Mauricio F. Landaeta. Niche partitioning and morphospace in early stages of two sympatric Diogenichthys species (Myctophidae). Journal of Fish Biology, 95(5):1275–1285, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rabaoui:2019:DBM

[3055] Lotfi Rabaoui, Lamia Yacoubi, Daria Sanna, Marco Casu, Fabio Scarpa, Yu-Jia Lin, Kang-Ning Shen, Todd R. Clardy, Marco Arculeo, and Mohammad A. Qurban. DNA barcoding of marine fishes from Saudi Arabian waters of the Gulf. Journal of Fish Biology, 95(5):1286–1297, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mas:2019:EPD

[3056] Dalmiro Borzone Mas, Patricio F. Alvarenga, and Pablo A. Scarabotti. Ecological and phylogenetic determinants of life-history patterns among ten loricariid species. Journal of Fish Biology, 95(5):1298–1310, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2019:DEA

[3057] Fengjiao Ma, Yanping Yang, Min Jiang, Denghua Yin, and Kai Liu. Digestive enzyme activity of the Japanese grenadier anchovy Coilia nasus during spawning migration: influence of the migration distance and the water temperature. Journal of Fish Biology, 95(5):1311–1319, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wojan:2019:ECU

[3058] Erin M. Wojan, Nalana C. Carreiro, David A. Clendenen, Heather M. Neldner, Crystal Castillo, Susan M. Bertram, and Gita R. Kolluru. The effects of commonly used anaesthetics on colour measurements across body regions in the poeciliid fish, Girardinus metallicus. Journal of Fish Biology, 95(5):1320–1330, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larsen:2019:UOM

[3059] Matthew E. Larsen, Daniel C. Abel, Derek P. Crane, Scott L. Parker, Paul H. Yancey, Bryan A. Keller, and Dean R. Grubbs. Unique osmoregulatory morphology in primitive sharks: an intermediate state between holocephalan and derived shark secretory morphology. Journal of Fish Biology, 95(5):1331–1341, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nevatte:2019:GHE

[3060] Ryan J. Nevatte, Barbara E. Wueringer, Michael R. Gillings, and Jane E. Williamson. Genetic and historical evidence of common sawsharks Pristiophorus cirratus in the waters of southern Queensland. Journal of Fish Biology, 95(5):1342–1345, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pimentel:2019:NRS

[3061] Caio R. Pimentel, Ciro C. Vilar, Fernanda A. Rolim, Maria L. Abieri, and Jean-Christophe Joyeux. New records of the snow bass Serranus chionaraia (Perciformes: Serranidae) confirm an established population in the Brazilian Province. Journal of Fish Biology, 95(5):1346–1349, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyuji:2019:EGP

[3062] Mitsuo Nyuji, Ikki Yamamoto, Kazuhisa Hamada, Yukinori Kazeto, and Koichi Okuzawa. Effect of GnRHa on plasma levels of Fsh and Lh in the female greater amberjack Seriola dumerili. Journal of Fish Biology, 95(5):1350–1354, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gache:2019:EAT

[3063] Camille Gache, Frédéric Bertucci, Ana Sofía Guerra, Maelle Calandra, Tristan Berr, Julie Lafaye, Hendrikje Jorissen, Maggy Nugues, Janine Cossy, and David Lecchini. Effects of Asparagopsis taxiformis metabolites on the feeding behaviour of post-larval Acanthurus triostegus. Journal of Fish Biology, 95(5):1355–1358, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2019:VFS

[3064] Paul J. B. Hart. Vanishing fish: shifting baselines and the future of global fisheries. Journal of Fish Biology, 95(5):1359, November 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2019:IIl

[3065] Anonymous. Issue information. Journal of Fish Biology, 95(6):1361– 1362, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2019:LSS

[3066] Michel J. Kaiser. The last supper: standardizing stomach contents analyses is critical for understanding foodweb dynamics. Journal of Fish Biology, 95(6):1363, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amundsen:2019:FST

[3067] Per-Arne Amundsen and Javier Sánchez-Hernández. Feeding studies take guts — critical review and recommendations of methods for stomach contents analysis in fish. Journal of Fish Biology, 95(6):1364–1373, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2019:ABL

[3068] Bradley R. Moore. Age-based life history of humpback red snapper, Lutjanus gibbus, in New Caledonia. Journal of Fish Biology, 95(6):1374– 1384, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orbach:2019:SST

[3069] Dara N. Orbach, Meghan Donovan, and Craig F. Purchase. Sexually selected traits are larger and more variable in male than female beachspawning capelin, Mallotus villosus. Journal of Fish Biology, 95(6):1385– 1390, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murase:2019:FAI

[3070] Iki Murase and Kei’ichiro Iguchi. Facultative amphidromy involving estuaries in an annual amphidromous fish from a subtropical marginal range. Journal of Fish Biology, 95(6):1391–1398, December 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Axworthy:2019:SBI

[3071] Jeremy B. Axworthy, Joseph M. Smith, Martina S. Wing, and Thomas P. Quinn. Sex biased individual variation in movement patterns of a highly mobile, near-shore marine planktivore, the reef manta ray Mobula alfredi. Journal of Fish Biology, 95(6):1399–1406, December 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palm:2019:RGD

[3072] Stefan Palm, Jerker Vinterstare, Jan Eric Nathanson, Alexandros Triantafyllidis, and Erik Petersson. Reduced genetic diversity and low effective size in peripheral northern European catfish Silurus glanis populations. Journal of Fish Biology, 95(6):1407–1421, December 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krylov:2019:DCI

[3073] Viacheslav V. Krylov, Yulia V. Chebotareva, and Yuri G. Izyumov. Delayed consequences of the influence of simulated geomagnetic storms on roach Rutilus rutilus embryos. Journal of Fish Biology, 95(6):1422–1429, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raoult:2019:RUG

[3074] Vincent Raoult, Matt K. Broadhurst, Vic M. Peddemors, Jane E. Williamson, and Troy F. Gaston. Resource use of great hammerhead sharks (Sphyrna mokarran) off eastern Australia. Journal of Fish Biology, 95(6):1430–1440, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Safi:2019:EAW

[3075] Hamid Safi, Yangfan Zhang, Patricia M. Schulte, and Anthony P. Farrell. The effect of acute warming and thermal acclimation on maximum heart rate of the common killifish Fundulus heteroclitus. Journal of Fish Biology, 95(6):1441–1446, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mouchlianitis:2019:PSO

[3076] Foivos A. Mouchlianitis, Ana F. Belo, Ana R. Vieira, Bernardo R. Quintella, Pedro R. Almeida, and Kostas Ganias. Primary and secondary oocyte growth dynamics in anadromous semelparous Allis shad Alosa alosa. Journal of Fish Biology, 95(6):1447–1456, December 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deVilliers:2019:CVC

[3077] Nina M. de Villiers, Cassandra Barker, Louw Claassens, and Alan N. Hodgson. Conservation value of Codium tenue habitat for the endangered Knysna seahorse Hippocampus capensis. Journal of Fish Biology, 95(6): 1457–1464, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hook:2019:RSS

[3078] Samantha A. Hook, Charlotte McMurray, Daniel M. Ripley, Natasha Allen, Timo Moritz, Bianka Grunow, and Holly A. Shiels. Recognition software successfully aids the identification of individual small-spotted catsharks Scyliorhinus canicula during their first year of life. Journal of Fish Biology, 95(6):1465–1470, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weinrauch:2019:AAA

[3079] Alyssa M. Weinrauch, Tamzin A. Blewett, Chris N. Glover, and Greg G. Goss. Acquisition of alanyl-alanine in an Agnathan: Characteristics of dipeptide transport across the hindgut of the Pacific hagfish Eptatretus stoutii. Journal of Fish Biology, 95(6):1471–1479, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

daGraca:2019:NSM

[3080] Weferson J. da Graça, Renata R. Ota, and Wladimir M. Domingues. A new species of miniature Characidium (Characiformes: Crenuchidae) from the upper Rio Paraguay basin, Mato Grosso state, Brazil. Journal of Fish Biology, 95(6):1480–1485, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McQueen:2019:EFL

[3081] Kate McQueen, Monica Mion, Annelie Hilvarsson, Michele Casini, Hans J. Olesen, Karin Hüssy, Krzysztof Radtke, and Uwe Krumme. Effects of freezing on length and mass measurements of Atlantic cod Gadus morhua in the Baltic Sea. Journal of Fish Biology, 95(6):1486–1495, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haugen:2019:OPS

[3082] Janne B. Haugen and Yannis Papastamatiou. Observation of a porbeagle shark Lamna nasus aggregation at a North Sea oil platform. Journal of Fish Biology, 95(6):1496–1499, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hauser:2019:GSE

[3083] Frances E. Hauser, João P. Fontenelle, Ahmed A. Elbassiouny, Nicholas E. Mandrak, and Nathan R. Lovejoy. Genetic structure of endangered lake chubsucker Erimyzon sucetta in Canada reveals a differentiated population in a precarious habitat. Journal of Fish Biology, 95(6):1500–1505, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2019:AHP

[3084] Takaomi Arai, Hussein Taha, Razikin Amalina, Yoshiyuki Iizuka, and Chih-Wei Chang. Anadromy and heterogeneous population of a tropical shad Tenualosa ilisha in Malaysia, as revealed by otolith microchemistry and molecular evidence. Journal of Fish Biology, 95(6):1506–1511, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klinard:2019:LTR

[3085] Natalie V. Klinard, Jordan K. Matley, Aaron T. Fisk, and Timothy B. Johnson. Long-term retention of acoustic telemetry transmitters in temperate predators revealed by predation tags implanted in wild prey fish. Journal of Fish Biology, 95(6):1512–1516, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hung:2019:PSP

[3086] Tien-Chieh Hung, Marlin Rosales, Tomofumi Kurobe, Troy Stevenson, Luke Ellison, Galen Tigan, Marade Sandford, Chelsea Lam, Andrew Schultz, and Swee Teh. A pilot study of the performance of captivereared delta smelt Hypomesus transpacificus in a semi-natural environment. Journal of Fish Biology, 95(6):1517–1522, December 2019. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tao:2019:DEA

[3087] Min Tao, Hong Hu, Lu Huang, Shengnan Li, Liujiao Yuan, Tian Zhou, Can Song, Rurong Zhao, and Shaojun Liu. Differential expression of activin β_{A} and β_{B} genes in female allotriploid and diploid red crucian carp Carassius auratus red var. Journal of Fish Biology, 95(6):1523–1529, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:2019:SRS

[3088] Emmett M. Johnston, Paul A. Mayo, Paul J. Mensink, Eric Savetsky, and Jonathan D. R. Houghton. Serendipitous re-sighting of a basking shark Cetorhinus maximus reveals inter-annual connectivity between American and European coastal hotspots. Journal of Fish Biology, 95(6):1530– 1534, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weideli:2019:EAS

[3089] Ornella C. Weideli, Jeremy J. Kiszka, Philip Matich, and Michael R. Heithaus. Effects of anticoagulants on stable-isotope values (δ ^{13}C and δ ^{15}N) of shark blood components. Journal of Fish Biology, 95(6):1535– 1539, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghedotti:2019:FEI

[3090] Michael J. Ghedotti, W. Leo Smith, and Matthew P. Davis. The first evidence of intrinsic epidermal bioluminescence within ray-finned fishes in the linebelly swallower Pseudoscopelus sagamianus (Chiasmodontidae). Journal of Fish Biology, 95(6):1540–1543, December 2019. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).